A Complete Bibliography of Computer Graphics Forum: 2010–2019

Nelson H. F. Beebe University of Utah Department of Mathematics, 110 LCB 155 S 1400 E RM 233 Salt Lake City, UT 84112-0090 USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

 E-mail: beebe@math.utah.edu, beebe@acm.org, beebe@computer.org (Internet)
             WWW URL: http://www.math.utah.edu/~beebe/

10 June 2022 Version 3.00

Title word cross-reference

1 [TWSM17]. 2 [AH11, BSW^{+}12, BCK^{+}12, GBS19, GST14, GT16b, GLX^{+}16, HCGW14, HRRR18, HL15, HYL^{+}15, JWH19a, KKBJ16, KCL^{+}18, KG19, LJB^{+}12, LWS18, MBES16, NBHN17, OM13, OW19, OGHT10, PCX^{+}18, PW12, PPBT12, PEPM12, RGG18, SW10, Sch11, SK16a, SSS^{+}12, STP17, WZL^{+}12, WRS^{+}13, WGS10, YFW12, ZRLS19, dGCSAD11, dGLB^{+}14]. 3 [ADMAS18, ASH15, Att15, ATF12, AH11, AKM16, ARG^{+}16, BSY^{+}19, BCBB16, BCK^{+}12, BLS^{+}17, BHMT13, BdM14, BL18, CS16, COO15, CYY^{+}11, CNKI13, CHWL19, CACB18, DDÖ^{+}17, DS11a, DGR^{+}14, DER^{+}10, DMS14, EHH19, FHE19, FW17, GS14, GRT14, GLX^{+}16, GTB14, HT11, HMTH13, HRRR18, HL14, HL15, HMW^{+}15, HBO^{+}10, HFL12, HKM15, HCSC16, HAGO19, JH15, JJKL18, JLW10, JL19, KZB19, KKBJ16, KFW19, KLTZ16, KCH^{+}18, KGP^{+}12, LSF^{+}11, LT17, Lav11, LCSL18, LOM^{+}18, LDGN15, LLLY19, LGK16, LAFT12, LTX^{+}14, LTR19, LEM^{+}17, LN17, MHS^{+}14, MGY^{+}18, MCHW18, MGG10b, MbMYR15, MCG^{+}19, MPWC13, MEKM17, NWHWfD16, NW17, NRM^{+}12, NRP11, NREM14, ORT18, PBMG15, PSCC18, PB11, PEP^{+}11b, PGK10, POG13, PS10, PBC^{+}16, RXX^{+}17, RKSA17, RI17, RAMG15, RBRY19, RLYL14, SLB^{+}18, SY14a, SYM10, STKD12, SXY^{+}11, SN12, SK17]. 3 [SGS14, STC^{+}16, SJW^{+}11, SJWS13, SBL12, TDS^{+}16, TWS^{+}11, TSW^{+}19, TTB12, TBTB12, VGB^{+}14b, WL10, WCT^{+}15, WZK16, WLL^{+}17, WLJ^{+}18, WBAI19, WMXC19, WXW18, XXLX14, XXS^{+}15, XSQ13, XSX^{+}14, XXY^{+}18, YLLL15, YHL^{+}16, YL11, ZZH15, ZZWC16, ZKWG16, ZWHK16, ZTG^{+}18, ZSG^{+}18, ZJC13, vMRBPM17]. 4 [AKS^{+}19, CVCH14, CRC^{+}15, HRS18, KPG^{+}16, KBvP^{+}17, KGGP18, LWL^{+}16a, MKP^{+}16, WH17a, ZFG^{+}17, dHvPJV14, vPJtHRV12]. ^{F} [vdCvW16]. ^{I} [vdCvW16]. A [KP15]. α [CD10]. B [CBV^{+}14]. C^{1} [WGS10]. C^{2} [KP18]. L_{0} [WZL^{+}17]. ℓ_{0} [BSH15]. G [MIGMM17]. G^{2} [KP11, SV14]. k [FKSS13, MSAP15]. kd [CCI13, HH11]. L_{0} [WTL15]. L_{1} [WZCF15]. N [DVPSH14, TPSH14a]. O(n) [Gov19]. t [RBMS17]. τϵχ́νη [Duc14].

-Buffers [MIGMM17]. -Continuity [WGS10]. -D [EHH19, PCX^{+}18, SW10, SY14a]. -Distribution [RBMS17]. -Means [FKSS13]. -Minimization [BSH15]. -PolyVector [DVPSH14]. -rep [CBV^{+}14]. -shapes [CD10]. -Sided [TPSH14a]. -Sparse [MSAP15]. -Tables [vdCvW16]. -Tensor [dGLB^{+}14]. -tree [CCI13, HH11]. -Valent [RKSA17].

11th [AJL^{+}11].

2009 [CSLG10, PS10]. 2010 [LF11, WBP11]. 2011 [Kuh12, LSF^{+}11]. 2019 editorial v2 [CB19]. 2Brick [KLC^{+}15]. 2D [NSC14, HBW11, LH11, MHDG11]. 2D-D [NSC14].

31st [Ano10a]. 32nd [Ano11c]. 33rd [Ano12d]. 34th [Ano13f]. 360 [XTZ19]. 3D [Jac19, BLVD11, JTRS12]. 3D-Mesh [BLVD11]. 3D-Model [JTRS12]. 3D-Printed [Jac19]. 3DOR’09 [PS10]. 3DOR’10 [DS11a]. 3DOR’2011 [LSF^{+}11].

4DCT [AAS^{+}16]. 4PCS [MAM14].

60km [HLH^{+}16b]. 60sec [HLH^{+}16b]. 6DOF [BH15, LAFT12].

8th [AJL^{+}11].

À-Trous [HDL11]. Ablation [RWS^{+}10]. Absolute [GBS19, KRM^{+}15]. Absorption [ACAA^{+}19, BAO^{+}19]. Abstract [BBP10, HW10, HKSK18, PBK10]. Abstraction [BFG^{+}17, CNKI13, KNH^{+}18, KK11b, MDI^{+}18, PJR^{+}14, SAMS^{+}17, STKD12, SAAF18, vdZLBI11]. Abstractions [ARH12]. Accelerated [HPMB19, LBPH10, MS14, WSR^{+}17, MAG19]. Accelerating [BRDC12, BEM11, KTO11, LL18, RGG15]. Acceleration [AT10, GRDE10, LBPH10]. Accelerator [BAAM17]. access [ZFE16]. Accessible [RLH17a]. According [TX16]. Accuracy [CEX^{+}18, GKB^{+}11, KER^{+}14, SSKB15]. Accurate [BSH12, BXH10, CP10, DZC11, ENMGC19, FH18, Hol15, Pat17, SDMS15, SPSK13, TPSH14b, YZXW12, YWF^{+}19, DCV14, LN18, SPCR14]. ACM [CSLG10, DS11a, ID10, IC11, LSF^{+}11, PS10]. Acoustic [HHD^{+}12]. Acquisition [BPW14, CML^{+}12, CMF18, DSY10, GGG^{+}16a, KBÖ^{+}14, LDW^{+}10, LPG19, LDY10, MPM^{+}14, VF16, WZ15, YHL^{+}16]. Across [AKS^{+}19, BJG^{+}15, KZZM12]. Active [BZL^{+}18, CKH19, ZZZ^{+}19]. Activities [GSHM10]. Activity [APH^{+}12, CD18]. Activity-Led [APH^{+}12]. Actors [GVS^{+}15, KGRG17]. Actual [KUMY10]. Actuator [dSNV^{+}17]. Adaptable [LCC^{+}18, RSSL17, ZHK15]. Adaptation [GRC13, LGW18, FHE19]. Adaptations [ARM^{+}15]. Adapted [ZSS17]. Adapting [GLK16]. Adaptive [AHTAM14, And12, BYM18, BBH13, BXH10, BHU10a, CWW^{+}11, CZY11, DMNV12, FHHJ18, GPK^{+}12, GCSA13, IFL13, JZYP18, KL14, KH19, KBT^{+}12, KKR18, LGP14, LPD14, LIK19, LXY19, LMPS16, MWN^{+}17, MHK^{+}19a, NGM14, OK12, PGSD13, RCB11, REH^{+}11, RAMG15, RÖG17, SVLL10, SO12, SKSS14, SGEM16, TSdSK13, XSXM13, YIC^{+}11, ZZWC16, ZJL^{+}15, vTKP11, FCS^{+}16, SXLS19]. Adaptively [And10]. Adding [BB17, SSJ^{+}10, WYKR17]. Additional [ABK^{+}19]. Additively [Man16]. Adjacency [DRW15, MR17]. adjacent [SZ18]. Adjoint [HO17]. Adjustment [AP10a, LM10, LCG10]. Advanced [KG18, UBH14]. Advances [Cas12, GI18, JH15, KRP^{+}15, ZJL^{+}15]. Advantage [KMJE12]. Advected [WHT12]. Advection [AVBC16, KSW^{+}12, YMM10]. Advection-Based [AVBC16]. Advection-Diffusion [KSW^{+}12]. Advective [HWK15]. Advective-Diffusive [HWK15]. Adversarial [CLGS18, SLCL19, WGW^{+}19, WLZX19, XHW^{+}18, ZLZ^{+}18]. Aerial [ACA18, YXHH18]. Aesthetic [Gda17, KM16, OM13]. Aesthetics [DH16, ID10, IC11]. Affine [BLTD17, Vax12]. Affine-Kernel [BLTD17]. African [GSHM10]. Afrigraph [GSHM10]. after [MGN17, SHZD17]. Afterimages [RE12]. Again [GLG^{+}16]. Agent [LCM^{+}18]. Agent-Based [LCM^{+}18]. Agents [LPSV14, RPA^{+}15]. Agglomerative [DCYG19]. Aggregate [BNH^{+}16, KCH19]. Aggregates [SM14a]. Aggregation [HBW11, LH11, MMV^{+}13, MHDG11, ZAM^{+}16]. Agile [BH15, LMLG15, LMPD15]. Aging [IGAJG15]. Ahead [HCBW19]. Air [BJA^{+}15, EHH^{+}13]. Airline [RPMO13]. al [BPY16, ERA^{+}16, Man16]. Albedo [GMD10]. Albero [DPD^{+}17]. Album [ZH12]. Albums [RGM^{+}18]. Algebra [SAD^{+}16]. Algorithm [Cet19, EMP^{+}12, GKLS19, Gov19, LCDW16, LS15, OM13, PM16, SB13, SAAB11, TTN^{+}13, TBKP12, WZKP14, JCT14, LCLJ10, MSW19, ZJC13]. Algorithms [AGR19, BZL^{+}18, CDSS14, CJWL19, FMH16, FEM^{+}19, GI18, HH11, HSS17, KS13a, Meu19, SSO^{+}10, SLD^{+}17, WH17a, DFIM15, DMCN^{+}17, EMU17]. Aliasing [AGJ12, DFY14]. Aligned [CK14, DLY^{+}18, KCL^{+}18, NZH^{+}18, PWS12, TLRB18, WYKR17, WZG^{+}19]. Alignment [AKM16, CIE^{+}16, CRC^{+}15, ESKBC17, GVS^{+}15, SBC14, SDK^{+}15, NNRS15]. All-Frequency [IFL13, MC10b, OPP10, IFDN12]. All-Hex [GSZ11]. all-scale [LN17]. Alleged [RPSF15]. Alleviating [BMS^{+}10]. Allocation [DDH^{+}18]. Allocator [VH15]. Allometric [MKR11]. Along [BJG^{+}15]. Alpha [GO10]. Alternating [HCW17, KR19]. Alternative [EWK^{+}13]. Ambient [LK10, LWDB10, MSW10, MBDC15, SGG15, Tim13, YWC^{+}10]. Ambiguity [SJB^{+}17]. Ambiguous [CD10]. Ambrosio [BCGL18]. AmniVis [NLB^{+}13]. among [Man16]. Amplification [BDC18]. Anaglyph [LMSG16, SW11]. Analogy [MHS^{+}14]. Analogy-driven [MHS^{+}14]. Analyses [BMS^{+}10, XBL^{+}18]. Analysis [ASB^{+}17, AHG^{+}19, ACAA^{+}19, AYLM13, AFK^{+}14, AAS^{+}16, ABG^{+}12, ACC15, ACOHS18, AGCO13, BAT11, BSK16, BPW14, BH19, BDF^{+}14, BPFG11, BWM^{+}11, BCGS13, BSK^{+}17, BMPM12, BHMT13, BvLBS11, BTB13, BHU10b, BTM^{+}19, CT11, CHH^{+}19, CD18, CGM19, CBK^{+}17, CKGC14, CLCL11, CJZW12, CWL^{+}15, CKH19, DAF^{+}18, DW13, DI18, DPD^{+}15, EKFM12, ERHH11, EHH^{+}13, ER18, ESKBC17, EBC17, FR11, FKRW16, FEM^{+}19, GM19, GHX^{+}17, GMLMG12, GL12, GHWG14, HK12, HBRFP19, HRD^{+}15, HCGW14, HZRS18, HCW17, HG18, HMP^{+}12, HFL12, HSvK18, HKM15, HO17, JFCS17, JE13, JKK^{+}18, KCJM16, KBL19, KMAB15, KFH10, KBHM15, KLCF10, KCA^{+}16, KFM^{+}19, KASH13, KVD^{+}10, KKL^{+}16b, KTW^{+}13, KSKL13, KBÖ^{+}14, LCP^{+}12, LSS^{+}12, LZW^{+}13, LMG^{+}18a, LB19, LFS^{+}15, LWT^{+}15, LA15, LJC17, LJLH19, MJK17, MPM^{+}14, MGB^{+}12, MRCB18, MC10a, MLD^{+}18, MHHH15, NHL16]. Analysis [OAM^{+}18, OGW19, OT12, OBCCG13, OMPG13, PSPM12, PPW18, PGS^{+}16, Pat16, PPY^{+}16, PDC^{+}19, PPBT12, Pos11b, RCMM^{+}16, RPK^{+}12, RL14, RHM^{+}12, SHSK16, SPB^{+}17, SvW13, SXY^{+}11, SV10, SSM12, SMJ17, SÖA^{+}19, SvL16, SGB13, SS15b, SBM^{+}14, SJW^{+}11, SMB^{+}17, TFA^{+}11, TPRH11, VHG^{+}18, WXL^{+}11, WMZ12, WCB15, WHP^{+}11, WAF^{+}11, XDC^{+}13, XKHK17, YKS^{+}19, YPF19, YNM^{+}13, ZAM^{+}16, ZSL^{+}17, vLKS^{+}11, DFIM15, KGGP18, NGM14, NNN11, SNJ^{+}14]. Analysis-Oriented [MC10a]. Analytic [AGJ12, AWJ13, BH19, EHH19, HC14, MS13, MRMH12, MMG10, NBMJ14, Ste19, TBTB12]. Analytics [ARRO^{+}17, ALA^{+}18, BEF17, BLY^{+}11, BJA^{+}15, CWGvW19, CHH^{+}19, CBK^{+}17, CLY17, CE19, DI18, DPD^{+}17, DWT^{+}11, EASKC18, ERT^{+}17, HSK^{+}10, HJM^{+}11, KMJE12, KWS^{+}15, KHI^{+}19, LGH^{+}17, LRB^{+}15, MHK^{+}19b, OSR^{+}14, PZDD19, RSM^{+}16, RvdHD^{+}15, RCMA^{+}18, RMM15, SGS18, SHP^{+}19, WDC^{+}10, ZAM^{+}16, ZCW^{+}19]. Analyzing [BAO^{+}19, JL17, KMJE12, LKZ^{+}15, LMA^{+}18, RXX^{+}17, SHS^{+}17, WDM^{+}12, WH17b]. Anatomical [DMNV12, GRP10, ZHK15]. Anatomy [NJB^{+}11]. Anatomy-Guided [NJB^{+}11]. anchored [SHS^{+}17]. Anchors [SFLP18, WSK^{+}19]. Android [LFC14]. Aneurysm [SLB^{+}18, vPGL^{+}14]. Aneurysms [LMW^{+}19, MVB^{+}17, NJB^{+}11, NLB^{+}13]. Angiography [SSM12]. Angle [AVR10, Gam16, VR12, SHD16]. Angles [SK16b, XYLY19]. Animal [DHS^{+}13, GKPL11, SvL16, WPP13]. Animals [KKBJ16]. Animated [BSAP11, GEY12, GRR^{+}16, KSO10, KCH19, KFA^{+}10, KFA^{+}14, LS10a, LGP14, LLD10, SC16, SO10, TTB12, XSQ13, YYG18]. Animating [GRP10, PNVS17, SJ13a]. Animation [AW13, ARG^{+}16, AO13, BYL16, BSC16, BS19, BTST12, BM15, BBL12, CH12, CTL13, CKGC14, CZZ^{+}18, CYI^{+}12, CNKI13, CKE^{+}12, ECN14, GP12, GCY^{+}14, HMLP13, HENfSYS16, HSmCY13, HSK14, HLL16, JCK^{+}13, JK13, KL14, KSO10, KL19, KYC16, KS12b, LF11, LCSCO10, LXFW11, LAFT12, MYGY19, MPP^{+}13, MGC^{+}16, MWTI19, OAO11, PHTB12, PMG13, RC19, RPPD17, SOC19, SKZF11, SCR^{+}18, SMM13, SHS13, SBC^{+}17, SJF11, TCLK12, TTN^{+}13, TWT^{+}16, TAAP^{+}16, VVE^{+}10, WPP13, WWS^{+}15, WDAH10, XWG^{+}13, XLL^{+}10, YLHQ12, YMJ^{+}17, YLRC10, ZLYL17, ZXTD10, ZLKW13, ZRJ^{+}15, vBE11, AFHdL14, CIPT14, GDAU14, HNJ^{+}14, JZW14, RTK^{+}14, TPSH14b, TSK14, VB14a, VF14, WGO^{+}14]. Animation-Aware [MPP^{+}13]. Animations [BBT11, FLJ^{+}14, MP10, NC10, GDAU14, VF14]. Anisotropic [BPY16, BMR^{+}16, CG16a, CHK13, FV14, MYLZ16, MRMH12, QCW^{+}18, SGW12, VF16, XYLY19, ZZCJ14, RGB^{+}14]. Anisotropy [OBH^{+}11]. ANN [TSPP16]. Annotated [KLTZ16, RAMG15, WTH^{+}13]. Annotating [WCM15]. Annotation [HBRFP19, WTLY12]. Anomaly [SBM^{+}14]. Anonymized [TLFC16]. Anti [AGJ12, DFY14]. Anti-Aliasing [AGJ12, DFY14]. Antialiased [EHH19]. Antialiasing [JESG12]. Anticipating [OGW19]. Any [MRL10, MFL13, Kur15]. AOI [BKW13]. Aorta [KPG^{+}16, MKP^{+}16]. Apertures [MPCG12, MRD12]. Appearance [ACG^{+}17, BSH12, BBDA10, CRC^{+}15, DKY16, DHI^{+}15, GMM^{+}12, GG15, GCGP18, HLR^{+}11, JRJ11, KRP^{+}15, LAS^{+}11, LL19, LWB14, LPG19, MWS^{+}16, MRMH12, MSHD15, MES^{+}11, NSR17, NMM^{+}19, OVB^{+}15, PCDS12, PdMJ14, SPN^{+}16, SKZ11, VAW^{+}10, VF16, WRK^{+}16, WW11, WDR11, WZ15, YLD^{+}18, CVCH14, IMDN14, NNRS15]. AppFusion [WZ15]. Application [Deb18, HSK^{+}10, HLJ^{+}13, LFC14, YB18]. Application-Specific [Deb18]. Applications [AAS^{+}16, BAA^{+}16, BPY16, BCBB16, CDSS14, CG16b, DPD^{+}15, EHH^{+}13, FPC^{+}16, FMH16, GI18, HWU^{+}19, HBO^{+}10, HGB^{+}10, HSvK18, KH19, KCA^{+}16, LCDW16, MWS^{+}10, MPWC13, MLK^{+}13, NWHWfD16, NK16, OLF^{+}14, PP10, PW13, PDW^{+}14, RPMO13, SBD15a, SSB^{+}14, VBHH13, WDC^{+}10, WJB^{+}13, WBS^{+}13, WD11, YNM^{+}13, ZWRH14, ZTG^{+}18, PHM^{+}14]. Applied [WH17a, BHT19, KDCM14, YMM10]. Applying [SHP^{+}19, XADR13]. Approach [AHMAM15, ABCCO13, BO19, BSAP11, BLW11, BBBL11, CWGvW19, CC14, DPD^{+}17, ELM^{+}12, FFN18, GLHH13, GD10, HJM^{+}11, HJS^{+}17, JBG17, JWL^{+}13, KPRN11, KPNS10, KS10, KK18, Lai13, LDdLRB16, LMW^{+}19, LZQ13, Mér11, NBCW^{+}11, OSR^{+}14, PR19, PZDD19, PC12, PNVS17, RCB^{+}17a, Ros13, SW10, SKZ11, SB19, Tak19, TMRL14, WHWB16, WLL^{+}17, WLX19, WT11, WC16, WCH^{+}15, YKM12, YWF^{+}19, dGCSAD11, vdEvW13, GBM^{+}19, PCBL16]. Approaches [AHG^{+}19, CGM19, KCA^{+}16, SDA^{+}18]. Approximate [ENSD12, GRT18, HLS12, KS12a, RLH^{+}17b, SGS14, YB18]. Approximately [KLAB15]. Approximating [CG17, EDPB15, HMA15, NL13]. Approximation [BTP13, BSH15, CDSS14, KER^{+}14, LK13, LPG13, MS14, MRL10, NBH18b, NL13, PPL13, PCDS12, SDS^{+}16, SMP13, XTZ19, YYG18, HP11]. Approximations [BF15]. April [LSF^{+}11]. ARAPLBS [TE18]. Arbitrary [AASB14, CYC15, DHI^{+}15, HH10, KFA^{+}14, LK13, MG10, MMP16, NS11, WWH^{+}10, CVDL16, VF14]. Arc [BSK^{+}13]. Archaology [AJL^{+}11]. Architectural [BKD^{+}17, Hua17, MHA17, SMS^{+}17, ZCX19]. Architecture [BAAM17, BP13, VOS^{+}10, WDM^{+}12, DPW11, GGM12, SKK^{+}14a]. Architectures [LCDW16]. Arcs [SK16b, SWG16]. Area [CVJ15, GL12, GUK^{+}17, KYM19, NM14, NPW10, UFK13, JZW14]. Area-Preserving [GUK^{+}17, UFK13]. Areal [BMS^{+}10, RHM^{+}12]. Areas [SK16b]. Arm [NVH^{+}13]. Arm-Muscle [NVH^{+}13]. Arousal [KKL16a]. Arrangements [CML^{+}12, MMP16, ZCOAM14]. Arranging [YH13]. Array [JEK^{+}19]. Arrays [BTS^{+}17b, FKR13, HGO18, PSHZ^{+}15]. Art [AHG^{+}19, AMA^{+}16, BBT11, BFR17, BHP15, BCMP18, BRWM18, BCD^{+}12, CLLC15, DFIM15, ERT^{+}17, FW17, GP12, GGG^{+}16a, GT18, KRP^{+}15, KKF^{+}17, KBG^{+}15, KKL^{+}16b, KPK18, KLC^{+}15, KYC16, LCCC13, LHT17, LKG^{+}16, LGH^{+}17, MFPA15, MGM^{+}19, MKRE16, NSG11, NMSL19, NK16, Pat16, PP10, PVS^{+}18, PPF^{+}11, RDGK12, RGG^{+}14, SLE17, SRM^{+}19, SPN^{+}16, SLKL14, SB19, TDS^{+}16, TAEE15, VCD^{+}16, WWS^{+}15, YIC^{+}12, ZTG^{+}18, ZSG^{+}18, ZJL^{+}15, vLKS^{+}11, PDC^{+}19]. Art-photographic [SLKL14]. arterie [KGGP18]. Arteries [DHS^{+}13]. Artery [GOH^{+}10, GHB^{+}17, KPG^{+}16]. Articles [CKH19, KvLB14]. Articulated [GHK^{+}10, JLW10, LG13, MCH13, RF15, TST^{+}15, WLZH17, WBM^{+}18, YLX^{+}16]. Articulated-ICP [TST^{+}15]. Articulated-Motion-Aware [WLZH17]. articulation [TSK14]. Artifact [MM19]. Artifacts [BKM^{+}19, GL10b, TMHD12, VCL^{+}11]. Artificial [dSNV^{+}17]. ArtiSketch [LG13]. Artistic [CJZW12, KPD10, KS10, LLC13, SPN^{+}16]. Artistically [CS16]. As-Conformal-As-Possible [YMYK14]. As-Killing-As-Possible [SBCBG11b]. As-Rigid-As-Possible [WCX^{+}13]. As-Similar-As-Possible [CG16b]. ASCII [MFPA15]. Aspheric [JKL^{+}16]. Assemblage [GDG12]. Assembly [Jac19, MRS18, NJGW15, TWT^{+}16, XXM^{+}13, Ano10a, Ano11c, Ano12d, Ano13f]. Assess [SSM12, BHT19]. Assessing [KH18]. Assessment [BCBB16, FAFM19, KPG^{+}16, Lav11, LLX^{+}11, MTM12, RWS^{+}10, RPSF15, vPGL^{+}14]. Asset [LTR19]. assets [LN17]. Assignment [KYKL14]. Assisted [BAAR14, BvLBS11, KWC^{+}12, LDGN15, MMF10, JHT14, ZWJ^{+}19]. Association [Ano10c, Ano11e, Ano12f, Ano13i, GSHM10]. Associative [KRD^{+}15]. Astronomical [WLM13]. Asymmetric [SK16a]. Asynchronous [SKZF11]. Atlas [LBJ^{+}16, NS11]. Atlases [HWA^{+}10]. Atmospheric [AHG^{+}19]. Atomic [LBH12a]. Atomistic [FKE13]. Attention [HBO^{+}10, KHW13, LCSL18, WWV17]. Attention-Guiding [KHW13]. Attenuators [BKB^{+}12]. Attractors [YMM10]. Attribute [BHMT13, CDA^{+}14, DKB^{+}16, MFL13, SSGM17]. Attribute-Based [CDA^{+}14]. Attribute-preserving [SSGM17]. Audio [GVS^{+}15]. Audiovisual [DDH^{+}18]. Auditors [Ano10c, Ano11e, Ano12f, Ano13i]. Auditory [HHD^{+}12]. Augmentation [BBIG17]. Augmented [BAAR13, BBCW10, EWK^{+}13, KAS^{+}19, LE13, MS10b, MJK11, OKG^{+}10, PHE^{+}11, RGSK10, SYC10, XZP^{+}13, HM15, YLD^{+}18]. Augmenting [NW13, WBAI19]. August [CSLG10, IC11]. Aura [JKL18]. author [LTC18]. Authoring [GPGB11, GPG^{+}16, HBDP17, RJT18, SH14a, WPHC16]. Authorship [SRG16]. Autocorrelation [AKM16]. autoencoder [NW17]. Automated [LLN^{+}14, MMNG17, vdKdJP^{+}19]. Automatic [ATCO^{+}10, BAAR13, BMWW14, CTHAM10, DBLW15, ELM^{+}12, GSE^{+}14a, GOH^{+}10, HCBW19, KLW12, KvLB14, KGRG17, LSZ^{+}18, LCUR14, LTX^{+}14, MEKM17, NS11, PGVG19, SM11, SML15, SHJ^{+}16, ŠBM^{+}10, WK12b, WXR^{+}16, WBSH^{+}13, YLRC10, ZCX19, vMRBPM17, MKP^{+}16]. Automatically [BTST12]. Automating [WZL^{+}12]. Automultiscopic [RHS^{+}12]. Autonomous [ELPH19]. autostereoscopic [CHA^{+}14]. AutoStyle [LCUR14]. Avalanches [CEG^{+}18]. AVOCADO [SLSG16]. Avoidance [BP17]. Awake [CWGvW19]. Award [Ano15a, Bru11, Can11, Eis11]. Aware [BPFG11, BFR17, BJFadH19, Cet19, CSD11, CG19, CLLC15, CK13, CBSS17, DSWH17, GDML13, GRE11, GPR^{+}15, HSmCY13, JWS12, JKJL18, JWL^{+}13, KLW12, KYC16, KWW^{+}14, LBG16, LGH13, LJZX15, LWS^{+}16, LMHH14, LMLF15, MS12, MPP^{+}13, NKB14, PTP^{+}15, PTA^{+}11, SLA15, SRWS10, TOZ^{+}11, TWT^{+}16, Tok15a, TMH11, WJDZ14, WTL15, WLZH17, ZSJT19, ZOA^{+}18, ZDJ16, DCNP14, HZL19, JLKL16, KKTD17, KDCM14, LM10, LFA^{+}15, PSP^{+}14, SP13, YWM15, YCL^{+}17, CACB18]. Awareness [LTC18, MbMYR15, NJC^{+}19]. Axes [CHWL19, RSSL17]. Axis [HCGW14, LKF12, PWS12, RAV^{+}19, TLRB18, WYKR17, YYG18]. Axis-Aligned [PWS12, TLRB18, WYKR17].

B [HQH15, JLW10, RC18]. B-Mesh [JLW10]. B-Spline [RC18]. B-Splines [HQH15]. Back [GLG^{+}16]. Backward [DBK11]. Bacterial [KMA^{+}19]. bag [LBBC14]. bag-of-features [LBBC14]. Baking [KBS11b]. Balance [dSNV^{+}17, HNJ^{+}14]. Balancing [SHQL18]. Ballet [Ano19a]. Balloons [STBG12]. Balls [CD10, CDSS14, LK13]. Bamboo [NKSI16]. Bamboo-Copters [NKSI16]. Band [DBLW15, FAW^{+}16, SWT^{+}18]. Band-Limited [DBLW15]. Bandlimiting [YB18]. Bar [HSBW13, SHK15, WPHC16]. Barycentric [CG16c, Rus10, WBCGH11, YS19]. Bas [JSLW14, ZZWC16]. Bas-Relief [JSLW14, ZZWC16]. Base [JLW10]. Baseball [ODS18]. Based [ASL^{+}19, AMTMH12, AIAT12, ATO17, AWCO10, AVBC16, BDA^{+}17, BSW10, BEJM15, BKY^{+}16, BDF^{+}14, BF15, BMO^{+}14, BHP15, BÖK11, BHH13, BSK^{+}17, BFG^{+}17, BCD^{+}12, BvLBS11, BB12b, BHU10b, BMS^{+}10, CG16a, CYC15, Cet19, CC14, CDA^{+}14, CSD11, CZZ^{+}18, CWL19, CCTL12, CDS16, DGP17, DKL10, DLGY12, DFY14, FWPS11, FV14, FW17, FP15, FE17, FB11, GSP19, GD10, GMM^{+}12, GTB^{+}13, GPRS14, HENfSYS16, HFM10, HMTH13, Hd14b, HFE13, HL18, HLL^{+}19, HWF^{+}17, HHC^{+}13, IGAJG15, IEH^{+}14, IMAW15, JL17, JZYP18, KLMK19, KS18, KTO11, KWS^{+}15, LDW^{+}10, LS10a, LBK14, LAS^{+}11, LK17, LCP^{+}12, LTKD15, LPG19, LMPD15, LFS^{+}15, LLB^{+}10, LVW^{+}15, LS15, LCM^{+}18, MWN^{+}17, MGG10b, MLP^{+}10, MDBS14, MVB^{+}17, MMO16, MRS12, MESG11, MWW12, NBH18a, NNN11, NKB14, NC16, ÖKB10, PdMJ14, PB11, PW17, PP10, PPF^{+}11]. Based [PNVS17, QCW^{+}18, RCB^{+}17a, RBMS17, RPLH11, RHL12, RÖM^{+}15, RPPD17, RF15, SOC19, SD10a, SAG^{+}13, SK17, SHP^{+}19, SMG10, SJT^{+}19, TPSH14a, TGM12, TE10, TTB12, TE18, TDDD18, VCL^{+}11, VS10, WL10, WZL^{+}12, WHB^{+}13, WMB15, WLL^{+}17, WESW17, WLM13, WDK^{+}13, WTLL13, WWD15, XWL^{+}15, XXM^{+}13, XZP^{+}13, XWY^{+}15, YYL^{+}16, YWF^{+}19, YWC^{+}10, YZL17, YBK^{+}12, YWTY12, ZDM^{+}14, ZCC14, ZZT15, ZDJ16, ZWY^{+}13, ZLW15, ZXW19, ZCBK12, dHvPJV14, vFG11, vTKP11, AASB19, ARLC^{+}13, ACOM12, AFHdL14, ARB^{+}18, AH11, BBK^{+}19, BLY^{+}11, BBB^{+}18, BWH^{+}11, BKM^{+}19, BHS^{+}17, BP19, CHH^{+}19, CD18, CAS^{+}19, CCM19, CIPT14, DWR10, DZCC19, DKWB18, DMS14, DBS^{+}11, DHI^{+}15, DMCN^{+}17, EPAS11, FH18, FLJ^{+}14, GHK^{+}10, GGM12, GSC18, GCY^{+}14, GCGP18, GTB14, HWA^{+}10, HNJ^{+}14, HLH^{+}16a, HGA^{+}10, HHGJ15, HKSK18, IY10, IYS^{+}13, IUDN10, JTRS12, JC10, JKK^{+}18, KMS^{+}13, KMAB15]. based [KS10, KYKL14, KK18, KO19, KGK19, KMM^{+}18, KTW^{+}13, KB12, LMM10, LVPI18, LCCC13, LOM^{+}18, LP15, LCB^{+}18, LEE17, Lie17, LLM^{+}17, LYG15, LCMP19, LFA^{+}15, LWX^{+}15, MW11, MH13, MRS18, MKRE16, NGM14, NS19, NB15, OJS^{+}11, PSPM12, PJJ^{+}11, PEP^{+}11a, PC12, RKR^{+}16, RZS10, RLGH15, RJT18, RMG18, RBDD18, RSK10, RSK13, SSKB15, SCN^{+}16, SLE18, Sch11, SXY^{+}11, SGG15, STP17, Szy11, TTN^{+}13, TWT^{+}16, TWC^{+}16, TSK14, VVC^{+}11, VMH^{+}13, WZC^{+}11, WFZ^{+}15, WTMT18, WLX19, WLS13, WGS10, WT11, WFLW18, XWG^{+}13, XWY19, XLL^{+}10, YCL^{+}17, ZLKW13, vDHO16, CSLG10, RGG^{+}14, BO19, TDF^{+}15]. Bases [HA17, KBB^{+}13]. Basic [DI18, KP11]. Basis [IYS^{+}13, MGV11, SMG10]. Bayesian [BB17, DRA10, WYD^{+}13]. BCC [VCRG14]. Be [GTK^{+}12, KP15, KSBC12]. Beam [DBK11, HCJ13, LWY^{+}11, NNDJ12, SDS^{+}16]. Beams [MGJ^{+}19]. before [SMJ17]. Behavior [LBK14, Ros13, WH17a, vdCAvW14]. Behaviors [CKH19, HRD^{+}15, LJK^{+}12]. Behaviour [RPA^{+}15, VAW^{+}10, WLZ13, ZCT18]. Behaviours [LCM^{+}18]. Beltrami [HP11, NBH18b, PPH^{+}13, QCW^{+}18, ZLW15]. Benchmark [MDBS14].

Bending [WLZ13, ZLW^{+}16, ZR13]. BendyLights [KPD10]. Bertin [AKMM11]. Better [DH16]. Between [GSTOG16, HB19, KFA^{+}14, LK19, LTR19, OBCCG13, SNB^{+}12, VVE^{+}10, EBC17, EHA^{+}19, JKL18, PCBL16, SBC16, SVBC19, SZ18, TMRL14, VM12]. BetweenIT [WNS^{+}10]. Beyond [Cas12, KASH13]. Bézier [SS15a, VSK16]. Bi [GSDC17, Hua17, ZSW^{+}10b, KP11, KP15]. Bi-cell [ZSW^{+}10b]. bi-cubic [KP11, KP15]. Bi-Directional [Hua17]. Bi-Layer [GSDC17]. Bias [ENSD12, GUS12]. Biases [HF16]. Bidirectional [HFM10, JA18, LZ10, NB15, PBPP11, PRDD15, RSK13, SKZ13, SRK13]. Bifurcation [MBES16]. Big [LJH13]. Biharmonic [HQH15, Rus11, WPG12]. Bijective [SHF13]. Bilateral [BCCS12, IYS^{+}13, LWS^{+}16, YBY10, vKZH13]. Bilinear [MS14]. Billboards [GHK^{+}10]. Bimodal [DDH^{+}18]. Binary [BBDA10, CCLN10, SPSK13]. Binocular [ZHLW18, DKR^{+}14]. Biological [BLY^{+}11, DTS^{+}14, FSTR13, GKPL11, HVH^{+}16, HPvU^{+}16, KRS^{+}13, LDB11, LMP13, LMS^{+}16, LBH12a, SCD^{+}16]. Biologically [LWPL15]. Biologically-Inspired [LWPL15]. Biomedical [BHRD^{+}15, BJG^{+}15, KKL^{+}16b, PBC^{+}16, PKE15, RvdHD^{+}15, SLSG16, LPSV14, ZH14, dHvPJV14, vPGL^{+}14, vdCAvW14]. Biomolecular [KKF^{+}17, KRS^{+}13, KKL^{+}16b]. Biophysically [IGAJG15]. Biophysically-Based [IGAJG15]. Biopsy [HMP^{+}12]. Biorthogonal [HA17, RLH^{+}18]. Bipartite [CLCL11, PFH^{+}18]. Biped [LK17]. Bipedal [FvdP15]. Bird’s [KHI^{+}19]. Bird’s-Eye [KHI^{+}19]. Birefringence [Ste19]. Birefringence-Induced [Ste19]. Bisector [ZCK17]. Bit [OBGB11]. Bitmap [WCB15]. Bitmaps [RLMB^{+}14]. Bivariate [CGT^{+}15, SARZL10, ZRLS19]. Black [LM15, SO12]. Blackwellization [SKMS18]. Bladder [RCMA^{+}18, RCMA^{+}18]. Blend [TE18]. BlendForces [BSC16]. Blending [BBCW10, GAK10, KG18, OK12]. Blendshape [KBB^{+}17, LXFW11]. Block [KGR^{+}16]. Blocks [KBW^{+}12, LCDW16, GGM12]. Blog [HC14]. Blood [BBB^{+}18, LGP14, NJB^{+}11, YBK^{+}12, dHvPJV14, vPJtHRV12, vPGL^{+}14, KGGP18]. Blood-Flow [dHvPJV14, vPJtHRV12]. Blooming [ZFG^{+}17]. Blue [CG12, HB19, KS11, KS12a, PCX^{+}18, RRSG16, YGJ^{+}14]. Blue-Noise [PCX^{+}18, RRSG16, YGJ^{+}14, CG12]. Blur [HWK^{+}10, JKL13, KSP^{+}18, LDG19, MTAM12, MVH^{+}14, NSG11, SGG16, TMH11, WWT^{+}16]. Blur-Aware [TMH11]. Bodies [Aan18, FP15, LWPL15, LYG15, NL18, WWD15, ZZT15]. Body [ARB^{+}18, BET14, DLGY12, HA17, JKJL18, KÓOH13, LBRM18, STC^{+}16, vBE11]. Body-Supporting [LBRM18]. Body-Worn [KÓOH13]. Bokeh [JKL^{+}16, MRD12]. Bone [ZHK15]. Bones [TE18]. Book [AKS^{+}19, CHWL19, EHH19, FHE19, HAGO19, JL19, LLLY19, LTR19, WMXC19]. Books [RL19]. Booleans [PCK10]. Boosting [OJS^{+}11, ZDM^{+}14]. Botanical [NPDD11, WLJ^{+}18]. Bottlenecks [ABW^{+}15]. Bounce [JR16, LWLD11]. Boundaries [Hol15, MSWI12, SGSP15]. Boundary [BXH10, BLVD11, CK10a, DBS^{+}11, FM15, GDML13, HEW15, LJB^{+}12, MIW13, SCN^{+}16, TL16, TDNL18, ZT10]. Boundary-Aware [GDML13]. Boundary-based [DBS^{+}11]. Bounded [GMD10, MM18]. Bounding [BHH13, BM15, CTHAM10, GHB15, MB18, ND12, VKJ^{+}17, VHB16, VMTS10]. Bounds [She12, WTTM15]. Bowl [MNP^{+}17]. Box [Att15, CR16b, CDA^{+}14]. BqR [PSC10]. BqR-Tree [PSC10]. Bracketing [GPR^{+}15]. Brain [ACAA^{+}19, HMP^{+}12, SEI10, Thi11]. Brainbows [WOH13]. Branch [SSW14a, ZBM^{+}17]. Branching [LKD^{+}17, SPK10]. BRDF [BPV18, BÖK11, BNH10, CCM19, DHI^{+}15, GMD10, GGG^{+}16a, HFM16, LXY19, NKLN10, ÖKB10, PR12, RGB^{+}14, SARZL10, XWZB17]. BRDF-Oriented [LXY19]. BRDFs [ACG^{+}17, FV14, HLR^{+}11, KGK19, MHRK19, SSGM17, VF16, XWZB17]. Breadth [GL10a]. Breadth-First [GL10a]. Breakdown [HGH^{+}11]. Brick [KLC^{+}15]. Bridge [WDC^{+}10]. Bridging [YKS^{+}19]. Broad [SR19]. Broad-Phase [SR19]. brochures [BMWW14]. Broken [CK11a]. Browsing [KRD^{+}15, KFLCO13]. Brush [TZD11]. Brushables [LFA^{+}15]. Brushes [VVC^{+}11, ZT10]. Brushing [FH18, KFH10, RSM^{+}16]. BSDF [PdMJ14]. BSDFs [Hd14b, RBMS17]. BSSRDF [MES^{+}11]. BTF [HHNČ19, RJGW19, WDR11]. Bubble [HGH^{+}11]. BubbleNet [MSFM16]. Bucket [KKS^{+}17]. Budgets [DBS^{+}18]. Buffer [AMAM13, LJBA13]. Buffers [MIGMM17]. Building [ALA^{+}18, KBW^{+}12, KWS16, NBA18, SK17, WFLW18]. Buildings [CML^{+}12, KMK12, LWB14, MWW12, VKW^{+}12]. Bulk [ASB^{+}17]. Bulk-Heterojunction [ASB^{+}17]. bundle [TSdSK13]. Bundles [HGRS^{+}17, JGH11, OVV10, TE10, LBA10]. Bundling [HET12, LHT17, ZSJT19, ZWHK16]. Buoyancy [WW16]. Businesses [WHS^{+}18]. BVH [ÁSK14, GD16, GM19, GPP^{+}10, HMB17, HPMB19, WTMT18]. BVH-based [WTMT18]. BVHs [DKY16]. By-example [ZLL13].

C [KDCM14]. C-LOD [KDCM14]. Cable [MCJM18]. Cache [BRDC12]. Caching [GKPS12, JR16, KVS^{+}14, LLD10, PGSD13, RCB11, SNRS12]. Cage [TTB12]. Cage-Based [TTB12]. CageR [TTB12]. Calabi [SLZ^{+}19]. Calculation [FM15, Gov19]. Calculus [GKLS19]. Calligraphy [XWG^{+}13]. CAMA [TWT^{+}16]. Cambial [KSG^{+}15]. Camera [BTS^{+}17b, ESKT15, GL10b, GCW15, GBS19, LK17, LLCZ16, LLB^{+}10, PSHZ^{+}15, SM11, SDHL11, WYD^{+}13, ZZ17]. Cameras [CSC^{+}18, KBKL10, LTX^{+}14, XXY^{+}18, ZSG^{+}18, KBÖ^{+}14]. Camouflage [DJM12]. Can [KP15, KSBC12]. Canada [IC11]. Cancellation [SCCN11]. Cancer [CWS^{+}17, LSS^{+}12, MK11]. Candle [BCRA12]. Canvas [ZLDM16]. Capacity [WGW^{+}19]. Caps [UG18]. Capstone [Grö11]. Capture [AVR10, ACOHS18, CZY11, CWL19, CKHL11, CKH19, DAD^{+}19, DBB^{+}18, DKY16, ESKT15, FHE19, FKR13, FHW^{+}11, FGT^{+}16, HBLB17, KMB^{+}17, KRP^{+}15, NSR17, NVH^{+}13, VF16, WRK^{+}16, WLI^{+}12, WYY13, VB14a].

Captured [CTL13, KZB19]. Capturing [HFM16, LWS^{+}13, PSCN10, SGMG17, WJG^{+}16]. Car [MJK11]. Carbon [RPK^{+}12]. Cardiac [KPG^{+}16, KBvP^{+}17, KGGP18]. Caricature [WL10]. Carlo [BMDS19, BEEM15, BRM^{+}16b, BB17, DWR10, GAM17, GKT16, HCJ13, HD14a, HB19, JEK^{+}19, KS13a, KVS^{+}14, MMG18, MJL^{+}13, MIGMM17, MGN17, NGHJ18, SHZD17, SMJ17, SÖA^{+}19, VAN^{+}19, ZHD18, ZJL^{+}15]. Carotid [DHS^{+}13]. Cars [CCC^{+}14]. Cartogram [AKV15]. Cartograms [CBC^{+}15, NK16]. Cartography [LJLH19]. Cartoons [SSJ^{+}10]. Carving [WPW^{+}11]. Cascaded [SPB^{+}17, WGH^{+}19]. Case [ARLC^{+}13, GKLS19, GBG^{+}14, KASH13, KFA^{+}10]. Cases [RPSF15]. Cast [LK10]. Casted [FKE13]. Casting [KWN^{+}14, XYM13]. Categorical [BTB13]. Category [hKTL^{+}17]. Category-Specific [hKTL^{+}17]. Catmull [BHU10b, Cas12, Csé18]. Causal [BHR17]. Causality [YPF19]. Caustic [GPGSK18]. Caustics [FSES14, GRR^{+}16, LMSG16, PJJ^{+}11]. Cave [SBG17]. Cavities [KFR^{+}11, KKL^{+}16b, SLD^{+}17]. CC [VCRG14]. Cell [ABW^{+}15, AASB14, FKE13, HPvU^{+}16, KER^{+}14, PKE15, PKE17, WH17a, WZL^{+}12, WMRSF15, YBK^{+}12, LCLJ10, ZSW^{+}10b]. Cell-Based [WZL^{+}12, YBK^{+}12]. Cells [HHA17]. Centered [BZL^{+}18, CKS^{+}15, ELM^{+}12, GRC13, HA17, RC18, WTLY12]. Centre [SPT14]. Centred [LW17]. Centres [HSK^{+}10]. Centric [CY14]. Centroidal [CCW12, LLW12, WHWB16]. Centroids [HHA17]. Cerebral [GHB^{+}17, MVB^{+}17, NJB^{+}11, NLB^{+}13, vPGL^{+}14]. CFGExplorer [DI18]. CGF [HGH^{+}11]. CGI [NREM14]. Chain [RKN10]. Chains [EASKC18, RL14]. Challenges [DAF^{+}18, OJMN^{+}19, vLKS^{+}11]. Chamber [SBG17]. Change [JWL^{+}13, SWG16]. ChangeCatcher [LTC18]. Changes [BBL12, GG15, JPK13, OVB^{+}15, PCBS16, WKM15]. Channel [BRB^{+}13, CSŠ18, FMH16]. Channels [CAB^{+}16]. Chaotic [LRB^{+}15]. Character [ARG^{+}16, BV19, CKGC14, ECN14, GP12, GCY^{+}14, HSK14, HLL16, ISYM15, KL14, LCMP19, LWX^{+}15, MYGY19, MKR11, MMG10, PMG13, RPPD17, SLHC12, SBC^{+}17, TAAP^{+}16, ZCK17, CVCH14, CTL13, RTK^{+}14]. Character-Object [ZCK17]. Characteristic [WHT12]. Characteristics [CCH^{+}14]. Characterization [GL12, Gia18, LSS^{+}12, RvdHD^{+}15, SLCO19]. Characterizing [BH19]. Characters [AMYB17, ARG^{+}16, BCB^{+}15, BSAP11, BTST12, FvdP15, ISYM15, JKL18, KL14, LWX^{+}15, PP10, RF15]. Charge [KFM^{+}19]. Chart [CEX^{+}18, CP10, PH17, WPHC16, XBL^{+}18]. Charts [CAB^{+}16, HSBW13, NS11, SK16a, SHK15, SK16b, WPHC16]. CHC [MBJ^{+}15]. Check [LA11]. Checking [OP10]. CheckViz [LA11]. Chemical [BWH^{+}11, BJG^{+}15]. Chinese [TXL^{+}19, XWG^{+}13]. Chip [ZFE16]. Chisel [BH15]. Chromatic [GTL^{+}18]. ChronoCorrelator [vDvdEvW19]. Cinemagraph [BAAR13]. Cinematic [GLCC18]. Circle [MMNG17]. Circular [BCN11a, BSK^{+}13, CLLC15, KBT^{+}12]. Circumgalactic [BAO^{+}19]. Cities [CML^{+}12, IMAW15, KMK12, MPBM^{+}17, MWW12, NJGW15, VKW^{+}12, SKK^{+}14b]. City [CY14, KMK12, KHI^{+}19, LSWW11, STKD12]. Clark [BHU10b, Cas12]. Class [ENMGC19, KP18, SGW12, WLN^{+}17, BMM^{+}15, KP15]. class-specific [BMM^{+}15]. Classic [LVJ10]. Classification [ACC15, BV19, CLCL11, EPAS11, HPMB19, KWD14, KMAB15, LA15, MKP^{+}16, SS15b, WCB15, MPM^{+}14]. Classifications [KK17]. Classifiers [SAMG14, SS15b]. Clean [HL14]. Cleaning [ACOHS18]. Clear [HCBW19, HKL17]. Clever [VGB14a]. Click [SGSP15]. Clicks [OGW19]. Clickstream [DRM19]. Climate [KBL19, PDW^{+}14, BHT19]. Climbing [HBDP17, NBH18a, WC16]. Clipping [BBBV12]. Cloning [ASL^{+}19, TSS^{+}11]. Closed [GAK10, PSP10, TSK14]. Closed-Form [PSP10, GAK10]. Closely [EDPB15, LJL^{+}18]. Closest [AMT^{+}12, AW13, BTP13, SBC16]. Cloth [BYL16, HENfSYS16, HEW15, JCK^{+}13, LeYO^{+}10, LL18, LMMCO17, MYLZ16, MBT^{+}12, TTN^{+}13, TWT^{+}16, WWF^{+}18, WLZ13, YMJ^{+}17, IMDN14]. Cloth-Fluid [HEW15]. Clothing [SOC19]. Clothoid [BLP10]. Clothoids [BD12]. Cloths [BD12, LJK^{+}12, MBT^{+}12, WK12a]. Cloud [LGW18, MC17, RDK13, SY12b, SHS13, YLH^{+}14, YLX^{+}16, ZLYL17, Kur15]. Clouds [BTS^{+}17a, BM12, BM16, CCLN10, CACB18, DSY10, GKOM18, JBG19, KJT14, LIK19, MEMO14, MAM14, PCBS16, QCW^{+}18, RL14, RÖPG18, SPB^{+}19, SYM10, SSG17, SSW14b, TOZ^{+}11, WXL^{+}13, WPW^{+}11, WXW18]. Cloudscapes [WCGG18]. Clue [ZDJ16]. Clue-Map [ZDJ16]. Clues [LA11]. CLUST [ZCT18]. Cluster [AASB19, HV10, LCB^{+}18, STMT12, TPRH11, WLN^{+}17]. Cluster-based [LCB^{+}18]. Clustered [GM19]. Clustering [DCYG19, FKSS13, GMLMG12, HFL12, JGH11, KBL19, LKC^{+}12, LTB19, LWS18, LJBA13, WZL^{+}12, XL10, YYL^{+}16, ZLMM16, vPJtHRV12]. Clusterings [APP10, RL16]. Clusters [CDA^{+}14, MBES16, RL16, ZLMM16]. ClustMe [AASB19]. Cluttered [CKHL11, MPM^{+}14, PGVG19]. CNN [FH18, XWY19]. CNN-based [FH18, XWY19]. CNNs [LSZ^{+}18, YLD^{+}18]. Co [AKM16, CSLG10, GLK18, HFL12, LT17, MRM^{+}18, NNRS15, SCD^{+}16, WFZ^{+}15, YS19, YLX^{+}16]. Co-Alignment [AKM16, NNRS15]. Co-Completion [GLK18]. Co-Dimension [YS19]. Co-estimation [WFZ^{+}15]. Co-occurrence [SCD^{+}16]. Co-representation [MRM^{+}18]. Co-Segmentation [HFL12, YLX^{+}16]. Co-Sponsored [CSLG10]. Co-variations [LT17]. Coarse [CDSS14, CLS16, SY11, SY13, SJ13a, ZWY^{+}13]. Coarse-Graining [CDSS14]. Coarse-to-Fine [SY11, SY13, ZWY^{+}13]. Code [LLC13]. Coded [CKL14, MPCG12, SBB14]. Codes [EPAS11]. Coding [FO12, KMD^{+}17, PHK^{+}10, SHG^{+}16, SGMG17, SO10, vTKP11, LBBC14]. Coffee [MGN17, SHZD17, SMJ17]. Cognition [CWB^{+}14, MSK14, OAJ14]. Cognitive [APM^{+}11]. Coherence [BBT11, JvdGMR19, KK11b, MSW10, MHDG11, MWW12, PC12, SYM^{+}12, XSQ13]. Coherence-Based [MWW12]. Coherence-Enhancing [KK11b]. Coherent [BBH13, BLV^{+}10, CS16, CGS16, EPCV15, GOPT11, GRDE10, KTO11, KER^{+}14, LAE^{+}12, LHNS18, MO10, MBES16, MBJ^{+}15, PPM^{+}16, RSD^{+}12, SW10, TCM10, ZAD15, SPCR14]. Cohort [RCMA^{+}18]. Collaboration [IC11, ID10]. Collaborative [BB17, JSH^{+}13, MHK^{+}19b, RDK13, TSPP16, XBL^{+}18]. Collage [GTZM10, ZH12]. Collection [CDSS14, SBC14]. Collections [AEWQ^{+}15, AKZM14, AKM16, CWG11, CRA^{+}17, DMS14, FDH^{+}15, HG13, KGAC15, KFLCO13, LLC11, LCUR14, LVW^{+}15, NBCW^{+}11, NNRS15, OSR^{+}14, PTT^{+}12, PDC^{+}19, RLGH15, SGB13, YXX14, ZCOAM14, vdCvW17, HAGO19]. Collective [XWY^{+}15]. Collision [FWPS11, JKJL18, LMM10, LPH^{+}15, LLC^{+}15, PKS10, SR19, TWT^{+}16, VMTS10, WTTM15, WLT^{+}17, WTMT18, WDZ17, WC14, YMJ^{+}17]. Collision-Aware [JKJL18]. Collisions [CK10a, PKS10, VMTS10]. collocated [RAmB^{+}19]. Color [AP10a, AAB^{+}10, AGJ12, BMS^{+}10, CKL14, FLJ^{+}14, GRC13, HGO18, HHGJ15, HZMH14, JYS19, JCW11, JESG12, KYKL14, KM16, KG18, LLC19, MO10, NKB14, NPCB17, OVB^{+}15, RMZ13, SO12, SSB^{+}14, TBKP12, TDMS14, WWL^{+}19, WK12b, WZL^{+}17, WTLL13, XADR13, YXX14, YWF^{+}19, YLL15, EKB14, HL14, RCM^{+}14]. Color-Pair [YWF^{+}19]. Colored [PK10, FLJ^{+}14]. Colorfields [SSSG16]. Colorimeter [GBS19]. Coloring [NPCB17, WZL^{+}17]. Colorization [LT12, JCT14]. colormaps [EKB14]. Colors [HZMH14, KNL^{+}15, KSG^{+}15, LFK^{+}13, SGSP15, YLL15]. Colour [AP10a, FPC^{+}16, LM10, LCG10, LKSD17, LS15, WDK^{+}13, XWL^{+}15]. Colourization [DRA10]. Colours [MSHD15, OYH18]. column [LLLY19]. Combination [BP17, WMXC19]. combinations [SNJ^{+}14]. Combinatorial [PBPP11, SY11]. Combined [WYKR17]. Combines [ARH12]. Combining [GLCC18, KP18, KKTD17, SJ19]. Combustion [GBG^{+}14, JK13, OLF^{+}14, OAM^{+}18]. Comfortable [FFN18]. Comics [CNKI13, McC11]. Comments [SHS^{+}17]. Communication [SMM13, YKS^{+}19]. Communicative [HRD^{+}15, SBD^{+}15b]. Communities [CDA^{+}14, VBAW15]. Community [DWT^{+}11, RTJ^{+}11]. Commutativity [NO17]. Compact [CCC17, CP11, GLLR11, KBK^{+}10, UKCB15, HLJ^{+}13]. Compacting [HL19]. Compaction [HLJ^{+}13]. Compactly [HA17]. Comparative [ASB^{+}17, ARH12, BvLBS11, DHS^{+}13, FL19, GT15, KARC15, MVB^{+}17, MWS^{+}10, OSR^{+}14, PPBT12, RLP10, SKR^{+}14, SS15b, vPGL^{+}14, EMU17, WGO^{+}14]. Compare [AKMM11]. Comparing [RL16, SSK16, WPHC16, vdCvW17]. Comparison [BSBE17, DWT^{+}11, FAFM19, HFM16, KWS^{+}15, KZZM12, MTM12, NNN11, PDW^{+}14, SSW14a, VCRG14, VHB16, WHS^{+}18]. Comparisons [SLB^{+}18]. Compatibility [LTR19, ZCOM13]. Compensated [PM19]. Compensating [MSK14]. Compensation [ENSD12, LMG^{+}18b, MK15, SYC10]. Competitive [KMJE12]. Compiler [CTHAM10]. Complete [AH11]. Completion [GS14, GLK18, HTG14, IS15, LL19, LZL^{+}15, MSAP15].

Complex [ASB^{+}17, BWH^{+}11, BHMT13, CG16c, DVPSH14, GKPS12, HPH10, KKSS15, KFM^{+}19, LDY10, MMG18, McD10, MPBM^{+}17, NSRS13, PMD12, RKN10, SCCN11, WBCGH11, DGR^{+}14, GSC18, LN17]. Complexes [AASB14, SN12, DFIM15, LCLJ10, WIFD13]. Complexity [JS10, LFS^{+}15, SvW13, TWD^{+}13, Tim12]. Compliant [LAS^{+}11]. Component [YKS^{+}19, ZFCO^{+}11, ZXTD10, BTST12, Lie17]. component-based [Lie17]. Component-wise [ZFCO^{+}11]. Components [SBL12]. Compositable [LWL^{+}16b]. Composite [BPV18, CKS^{+}16, GSDC17, PTO10, SHF13]. Compositing [DZC11, EKFM12, GWO^{+}10, GTZM10, LCWCO10, RLH^{+}17b, RWSG13]. Composition [GLGW12, LCWCO10, LMLF15, SML15, TXGW19, WTLL13]. Composition-Aware [LMLF15]. Compound [RMF12]. Compoundly [CD10]. Comprehensible [KGAC15]. Comprehensive [FHW^{+}11, PRS15, YZL17, ZWJ^{+}19]. Compressed [And10, NVT^{+}14, RGG^{+}14, SBE16b, XSE14]. Compressing [VMHB14]. Compression [AMAM13, And12, CH11, DKB^{+}16, HML^{+}19, HFM10, JBG17, JKJL18, KGR^{+}16, LPG10, MP12a, PM19, RJGW19, SBE16a, VS10, WDR11, BHT19, CVDL16, DCV14, GGM12, PHM^{+}14, VB14a, VD18]. compression-domain [GGM12]. Compressive [JCK16, LA15, MKU15, SD10b]. Computation [HA19, Hol15, KS11, Pat17, SKZ13, SN12, MVPG11]. Computational [BS12a, BRWM18, DBS^{+}18, DRF12, EMP^{+}12, EWK^{+}13, FO12, GTM^{+}12, HKW12, Jac17, LLW12, MCM^{+}12, MP12b, RE12, SLE17, She12, STBG12, TSW^{+}19, WW16, WILH11, YZC18, ID10, IC11]. Computations [RPWO19]. computed [LWLD11]. Computer [ALCS18, BET14, BMO^{+}14, CTHAM10, DCGG11, GSHM10, GMD10, GP16, KH19, KWC^{+}12, KBKL10, LF11, LF10, MWN^{+}17, MSS^{+}10, OP10, PTO10, PCR11, SDA^{+}18, SRH^{+}11, ZWJ^{+}19]. Computer-assisted [ZWJ^{+}19]. Computer-Suggested [SRH^{+}11]. Computing [APH^{+}12, BAA^{+}16, CVJ15, CYY^{+}11, CSaLM13, HS17, HR10, JSH^{+}13, NB12, NBHN17, RK10, SRM^{+}19, SD10b, WMZ12, WT11, YFL19, YL11]. ConcaveCubes [LCB^{+}18]. Concentric [XLH^{+}13]. Conceptualizing [DCK12]. Concurrent [KBLE19, YHGT10, ZFE16]. Condenser [LDW^{+}10]. Condenser-Based [LDW^{+}10]. Condition [KS14, OHG11]. Conditional [CLGS18]. Conditions [PCDS12, RGSK10, TDNL18, SZ18]. Cone [CNS^{+}11, KB12]. Cones [WLT^{+}17]. Configurable [SSSG16]. Configuration [LT17]. Confinement [HL13]. Confocal [SPB^{+}17]. Conformal [CG17, SBC16, SLZ^{+}19, WLZT18, YMYK14]. Congestion [BJB^{+}18]. Congruences [WJB^{+}13]. Conjoined [CLJ^{+}15]. Connected [CPZ^{+}15]. ConnectedCharts [VM12]. Connecting [OM13, OW19]. Connection [NB15, YXX14]. Connections [CDS10, PRDD15, ZLZ^{+}18]. Connectivity [KLS^{+}17, PW13, SSE^{+}14, SMB^{+}17, VR12]. Connolly [MJK17]. Conoid [KTN10]. consecutive [DMS14]. Consensus [MW18b, RBC14, ZST^{+}10, HDV19]. Conserving [dFH^{+}11]. Considerations [KK17, KSS^{+}15]. Considering [GLGW12]. Consistency [LFYX19, NPCB17]. Consistent [ADG19, BIZ18, BTS^{+}17b, BSL18, CRA^{+}17, FACO17, FNH^{+}17, GLCC17, GG15, HG13, JBG19, PSDB^{+}10, SJP^{+}13, SCF10, TL19, WKBB18, GFK^{+}19]. Consolidation [RÖPG18, WXL^{+}13]. Constant [CRGZ10, GUS12, Szy11, SBL12, WDZ17]. Constellations [XBL^{+}18]. Constrainable [JCK^{+}13]. Constrained [AMR^{+}17, ATK17, ATF12, BHMT13, CCW12, CG12, DBD^{+}13, GBKS18, LHNS18, MHA17, RKSA17, SC10, YFL19, YLG^{+}18, ZFA^{+}16, ZXZ^{+}17]. Constraint [MC19, VRBC17]. Constraints [HK12, HBH18, RLGH15, SJ13b, TSK14]. Constructing [BW13, IEK^{+}14, KK17, KLC^{+}15, WXW18, PCBL16]. Construction [AYLM13, BLD14a, FBL16, Gro16, HMB17, HHH12, KWS16, KS18, LCSS19, LA13, LCY^{+}11, QYZ17, SE19, SFL19, SJWS13, VKJ^{+}17, YHGT10, ZCOM13, DGR^{+}14, GD16]. Constructive [ARRO^{+}17]. Consumer [LK17, XXY^{+}18]. Consumer-level [XXY^{+}18]. Contact [HENfSYS16, HEW15, KL14, TWT^{+}16, ZLM^{+}15, dSNV^{+}17, TAAP^{+}16]. Contact-Aware [TWT^{+}16]. contacts [TSK14]. Content [ARM^{+}15, DER^{+}10, FW17, KLW12, LMLG15, LSWW11, MBRHD18, PSP^{+}14, SGMG17, WDK^{+}13]. Content-Aware [KLW12, PSP^{+}14]. Content-Based [WDK^{+}13]. Content-Independent [LMLG15]. Context [BAO^{+}19, BTM^{+}19, GRE11, HVP^{+}19, KDCM14, KGRG17, LCSCO10, MbMYR15, OJS^{+}11, SEI10, XXM^{+}13, ZCL18, ZLDM16, vTKP11, BCN11b]. Context-Aware [GRE11, KDCM14]. Context-Based [XXM^{+}13, vTKP11]. Context-Dependent [LCSCO10]. Context-inferred [ZLDM16]. Contexts [RHM^{+}12]. Contextual [JKK^{+}18, KSP^{+}18, MGB14]. Contingent [SGEM16]. Continuity [GP16, SYT^{+}13, WGS10, SZ18]. Continuous [BPFG11, COC15, GAM17, HBW11, LPD14, LMSF19, MFL13, PJR^{+}14, PSO18, RJT18, SFFP15, WTTM15, WT11, WC14, vdZLBI11, CH12, TWSM17]. Continuum [SWML10]. Contour [IY10, IYS^{+}13, NBHN17]. Contour-based [IY10, IYS^{+}13]. Contouring [MSS11, SSO^{+}10]. Contours [BSL18, FKRW16, Fre18]. ConToVi [EAGA^{+}16]. Contracting [LW17]. Contraction [GHB15]. Contrast [HW16, HDL11, LM10, LM15, MO10, MSK14, MK15, NWHWfD16, ZHLW18]. Contrast-aware [LM10]. Contrast-Enhanced [LM15]. Contribution [Can11]. Control [BV19, DI18, FLJ^{+}14, GLCC17, HENfSYS16, HL18, KPRN11, KPD10, LK17, LTKD15, LYG15, LLC^{+}15, LCY^{+}19, RCMM^{+}16, RPPD17, SBE19a, SHW^{+}18, SKMS18, VVE^{+}10, VSK16, WK12a, dSNV^{+}17, AFHdL14, HNJ^{+}14, RTK^{+}14, VD18]. Controllable [AP10a, BP19, CBTB16, GMM15, HL13, RHLH18]. Controlled [JWS12, KSG^{+}15, KFA^{+}14, ZSL^{+}17]. Controller [FvdP15, GCY^{+}14, HKSK18]. Controllers [WPP13, ZFCO^{+}11]. Controlling [LDG19, YYZZ18]. Controls [GLCC18]. Convergence [LPH^{+}15, SMJ17]. Conversation [EAGA^{+}16]. Conversational [EASKC18]. Conversations [EASG^{+}17, HC14]. Conversion [BCK^{+}12, MEKM17, CHA^{+}14]. Convex [AGCO13, GCW15, MHA17]. Convey [KCH19]. Conveying [ARM^{+}15, BJG^{+}15, RRS12]. ConVis [HC14]. Convolution [LKEP14, MESG11, SFFP15, TLH19]. Convolution-Based [MESG11]. Convolutional [ACA18, hKTL^{+}17, KG19, LLLY19, LIK19, MBRHD18, NAM^{+}17, RÖPG18, SHG^{+}16, SGMG17, YSM19, YYZZ18, BMM^{+}15]. Convolutions [WLZX19]. convolved [SNRS12]. Cooperation [LSF^{+}11, PS10, DS11a]. Coordinate [KZZM12]. Coordinated [BSK16, BRM^{+}16a, HSH16, RSM^{+}16, SS16]. Coordinates [CGH18, DCK12, DVPSH14, GRPF16, HBW11, KARC15, MS10a, MR12, NVT^{+}14, PR19, Rus10, SHF13, SKPSH13, VB14b, WLN^{+}17, WPG12, YS19, dGLB^{+}14, YI10]. Copters [NKSI16]. Copula [ÖKB10]. Copula-Based [ÖKB10]. Core [BLD14a, Bik12, DC10, GG14, IABT11, KTO11, LCDW16, ORT18]. CoreFlow [LKD^{+}17]. Corner [BdM14, SNA17]. Corners [BW13, CSŠ18]. Coronary [GOH^{+}10, SSM12]. Corotated [KKB18]. Corotational [CFS14]. Correct [Gro16]. Correcting [CWGvW19, EPAS11]. Correction [DBB^{+}18, UBH14, WYY13, XWY19, ZNZ19]. Correctness [OP10]. Correlated [SJWS13, VR12]. CorrelatedMultiples [LHNS18]. Correlation [AWCO10, GHX^{+}17, KWS^{+}15, LWS18, PPH12, PW12, RB10, SBLC17]. Correlation-Based [AWCO10, KWS^{+}15]. Correlations [FKRW16, FSTR13, RÖG17, ST18, SÖA^{+}19]. Correspondence [ATCO^{+}10, BS12b, COC15, ELC19, EHA^{+}19, GSTOG16, HMW^{+}15, KBW^{+}12, KKBL15, KO19, LRB^{+}16, LCY^{+}11, LKF12, RCB^{+}17b, RMC17, SY11, SY13, SY14a, SY14b, SNB^{+}12, WLZT18, YYL^{+}16, YSC^{+}18, vKTS^{+}11, vKZHCO11]. Correspondences [DZD^{+}16, GSTOG16, GSDG18, GLHH13, KBW^{+}12, KBB^{+}13, KSKL13, LK19, LKF12, PBB^{+}13, SCF10, TMRL14]. Cosine [RC18]. Cosine-Weighted [RC18]. Cosmetics [KK18]. Cosmic [BAO^{+}19]. Cosserat [SMSH18]. Cost [ACKM16, ESKD14, SB19, ZFE16]. Count [KKR18]. Countershading [TMHD12]. Coupled [ADG19, EHT18, HSK14, KBB^{+}13, dHvPJV14, AKZM14]. Coupling [Aan18, SZMTW15, TL19, YLHQ12]. Course [Duc14]. Cover [ASW14, HGH^{+}11, MNP^{+}17, SJ13c, ZJ18, vFG11, NRP11]. Covered [CEG^{+}18]. COVRA [GGM12]. CPH [UKCB15]. CPU [ÁSK14, JZYP18, PKS10, PBPP11, SR19]. CPU/GPU [PKS10, PBPP11]. CPUs [IABT11, KWN^{+}14]. Crack [CBV^{+}14, DC10]. Crack-Free [DC10, CBV^{+}14]. Cranial [HRRR18]. Crawl [PM16]. crdbrd [HBA12]. Crease [BAT11]. Creases [KSD14a, KSD14b]. Creating [FDH^{+}15, GKB12, GRP10, OAO11, RGM^{+}18, ZYF10]. Creation [KrJC^{+}11, LCCC13, WLL^{+}17]. Creative [APH^{+}12]. CRFs [CPZ^{+}15]. Critical [GST14, LS16, WRS^{+}13]. Cropping [SML15]. Cross [DMNV12, HHCJ18, HHD^{+}12, KBT^{+}12, SBC14, ZT10]. Cross-Boundary [ZT10]. Cross-Collection [SBC14]. Cross-Modal [HHD^{+}12]. Cross-Sections [KBT^{+}12, DMNV12, HHCJ18]. Crossing [SAAB11]. Crowd [BSK16, BPA16, CKGC14, CC14, JVS^{+}12, JPCC14, KWD14, LCSCO10, LLD10, LCM^{+}18, SHW^{+}18, XWY^{+}15, ZCT18, DMCN^{+}17, LJK^{+}12, WGO^{+}14]. Crowd-sourced [KWD14]. Crowded [GS14, GTK^{+}12]. Crowds [BD12, LJK^{+}12, MBT^{+}12, WK12a]. Crowdsourcing [BMB^{+}18, OJ15]. Cryptography [CJFH14, LMHH14, SY14b, SGS14, ZZCJ14]. Crystal [RHLH18]. CSG [WXW18]. CT [DHS^{+}13, SSM12]. Cube [BW13, NRP11]. Cubemap [XTZ19]. Cubes [BDA^{+}17, PWH11]. Cubic [AASB14, HA17, RC18, KP11, KP15]. CUDA [Ros13]. Cue [KLK17, RHL12]. Cues [PD16]. Culling [AMTMH12, GL10a, GPP^{+}10, GRDE10, KBK^{+}10, MBJ^{+}15, TH17, WLT^{+}17]. Cultural [AJL^{+}11, HBRW^{+}12, PPY^{+}16, SCP^{+}17]. Cumulus [ZLYL17, YLH^{+}14]. Curation [YDP19]. Curl [RLH^{+}17b]. Curl-free [RLH^{+}17b]. Curvature [CBLW19, CK11b, KSBC12, SYT^{+}13, SZ18, TAOZ12, VVP^{+}16, ZWC^{+}10, MM18]. Curve [GLK16, GLK18, GBKS18, HHCJ18, LH11, LZW^{+}13, LW17, LW18, OMW16, PM16, SV14, TBTB12]. Curved [DHP^{+}19]. Curves [AJA11, AJC11, BLW11, BvTH16, CGS16, DSL15, GLK16, GKKT13, Jes16, KSD14b, Kur15, LTKD15, LKF12, MS13, PJSH15, VSK16, WHT12]. Curvicircular [MMV^{+}13]. Curvilinear [AO13, ZLL13]. Curving [KPD10]. CustomCut [GLX^{+}16]. Customizable [BP13]. Customized [GSE^{+}14a, GLX^{+}16]. Cut [Jac19, WMRSF15]. Cut-Cell [WMRSF15]. Cutaway [LMS^{+}16]. Cutout [HZZ11, TZD11]. Cuts [LDdLRB16, LVJ10, WWD15]. Cutting [FLL11, HL14, JZYP18, RCM^{+}14, RLYL14, ZWC^{+}10]. Cuttlefish [WWL^{+}19]. CV3 [FAFM19]. CVs [FAFM19]. Cyber [MSFM16]. Cycle [BFG^{+}17, GFK^{+}19]. cycle-consistent [GFK^{+}19]. Cylinder [Gam16]. Cytosplore [HPvU^{+}16].

D [AKS^{+}19, CHWL19, DS11a, EHH19, FHE19, HAGO19, JL19, KPG^{+}16, LSF^{+}11, LLLY19, LTR19, NSC14, PS10, WMXC19, ADMAS18, ASH15, Att15, ATF12, AH11, AKM16, ARG^{+}16, BSW^{+}12, BIZ18, BSY^{+}19, BCBB16, BCK^{+}12, BLS^{+}17, BHMT13, BdM14, BL18, CJXH17, CS16, COO15, CVCH14, CRC^{+}15, CYY^{+}11, CNKI13, CACB18, DDÖ^{+}17, DGR^{+}14, DER^{+}10, DMS14, FW17, GS14, GBS19, GST14, GRT14, GT16b, GLX^{+}16, GTB14, HT11, HCGW14, HMTH13, HRRR18, HL14, HL15, HMW^{+}15, HBO^{+}10, HYL^{+}15, HRS18, HFL12, HKM15, HCSC16, JWH19a, JH15, JJKL18, JLW10, KZB19, KKBJ16, KFW19, KLTZ16, KCL^{+}18, KG19, KCH^{+}18, KGP^{+}12, KBvP^{+}17, KGGP18, LT17, Lav11, LCSL18, LOM^{+}18, LJB^{+}12, LDGN15, LWL^{+}16a, LWS18, LGK16, LAFT12, LTX^{+}14, LEM^{+}17, LN17, MHS^{+}14, MGY^{+}18, MCHW18, MBES16, MGG10b, MbMYR15, MCG^{+}19, MKP^{+}16, MPWC13, MEKM17, NWHWfD16, NW17]. D [NRM^{+}12, NRP11, NREM14, NBHN17, OM13, OW19, ORT18, OGHT10, PBMG15, PSCC18, PB11, PCX^{+}18, PW12, PPBT12, PEP^{+}11b, PEPM12, PGK10, POG13, PBC^{+}16, RXX^{+}17, RI17, RAMG15, RGG18, RBRY19, RLYL14, SLB^{+}18, SW10, SY14a, SYM10, Sch11, SK16a, STKD12, SXY^{+}11, SN12, SK17, SGS14, STC^{+}16, SSS^{+}12, STP17, SJW^{+}11, SJWS13, SBL12, TDS^{+}16, TWS^{+}11, TSW^{+}19, TTB12, TBTB12, TWSM17, VGB^{+}14b, WH17a, WL10, WZL^{+}12, WRS^{+}13, WCT^{+}15, WZK16, WLL^{+}17, WLJ^{+}18, WBAI19, WGS10, WSSC11, WXW18, XXLX14, XXS^{+}15, XSQ13, XSX^{+}14, XXY^{+}18, YFW12, YLLL15, YHL^{+}16, YL11, ZZH15, ZZWC16, ZFG^{+}17, ZRLS19, ZKWG16, ZWHK16, ZTG^{+}18, ZSG^{+}18, ZJC13, dGCSAD11, dHvPJV14, vPJtHRV12, vMRBPM17]. D-Printing [KFW19]. Dashboard [MSFM16]. Dashcam [CCC^{+}14]. Data [AKMM11, ARH12, AFK^{+}14, AAS^{+}16, AECOK16, BDA^{+}17, BMH^{+}12, BLY^{+}11, BWH^{+}11, BSW^{+}14, BHR^{+}19, BSY^{+}19, Bik12, BKR^{+}17, BBBL11, BBL12, BvLBS11, BTB13, CGT^{+}15, CGM19, CKGC14, CC14, CLJ^{+}15, CWL^{+}15, CYI^{+}12, CNKI13, CWM19, CDS16, CMF18, CKE^{+}12, CKS^{+}16, CR16b, DKG15, DAF^{+}18, DRM19, DH16, DPD^{+}15, DMSL11, DBS^{+}11, Duc14, EAGA^{+}16, EGG^{+}15, EIKM16, EHH^{+}13, ECN14, FR11, FDL14, FL19, FMH16, GLHH13, GCLX17, GLK18, GHB^{+}17, GCZ^{+}12, GKPL11, GRPF16, GSW12, HENfSYS16, HSK^{+}10, HJM^{+}11, HVH^{+}16, HSmCY13, HSH16, HPvU^{+}16, HV10, HPH10, HLJ^{+}13, HJS^{+}17, IS15, KWD14, KZB19, KCJM16, KFH10, KKS^{+}12, KH18, KMJE12, KCA^{+}16, KPG^{+}16, KBvP^{+}17, KVD^{+}10, KFR^{+}11, KHI^{+}19, KTW^{+}13, KZZM12, LDB11, LSBP18, LMS^{+}16, LKZ^{+}15, LT16, LOM^{+}18, LSS^{+}12, LCB^{+}18, LMG^{+}18a, LFK^{+}13, LFS^{+}15, LJH13, LWBP14, LWT^{+}15, LBJ^{+}16]. Data [LCDW16, LGW18, MK11, MHK^{+}19a, MHK^{+}19b, MRL^{+}17, MKP^{+}16, MKSS12, MBT^{+}12, MLD^{+}18, NNN11, NJB^{+}11, NHG19, OJMN^{+}19, OGW19, POS^{+}11a, PSPM12, PLL11, PEP^{+}11a, PVS^{+}18, PHL^{+}16, PSC10, PEP^{+}11b, PEPM12, PDW^{+}14, Pos11b, PKRJ10, PBC^{+}16, RW18, RXX^{+}17, RTK^{+}14, RL16, RPLH11, RPMO13, RBRY19, RSSL17, RSK12, SSW14a, SSKB15, SML15, SCD^{+}16, SBG17, SPB^{+}17, STMT12, SA15, SY12b, SHW^{+}18, SV10, SAAF18, SK16b, SvL16, SBS^{+}17, SWG16, SGG15, SGG16, SLSG16, SRG^{+}19, SSS^{+}12, SSSG16, TFA^{+}11, TWC^{+}16, TLFC16, TCM10, TPRH11, VSG^{+}13, VB14a, VF14, VM12, WHC15, WDM^{+}12, WYZC13, WXR^{+}16, WLL^{+}17, WLN^{+}17, WBAI19, WG11, WBS^{+}13, WHP^{+}11, WGO^{+}14, WCH^{+}15, XSE14, XKHK17, YWS^{+}14, YKS^{+}19, YNM^{+}13, YLRC10, ZFAQ13, ZFA^{+}16, ZLW^{+}16, ZAM^{+}16, ZM16, ZCW^{+}19, ZZ17, vDHO16, vGPNB17, vdEvW13, vdCvW16, vdZLBI11, BHT19, CFGL16]. data [KGGP18, NDD14, RK10]. Data-Driven [CKGC14, CC14, ECN14, GLHH13, GCLX17, GCZ^{+}12, HSmCY13, IS15, MHK^{+}19b, MBT^{+}12, RXX^{+}17, RBRY19, SML15, SHW^{+}18, WLL^{+}17, XKHK17, ZLW^{+}16, ZZ17, AECOK16, CLJ^{+}15, EIKM16, HPH10, LOM^{+}18, SA15]. Data-guided [HENfSYS16]. Data-Parallel [MKSS12]. Database [LDGN15, OAO11, SK17, LTR19]. Database-Assisted [LDGN15].

Databases [MHHH15]. Dataflow [VOS^{+}10]. Datasets [AHG^{+}19, HPvU^{+}16, YDP19, ZC18]. Deblurring [BCN11a, CCTL12, EBC17, MPCG12]. Debugging [HSH16]. Decadal [KBL19]. Decades [MLP^{+}10]. Decals [dGWB^{+}14]. Decay [KRB11]. Decimation [SLA15]. Decision [BMPM12, CCH^{+}14, CWS^{+}17, KMJE12, LCP^{+}12, WKS^{+}14, WSK^{+}19]. Decision-Making [WSK^{+}19]. Decoder [LIK19]. Decolorization [YLL15]. Decomposing [IRWM17, KG18]. Decomposition [AGCO13, BÖK11, BMB15b, CG19, DSH^{+}17, DJM12, ERHH11, HYZ^{+}14, JTRS12, LJKL17, LVV18, MP12a, NKLN10, SSW14a, SJF11, TLRB18, WLZH17, WLX19, YL11, ZT10, ZRJ^{+}15]. Decompositions [BKPB17, Szy11]. Decompression [JBG17, MKSS12, OBGB11]. Deconstructed [SCM^{+}19]. Decorative [STG16]. Decoupled [GTG17, WWT^{+}16, ZM16]. Deep [ASL^{+}19, APH^{+}12, BMDS19, BM16, CWGvW19, CWL19, DAD^{+}19, EIKM16, ESKBC17, Fre18, GZH^{+}19, HGO18, HMP^{+}12, HMRR19, KR19, hKTL^{+}17, KKK18, KSP^{+}18, KAT^{+}19, KKR18, LVV18, LGK16, LPSB18, MBRHD18, NAM^{+}17, VAN^{+}19, XHW^{+}18, ZLZ^{+}18, HKM15, NW17]. deep-learned [HKM15]. Deep-learning [HMRR19]. DeepGarment [DDÖ^{+}17]. DeepProp [EIKM16]. Defect [CK11a]. Defect-Tolerant [CK11a]. Defects [AAS^{+}16]. Deferred [ENSB13, SBF15, TBP18]. Defiltering [WGH^{+}19]. Definite [KASH13]. Defocus [JKL13, KSP^{+}18, MPCG12, MTAM12, MVH^{+}14, WWT^{+}16]. Deformable [AWO^{+}10, BvTH16, CKHL11, FMD^{+}19, GGV^{+}19, HK16, JZYP18, KKBJ16, LPH^{+}15, MMO16, PKS10, PB11, RBC14, SKR^{+}14, SSB13, SCF10, WC14, WWD15, YYG18, BMM^{+}15]. Deformation [BD12, CRY11, CGS16, ELC19, GSZ11, GSDC17, HK16, HYZ^{+}14, ISYM15, JWS12, JZW14, KKBJ16, KS14, KSO10, KWW^{+}14, LJK^{+}12, LLC10b, MS11b, MBT^{+}12, MWW16, PWS12, RKSA17, SKNS15, SK17, SBCBG11b, SVWG12, UBH14, VRBC17, WK12a, WDAH10, WWH^{+}14, YLHQ14, ZXTD10, GFK^{+}19, RRS12, TAAP^{+}16]. Deformations [FSTR13, FB11, HCW17, NVH^{+}13, SBO18, SHF13, SKPSH13]. Deformed [PTW13, SLHC12]. Deforming [AW13, GB10, CH12]. Degenerate [CFS14]. Deghosting [TAEE15, TAEE16]. Degradation [HM15]. Degree [SS15a, SHD16]. Delaunay [CCW12, DLS10, DS11b, SSE^{+}14, ZSW^{+}10b]. Delay [WBFvL17]. demand [GLX^{+}16]. Demonstration [SE19]. Demosaicing [HGO18]. Dendritic [RHLH18]. Denoising [BRM^{+}16b, BB17, EBC17, HDL11, KS13a, LZFH18, MC17, MJL^{+}13, MVH^{+}14, RDK13, RMZ13, VAN^{+}19, WZCF15, ZWY^{+}13]. Dense [DRW15, LK19, NJC^{+}19, SY11, ZLZ^{+}18, CSFP12, WHD17]. DenseCut [CPZ^{+}15]. Densely [CPZ^{+}15]. Density [DWR10, ERA^{+}16, EHT18, FM15, GUS12, HGNH17, HET12, LH11, LGW18, RÖG17, WZL^{+}12]. Density-based [DWR10]. Dependency [DG12]. Dependency-Free [DG12]. Dependent [BPKB14, GPK^{+}12, GKKT13, HGRS^{+}17, HS19, KBK13, LCSCO10, LRB^{+}15, NPDD11, SW10, SW17, SBD15a, vDHO16]. Depiction [BCRA11]. deposition [RCM^{+}14]. Depth [AMTMH12, AMAM13, BEJM15, BTS^{+}17b, CWW^{+}11, CCC^{+}14, FO12, HM15, IEK^{+}14, JSLW14, KRMS13, LK17, LEE17, LFS^{+}15, LCD10, LTX^{+}14, MRD12, NSC14, NDD14, PHE^{+}11, RWW16, RHL12, SGM^{+}11, SSJ^{+}10, XXS^{+}15, XXY^{+}18, YWY10, PHM^{+}14]. Depth-and-Normal [JSLW14]. Depth-augmented [HM15]. Depth-of-Field [CWW^{+}11]. Deraining [LQS^{+}19]. Derived [SLSG16, ACAA^{+}19]. Description [GUS12]. Descriptions [JS10, SKSS14]. Descriptive [BDA^{+}17]. Descriptor [AKM16, BvLBS11, MGG10b, NO17, ZYF13]. Descriptors [BMM^{+}15, BMR^{+}16, COO15, HKM15, RBRY19, RÖM^{+}15, LBBC14]. Desertscape [PPG^{+}19]. Design [AWO^{+}10, ADN^{+}17, BRM^{+}16a, BFR17, BLS^{+}17, CE19, CKE^{+}12, CCH^{+}14, FvdP15, FKR13, GEY12, GLX^{+}16, HRS^{+}16, HGO18, HSH16, Hua17, IMAW15, Jac17, Jac19, KGP^{+}12, KW18, LT17, LBRM18, LZY^{+}17, LLM^{+}17, LHH^{+}13, MKMA19, MKR11, MJK11, MK15, MEKM17, NKSI16, NW17, NBA19, OPP10, RSM^{+}16, RLGH15, SS16, SLE17, SGS18, SH14a, SH14b, SPB^{+}17, STG16, SP13, SA15, SPF^{+}19, SHK15, STBG12, SMS^{+}17, TSW^{+}19, UMK19, VCD^{+}16, WHC15, WZL^{+}12, WK12b, WTH^{+}13, WFLW18, WBFvL17, YCXW17, ZLW^{+}16, ZFE16, ZLDM16, vdKdJP^{+}19, CLGS18, KP11, MVLS14, SBD^{+}15b, YZC18, XXM^{+}13]. Design-to-Assembly [Jac19]. Designing [DI18, DVPSH14, GDAU14, GCW15, KCH19, SS16, STG16, SPV^{+}10, TLM16, ZSW10a]. Designs [ZCL18, RLYL14]. Desktop [JSH^{+}13, MCG^{+}19]. Destination [BBBL11, ZFA^{+}16, ZSJT19]. destinations [vDHO16]. Detail [BHS^{+}17, BCN11b, DGP17, GRE11, GP18, HREB11, JKL13, KDCM14, KBK13, LPD14, OJS^{+}11, PJR^{+}14, SBD15a, SRK13, WT11, WYY13, ZCH^{+}17, ZBM^{+}17, SLKL14]. Detail-Preserving [DGP17]. Detailed [NJGW15]. Details [BDC18, GD10, HK16, ZHLW18, ZYF10, GGG^{+}16b]. Detangler [RMM15]. Detected [PSCC18]. Detecting [WKM15]. Detection [ARRO^{+}17, ABW^{+}15, BMH^{+}12, BPW14, BFG^{+}17, FWPS11, GL12, JL17, KZB19, KBWS13, KSP^{+}18, KLAB15, LMW^{+}19, LPH^{+}15, LLC^{+}15, OLA16, PKS10, PCBS16, RBC14, SCN^{+}16, SR19, SAD^{+}16, SLD^{+}17, SGS14, SBM^{+}14, SJWS13, VMTS10, WTTM15, WTMT18, WMXC19, WDZ17, WC14, YMJ^{+}17, ZDM^{+}14, ZVE^{+}14, MPM^{+}14]. Detector [CWL^{+}15]. Determining [OM13]. Deterministic [AGR19]. Develop [SJL15]. Developable [SVWG12]. Developing [GKLS19]. Development [LTC18]. Devices [KKTD17, RKR^{+}16, UMM^{+}10]. Dexterous [BYL16]. Diagnosis [ZVE^{+}14]. Diagram [EDPB15, Man16, QYZ17, QCW^{+}18]. Diagrams [BPY16, DvKSW12, GEY12, HHA17, HET12, LBK14, MWTI19, NL13, NB12, XLS^{+}14, SNA17, SBD^{+}15b]. Diamond [WD11]. Dichromats [MO10]. Dictionary [LOM^{+}18]. Dictionary-based [LOM^{+}18]. Difference [GG15, HCO18]. Differences [HAGO19, PSCC18]. Different [AMYB17, ARM^{+}15, CCH^{+}14, HV10, MBDC15, SSM12, PCBL16, RAmB^{+}19].

Differential [BM16, BHU10b, CDS10, FSES14, JTSZ10, MS10a, Rus10, VVP^{+}16, YI10]. Differentials [EBR^{+}14]. Difficulty [ZDJ16]. Diffraction [DTS^{+}14, SLE18]. Diffusion [BS12b, BMR^{+}16, BLW11, HCJ13, HHA17, Jes16, KSW^{+}12, KPS^{+}14, PJSH15, ZCH^{+}17, vFG11, HGA^{+}10]. Diffusion-Based [vFG11]. Diffusive [HWK15, KK18]. Digital [CG16b, DJM12, GGP^{+}19, GBS19, JFCS17, JSH^{+}13, SCP^{+}17, VGB14a, ZH12, CFGL16]. Dihedral [VR12]. Dimension [ATW15, FR11, LF10, MDI^{+}18, YS19]. Dimension-reduced [ATW15]. Dimensional [ABD10, BTB13, FR11, FDL14, GHWG14, JH19, KKS^{+}12, KZZM12, LZQ13, LWBP14, LBJ^{+}16, LHNS18, MC14, PSPM12, PHL^{+}16, STMT12, SGG16, TLRB18, WLN^{+}17, vdCvW16, EHH^{+}13, JPN15, KARC15, Kur15, LKZ^{+}15, LWT^{+}15, TWSM17, ZCW^{+}19]. Dimensionality [PSPM12, RL15]. Dimensioning [KWM15]. Dimensions [EMP^{+}12, KZZM12, LK13, TWMSK18]. DimSUM [MDI^{+}18]. Dipoles [BSW^{+}12]. Dirac [LJC17, YDT^{+}18]. Direct [AFK^{+}14, BELD13, BD16, DKWB18, ER18, JKK^{+}18, KPD10, KVS^{+}14, KGP^{+}12, KWN^{+}14, LWBP14, LKG^{+}16, OT11, OLF^{+}14, RGG^{+}14, SJ19, SGM^{+}11, ŠPBV10, VCRG14, WK12b, XSE14, ZKWG16, SNJ^{+}14, WZC^{+}11]. Direct-Touch [KGP^{+}12]. Directable [MSGT18]. Directed [SAAB11, ZWHK16]. Directing [GLCC18]. Direction [MYGY19, WZK16]. Directional [CIE^{+}16, Hua17, LFA^{+}15, VCD^{+}16, MFW18]. Director [AWCO10, MCB16]. Dirichlet [LPG10, SGB13]. Disassembling [Att15]. Disassembly [KKSS15]. Discovering [BLY^{+}11, CDM^{+}17]. Discovery [PZDD19, SBCBG11a]. Discrepancy [PCX^{+}18]. Discrete [BCBSG10, BDS^{+}12, CDS10, FB11, HRWW12, HZRS18, LGH13, MSS11, PPH^{+}13, SLZ^{+}19, VMH^{+}13, WLT12, WJB^{+}13, WWH18, YMJ^{+}17, ZXW19, dGLB^{+}14, DFIM15, WIFD13, YDT^{+}18]. Discretization [QPCRM19]. Discretizations [SJP^{+}13]. Discrimination [NWHWfD16]. Discriminative [SXY^{+}11, WLN^{+}17]. Diseases [ZVE^{+}14]. Disk [EMP^{+}12, ERA^{+}16, ENMGC19, Gam16, Yuk15]. Disks [EMA^{+}13]. Disorders [BW17]. Disparity [KRMS13, RHL12, SBE19b, DKR^{+}14]. Dispersion [HHGJ15]. Dispersion-based [HHGJ15]. Displacement [JH12, ZR19, vGPNB17]. Display [CLGS18, HLR^{+}11, KKTD17, KAS^{+}19, LMLG15, LMG^{+}18b]. Displays [DER^{+}10, ESKD14, GRPF16, HF16, MbMYR15, RHS^{+}12, SM10, WRK^{+}16, WHL10, CHA^{+}14]. Disruption [RPMO13]. Dissection [TSW^{+}19]. Dissimilarity [KMAB15]. Dissipation [GBG^{+}14, SBO18]. Distance [AMAM13, BDF^{+}14, BF15, BFG^{+}17, CG16a, CT11, CSŠ18, LMM10, LPK13, MRL10, PPL13, PGK10, SFFP15, WCX^{+}13]. Distance-Based [BFG^{+}17]. Distances [CK11a, Pat16, Pat17]. distill [GBM^{+}19]. Distilled [AEWQ^{+}15]. Distinctive [SDA^{+}18]. Distortion [LWBP14, VP11, ZR19, VR12, VB14a]. Distortion-Free [ZR19]. Distortion-Guided [LWBP14]. Distortions [ČHM^{+}13]. Distributed [BAA^{+}16, GG14, LS16, MHHH15, MKRE16, TIK17, UWA^{+}19].

Distributing [HB19, SHQL18]. Distribution [GGV^{+}19, Hd14b, KH18, RBMS17, WYKR17, BSH12]. Distributions [ACKM16, ENMGC19, PSP10, SDS^{+}16]. Dithering [LLC19]. DIVA [KQR^{+}19]. Divergence [ELC19, KS14]. Divergence-Free [ELC19, KS14]. Diverse [ESKD14]. DMAT [YYG18]. DNA [MDI^{+}18]. Document [CY11, CWG11, EAGA^{+}16, HC14, KvLB14, LKC^{+}12, OSR^{+}14, PTT^{+}12, SSDK12, SSSG16, WPW^{+}11, ZAM^{+}16]. Documents [LKC^{+}12, PTW13, PTT^{+}12, SSDK12, SSS^{+}12, PSP^{+}14]. Domain [BYM18, BCCS12, HOB^{+}19, HGNH17, HGP^{+}19, LL18, MVZ16, RLH^{+}17b, GGM12]. Domains [SBC16, SCM^{+}19]. Domes [SM11]. Dominant [GSW12, PP11, PW13, KP19, ZSW10a]. Donut [SK16b]. Doughnut [CEX^{+}18]. Downsampling [TMH11, LN18]. dPSO [VMH^{+}13]. dPSO-Vis [VMH^{+}13]. Drawing [BZBM^{+}16, CLJ^{+}15, CLLC15, FWX^{+}13, FCS^{+}16, GZH^{+}19, IIX^{+}19, KHKS12, KWÖG18, LTKD15, LFA^{+}15, PFC15, WLL^{+}17, YSM19, YKM12, ZLDM16]. Drawing-Style [WLL^{+}17]. Drawings [CBC^{+}15, DKMT18, KÖS^{+}15, NS19, WK17, XSQ13]. Dressed [STC^{+}16]. Dressed-Human [STC^{+}16]. Driven [BS19, Bru19, CKGC14, CC14, ECN14, FB11, GLHH13, GCLX17, GCZ^{+}12, HSmCY13, IS15, JKK^{+}18, KFA^{+}14, LWBP14, LRB^{+}15, MHK^{+}19b, MBT^{+}12, NGDA16, RXX^{+}17, RBRY19, RSK12, SML15, SHW^{+}18, TCGK15, VS10, WK12b, WLL^{+}17, XKHK17, ZLW^{+}16, ZZ17, AECOK16, CLJ^{+}15, EIKM16, HPH10, LOM^{+}18, LMPS16, MHS^{+}14, MRM^{+}18, MFNP13, MBM13, RGG15, SA15, WSR^{+}17, ZR13]. Droplet [YLHQ14]. Droplet/Spray [YLHQ14]. Drosophila [SMB^{+}17]. Drug [KQR^{+}19]. Drug-Drug [KQR^{+}19]. Drying [JPK13]. DTI [SEI10]. Dual [DRW15, IEGT17, RCM^{+}14, SDKG18, Tak19, WHLW19, ZNZ19, WIFD13]. Dual-color [RCM^{+}14]. Duals [LMP^{+}10]. Dubbed [GVS^{+}15]. DuctTake [RWSG13]. During [GLGW12, OGW19]. Dwivedi [MHD16]. Dynamic [ACS^{+}17, BHRD^{+}15, BSC16, BBDW17, BBB^{+}18, BLV^{+}10, Bru19, BKW13, BBP10, CKL14, CNKI13, CLL^{+}13, DPF16, GRE11, GEY12, GPRS14, HLR^{+}11, ISYM15, IFDN12, JL17, JVS^{+}12, JvdGMR19, KKK18, KKPL19, KFLCO13, LBPH10, LWT^{+}15, MBRHD18, MC10b, MWTI19, NPW10, RTJ^{+}11, REH^{+}11, SPB^{+}19, SAMS^{+}17, SY14b, SHG^{+}16, SWG16, SRG^{+}19, VS10, VP11, VBAW15, VH15, WWL^{+}19, WKM15, WG12, YWC^{+}10, YL10, YWY10, EMU17, LCMP19, VMHB14]. Dynamically [ZZL^{+}17, CBV^{+}14]. Dynamics [ATK17, BET14, BTM^{+}19, FSTR13, FP15, FMD^{+}19, GRDE10, HENfSYS16, KHI^{+}19, LMA^{+}18, NHL16, PDV^{+}15, SMSH18, SHS^{+}17, TFA^{+}11, WBS^{+}13, ZTW^{+}12, OKK13]. DYVERSO [ATO17].

EACS [BP17]. Easy [CY11, SCR^{+}18, WHS^{+}18]. EasyXplorer [WCH^{+}15]. EcoBrush [GLCC17]. Ecology [SvL16]. Ecosystems [GLCC17]. Eddies [LSB^{+}17, SJ13a, WHP^{+}11]. Eddy [WPH^{+}12]. Edge [BdM14, HDL11, HGRS^{+}17, JGH11, LHT17, LJZX15, LFA^{+}15, TE10, WTL15, ZSJT19, ZWHK16]. Edge-Aware [WTL15, LFA^{+}15]. Edge-Optimized [HDL11]. Edges [BLVD11, BW13, SAAB11, LBA10]. Edit [BHW11, EIKM16, RKGS18]. editable [CH12]. Editing [BKY^{+}16, BHW11, BCK^{+}12, BKPB17, CCM16, CJZW12, CG16b, CCTL12, DRA10, DJM12, EKFM12, FACO17, FCS^{+}16, GCLX17, GMLMG12, GSE^{+}14a, GG15, GLGW12, HMTH13, HM15, HZZ11, HSK14, Jes16, KMS^{+}13, KMAB15, KPD10, KrJC^{+}11, LAS^{+}11, LJZX15, Lie17, LSL^{+}19, LLSC13, MBM13, MIW13, NKLN10, NSRS13, OPP10, PW13, RLH^{+}17b, SMH10, SBE19b, SPN^{+}16, SHJ^{+}16, SARZL10, TZD11, TRAW12, WDR11, ZH12, ZLDM16, dGWB^{+}14, DCNP14, LGC^{+}19, NGM14]. Editorial [CR16a, CR17, CB18, DZ15, GR11, JAP10, RD12, RD13, RD14a, RD14b, SG10]. Editors [Pet10, Pet10]. Edits [LJH10]. Education [SDA^{+}18]. EEG [APM^{+}11, JvdGMR19]. Effect [CEX^{+}18, GCGP18, HK16, MK15, WWV17]. Effective [BP17, Gia18, KGM^{+}10, LF10, TLM16, VF16, WCT^{+}15, XBL^{+}18]. Effectiveness [APM^{+}11, KH18]. Effects [AMT^{+}12, HHRZ12, JMV^{+}15, KH18, MSK14, OKG^{+}10, RKN12, SGM^{+}11, SKWL13, VWH18, WSK^{+}19, WYKR17]. Efficiency [CCI13]. Efficient [Aan18, AMTMH12, Att15, AKM16, AO13, BAAM17, BHW11, Bik12, CC14, CDA^{+}14, CWW^{+}11, CGH18, DZD^{+}16, DGR^{+}14, FFN18, FBL16, FM15, GPGSK18, GGG^{+}16b, HD14a, HA19, HZZ11, HHH12, IGMK^{+}16, IEK^{+}14, JH19, JKL^{+}16, KS11, KSO10, KBS11b, KrJC^{+}11, KCL^{+}18, KL19, LWS^{+}16, LLC^{+}15, LVW^{+}15, MCH13, MJBC13, Meu19, MK15, MRS12, MGN17, NS11, OM13, OBGB11, POS^{+}11a, PBPP11, Pat17, QYZ17, RWW16, RKR^{+}16, RLH^{+}17b, RPMO13, RS19, SBE16a, SLE18, SSO^{+}10, SKZ13, SDS^{+}16, SO10, TDNL18, TMRL14, TE18, TSdSK13, TSK14, TZD11, VGB14a, VB14b, VH15, WZC^{+}11, WZKP14, WTTM15, WLT^{+}17, WTMT18, WBS^{+}13, WCB15, WTH^{+}13, XL10, XXZC13, ZCZL13, ZZH15, TPSH14b]. Efficiently [MRD12]. EG [DS11a, LSF^{+}11, PS10]. Eigenproblems [NBH18b]. Elastic [Cet19, EHA^{+}19, KSKL13, LLC10b, PGBT18, SBO18, ZZT15, MWCS13]. Elasticity [TE18, ZLW^{+}16, ZZL^{+}17]. Elasticity-Based [TE18]. Elasto [ZLKW13]. Elasto-Plastic [ZLKW13]. Elastoplastic [CZZ^{+}18]. Electors [ATCO^{+}10]. Electronic [CKL14]. Electroporation [KFR18]. Electrostatic [SGBW10, SGW12]. Elegantly [GLGW12]. Element [BHU10b, KWSH^{+}13, MRS18, XGDC17, YFW12]. Elements [BEF17, CK11b, CFS14, GBG^{+}14, HZZ11, JTSZ10, PSC10, ÜFE10]. Elevation [SS15a]. Elimination [Yuk15]. Ellipses [GUK^{+}17]. Ellipsis [SH14a]. Ellipsoid [XDY18]. Embedded [BGCP11, BXH10, GBW16, KFA^{+}10, AH11]. Embedding [ABK^{+}19, CK14, KCL^{+}18, PHL^{+}16, SSSSW13, SDKG18]. Embeddings [BLTD17, HG18, HVP^{+}19, LJLH19]. Embellishment [LLC13]. Embellishments [SHK15]. Embodied [Thi11]. Emission [WW11]. Emissive [KYM19, SHZD17]. Emotion [ACC15, KKL16a]. Emotions [ACC15]. Emphasis [BRM^{+}16a, GOH^{+}10, HPK^{+}16]. Empirical [AKMM11, SBLC17]. Empirically [KARC17]. Employing [PEPM12].

Empty [GKL17, Man16]. Enabled [VOS^{+}10, OAJ14]. Enabling [ACS^{+}17, CJP^{+}19]. Enclosing [LK13]. Encoding [AMAM13, HWA^{+}10, KGR^{+}16, POCM19, VP11, XTZ19, VB14b]. Encodings [KH18, PH17, SK16b]. Encounters [BMH^{+}12]. End [APH^{+}12]. Energy [ACAA^{+}19, Cet19, KSS^{+}15, MMO16, PTA^{+}11, SGB13, ZLM^{+}15, dFH^{+}11, HP11]. Energy-Based [MMO16]. Energy-Conserving [dFH^{+}11]. Energy-scale [PTA^{+}11]. enforced [JZW14]. enforcement [TSK14]. Engineering [ABW^{+}15, DPD^{+}15, KBHM15, LSS^{+}12, LMK^{+}15, PPBT12, PH17, RPK^{+}12, TTB12, DMS14]. England [ID10]. Enhanced [CNCO15, JESG12, LM15, PSCC18, RAMG15, TGM12, XCDR10]. Enhancement [BCN11a, CZY11, HA11, HDL11, KLW12, LLC11, MO10, RHL12, SLCL19, ZCC14, ZBW11, HHS14, SLKL14]. Enhancements [TMHD12]. Enhancing [KK11b, LCC^{+}18, PW17, QWZ^{+}19, SK10, SGG16, VMTS10, WYD^{+}13]. Enjoyment [SSK16]. Enriched [KL19, ZZL^{+}17]. Enriching [KBB^{+}17, vDvdEvW19]. Ensemble [KBL19]. Ensembles [FKRW16, HW10, KWS^{+}15, LMK^{+}15, MW18b, PPBT12, WKS^{+}14, ZCH^{+}17]. entangled [GGG^{+}16b]. Entities [ACC15, SJB^{+}17]. Entity [EASG^{+}17]. Entropy [HVH^{+}16, KS13b]. Environment [BCD^{+}13, BNJ15, BCRA11, KL14, KÓOH13, MRMH12, PMD12, SH14b, SPV^{+}10, IMDN14, TAAP^{+}16]. Environment-Adaptive [KL14]. Environmental [GPG^{+}16]. Environments [CKHL11, DDH^{+}18, FVHK17, JH15, MHA17, MCG^{+}19, PDV^{+}15, SCCN11, DGR^{+}14, JPCC14, LCMP19, EHH19]. EnvyDepth [BCD^{+}13]. EPDiff [AVBC18]. Epigenomic [YNM^{+}13]. equalities [SSJ^{+}10]. Equalizer [LMS^{+}16]. Equation [PHTB12, HGA^{+}10]. Equations [CRGZ10, YLG^{+}18]. Equivalence [ZRLS19]. equivalent [vDHO16]. Erosion [CBC^{+}16]. Erratum [Ano13a, Ano13b, Ano15b]. Error [BEEM15, CJWL19, DCYG19, EPAS11, JKJL18, MMG10, NIDN16, SNJ^{+}14, VR12, VD18, WTTM15]. Error-Correcting [EPAS11]. Errors [GKPL11, HB19, KUMY10]. Estimates [LH11, SSSSW13, ZHD18]. Estimating [JCW11]. Estimation [AAS17, BEEM15, BM12, BM16, CEX^{+}18, CACB18, DDÖ^{+}17, GUS12, HDF15, HGNH17, HWL18, HET12, JMD15, KUMY10, LWDB10, LJL^{+}18, LMGH^{+}13, MBT^{+}12, MMO16, MES^{+}11, NIDN16, SW17, STC^{+}16, VVP^{+}16, WHD17, YWF^{+}19, ZNZ19, ZRJ^{+}15, vMRBPM17, SD10b, WFZ^{+}15, WGO^{+}14]. Estimators [SKGM^{+}17, WYD^{+}13, SNJ^{+}14]. Eulerian [JS10, PCBL16]. EUROGRAPHICS [Ano10a, Ano11c, Ano12d, Ano13f, Ano13i, Ano10c, Ano11e, Ano12f, AJL^{+}11, Bru11, Can11, CSLG10, Eis11, ID10, IC11, Kuh12, Ano10b, Ano11d, Ano12e, Ano13g, DS11a, LSF^{+}11, PS10, WBP11]. EvalBench [AHR13]. Evaluating [BHT19, CCH^{+}14, GHX^{+}17, HVH^{+}16, HYL^{+}15, JVS^{+}12, KK17, OJ15, SDK^{+}15, WLS13, HDV19]. Evaluation [AKMM11, ARH12, AHR13, APM^{+}11, APP10, BGCP11, BH19, BELD13, BMB^{+}18, BCBL13, CRY11, CK10a, CE19, DBS^{+}18, EWMU13, HV10, JC10, JKLS10, KDC17, KHW13, LCSCO10, MBDC15, MK15, MAAG12, OJ15, RGSK10, RB10, RLP10, RL15, SLB^{+}18, SSKB15, SA15, SHK15, SBD^{+}15b, TAEE15, UMM^{+}10, WHC15, WVV11, YCXW17, BLD14b, WGO^{+}14]. Even [BHW17]. Event [CWM19, HDF15, KCJM16, LKD^{+}17, WHD17]. Events [GPK^{+}12, PZDD19, SCD^{+}16, vDvdEvW19]. Everywhere [IRWM17, LSR17]. Evolution [Duc14, GLG^{+}16, RTJ^{+}11, SKZF11, SLSG16, TOZ^{+}11, VBAW15, WBT19]. Evolutionary [BCBL13, RC19]. evolving [CEG^{+}18]. EWA [HWK^{+}10]. Exact [CK10b, Kaz15, MC19, MAAG12, SHCB15]. Exaggeration [TX16]. Examining [QPCRM19]. Example [CYC15, FLJ^{+}14, FB11, GLCC18, GMM^{+}12, KS10, LP15, LPG19, LFA^{+}15, NGDA16, PP10, RJT18, RÖM^{+}15, RSK13, SLB^{+}18, SD10a, SJT^{+}19, ZZT15, ZLW15, ZLL13]. Example-Based [CYC15, GMM^{+}12, LPG19, PP10, SD10a, SJT^{+}19, ZZT15, ZLW15, FLJ^{+}14, KS10, LP15, LFA^{+}15, RJT18, RSK13]. Example-Driven [FB11, NGDA16]. Examples [PNVS17]. Excel [WPHC16]. Exchange [WPH^{+}12]. Excitable [NHL16]. Exemplar [IEKM16, ZCC14]. Exemplar-Based [ZCC14]. Exhibition [YH13]. Expanded [MW18a]. ExpandNet [MBRHD18]. Expansion [MBRHD18]. Expectation [JRJ11]. Expectation-Maximization [JRJ11]. Experiences [MRL^{+}17]. Experimental [MTVJ11]. Experiments [Meu19]. Expert [BCBL13]. Explanation [SvW13]. ExPlates [JE13]. Explicit [DGP17, FWSH19, LTKD15, VM12, YWTY12, AVBC18]. Exploiting [RBDD18]. Exploration [ABW^{+}15, AFK^{+}14, BGCP11, BPFG11, BSW^{+}14, BCBL13, BBBL11, BBL12, BBP10, CSG^{+}18, EAGA^{+}16, EASG^{+}17, FAFM19, FMH16, GHB^{+}17, GLG^{+}16, HGRS^{+}17, HVP^{+}19, JE13, JvdGMR19, KRD^{+}15, KQR^{+}19, KvLB14, KFR^{+}11, KTW^{+}13, LWBP14, LWT^{+}15, ME13, Man16, NJB^{+}11, NLB^{+}13, OSR^{+}14, OJMN^{+}19, PTW13, PEP^{+}11a, PFH^{+}18, PEPM12, RvdHD^{+}15, RCMA^{+}18, RPK^{+}12, RAMG15, SMH10, WG11, WCH^{+}15, YNM^{+}13, ZH14, vdEvW13, AKZM14, DGR^{+}14, MAG19, DKG15]. Explorative [BBB^{+}18]. Exploratory [ADN^{+}17, BH19, CHH^{+}19, KGP^{+}12, OGW19, SPF^{+}19, SV10, SvL16, TCM10, WHC15, YPF19]. Explore [BHRD^{+}15, SWG16, TLFC16]. ExploreMaps [DGR^{+}14]. Explorer [RHM^{+}12]. Exploring [AAB^{+}10, BCD^{+}10, BJA^{+}15, CWG11, DBD^{+}13, DRW15, HRS^{+}14, HC14, JTRS12, KWD14, KFLCO13, KW18, LT17, LXFW11, LBJ^{+}16, MRL^{+}17, PEP^{+}11b, PDV^{+}15, RL16, SMSL17, SSM12, SHS^{+}17, TLRB18, WLN^{+}17, ZOM19, vdCvW16]. Exponent [GKT16]. Exponential [HA19, PR19, SFY13]. Exposure [GTM^{+}12, GPR^{+}15, HA11, LP15, MPBM^{+}17, SLE17, SBE19a, SLWSS15, ZNZ19, SBB14]. Exposures [KR19, LDG19]. Expression [MD19]. Expressive [IIX^{+}19]. Expressiveness [KRMS13, LHH^{+}13, WTLL13]. Extended [SSW14a, SWT^{+}18, TGK^{+}17, WZCF15]. Extending [DBK11]. Extensions [FPC^{+}16]. External [BNN19, KLK17]. Extinction [GDML13]. Extracted [CS16]. Extracting [BPKB14, CJXH17, DHI^{+}15, EIKM16, KBW^{+}12, KCH^{+}18, LKD^{+}17].

Extraction [BAT11, GSW12, GT18, GLX^{+}16, HLH^{+}16a, JWH19a, KG19, KRS^{+}13, LW17, LW18, LKG^{+}16, MBES16, MPWC13, POS^{+}11a, PTA^{+}11, SBD15a, WL10]. Extraordinary [MM19, MM18]. Extrema [JWS12]. Extremal [GSW12, KCH^{+}18, WMZ12]. Extreme [XXLX14]. Extremities [SY14a]. Extrinsic [LJC17, WMXC19, YDT^{+}18]. Extruded [CJFH14, SM10]. Eye [BBGB19, BSBE17, BKR^{+}17, BKW13, GRC13, HYL^{+}15, KAS^{+}19, KHI^{+}19, RPA^{+}15, WBM^{+}18]. Eye-Centered [GRC13]. Eyeglass [KTN10].

F [vdCvW16]. Fabricating [CSaLM13]. Fabrication [ASH15, Att15, BFR17, BCMP18, BRWM18, BLS^{+}17, FKR13, HL15, HMA15, HBA12, Jac19, LMSG16, LZY^{+}17, MEKM17, NKSI16, PDP^{+}15, SP13, SCP^{+}17, UTZ16, VGB14a, WW16, WZK16, YWM15, YCXW17, ZZWC16, ZLW^{+}16, ZXZ^{+}17, ZKWG16]. Fabrication-Aware [BFR17, SP13]. Fabrics [SKZ11]. Facade [CML^{+}12, DCNP14, MWW12]. Facades [AYLM13, HWA^{+}10, KBK13, RBDD18, WXR^{+}16, PW17, IMAW15, VCL^{+}11]. Face [FNH^{+}17, GVS^{+}15, RC18, SLCO19, SDK^{+}15, SK17, TX16, WL10, ZTG^{+}18]. Face-Centered [RC18]. Faces [BSC16, BBIG17, CLG^{+}18, FGT^{+}16, GVS^{+}15, KÖS^{+}15, LLLY19, ZSW10a]. facet [BSH12]. Faceted [BRM^{+}16a, FKR13]. Facets [YS19]. Facial [BSC16, BS19, DBB^{+}18, DYFX19, FHW^{+}11, FGT^{+}16, GTB^{+}13, KÖS^{+}15, KMB^{+}17, KRP^{+}15, KBB^{+}17, LJZX15, LXFW11, MD19, RC19, SRH^{+}11, SCR^{+}18, SK17, TFA^{+}11, TX16, WL10]. Facilitating [BMPM12]. Factored [CML^{+}12]. Factorization [MC10b]. Factorized [WHB^{+}13]. Factors [DAF^{+}18, MRL^{+}17, SGS18, STMT12]. Fairing [CBLW19]. Families [LT17, HKM15]. Family [YS19]. Farthest [BJFadH19, CG12, YGJ^{+}14]. Fashion [SSE^{+}14, CLGS18]. FashionGAN [CLGS18]. Fast [ABD10, ATCO^{+}10, BCCS12, BAT11, BHH13, BM12, CK11b, Csé18, FWPS11, FH18, GL10a, GPP^{+}10, HCW17, HCW19, IUDN10, KS12a, KSO10, Kaz15, KLAB15, KKB18, KS12b, LK10, LK13, LWDB10, LMP^{+}10, LPH^{+}15, LW17, LGZ^{+}16, MS16, MFPA15, MAM14, NS19, NBH18b, NL18, PKS10, PEPM12, POCM19, QYZ17, SR19, SZMTW15, SPD14, SO10, SJT^{+}19, TSPP16, VKJ^{+}17, WWH18, WLM13, WTYH18, XSXM13, XWY19, YXX14, YYZZ18, YBY10, ZWC^{+}10, ZLSW17, ZSW^{+}10b, GD16, VR12]. Fast-forward [ZLSW17]. Faster [BRM^{+}16b, LCD10, MS16, NB12, SGEM16]. Fat [BJB^{+}18]. Fat-Tree [BJB^{+}18]. FaVVEs [BRM^{+}16a]. FCC [VCRG14]. Feathering [LLC^{+}15]. Feature [BYM18, BHU10a, CS16, CLLC15, DGQ^{+}12, ELM^{+}12, GSP19, GLK16, GLK18, GGP^{+}15, HTG14, HLH^{+}16a, HGA^{+}10, JH12, JKK^{+}18, KMD^{+}17, KYC16, LBG16, LB19, LPG10, LKG^{+}16, LRB^{+}15, MFNP13, MVZ16, MMV^{+}13, POS^{+}11a, PHK^{+}10, PTA^{+}11, RMZ13, SW17, SYM10, TX16, WL10, WYZC13, XXS^{+}15, XADR13, YXX14, YYL^{+}16, ZWY^{+}13, AFHdL14, WZC^{+}11]. Feature-Aware [CLLC15, KYC16, LBG16]. Feature-Driven [LRB^{+}15]. Feature-Oriented [YXX14]. Feature-Preserving [BHU10a, DGQ^{+}12, HTG14, JH12, KMD^{+}17, WYZC13, ZWY^{+}13]. Features [ABCJ10, AMS16, ABK^{+}19, AKZM14, BLY^{+}11, BBB^{+}18, BHS^{+}17, BPKB14, BM12, CJXH17, DMS14, EIKM16, KSP^{+}18, KCH^{+}18, KRS^{+}13, PSCN10, PPH12, RHM^{+}12, Sch11, SJW^{+}11, WVV11, ZCOAM14, vdCvW16, LBBC14]. Featuring [Pos11b]. Feedback [AFHdL14, MTVJ11, XWY^{+}15]. Fellows [Ano10b, Ano11d, Ano12e, Ano13g]. FEM [BHU10b, CFS14, JTSZ10, KKB18, SSB13, YI10, YLHQ12]. Few [MSZ^{+}18]. Few-shot [MSZ^{+}18]. Fewer [OMW16]. Fiber [AAS^{+}16, CGT^{+}15, FEM^{+}19, KGM^{+}10, LZB17, OVV10, PEPM12, RPK^{+}12, SSSSW13, SEI10]. Fiber-Level [LZB17]. Fibers [ACG^{+}17]. Fibonacci [MBR^{+}13]. Fibrous [KK14]. Fictional [RGM^{+}18, SJB^{+}17]. Fidelity [DMAL10, JVS^{+}12, WZCF15]. Field [AMTMH12, AH11, ABCCO13, BEJM15, BB12a, CWW^{+}11, COC15, DZD^{+}16, DLY^{+}18, DZC11, FKSS13, FL19, FKS^{+}10, FKR13, GEZ^{+}17, GT16b, GG17, HML^{+}19, HWK15, HRS18, JWH19a, MRD12, NSRS13, NZH^{+}18, NJC^{+}19, OKG^{+}10, OGHT10, PPM^{+}16, PM19, RHLH18, SW10, SGM^{+}11, TW10, VSG^{+}13, VCD^{+}16, WLZT18, WZG^{+}19, WCX^{+}13, YWY10, ZZH15, BB14, CGT^{+}15, CFGL16, EGG^{+}15, SBB14]. Field-Aligned [DLY^{+}18, NZH^{+}18, WZG^{+}19]. Field-Coherent [PPM^{+}16]. Fields [BCBSG10, BPKB14, BB12a, BOB13, BSEH17, CYY^{+}11, CGBG13, CSŠ18, DLD12, DVPSH14, FKSS13, GPK^{+}12, GRT18, GKKT13, GST14, GKT16, HLH^{+}16a, HMA15, KG19, KCH^{+}18, LS16, LWS18, LKG^{+}16, NVT^{+}14, OBH^{+}11, ORT18, OT12, PPH12, PRW11, PW12, RSK12, SW17, SFFP15, Sch11, SFL^{+}16, SBCBG11b, Szy11, SBL12, Tim12, TRAW12, VB14b, WRS^{+}13, WGS10, WHT12, ZRLS19, dGLB^{+}14, vPJtHRV12, GRT14]. Figure [YPL19]. Figures [CYI^{+}12]. Filament [KGM^{+}10, HL14]. Filament-Surface [KGM^{+}10]. Filling [SB13]. Film [HBRFP19, NDG17, TDMS14]. Films [NREM14, XADR13]. Filter [BCCS12, HGO18, MS14, PTO10]. Filterable [WWH18]. Filtered [AHMAM15, LTB19]. Filtering [ABD10, BEM11, BEJM15, BBB^{+}18, BTS^{+}17b, CLC12, FWSH19, FR11, GO15, IGMK^{+}16, JLKL16, KVS^{+}14, KWN^{+}14, KTW^{+}13, KK11b, LZFH18, LWS^{+}16, MS16, MJBC13, MIGMM17, RSD^{+}12, SFFP15, SDMS15, SGYF11, SBS^{+}17, SXLS19, SCQ^{+}19, Tok15a, TSPP16, VCRG14, WMB15, WHLW19, YBY10, YWY10, ZCZL13, ZDZ^{+}15, ZSW^{+}10b, ZWY^{+}13, ZRJ^{+}15, LN17]. Filters [BTS^{+}17b, LK17, MS13, PK15, RI17, WYKR17]. Final [MW11]. Financial [KCA^{+}16, ST18]. Finding [HKL17, LKF12, PLL11]. Fine [KKS^{+}17, SY11, SY13, ZWY^{+}13]. Fine-Grained [KKS^{+}17]. Finger [LAFT12, WWS^{+}15]. Fingering [SKK10]. Fingerprint [OKK13]. Finite [BHU10b, CK11b, GT15, GKT16, JTSZ10, KWSH^{+}13, MRS18, ÜFE10, ZHD18]. Finite-Elements [CK11b]. Finite-Sample [ZHD18]. Finite-Time [GT15, GKT16]. Fire [SKWL13]. Firefly [ZHD18]. First [APH^{+}12, BRM^{+}16b, CCC^{+}14, GL10a]. First-order [BRM^{+}16b]. First-Person-Views [CCC^{+}14]. Fisheries [BMPM12]. FitConnect [OW19]. Fitted [GPP^{+}10, OW19]. Fitting [ABCJ10, BTP13, FKSS13, KWW^{+}14, LK13, LWL^{+}16a, MHRK19, NL13, PPL13, TSB16, VSG^{+}13, BSH12]. fixed [VR12]. Flame [OAM^{+}18]. Flare [LE13]. Flash [MJL^{+}13]. Flat [MEKM17]. Flat-Foldable [MEKM17]. Flattening [KMM^{+}18, PTW13]. Flattening-based [KMM^{+}18]. Fleshing [ZR13]. Flexibility [BLY^{+}11]. Flexible [ABCJ10, BAAM17, BXH10, BSAP11, DAD^{+}19, DZC11, PFC15, SR19, SVWG12, SB19, SFLP18, WK12a, WCH^{+}15, vBE11]. FlexyFont [PFC15]. Flicker [WWV17]. Flight [HA11, KBKL10]. Flights [PSC10]. FLIP [CIPT14, FAW^{+}16, SWT^{+}18]. Flips [SY13]. Floating [MSS^{+}10, SFLP18, SMG10]. Floating-Point [MSS^{+}10]. Flood [CKS^{+}15, CBKK^{+}19, WKS^{+}14]. floor [PGVG19]. Flow [Aan18, BGCP11, BBB^{+}18, BPKB14, BP19, BBL12, CRC^{+}15, CZY11, CK11b, CGS16, COC15, CKS^{+}16, DI18, DYFX19, ELM^{+}12, ELPH19, FE17, GOPT11, GSE^{+}14b, GT15, GKT16, GT16a, GG17, HYL^{+}15, HWK15, JS10, JWC^{+}11, KSW^{+}12, KSBC12, KS18, KFM^{+}19, LGP14, LL19, LS16, MRL^{+}17, MLP^{+}10, MKP^{+}16, MC14, NJB^{+}11, OBH^{+}11, OJMN^{+}19, POS^{+}11a, PTA^{+}11, PPF^{+}11, SGRT12, SEG^{+}14, SLZ^{+}19, SBL12, TBKP12, WHT12, WBT19, WPH^{+}12, XWY19, ZHQH17, ZWC^{+}10, ZAD15, dHvPJV14, vPJtHRV12, vPGL^{+}14, CSFP12, KGGP18, LCMP19, PHE^{+}11]. Flow-Based [FE17]. Flow-Induced [GG17]. Flow-Orthogonal [SGRT12]. Flower [YGCO^{+}14]. Flowers [BSCH18, ZFG^{+}17]. Flows [ATW15, BvTH16, GT16a, HRWW12, HWK15, KSBC12, LJB^{+}12, RGG18, TAOZ12, TDF^{+}15, WESW17, WMRSF15]. Flowstrates [BBBL11]. Fluctuations [DPF16]. Fluid [AMT^{+}12, BSW10, CK13, EHT18, FAW^{+}16, GBW16, HEW15, HJS^{+}17, KPNS10, KSW^{+}12, KAT^{+}19, LCY^{+}19, OAO11, SCN^{+}16, SJ13a, SHQL18, SZMTW15, TL19, WMRSF15, WBT19, WTYH18, ZYF10, dHvPJV14]. Fluid-Solid [TL19]. Fluids [AIAT12, ATO17, ATW15, AWO^{+}14, BXH10, CK13, CBC^{+}16, DGP17, GGV^{+}19, GDGP16, HLL^{+}12, IPKK13, IEGT17, KPNS10, KAT^{+}19, NC10, SCN^{+}16, SKK10, TDF^{+}15, TL16, WMRSF15, WKBB18, WYY13, YLHQ12, YWTY12, CIPT14, YLCH18]. Fluorescence [JWH^{+}19b, MFW18]. Fluvial [CBC^{+}16]. Flux [BSW^{+}12, KPS^{+}14]. Flux-Limited [KPS^{+}14]. Fly [LZB17, OAM^{+}18, ZOA^{+}18, DJSJ19, SKK^{+}14b]. Foams [KL19]. Focal [SGG16]. Focus [BTM^{+}19, GRE11, HVP^{+}19, KGRG17, MbMYR15, OJS^{+}11, TM13, BCN11b]. Foldable [MEKM17]. Folding [ZIM13, ZCBK12]. Foldover [SFL19]. Foldover-Free [SFL19]. Follow [OGW19]. following [LJK^{+}12]. Font [ZCL18]. FontRNN [TXL^{+}19]. Fonts [BBK^{+}19, TXL^{+}19, ZCL18]. Footprint [WH17a]. Force [Gov19, SAAB11, ZWHK16]. Force-calculation [Gov19]. Force-Directed [SAAB11, ZWHK16]. Forecast [WFZ^{+}15]. Forecasting [DPD^{+}15, DPD^{+}17, FKRW16, RCMM^{+}16, RG19]. Forecasts [SWG16]. Forest [KS18, ZBM^{+}17]. Forests [BN12]. Form [PSP10, GAK10, HMA15]. Formalizing [HPK^{+}16]. Format [GPM^{+}18]. Formation [IPKK13, XWY^{+}15]. Formed [NKSI16]. Forms [WLT12]. Formulation [PGBT18, TDF^{+}15]. Forward [GT16a, LDdLRB16, WHD17, ZLSW17]. forward-scattering [WHD17]. Foundations [BRB^{+}13, LFK^{+}13, LJH13]. Four [HTG14, LZQ13, MTM12]. Four-Dimensional [LZQ13]. Fourier [SLE18]. Foveated [LCC^{+}17, WRK^{+}16]. Fractal [SB13]. Fractals [Gda17]. Fractional [LPD14]. Fractions [KPNS10]. Fracture [CZZ^{+}18]. Fractured [GMM^{+}12]. Fragment [PTO10]. Fragment-Parallel [PTO10]. Frame [DBS^{+}18, KG19, LWL^{+}16b, PC12, SKWL13]. Frame-to-Frame [PC12]. FrameBuffer [UWA^{+}19]. Frames [BPKB14, HB19]. Framework [ABW^{+}15, AMS16, BDA^{+}17, BSC16, BDC18, CKGC14, CYY^{+}11, CE19, DKMT18, FWPS11, GTL^{+}18, HCSC16, JZYP18, KL19, LBJ^{+}16, NIDN16, PEP^{+}11b, PJSH15, RPLH11, SSE^{+}14, TIK17, Vax14, XSE14, YMJ^{+}17, ZCZL13, ZAM^{+}16, BRM^{+}16a, EPAS11, EKB14, Lie17, MAG19, YDT^{+}18]. Free [DG12, DC10, ELC19, GGV^{+}19, HMA15, KS14, KFW19, LGZ^{+}16, NKSI16, PPH^{+}13, SFL19, SKTM11, YIC^{+}11, ZR19, CBV^{+}14, RRS12, RLH^{+}17b]. Free-form [HMA15]. Free-Formed [NKSI16]. Freeform [JWWP14, MIW13, XDY18, BPW14, DPW11]. Freehand [ZBW11]. Frequencies [BKW13]. Frequency [BDC18, IFL13, MC10b, OPP10, TM13, WWV17, IFDN12]. Fresnel [EKFM12]. Friendly [KGR^{+}16]. Front [Ano16g, Ano16a, Ano16b, Ano16c, Ano16d, Ano16e, Ano16f, Ano17a, Ano17b, Ano17c, Ano17d, Ano17e, Ano17f, Ano17g, Ano18a, Ano18b, Ano18c, Ano18d, Ano18e, Ano18f, Ano18g, Ano18h, Ano19b, Ano19c, Ano19d, Ano19e, Ano19f, Ano19g, Ano19h, Ano19i, CCC^{+}14, CSC^{+}18, KSCN19, SGRT12]. Frontmatter [Ano15a, Ano15f, Ano15c, Ano15d, Ano15e]. Fruit [KRB11]. FTP [YSC^{+}18]. FTP-SC [YSC^{+}18]. Full [MMRO13, YHL^{+}16]. Fullsphere [MC10b]. Fully [ACA18, LRBB17, SKZF11, WBM^{+}18]. Fully-Implicit [SKZF11]. Function [AVBC16, BF15, CNCO15, GOH^{+}10, HFM10, JH19, KPG^{+}16, PPBT12, PGK10, PSO18, SPB^{+}17, WZL^{+}12]. Functional [AWO^{+}14, BCGL18, DPW11, GSTOG16, GBKS18, HCO18, LRB^{+}16, MWS^{+}10, NMR^{+}18, RPWO19, RCB^{+}17b, RMC17, RLB^{+}19, SVBC19, SSM12, WGBS18, ZCOM13]. Functionality [HSvK18]. Functions [AGJ12, BDF^{+}14, HW10, HH10, IYS^{+}13, JWS12, LLC^{+}10a, LKG^{+}16, MGV11, MSS11, MRL10, NL13, NBHN17, PSP10, PDV^{+}15, RSK13, SKZ13, SRK13, SFL^{+}16, SS15b, SMG10, BCGS13, TWSM17]. Fundamental [WLT12]. fused [RCM^{+}14]. Fusion [DYFX19, EGG^{+}15, HW16, HA11, JJKL18, YCL^{+}17, ZCZL13, KBÖ^{+}14]. Future [OJMN^{+}19, vLKS^{+}11]. Fuzzy [LS15, SCF10, SDKG18, YSC^{+}18].

G [KP19]. G-splines [KP19]. galleries [LSN^{+}14]. Game [MVLS14, MTVJ11]. Gamma [BSH12]. Gamut [JWH^{+}19b, NKB14, PR12, SSGM17]. Gamut-Based [NKB14]. Gap [YKS^{+}19]. Garment [DDÖ^{+}17, ZCF^{+}13]. Gas [ZHQH17]. Gathering [LWLD11, MW11]. Gauss [Cet19, YBY10]. Gaussian [HWF^{+}17, IFDN12, LS16, Tok15b, WFZ^{+}15, XL10, YZXW12]. Gaussian-Based [HWF^{+}17]. Gaussian-Distributed [LS16]. Gaze [BKW13, HBO^{+}10, MBM13, OAJ14, PMG13, RKGS18, RPA^{+}15, SGEM16, WBM^{+}18]. Gaze-Contingent [SGEM16]. Gaze-driven [MBM13]. Gaze-enabled [OAJ14]. GazeDirector [WBM^{+}18]. GEARS [WZKP14]. GEMSe [FMH16]. Genealogical [RHM^{+}12]. General [AMS16, Ano10a, Ano10c, Ano11c, Ano11e, Ano12d, Ano12f, Ano13f, Ano13i, BEEM15, BÖK11, HVVR18, HCSC16, HLJ^{+}13, KS13a, LT16, LWS^{+}13, LBJ^{+}16, MR17, PCDS12, PSP10, RÖG17, SV14, SDHL11, WZKP14, WHT12]. General-Purpose [HLJ^{+}13]. generalization [vDHO16]. Generalized [AMAM13, BDA^{+}17, CG16b, DLGY12, EDPB15, HWU^{+}19, Jac17, Jes16, KPK10, MCM^{+}12, PKS11, PGK10, STKD12, XTZ19]. Generalizing [CGT^{+}15]. Generated [UKCB15]. Generating [DJSJ19, LVJ10, LLC19, MH17, RLGH15, TXL^{+}19, WTLL13, Yuk15, LLLY19]. Generation [AKV15, BYM18, BSK^{+}13, BLP^{+}13, CNCO15, CG16a, CGH18, CBC^{+}16, CEG^{+}18, DCYG19, GPMG10, GTL^{+}18, Gda17, GD10, GSZ11, IIX^{+}19, JYS19, JSLW14, KS12a, KWC^{+}12, KÖS^{+}15, KMA^{+}19, KGRG17, LMP^{+}10, LEE17, NREM14, NS11, OJMN^{+}19, RPA^{+}15, RSK10, SYM10, SD10a, SCR^{+}18, ŠBM^{+}10, SB19, VKW^{+}12, VGB14a, WL10, YCL^{+}17, YWY10, ZZWC16, ZCX19, vFG11, vdKdJP^{+}19, BMWW14, HGA^{+}10, SKK^{+}14b, SKK^{+}14a]. Generative [CLGS18, KAT^{+}19, QWZ^{+}19, SLCL19, WGW^{+}19, WPY^{+}19, WLZX19, YDP19, ZLZ^{+}18, HKM15, NW17]. Generic [LGC^{+}19]. Genetic [Deb18, HSS17, LS15]. Geneva [Ano10c, Ano11e, Ano12f, Ano13i]. Genome [KMA^{+}19, SCD^{+}16]. Genomic [NHG19]. Genomics [LSS^{+}12, MWS^{+}10]. Genotype [GKPL11]. Genus [LK19]. Genus-Zero [LK19]. Geo [DBS^{+}11, vGPNB17]. Geo-information [DBS^{+}11]. Geo-referenced [vGPNB17]. GeoBrush [TSS^{+}11]. Geodesic [CK11a, HZRS18, MR12, WK12a, XYM13, XLS^{+}14, ZWC^{+}10]. Geodesics [HRWW12, SCF10, ZZCJ14]. Geodesy [CHK13]. Geographic [CY14, LBK14, WKS^{+}14]. Geographical [LCB^{+}18, ZLMM16]. Geological [DKG15, RMH^{+}18]. Geometric [ATK17, BHS^{+}17, BMR^{+}16, BvTH16, CK11a, CD10, CDSS14, FLL11, FSTR13, GUK^{+}17, HK12, KZB19, KBK13, LMW^{+}19, MLP^{+}10, PCBS16, PSCC18, PPY^{+}16, SLD^{+}17, TSS^{+}11, VKJ^{+}17, WLX19, WMRSF15, YLX^{+}16, BCGS13, BHU10b, VMHB14]. Geometrical [PRW11, SPT14]. Geometrically [SHCB15, TL19, YYL^{+}16]. Geometries [JL17, Kaz15, RXX^{+}17, WH17b]. Geometry [ABCJ10, BPW14, BS12a, BDS^{+}12, BGS10, BNH10, BBDA10, CVDL16, CSD11, CCLN10, CCW12, CG19, CK13, CH11, CDS10, DKB^{+}16, DRF12, DFY14, EMP^{+}12, EBGM12, FWX^{+}13, FGT^{+}16, GPK^{+}12, HRS^{+}14, HM15, HMA15, HREB11, IEKM16, Kaz15, KMHG13, Kim15, KBÖ^{+}14, LT17, Lai13, LGH13, LTB19, LZQ13, LJBA13, LPG10, LCLJ10, LLW12, MS10a, MS12, MCM^{+}12, MBG^{+}12, MSWI12, MPM^{+}14, MPWC13, NWHWfD16, Pat16, PHK^{+}10, PPM^{+}16, PKS11, PK15, Rus10, SCM^{+}19, SY12b, She12, SLHC12, SFL^{+}16, SYT^{+}13, SJW^{+}11, TSS^{+}11, VS10, WKM15, WLJ^{+}18, WGH^{+}19, WCX^{+}13, WDAH10, YH13, ZYF13, ZSW^{+}10b, ZWY^{+}13, ZIM13, BLD14b, VB14b, YDT^{+}18]. Geometry-Aware [CK13]. Geometry-Driven [VS10]. Geosemantic [SAG^{+}13]. Geospatial [AKV15, BMH^{+}12, BMS^{+}10, CBC^{+}15, CKS^{+}15, CKS^{+}16, DKG15, RHM^{+}12, SvL16, WTLY12, vDHO16]. Germany [PS10]. Gestalt [KNH^{+}18]. Gestaltlines [BNRS13]. Gestures [ATF12, LAFT12, SHW^{+}18]. GGX [TH17]. Ghosted [BGCP11]. Ghosting [SW11]. Gigaray [BOB13]. GigaSample [TRAW12]. GIS [NDD14]. Glass [AAS^{+}16]. Glint [CCM19]. Global [BRDC12, BEM11, BEEM15, BEJM15, BLK11, BB12b, BMB15a, BMB15b, CLC12, DKL10, DVP19, ENSB13, ESRT13, FKRW16, GRC13, GRR^{+}16, HGRS^{+}17, HREB11, JMD15, KLCF10, KKPL19, KTO11, KLAB15, LDdLRB16, MAM14, NKLN10, NNDJ12, PW12, REH^{+}11, RDGK12, Ros13, SW17, SNRS12, SY12b, SYC10, SCCN11, SSM12, SHD15, TSdSK13, WHB^{+}13, WMB15, WH17b, WWH18, WHP^{+}11, XXY^{+}18, YWC^{+}10]. Globally [FACO17, JBG19, PSDB^{+}10, SSG17, SEG^{+}14]. Globes [YJD^{+}18]. Gloss [BPV18]. Glossy [DBK11, LWLD11, Tok15b, XWB15]. Glowing [WW11]. Glycogen [ACAA^{+}19]. Glycogen-derived [ACAA^{+}19]. Glyph [JKLS10, LCP^{+}12, MVB^{+}17, SK10, YWS^{+}14]. Glyph-Based [LCP^{+}12, MVB^{+}17]. Glyphs [CAB^{+}16, GRT19, KWD14, SK16a, SK16b, WPHC16, HSJW14]. Goal [PJJ^{+}11]. Goal-based [PJJ^{+}11]. gons [KP19]. Good [KP18]. Governing [NHL16]. GPGPU [EPAS11]. gProximity [LMM10]. GPU [ÁSK14, AWJ13, BHP15, BJFadH19, Cet19, CYY^{+}11, DHP^{+}19, GM19, GL10a, HGRS^{+}17, JBG17, JH12, JSLW14, JZYP18, KKS^{+}17, KDCM14, KGR^{+}16, LMM10, LWY^{+}11, LMMCO17, MW11, MS14, MAG19, MBG^{+}12, MRS18, OBGB11, PKS10, PBPP11, PC12, PGKS17, RLH17a, RPLH11, RGG^{+}14, SKNS15, SSO^{+}10, SKK^{+}14b, SKK^{+}14a, TTN^{+}13, TWT^{+}16, VHB16, WBS^{+}13, WT11, WTYH18, YHGT10, ZSS17, ZFE16]. GPU-accelerated [MAG19]. GPU-Adapted [ZSS17]. GPU-Based [RGG^{+}14, BHP15, RPLH11, LMM10, MW11, MRS18, PC12, TTN^{+}13, TWT^{+}16]. GPU-Friendly [KGR^{+}16]. GPUs [HHRZ12, HLJ^{+}13, KBS11a, LCD10, MRAS17, VH15]. GrabCut [CPZ^{+}15]. Gradient [BYM18, CRGZ10, CGBG13, DMCN^{+}17, HOB^{+}19, HGNH17, HGP^{+}19, LJKL17, LTKD15, LKSD17, MVZ16, MJBC13, NNN11, PBE18, RLH^{+}17b, RLMB^{+}14, RLGH15, SBE19b, VK18, WTL15, WZL^{+}17, XWL^{+}15]. Gradient-Based [NNN11, DMCN^{+}17, RLGH15]. Gradient-Domain [BYM18, HOB^{+}19, HGNH17, HGP^{+}19, MVZ16, RLH^{+}17b]. Gradient-Guided [SBE19b]. Gradient-Index [CRGZ10].

Gradient-Preserving [WZL^{+}17]. Gradients [HYZ^{+}14]. Graffinity [KLS^{+}17]. Grain [NDG17, TDMS14]. Grained [KKS^{+}17]. Graining [CDSS14]. Grammar [HWA^{+}10, KWS16]. Grammar-based [HWA^{+}10]. Grammars [BHMT13, HLL16, MBG^{+}12, VRBC17]. Graph [ACS^{+}17, BBDW17, CC14, CDS16, GRE11, GSE^{+}14a, Gov19, HBH18, HET12, KKTD17, KRM^{+}17, LLLY19, LVJ10, LFC14, LGW18, OAJ14, OJ15, PMD12, SLH^{+}18, SBM^{+}14, STP17, WT11, XDC^{+}13, YKS^{+}19, MSW19]. Graph-Based [CDS16, STP17, WT11]. GraphDice [BCD^{+}10]. Graphics [ALCS18, AJL^{+}11, BET14, BMO^{+}14, BCD^{+}12, CTHAM10, DCGG11, DHP^{+}19, GSHM10, GMD10, GP16, HČA^{+}12, JH12, JVS^{+}12, JCW11, KH19, KBKL10, KDCM14, Kuh12, LF10, LBH12b, MWN^{+}17, MSS^{+}10, MH13, OP10, PJJ^{+}11, PTO10, PCR11, RE12, SDA^{+}18, SRH^{+}11, SGYF11, SLD^{+}17, TGM12, TMHD12, VM12, Lie17]. Graphs [APP10, BHR17, BCD^{+}10, BHW17, CLWM11, DPF16, DvKSW12, DRW15, DSL15, GEY12, GSE^{+}14a, GOB^{+}10, HET12, KRD^{+}15, KS13b, KLS^{+}17, KHKS12, KS10, Kur15, LBA10, LLW12, LWM^{+}17, MSDK12, MSFM16, NHL16, NB12, OJ15, PGS^{+}16, PFH^{+}18, PMD12, RTJ^{+}11, RZS10, RMM15, RMF12, SSW14a, SSKB15, SSK16, SCR^{+}18, SAAB11, SBD^{+}15b, TCLK12, TLFC16, TE10, VBAW15, VBW17, VM12, WDM^{+}12, WT17, WK17, XDC^{+}13, ZWHK16, vLKS^{+}11, DGR^{+}14]. GraphUnit [OJ15]. Grasp [KS12b, KKTD17]. Grasping [PGGM10]. Grassmannian [LBJ^{+}16]. great [KGGP18]. Greedy [CDSS14]. Green [SO12]. Grid [FWPS11, FDH^{+}15, IPKK13, KH19, PTP^{+}15, UBH14, YLHQ14]. Grids [AO13, DC10, KBS11a, KW18, LS16, MS10b, PGKS17, XYM13, ZSS17]. Ground [CMF18, SDK^{+}15]. Grounded [CBK^{+}17]. Group [CDA^{+}14, GKB12, NPCB17, RCB^{+}17a, SSK16, VBW17, WH17b]. Group-in-a-Box [CDA^{+}14]. Group-Theme [NPCB17]. Groups [CDA^{+}14, DRW15, JPN15]. Growing [MGB^{+}12]. Growth [KSG^{+}15, YLG^{+}18]. Guaranteed [GOPT11]. Guarantees [KLAB15]. Guarding [YL11]. Guest [JAP10]. Guidance [CGM19, FWX^{+}13, LKZ^{+}15, SJT^{+}19]. Guided [BEM11, BŠMM11, CJFH14, DW13, DBB^{+}18, DYFX19, FMH16, HLL^{+}19, KP18, KMAB15, KMHG13, KSKL13, LWBP14, NJB^{+}11, NZH^{+}18, RC19, SBE19b, ZTW^{+}12, ZDZ^{+}15, GD16, HENfSYS16, KBL19, LLSC13, RKR^{+}16, SMH10, SCQ^{+}19, VB14b, WMXC19, ZH14]. Guidelines [SLB^{+}18]. GuideME [ZH14]. Guiding [HEV^{+}16, KHW13, MGN17, NNRS15, NC10, WWV17, XM15]. GWCNN [ESKBC17].

Hair [CBTB16, HBLB17, JL18, KB12, MSGT18, OXKP12, PKS11, QWZ^{+}19, RKN12, WWL^{+}13, WLI^{+}12, ZTW^{+}12, dFH^{+}11]. HairControl [MSGT18]. Hairstyles [RKN10]. Half [AMTMH12, HKD15]. Half-Space [AMTMH12]. Halftoning [AP10a, LM10, LCG10, SGBW10, SGW12]. Hallucinating [ZCW^{+}15]. Halos [TMHD12]. Hand [FLJ^{+}14, LLCZ16, PGGM10, TST^{+}15, WWS^{+}15]. Hand-colored [FLJ^{+}14]. Hand-held [LLCZ16]. Handling [FM15, HEW15, LPH^{+}15, MFW18, TL16, TDNL18, TWT^{+}16]. Handwriting [CLJ^{+}15]. Hard [SWP11, TSK14]. Hardware [And10, BH15, GI18, LBG16, LMLG15, LMPD15, MMF10, NKF^{+}16, SM10, SDHD17, TGM12, VKJ^{+}17, VOS^{+}10]. Hardware-Assisted [MMF10]. Hardware-Based [LMPD15]. Harmonic [HCW17, HCW19, LPG10, SFL^{+}16, KBB^{+}13]. Harmonics [MRCB18]. Harmonization [LPSB18, MWS^{+}16]. Harmony [YWF^{+}19]. Hatching [USSK11]. Having [SM14a]. Hazards [CKS^{+}15]. Hazy [BPV18]. HCCMeshes [KBK^{+}10]. HCI [RPA^{+}15]. HDR [EWMU13, GPR^{+}15, KR19, MBDC15, SBB14, SHG^{+}16, TAEE15, TAEE16, ZBW11]. HDR-Video [EWMU13]. Head [WRK^{+}16]. Head-Mounted [WRK^{+}16]. Heat [AAS17, BPVR11, BJG^{+}15, DLL^{+}10, MGG^{+}10a, OMMG10, WPH^{+}12]. Heavy [CBKK^{+}19]. Height [HMA15, TW10, Tim12]. held [LLCZ16]. Helicoids [PKS11]. Hemispherical [MBB19]. Hemodynamics [NLB^{+}13]. Heritage [HBRW^{+}12, PPY^{+}16, SCP^{+}17, AJL^{+}11]. Heritge [AJL^{+}11]. Hermite [IYS^{+}13, MGV11]. Hero [WND^{+}14]. Heterogeneous [BAAM17, BDF^{+}14, ENSD12, KRD^{+}15, LSS^{+}12, RGG18, SDS^{+}16, SHZD17, WWH^{+}10, ZM16]. Heterojunction [ASB^{+}17]. Heuristic [SSS^{+}12]. Hex [GHX^{+}17, GSZ11]. Hex-mesh [GHX^{+}17]. Hexahedral [CAS^{+}19, GSP19, LMPS16, XGDC17]. Hexahedralization [Tak19]. Hidden [BKM^{+}19, MCB16, ZDJ16]. Hiding [WGW^{+}19]. HidingGAN [WGW^{+}19]. Hierarchical [AP10b, BL18, CNKI13, DHP^{+}19, FWPS11, GPGB11, GRT14, HMB17, HGRS^{+}17, HVP^{+}19, JRJ11, JYS19, JGH11, KBL19, KKS^{+}17, KBK^{+}10, LMM10, LWS18, LHH^{+}13, MS11b, MBJ^{+}15, PP11, PHL^{+}16, PM19, SVBC19, SMG10, TF15, UKCB15, VK18, VRBC17, VBHH13, WHWB16, ZH12, vPJtHRV12]. Hierarchical-Culling [KBK^{+}10]. Hierarchies [Áfr12, BHH13, BM15, DWT^{+}11, GHB15, ND12, SM11, SBE16a, SBE16b, SPH11, VKJ^{+}17, VHB16, WD11, WAF^{+}11]. Hierarchy [KBLE19, MB18, SMP13, WXL^{+}11]. HiFiVE [SSSSW13]. High [ABD10, AFK^{+}14, ASH15, BAT11, BAA^{+}16, BEM11, BHU10a, CKL14, Csé18, DW13, DER^{+}10, DMAL10, EMP^{+}12, EHH^{+}13, FR11, FGT^{+}16, GO15, GBW16, GPRS14, HBH18, JPN15, KARC15, KP19, KKK18, KBK13, LCC^{+}17, LTB19, LKZ^{+}15, LWBP14, LWT^{+}15, LBJ^{+}16, MBRHD18, MSW10, PSPM12, PHL^{+}16, REH^{+}11, SPB^{+}19, SBE16b, SHG^{+}16, SGMG17, SGYF11, ST18, SGG16, SKTM11, TTN^{+}13, TM13, TLRB18, TSPP16, ÜFE10, WWV17, WLN^{+}17, WGW^{+}19, WLI^{+}12, XSE14, ZLYL17, vdCvW16, EMU17, SHD16]. high-degree [SHD16]. High-Dimensional [ABD10, FR11, LWBP14, LBJ^{+}16, PHL^{+}16, SGG16, TLRB18, vdCvW16, EHH^{+}13, JPN15, KARC15, LKZ^{+}15, LWT^{+}15]. High-Frequency [TM13]. High-Level [HBH18]. High-Order [GO15]. High-Quality [BHU10a, Csé18, DW13, MSW10, REH^{+}11, SBE16b, TSPP16, ÜFE10, BAT11, BEM11]. High-refresh-rate [DER^{+}10]. High-Resolution [FGT^{+}16, GBW16, TTN^{+}13, KKK18, LCC^{+}17, ZLYL17]. High-Speed [LTB19, SGMG17, WLI^{+}12]. Higher [Muñ14, POS^{+}11a, SK10, Sch11, ÜFE10]. Higher-Order [POS^{+}11a, SK10, Sch11, ÜFE10]. Highlight [FZS^{+}19]. Highlighting [GP18]. Highlights [LWDB10, RGB^{+}14]. Highly [CR16b, SPV^{+}10]. Hilbert [SSSSW13]. Histogram [KMAB15, LFS^{+}15, WCB15]. Histograms [LS15]. Histology [GHB^{+}17]. Historical [BMH^{+}12, KMK12, PTW13, PSP^{+}14]. History [GLG^{+}16, IGMK^{+}16, SLSG16]. HMLFC [PM19]. Holistic [KD13]. Hollow [LLC10b]. Homogeneous [MSZ^{+}18, MESG11, RRS12]. Homography [HWL18]. homological [DFIM15]. homologically [Kur15]. Homologies [CVJ15]. Homology [RL15, RL16]. Homotopic [DSL15]. Homunculus [RRS12]. Honeycomb [JWWP14]. HPR [MEMO14]. Human [AWCO10, BW17, BBIG17, CYC15, CYI^{+}12, DAF^{+}18, DLGY12, ESKT15, ECN14, JL17, KSCN19, LWS^{+}13, MSZ^{+}18, NMM^{+}19, PLL11, RZS10, STC^{+}16, WPY^{+}19, YBK^{+}12, vMRBPM17]. Human-Object [WPY^{+}19]. Humanities [HRD^{+}15, JFCS17, MHHH15, RPSF15, RAMG15]. Humanoid [NBH18a]. Humans [VVE^{+}10]. Hunting [ST18]. Hurricane [LMK^{+}15]. Hybrid [BSY^{+}19, HCJ13, KPNS10, KFA^{+}10, LFS^{+}15, LS15, LN17, MCHW18, PBPP11, PCK10, RMF12, SKK10, TDNL18, UBH14, YLHQ14, KBÖ^{+}14]. Hydrographics [PDP^{+}15]. Hyperbolic [RKSA17]. Hyperelastic [MMO16, XHC^{+}18]. HyperMoVal [PBK10]. Hypersliceplorer [TWMSK18]. Hypothesized [KQR^{+}19]. Hysteresis [WLZ13].

I-SI [WDM^{+}12]. Ice [IPKK13, IUDN10]. ICP [TST^{+}15]. Ideation [ADN^{+}17]. Identification [HV10, MMNG17]. Identifying [LGK16]. IGM [BAO^{+}19]. II [ABCJ10, CD10, CBV^{+}14, DLS10, DER^{+}10, EKB14, HWK^{+}10, HHS14, IY10, JCT14, JPCC14, KS10, LSN^{+}14, LCLJ10, MDD^{+}10, PHM^{+}14, RTK^{+}14, SPK10, TW10, VB14a, VF14, WGO^{+}14, ZSW^{+}10b, ZH14, vdCAvW14]. III [CVCH14, GDAU14, JZW14, PSP^{+}14, SPCR14, SLKL14, TPSH14b]. IISPH [CIPT14]. Illuminant [NKB14]. Illuminated [JVS^{+}12]. Illumination [ACG^{+}17, BCD^{+}13, BRDC12, BEM11, BEEM15, BEJM15, BELD13, BD16, BB12b, BMB15a, BMB15b, CJZW12, CLC12, CSC^{+}18, CNS^{+}11, DKL10, DVP19, ENSB13, GKPS12, GRC13, GRR^{+}16, HVVR18, HREB11, IFL13, JSYR14, JMD15, KS13a, KD13, KPD10, KTO11, KYM19, LDdLRB16, LVV18, LWS^{+}13, MR17, MBR^{+}13, MC10b, NKLN10, NNDJ12, RGG15, REH^{+}11, RDGK12, SJ19, SNRS12, SYC10, SCCN11, SHD15, SJ13b, SSGM17, TSdSK13, Tok15b, Tok15a, WHB^{+}13, WMB15, WWH18, XSXM13, XXZC13, XWZB17, YWC^{+}10, ZZWC16, ZM16, ZZLX17, ZNZ19, SNJ^{+}14]. Illumination-driven [RGG15]. Illumination-Varying [LVV18]. IlluminationCut [BMB15a]. Illustrate [CRY11]. Illustrating [SEI10]. Illustration [FLJ^{+}14, JBMC10, LKEP14, LCUR14, RLMB^{+}14].

Illustrations [WT11]. Illustrative [ABG^{+}12, BBBV12, CSFP12, CYY^{+}11, HGH^{+}11, LMP13, LVPI18, LPSV14, LSB^{+}17, OVV10, RMH^{+}18, STKD12, vdZLBI11]. Image [ASW14, ARM^{+}15, AGJ12, AEWQ^{+}15, AECOK16, BCN11a, BCD^{+}13, BEM11, BKM^{+}19, BHW11, BCK^{+}12, BKPB17, BP19, BMS^{+}12, ČHM^{+}13, CG16a, CJFH14, CRC^{+}15, CSD11, CJZW12, CG16b, CPZ^{+}15, CCTL12, CNKI13, DTV15, DRA10, DDÖ^{+}17, DAD^{+}19, DSH^{+}17, DYFX19, DJM12, EWMU13, EKFM12, EIKM16, FWSH19, FDH^{+}15, FACO17, GMLMG12, GO15, GCW15, GD10, GTK^{+}12, GLGW12, HW16, HA11, HDL11, HFM16, HM15, HČA^{+}12, HFE13, HZF10, HZZ11, HZMH14, HWL18, HGH^{+}11, IEH^{+}14, IEK^{+}14, IEKM16, IY10, Jes16, JESG12, JWL^{+}13, KS13a, KWSH^{+}13, KS10, KrJC^{+}11, KGAC15, KKL16a, KGK19, KK11b, LG19, LLC11, LCCC13, LVV18, LJH10, LZL^{+}15, LWL^{+}16b, LB19, LCG10, LJZX15, LQS^{+}19, LEE17, LLC13, LWS^{+}16, LSZ^{+}18, LLB^{+}10, LLX^{+}11, LCUR14, LS15, LMGH^{+}13, LLSC13, MWS^{+}16, MEMO14, MTM12, MVZ16, MFPA15, MNP^{+}17, MKU15, MJL^{+}13, MRS12, NNRS15, NPCB17, NPW10]. Image [NBA18, NREM14, OAO11, PPW18, PWS12, PSHZ^{+}15, PDC^{+}19, PCR11, PK15, PBC^{+}16, QWZ^{+}19, RWW16, RKR^{+}16, RSD^{+}12, RBDD18, SMH10, SKZ13, SC10, SO12, SHJ^{+}16, SDK^{+}15, SPK10, SKWL13, SLCL19, SJ13c, SGEM16, SPT14, SPSK13, SGSP15, SFLP18, TXGW19, TBKP12, TE10, TDDD18, TZD11, TMH11, TCGK15, VCL^{+}11, VSG^{+}13, WZH13, WKM15, WLX19, WZG^{+}19, WWL^{+}13, WLM13, WHL10, XL10, XXZC13, XSQ13, XADR13, YXX14, YWC^{+}10, YZL17, YLD^{+}18, YLH^{+}14, ZCW^{+}15, ZH12, ZCZL13, ZZH15, ZCF^{+}13, ZJ18, ZFJ^{+}16, ZLZ^{+}18, vdCvW17, BHT19, DZD^{+}16, HHS14, LGC^{+}19, HRRR18]. Image-Based [CSD11, GD10, LLB^{+}10, MRS12, TE10, VCL^{+}11, WLM13, BKM^{+}19, BP19, KGK19, RKR^{+}16, RBDD18]. Image-Guided [CJFH14]. Image-Space [NPW10, SGEM16]. Image/Video [DRA10]. Imager [AVR10]. Imagery [ACA18, LTK12, MK11]. Images [ABD10, AVR10, ARM^{+}15, AECOK16, BKB^{+}12, BKY^{+}16, BBAM12, BRWM18, ČHM^{+}13, CJXH17, DZD^{+}16, DJM12, EIKM16, FACO17, FZS^{+}19, GMLMG12, GEZ^{+}17, GCZ^{+}12, GPRS14, GG15, HMTH13, HM15, IEKM16, IRWM17, Jes16, KKBJ16, KWSH^{+}13, KLTZ16, KG18, KYC16, LJKL17, LJH10, LM15, LJL^{+}18, LEE17, Lie17, LCUR14, LMLF15, LLC19, LDY10, MES^{+}11, OVB^{+}15, PK10, PH17, PK15, PNVS17, RLH^{+}17b, RWSG13, SKWL13, SLKL14, SPSK13, TM13, TX16, TAEE16, WWV17, WKM15, WCM15, XXS^{+}15, YCL^{+}17, YFWR11, ZCW^{+}15, ZZH15, ZLYL17, ZDJ16, ZVE^{+}14, PGVG19, BB14, CVCH14, CHA^{+}14, DKR^{+}14, EKB14, HHS14, JCT14, PHM^{+}14, PSP^{+}14, SBB14, SPCR14]. Imaging [CKL14, DHS^{+}13, HA11, HHNČ19, HZMH14, KYKL14, KB12, NKB14, PKE17, RKN12, SHG^{+}16, SSB^{+}14, TDMS14, TAEE16, WILH11, WLI^{+}12, YXX14, YMS10, ZTW^{+}12, ZBW11, BB14]. Imitating [SLC^{+}19]. imMens [LJH13]. Immersive [FFN18, MCG^{+}19, SPV^{+}10]. Immune [HPvU^{+}16]. Impact [BJA^{+}15, CCH^{+}14, SHK15, WBFvL17, ZLMM16, BHT19, ZLM^{+}15].

Impaired [CJP^{+}19]. Imperfect [BBH13, JKL^{+}16, REH^{+}11]. Implementation [IABT11, OP10]. Implicit [BBCW10, EBGM12, KFA^{+}10, MS10b, PP11, PGBT18, QPCRM19, SKZF11, SYC10, TDF^{+}15, VVC^{+}11, WKBB18, ZLKW13, dGWB^{+}14]. Implicits [MGV11]. Importance [BMDS19, GKPS12, Hd14b, HEV^{+}16, HH10, KS11, KF12, LXY19, LPG13, MW11, OXKP12, SHSK16, WK12b, ZZ19, RRS12]. Impostors [ABCN10]. ImPrEd [SAAB11]. Improve [HBO^{+}10, OAJ14, OP10]. Improved [CWY11, CSŠ18, FC10, HKD15, Jes16, LW18, NPDD11, NC10, NMR^{+}18, PR12, RSSL17, SAAB11, WZK16, WLX19, ZHLW18, ZHD18]. Improvement [AMR^{+}17]. Improving [Bik12, CCI13, GKB^{+}11, GLGW12, HLS12, HL14, LYG15, LS15, MHD16, NBCW^{+}11, Ren16, SHD15]. IMUs [vMRBPM17]. In-Core [Bik12]. in-Front-of [CCC^{+}14]. In-Kernel [HLJ^{+}13]. In-situ [WAF^{+}11]. Inaccurate [SPSK13]. Inbetweening [WNS^{+}10]. Incomplete [DLL^{+}10, SY12b, TOZ^{+}11]. Incompressible [Aan18, HLL^{+}12, KS14, PGBT18, SZMTW15, TDNL18, ZHQH17, CIPT14]. Inconsistent [HHCJ18]. Incorporating [AMS16]. Increase [SSKB15]. Increasing [HHNČ19, LTC18]. Incremental [LSR17, LQS^{+}19, MWW16, PPT^{+}19]. Independent [KKS^{+}12, LMLG15, MSWI12, NDG17, NPW10, SVLL10, SBF15, YHGT10, NB12]. Index [CRGZ10, DLGY12]. Indexing [AKMM11, MAM14, WCB15]. Indicator [MSS11, WPH^{+}12]. Indices [SBLC17]. Indirect [BHR17, CLC12, CNS^{+}11, GLCC18, KYM19, LK10, LWDB10, Tok15b, Tok15a, YWC^{+}10]. Individual [SK16b]. Individualized [WL10]. Indoor [JL19, KMHG13, LTX^{+}14, SHL^{+}14, SLC^{+}19, WCM15, MPM^{+}14]. Induced [GG17, RCMA^{+}18, Ste19]. Inertia [NS19]. Inertia-based [NS19]. Inertial [GT15, GT16a, GT16b, GG17, RGG18, vMRBPM17]. Inexact [YLD^{+}18]. Inexpensive [ACV^{+}14]. Inextensible [SHCB15]. Inference [SBC14]. inferred [ZLDM16]. infinite [SKK^{+}14b]. Inflation [GHB^{+}17]. Information [ABK^{+}19, Ano11a, Ano11b, Ano12a, Ano12b, Ano12c, Ano13c, Ano13d, Ano13e, Ano14a, Ano14b, Ano14c, Ano15i, Ano15g, Ano15h, Ano16j, Ano16k, Ano16h, Ano16i, Ano17h, Ano17i, Ano17j, Ano18i, Ano18j, Ano19j, Ano19k, BBR^{+}16, BBK^{+}18, BMB^{+}18, BRB^{+}13, DAF^{+}18, FFN18, HPK^{+}16, ML17, MSK14, RL19, RLH17a, RMZ13, TLM16, WDM^{+}12, WGW^{+}19, WLS13, XWY^{+}15, ZC18, DBS^{+}11, GBM^{+}19]. Information-Based [XWY^{+}15]. Information-Theoretic [BRB^{+}13, ZC18]. Informative [FDH^{+}15, NO17]. Informed [ZCC14]. InfoVis [BNRS13, HSBW13, vdEvW13]. Infrastructures [BAA^{+}16]. Inherent [WJDZ14]. Inhomogeneous [KYM19, SKTM11, SKGM^{+}17, SKMS18, YIC^{+}11, ZSL^{+}17]. Injective [HCW19, JHT14, SKPSH13]. Ink [LCCC13, SO12, HLJ^{+}13]. InK-Compact [HLJ^{+}13]. Inpainting [LZL^{+}15, WLZX19]. Input [SPSK13]. Inputs [LLCZ16]. Ins [KFA^{+}10]. Insect [LWPL15, WJDZ14]. Insects [GCY^{+}14]. Insertion [BHH13]. Insertion-Based [BHH13]. Insets [GRE11].

InsideInsights [MHK^{+}19b]. Inspection [RKRD12]. InSpectr [AFK^{+}14]. Inspired [LWPL15, SW10]. Instance [LIK19]. Instancing [FKE13]. Instant [DGGK11, FGT^{+}16, LDW^{+}10, LJH10, MGY^{+}18, YWHB18]. Integer [WSSC11]. Integral [GKKT13, IFDN12, LKEP14, MBB19, ZBQC13]. Integrals [MBR^{+}13]. Integrated [GHB^{+}17, GRPF16, JKJL18, SMS^{+}17]. Integrating [CBK^{+}17, ERT^{+}17, MHK^{+}19b]. Integration [GT16a, MLP^{+}10, NBMJ14, SKZF11, SHZD17, Ste19, ZZT15, ZOM19, ZLKW13, SD10b]. Integration-Based [MLP^{+}10]. Integrations [RLH^{+}18]. Intelligence [LPSV14]. Intelligent [EPAS11, PJJ^{+}11, SRH^{+}11, SGYF11]. Intended [CS16]. Inter [LTC18, PDW^{+}14]. Inter-author [LTC18]. Inter-Comparison [PDW^{+}14]. Interacting [LJL^{+}18, WC14, YLCH18]. Interaction [BRM^{+}16a, CTL13, CYI^{+}12, CKE^{+}12, GEY12, HSH16, HHD^{+}12, JH15, KKTD17, KGP^{+}12, LAFT12, NW13, PHTB12, RSM^{+}16, RPA^{+}15, SS16, TCLK12, WLL^{+}17, YLHQ12, YBK^{+}12, ZCK17]. Interactions [CTL13, KQR^{+}19, OGW19, SGSP15, WPY^{+}19, VHG^{+}18]. Interactive [Áfr12, ACAA^{+}19, ARH12, AYLM13, ATF12, BCB^{+}15, BEM11, BEJM15, BET14, BHR^{+}19, BJB^{+}18, BTM^{+}19, CK11a, CRGZ10, CJFH14, CCM16, CSG^{+}18, CLC12, CWL^{+}15, CFS14, CCH^{+}14, CKH19, CEG^{+}18, CBKK^{+}19, CNS^{+}11, DRA10, DRM19, DTS^{+}14, EGG^{+}15, EPCV15, ELPH19, FLL11, FE17, GLCC17, GBKS18, GP12, GKHF14, HK12, HBDP17, HG18, HBO^{+}10, HPvU^{+}16, IMDN14, JTRS12, JE13, JZYP18, JSYR14, KFH10, KWM15, KKSS15, KVS^{+}14, KÖS^{+}15, KMJE12, KS18, KNH^{+}18, KFR^{+}11, KRS^{+}13, KTW^{+}13, LKC^{+}12, LCP^{+}12, LCCC13, LBRM18, LCG10, LTK12, LHH^{+}13, LBH12a, LSWW11, LWBP14, LZ10, MJK17, MMG18, MLD^{+}18, MLK^{+}13, MWW12, NKSI16, NSM19, NPW10, NBA19, OJ15, PTW13, PEP^{+}11a, PP10, PBK10, POG13, Pos11b, PM19, RPK^{+}12, RWW16, RKRD12, RTK^{+}14, RLP10, RWS^{+}10, RDGK12, SH14b, SGM^{+}11, SRH17, SvW13, STKD12, SMSL17, SPH11, SARZL10]. Interactive [SPV^{+}10, SHS13, SWS12, SMS^{+}17, SJB^{+}17, SHS^{+}17, SMP13, SMB^{+}17, TSS^{+}11, Tak19, TLW^{+}19, Tok15a, TGK^{+}17, TWMSK18, TRAW12, TPRH11, TCGK15, ÜFE10, UTZ16, WDC^{+}10, WNS^{+}10, WCM15, WAF^{+}11, WZ15, XYM13, YWC^{+}10, YDP19, YNM^{+}13, ZWC^{+}10, ZWHK16, dGWB^{+}14, BSW^{+}14, CVCH14, GDAU14, LGC^{+}19, PSP^{+}14]. Interactively [BLY^{+}11]. Interchange [MRM^{+}18, ZFAQ13]. Interconnected [KK11a]. Interest [JKK^{+}18]. Interest-based [JKK^{+}18]. Interesting [BSW^{+}14]. Interface [BCD^{+}13, CGS16, IY10, LWX^{+}15, RC19, TDNL18, TZD11, WPH^{+}12, DKG15]. Interfaces [ATF12, CSLG10, HK12, JTRS12, MC10a, SWS12]. Intergalactic [BAO^{+}19]. Interiors [MMP16, WFLW18]. Interleaved [FC10]. Interlocking [YCXW17]. Internal [BPKB14, KLK17]. International [AJL^{+}11]. Interpolation [CRC^{+}15, Csé18, GCLX17, GP16, HCSC16, KMK12, LKSD17, MCH13, NC16, PCBL16, RC18, RJGW19, RSK10, RSK13, SV14, SC16, WWL^{+}13, WDAH10, YYL^{+}16, YFW12]. Interpret [LMA^{+}18].

Interpretation [ZCW^{+}19, ZLDM16]. Interrupted [AAS^{+}16]. Intersecting [KJT14, SJP^{+}13, SP13]. intersection [RRS12]. Intersections [CK10b, CCC17]. Interval [LJKL17]. Interventional [RWS^{+}10]. Interventions [KFR18]. Intervisibility [Tim12]. Intrinsic [BKPB17, BEKB15, DSH^{+}17, GMLMG12, GEZ^{+}17, GUS12, KLCF10, LVV18, PBB^{+}13, SBC14, SBCBG11a, TBW^{+}11, WH17b, ZHH^{+}15, YDT^{+}18, WMXC19]. Introduction [APH^{+}12]. Introductory [TLM16]. Intrusive [YHL^{+}16]. Intuitive [BCB^{+}15, CGS16, RPMO13]. Invariance [MYGY19]. Invariant [BV19, CIE^{+}16, ZBQC13]. Inverse [ALCS18, HSmCY13, HWF^{+}17, ŠBM^{+}10, SPK^{+}14, ZHQH17]. Inversion [Gda17, GT16a, LGZ^{+}16]. Inversion-Free [LGZ^{+}16]. Inverted [HSK14, HKSK18]. Inverted-Pendulum-based [HKSK18]. Invertible [ZZL^{+}17]. Investigating [Ros13, SE19, WSK^{+}19]. Investigation [BJB^{+}18]. Invisible [RNLL10]. Invited [Ake11, Ano15a, Pos11b, Thi11]. Iridescence [KGK19, Ste19]. Irradiance [JR16, KVS^{+}14, MRMH12, MC10b, RCB11]. Irregular [DC10, KMD^{+}17, MCHW18, PGKS17, RSK12, TLH19]. Irreversible [KFR18]. ISHair [OXKP12]. Iso [BHU10b, FKRW16]. Iso-Contours [FKRW16]. Iso-geometric [BHU10b]. IsoMatch [FDH^{+}15]. Isometric [GSTOG16, OMMG10, SY11, SY12a, SY13, TSB16, ABCJ10]. Isometry [CBSS17]. Isometry-Aware [CBSS17]. Isosurface [BM10, LCD10, MRL10, SBD15a]. Isosurfaces [BW13, CGT^{+}15, Gro16, MS10b, PRW11, WLS13]. Isosurfacing [LCDW16]. Isotropic [CCW12, CR16b, DLY^{+}18, ZWY^{+}13]. Isotropically [LW17]. Issue [Ano11a, Ano11b, Ano12a, Ano12b, Ano12c, Ano13c, Ano13d, Ano13e, Ano14a, Ano14b, Ano14c, Ano15i, Ano15g, Ano15h, Ano16j, Ano16k, Ano16h, Ano16i, Ano17h, Ano17i, Ano17j, Ano18i, Ano18j, Ano19j, Ano19k]. Items [vdCvW16]. Iteration [Gda17]. Iterations [DHI^{+}15]. Iterative [BTP13, BMS^{+}12, SBC16, WH17a, YYL^{+}16, MRL10, TBKP12]. iVisClustering [LKC^{+}12].

Japanese [vdKdJP^{+}19]. jets [KGGP18]. Jittered [CKK18]. Joint [FWSH19, HGO18, KVD^{+}10]. Jointly [DTV15]. Joints [Jac19, MCJM18, TE18]. Journalists [SRG^{+}19]. Journey [AKS^{+}19]. Julia [Kim15, SvW13]. June [ID10]. Just [WSCP13]. Just-in-Time [WSCP13].

kd [VHB16, SR19, XL10]. KD-Tree [XL10]. kd-Trees [VHB16, SR19]. kDet [WDZ17]. Kelp [DvKSW12]. Kernel [AAS17, BLTD17, GUS12, HLJ^{+}13, HET12, OMMG10, WGBS18]. Kernels [AAS17, BS12a, BLTD17, CG16c, FWSH19, LJC17, Pat16, Pat17, Ros13, Rus11, SR14]. Key [LLD10, PZDD19, ZC18]. Key-Pose [LLD10]. Keynote [McC11]. Killing [SBCBG11b, BCBSG10]. Kinect [WZ15]. Kinematic [BSK^{+}13, KVD^{+}10]. Kinematics [ALCS18, HSmCY13, HWF^{+}17]. Kit [HHH12]. Knot [KSD14a]. Knowledge [KARC17, PZDD19, vKTS^{+}11]. Kyrix [TLW^{+}19].

L [ŠBM^{+}10]. L-systems [ŠBM^{+}10]. L4RW [XLL^{+}10]. Labeling [BNN19, LCHB12, XXLX14]. Labelling [PPT^{+}19]. Lagrangian [AW13, CZY11, FKS^{+}10, GOPT11, JS10, KER^{+}14, MBES16, SW10]. Lamps [BBIG17]. Landmark [JL17, KSKL13, TBW^{+}11]. Landmark-Guided [KSKL13]. Landscape [HW10, ML17]. Landscaper [ADMAS18]. Landscapes [ADMAS18, ACV^{+}14, CEG^{+}18]. Langevin [CZY11]. Language [NW13, RHM^{+}12]. Languages [RHM^{+}12]. Laplace [HP11, NBH18b, PPH^{+}13, QCW^{+}18, ZLW15]. Laplacian [PEP^{+}11a, EKB14, LGW18, Pat16, Pat17, UMK19, VD18, YI10]. Laplacian-based [PEP^{+}11a]. Laplacians [HKA15, VMHB14]. Lapse [SSB^{+}14]. Large [Áfr12, ABCN10, BHP15, CSG^{+}18, CBC^{+}16, DWT^{+}11, GLCC17, GRDE10, HSK^{+}10, HJM^{+}11, HPvU^{+}16, KRD^{+}15, KLS^{+}17, KHKS12, KS18, KHI^{+}19, LAE^{+}12, LSS^{+}12, LCB^{+}18, LCDW16, MSDK12, MLD^{+}18, MHDG11, NPDD11, NSM19, PFH^{+}18, PEPM12, REH^{+}11, SM11, SSG17, SJ13a, SPH11, SAAF18, SK17, SMM13, SJWS13, TXL^{+}19, TE10, TBP18, TSdSK13, WSSC11, WAF^{+}11, YXX14, ZC18, vdEvW13, vLKS^{+}11, GDAU14, MPM^{+}14, TAAP^{+}16]. Large-Scale [ABCN10, BHP15, DWT^{+}11, GLCC17, KS18, KHI^{+}19, LSS^{+}12, MHDG11, SM11, SPH11, TBP18, WSSC11, WAF^{+}11, ZC18, SMM13, TXL^{+}19, GDAU14, MPM^{+}14]. Lassoing [DH16]. Latent [FMD^{+}19, GSDG18, HMRR19, LJLH19, WDM^{+}12, WBT19]. Latent-space [FMD^{+}19]. Lateral [HF16]. Lattice [ABD10, FDL14, HA17, ISYM15, MLD^{+}18, NZH^{+}18, RC18]. Lattice-Guided [NZH^{+}18]. Lattices [CR16b, VCRG14]. Lauren [Ano16l]. Layer [ESKBC17, GSDC17, HZF10, IMAW15, SDMS15, CSFP12, ISYM15, LCD10]. Layer-Based [IMAW15]. Layered [AHMAM15, BSW10, BKB^{+}12, IEK^{+}14, IRWM17, KTKM19, KYM19, MbMYR15, MVH^{+}14, RHS^{+}12, LLC19]. Layers [DVP19, KG18, LSWW11, RLMB^{+}14]. Layout [CK14, HBH18, JL19, WTLY12, WFLW18, ZOM19, MVLS14, MSW19]. Layouts [CDA^{+}14, CLS16, CCH^{+}14, FDH^{+}15, GSE^{+}14a, Gov19, KRM^{+}17, LSWW11, PPM^{+}16, RRP15, SLH^{+}18]. Laziness [XLL^{+}10]. Laziness-based [XLL^{+}10]. Lazy [MAAG12]. LBSN [TLFC16]. Learned [HMW^{+}15, MHRK19, HKM15]. Learner [SHS^{+}17]. Learning [BBK^{+}19, BKM^{+}19, BLVD11, BHS^{+}17, BZL^{+}18, BMM^{+}15, BM16, CWGvW19, ČHM^{+}13, CSC^{+}18, ERT^{+}17, FWSH19, GKOM18, HZMH14, IIX^{+}19, KKBJ16, KLTZ16, LJZX15, LGK16, LCY^{+}11, MSZ^{+}18, NBH18a, OGW19, PFC15, SOC19, SLCO19, SHS13, SS15b, SDKG18, SJW^{+}11, WPY^{+}19, WBT19, XXLX14, XXS^{+}15, XADR13, ZZ19, HMRR19, LBBC14, LTR19]. Learning-Based [SOC19]. Learnt [SM14b]. Least [BGS10, KBS11b, MM19, MS10a, MGB^{+}12, PSPM12, RAV^{+}19, KBÖ^{+}14]. Leaves [JPK13]. Led [APH^{+}12]. Left [SSM12]. Legacy [XZP^{+}13]. Legends [GKB12, RLP10]. LEGO [ZCX19]. length [JZW14]. Lens [BTS^{+}17b, BCN11b, HHH12, LE13, LK17, LZQ13, SMSL17, SDHL11, ZZ17]. Lens-Flare [LE13]. Lenses [JKL^{+}16, TGK^{+}17, HD14a, SHD16]. LeSSS [HMW^{+}15]. Letters [Pet10]. Level [AFK^{+}14, ABCN10, HFM10, HBH18, HREB11, JKL13, KBS11a, KVS^{+}14, KWN^{+}14, KDCM14, LPD14, LWBP14, LMMCO17, LZB17, NJB^{+}11, SBD15a, SRK13, STKD12, TSB16, VCRG14, WDM^{+}12, WLS13, XSE14, ZBM^{+}17, HK16, JWL^{+}13, LGC^{+}19, MVLS14, XXY^{+}18]. Level-of-Abstraction [STKD12]. Level-of-Detail [HREB11, KDCM14, SRK13]. Level-set-based [WLS13]. Levels [MBDC15, PJR^{+}14, WT11]. Levels-of-Detail [PJR^{+}14]. Library [AHR13]. LiDAR [PW17, WXR^{+}16, vKvLV13]. LiDAR-Based [PW17]. Lie [SAD^{+}16]. Lie-Algebra [SAD^{+}16]. Lifting [NSC14]. ligand [VHG^{+}18]. Light [BB12a, BOB13, BH15, CLG^{+}18, CJWL19, DZD^{+}16, DKH^{+}14, DKL10, DLD12, FKR13, GLCC18, GEZ^{+}17, GPK^{+}12, GP18, GPRS14, HML^{+}19, HKD15, HDF15, HEV^{+}16, HMRR19, HH10, HHS14, HR10, JVS^{+}12, JMV^{+}15, JA18, JG19, JR16, KD13, KPD10, LDdLRB16, LF10, LHH^{+}13, LMGH^{+}13, MDBS14, MGN17, MMRO13, NIDN16, NSRS13, NJC^{+}19, NGHJ18, OKG^{+}10, PdMJ14, PSP10, PDC^{+}19, PRDD15, PM19, RGG15, RKRD12, RSK12, SLE17, SHD15, SKGM^{+}17, TH17, VSG^{+}13, WHL10, WW11, YBK^{+}12, YWY10, YIC^{+}12, ZZH15, ZAD15, BB14, SBB14]. Light-Field [BB12a, BB14, SBB14]. Light-Transport [MGN17]. Lighting [BNH10, BN12, CLC12, CNS^{+}11, GLCC18, GKB^{+}11, GKPS12, IFDN12, KGK19, LK10, LWDB10, LHH^{+}13, NAM^{+}17, NPW10, PCDS12, RKRD12, SPN^{+}16, SPF^{+}19, SARZL10, TH17, WJB^{+}13, WHD17, WG12, XZP^{+}13, YWC^{+}10, YIC^{+}11, IMDN14, TPSH14b, JL19]. Lighting-by-Example [GLCC18]. Lights [Gam16, KYM19, LXY19, NNDJ12, Tok15b, VGS^{+}19, YZXW12]. Lightweight [GPGSK18, KÓOH13]. Like [AYWM14, PSCN10, GTZM10, LS10b]. Limit [HAGO19]. Limited [DBLW15, KPS^{+}14, MW18b]. Limited-Size [MW18b]. Limiting [MYLZ16]. Line [ABCJ10, AKMM11, BBB^{+}18, BZBM^{+}16, BD16, CML^{+}12, CBLW19, GZH^{+}19, IIX^{+}19, KWÖG18, KER^{+}14, LKEP14, LLW12, MHDG11, NS19, SJ19, SHZD17, SMJ17, SJW^{+}11, Tim13, VVC^{+}11, WJB^{+}13, WESW17, WT11, XSQ13, YKS^{+}19, DZD^{+}16, GRT14, HNJ^{+}14]. Line-Based [WESW17, VVC^{+}11]. Line-Drawing [BZBM^{+}16]. Line-Plot [MHDG11]. Line-Sweep [Tim13]. lineage [PKE15]. Linear [AJC11, AGJ12, HK12, IGMK^{+}16, LA11, LBH12b, LXFW11, LWY^{+}11, LBJ^{+}16, SSB13, TLRB18, TE18, WLT12, WLN^{+}17, WBS^{+}13, WSSC11, XWZB17, YYL^{+}16, GSC18, MJBC13, RAMG15]. Linearization [LP15]. Linearly [PGBT18]. Lines [AGCO13, CS16, GTG17, KSD14a, KGM^{+}10, MBES16, ORT18, OVV10, SGRT12, SEI10, vKvLV11]. Lines-of-sight [AGCO13]. LinesLab [SB19]. Link [BHR17, DRW15, GEY12, SSK16, SBD^{+}15b]. Linkage [NBA19]. Linked [RSM^{+}16, YHGT10]. Linking [ZOM19]. Lip [DBB^{+}18]. Liquid [ATW15, FAW^{+}16, GBW16, SWT^{+}18, UHT18]. List [YHGT10]. Literary [CWG11]. Literature [BH19, CY11, SJB^{+}17, OKK13]. Live [BBIG17, PKE17, SFLP18, ZZCJ14]. Live-Wire [SFLP18, ZZCJ14]. Living [NMM^{+}19]. Lluandudno [LSF^{+}11]. LMA [ACC15]. Load [APM^{+}11, SHQL18]. Lobe [Tok15a]. Lobe-Aware [Tok15a]. Local [BCD^{+}13, BDC18, BJA^{+}15, CSG^{+}18, DBD^{+}13, GKB^{+}11, GAK10, GKOM18, HW16, HCDC18, HDL11, HLL^{+}12, LZL^{+}15, LZFH18, LCG10, LFYX19, LSL^{+}19, PR19, PB11, PD16, PPH12, SBE19b, SKNS15, SKZ13, SJT^{+}19, VS10, VK18, WHS^{+}18, WCB15, ZHLW18, ZWRH14, ZFJ^{+}16, ZR13]. Locality [Bik12]. Localized [ČHM^{+}13, DLS10, DS11b, HYZ^{+}14, MRCB18, WLZH17, BMM^{+}15]. Locally [HCW19, IFL13, JHT14, MS11b, SKPSH13, KP19]. Location [KHI^{+}19, vDHO16]. Location-dependent [vDHO16]. Locomotion [BV19, ECN14, GCY^{+}14, HKSK18, MSZ^{+}18, WPP13, vBE11]. Locus [SPT14]. LOD [DC10, KDCM14, BP13]. LoDs [LN17]. London [ID10, AYWM14]. Long [BTM^{+}19, KTW^{+}13, LDG19]. Long-term [KTW^{+}13]. Looking [JKL^{+}16, MBM13, MIW13, NSRS13, SHD16]. Loops [SLWSS15, TSK14]. LOP [LW18]. Lossless [PHK^{+}10]. Lossy [VP11, BHT19]. Louisiana [CSLG10]. LoVis [ZWRH14]. Low [ACKM16, BCCS12, ESKD14, HK16, JH19, KVS^{+}14, KHKS12, KWN^{+}14, KSS^{+}15, KKR18, LZL^{+}15, LZFH18, MBRHD18, PCX^{+}18, ST18, SB19, SJF11, Tim12, VCRG14, WXL^{+}13, ZFE16, ZFJ^{+}16]. Low-Complexity [Tim12]. Low-Cost [ACKM16, ESKD14, SB19, ZFE16]. Low-Dimensional [JH19]. Low-Discrepancy [PCX^{+}18]. Low-Level [KVS^{+}14, KWN^{+}14, VCRG14, HK16]. Low-Memory [BCCS12]. Low-quality [WXL^{+}13]. Low-Rank [KHKS12, LZL^{+}15, LZFH18, SJF11, ZFJ^{+}16]. LSMAT [RAV^{+}19]. Luminance [ARM^{+}15, KRM^{+}15, MBDC15]. Lyapunov [GKT16]. Lyra [SH14b].

M&M [KWS16, LGZ^{+}16]. Machine [ERT^{+}17, XXLX14]. Made [CY11, PGGM10, WXL^{+}11, FCS^{+}16, LT17]. Magnetic [BSW^{+}12]. Magnification [LZQ13]. Mahalanobis [CG16a]. Majorization [KHKS12]. Makes [AYWM14]. Makeup [BBIG17, SRH^{+}11]. Making [BKB^{+}12, BBAM12, BMPM12, CCC^{+}14, CCH^{+}14, HBA12, REH^{+}11, WSK^{+}19, YIC^{+}12, ZZH15, ZCBK12]. Man [FCS^{+}16, LT17, WXL^{+}11]. Man-Made [WXL^{+}11, FCS^{+}16, LT17]. Management [BMPM12, HSK^{+}10, WDC^{+}10, WKS^{+}14]. Managing [MGB14]. Mandatory [GST14]. Manhattan [LWB14]. Manifold [BEJM15, Gro16, HDF15, LCY^{+}11, MRCB18, NVT^{+}14, RRP15, WL10]. Manifold-Based [WL10]. Manifolds [AH11, AAS17, KJT14, MGB^{+}12, RLB^{+}19, SSN18, YS19]. Manipulating [DKR^{+}14]. Manipulation [BYL16, BEKB15, CZY11, CKHL11, EDPB15, FHW^{+}11, GKB12, GCZ^{+}12, HZF10, JYS19, KKSS15, LTK12, LAFT12, LGZ^{+}16, MRM^{+}18, MCG^{+}19, NSRS13, OVB^{+}15, RHL12, SHL^{+}14, TBW^{+}11, WK12a, YFW12, ZFCO^{+}11, ZHH^{+}15, vKTS^{+}11, RGB^{+}14].

Manipulations [NNRS15]. Manner [CLJ^{+}15]. Mantle [OBH^{+}11]. Manual [SE19]. Manufacturing [BKB^{+}12, HMA15]. Many [DKH^{+}14, HREB11, KBLE19, LCDW16, LXY19, NIDN16, OMW16, SHD15, WKS^{+}14]. Many-Core [LCDW16]. Many-Light [DKH^{+}14, NIDN16, SHD15]. Many-View [HREB11, KBLE19]. Manycore [KWN^{+}14]. ManyLands [AKS^{+}19]. ManyLoDs [HREB11]. Map [And10, BNJ15, BJG^{+}15, FBL16, HO17, MDI^{+}18, RPWO19, RMC17, RHL12, SSKB15, SBE16a, SBC14, WLZT18, WTLY12, ZDJ16, PCBL16]. Map-Based [RHL12, SSKB15]. Mapping [AHMAM15, AASB14, CWY11, Deb18, EWMU13, FPC^{+}16, FBL16, GUS12, GDML13, GDG12, GPGSK18, GI18, GG14, HLS12, HCW19, JRJ11, JH12, JLKL16, KKPL19, KB12, LSMD15, LWS^{+}16, MS12, MP12a, ML17, MBDC15, MGC^{+}16, NC16, PDP^{+}15, PR12, SWP11, SC10, SFY13, SJL15, SFL19, UMM^{+}10, WWL^{+}19, WG12, XWZB17, YDF^{+}10, YLT19, ZFE16, ZHLW18, ZR19, EMU17, GSC18, RTK^{+}14, SSGM17]. Mappings [ARLC^{+}13, AVBC16, CG16c, JHT14, LA11, NMR^{+}18, SHF13, SKPSH13, SBC16, VMTS10, WBCGH11]. Maps [AAB^{+}10, ABK^{+}19, BCD^{+}13, BBH13, BCRA11, BBL12, BM10, CG17, CBSS17, CKS^{+}16, DLGY12, ESKBC17, EBC17, GBKS18, HA19, HG13, HCO18, JSLW14, JHT14, KFLCO13, KMK12, LT16, MR17, MH17, MRMH12, NBCW^{+}11, OBCCG13, RPK^{+}12, REH^{+}11, RLB^{+}19, SVLL10, SBC16, SDMS15, SBC14, SGYF11, SVBC19, SNB^{+}12, SGB13, TBP18, Vax12, WGBS18, WHLW19, WMZ12, WLS13, WTH^{+}13, YJD^{+}18, YMYK14, vKZH13, YFL19]. March [PS10]. Marching [LCDW16, Muñ14, PWH11]. Markerless [FHE19]. MarketAnalyzer [KMJE12]. Markets [ST18]. Markov [MCB16]. Mask [FO12]. Masked [BNJ15, HHGJ15]. Masking [TMHD12]. Mass [GKKT13, GT15]. Mass-Dependent [GKKT13]. Massive [BN12, ND12, PC12, TRAW12, ZFAQ13, ZFA^{+}16, MAG19]. Massively [VBHH13]. Matching [AYWM14, ATCO^{+}10, AVBC16, ADG19, BLP10, BS12b, CDM^{+}17, COC15, CRA^{+}17, DGP17, DLL^{+}10, GAWJ15, HMW^{+}15, HCSC16, HO17, KKBL15, LRBB17, LBBC14, NC16, NO17, OMMG10, OHG11, OMPG13, PCBL16, PSO18, RRP15, RKN10, SY11, SY12a, SC16, SXY^{+}11, SM14b, SL11, ŠBM^{+}10, SCF10, TMRL14, TBW^{+}11, WSSC11, XXZC13, YYL^{+}16, ZYF13, ZST^{+}10, ZFCO^{+}11, vKTS^{+}11, vKZH13, GFK^{+}19]. Matchings [LB19]. MatchPad [LCP^{+}12]. Material [BSH12, BCRA11, DHI^{+}15, FHHJ18, FVHK17, GOPT11, GMM^{+}12, KDCM14, LPG19, MRMH12, MSHD15, MC10a, NRM^{+}12, NSRS13, PCDS12, PR12, SPN^{+}16, SSN18, SKSS14, YZXW12, YWM15, ZZT15, ZXW19]. Material-aware [YWM15]. Materials [ABW^{+}15, ACOM12, BCRA11, DHI^{+}15, FEM^{+}19, GSDC17, HCJ13, HS17, KPD10, LLD12, LT12, LBH12a, LZB17, MG10, NKLN10, NRM^{+}12, OPP10, SARZL10, TLM16, XHC^{+}18, XDR11, ZLW15, dFH^{+}11]. Matrices [BLY^{+}11, BDF^{+}14, LAE^{+}12, OKK13, RS19]. Matrix [AT10, BBR^{+}16, DRW15, GKLS19, HR10, LZL^{+}15, MRAS17, MRS18, TWT^{+}16]. Matryoshka [Jac17]. Matte [BO19]. Matter [Ano16g, Ano16a, Ano16b, Ano16c, Ano16d, Ano16e, Ano16f, Ano17a, Ano17b, Ano17c, Ano17d, Ano17e, Ano17f, Ano17g, Ano18a, Ano18b, Ano18c, Ano18d, Ano18e, Ano18f, Ano18g, Ano18h, Ano19b, Ano19c, Ano19d, Ano19e, Ano19f, Ano19g, Ano19h, Ano19i, OVV10, SJ13a]. Matting [DZC11, EKFM12, GO10, JWL^{+}13, YZL17, SPCR14]. Maximal [EMP^{+}12, ERA^{+}16]. Maximization [ACOM12, JRJ11]. Maximum [KS13b, SK10]. MCFTLE [GKT16]. Me [FLJ^{+}14]. Mean [CK11b, KSBC12, SHF13, TAOZ12, XL10, YB18]. Mean-Curvature [CK11b, KSBC12]. Mean-shift [XL10]. Mean-Variance [YB18]. Meaningful [MSHD15]. meanings [RAmB^{+}19]. Means [FKSS13]. Measure [AASB19, NNN11, SY12b, VR12]. Measured [SSN18, BSH12, SSGM17]. Measurement [GTB^{+}13, HHNČ19]. Measurement-Based [GTB^{+}13]. Measurements [DSH^{+}17]. Measures [AKV15, SA15, BHT19]. Measuring [DCK13, JS10, KARC17, WK17]. Mechanical [UTZ16]. Mechanism [LSZ^{+}18]. Mechanisms [HL15]. Media [CRGZ10, CLY17, ENSD12, Hol15, KPS^{+}14, KF12, SDS^{+}16, SKTM11, SKGM^{+}17, SKMS18, VGS^{+}19, WDM^{+}12, WHD17, YIC^{+}11, GBM^{+}19]. Medial [BS12a, BS12b, CHWL19, DRF12, HCGW14, LW17, RAV^{+}19, YYG18]. Medical [CNCO15, DHS^{+}13, HMTH13, HMP^{+}12, KMM^{+}18, KVD^{+}10, KBT^{+}12, LSBP18, MK11, MHK^{+}19a, MMV^{+}13, MLK^{+}13, MP10, NJB^{+}11, NLB^{+}13, OJMN^{+}19, PEPM12, PBC^{+}16, RWS^{+}10, RPLH11, SMH10, SLB^{+}18, vPJtHRV12]. Medicine [vdCAvW14]. Medium [BAO^{+}19, McC11]. Meeting [Ano10c, Ano11e, Ano12f, Ano13i]. MegaViews [KBLE19]. Melting [IUDN10]. Members [Ano10c, Ano11e, Ano12f, Ano13i]. Membership [DvKSW12]. Membrane [BBBV12]. Memory [BCCS12, CCI13, KSS^{+}15, QYZ17, RWW16, Ros13, TSdSK13, VH15]. Memory-Efficient [QYZ17, RWW16]. Men [BMWW14, DGR^{+}14, YGCO^{+}14]. Merged [SBE16a]. Merging [BW13, CTL13, DMAL10, SSE^{+}14]. Mesh [AMR^{+}17, BLVD11, BPVR11, BLK11, BLP^{+}13, BCGL18, CK10b, CK11a, CYC15, CCW12, CFB16, CBSS17, CH11, DGP17, DRF12, DSWH17, FLL11, GSZ11, GLLR11, JLW10, JKL18, KMD^{+}17, LMM10, Lav11, LT12, LZFH18, LCHB12, MSS11, MSAP15, MKSS12, MLK^{+}13, MRAS17, NVT^{+}14, NL13, PHK^{+}10, PW13, PPH^{+}13, RGG15, Ren16, RMG18, SLA15, SYM10, SPD14, SBCBG11a, TWS^{+}11, TSS^{+}11, TPC^{+}10, VS10, VR12, VVP^{+}16, Vax14, WJB^{+}13, WGH^{+}19, WZCF15, XWL^{+}15, XXS^{+}15, XM15, YYG18, YL11, YWTY12, ZVD10, ZWC^{+}10, ZDZ^{+}15, ZT10, ZWY^{+}13, CH12, GHX^{+}17, LN17, RRS12, VB14b, VD18, VF14, YI10, ZZCJ14]. Mesh-Free [PPH^{+}13]. mesh-volume [LN17]. Meshes [BLK11, BSH15, BHU10a, CK10b, CG17, CLL^{+}13, DBD^{+}13, DG12, EHA^{+}19, GLLR11, HKA15, JBG17, JJKL18, JLW10, KSO10, KBK^{+}10, KFA^{+}10, KFA^{+}14, LBG16, LS10a, LKSD17, MYLZ16, MM19, MS11b, MR12, MRS18, NWHWfD16, PBMG15, PPL13, PP11, PW13, PTP^{+}15, POG13, QYZ17, RRP15, RKSA17, SFFP15, SKNS15, SC16, She12, SGS14, SS15a, SBCBG11a, SO10, TWS^{+}11, UKCB15, VP11, Vax12, VK18, WLT12, XLS^{+}14, ZSW10a, vTKP11, CVDL16, KP19, SNA17, VR12, VMHB14, WIFD13]. Meshing [BL18, CAS^{+}19, DLS10, DMSL11, GHX^{+}17, GSP19, JWWP14, LMPS16, MPP^{+}13, NZH^{+}18, QYZ17, SJP^{+}13, SNA17, TPSH14a, Tak19, XGDC17, YGJ^{+}14, ZJC13]. Meshless [AWO^{+}10, SM14b, YLHQ12]. Mesoscale [LSB^{+}17]. Mesoscopic [FKE13, LMS^{+}16]. Meta [CDA^{+}14]. Meta-Layouts [CDA^{+}14]. Metaballs [GPP^{+}10]. Metabolic [LDB11]. Metalights [FC10]. Metallic [BCRA12]. Metamorphosis [KPRN11, KFA^{+}14, WCX^{+}13]. Metering [GTM^{+}12, GPRS14]. Method [AMT^{+}12, AW13, CK13, FHHJ18, GBW16, HCJ13, HWU^{+}19, HCW19, HZN^{+}18, IPKK13, PCDS12, SSS^{+}12, SLZ^{+}19, UBH14, XWY19, ZXW19, JPCC14]. Methods [ABC11, BXH10, BAA^{+}16, BBR^{+}16, BMO^{+}14, BFR17, Cas12, DKH^{+}14, FPC^{+}16, HLH^{+}16a, KZB19, Kaz15, Kim15, KDC17, LCSS19, LK10, LOM^{+}18, MTM12, MSK14, NGHJ18, Ren16, SWP11, SYM^{+}12, SHD15, WBS^{+}13]. Metric [ESKBC17, HČA^{+}12, JC10, JWC^{+}11, Lav11, LGK16, PR12, RKSA17, TAEE16, Kur15, WGK^{+}19, SLCO19]. Metrics [BBK^{+}18, CLL^{+}13, GHX^{+}17, MMG10, JZW14]. Metro [CY14, WTLY12, WTH^{+}13]. Metropolis [CWY11, HH10]. MIC [ÁSK14]. micro [BSH12]. micro-facet [BSH12]. Microbe [DWT^{+}11]. Microblog [ZAM^{+}16]. microcylinder [IMDN14]. Microfacet [DHI^{+}15, Hd14b, JG19, KGK19, RBMS17]. Microfacet-Based [Hd14b, DHI^{+}15]. Microfacets [WWH18]. microflake [LN18]. Microgeometry [GTB^{+}13]. microscopic [LN18]. Microsoft [WPHC16]. Microsurface [ZXW19]. Microsurfaces [TH17]. Microtiles [KBW^{+}12]. Mid [MCM^{+}12]. Mid-structure [MCM^{+}12]. Min [STP17]. Min-path [STP17]. Mine [SPV^{+}10]. Minimal [LSR17, VF16]. Minimization [BPY16, BSH15, DCYG19, JKJL18, PPL13, WTL15]. Minimum [CVJ15, vGPNB17]. Minimum-Displacement [vGPNB17]. Mining [PZDD19, ZCT18]. Minkowski [CK10a]. MIP [MS12, XWZB17]. MIP-Mapping [MS12, XWZB17]. Mipmap [WHLW19]. MIQP [WFLW18]. MIQP-based [WFLW18]. Mirror [FKR13]. Miscible [SKK10]. Misinformation [WSK^{+}19]. Mission [BSK^{+}17]. Mixed [BSW^{+}14, JTSZ10, MCH13, KBG^{+}15]. Mixer [HZL19]. Mixing [MNP^{+}17, SKK10, RCM^{+}14]. Mixture [WFZ^{+}15, WDR11]. MLS [CGBG13, GB10]. Mobile [CSC^{+}18, JSH^{+}13, KKTD17, KDCM14, PSC10, RKR^{+}16, RPG16, SDHD17]. Mobility [LWL^{+}16a, SHL^{+}14]. Mobility-Trees [SHL^{+}14]. Möbius [BCK18, KLCF10]. Modal [AFK^{+}14, HDBRC17, HMW^{+}15, HHD^{+}12, CGT^{+}15, EGG^{+}15]. Model [AGR19, ACG^{+}17, ALA^{+}18, BSW10, BPV18, CT11, CRC^{+}15, CHH^{+}19, CLCL11, CBTB16, DLGY12, FHE19, GMD10, GSDC17, GCGP18, HENfSYS16, HM15, HVVR18, HBO^{+}10, Hua17, HAGO19, HSK14, HKSK18, IGAJG15, ISYM15, JTRS12, JG19, JYS19, KrJC^{+}11, KKL16a, LGH13, LMG^{+}18b, LS15, MEKM17, NPDD11, NDG17, ÖKB10, PSCN10, PC12, QWZ^{+}19, RE12, SXY^{+}11, SK17, ŠPBV10, SWG16, SJWS13, WFZ^{+}15, WWH18, WPY^{+}19, WW11, WDR11, WZCF15, YL10, YBK^{+}12, ZLM^{+}15, ZWRH14, dFH^{+}11, AKZM14, IMDN14, LN18, NW17, WJG^{+}16, XWG^{+}13]. Model- [NPDD11]. Model-Based [DLGY12, SK17]. Modeling [ADMAS18, ATCO^{+}10, ATF12, BLP10, BSK^{+}13, BŠMM11, BMR^{+}16, BDS^{+}12, CJFH14, CSLG10, CD10, CGS16, CRA^{+}17, DCNP14, DBD^{+}13, DGGK11, DSY10, EASKC18, FWX^{+}13, GMM15, GPGB11, GGP^{+}19, GGG^{+}16b, GLX^{+}16, HK12, HBDP17, HSS17, HRS^{+}16, HGA^{+}10, JTSZ10, JTRS12, JLW10, JPCC14, JL18, KTKM19, KRM^{+}15, KWM15, KSG^{+}15, KPK10, KK11a, KUMY10, LT17, LKC^{+}12, LOM^{+}18, LG13, LLC10b, LSN^{+}14, LLM^{+}17, LSZ^{+}18, LSWW11, LVW^{+}15, LMPS16, LDY10, LCHB12, MVLS14, MGG^{+}10a, MMO16, MWCS13, MWW12, NVH^{+}13, NBA18, PdMJ14, PKS11, PBB^{+}13, RCB^{+}17a, RJT18, SY12a, SSSSW13, STG16, SAG^{+}13, ŠBM^{+}10, SWS12, SKK^{+}14b, SKK^{+}14a, TBTB12, UHT18, UTZ16, VKW^{+}12, Vax12, WL10, WXR^{+}16, WLJ^{+}18, WLM13, WWS^{+}15, XM15, XHC^{+}18, YFW12, YLHQ14, YCXW17, YLD^{+}18, YLX^{+}16, ZTW^{+}12, ZLYL17, ZCL18, ZST^{+}10, ZCF^{+}13, ZXW19, GDAU14, GDGP16, PGVG19, RCM^{+}14]. Modelling [AJC11, BŠMM11, BWK14, EPCV15, FW17, GGP^{+}15, KK14, KFA^{+}10, KK11a, KBT^{+}12, LSWW11, LCM^{+}18, MbMYR15, MMRO13, NMM^{+}19, RKSA17, RMH^{+}18, SM14a, STBB14, SPK^{+}14, VAW^{+}10, Vax14, WLZ13, WC16, YLG^{+}18, YMS10, YLH^{+}14, ZHQH17, ZWJ^{+}19]. Models [Áfr12, ADMAS18, ABCN10, BKD^{+}17, BN12, CWGvW19, CRY11, CS16, CYC15, Cet19, CSaLM13, CP10, CP11, DLL^{+}10, DMNV12, GWO^{+}10, GRP10, HFE13, IGMK^{+}16, JL17, KWM15, LMS^{+}16, LWS^{+}13, LSL^{+}19, MWN^{+}17, MCM^{+}12, MMG18, MG10, MCB16, MBT^{+}12, MK15, NVH^{+}13, ND12, PC12, PBK10, PH13, RCMM^{+}16, SKR^{+}14, STKD12, SLHC12, TRAW12, WMXC19, WD11, WC14, WXW18, XCDR10, YDP19, ZBM^{+}17, ZHK15, ZR13, CBV^{+}14, DFIM15, DMS14, HKM15, LSN^{+}14, LEM^{+}17, MWCS13, RLYL14, TPSH14b, CHWL19]. Modes [NVT^{+}14]. Modifiable [BMS^{+}10]. Modification [KB12, LAS^{+}11]. Modifications [DBD^{+}13]. Modified [FM15, KSBC12]. Modifying [GVS^{+}15]. Modular [GSDG18]. Modulation [WHL10]. MoleCollar [BJG^{+}15]. Molecular [BWH^{+}11, Bru19, BTM^{+}19, CDSS14, FSTR13, GRDE10, HVVR18, KFR^{+}11, LKEP14, LPSV14, LMA^{+}18, LBPH10, PJR^{+}14, SAMG14, SKR^{+}14, SLD^{+}17, vdZLBI11]. Moments [KFH10]. Monochromatic [LLC19]. Monochrome [AASB19, SB19]. Monocular [MD19, ZTG^{+}18]. montage [HHS14, GWO^{+}10]. Monte [BMDS19, BEEM15, BRM^{+}16b, BB17, DWR10, GAM17, GKT16, HCJ13, HD14a, HB19, JEK^{+}19, KS13a, KVS^{+}14, MMG18, MJL^{+}13, MIGMM17, MGN17, NGHJ18, SHZD17, SMJ17, SÖA^{+}19, VAN^{+}19, ZHD18, ZJL^{+}15]. Monte-Carlo [KVS^{+}14, MMG18]. Morphing [AECOK16, CFB16, DSL15, GLHH13, GCLX17, LLN^{+}14, WWL^{+}13]. Morphological [JESG12]. Morphologies [AMYB17, ASB^{+}17, KFM^{+}19]. Morphology [JPK13]. Morse [DFIM15, SN12, Szy11, WIFD13]. Mosaic [CBC^{+}15]. Mosaics [LVJ10, ME13, MFPA15]. Motion [AMYB17, AWO^{+}10, AGR19, ARB^{+}18, ACOHS18, AWCO10, BCN11a, BSC16, BIZ18, BTS^{+}17b, BvTH16, CYC15, CZY11, CCTL12, CKHL11, CYI^{+}12, CKE^{+}12, DLGY12, EHT18, ESKT15, FHE19, FP15, FHW^{+}11, FGT^{+}16, GS14, GPR^{+}15, HWK^{+}10, HL18, HLL^{+}19, HZF10, HHRZ12, HLL16, JKL18, KRMS13, KÓOH13, KSP^{+}18, LMSG16, LDG19, LK17, LWPL15, LYG15, LZ10, MGC^{+}16, MVH^{+}14, NSG11, PHE^{+}11, PLL11, PP10, PM19, RZS10, RTK^{+}14, SCR^{+}18, SHW^{+}18, SHS13, SJF11, USSK11, VB14a, VF14, WLZH17, WWT^{+}16, WGO^{+}14, WLI^{+}12, YL10, ZZT15, ZHK15, ZRJ^{+}15, vBE11, HNJ^{+}14]. Motion-Compensated [PM19]. Motion-flow [PHE^{+}11]. Motional [FSTR13]. Motions [HSK14]. Motivated [HFM16, PBC^{+}16, DER^{+}10, TDMS14]. Motor [CD18]. Mounted [WRK^{+}16]. Mouse [OGW19]. Movement [BMH^{+}12, BSBE17, LRB^{+}15, SAMS^{+}17, SvL16, WVV11, ZFAQ13]. Movements [NBH18a, SK17, AFHdL14]. Movies [DH16]. Moving [KASH13, MS10a, MC17, ZSW^{+}10b, ZBW11, KBÖ^{+}14]. MPM [YLCH18, ZZL^{+}17]. MR [MK11, ZVE^{+}14]. MRI [KPG^{+}16, KBvP^{+}17, KGGP18, MKP^{+}16, dHvPJV14]. Multi [ÁSK14, ATO17, AFK^{+}14, ABCN10, BTST12, BTB13, BBP10, CSFP12, CGT^{+}15, CD10, CWL19, CSŠ18, CKK18, CDS16, CP10, DZD^{+}16, DAD^{+}19, DSH^{+}17, ENMGC19, EAGA^{+}16, EASG^{+}17, EGG^{+}15, FDL14, FvdP15, FL19, FMH16, FNH^{+}17, GDG12, GHWG14, HDBRC17, HFM10, HMW^{+}15, HWF^{+}17, HZL19, HLJ^{+}13, HJS^{+}17, IABT11, ISYM15, JSH^{+}13, KARC15, KH19, KKS^{+}12, KMB^{+}17, KVD^{+}10, KZZM12, LBG16, LWLD11, LeYO^{+}10, LMLG15, LWL^{+}16b, LJL^{+}18, LLLY19, LFYX19, LTK12, LCD10, LHNS18, LLC19, MJBC13, MbMYR15, MMP16, NJB^{+}11, NPCB17, NBA19, PBMG15, PSPM12, PDC^{+}19, RHS^{+}12, RLYL14, SM10, SM11, SHSK16, SGW12, STMT12, SDMS15, SHW^{+}18, SR14, SS15a, SGG16, TIK17, VSK16, WWL^{+}19, WPY^{+}19, WDAH10, WLI^{+}12, XSXM13, XXS^{+}15, XBL^{+}18, YFW12, YLHQ14, YWC^{+}10, YMS10, ZLM^{+}15, ZCW^{+}19, ZXTD10, BRM^{+}16a, HL14, NGM14, PGVG19, TWSM17, VF14]. Multi-Bounce [LWLD11]. Multi-BVH [ÁSK14]. Multi-Channel [CSŠ18, FMH16]. Multi-Chart [CP10]. Multi-Class [ENMGC19, SGW12]. Multi-column [LLLY19]. Multi-Component [ZXTD10, BTST12]. Multi-Core [IABT11]. Multi-Dimensional [BTB13, FDL14, GHWG14, KKS^{+}12, KZZM12, LHNS18, PSPM12, STMT12, SGG16, ZCW^{+}19, TWSM17]. Multi-element [YFW12]. Multi-Field [FL19, CGT^{+}15, EGG^{+}15]. multi-filament [HL14]. Multi-Frame [LWL^{+}16b]. Multi-Grid [KH19]. Multi-Image [DAD^{+}19, NPCB17, YWC^{+}10, DZD^{+}16]. Multi-impact [ZLM^{+}15]. Multi-Jittered [CKK18]. Multi-Joint [KVD^{+}10]. Multi-Kernel [HLJ^{+}13]. Multi-Layer [SDMS15, CSFP12, ISYM15, LCD10].

Multi-Layered [MbMYR15, RHS^{+}12, LLC19]. Multi-Level [ABCN10, HFM10, NJB^{+}11]. Multi-Light [PDC^{+}19]. Multi-Modal [AFK^{+}14, HDBRC17, HMW^{+}15, CGT^{+}15, EGG^{+}15]. Multi-Party [EAGA^{+}16, EASG^{+}17]. Multi-Perspective [YMS10, LTK12, BRM^{+}16a]. Multi-Phase [ATO17, HJS^{+}17]. Multi-Pose [NBA19]. Multi-Projection [TIK17]. Multi-Projector [SM10, SM11]. Multi-Resolution [LBG16, LeYO^{+}10, VF14]. Multi-room [MMP16, PGVG19]. Multi-Run [FL19]. Multi-sample [SHSK16]. Multi-Scale [DSH^{+}17, SR14, WWL^{+}19, WDAH10, XSXM13, CD10, GDG12, MJBC13, YLHQ14, NGM14]. Multi-Sided [SS15a, VSK16]. Multi-Skilled [FvdP15]. Multi-Spectral [LMLG15]. Multi-Step [WPY^{+}19]. Multi-Style [HZL19, RLYL14]. Multi-task [KARC15]. Multi-Texturing [PBMG15]. Multi-Touch [JSH^{+}13, SHW^{+}18]. Multi-User [XBL^{+}18]. Multi-Variate [CDS16, HWF^{+}17, KKS^{+}12, KZZM12, PSPM12, STMT12]. Multi-View [BBP10, CWL19, FNH^{+}17, KMB^{+}17, LFYX19, WLI^{+}12, XXS^{+}15, LJL^{+}18, SM10]. Multibody [ATK17]. Multichannel [JvdGMR19]. Multicore [KWN^{+}14, VOS^{+}10]. Multidimensional [FR11, JPN15, LT16, LMG^{+}18b, NNN11, PEP^{+}11a, PEP^{+}11b, RW18, WKS^{+}14, YWS^{+}14]. Multidirectional [ŠPBV10]. Multifaceted [PDW^{+}14]. Multifield [HHC^{+}13, NNN11]. Multifields [MFL13]. Multifocal [CGM19]. Multigrid [JCK^{+}13, WWF^{+}18, WMRSF15]. Multilayer [MGM^{+}19]. Multilevel [TL16]. Multimodal [LSBP18]. MultiPiles [BHRD^{+}15]. Multiplanar [CSaLM13]. Multiplane [WZH13]. Multiple [AWO^{+}10, BKB^{+}12, BBAM12, CCM19, DWT^{+}11, FKSS13, GRT18, JR16, KPNS10, KPS^{+}14, KSD14a, KFR18, KJT14, LMW^{+}19, LJZX15, LWS18, LXY19, LMGH^{+}13, LPG13, RAmB^{+}19, SS16, SY14b, SHSK16, SDS^{+}16, SKMS18, TMRL14, WHD17, WYKR17, ZAM^{+}16, KBÖ^{+}14, TWMSK18, ZJC13]. Multiple-Bounce [JR16]. Multiple-Scattering [CCM19]. Multiples [BBL12, LHNS18, vdEvW13, BRM^{+}16a]. Multiplex [RMM15]. Multiplexing [CWB^{+}14]. Multiresolution [MS10b, NPW10, SBE16a, SBE16b, SMP13, vTKP11]. Multisampling [JCK16]. Multiscale [CWW^{+}11, GL12, Lav11, Mér11, PFH^{+}18, Rus11]. Multitouch [ATF12]. Multivariate [AKMM11, BPFG11, BSW^{+}14, BHR^{+}19, BCD^{+}10, BFG^{+}17, HV10, KFH10, KKS^{+}12, ME13, NMSL19, PKRJ10, RL14, RL16, RMH^{+}18, RSSL17, SV10, WCH^{+}15, ZLMM16, ZH14, DKG15]. Multiview [FHE19, TF15, CHA^{+}14]. Mumford [BCGL18]. Munich [PS10]. Muscle [KK14, NVH^{+}13, RPPD17]. Muscle-Based [RPPD17]. Museum [LBK14]. Music [WBSH^{+}13]. Mutable [MWCS13]. Mutual [XWY^{+}15]. MyEvents [PZDD19].

Named [EASG^{+}17, SJB^{+}17]. Named-Entity [EASG^{+}17]. Names [KvLB14]. Nanometric [ACAA^{+}19]. Nanoscopic [LBH12a]. Nanostructures [DTS^{+}14]. Narrative [MRL^{+}17, SH14a, ZH12]. Narrow [FAW^{+}16, SWT^{+}18, SDS^{+}16]. Natural [BCD^{+}13, DGGK11, GPGB11, HT11, HZMH14, KRB11, NW13, PKS11, GGG^{+}16b, JCT14]. Naturalness [SLCL19, VVE^{+}10]. Naturalness-Preserving [SLCL19]. Nature [BMWW14, DGR^{+}14, HFM16, WJG^{+}16, YGCO^{+}14]. Navigating [BBP10, HW10, PBK10]. Navigation [ACS^{+}17, CLWM11, GRE11, LLB^{+}10, MSDK12, WBAI19, ZBM^{+}17, LCMP19]. Near [FDL14, FGT^{+}16, KAS^{+}19, NLB^{+}13, TSB16, WLZT18, MM18]. Near-Eye [KAS^{+}19]. Near-Instant [FGT^{+}16]. Near-Isometric [TSB16]. Near-Wall [NLB^{+}13]. Nearly [SBL12]. Nebulae [WLM13]. Needle [HMP^{+}12, KFR18]. Neighbor [KCL^{+}18, PHL^{+}16]. Neighbors [PM16]. Neighbourhood [SL11, VS10]. Neighbourhoods [OW19]. NEREx [EASG^{+}17]. Nested [KP18, LWM^{+}17]. Nesting [Jac17]. netflower [SRG^{+}19]. Nets [CLGS18]. Network [BBR^{+}16, CDA^{+}14, CLWM11, DAD^{+}19, HZL19, JGH11, KAT^{+}19, LQS^{+}19, MBRHD18, MGM^{+}19, NHL16, RG19, SRG^{+}19, TXL^{+}19, WGW^{+}19, YYZZ18, YC19]. Networks [ACA18, BHRD^{+}15, BWK14, BCD^{+}10, BJB^{+}18, CBLW19, CLWM11, DPF16, DGGK11, DRW15, GPGB11, GLK16, GCW15, HGRS^{+}17, HHCJ18, JvdGMR19, KRD^{+}15, hKTL^{+}17, KWÖG18, KG19, KvLB14, KFM^{+}19, LDB11, MSFM16, NAM^{+}17, NHL16, NMSL19, PGS^{+}16, POCM19, RTJ^{+}11, RMM15, RÖPG18, SSK16, SV14, SAAB11, SLCL19, SBM^{+}14, TLFC16, UHT18, WLZX19, XHW^{+}18, YSM19, ZZ17, ZWHK16, BMM^{+}15, OKK13, vDHO16, LLLY19]. Neural [hKTL^{+}17, KWÖG18, KG19, MBRHD18, NAM^{+}17, POCM19, RJGW19, RG19, RÖPG18, SMB^{+}17, TXL^{+}19, UHT18, YYZZ18, YC19, ZZ17]. NeuroLens [ZZ17]. Neuromatrix [Thi11]. Newspaper [KvLB14]. Next [HDF15, WHD17]. No [HČA^{+}12]. No-Reference [HČA^{+}12]. Node [BHR17, GEY12, SSK16, SBD^{+}15b]. Node-Link [BHR17, GEY12, SSK16, SBD^{+}15b]. Node-Link-Group [SSK16]. Nodes [SAAB11]. Noise [BLV^{+}10, BB17, FLJ^{+}14, GLM17, GCSA13, HB19, KS11, KS12a, KS13a, LLC^{+}10a, MIGMM17, PCX^{+}18, RRSG16, WJDZ14, WYKR17, YGJ^{+}14, CG12]. Noise-Adaptive [GCSA13]. Noise-Aware [WJDZ14]. Noisy [FLJ^{+}14, GLK18, KJT14, OW19, WYZC13]. Non [ABG^{+}12, BPVR11, CRGZ10, CYY^{+}11, CGBG13, CCTL12, CJP^{+}19, GSTOG16, GO15, HHCJ18, JCW11, KSBC12, KPK10, KBT^{+}12, LA11, LZL^{+}15, LZFH18, LMPD15, LRB^{+}16, LXFW11, LWY^{+}11, MJBC13, MRL10, OHG11, RAMG15, SM14a, SKZ13, SSB13, SSJ^{+}10, TMRL14, TBKP12, USSK11, VVC^{+}11, WNS^{+}10, WT11, XCDR10, YKM12, YLLL15, ZYF13, ZST^{+}10, ZFJ^{+}16, RRS12, XXY^{+}18]. Non-Circular [KBT^{+}12]. Non-Constant [CRGZ10]. non-homogeneous [RRS12]. Non-Isometric [GSTOG16]. Non-iterative [MRL10, TBKP12]. Non-Linear [LA11, LXFW11, LWY^{+}11, SSB13, MJBC13, RAMG15]. Non-Local [LZL^{+}15, LZFH18, SKZ13, ZFJ^{+}16]. Non-Oriented [CGBG13]. Non-parallel [HHCJ18]. Non-Periodic [SM14a]. Non-Photorealistic [ABG^{+}12, CYY^{+}11, JCW11, LMPD15, USSK11, VVC^{+}11, WT11, YKM12, SSJ^{+}10, WNS^{+}10, XCDR10]. Non-Rigid [LRB^{+}16, OHG11, TMRL14, ZYF13, BPVR11, YLLL15, ZST^{+}10, XXY^{+}18].

Non-singular [KSBC12]. Non-Terminal [KPK10]. Non-uniform [CCTL12]. Non-Uniformly [GO15]. Non-Visual [CJP^{+}19]. Nonlinear [HK12, WWF^{+}18, YLHQ12]. Nonlinearly [BRM^{+}16b]. Nonparametric [PH13]. Nonrigid [BS12b]. Nonuniform [WH17a]. NoRM [HČA^{+}12]. Normal [BM12, BM16, CACB18, JBG19, JSLW14, LZFH18, MSS^{+}10, SSG17, SCQ^{+}19, WLT^{+}17, ZDZ^{+}15, ZWY^{+}13, CFGL16]. Normalization [LIK19, SY12a]. Normals [Hd14b]. Note [Ano13h]. Novel [ABW^{+}15, GPK^{+}12, LL19, LSR17, LZQ13, TZD11, WYY13]. Novel-View [GPK^{+}12, LSR17]. NPR [CS16, HLH^{+}16b, KM16, TLM16]. Number [OHG11]. Numbers [SL11]. Numerical [Kaz15, Kim15, KUMY10, OLF^{+}14, AVBC18]. NURBS [SYT^{+}13].

Object [BSL18, CKS^{+}15, DS11a, DZC11, EBGM12, GCZ^{+}12, GTK^{+}12, IKL^{+}10, IEH^{+}14, KFW19, LSF^{+}11, LDGN15, MBM13, MPM^{+}14, MSAP15, MCG^{+}19, MRS12, PS10, SSB13, SJWS13, WPY^{+}19, XXY^{+}18, YCL^{+}17, ZZWC16, ZCK17, ZXZ^{+}17, JCT14]. Object-Centered [CKS^{+}15]. Object-Space [MRS12]. Objective [TAEE16]. Objectives [SPF^{+}19]. Objects [AWO^{+}10, Att15, ATF12, BvTH16, EDPB15, FCS^{+}16, GPG^{+}16, HK16, HCSC16, Jac17, JZYP18, KKBJ16, KKSS15, KMHG13, KKPL19, KS12b, LPH^{+}15, LAFT12, LTR19, MCHW18, MGG10b, MMO16, PDP^{+}15, RAMG15, TXGW19, UTZ16, WWH^{+}10, WXL^{+}11, WK12a, WDZ17, WWT^{+}16, WW11, YH13, ZXTD10, NW17, ZKWG16]. Oblivious [SSE^{+}14, DLL^{+}10]. Obscurance [Tim13]. Observation [BRB^{+}13]. Observer [WWV17]. Obstacles [Man16]. Obtuse [XYLY19]. Occluded [TXGW19]. Occluder [SSLL14]. Occluding [ABK^{+}19]. Occlusion [IH11, LK10, LWDB10, MSW10, NJC^{+}19, ŠPBV10, SGG15, YWC^{+}10]. occurrence [SCD^{+}16]. Ocean [AHG^{+}19, BNH10, DCGG11, LSB^{+}17, WHP^{+}11]. OCT [GHB^{+}17]. Octree [GSP19, JZYP18, KKK18, LMP^{+}10]. Octree-Based [GSP19, JZYP18]. Octrees [BLD14a]. off [VVE^{+}10]. Offline [BMDS19]. Offs [BMPM12]. Oil [SDHD17]. Olfaction [HBRD^{+}18]. Omnidirectional [XLH^{+}13]. Omniscient [LPSV14]. On-demand [GLX^{+}16]. On-line [DZD^{+}16, HNJ^{+}14]. On-Site [LPG19]. On-The-Fly [OAM^{+}18, LZB17, ZOA^{+}18, DJSJ19, SKK^{+}14b]. One [OMMG10, SHSK16, Kur15]. one-dimensional [Kur15]. One-sample [SHSK16]. Online [AWCO10, JKJL18, RWW16, SHS^{+}17, WHS^{+}18, XXY^{+}18, ZZZ^{+}19]. Ontological [CE19]. Ontology [FW17, XWG^{+}13]. Ontology-Based [FW17, XWG^{+}13]. Opacity [GSE^{+}14b, GTG17, GRT14, WZC^{+}11]. Opaque [IPKK13, WWT^{+}16]. Operations [DI18, KCH19, LMM10]. Operator [ABCCO13, BEKB15, GRT18, HS19, KKB18, LJC17, MEMO14, PPH^{+}13, QCW^{+}18, HP11]. Operators [BNH^{+}16, BEKB15, BHU10b, EWMU13, HCO18, JTSZ10, MBDC15, UMM^{+}10, YI10, YDT^{+}18]. Opportunities [DAF^{+}18]. Optical [DYFX19, IGAJG15, TBKP12, VWH18, WHL10, LCMP19]. Optics [CRGZ10, HHH12, MTAM12]. Optimal [BSH15, FDL14, HW16, KPRN11, MBB19, Mér11, Meu19, SSG17, SEG^{+}14, She12, YIC^{+}11, dGCSAD11, MM18]. Optimality [HHC^{+}13]. Optimising [VP11]. Optimization [ADG19, BHH13, BLK11, CG12, ESKT15, GP18, GSE^{+}14b, GTG17, GLGW12, IEGT17, JHT14, JL19, KNL^{+}15, Kim15, LMW^{+}19, LEE17, LMLF15, MCHW18, MJBC13, MSAP15, MB18, NBCW^{+}11, NSRS13, PW13, POG13, PSO18, RZS10, RS19, SBC^{+}17, TE18, TIK17, VMH^{+}13, WJB^{+}13, WCT^{+}15, WW16, XWL^{+}15, YGJ^{+}14, YCXW17, YWHB18, ZXZ^{+}17, GSC18, GRT14, ZKWG16]. Optimization-Based [XWL^{+}15, RZS10]. Optimized [BM15, DTV15, HDL11, MPCG12, NREM14, TWS^{+}11, VB14a]. Optimizer [VGB^{+}14b]. Optimizing [BCRA11, CHA^{+}14, CTHAM10, KRMS13, LCWCO10, MWTI19, RGB^{+}14, WZK16]. Orbit [SAD^{+}16]. Order [GO15, GRT19, IH11, KASH13, LPG13, Muñ14, NM14, NPW10, ORT18, POS^{+}11a, RW18, SVLL10, SBF15, SK10, Sch11, SK16a, ÜFE10, YHGT10, BRM^{+}16b, MRL10]. Order-Independent [SBF15]. Orderability [CAB^{+}16]. Ordered [CAB^{+}16]. Ordering [DH16, FR11, WTMT18]. Organic [ABW^{+}15]. Organization [PEP^{+}11a]. Organizing [AAB^{+}10, SWS12]. Orientation [CCLN10, JBG19, PD16, RHLH18, WSSC11]. Orientations [SSG17, RGB^{+}14]. Oriented [CGBG13, MC10a, PHK^{+}10, SSW14b, YXX14, KBK^{+}10, LXY19]. Origami [KTKM19, MDI^{+}18]. Origin [BBBL11, ZFA^{+}16, ZSJT19]. Origin-Destination [BBBL11, ZFA^{+}16, ZSJT19]. Original [TLFC16]. Orleans [CSLG10]. Ornamental [ZWXL17]. Orthogonal [DKMT18, JEK^{+}19, SGRT12, MVPG11]. Oui [ZCW^{+}19]. our [Thi11]. Out-of-Core [BLD14a, DC10, GG14, KTO11]. Outdoor [CLG^{+}18, ESKT15, GPK^{+}12, WXL^{+}13, XZP^{+}13]. Outer [KTKM19]. Outlier [BFG^{+}17, DWR10, HOB^{+}19, ZCW^{+}19]. Output [BSAP11, EPAS11, SBS^{+}17, GGM12]. Output-Sensitive [BSAP11, SBS^{+}17, GGM12]. Outstanding [Can11]. Overall [ZHLW18]. Overlap [Meu19, SSS^{+}12, vGPNB17]. Overlapping [DSWH17]. Overloaded [DKMT18]. Overview [BCN11b, GRE11, OJS^{+}11, ZCH^{+}17, LEM^{+}17].

Packable [WPGSH18]. Packaging [YH13]. Packet [GL10a]. Packing [Att15, ERA^{+}16, MCHW18, NS11, XDY18]. PackMerger [VGB^{+}14b]. Padding [CAS^{+}19]. PAG [CC14]. Paint [FLL11, SGSP15]. Painted [LCC^{+}18]. Painterly [LPSB18]. Painting [CLJ^{+}15, CLLC15, HZF10, KKK18, LMPD15, LFA^{+}15, PFC15, SLE17, SKNS15, SDHD17]. Paintings [WTLL13, ZCC14]. Pair [BMB15b, YWF^{+}19]. Pairs [WKM15]. Pairwise [CSC^{+}18, PD16]. Palette [JYS19, WLX19]. Palette-based [WLX19]. Palettes [KM16]. Pan [TLW^{+}19]. Pan/Zoom [TLW^{+}19]. Panoptic [YC19]. Panorama [BB14, BCK^{+}12]. Panoramic [LWL^{+}16b, PSHZ^{+}15, DGR^{+}14, PGVG19]. Paper [Lar10, PdMJ14, RLYL14, SRH17]. Papercraft [XCDR10]. Papers [ML17]. PaperVis [CY11]. Papilio [LFC14]. Paraboloid [WHLW19]. Paradigms [SE19]. Parallax [Ake11, LMSG16]. Parallel [AIAT12, CG12, DCK12, DG12, FP15, GMM15, GRT18, GRPF16, HBW11, HMB17, HREB11, IABT11, JBG17, JBG19, JZYP18, JR16, KARC15, KVS^{+}14, KMA^{+}19, KZZM12, Kuh12, LCDW16, MB18, MKSS12, MRAS17, POS^{+}11a, PTO10, SN12, SMM13, SKK^{+}14a, VBHH13, VOS^{+}10, WWF^{+}18, WDZ17, HHCJ18]. Parallel-Coordinates [GRPF16]. Parallelization [LL18]. Parameter [AGR19, BPFG11, BvLBS11, ERHH11, GKHF14, GAWJ15, LBH12b, LRB^{+}15, SBC^{+}17, TFA^{+}11, WGO^{+}14]. Parameter-Dependent [LRB^{+}15]. Parameterisation [SJ13b]. Parameterization [CK14, CBSS17, GDG12, HLS12, MS12, MP12a, NRP11, Sch13, SSN18, SLZ^{+}19, PSP^{+}14]. Parameterization-Aware [MS12]. Parameterizations [CK11a, GUK^{+}17]. Parameterized [KAT^{+}19, KSKL13, LPG10]. Parameterizing [AKZM14]. Parameters [CEX^{+}18, DHI^{+}15, ESKT15, JCW11, SSM12]. Parametric [GMM15, KWM15, WDR11]. Parametrization [AMS16, CIE^{+}16, CLS16, CBSS17, ESKBC17, MC19, SGRT12, UFK13, YFWR11]. Parametrizations [CIE^{+}16]. Parametrized [TE18]. Parcels [VKW^{+}12]. Pareto [HHC^{+}13]. Parity [SY12b]. Part [JTRS12, LVW^{+}15, RMG18, XXM^{+}13, ZCOM13, vKTS^{+}11, DMS14, NW17, ZCOAM14]. Part-Based [LVW^{+}15, JTRS12, RMG18, DMS14]. part-segmented [NW17]. Partial [CRA^{+}17, GAWJ15, LRBB17, MH13, PSPM12, RLH^{+}18, RCB^{+}17b, SY14a, SGS14, SDKG18, vKZH13]. Partially [TXGW19]. Participating [ENSD12, Hol15, KPS^{+}14, KF12, SKTM11, SKGM^{+}17, SKMS18, VGS^{+}19, YIC^{+}11]. Particle [AIAT12, BSW10, CZY11, DGP17, IPKK13, IUDN10, KS18, SCN^{+}16, SMM13, SGG15, UBH14, VMH^{+}13, WAF^{+}11, YLHQ14, YWTY12, ZLKW13]. Particle-Based [AIAT12, BSW10, DGP17, IUDN10, SCN^{+}16, SGG15, ZLKW13]. Particle-Grid [IPKK13, UBH14, YLHQ14]. Particles [ELPH19, GT15, HYZ^{+}14, KS14, RGG18, UBH14, WWH^{+}14, YLHQ14]. Partition [DLGY12, MHA17]. Partitioned [GMAG15]. Partitioning [Cam17, EBGM12, FP15, KFW19, YIC^{+}11, GD16, VF14]. Partitions [MS10b]. Parts [GLX^{+}16, SJWS13, TMRL14]. Party [EAGA^{+}16, EASG^{+}17]. Pass [LCD10]. Passive [LPSV14]. Patch [CCM19, FACO17, RDK13, RSK10, VSK16]. Patch-based [RSK10]. Patch-Collaborative [RDK13]. Patch2Vec [FACO17]. Patches [AJA11, Cam17, DSWH17, GGV^{+}19, LCC^{+}18, MVZ16, SS15a, TPSH14a, SZ18]. PatchMatch [BO19]. PatchMatch-based [BO19]. Path [APP10, BYM18, BMDS19, Bik12, BNH^{+}16, CGH18, CBTB16, ENSB13, HOB^{+}19, HEV^{+}16, HKL17, KD13, KF12, LMK^{+}15, MHK^{+}19a, MMG18, MFW18, MGN17, NB15, PBPP11, PBE18, PRDD15, RI17, SKTM11, YIC^{+}11, ZRJ^{+}15, NDD14, STP17]. Path-Preserving [APP10]. Path-Searching [RI17]. Path-space [ZRJ^{+}15]. Path-Tracing [PBPP11]. Pathfinder [PGS^{+}16]. Pathline [HSJW14, MWS^{+}10]. Pathology [JSH^{+}13]. Paths [BLP10, HKD15, HDF15, PGS^{+}16, PRDD15]. Pathway [HMP^{+}12, LDB11]. Pathways [BWH^{+}11]. Pattern [TPSH14a, TLH19, WESW17, ZWRH14, ZCT18]. Pattern-Based [TPSH14a]. Patterns [AASB19, AAB^{+}10, BHRD^{+}15, BCBSG10, BSBE17, CSG^{+}18, DJSJ19, Gda17, Gia18, LKD^{+}17, MBES16, MSFM16, RL16, STG16, XCDR10, ZFAQ13, ZFA^{+}16, ZLMM16, dGWB^{+}14, ZLL13]. PC [KPG^{+}16, KBvP^{+}17, KGGP18, MKP^{+}16]. PC-MRI [KGGP18, MKP^{+}16]. PCA [GKB^{+}11, NKLN10]. PCP [HV10]. PCPN [GKOM18]. PCs [SPH11]. PDE [BF15, LJZX15]. PDE-Based [BF15]. Peak [HJM^{+}11]. Peak-Preserving [HJM^{+}11]. Pedigree [GKPL11]. Peeling [LCD10]. Pen [BH15]. Pencil [YSM19, YKM12]. PencilArt [GTL^{+}18]. Penciling [GTL^{+}18]. Pendulum [HSK14, HKSK18]. Penetration [NL18]. Per-Vertex [MTAM12]. Perceived [DBS^{+}18, JVS^{+}12, VP11]. Perceiving [SBLC17]. Perception [CWB^{+}14, ER18, FVHK17, GEY12, HK16, HDBRC17, HHD^{+}12, JKLS10, KRMS13, KRM^{+}15, LHH^{+}13, MG10, MC14, MSK14, MK15, OAJ14, RGG15, RB10, RPA^{+}15, SW11, SGRT12, SCCN11, SLH^{+}18, TX16, VCL^{+}11, VR12, WSR^{+}17, WTLL13, XCDR10, YMM10, YLRC10]. Perception-driven [WSR^{+}17]. Perceptual [BGCP11, CAB^{+}16, CLL^{+}13, HČA^{+}12, JVS^{+}12, KNH^{+}18, LBH12b, MG10, PHM^{+}14, RE12, SCCN11, TMHD12, HHS14]. Perceptually [DER^{+}10, HFM16, KYKL14, LAS^{+}11, LBH12b, MPCG12, PBC^{+}16, TDMS14]. Perceptually-based [KYKL14]. Perceptually-motivated [DER^{+}10, TDMS14]. Perfect [HKA15, RRP15]. Perfopticon [MHHH15]. Performance [ASL^{+}19, ABW^{+}15, ASB^{+}17, BS19, BAA^{+}16, CWL19, DBB^{+}18, FHW^{+}11, GKB^{+}11, HV10, LK17, LCP^{+}12, LS15, Ros13, SMM13, SMP13, VHB16, WLI^{+}12]. Performance-Based [LK17]. Performance-Driven [BS19]. Performances [LWS^{+}13]. Performative [PMG13]. Performer [ACC15]. Peridynamic [XHC^{+}18]. Peridynamics [CZZ^{+}18]. Peridynamics-Based [CZZ^{+}18]. Periodic [SM14a]. Permissions [LFC14]. Permuting [HB19]. Permutohedral [ABD10]. Persistence [GSW12]. Persistent [DLL^{+}10, RL15, RL16, Kur15]. Person [BBGB19, CCC^{+}14]. Person-Specific [BBGB19]. Personal [PZDD19, RGM^{+}18, vdCvW17]. Personalized [MK15]. Persons [LJL^{+}18]. Perspective [CGM19, IEH^{+}14, SSK16, YMS10, BRM^{+}16a, LTK12]. Perspectives [BCBB16]. PET [DHS^{+}13]. PET/CT [DHS^{+}13]. Phase [ATO17, AKS^{+}19, HJS^{+}17, NSM19, PSP10, PNVS17, RGG18, QWZ^{+}19, SR19]. Phase-Based [PNVS17]. Phenomena [BP19, DGGK11, GPGB11, HT11, KRB11, PKS11]. phenomenological [WJG^{+}16]. Phenotyping [HPvU^{+}16]. PHOG [BCGS13]. Phong [OT11]. Phosphorescent [NSR17]. Photo [GKB12, GCZ^{+}12, GLGW12, IRWM17, KLW12, LTK12, LCWCO10, RHL12, ZH12, YGCO^{+}14]. Photoelasticity [BES15]. Photograph [DSY10, RGM^{+}18, XZP^{+}13, XDR11, ZCC14]. Photographic [EWK^{+}13, SLKL14]. Photographs [IIX^{+}19, LP15, YKM12]. Photography [ABD10, AVR10, BAAR13, CG16b, EWK^{+}13, FO12, GTM^{+}12, GLCC18, HKW12, LE13, MP12b, PK10]. Photometric [BCGS13, SSB^{+}14, ZLSW17]. Photon [CWY11, FSES14, GUS12, GPGSK18, GG14, HCJ13, HGNH17, JRJ11, MGJ^{+}19, SJ13b, SJL15, WG12, YWC^{+}10]. Photons [Duc14]. Photorealistic [ABG^{+}12, CYY^{+}11, JCW11, KBG^{+}15, LMPD15, USSK11, VVC^{+}11, WT11, YKM12, SSJ^{+}10, WNS^{+}10, XCDR10]. Photos [CSC^{+}18, GKB12, LTK12, MJH^{+}17, SLC^{+}19]. Photovoltaic [ABW^{+}15]. Physical [ABW^{+}15, BS19, DPD^{+}15, KBHM15, KBB^{+}17, LLC^{+}12, LMK^{+}15, SB19]. Physically [AFHdL14, BKY^{+}16, CIPT14, GCGP18, HNJ^{+}14, MWN^{+}17, MSHD15, NBH18a, PdMJ14, SHP^{+}19, TSK14, WKBB18, WW11, WWD15, MDBS14, KK18]. Physically-Based [PdMJ14, AFHdL14, CIPT14, GCGP18, HNJ^{+}14, TSK14, KK18]. Physics [ARB^{+}18, GP12, HMLP13, LCCC13, WBT19, ZDM^{+}14, dHvPJV14, HNJ^{+}14]. Physics-Based [ZDM^{+}14, dHvPJV14, ARB^{+}18, LCCC13, HNJ^{+}14]. PhysioEx [KCJM16]. Physiological [KCJM16]. Picking [RGSK10]. Pictograms [LKZ^{+}15]. Pictorial [RHL12]. Picture [GKB12, USSK11, vdKdJP^{+}19]. PICTuReVis [vdCvW17]. Pie [SK16b]. Piece [PEP^{+}11a]. Pieces [SP13, YCXW17]. Piecewise [BdM14, CFGL16, GSC18, MMP16, NBHN17, Szy11, SBL12]. Piecewise-Planar [BdM14, MMP16]. Pigmentation [BW17]. Pigments [GCGP18]. Pile [SM14a]. Piling [BHRD^{+}15]. pipeline [LEM^{+}17]. Pipelines [SBF15]. Pix [BBAM12]. Pixel [DFY14, HB19, IGMK^{+}16, JMD15, KCL^{+}18, KLC^{+}15, KYC16, OJS^{+}11, TBP18, YIC^{+}12, LGC^{+}19, KLC^{+}15]. Pixel-Aligned [KCL^{+}18]. Pixel-based [OJS^{+}11]. pixel-level [LGC^{+}19]. Pixel-Precise [TBP18]. Pixels [EMA^{+}13, MS13, PK10, TM13]. PixelSNE [KCL^{+}18]. Placement [KFR18]. Plagiarism [RPSF15]. Planar [BPY16, BdM14, CG17, DSL15, HCW17, HBA12, NC16, OLA16, SP13, SBC16, SSLL14, SBCBG11b, ZSW10a, vKvLV13, MMP16]. Planarization [POG13]. Plane [AMAM13, VMTS10, vKvLV11]. Planets [CPGG19, DGGK11]. Planned [TBW^{+}11]. Planner [HMP^{+}12]. Planning [AGR19, BSK^{+}17, CY14, HMP^{+}12, HLL^{+}19, KKSS15, LSR17, MP10, SRM^{+}19, SMS^{+}17, WBFvL17]. Plans [WKS^{+}14, PGVG19]. Plant [YHL^{+}16]. Plants [HBDP17, WC16]. Plaque [GOH^{+}10]. Plastic [ZLKW13]. Platforms [BAA^{+}16, JSH^{+}13]. Plausible [CCM16, GVS^{+}15, KTKM19, WW11]. Plays [ODS18]. Plenoptic [GL10b, WILH11]. Plot [BFG^{+}17, MHDG11]. Plots [AKMM11, GKL17, GRPF16, KARC15, KZZM12, LAE^{+}12, RSSL17, SMSL17, TWD^{+}13, WZL^{+}12]. Poetry [ARLC^{+}13]. Point [AP10b, AMT^{+}12, AW13, BSW^{+}12, BTS^{+}17a, BTP13, BM12, BM16, BJFadH19, BB12b, CCLN10, CG12, CACB18, DKL10, DGQ^{+}12, DMSL11, DvKSW12, FHHJ18, GCSA13, GKOM18, HCDC18, JBG19, JK13, KS11, KS12a, KMJE12, KTO11, KJT14, LA13, Lar10, LK19, LCG10, LIK19, LGW18, MEMO14, MW11, MBR^{+}13, MC17, MAM14, MSS^{+}10, OMMG10, PCBS16, QCW^{+}18, RL14, RDK13, RÖG17, RÖPG18, SPB^{+}19, SJ19, SYM10, SSG17, SSW14b, SY12b, SHS13, SCQ^{+}19, TOZ^{+}11, TLH19, WYZC13, WHB^{+}13, WXL^{+}13, WMB15, WXR^{+}16, WXW18, YGJ^{+}14, YLX^{+}16, ZCBK12, DCV14, Kur15, NGM14]. Point-and-Swipe [LCG10]. Point-Based [BB12b, KTO11, WMB15, NGM14]. Point-Cloud [RDK13, SHS13]. Point-Folding [ZCBK12]. Point-Light-Based [DKL10]. Point-Sampled [AP10b]. Point-to-Point [LK19]. Pointcloud [MAM14]. Pointerless [LMP^{+}10]. Pointillism [WTLL13]. PointProNets [RÖPG18]. Points [COO15, CCM16, EMA^{+}13, FDL14, GST14, GTG17, HTG14, LS16, MS13, OM13, SSW14b, TM13, WRS^{+}13, ZSW^{+}10b, vKvLV11]. Pointsets [MDD^{+}10]. Pointwise [RMC17]. Poisson [LS10a, EMP^{+}12, HLL^{+}12, Kaz15, WMRSF15, Yuk15]. Poisson-Based [LS10a]. Poisson-Disk [EMP^{+}12]. Polar [CLCL11, MM19, MR12, SYT^{+}13]. Polycube [CLS16, CAS^{+}19, ZLW^{+}19, FBL16, GSZ11, YFL19]. Polycube-based [CAS^{+}19]. PolyCube-Map [FBL16]. PolyCube-Maps [YFL19]. Polygon [DBK11, HKA15]. Polygonal [CK10a, CK10b, CG17, KYM19, MCM^{+}12, MRD12, MR12, MEKM17, SFFP15, WDZ17, YMM10, YI10]. Polygons [HKA15, KP15, RRP15, ZJC13]. Polyhedra [HL19, MVPG11]. Polyhedral [KFA^{+}14, Vax12, Vax14]. Polyline [XLS^{+}14]. Polyline-sourced [XLS^{+}14]. Polymers [AAS^{+}16, RPK^{+}12]. Polynomial [HHH12, MS13, MRS18, NBHN17, PPL13, XTZ19]. Polynomials [DVPSH14, SHD16]. Polytopes [AGJ12]. PolyVector [DVPSH14]. Poorly [VSG^{+}13]. pop [RLYL14, LJBA13]. pop-up [RLYL14]. Popular [SLC^{+}19]. populating [JPCC14]. Population [MKMA19]. Populations [SCD^{+}16]. PORGY [PMD12]. Porosity [RPK^{+}12]. Portal [BNJ15]. Portal-Masked [BNJ15]. Portrait [SHJ^{+}16, WWL^{+}13, ZCC14]. Portraits [BAAR13, CJZW12, LCC^{+}18]. Pose [DLL^{+}10, ESKT15, HFE13, LJL^{+}18, LLD10, NBA19, VF14, vMRBPM17, CH12]. Pose-oblivious [DLL^{+}10]. Poser [vMRBPM17]. Poses [KL14]. Posing [LWX^{+}15]. Position [ATO17, BMO^{+}14, Cet19, DKWB18, FP15, RF15]. Position-Based [ATO17, BMO^{+}14, Cet19, RF15, DKWB18]. Positional [PRW11]. Positioning [ACS^{+}17, TAAP^{+}16]. Positive [KASH13]. Positive-Definite [KASH13]. Possible [CG16b, CKS^{+}15, SBCBG11b, WCX^{+}13, YMYK14]. Post [AVR10, MRD12, RWS^{+}10]. Post-Capture [AVR10]. Post-Interventional [RWS^{+}10]. Post-Process [MRD12]. Power [ACKM16, DHI^{+}15, ZOA^{+}18]. Power-Aware [ZOA^{+}18]. Practical [ACKM16, BBGB19, CHK13, CBTB16, GBW16, HKW12, KK18, LE13, LMG^{+}18b, LCDW16, MGN17, NSR17, RGSK10, SFL19, TPC^{+}10]. Practice [CBK^{+}17, MTVJ11]. Pre [LWLD11, SNRS12, LN17]. Pre-computed [LWLD11]. Pre-convolved [SNRS12]. pre-filtering [LN17]. Prealiasing [VCRG14]. Precise [TBP18]. Precision [Jac19, KCL^{+}18, MRS12]. Precision-Cut [Jac19]. Precomputed [CRC^{+}15, LPK13, SBE16b]. Precomputing [HR10]. Preconditioning [CBSS17]. Predict [BKM^{+}19, ČHM^{+}13]. Predicted [SGYF11]. Predicting [FVHK17]. Prediction [BPFG11, CH11, FKRW16, HJM^{+}11, RCMM^{+}16, TSdSK13, ZCL18]. Predictions [CWGvW19, CD18, KBL19, LMK^{+}15, RG19]. Predictive [CMF18, HENfSYS16, HBRW^{+}12, LGH^{+}17, OVB^{+}15, SKZ11, SO10, HNJ^{+}14]. Predictors [VS10]. Preferences [KLK17]. Premixed [OLF^{+}14]. Preparatory [ZLDM16]. Preprocessing [BHR^{+}19]. Prescribed [BKB^{+}12, GLK16]. Prescription [vdCAvW14]. presence [TSK14]. Presentation [CKE^{+}12]. Presentations [KLK17]. Preservation [LPG10, NO17, NMR^{+}18, VRBC17, ZR13]. Preserve [JKL18]. Preserving [APP10, BHU10a, CWM19, CHWL19, DGP17, DCK13, DGQ^{+}12, GSP19, GUK^{+}17, HJM^{+}11, HTG14, JH12, JLKL16, KMD^{+}17, LDB11, SYM10, SLCL19, UFK13, WYZC13, WCT^{+}15, WZL^{+}17, WSSC11, WPW^{+}11, WWH^{+}14, YFW12, YSC^{+}18, YLL15, ZFCO^{+}11, ZYF10, ZWY^{+}13, ZFJ^{+}16, AVBC18, EKB14, SXLS19, SSGM17]. Presorting [EBGM12]. Pressure [ATW15, WBAI19]. Prevents [SAAB11]. Preview [XWY19, LLD12]. Primal [IEGT17, UKCB15, WIFD13]. Primal-Dual [IEGT17]. primal/dual [WIFD13]. Primary [ZZ19]. Primitive [BLV^{+}10, CZY11, LPK13]. Primitives [DHP^{+}19, GGP^{+}15, HWU^{+}19, KZB19, SBCBG11a]. Principal [CBLW19, HZRS18]. Print [VGB^{+}14b]. Printable [BLS^{+}17, ZKWG16]. Printed [Jac19, XCDR10]. printers [RCM^{+}14]. Printing [ADMAS18, HL14, KFW19, PPW18, RCM^{+}14, RLYL14, WCT^{+}15, WZK16]. prints [HL14, LEM^{+}17]. Prior [vKTS^{+}11]. Priority [KKS^{+}17]. Priors [RXX^{+}17]. Prism [SJ13c]. Prisms [HHGJ15]. Privacy [CWM19, DCK13]. Privacy-Preserving [DCK13]. Proactive [LCM^{+}18]. ProactiveCrowd [LCM^{+}18]. Probabilistic [AGR19, BELD13, BLD14b, DPD^{+}17, HJS^{+}17, PPH12, PRDD15, PWH11, RLGH15, SMH10, SS15b, Lie17]. Probably [KLAB15]. Probe [BMS^{+}10]. Probe-Based [BMS^{+}10]. Probes [CLG^{+}18, MJK17, MS16]. ProbExplorer [SMH10]. Problem [BMS^{+}10]. Problems [RS19]. Procedural [BŠMM11, BWK14, BKD^{+}17, BP13, CPGG19, DGGK11, DBLW15, EPCV15, GMM15, GPMG10, GPGB11, GKHF14, GD10, GDG12, GPG^{+}16, HWA^{+}10, HSS17, Hua17, IMAW15, KWM15, KKS^{+}12, KS18, KPK10, KK11a, KMK12, KBK13, LLC^{+}10a, LLD12, LLM^{+}17, LSWW11, LSL^{+}19, LZB17, NGDA16, NBA18, PDG^{+}19, RJT18, STBB14, ŠBM^{+}10, SPK^{+}14, VKW^{+}12, WCGG18, WC16, XM15, YB18, GDGP16, LSN^{+}14, DCNP14, JPCC14, LSN^{+}14, MVLS14, SKK^{+}14b, SKK^{+}14a]. Process [KK17, MRD12, MJK11, SBD^{+}15b]. Processes [EWK^{+}13, Gda17, OAM^{+}18, MSW19]. Processing [AGJ12, ABCCO13, BCD^{+}13, BPVR11, BS12a, BCK^{+}12, BLK11, BLP^{+}13, BCGL18, BvTH16, BSEH17, ČHM^{+}13, CG16a, CK10b, CK11a, CCW12, CG16b, CCTL12, CNKI13, DZD^{+}16, DTV15, DJM12, EMP^{+}12, EWMU13, FLL11, GO15, GCW15, GLGW12, GLLR11, JBG17, JH12, Jes16, JESG12, JWL^{+}13, Kaz15, Kim15, KKL16a, KBvP^{+}17, LMM10, LCCC13, LWL^{+}16b, LLC13, LLW12, LLSC13, LCHB12, MWS^{+}16, MSS11, MCM^{+}12, MBG^{+}12, MRAS17, NVT^{+}14, Pat16, PSHZ^{+}15, PK15, RTK^{+}14, SY12a, SCM^{+}19, SY12b, SO12, SLHC12, SHJ^{+}16, SBCBG11a, SKWL13, TWS^{+}11, TSS^{+}11, TBKP12, TPC^{+}10, TBTB12, VB14a, VF14, VCD^{+}16, VBHH13, WZH13, WWL^{+}13, WLM13, WHL10, WGO^{+}14, XXZC13, XSQ13, XKHK17, XADR13, YFW12, YL11, ZVD10, ZH12, ZCZL13, ZCF^{+}13, ZFJ^{+}16, LEM^{+}17, YDT^{+}18]. Processor [MH13]. Product [HEV^{+}16, LMLF15, NMR^{+}18, RLB^{+}19]. Production [CBTB16, ENSB13, USSK11]. Products [IFL13]. Profile [JKK^{+}18]. Profiling [DWT^{+}11]. Prog&Play [MTVJ11]. Program [DI18, YB18]. Programmable [HHRZ12, MH13, MGC^{+}16, SBF15]. Programming [Deb18, GCW15, HG13, MTVJ11, RLGH15, SY14b, WSSC11]. Programs [RJT18]. Progressive [BEF17, BJFadH19, CVDL16, CWY11, CKK18, DKL10, DW13, DG12, GG14, GG15, GRR^{+}16, HBW11, HMB17, JRJ11, JWL^{+}13, LJBA13, MGJ^{+}19, NNDJ12, PHK^{+}10, WG12]. Projected [FL19, MMF10, MFL13, UG18]. Projection [Ake11, AASB14, BDF^{+}14, BBIG17, DGP17, DBS^{+}11, GL12, GI18, HHGJ15, KSCN19, LT16, LMG^{+}18b, MW11, PEP^{+}11a, RLH^{+}17b, SCCN11, SPT14, TIK17, WYY13]. Projection-Based [BDF^{+}14]. Projections [BDS^{+}12, HCW17, JPN15, LKZ^{+}15, LMLG15, LWT^{+}15, LBJ^{+}16, MFNP13, MFL13, PCX^{+}18, PEP^{+}11b, PEPM12, TLRB18, WG11]. Projective [LT16, RRSG16, SMSH18, Vax14, XXS^{+}15]. Projector [SM10, SM11]. Propagation [BO19, BHW11, EIKM16, IEK^{+}14, IEKM16, LJH10, MMRO13, VD18]. Properties [BM16, BJG^{+}15, GKOM18, IGAJG15, JKLS10, SLH^{+}18, VVP^{+}16]. Proportion [CEX^{+}18]. prose [OKK13]. Prostate [MK11]. Protein [ABCJ10, BLY^{+}11, BJG^{+}15, HVH^{+}16, KFR^{+}11, SAMG14, SLD^{+}17, VHG^{+}18]. protein-ligand [VHG^{+}18]. Prototyping [RPLH11]. Provenance [BH19, SASF11, SLSG16, DMS14]. proxies [BLD14b]. Proximal [HRD^{+}15]. Proximity [LBK14, LMA^{+}18, MRS12, TCLK12]. Proximity-Based [LBK14]. Proxy [RKR^{+}16]. Proxy-guided [RKR^{+}16]. Pruning [NPDD11]. Pseudo [DH14]. Pseudo-Spline [DH14]. Psychology [TX16]. Publisher [Ano13h]. Pulmonary [KPG^{+}16]. Purkinje [ARM^{+}15]. Purpose [HLJ^{+}13]. Puzzle [GTZM10]. Puzzle-like [GTZM10]. Puzzles [LRB^{+}16, TSW^{+}19, vdKdJP^{+}19]. Pyramid [LFYX19]. Pyramids [And12].

QEM [YFWR11]. QR [LLC13]. Quad [BLP^{+}13, BL18, CK14, PP11, PW13, PPM^{+}16, RRP15, SLE18, ZSW10a, KP19, TPC^{+}10]. Quad-based [SLE18].

Quad-Dominant [PP11, PW13, ZSW10a, KP19]. Quad-Mesh [BLP^{+}13]. Quadrangulation [HZN^{+}18, MPP^{+}13, TPSH14a, WZG^{+}19]. Quadratic [DCYG19, GCW15, KKBL15, ZLM^{+}15]. Quadratures [MBB19]. Quadrics [LTB19]. QuadriFlow [HZN^{+}18]. Quadrilateral [BLK11, Cam17, MYLZ16, MM19]. Quads [HKA15, KP15, RRP15]. Qualitative [ARH12, BBL12, CBK^{+}17, ER18, NLB^{+}13]. Quality [AASB19, BBK^{+}18, BHU10a, Csé18, DBS^{+}18, DW13, EHH^{+}13, GHX^{+}17, HČA^{+}12, JC10, JWC^{+}11, Lav11, LLX^{+}11, MS11a, MTM12, MSW10, PHE^{+}11, REH^{+}11, SBE16b, SSO^{+}10, SA15, SGYF11, SDK^{+}15, TSPP16, TAEE16, ÜFE10, WZK16, ZBW11, BHT19, BAT11, BEM11, KP19, WXL^{+}13]. Quantification [MMNG17, SMB^{+}17]. Quantifying [CJWL19]. Quantitative [ARH12, AKV15, ER18, RSM^{+}16, SLB^{+}18, YYL^{+}16, KGGP18]. Quantization [BM15, FDL14, GMAG15, HFM10, LJBA13, MBJ^{+}15, PGSD13]. Quantized [BB12b, LJB^{+}12, VKJ^{+}17]. Quantizing [CCC17]. Quasar [BAO^{+}19]. Quasi [GAM17, KBB^{+}13]. quasi-harmonic [KBB^{+}13]. Quasi-Monte [GAM17]. Quasiconformal [WMZ12]. Quaternion [Kim15]. Quaternions [McD10]. Queries [AEWQ^{+}15, LMM10, LPK13, MRS12]. Query [MW18a, MHHH15]. Querying [LJH13]. Queuing [KKS^{+}17]. Quotient [OMPG13].

Radial [DBS^{+}11, GL12, IYS^{+}13, MGV11, RSSL17, SMG10]. Radiance [SNRS12]. Radiation [SRM^{+}19]. Radiofrequency [RWS^{+}10]. Radiometric [LP15, LMG^{+}18b]. Radiotherapy [RCMM^{+}16]. Rain [CBKK^{+}19, LQS^{+}19]. Raman [SPB^{+}17]. Random [ERA^{+}16, Gov19, SB13, SR14, ZFE16]. Random-access [ZFE16]. randomized [JCT14]. Randomly [LQS^{+}19]. Range [CKL14, CACB18, GPRS14, MBRHD18, SPB^{+}19, SHG^{+}16, EMU17]. Rank [KHKS12, LZL^{+}15, LZFH18, SJF11, ZFJ^{+}16]. Ranking [AASB19]. RANSAC [HDV19, LWL^{+}16a]. Rao [SKMS18]. Rapid [LJBA13, RPLH11]. Raster [YFWR11]. Rasterisation [LCD10]. Rasterization [AMTMH12, DHP^{+}19, MS11a, MS13, MTAM12]. Rasterized [ND12]. Rate [DBS^{+}18, OBGB11, SBO18, VB14a, DER^{+}10]. Rate-distortion [VB14a]. Rating [KM16]. Rational [KP11]. Rationale [SMvdWvW15]. Rationalization [ZCBK12]. Raw [DMSL11, GCSA13, GKOM18, JBG19, MDD^{+}10, WXR^{+}16, WXW18]. Ray [Áfr12, ÁSK14, BAAM17, BLW11, CRGZ10, CCI13, DKY16, EBR^{+}14, FKE13, GL10a, GPP^{+}10, GAM17, GHB15, HWU^{+}19, HH11, HPMB19, HHH12, IH11, JKL^{+}16, JKK^{+}18, KBS11a, KBK^{+}10, KWN^{+}14, KSS^{+}15, LK10, MW11, MHA17, MBJ^{+}15, MJL^{+}13, Muñ14, NM14, NB15, OT11, PBPP11, PGKS17, PJSH15, TSdSK13, UWA^{+}19, VF16, VGS^{+}19, VHB16, WRK^{+}16, XYM13]. Ray-bundle [TSdSK13]. Ray-Cast [LK10]. Ray-Casted [FKE13]. Ray-Tracing [HHH12]. Raycasting [BES15, Fre18]. Rays [IH11, KPD10, HRRR18]. Raytracing [GPP^{+}10, LCD10]. RBF [KWN^{+}14]. Re [LWL^{+}16b, ZHD18, TAAP^{+}16]. Re-Compositable [LWL^{+}16b]. re-positioning [TAAP^{+}16]. Re-Weighting [ZHD18]. Reactions [LPSV14]. Reactive [HSH16]. Readability [APP10]. Reading [CKH19, MRL^{+}17, OAJ14]. Ready [HBLB17]. Real [ATK17, AMT^{+}12, BSW10, BPA16, BHW11, BM15, BLW11, BNH10, BN12, CRC^{+}15, CCI13, DER^{+}10, FZS^{+}19, GO10, GMAG15, HML^{+}19, HCBW19, HREB11, HR10, HKSK18, IGMK^{+}16, ISYM15, IFDN12, JKLS10, JKL13, JSLW14, JZYP18, KMHG13, KSCN19, KMB^{+}17, KGK19, KK14, KYM19, LDdLRB16, LMP13, LKEP14, LE13, LCC^{+}17, LCP^{+}12, LDGN15, LLD10, LJH13, LTR19, MD19, MO10, MRM^{+}18, MW11, MBM13, MBJ^{+}15, MC10b, MPBM^{+}17, NKLN10, NKF^{+}16, OT11, PDP^{+}15, PBK10, RHL12, SPB^{+}19, SWP11, SYM^{+}12, SBD15a, SM14b, SGEM16, SDHD17, TST^{+}15, Tok15b, VVE^{+}10, VWH18, WWH^{+}10, WMB15, WRK^{+}16, WWH^{+}14, XLL^{+}10, XWZB17, YHGT10, YCL^{+}17, YWY10, ZZT15, ZFE16, ZM16, ZRJ^{+}15, HNJ^{+}14, RTK^{+}14, WJG^{+}16, EHH19]. Real-Time [AMT^{+}12, BSW10, BPA16, BM15, BLW11, GO10, GMAG15, HCBW19, HREB11, HR10, IGMK^{+}16, ISYM15, JKL13, JZYP18, KMHG13, KSCN19, KMB^{+}17, KK14, LDdLRB16, LMP13, LKEP14, LE13, LCP^{+}12, LDGN15, LLD10, MD19, MO10, MW11, MBJ^{+}15, MC10b, MPBM^{+}17, NKF^{+}16, OT11, PBK10, RHL12, SPB^{+}19, SWP11, SYM^{+}12, SBD15a, SDHD17, TST^{+}15, VWH18, WRK^{+}16, WWH^{+}14, XWZB17, YHGT10, ZZT15, ATK17, BNH10, BN12, CRC^{+}15, CCI13, DER^{+}10, HKSK18, IFDN12, JSLW14, KGK19, KYM19, LCC^{+}17, LJH13, MRM^{+}18, NKLN10, SM14b, SGEM16, Tok15b, WWH^{+}10, WMB15, XLL^{+}10, YCL^{+}17, YWY10, ZFE16, ZM16, ZRJ^{+}15, HNJ^{+}14, RTK^{+}14, WJG^{+}16, EHH19]. Real-World [LTR19, PDP^{+}15, FZS^{+}19]. Realism [BKD^{+}17, LCC^{+}18]. Realistic [ARB^{+}18, ACV^{+}14, BNH^{+}16, BNH10, BN12, DKH^{+}14, GMM15, GLHH13, HČA^{+}12, JRJ11, KÖS^{+}15, LJK^{+}12, MGG^{+}10a, MMG18, NIDN16, NDD14, SHSK16, SKZ11, WWT^{+}16, WW11, XLL^{+}10, ZCT18, ZHK15, HD14a]. Reality [AJL^{+}11, BBCW10, CSaLM13, CKE^{+}12, GTB^{+}13, KKK18, KAS^{+}19, MS10b, MJK11, PTW13, RGSK10, SP13, SYC10, YJD^{+}18, KBG^{+}15]. Realtime [BS19, CPZ^{+}15, KBK13, XWB15, YLHQ12]. Rear [CSC^{+}18]. Rearrangeable [YIC^{+}12]. Reasoning [KNH^{+}18, SR19]. Rebalance [XLL^{+}10]. Recall [SSKB15]. Recognition [DLGY12, SBG17, SRG16, ZDJ16]. Recognition-Difficulty-Aware [ZDJ16]. Recoloring [HZMH14, HWL18, KKL16a, NPCB17, WLX19]. Recolourization [DRA10]. Recombination [JTRS12]. Recommendation [WLL^{+}17]. Recommendations [CLWM11, GOB^{+}10, LCSS19, RAMG15]. Reconstructability [JWC^{+}11]. Reconstructing [BSCH18, RK10]. Reconstruction [AIAT12, ACV^{+}14, BPW14, BEEM15, BTS^{+}17a, BLK11, BdM14, CT11, CD10, CCLN10, CWW^{+}11, CLCL11, CCC^{+}14, CFGL16, DLS10, DGQ^{+}12, DMAL10, DFY14, ECN14, FEM^{+}19, GS14, GCSA13, Gro16, GZH^{+}19, GLLR11, HTG14, IKL^{+}10, IEK^{+}14, JJKL18, JSLW14, KBÖ^{+}14, LA13, LDGN15, LEE17, LCD10, LTX^{+}14, MSS11, MVZ16, MPM^{+}14, MC10a, MKU15, MRL10, MWW16, MDD^{+}10, MVH^{+}14, MMP16, MWA^{+}13, NSC14, NJGW15, NDD14, NJC^{+}19, OMW16, PM16, PPT^{+}19, PSDB^{+}10, PGK10, RWW16, RLH^{+}18, RI17, SYM10, SKZ13, SSW14b, SXY^{+}11, SBCBG11a, STC^{+}16, SMG10, TOZ^{+}11, WLT12, WYZC13, WXL^{+}13, WGS10, WCM15, XSX^{+}14, XXY^{+}18, YCL^{+}17, YHL^{+}16, ZFG^{+}17, ZZZ^{+}19, ZHK15, ZTG^{+}18, ZSG^{+}18, ZJL^{+}15, dGCSAD11, vKvLV13, YGCO^{+}14]. Recording [WLI^{+}12, SBB14]. Records [RCB11]. Recovering [BCD^{+}13, BEKB15, PH17]. Recovery [ACKM16, AP10b, DZCC19, RMC17, WGH^{+}19, HNJ^{+}14]. Rectangles [UFK13]. Rectification [ACOM12]. rectilinear [MAG19]. Recurrent [SBL12, TXL^{+}19, YYZZ18]. Recurring [ZCOAM14]. Recursive [GO15, RS19]. Redirection [WBM^{+}18]. Reduced [BNH^{+}16, CGH18, FMD^{+}19, MKU15, SO12, YWHB18, ATW15]. Reducing [GL10b, HF16]. Reduction [LS10a, LMG^{+}18a, LPG13, MIGMM17, PSPM12, RGG15, RL15, XSE14]. Reevaluating [RI17]. Reference [BKM^{+}19, BPKB14, HČA^{+}12, KG19]. referenced [vGPNB17]. refinable [KP19]. Refinement [DRM19, DLS10, DS11b, HMB17, KP18, UKCB15, VK18, ZWRH14]. Refinery [KRD^{+}15]. Refining [IY10, YDP19]. Refitting [VKJ^{+}17]. Reflectance [BB12a, BCRA11, DLD12, FGT^{+}16, GPK^{+}12, NSRS13, PdMJ14, PR12, RSK12, XDR11, dFH^{+}11]. Reflectances [OYH18, BSH12]. Reflection [EKFM12, MS16, MMRO13]. Reflections [LWLD11, XWB15]. reflective [DKR^{+}14]. Reflectometry [LDW^{+}10, RPG16]. Reforming [YWM15]. Reformulating [XHC^{+}18]. Refracted [YIC^{+}12]. Refraction [BSW10, CRGZ10, PSP10]. Refractive [CRGZ10, DZC11, Hol15, DKR^{+}14]. refresh [DER^{+}10]. Region [GSTOG16, GM19, RBC14]. Regional [FHHJ18, SSM12]. Regions [LKEP14, MM19, MRM^{+}18]. Register [VH15]. Registration [ABCJ10, BCK18, CCTL12, MAM14, PR19, PB11, PD16, SM10, SM11, TST^{+}15, TF15, XXY^{+}18, YLLL15, YMYK14, ZST^{+}10, HDV19]. Regression [BRM^{+}16b, JL17, PBK10, SMSL17]. Regression-Based [JL17]. Regressors [DTV15]. Regular [AYLM13, CG17, CBLW19, PRS15, WHWB16]. Regularization [BdM14, JG19, KD13, OLA16, RPWO19, ZFJ^{+}16]. Regularized [RMC17]. Regularizing [MVZ16]. Reinforced [AAS^{+}16, RPK^{+}12, LZY^{+}17]. Reinsertion [MB18]. Reinterpretable [AVR10]. Rejection [DWR10]. Relating [HMW^{+}15]. Relations [ASB^{+}17, BSW^{+}14, XDC^{+}13]. Relationship [EASG^{+}17, LXFW11]. Relationships [BHR17, BSW^{+}14, CTL13, HRD^{+}15, JKL18, KWD14, KGM^{+}10, PMD12, VM12, WDM^{+}12, YWS^{+}14]. Relativistic [JMV^{+}15]. Relax [LBRM18]. Relaxation [KKBL15, SJ13b]. Relevant [LAE^{+}12, PCR11]. Reliability [SBLC17]. Reliable [KBT^{+}12, SCQ^{+}19, WLT^{+}17]. Relief [ABG^{+}12, ABCN10, Gia18, JSLW14, KWC^{+}12, ZWJ^{+}19, ZZWC16]. Reliefs [ASH15]. Relightable [LWS^{+}13]. Relighting [IFL13, NKLN10, SPB^{+}19, VSG^{+}13]. Remeshing [CCW12, DLY^{+}18, JHT14, PRS15, XYLY19, YGJ^{+}14]. Remeshing-assisted [JHT14]. Reminiscence [PZDD19]. Remote [PHE^{+}11]. Removal [ABC11, FZS^{+}19, GTK^{+}12, HOB^{+}19, KSCN19, MWS^{+}16, Meu19, SSS^{+}12, SCQ^{+}19, TIK17, WLZX19, XSXM13, XXZC13, ZZLX17, vGPNB17]. Removing [BB17, KS13a, WYD^{+}13, WYKR17]. Rendered [ČHM^{+}13, LCSL18]. Rendering [AFK^{+}14, ABG^{+}12, AHTAM14, ACV^{+}14, BSW10, BEEM15, BSAP11, BPA16, BKM^{+}19, BOB13, BRM^{+}16b, BM15, BNH^{+}16, BB17, BLW11, BNH10, BN12, BES15, CNCO15, CRGZ10, CSD11, CWW^{+}11, CYY^{+}11, CCM19, CBTB16, CCI13, CMF18, CNS^{+}11, DKH^{+}14, DCGG11, DWR10, DMNV12, DER^{+}10, DHP^{+}19, DBLW15, DC10, DFY14, EBR^{+}14, ENSD12, ESKD14, ER18, FC10, FV14, FLJ^{+}14, FSES14, GD16, GM19, GKB^{+}11, GDML13, GKPS12, GMAG15, GHK^{+}10, GPGSK18, GRC13, GPRS14, GKT16, GRR^{+}16, HML^{+}19, HD14a, HBRW^{+}12, HDBRC17, HBRD^{+}18, HWK^{+}10, HFE13, HREB11, HGP^{+}19, HR10, HHRZ12, HHD^{+}12, IGMK^{+}16, IRWM17, IFDN12, IMDN14, JA18, JEK^{+}19, JKL13, JCW11, JSYR14, JKK^{+}18, KS13a, KVS^{+}14, KCH^{+}18, KWN^{+}14, KBLE19, KDCM14, KBK13, KUMY10, KKR18, LWDB10, LMP13, LVPI18, LE13, LSR17, LMPD15, LBH12a, LLD10, LWBP14]. Rendering [LA15, LKG^{+}16, LPG13, MEMO14, MW11, MS14, MS16, MBM13, MVZ16, MFPA15, MBJ^{+}15, MMF10, MDBS14, MSHD15, MH13, NIDN16, NSG11, NDG17, NPCB17, NPW10, NKF^{+}16, OT11, OKG^{+}10, OP10, PHE^{+}11, PBPP11, PC12, PM19, RHS^{+}12, RGB^{+}14, RKRD12, RRS12, RKR^{+}16, RLH17a, RSD^{+}12, RPLH11, RGG^{+}14, RBDD18, RMZ13, SHSK16, SBE16a, SYM^{+}12, SBF15, SGM^{+}11, SKZ11, SRK13, SDS^{+}16, SHZD17, SHP^{+}19, SÖA^{+}19, ŠPBV10, SKSS14, SGG15, SGEM16, SSJ^{+}10, TDDD18, TW10, TSdSK13, Tok15b, USSK11, VCRG14, VCL^{+}11, VSG^{+}13, VWH18, VVC^{+}11, VGS^{+}19, VBHH13, WWH^{+}10, WK12b, WZKP14, WMB15, WTL15, WZL^{+}17, WSR^{+}17, WG12, WNS^{+}10, WWT^{+}16, WND^{+}14, WT11, WDR11, WYKR17, XZP^{+}13, XSE14, XWB15, XCDR10, YZXW12, YSM19, YDF^{+}10, YKM12, YMS10, YWY10, YIC^{+}11, ZFE16, ZOA^{+}18, ZRJ^{+}15, ZJL^{+}15, CBV^{+}14, GGM12, KBG^{+}15, SD10b, SKK^{+}14b, WZC^{+}11]. rendering [WJG^{+}16]. Renderings [BRM^{+}16b, VAN^{+}19]. Reordering [BBR^{+}16]. rep [CBV^{+}14]. Repairing [HHCJ18]. Repeated [HZZ11]. Repertory [KW18]. Repetition [ACOM12]. Repetitions [RBDD18]. Repetitive [RÖM^{+}15, dGWB^{+}14]. Rephotography [LLC11]. Replaceable [LVW^{+}15]. Replacement [HCBW19]. Reply [EASKC18]. Reply-Chains [EASKC18]. Report [ASW14, Ano13i, Ano16l, BBT11, BCD^{+}12, MNP^{+}17, MKRE16, SRM^{+}19, TDS^{+}16, Ano10c, Ano11e]. Reporting [MHK^{+}19b]. Reports [Ano10a, Ano10b, Ano10d, Ano10c, Ano11c, Ano11d, Ano11e, Ano12d, Ano12e, Ano12f, Ano13f, Ano13g, Ano13h, Ano13i, AJL^{+}11, BCRA12, CSLG10, DS11a, DFIM15, GGG^{+}16a, ID10, IC11, KRP^{+}15, KBG^{+}15, Kuh12, LSF^{+}11, LF11, Pat16, PGGM10, PS10, SJ13c, TDS^{+}16, TAEE15, VCD^{+}16, WWS^{+}15, WBP11, ZJL^{+}15, Ano12f]. Repositioning [IEH^{+}14]. Representation [BBK^{+}19, BÖK11, CDS16, DZCC19, FW17, FACO17, GGG^{+}16a, GLLR11, HVH^{+}16, HO17, JYS19, Jes16, KrJC^{+}11, LCSS19, LDB11, LJB^{+}12, LHH^{+}13, MMNG17, MRL10, MFL13, MRAS17, OLF^{+}14, RLB^{+}19, SARZL10, UKCB15, WHC15, WH17b, WLZT18, XLH^{+}13, CH12, GGM12, GDGP16, MRM^{+}18, WIFD13]. Representations [BFR17, CR16b, HMW^{+}15, HSvK18, SDKG18]. Representative [JPN15, MW18b]. Representatives [SSS^{+}12]. Representing [FACO17, Lie17, RLH^{+}17b]. Reproducible [SLSG16]. Reproducing [OYH18]. Reproduction [KK18, NSR17]. RepSnapping [HZZ11]. Resampling [AMR^{+}17, GB10, NB15]. Rescue [BSK^{+}17]. Research [BSW^{+}14, EHH^{+}13, Lar10, OJMN^{+}19, SLSG16, vLKS^{+}11]. Researcher [Bru11, Eis11]. Residue [LMA^{+}18]. Resolution [AVR10, DBS^{+}18, FGT^{+}16, GBW16, HCDC18, HHNČ19, KCL^{+}18, LBG16, LeYO^{+}10, NDG17, NB12, SJB^{+}17, SKTM11, TTN^{+}13, ZLZ^{+}18, ZBW11, ACA18, DTV15, KKK18, LCC^{+}17, VF14, WGK^{+}19, ZLYL17]. Resolution-Independent [NDG17, NB12]. Resolved [JMV^{+}15]. Resonant [LFK^{+}13]. Resource [DDH^{+}18]. Respect [BKM^{+}19]. Response [DW13, RCMM^{+}16]. Responsive [XLL^{+}10, YL10]. restoration [PSP^{+}14]. Restricted [NL13]. Restricting [SNA17]. Restructuring [WTMT18]. Results [MW18a]. Retailoring [CR16b]. Retargeting [AMYB17, ARB^{+}18, BB12a, BMS^{+}12, CYC15, GTK^{+}12, HCSC16, JKL18, LLX^{+}11, PWS12, RKGS18, RSD^{+}12, WBSH^{+}13]. Rethinking [YLT19]. Retouching [IRWM17]. Retracted [KL19]. Retrieval [CYI^{+}12, DS11a, GKHF14, LSF^{+}11, LDGN15, LS15, PS10, SXY^{+}11, SJF11, TCLK12, WLL^{+}17, ZCK17, BCGS13]. Reusable [MP10]. Reverse [DMS14, PH17, TTB12]. Reverse-Engineering [PH17]. Review [AKS^{+}19, BH19, CGM19, CY11, CHWL19, EHH19, FPC^{+}16, FHE19, FW17, GGP^{+}19, GP12, HH11, HAGO19, JL19, LLLY19, LTR19, RPA^{+}15, WMXC19, EMU17]. Reviewers [Ano10d, Ano11f, Ano12g, Ano13j, Ano14d, Ano15j, Ano16m, Ano17k]. Reviews [WHS^{+}18]. Revisited [BBCW10, BFG^{+}17, KKF^{+}17]. Rewriting [MH13, PMD12]. Reyes [LPD14]. Reyes-Style [LPD14]. RGB [BIZ18, CJXH17, KLTZ16, MD19, ZSG^{+}18]. RGB-D [BIZ18, KLTZ16, ZSG^{+}18]. RGBD [LEE17, LTX^{+}14, WCM15, YCL^{+}17]. RGBZ [RSD^{+}12]. Rib [LZY^{+}17]. Rib-reinforced [LZY^{+}17]. Ribbon [HVH^{+}16, SHCB15]. Rich [JBMC10, SHD15]. Rich-VPLs [SHD15]. Rig [SBC^{+}17]. Rigging [BCB^{+}15, BBGB19, BTST12, LCY^{+}19]. Rigging-Skinning [LCY^{+}19]. Right [AKMM11]. Rigid [Aan18, BET14, JKJL18, KLAB15, LRB^{+}16, MS11b, NL18, OHG11, PKS10, TMRL14, TF15, WCX^{+}13, ZYF13, BPVR11, YLLL15, ZST^{+}10, XXY^{+}18].

Rigs [KBB^{+}17]. River [DGGK11]. Rivers [BKW13]. Riverscapes [PDG^{+}19]. Road [BWK14, GPGB11, NDD14, vDHO16]. Roads [GPMG10, IMAW15, LBA10, NJGW15, NGDA16]. Roadside [ACV^{+}14]. Robot [LMPD15]. Robotics [RPA^{+}15]. Robust [ARB^{+}18, BEEM15, BPVR11, BM12, BM16, CK10b, CFS14, HDL11, HL13, HZN^{+}18, JWH19a, JG19, JMD15, KD13, KG19, KO19, KPG^{+}16, KJT14, LK13, LWL^{+}16b, LGZ^{+}16, MYGY19, MJL^{+}13, MDD^{+}10, PWS12, RBC14, RMZ13, SXY^{+}11, SJ13b, TST^{+}15, Tak19, TE18, VSG^{+}13, VVP^{+}16, WTTM15, ZWC^{+}10, ZYF13, ZNZ19, ZRJ^{+}15, ZVE^{+}14, dGCSAD11, LCLJ10]. Robustness [WRS^{+}13, dSNV^{+}17]. Rods [DKWB18, Jac19, SMSH18]. RodSteward [Jac19]. ROF [WZCF15]. Roles [CKE^{+}12]. Rolled [SSS^{+}12]. Rolled-out [SSS^{+}12]. Rolling [SCQ^{+}19]. Rolling-guided [SCQ^{+}19]. Rom [Csé18]. room [MMP16, PGVG19]. Rotational [MGG10b, SFL^{+}16, AFHdL14]. Rough [CMF18]. Route [WTLY12, ZSJT19]. Route-Aware [ZSJT19]. Routine [ZVE^{+}14]. Routing [LBA10]. RPCA [MC17]. RT [MBJ^{+}15, RCMA^{+}18]. RT-Induced [RCMA^{+}18]. RTSAH [IH11]. Rubber [STBG12]. Rule [ARLC^{+}13, CNCO15, LLM^{+}17]. Rule-based [ARLC^{+}13, LLM^{+}17]. Rule-Enhanced [CNCO15]. Rules [GLCC18, PFC15]. Run [FL19]. Runner [RCMA^{+}18]. Runtime [XWY^{+}15, YWHB18].

SAFE [DCNP14]. SafeGI [OP10]. Safety [SMvdWvW15]. SAH [BM15, GD16]. Sale [KMJE12]. Salience [JC10]. Salience-based [JC10]. Saliency [JKK^{+}18, WCT^{+}15, YCL^{+}17]. Saliency-aware [YCL^{+}17]. Saliency-Driven [JKK^{+}18]. Saliency-Preserving [WCT^{+}15]. Salient [BPVR11, hKTL^{+}17, LKEP14]. Sample [BEJM15, CKK18, GCY^{+}14, JKL13, KKR18, MBB19, SÖA^{+}19, Yuk15, ZZ19, ZHD18, SHSK16]. Sample-Based [BEJM15, GCY^{+}14]. Sampled [AP10b, AAS17, GO15, VSG^{+}13, DCV14]. Samples [CTL13, OMW16, OW19, RSK12, SJ19, ZHD18]. Sampling [BMDS19, BBH13, BHW11, BELD13, BD16, BNJ15, BJFadH19, CWW^{+}11, CWY11, CG12, Csé18, DLS10, EMP^{+}12, EJFadH13, FV14, Gam16, GO10, Gov19, Hd14b, HEV^{+}16, HJS^{+}17, IGAJG15, JEK^{+}19, KS11, KS12a, KF12, KKR18, LK11, LYG15, LXY19, LPG13, MHK^{+}19a, MHD16, MKU15, MFL13, NPCB17, OXKP12, PBE18, RRSG16, RÖG17, SHSK16, SMJ17, SGEM16, SNJ^{+}14, SKTM11, TBW^{+}11, UFK13, UG18, VF16, WZL^{+}17, WWT^{+}16, WND^{+}14, WAF^{+}11, YZL17, YIC^{+}11, Yuk15, ZJL^{+}15]. Sampling-Based [YZL17, LYG15]. Sanity [LA11]. SAR [SRG16]. Satellite [ZLYL17]. Satisfaction [MC19]. SATO [NM14]. Scaffold [JE13]. Scalability [MS11a, PHE^{+}11, ZBW11]. Scalable [BMH^{+}12, DKH^{+}14, FP15, HHRZ12, HZN^{+}18, KBWS13, KMD^{+}17, KBLE19, LDdLRB16, PKS10, PHE^{+}11, SM10, SR19, SPD14, SO10, TW10, UWA^{+}19, VGS^{+}19, WDM^{+}12, WBSH^{+}13, ZCZL13]. Scalar [GST14, HW10, HLH^{+}16a, HHC^{+}13, LS16, LWS18, LKG^{+}16, PRW11, PW12, SSW14a, SW17, WGS10, MAG19]. Scalar-Valued [HW10]. Scale [ADMAS18, ABCN10, BHP15, CIE^{+}16, CBC^{+}16, DMSL11, DSH^{+}17, DWT^{+}11, GLCC17, GLK16, HSBW13, JLKL16, KS18, KHI^{+}19, LS10b, LSS^{+}12, LCB^{+}18, LBH12a, MGB^{+}12, MDI^{+}18, MHDG11, SY12a, SM11, SPH11, SR14, SXLS19, TLW^{+}19, TBP18, WWL^{+}19, WSSC11, WDAH10, WAF^{+}11, XSXM13, ZC18, CD10, DJM12, GDAU14, GDG12, LN17, MJBC13, MPM^{+}14, NGM14, PTA^{+}11, SMM13, TXL^{+}19, YLHQ14, ZBQC13]. Scale-adaptive [SXLS19]. Scale-aware [JLKL16]. Scale-Invariant [CIE^{+}16, ZBQC13]. Scale-like [LS10b]. Scale-Space [MGB^{+}12]. Scale-Stack [HSBW13]. Scales [LS10b, LMSF19, ZAM^{+}16]. Scaling [ACS^{+}17, LHNS18, MKR11, SSDK12]. Scan [DMAL10]. Scanned [BL18]. Scanners [MDBS14]. Scanning [KMHG13]. Scans [CACB18, SJWS13, MPM^{+}14]. Scatter [KARC15, SMSL17]. Scattering [BSW10, BNH^{+}16, CRGZ10, CCM19, DBK11, HCJ13, Hol15, JZJ^{+}15, KPS^{+}14, MESG11, OXKP12, PSP10, SDS^{+}16, SKMS18, WHD17]. Scatterplot [CSG^{+}18, LCSS19, LAE^{+}12]. Scatterplots [AASB19, FH18, HBW11, KZZM12, RB10, SBLC17, SGG16]. Scene [ACS^{+}17, BSK16, CSC^{+}18, CGS16, GG15, HZZ11, KLTZ16, KBLE19, LTX^{+}14, LMLF15, SLC^{+}19, VBHH13, WKM15, XZP^{+}13, ZZZ^{+}19]. Scene-View [KBLE19]. Sceneries [MGG^{+}10a]. Scenes [ASH15, BHMT13, DKB^{+}16, EPCV15, ENSD12, GS14, GTK^{+}12, GPG^{+}16, HCSC16, IFDN12, JVS^{+}12, KBWS13, LMLF15, MGY^{+}18, MC10b, NPDD11, NSRS13, PSC10, RKRD12, REH^{+}11, SHL^{+}14, TSdSK13, WXL^{+}13, WG12, YWC^{+}10, ZCW^{+}15, ZLYL17, ZBW11, vKvLV13, GGG^{+}16b, KBG^{+}15, JL19]. Scheduling [KKS^{+}17]. Scheimpflug [HHNČ19]. Scheme [DRA10, LKSD17, LCY^{+}19, MHD16, WTMT18, WHT12, AVBC18]. Schemes [MM19, QPCRM19, RL15, MM18]. Scholarship [CWG11]. Science [FEM^{+}19, HS17, LSS^{+}12, PPBT12, RPK^{+}12]. Sciences [ABW^{+}15, DPD^{+}15, KBHM15, LLC^{+}12, LMK^{+}15]. Scientific [KGP^{+}12, SASF11, WT17]. Sclow [GKL17]. Scope [BJB^{+}18]. Scratch [WPP13]. Scratched [VWH18]. Screen [HB19, KCL^{+}18, KGRG17, MSW10, NAM^{+}17, UMM^{+}10, XWB15]. Screen-Resolution [KCL^{+}18]. Screen-Space [MSW10]. Scribble [CLLC15, LLC19]. Scribbles [NSS^{+}12]. Sculpting [PW17, JPCC14, MWCS13]. Sculpture [BCRA12]. Sculptures [KLC^{+}15, ZCX19]. SD [SLHC12]. Seam [WPW^{+}11]. Seamless [BNH10, GWO^{+}10, LLCZ16, MC19, PBMG15, PJR^{+}14, WHLW19]. Seams [RNLL10]. Search [BSK^{+}17, LZ10, SML15, WESW17, YWHB18]. Searching [RI17, WCB15]. Seasonal [HJM^{+}11]. Seated [HMP^{+}12]. Second [GRT19, HL13, KASH13, LPG13, MRL10, ORT18, SK16a]. Second-Order [GRT19, KASH13, LPG13, ORT18, SK16a, MRL10]. Section [AJA11, DPW11, DS11b, FWPS11, GSZ11, HP11, KP11, Lav11, MVPG11, Mér11, NBCW^{+}11, NRP11, OHG11, PB11, Rus11, SY11, SBCBG11b, TOZ^{+}11, VP11, WBCGH11, WSSC11, dGCSAD11, vKvLV11]. Sections [KBT^{+}12, SSLL14, DMNV12, HHCJ18]. Security [MSFM16, SMvdWvW15]. Seeding [ELM^{+}12]. Seeds [HB19]. Seen [GTK^{+}12]. Segment [SMJ17]. Segmentation [BLVD11, BPVR11, BIZ18, DSWH17, FMH16, HMTH13, HPH10, HFL12, IY10, IYS^{+}13, Jac17, JWL^{+}13, KWÖG18, KJT14, LT17, LMW^{+}19, LWL^{+}16b, LCHB12, NW17, NSS^{+}12, RBC14, RMG18, SMH10, SHJ^{+}16, SHS13, STP17, SFLP18, TWS^{+}11, VGB14a, XXLX14, XSX^{+}14, YC19, YL11, YLX^{+}16, ZWC^{+}10, ZT10, ZVE^{+}14, DFIM15, VF14, ZZCJ14]. Segmented [JJKL18, NW17]. Segmentifier [DRM19]. Segments [LLW12]. Seidel [Cet19]. Select [SV10, SLC^{+}19]. Selecting [LAE^{+}12, LFK^{+}13, WGK^{+}19]. Selection [BSY^{+}19, BvLBS11, ESRT13, FE17, HREB11, hKTL^{+}17, SPSK13, TWC^{+}16, ZC18, GTB14]. Selections [MGB14]. Selective [CAS^{+}19, HDBRC17, HBRD^{+}18, RKRD12, SS15a]. Self [AAB^{+}10, ACOHS18, BBAM12, CK10b, CZY11, DCV14, KNL^{+}15, QWZ^{+}19, SJP^{+}13, SWS12, TW10, WLT^{+}17, WC14, YLD^{+}18, LN18, RRS12]. Self- [CK10b]. Self-Adaptive [CZY11]. Self-augmented [YLD^{+}18]. Self-Collision [WLT^{+}17, WC14]. Self-Enhancing [QWZ^{+}19]. Self-Intersecting [SJP^{+}13]. self-intersection-free [RRS12]. Self-Organizing [AAB^{+}10, SWS12]. Self-Shadowing [TW10, LN18]. Self-similarity [ACOHS18, DCV14]. Selfie [LWL^{+}16b]. Semantic [BDF^{+}14, HMW^{+}15, HZL19, JJKL18, KWÖG18, PTT^{+}12, SML15, WPW^{+}11, ZZZ^{+}19]. Semantic-aware [HZL19]. Semantic-Preserving [WPW^{+}11]. Semantically [JBMC10, JJKL18, LFK^{+}13]. Semantically-Resonant [LFK^{+}13]. Semantically-Rich [JBMC10]. Semantizing [BHMT13]. Semi [ABCJ10, AYLM13, AW13, ER18, GSE^{+}14a, KvLB14, KSD14b, LLN^{+}14, LCHB12, MKP^{+}16, MC14, MEKM17, NBMJ14, PSPM12, PRS15, RLMB^{+}14, SD10a, SKZF11]. Semi-Analytic [NBMJ14]. Semi-Automated [LLN^{+}14]. Semi-Automatic [GSE^{+}14a, KvLB14, MEKM17, MKP^{+}16]. Semi-Implicit [SKZF11]. Semi-isometric [ABCJ10]. Semi-Lagrangian [AW13]. Semi-Regular [AYLM13, PRS15]. Semi-sharp [KSD14b]. Semi-Stochastic [SD10a]. Semi-Supervised [PSPM12, LCHB12]. Semi-Transparent [ER18, MC14, RLMB^{+}14]. Semiautomatic [ZH14]. Semidefinite [HG13]. Senescence [KRB11]. Sensing [KÓOH13, WBAI19, XSE14]. Sensitive [BSAP11, SBS^{+}17, GGM12]. Sensitivity [BMPM12, DW13, NWHWfD16]. Sensor [BCN11a, CD18, SBM^{+}14, WZ15, CACB18]. Sensor-aware [CACB18]. Sensor-based [CD18]. Sensorimotor [HL18]. Sentiment [KPK18]. Separable [JZJ^{+}15]. separate [CH12]. Separated [BMB15b]. Separation [GOPT11, GT15, STMT12, WLN^{+}17]. Separatrix [GSW12]. Sequence [BHW17, CFB16, CWM19, YLL15]. Sequences [AECOK16, CKK18, DH16, GRPF16, HKL17, KR19, KCJM16, LVV18, LKD^{+}17, PCDS12, PCX^{+}18, SCD^{+}16, SWG16, YLX^{+}16, CH12]. Series [AKMM11, ARH12, BHR^{+}19, BBL12, DH16, GRPF16, HJM^{+}11, KCJM16, MLD^{+}18, SAAF18, SWG16, SBM^{+}14, TFA^{+}11, WG11, vDvdEvW19, KWS^{+}15, LRB^{+}15, RL15]. Serious [MTVJ11]. Session [AW13, AO13, BCD^{+}13, ČHM^{+}13, CTL13, CNKI13, CCI13, EWMU13, FWX^{+}13, GDML13, GRC13, HSmCY13, JCK^{+}13, JK13, JWL^{+}13, KMHG13, LCCC13, LJBA13, LLC13, LPG13, LLSC13, MH13, RMZ13, SRK13, SYT^{+}13, SHS13, SKWL13, TTN^{+}13, TSdSK13, WZH13, WCX^{+}13, WWL^{+}13, WLM13, XWG^{+}13, XXZC13, XZP^{+}13, XSQ13, XADR13, YH13, ZYF13, ZCZL13, ZLKW13, ZCF^{+}13, ZWY^{+}13, ZIM13]. Set [AMA^{+}16, DvKSW12, HCDC18, JK13, Kim15, KTW^{+}13, LA13, LZW^{+}13, LDY10, MWTI19, SDKG18, TSB16, XDC^{+}13, WLS13, vKvLV11]. Set-based [KTW^{+}13]. SetCoLa [HBH18]. Sets [BDF^{+}14, BSW^{+}14, BM12, BTB13, CG12, DGQ^{+}12, DMSL11, GCSA13, KS11, KS12a, KWS16, MBR^{+}13, NSM19, RGM^{+}18, RL16, SBG17, SvW13, SEG^{+}14, Yuk15, GRT14]. Setup [ARG^{+}16]. Seurat [WTLL13]. SfM [WKM15]. shade.js [SKSS14]. Shaded [Jes16]. Shader [CTHAM10, MRD12, RPLH11, YB18, YWHB18]. Shaders [DBLW15]. Shading [ABG^{+}12, BBDA10, DZCC19, ENSB13, FC10, GRDE10, HSS19, KB12, LTKD15, NAM^{+}17, SBF15, ŠPBV10, TH17, WJB^{+}13, ZM16, AHTAM14]. Shading-based [DZCC19]. Shadow [AHMAM15, BBH13, BKB^{+}12, DVP19, FWX^{+}13, GMAG15, KSCN19, LSMD15, MWS^{+}16, MSWI12, MIW13, NM14, OPP10, REH^{+}11, SVLL10, SBE16a, SWP11, SDMS15, SGYF11, SFY13, TIK17, WLZX19, XLH^{+}13, XSXM13, XXZC13, YDF^{+}10, ZZLX17, BBAM12]. Shadowing [BKB^{+}12, BBAM12, HBA12, TW10, YIC^{+}12, ZCBK12, LN18]. Shadows [BBH13, BNJ15, KPD10, LSMD15, MAAG12, NKLN10, NPW10, OPP10, SVLL10, SBE16b, SDMS15, SFY13, SARZL10, Tim13, WZKP14, YHGT10]. Shah [BCGL18]. Shake [WYD^{+}13]. Shallow [PHTB12]. Shape [AWO^{+}10, AYLM13, AGCO13, ATCO^{+}10, AKZM14, AKM16, ADG19, BLP10, BSK^{+}13, BDC18, BWM^{+}11, BD12, BCGS13, BCBB16, BMM^{+}15, BEKB15, BDS^{+}12, BHMT13, BvTH16, BBP10, COO15, CRA^{+}17, DDÖ^{+}17, DBD^{+}13, DMS14, EDPB15, ELC19, ESKBC17, EBC17, FHE19, FSTR13, FCS^{+}16, GSDG18, GLHH13, GCLX17, GL12, GCSA13, GAWJ15, GKOM18, HCW17, HBA12, HPH10, HFL12, HSvK18, HG13, HKM15, HO17, HCO18, HAGO19, JWS12, JTRS12, KWS16, KP18, KTKM19, KSO10, KKSS15, Kim15, KFLCO13, KO19, KBB^{+}13, KNH^{+}18, KRS^{+}13, KSKL13, LA13, LT17, LGH13, LMP13, LKEP14, LJK^{+}12, LJL^{+}18, LRBB17, LPG10, LVW^{+}15, LGZ^{+}16, LJC17, LCHB12, MCH13, MGG10b, MBG^{+}12, MSAP15, MGB^{+}12, MRCB18, MBT^{+}12, MWCS13, MEKM17, NSC14, NBCW^{+}11, NNRS15, NC16, NO17, OLA16, OM13, OHG11, OMPG13, PB11, Pat16, PPT^{+}19, PBB^{+}13]. Shape [PSO18, RLGH15, RBRY19, RÖM^{+}15, RKN10, SY11, SY12a, SY13, SY14a, SY14b, SXY^{+}11, SBC14, SGB13, ŠBM^{+}10, SJW^{+}11, SJWS13, TBW^{+}11, TBTB12, WXL^{+}11, WK12a, WLL^{+}17, WG11, WMZ12, WSSC11, WDAH10, WYY13, WZCF15, XXLX14, XXS^{+}15, XSX^{+}14, XM15, XKHK17, XCDR10, YFW12, YL11, YLH^{+}14, ZWC^{+}10, ZDZ^{+}15, ZLW^{+}19, ZST^{+}10, ZT10, ZFCO^{+}11, ZCOM13, ZCOAM14, ZHH^{+}15, ZXTD10, ZLW15, vKTS^{+}11, vKZHCO11, CH12, DFIM15, GFK^{+}19, LBBC14, NW17, ZKWG16, vKvLV11, LGZ^{+}16]. Shape-appearance [NNRS15]. Shape-Aware [JWS12, LGH13]. Shape-from-Operator [BEKB15]. Shape-Up [BDS^{+}12]. Shaped [KS12b, NS11]. ShapeGenetics [HSS17]. Shapes [ARB^{+}18, Att15, BPVR11, BS12b, BEKB15, CJFH14, COO15, CD10, EBC17, GSTOG16, HCGW14, HMW^{+}15, HFL12, HAGO19, JLW10, KLMK19, KFLCO13, LCSL18, LK19, LZW^{+}13, LGK16, LMHH14, LMPS16, MCH13, OBCCG13, RXX^{+}17, RBC14, SY14b, SGS14, STC^{+}16, SCF10, SDKG18, TTB12, WWH^{+}10, XXS^{+}15, YFW12, YWM15, YLLL15, ZHH^{+}15, ZZCJ14, dGCSAD11, BMM^{+}15, KP11, MWCS13, TWMSK18]. ShapeSynth [AKZM14]. Shaping [BDS^{+}12, MRL^{+}17]. Shared [GO10, HMW^{+}15, Ros13]. Sharing [CCC^{+}14]. Sharp [AMS16, BW13, BM12, CJXH17, KSD14b]. Sharpening [TM13]. Sheet [Tak19]. Shell [LZY^{+}17, ZLW^{+}16, ZXZ^{+}17, YCXW17]. Shells [AWO^{+}14, DH14, FB11, HRWW12, HRS^{+}14, HRS^{+}16, HZRS18, KSD14b, PTP^{+}15, UMK19]. shift [XL10]. Shifted [BSH12]. Shifting [ST18]. Shiny [BCRA12]. Shock [PK15]. Shortest [BLP10]. Shot [KBHM15, CKL14, MSZ^{+}18]. Shoulder [HPH10]. Shutter [CKL14]. ShutterApp [SBE19a]. SI [WDM^{+}12]. Sickness [HL18]. Sided [SS15a, TPSH14a, VSK16]. Sifted [EMA^{+}13]. SIGGRAPH [CSLG10, LSF^{+}11, PS10, DS11a, ID10, IC11]. sight [AGCO13]. Sightlines [BAO^{+}19]. signal [SD10b]. Signals [GO15, CH12]. Signature [DLL^{+}10]. Signatures [COO15]. Signed [CT11, CSŠ18, MRL10, SFFP15]. Significance [BRDC12]. Signing [MDD^{+}10]. Signposts [MSDK12]. Silhouette [CSD11]. Silhouette-Aware [CSD11]. Silhouettes [NBMJ14, STC^{+}16]. Similar [CG16b, WLL^{+}17]. Similarities [HCDC18, SBM^{+}14]. Similarity [ARRO^{+}17, BCBB16, BM10, CG16c, FL19, GAWJ15, GLGW12, GTB14, PB11, PR12, SW17, TWC^{+}16, ACOHS18, DCV14]. Similarity-based [GTB14]. SimilarityExplorer [PDW^{+}14]. Simple [EMP^{+}12, KZB19, PM16, Ren16, VSK16, LCLJ10]. SimpleFlow [TBKP12]. Simpler [NB12]. Simplex [MRS18, WD11]. Simplicial [LS16, MS10b, RL14, YS19]. Simplification [CLS16, CHWL19, CP10, DC10, JWH19a, KWS16, KRM^{+}15, MG10, RGG15, SSLL14, YWHB18, dGCSAD11, TPC^{+}10]. Simplified [TE10]. Simplifying [MMG10]. SimSelect [GTB14]. Simulated [BAAR13, EWK^{+}13, GP12, LE13]. Simulating [MRD12, RKN10, ZCT18]. Simulation [AO13, BSK16, BS19, BET14, BMO^{+}14, Cet19, CC14, CZY11, CKHL11, DCGG11, EWK^{+}13, FAW^{+}16, FM15, FMD^{+}19, FHW^{+}11, GBW16, HBLB17, HL18, HLL^{+}19, HJS^{+}17, IUDN10, JPK13, JL18, KPNS10, KRB11, KS18, KBB^{+}17, LeYO^{+}10, LWPL15, LMG^{+}18a, LL18, LCY^{+}19, LCM^{+}18, MR17, MGG^{+}10a, MKMA19, MMG18, MBT^{+}12, MSGT18, MGN17, MPBM^{+}17, MESG11, NGHJ18, OLF^{+}14, PPG^{+}19, PBK10, RHLH18, RKN12, SCN^{+}16, SJ13a, SWML10, SHCB15, SKK10, SM14b, STBG12, SKWL13, SDHL11, TTN^{+}13, TWT^{+}16, WJDZ14, WWF^{+}18, WMRSF15, WHT12, WHP^{+}11, WLZ13, WAF^{+}11, WYY13, WWD15, WCH^{+}15, XZP^{+}13, YLHQ14, YMJ^{+}17, YBK^{+}12, YWTY12, YPL19, ZTW^{+}12, ZLM^{+}15, ZLYL17, ZZ17, ZLKW13, ZYF10, dHvPJV14, BHT19, DMCN^{+}17, LJK^{+}12, YLCH18]. Simulation-Ready [HBLB17]. Simulations [ATW15, BTM^{+}19, CFS14, CBKK^{+}19, FKE13, FAW^{+}16, JKJL18, KAT^{+}19, RGG15, SWT^{+}18, SHQL18, WTYH18, WGO^{+}14]. Simulator [KTN10, SSB13]. Simultaneous [ESKT15, ZBW11]. Single [ARM^{+}15, BCN11a, CKL14, DDÖ^{+}17, DSY10, EHT18, EIKM16, ECN14, HFM16, HPvU^{+}16, Hol15, IEH^{+}14, IEK^{+}14, IEKM16, IRWM17, KTKM19, KrJC^{+}11, LVV18, LQS^{+}19, LSZ^{+}18, LCD10, LMGH^{+}13, MWS^{+}16, MES^{+}11, NBA18, OAO11, PSP10, SM11, SPT14, YXHH18, YLD^{+}18, YLH^{+}14, ZCW^{+}15, ZCF^{+}13, YGCO^{+}14, HRRR18]. Single-Cell [HPvU^{+}16]. Single-Image [LVV18, HRRR18]. Single-Pass [LCD10]. Single-shot [CKL14]. Single-View [KTKM19]. Singles [vdEvW13]. Singular [DGQ^{+}12, KSBC12]. Sit [LBRM18]. Site [LPG19]. Situ [BAA^{+}16, SEI10, AAS^{+}16, WAF^{+}11]. Size [KGR^{+}16, MW18b, She12, YYZZ18]. Sizes [XGDC17]. Skating [YPL19]. Skeletal [LLSC13]. Skeletex [MRM^{+}18]. Skeleton [KMS^{+}13, KFA^{+}14, LW18, LMPS16, MRM^{+}18, MVPG11, NB15, TE18, WMXC19, ZST^{+}10, ZHH^{+}15, Kur15]. Skeleton-based [KMS^{+}13]. Skeleton-Driven [KFA^{+}14, LMPS16]. Skeleton-guided [WMXC19]. Skeleton-Intrinsic [ZHH^{+}15]. Skeleton-texture [MRM^{+}18]. Skeletons [LW17, TAOZ12, TDS^{+}16, TBTB12]. Sketch [FCS^{+}16, HMTH13, LCC^{+}18, LG13, LWX^{+}15, NSS^{+}12, SXY^{+}11, SAG^{+}13, TCGK15, WLL^{+}17, YXHH18, ZLDM16, CSLG10, XXM^{+}13]. Sketch-Based [HMTH13, SAG^{+}13, WLL^{+}17, LWX^{+}15, SXY^{+}11, CSLG10]. Sketch-Driven [TCGK15]. Sketch-to-Design [XXM^{+}13]. Sketches [ADN^{+}17, BSCH18, GLX^{+}16, ZLDM16]. Sketching [BLP10, KNH^{+}18, LG13, SLE17, Sch13, SAG^{+}13, WHC15]. SketchSoup [ADN^{+}17]. Sketchy [KÖS^{+}15]. Skies [HCBW19]. Skiing [HLL^{+}19]. Skill [HGB^{+}10, HLL^{+}19]. Skill-Guided [HLL^{+}19]. Skilled [FvdP15]. Skills [GCY^{+}14]. Skin [BW17, IGAJG15]. Skinned [JKL18, TPSH14b]. Skinning [KSO10, LCY^{+}19, RF15, TE18]. Skins [IGAJG15, MMG10, MKU15, Yuk15]. Skip [ZLZ^{+}18]. Sky [GPRS14, HCBW19]. Sleep [CWGvW19]. Slice [WLZX19, ZH14, SV10]. Slice-guided [ZH14]. Sliceplorer [TWSM17]. Slices [HBA12, TWSM17]. Slicing [WCT^{+}15]. Sliding [HBA12]. Slopes [RBMS17]. SMAA [JESG12]. Smale [SN12]. Small [BHRD^{+}15, BBL12, DHS^{+}13, LHNS18, PGGM10, SJ13a, UMM^{+}10, vdEvW13, BRM^{+}16a]. Small-Multiples [BBL12]. Smallest [LK13]. SmallWorlds [GOB^{+}10]. Smart [MAM14, NSS^{+}12, ZCOM13, ZLDM16]. SmartAnnotator [WCM15]. Smartphone [ECN14]. Smoke [AO13, BXH10, KPNS10, NC10, PSCN10, SKK10, WYY13]. Smooth [AECOK16, CT11, JWS12, Jes16, MSWI12, SLWSS15, CFGL16, SBD^{+}15b]. Smooth-Shaded [Jes16]. Smoothed [HENfSYS16, HA19]. Smoothing [FWSH19, LTB19, LWS^{+}16, MLK^{+}13, WTL15, WGS10, YB18, ZFJ^{+}16]. Snakes [BSK^{+}13]. Snapping [SAG^{+}13, ZWC^{+}10]. Snapshots [MGB14]. SNE [KRM^{+}17]. Snooker [HGB^{+}10]. Snow [CEG^{+}18, vFG11]. Snow-Covered [CEG^{+}18]. Soccer [SAMS^{+}17]. Social [BCD^{+}10, CLY17, GOB^{+}10, HRD^{+}15, JGH11, KvLB14, KHI^{+}19, RTJ^{+}11, RPA^{+}15, TX16, WDM^{+}12, OKK13]. Soft [HK16, KKBJ16, KARC17, LSMD15, MP12b, MMO16, MAAG12, RF15, SDMS15, SGYF11, SFY13, SNB^{+}12, SGB13, XLH^{+}13, YDF^{+}10, ZIM13]. Software [AHR13]. Solar [ABW^{+}15, MPBM^{+}17]. Solid [AMS16, Gam16, SCM^{+}19, SZMTW15, TL16, TL19, WW11, WC14]. Solids [BHU10b, CZZ^{+}18, PGBT18, ZLKW13, YLCH18]. Solution [ATO17, GUS12, MSGT18, PSP10, RS19]. Solve [WBS^{+}13]. Solver [Aan18, ATW15, DKWB18, KH19, TL16, TDNL18, TL19, WMRSF15, WKBB18]. Solvers [Kaz15]. Sorted [ENSB13]. Sorting [GL10a]. SoS [ML17]. Sound [JBMC10, STG16]. SoundRiver [JBMC10]. Source [GT16a, LMGH^{+}13]. sourced [KWD14, XLS^{+}14]. Space [AJA11, AVR10, AMTMH12, AKS^{+}19, AAB^{+}10, BDA^{+}17, BvTH16, BSL18, BKW13, BMS^{+}10, CDA^{+}14, CCI13, DMSL11, EBGM12, EAGA^{+}16, GKL17, HKD15, HRWW12, HRS^{+}14, HRS^{+}16, HZRS18, HB19, HMRR19, HFE13, JH19, KD13, KNH^{+}18, KW18, LS10b, LMP^{+}10, LLB^{+}10, LJLH19, MCHW18, MO10, MBES16, MEMO14, MFNP13, Man16, MHA17, MSW10, MGB^{+}12, MRS12, NAM^{+}17, NSM19, NPW10, PSDB^{+}10, RGG18, SSSSW13, SB13, SGEM16, TFA^{+}11, WG11, WWT^{+}16, WBT19, WTH^{+}13, XWB15, YLX^{+}16, YWHB18, YIC^{+}11, ZSW10a, ZFAQ13, ZLW^{+}19, ZST^{+}10, ZZ19, ZLW15, AFHdL14, AHTAM14, FMD^{+}19, JPCC14, Kur15, ZRJ^{+}15]. Space-Efficient [CDA^{+}14]. Space-in-Time [AAB^{+}10]. Space-Time [MBES16]. Spaced [EDPB15]. Spaces [BPFG11, BvLBS11, BBP10, CSG^{+}18, ERHH11, GSDG18, GHWG14, HW10, HMW^{+}15, MCH13, MFL13, OMPG13, PBK10, SBC14, SDKG18, TFA^{+}11, VAW^{+}10, vBE11]. Spacetime [CTL13, TCLK12]. Sparse [BLD14a, BTP13, CCM16, CWL19, CRA^{+}17, GDGP16, HR10, HYZ^{+}14, LJH10, LMMCO17, LDY10, MSAP15, MRAS17, OLF^{+}14, PBB^{+}13, RXX^{+}17, RSK12, RK10, SHD16, SRK13, SD10b, SHG^{+}16, SGMG17, SHQL18, SBC^{+}17, SCF10, SSJ^{+}10, WLZH17, WBS^{+}13, WDR11, WTYH18, XSX^{+}14, YLLL15, ZFJ^{+}16, vMRBPM17, GGM12, LBBC14]. Sparsely [HR10]. Sparsity [DSH^{+}17]. Spatial [BCCS12, BSY^{+}19, EGG^{+}15, FKRW16, FL19, HHNČ19, JKL18, MR17, MJK11, MGB14, MFL13, PPH12, SC16, SR19, WCH^{+}15, GD16]. Spatializing [JE13]. Spatially [KKTD17, LHNS18, MHRK19, PR12, SYC10, WTH^{+}13, SPCR14]. Spatially-aware [KKTD17]. Spatio [BJG^{+}15, CLC12, DPD^{+}15, KBÖ^{+}14, SBE19a, ZZLX17, Fre18]. Spatio-Temporal [CLC12, DPD^{+}15, KBÖ^{+}14, SBE19a, ZZLX17, Fre18]. Spatio-Temporo-Chemical [BJG^{+}15]. Spatiotemporal [AAB^{+}10, BMH^{+}12, BBBL11, DBS^{+}18, DBS^{+}11, LBK14, RSD^{+}12, RWSG13, WHP^{+}11, vPJtHRV12]. SPC [SO10]. Special [GCGP18]. Specialized [GHB15]. Specific [BBGB19, Deb18, GRP10, hKTL^{+}17, LMK^{+}15, BMM^{+}15, WGK^{+}19]. Specification [GOH^{+}10, SE19, MVLS14, WZC^{+}11]. Spectra [AAS17, BLTD17, LJC17, RK10]. Spectral [ACKM16, AFK^{+}14, BSEH17, BLTD17, CMF18, EBR^{+}14, HKW12, HCGW14, JH19, JWH^{+}19b, LMLG15, LRBB17, MRCB18, MFW18, OYH18, Pat16, Pat17, PBE18, QPCRM19, RDK13, Ste19, SDHL11, WND^{+}14, ZVD10, BMM^{+}15]. Spectralization [RK10]. Spectroscopy [JCK16, MK11, SPB^{+}17]. Specular [FZS^{+}19, IKL^{+}10, Tok15a]. Specularity [ABC11]. Speed [LTB19, SGMG17, WLI^{+}12]. SPH [FM15, HLL^{+}12, IABT11, KS14, OK12, PGBT18, SZMTW15, TDF^{+}15, TL16, TDNL18, WKBB18, ZYF10]. SPH-based [TDF^{+}15]. Sphere [DKY16]. Spherical [EJFadH13, GUK^{+}17, IFDN12, Man16, MBR^{+}13, NBMJ14, Tok15b, UFK13, UG18, XLH^{+}13, YZXW12]. Spherically [KSKL13]. Spherically-Parameterized [KSKL13]. Spine [ZR13]. Spine-driven [ZR13]. Spiral [HT11]. Splash [UHT18]. Splattering [LCCC13]. Splatting [FSES14, HBW11, HWK^{+}10]. Splicing [vBE11]. Spline [BSH15, Csé18, DH14, RC18]. Splines [BLP10, BSH15, CR16b, HRS^{+}16, HQH15, PTP^{+}15, SS15a, KP11, KP19]. Split [KGRG17, LEE17, GD16]. Split-Depth [LEE17]. Splitting [KKB18]. SPLOMs [YWS^{+}14]. Sponsored [CSLG10, IC11, ID10]. Sport [GPK^{+}12]. Sports [KÓOH13, LCP^{+}12, PVS^{+}18]. Spray [YLHQ14]. Springs [WPGSH18]. SQuad [GLLR11]. Squared [PPL13]. Squares [BGS10, KBS11b, MS10a, MGB^{+}12, PSPM12, RAV^{+}19, KBÖ^{+}14]. Squeezing [CKHL11]. SSD [CT11]. Stability [GBG^{+}14, HCO18, STP17]. Stabilization [BAAR14, WZH13, XHW^{+}18, ZLSW17]. Stable [COO15, GSTOG16, PK15, SZMTW15, Szy11, MSW19]. Stack [HSBW13]. Stacking [HL19, MP12b]. Stackless [ÁSK14]. Stacks [GTM^{+}12]. Stage [SPF^{+}19]. Staggered [WLI^{+}12]. Staging [CWS^{+}17]. Stamps [MGC^{+}16]. Stand [KFA^{+}10]. Stand-Ins [KFA^{+}10]. Standard [RLP10]. Standardized [MMNG17]. Star [WLN^{+}17, YL11, HS17, Pat16]. Stardust [RLH17a]. Starting [Lar10]. State [AHG^{+}19, AMA^{+}16, BBT11, BFR17, BHP15, BCMP18, BCD^{+}12, DFIM15, ERT^{+}17, FW17, GP12, GHWG14, GGG^{+}16a, GT18, HSK^{+}10, KRP^{+}15, KKF^{+}17, KBG^{+}15, KKL^{+}16b, KPK18, LHT17, LKG^{+}16, LGH^{+}17, MGM^{+}19, MKRE16, NSG11, NMSL19, NK16, Pat16, PP10, PVS^{+}18, PDC^{+}19, PPF^{+}11, RDGK12, RGG^{+}14, SRM^{+}19, SPN^{+}16, TDS^{+}16, TAEE15, VCD^{+}16, WWS^{+}15, ZTG^{+}18, ZSG^{+}18, ZJL^{+}15, vLKS^{+}11]. State-of-the-Art [AMA^{+}16, BBT11, BHP15, FW17, GP12, LGH^{+}17, MKRE16, RGG^{+}14, SRM^{+}19, TDS^{+}16, vLKS^{+}11, PDC^{+}19]. Static [CKE^{+}12, CLL^{+}13, ODS18, PBMG15, SBD^{+}15b]. Statics [PTP^{+}15]. Stationarization [MJH^{+}17]. Stationary [ZXW19]. Statistical [CRY11, JMD15, NVH^{+}13, PCDS12, WLJ^{+}18]. Statistics [PKRJ10, PCR11, RSM^{+}16, RW18, VVP^{+}16, YB18]. Statutory [Ano10c, Ano11e, Ano12f, Ano13i]. STD [RBMS17]. Steady [GT16b, GG17, TSW^{+}19]. Steering [AGR19, LCM^{+}18, DMCN^{+}17]. Stenciling [STG16]. Step [KB12, WPY^{+}19]. Stepping [FHHJ18]. Steps [ZC18]. Stepwise [MWTI19]. Stereo [BSL18, FNH^{+}17, LFYX19, SW11, SSB^{+}14, CHA^{+}14, PHM^{+}14]. Stereo-Consistent [BSL18]. stereo-to-multiview [CHA^{+}14]. Stereoscopic [KrJC^{+}11, NREM14, TDMS14, ZZH15]. Stereoscopy [ZCW^{+}15]. Stick [CYI^{+}12]. Sticks [YIC^{+}12]. Stiff [DKWB18]. Stiffness [ATK17]. Still [SKWL13]. Still-Frame [SKWL13]. Stippling [MGY^{+}18]. Stitching [GWO^{+}10, GTZM10, LCC^{+}17, LWL^{+}16b, LLCZ16, LCWCO10]. Stochastic [AMAM13, AHMAM15, AAS17, CWY11, DJSJ19, GD10, GG14, KCL^{+}18, LK11, LSMD15, MTAM12, NDG17, PHL^{+}16, SD10a, TH17, WWH18, WG12, AHTAM14, SNJ^{+}14]. Stone [XDR11]. Stones [PGGM10]. Stories [RGM^{+}18]. Storm [LMK^{+}15]. Story [GLG^{+}16, MRL^{+}17, RGM^{+}18, SLSG16]. Storyboard [ZC18]. Storyboards [HMLP13]. Storytelling [GLG^{+}16, ZOM19]. Straightforward [BCCS12]. Strain [MYLZ16, OBH^{+}11, SBO18]. Strange [YMM10]. Strategies [GB10, HKL17, SNJ^{+}14]. Strategy [AMR^{+}17, BP17, dSNV^{+}17]. Stratified [LK11, UG18]. StratomeX [LSS^{+}12]. Stream [ELM^{+}12, ESRT13, HLJ^{+}13, MH13, SRWS10, SGRT12, SEG^{+}14, WESW17]. Streamgraphs [DH16]. Streaming [DAF^{+}18, HSS19, PHE^{+}11, SRK13, SBS^{+}17, TTN^{+}13, VOS^{+}10]. Streaming-Enabled [VOS^{+}10]. Streamlines [LMP13, MC14]. Streams [KCJM16, RWW16, GBM^{+}19]. strengthening [ZKWG16]. Stress [KHKS12, MVB^{+}17, ZXZ^{+}17]. Stress-Constrained [ZXZ^{+}17]. String [BRWM18, KLMK19]. String-Based [KLMK19]. Stroke [BBK^{+}19, LCY^{+}11, LLSC13, Sch13, SRG16, YYZZ18, YSC^{+}18]. Stroke-based [BBK^{+}19]. Stroke-guided [LLSC13]. Strokes [KS10, LLC19]. Structural [ERHH11, RL14, RJT18, SSW14a, SMS^{+}17, UMK19]. Structurally [STG16]. Structurally-Sound [STG16]. Structure [ASB^{+}17, AP10b, BLK11, CCM16, DCNP14, FVHK17, FO12, FCS^{+}16, HVH^{+}16, JLKL16, KO19, LJKL17, LZY^{+}17, LWS^{+}16, LWBP14, LMHH14, LLSC13, MLD^{+}18, PSC10, RBC14, RÖM^{+}15, SM14a, SLA15, SXLS19, TLRB18, VGB14a, VSK16, XXLX14, XSX^{+}14, YFW12, ZSS17, ZFCO^{+}11, ZFJ^{+}16, AVBC18, EKB14, MCM^{+}12]. Structure-adaptive [FCS^{+}16]. Structure-Aware [LWS^{+}16, SLA15, DCNP14]. Structure-based [KO19]. Structure-Driven [LWBP14]. Structure-Performance [ASB^{+}17]. Structure-Preserving [JLKL16, ZFCO^{+}11, ZFJ^{+}16, SXLS19, AVBC18, EKB14]. Structure-Texture [LJKL17]. Structured [CDM^{+}17, KLMK19, MDBS14, RSM^{+}16, RPWO19, RS19, MWCS13, ZLL13].

Structures [ABCJ10, ACAA^{+}19, BGCP11, Duc14, ER18, FR11, GOPT11, GSW12, JWWP14, KLTZ16, KMA^{+}19, KKF^{+}17, KK11a, KBT^{+}12, KER^{+}14, LS10b, LMP^{+}10, LGK16, LBH12a, MBES16, MDI^{+}18, NHL16, PLL11, PW12, RTJ^{+}11, RL16, SW10, SAD^{+}16, SPK10, TDDD18, TPRH11, VBW17, WBS^{+}13, XDY18, ZSW10a, ZLMM16, ZCBK12, ZAD15]. Structuring [HKL17, LA13]. Student [RBMS17]. Studies [DKMT18, ZLDM16]. Study [ARLC^{+}13, APM^{+}11, BBL12, CEX^{+}18, CWM19, CCH^{+}14, GKLS19, GBG^{+}14, KARC15, KGP^{+}12, NW13, PRS15, RCMA^{+}18, SH14a, SH14b, SBLC17, SS15b, TLM16, UMM^{+}10, WHC15, WBFvL17, YPF19]. Style [AMYB17, BV19, GTL^{+}18, HZL19, JL18, LZW^{+}13, LPD14, LGK16, LCUR14, LTR19, NRM^{+}12, WLL^{+}17, YYZZ18, MHS^{+}14, RLYL14]. StyleBlit [SJT^{+}19]. Styles [MSZ^{+}18, XADR13]. Stylization [BZBM^{+}16, BLV^{+}10, CS16, FLJ^{+}14, HHGJ15, KLC^{+}15, LWB14, LM15, LCUR14, RLMB^{+}14, SHJ^{+}16, SJT^{+}19]. Stylized [BBT11, BCMP18, GRR^{+}16, LG19, PMG13, VVC^{+}11, YKM12]. Sub [DFY14, JKL13, UBH14]. Sub-Grid [UBH14]. Sub-Pixel [DFY14]. Sub-Sample [JKL13]. Subdivision [BPY16, BGS10, BHU10a, Cas12, DH14, DZCC19, DKY16, GZH^{+}19, HLS12, KP18, KSD14a, KSD14b, MM18, MM19, MRAS17, SVBC19, VK18, BHU10b]. Subdivision-based [DZCC19]. Subgrid [KER^{+}14]. Subject [GRP10]. Subject-Specific [GRP10]. Subjective [DBS^{+}18, KWD14, KDC17, MTM12]. Sublinear [TBKP12]. Subneighborhoods [ZWY^{+}13]. Subpixel [ESKD14, JESG12]. Subregion [XXZC13]. Subsampling [BCCS12]. Subspace [BWM^{+}11, FR11, HCW19, HFL12, LWT^{+}15, MWW16, SJF11, ZHQH17, ZZT15]. Subspaces [MKU15, NNRS15]. Substructures [LVW^{+}15, ZCOM13]. Subsurface [HCJ13, JZJ^{+}15, MESG11]. Subtype [LSS^{+}12]. Succinct [JYS19]. Suggested [SRH^{+}11]. Suggestion [KM16]. Suggestions [RLGH15]. Suggestive [LWX^{+}15]. Summarization [XBL^{+}18]. Summarized [YKS^{+}19]. Summarizing [ODS18]. Summary [KS13b, PKRJ10, SGS18]. Sums [CK10a]. Super [ACA18, BD12, DTV15, HCDC18, MAM14, SLHC12, ZLZ^{+}18]. Super-Clothoids [BD12]. Super-Deformed [SLHC12]. Super-Resolution [HCDC18, ZLZ^{+}18, ACA18, DTV15]. Superfacets [SPD14]. Superimposed [LMLG15]. Supernova [ASW14]. Superpixel [CG16a, DCYG19]. Supervised [LBBC14, PSPM12, LCHB12]. Supervision [YLD^{+}18]. Support [BSK^{+}17, BMPM12, CBK^{+}17, CWS^{+}17, GKLS19, KFW19, KMJE12, LCP^{+}12, PZDD19, RLH17a, RWS^{+}10, VGB14a, VMTS10, WKS^{+}14]. Support-Free [KFW19]. Supported [HA17, RGSK10]. Supporting [BEF17, CE19, KFR18, LBRM18, LCB^{+}18, MbMYR15, SV10]. Surface [AJA11, AIAT12, ABG^{+}12, BSK^{+}13, BBB^{+}18, BTS^{+}17a, BLK11, CT11, CIE^{+}16, Cas12, CCLN10, CLCL11, CCW12, CBLW19, DRF12, DZCC19, DLY^{+}18, ELM^{+}12, GGV^{+}19, GLLR11, HCJ13, HTG14, HWK^{+}10, HLR^{+}11, JTSZ10, JK13, KL19, KRS^{+}13, KGM^{+}10, LA13, LS10b, LGP14, LKEP14, LVPI18, LMP^{+}10, LMA^{+}18, LKF12, MRM^{+}18, MSS11, MMG18, MDD^{+}10, MES^{+}11, NM14, NSRS13, PPL13, PSCN10, PDC^{+}19, PGK10, RLH^{+}18, Ren16, RPG16, RSK12, SLB^{+}18, SV14, SAMG14, SSW14b, SGRT12, SYT^{+}13, SBCBG11a, SCQ^{+}19, SMG10, TOZ^{+}11, TBTB12, VSG^{+}13, WLT12, WYZC13, WZK16, WLZT18, WDR11, XYLY19, YMM10, YFL19, YLD^{+}18, YMYK14, ZSW10a, ZYF13, ZCK17, HDV19, NGM14, PSP^{+}14, SZ18]. Surface-based [LVPI18]. Surface-Like [PSCN10]. Surfaces [AP10b, AMT^{+}12, AW13, ABCCO13, AWO^{+}14, AVBC16, BAT11, BCBSG10, BBCW10, BW13, BGS10, BvTH16, Bru19, BL18, CHK13, Cam17, CGT^{+}15, CLS16, CMF18, CBV^{+}14, CDS10, DW13, DH14, DGQ^{+}12, DKY16, DBK11, EKFM12, ESRT13, GLK16, GB10, GBW16, GSE^{+}14b, GTG17, HCDC18, HLS12, HRS^{+}14, JKL^{+}16, KP18, KSD14a, KSD14b, KSKL13, LBRM18, LW17, LMG^{+}18b, LPD14, LBPH10, LPG10, MJK17, MS10b, MRMH12, NGM14, OAM^{+}18, PKS10, Rus10, SJP^{+}13, SM10, SV14, SKR^{+}14, SRWS10, SHD16, SEG^{+}14, STG16, SYC10, SVBC19, SPD14, SLD^{+}17, SNB^{+}12, SVWG12, Szy11, VSK16, VWH18, WLS13, WC14, XLS^{+}14, XDY18, YWTY12, ZBQC13, dGWB^{+}14, vKvLV13, DCV14, GTB14, HP11, HKM15, KP15, KBÖ^{+}14, PCBL16, DCV14, GTB14, MHS^{+}14, VMHB14]. Surgery [MP10, TWS^{+}11]. SurgeryCuts [ABK^{+}19]. Survey [ALCS18, ABC11, BPA16, BBDW17, BNN19, BMO^{+}14, BTS^{+}17a, BKR^{+}17, BLP^{+}13, Cam17, Cas12, DCGG11, HLH^{+}16a, HGP^{+}19, IS15, JSYR14, KWD14, KZB19, KWC^{+}12, KCA^{+}16, KBvP^{+}17, KMM^{+}18, LLC^{+}10a, LSBP18, LVPI18, LOM^{+}18, ML17, MCG^{+}19, MWA^{+}13, OJMN^{+}19, PPY^{+}16, PCR11, PBC^{+}16, PKE17, RL19, RMG18, RBRY19, SWP11, SCP^{+}17, SLD^{+}17, STBB14, TGK^{+}17, TAEE15, VBW17, WT17, WWD15, YZL17, ZWJ^{+}19, vKZHCO11]. Surveys [ML17]. SVBRDF [DAD^{+}19, PCDS12, ZXW19]. Swarm [VMH^{+}13, WJDZ14]. Swarms [LWPL15]. Sweep [AMS16, Tim13]. sweeps [BPW14]. Swept [CK10a, DKY16, MCM^{+}12]. Swipe [LCG10]. Symbols [KPK10]. Symmetric [CML^{+}12, GRT19, JWH19a, JKLS10, Kaz15, KLAB15, KASH13, SY13, Sch11]. Symmetries [BWM^{+}11, GAK10, WH17b, WMXC19]. Symmetrization [ZHH^{+}15]. Symmetry [BCBSG10, DSWH17, GAK10, GL12, KBWS13, KLCF10, KLAB15, KWW^{+}14, LKF12, MGG10b, MPWC13, SFL^{+}16, SAD^{+}16, SGS14, WXL^{+}11, WMXC19, WK17, WWH^{+}14, ZYF13]. Symmetry-Aware [DSWH17, KWW^{+}14]. Symmetry-Preserving [WWH^{+}14]. Symposium [AJL^{+}11, Kuh12, LF11]. Synopsis [HLH^{+}16b]. Syntactic [PLL11]. Synthesis [AYLM13, BW17, BMM^{+}15, COO15, CLJ^{+}15, CWL19, ENMGC19, GGV^{+}19, GMM15, GPK^{+}12, GMM^{+}12, GTB^{+}13, GSDC17, HČA^{+}12, HKM15, JKL^{+}16, KLMK19, KÓOH13, KLTZ16, KKS^{+}12, KDC17, KS12b, LGH13, LL19, LSR17, LZ10, LLSC13, PHTB12, PP10, PR12, PFC15, QWZ^{+}19, RZS10, RÖM^{+}15, RSK13, SLE18, SD10a, SPK10, SGB13, TGM12, TLH19, TCGK15, VCD^{+}16, WSCP13, WYZB14, WBSH^{+}13, XWG^{+}13, YL10, ZZH15, ZFE16, ZLL13, ZSL^{+}17, AKZM14]. Synthesize [KS10]. Synthesizing [CTL13, HSK14, ZWXL17]. Synthetic [FW17, HL13, KMB^{+}17, LDG19, SKZ13, SJ13a, WOH13]. System [ADMAS18, AJC11, BCD^{+}10, BSK^{+}17, CD18, DI18, HC14, Jac19, JSH^{+}13, JLW10, KSCN19, KMJE12, LG13, LLD10, LLC^{+}15, NKSI16, NLB^{+}13, NSM19, OP10, RGSK10, SPH11, WDC^{+}10]. Systems [BWH^{+}11, CE19, CTHAM10, HBO^{+}10, HHRZ12, HHH12, OP10, PTO10, PMD12, TIK17, VOS^{+}10, WBS^{+}13, YZXW12, ŠBM^{+}10].

T [BM15, KP19, KRM^{+}17]. T-gons [KP19]. T-SAH [BM15]. t-SNE [KRM^{+}17]. Table [BBR^{+}16, SV10]. Tableau [BH19]. Tables [vdCvW16]. Tablet [SPH11]. Tablets [SS16]. Tablorer [SPH11]. Tactile [WBAI19]. Tagged [CWG11]. Tagging [KvLB14]. Talks [Ake11, Ano15a, Pos11b, Thi11]. TAMRESH [SMP13]. Tandem [OLA16]. Tangent [ABCCO13, MRMH12, WHT12]. Tangential [BSEH17, HCW17]. Tangible [BSY^{+}19, WPHC16]. Tangram [LBK14]. Task [KK17, KH18, KARC15, WGK^{+}19]. task-specific [WGK^{+}19]. Tasks [LCSS19, MbMYR15, NHG19, OAJ14, RI17]. Taxonomy [BBDW17, BNN19, BKR^{+}17, SSW14b, STMT12]. Taylor [CH11]. Teaching [McD10, SASF11]. tearing [SRH17]. Technical [Can11]. Technique [KTW^{+}13, SMM13, TCM10, WYY13]. Techniques [AKV15, ALCS18, BGCP11, BBR^{+}16, BNN19, BNRS13, CRY11, CBTB16, DCGG11, GD16, GP16, GUK^{+}17, HPK^{+}16, HSBW13, HSH16, HHRZ12, JSYR14, KMM^{+}18, KF12, KGM^{+}10, LHT17, LMG^{+}18a, MVZ16, MJK11, NKF^{+}16, NHG19, SCP^{+}17, SJL15, VKJ^{+}17, YZXW12, ZDM^{+}14, vdEvW13]. Technologies [KAS^{+}19]. Tectonic [CBC^{+}16, CPGG19]. Teichmüller [NC16]. Tele [GBS19]. Tele-Colorimeter [GBS19]. Template [CDM^{+}17, FV14, KWW^{+}14, NJGW15]. Template-Based [FV14]. Temporal [BHRD^{+}15, BDA^{+}17, BBT11, BBBL11, BBL12, CLC12, DER^{+}10, DPD^{+}15, DBS^{+}11, FLJ^{+}14, FE17, HRD^{+}15, HJM^{+}11, IGMK^{+}16, MO10, MHK^{+}19a, MSW10, OK12, PCBS16, RPMO13, SYM^{+}12, SR19, TPRH11, WG11, WBT19, XSQ13, KBÖ^{+}14, SBE19a, ZZLX17, Fre18]. Temporal-Coherent [MO10]. Temporally [BBH13, BIZ18, CS16, FHHJ18, SPCR14]. Temporo [BJG^{+}15]. Tendrils [WC16]. Tensor [AH11, BÖK11, JWH19a, JKLS10, MVB^{+}17, ORT18, SK10, Sch11, SMP13, dGLB^{+}14]. Tensors [FKS^{+}10, GRT19, JKLS10, KASH13, SRWS10, SK10, SK16a, ZCH^{+}17]. term [KTW^{+}13]. Terminal [KPK10]. Terms [VCRG14]. Ternary [MRAS17]. Terrain [ACA18, CDS16, CBC^{+}16, GMM15, GGP^{+}19, GGP^{+}15, TGM12, TRAW12, HGA^{+}10]. Terrains [CBC^{+}16, GDGP16]. tesselated [CBV^{+}14]. Tessellated [HSS19]. Tessellation [JH12, LBG16, LPD14, LLW12, MSWI12, NKF^{+}16, OT11]. Tessellation-Independent [MSWI12]. Tessellations [WHWB16]. Test [AMTMH12]. Tests [AAS^{+}16]. Tetrahedra [MMF10]. Tetrahedral [BXH10, BHU10a, NZH^{+}18, WIFD13]. Tetrahedron [DKY16]. TexNN [POCM19]. Text [ARLC^{+}13, ARRO^{+}17, CBK^{+}17, CY11, CWG11, EAGA^{+}16, EASKC18, EHH19, HRD^{+}15, HC14, JFCS17, KvLB14, LKC^{+}12, MHHH15, NW13, OKK13, OSR^{+}14, PTT^{+}12, RPSF15, RAMG15, SSDK12, SSS^{+}12, SSSG16, WPW^{+}11, ZAM^{+}16, ZOM19]. TextDNA [SSSG16]. Textile [ACG^{+}17]. Textiles [LZB17]. Texton [GLM17]. Textual [AEWQ^{+}15]. Texture [ACOM12, And12, Csé18, GGV^{+}19, GD10, GDG12, HWA^{+}10, HFM10, HLS12, JLKL16, JCW11, KPRN11, KNL^{+}15, KKS^{+}12, KKPL19, KDC17, KGR^{+}16, LGH13, LJKL17, LWS^{+}16, LFA^{+}15, MP12a, MJH^{+}17, NS11, OVB^{+}15, OBGB11, PDP^{+}15, POCM19, RSK10, RSK13, SD10a, SKZ13, SRK13, SC10, SL11, SXLS19, SCQ^{+}19, TGM12, WSCP13, YLT19, ZFE16, AHTAM14, LLD12, MRM^{+}18, VB14b, WGK^{+}19]. Texture-Based [TGM12]. Texture-by-Numbers [SL11]. Textured [CP10, LMG^{+}18b, BCGS13, CVDL16]. Textures [ACOM12, And10, BAAR14, FBL16, GKHF14, GSDC17, HWA^{+}10, JLKL16, KNL^{+}15, KBS11b, KSG^{+}15, LLD12, LT12, LLN^{+}14, LWS^{+}16, LZB17, MGC^{+}16, MJH^{+}17, NRM^{+}12, NS11, OBGB11, SGSP15, YLL15, ZFE16, ZSL^{+}17, CVCH14]. Textures/Mapping [FBL16, JLKL16, LWS^{+}16, MGC^{+}16, ZFE16]. Texturing [DKB^{+}16, GWO^{+}10, GDG12, KGR^{+}16, MGC^{+}16, PBMG15]. Their [CDA^{+}14, SHS^{+}17, CTL13, NWHWfD16]. Theme [NPCB17]. Theoretic [BRB^{+}13, ZC18]. Theory [CBK^{+}17, GMD10, LF10, MSS^{+}10]. Therapy [CWS^{+}17, RWS^{+}10, SRM^{+}19]. There [DH16]. Thermal [HSK^{+}10]. Thickness [ZLW^{+}16, ZXZ^{+}17]. Thin [TDDD18]. Thingi10K [ZJ18]. thinning [LCLJ10]. Thousand [BHW17]. ThreadReconstructor [EASKC18]. Three [MC14, YKS^{+}19]. Three-Component [YKS^{+}19]. Three-Dimensional [MC14]. Threshold [KRM^{+}15]. throughfall [WJG^{+}16]. Thumbnail [LSN^{+}14]. Thumbnails [KGAC15, SSDK12]. Ties [CDA^{+}14]. Tight [KKBL15, RLGH15, WTTM15, WNS^{+}10]. TightCCD [WTTM15]. Tile [AMTMH12, MH17]. Tileable [MJH^{+}17]. Tiles [DJSJ19, WPHC16]. Tiling [KWS16]. Tilings [SD10a]. Time [AVR10, AKMM11, ARH12, AAB^{+}10, AMT^{+}12, BHRD^{+}15, BDA^{+}17, BSW10, BPA16, BHR^{+}19, BPKB14, BHW11, BM15, BLW11, BBL12, BKW13, CEG^{+}18, DH16, FHHJ18, FL19, GO10, GMAG15, GAM17, GRPF16, GT15, GKT16, HA11, HML^{+}19, HCBW19, HJM^{+}11, HRWW12, HGRS^{+}17, HREB11, HR10, IGMK^{+}16, ISYM15, JMV^{+}15, JKL13, JZYP18, KCJM16, KMHG13, KSCN19, KMB^{+}17, KK14, KBKL10, KWS^{+}15, KER^{+}14, LDG19, LDdLRB16, LMP13, LKEP14, LE13, LCP^{+}12, LDGN15, LL18, LLB^{+}10, LLD10, LMK^{+}15, LRB^{+}15, MD19, MO10, MBES16, MW11, MBM13, MBJ^{+}15, MC10b, MLD^{+}18, MPBM^{+}17, NKF^{+}16, OT11, PBK10, RL15, RHL12, SPB^{+}19, SW10, SW17, SWP11, SYM^{+}12, SBD15a, SKZF11, SSB^{+}14, SAAF18, SWG16, SBM^{+}14, SDHD17, SHS^{+}17, TST^{+}15, TFA^{+}11, TCM10, VVE^{+}10, VWH18, WSCP13, WRS^{+}13, WG11, WRK^{+}16, WDZ17, WWT^{+}16, WTH^{+}13]. Time [WWH^{+}14, XWZB17, YHGT10, YLRC10, ZFAQ13, ZZT15, ZC18, vDvdEvW19, ATK17, BNH10, BN12, CRC^{+}15, CH12, CCI13, DER^{+}10, HNJ^{+}14, HKSK18, IFDN12, JSLW14, JPCC14, KGK19, KYM19, LCC^{+}17, LJH13, MRM^{+}18, MAG19, NKLN10, PSDB^{+}10, RTK^{+}14, SM14b, SGEM16, Tok15b, WWH^{+}10, WMB15, WJG^{+}16, XLL^{+}10, YCL^{+}17, YWY10, YLX^{+}16, ZFE16, ZM16, ZST^{+}10, ZRJ^{+}15, DKG15, EHH19]. Time-anchored [SHS^{+}17]. Time-Continuous [GAM17]. Time-Dependent [BPKB14, HGRS^{+}17, SW10, SW17]. Time-Discrete [HRWW12]. Time-Domain [LL18]. Time-evolving [CEG^{+}18]. Time-in-Space [AAB^{+}10]. Time-Lapse [SSB^{+}14]. Time-Of-Flight [HA11, KBKL10]. Time-Resolved [JMV^{+}15]. Time-Series [AKMM11, ARH12, BBL12, GRPF16, SBM^{+}14, TFA^{+}11, WG11, KWS^{+}15, LRB^{+}15, RL15]. Time-Specific [LMK^{+}15]. Time-Varying [FL19, WRS^{+}13, YLRC10, ZC18, TCM10, MAG19, DKG15]. TimeArcs [DPF16]. Timeline [DMS14, ODS18]. Tissue [RvdHD^{+}15]. Tissues [NMM^{+}19]. TOC [Ano16k]. Toleranced [CD10]. Tolerant [CK11a]. Tomographic [FEM^{+}19]. Tomographies [GBM^{+}19]. Tomography [HRRR18]. Tone [CLLC15, Deb18, EWMU13, GKHF14, MBDC15, SLCL19, UMM^{+}10, ZHLW18, EMU17]. Tone- [CLLC15]. Tone-Mapping [MBDC15, EMU17]. Tool [FEM^{+}19, HBRFP19, HAGO19, MWS^{+}10, PDW^{+}14, SJL15, WNS^{+}10, WCM15, YNM^{+}13]. Tools [BCB^{+}15, HMTH13, NHG19]. Topic [EAGA^{+}16, GBM^{+}19, LKC^{+}12]. Topic-Space [EAGA^{+}16]. Topical [WDM^{+}12]. Topical-Level [WDM^{+}12]. Topographic [And10]. Topological [BSK16, COO15, CDA^{+}14, HW10, Lai13, LA11, NHL16, PSO18, Sch11, TWS^{+}11]. Topologically [Gro16, LKSD17]. Topologies [KFA^{+}14, PCBL16]. Topology [AH11, BSW^{+}12, BWH^{+}11, DGP17, FKS^{+}10, FNH^{+}17, GST14, GT16b, GG17, GZH^{+}19, GBG^{+}14, HLH^{+}16a, HSmCY13, HRS18, HHC^{+}13, JWH19a, JWL^{+}13, KWS^{+}15, Lai13, LGH13, LJB^{+}12, LRB^{+}15, MRM^{+}18, MKSS12, OGHT10, PPF^{+}11, RL15, SW10, SSW14a, SV14, SRWS10, Sch11, SN12, Szy11, VMH^{+}13, WLJ^{+}18, WGS10, WIFD13, YSC^{+}18, CHWL19]. Topology-Based [PPF^{+}11, AH11, BWH^{+}11, HLH^{+}16a, Sch11, Szy11, VMH^{+}13, WGS10]. Topology-Change-Aware [JWL^{+}13]. Topology-driven [MRM^{+}18]. TopTom [GBM^{+}19]. Tori [CG19]. toric [SZ18]. Tortorelli [BCGL18]. Touch [BSY^{+}19, JSH^{+}13, KGP^{+}12, SHW^{+}18]. Touch/Tangible [BSY^{+}19]. TouchTone [LCG10]. Tour [CY14]. tourist [BMWW14]. Toxicity [RCMA^{+}18]. Trace [IIX^{+}19]. Traced [PJSH15]. Traceless [JKLS10]. Tracer [DHS^{+}13, MFW18]. Traces [SMM13]. Tracing [Áfr12, ÁSK14, BYM18, BMDS19, BAAM17, Bik12, BNH^{+}16, BLW11, CBTB16, CCI13, CNS^{+}11, DKY16, DBK11, ENSB13, GL10a, GPP^{+}10, GAM17, GHB15, HOB^{+}19, HWU^{+}19, HH11, HHH12, IH11, JKL^{+}16, KTN10, KBS11a, KBK^{+}10, KSS^{+}15, KZZM12, KF12, LWY^{+}11, MW11, MHA17, MHK^{+}19a, MBJ^{+}15, MMG18, MJL^{+}13, NM14, NB15, OT11, PBPP11, PGKS17, PBE18, PPM^{+}16, PRDD15, TSdSK13, UWA^{+}19, VF16, VHB16, WRK^{+}16, YWC^{+}10]. Tracing/Appearance [DKY16, VF16, WRK^{+}16].

Track [GVS^{+}15, CD18]. Tracking [BKR^{+}17, BH15, CK11b, EGG^{+}15, HBO^{+}10, HYL^{+}15, KAS^{+}19, KRS^{+}13, KER^{+}14, LWL^{+}16a, LWM^{+}17, MBM13, MSGT18, OAM^{+}18, PEPM12, SY13, SW17, SK17, TST^{+}15, TSB16, WH17a, ZTG^{+}18]. Tracts [SEI10]. Trade [BMPM12, HGRS^{+}17, VVE^{+}10]. Trade-off [VVE^{+}10]. Trade-Offs [BMPM12]. Traffic [BJB^{+}18, BJA^{+}15, CWL^{+}15, PDV^{+}15, SMvdWvW15, SWML10, ZSJT19, GDAU14]. Trail [LHT17]. Trails [ZSJT19]. Train [WBFvL17]. Training [HGB^{+}10, KMB^{+}17, YDP19]. Trajectories [FKSS13, GHWG14, NSM19, AKS^{+}19]. Trajectory [KTW^{+}13, WVV11]. TrajectoryLenses [KTW^{+}13]. Transductive [XSX^{+}14]. Transfer [AP10a, ARG^{+}16, BHS^{+}17, BP19, CNCO15, CJZW12, GOH^{+}10, HZL19, LP15, LCUR14, LKG^{+}16, MGG^{+}10a, NRM^{+}12, NKB14, OVB^{+}15, SPB^{+}17, SS15b, TM13, WZL^{+}12, WZL^{+}17, WDK^{+}13, XSXM13, XXZC13, XWL^{+}15, YYZZ18, ZZLX17, ZXTD10, MHS^{+}14]. Transferring [PFC15]. Transfinite [CG16c]. Transform [GL12, RAV^{+}19, SLE18, XWY^{+}15, YYG18]. transformables [YZC18]. Transformation [ATF12, BWH^{+}11, GAWJ15, MD19, MFNP13, SO12]. Transformations [KLCF10, LLM^{+}17, EKB14]. Transforms [PB11]. Transient [MGJ^{+}19]. Transit [CY14]. Transit-Centric [CY14]. Transition [McC11]. Transitions [BNH10, KCH19, STKD12]. Translation [NW13]. Translucency [MR17]. Translucent [BCRA12, HWL18, WWH^{+}10, YZXW12]. Transparency [LK11, NPW10, SVLL10, SBF15, SGG15, YHGT10]. Transparent [CCC^{+}14, DZC11, ER18, IKL^{+}10, MC14, RLH17a, RLMB^{+}14, WWT^{+}16]. Transparent-Object [DZC11]. Transport [CJWL19, HKD15, HEV^{+}16, HMRR19, HH10, HR10, JMV^{+}15, JA18, JG19, JR16, KD13, LDdLRB16, LF10, Mér11, MGN17, NGHJ18, RGG15, RKRD12, SKGM^{+}17, WPH^{+}12, ZAD15, dGCSAD11]. Transportation [ZFA^{+}16]. Travel [WTLY12]. Travel-Route-Centered [WTLY12]. Traversal [ÁSK14, GL10a, HH11, HPMB19, IH11, KBLE19, NM14]. TrayGen [YH13]. Treatment [CFS14, vPGL^{+}14]. Tree [BJB^{+}18, HL19, SPH11, WYZB14, XL10, XM15, YLG^{+}18, ZBM^{+}17, CCI13, HH11, BJB^{+}18, PSC10]. Treemaps [NB12]. TreeMatrix [RMF12]. Trees [AT10, BLS^{+}17, CCLN10, ERHH11, KS13b, LPK13, LDY10, NBHN17, SJ13a, SR19, SHL^{+}14, SPK^{+}14, VHB16, WLJ^{+}18]. Trends [BCBB16, OJMN^{+}19]. Tri [JWL^{+}13]. Tri-level [JWL^{+}13]. Triangle [CLL^{+}13, DFY14, EHA^{+}19, GLLR11, JBG17, MKSS12, PRS15, QYZ17, Ren16, SBCBG11a, WLT12, XLS^{+}14, ZCBK12, VR12]. Triangle-Based [DFY14, ZCBK12]. Triangular [MYLZ16]. Triangulated [AH11]. triangulating [ZJC13]. Triangulation [LG19, SSE^{+}14, ZSW^{+}10b]. Triangulations [AMR^{+}17, SSE^{+}14, dGLB^{+}14]. Trip [HLH^{+}16b, WBFvL17]. Tristimulus [MSHD15, OYH18]. Trivariate [HQH15]. Trivial [CDS10]. Trous [HDL11]. Truly [MC19]. Truncated [KSD14a]. Truth [CMF18, SDK^{+}15]. Try [SOC19]. Try-On [SOC19]. TSS [DKY16]. Tube [HWU^{+}19]. Tubes [LLC10b]. Tubular [LMPS16]. Tumor [HMP^{+}12, RvdHD^{+}15, RCMM^{+}16]. Tumour [HMTH13]. Tunable [dSNV^{+}17]. Tuning [ERA^{+}16, KNL^{+}15]. Tunnel [BJG^{+}15]. Tuples [KZZM12]. Turbulence [HL13, LMSF19, SJ13a]. Turning [GBS19, MJH^{+}17]. Two [Aan18, CTL13, DJM12, GKHF14, HSK14, KBS11a, LK10, LAFT12, MLP^{+}10, PK10, QWZ^{+}19, Ren16, TL19, YLHQ12]. Two-Character [HSK14, CTL13]. Two-Colored [PK10]. Two-Finger [LAFT12]. Two-Level [KBS11a]. Two-phase [QWZ^{+}19]. Two-scale [DJM12]. Two-Tone [GKHF14]. Two-Way [TL19, YLHQ12, Aan18]. Type [OP10]. Typeface [PFC15]. Typefaces [ZWXL17].

ubiquitous [DGR^{+}14]. UI [BEF17]. UK [LSF^{+}11]. Ultrasound [MMG18]. Unbiased [SKGM^{+}17, VCRG14, YIC^{+}11]. Uncalibrated [SM11]. Uncertain [GRT19, GST14, OGHT10, OT12, PPH12, PRW11, PW12]. Uncertainty [BPFG11, BMPM12, DCK12, LMK^{+}15, OT12, PPH12, PW12, PH13, PKRJ10, SMH10, SSSSW13, STP17, TWD^{+}13, WH17a, WBFvL17]. Uncertainty-Aware [BPFG11]. Uncertainty-guided [SMH10]. Unconditionally [PK15]. Unconstrained [SLWSS15]. Uncontrolled [LWS^{+}13]. Uncovering [JPN15, OKK13]. Uncut [YXHH18]. Undersampled [BS12b]. Understanding [ACC15, BHR17, CKE^{+}12, HAGO19, LCSS19, RLP10, TFA^{+}11, ZZZ^{+}19]. Undistort [BCN11b]. Unfolding [TWS^{+}11]. Unfoliaged [LDY10]. Uni [MFW18]. Uni-directional [MFW18]. Unified [KL19, RCB^{+}17a, SLZ^{+}19, SGSP15, TL16, TWT^{+}16, WWT^{+}16, YC19, YMJ^{+}17, YDT^{+}18]. Uniform [DSWH17, GPM^{+}18, CCTL12]. Uniformization [LGW18]. Uniformly [GO15]. Unifying [MDI^{+}18, OVB^{+}15]. Union [MKU15]. Unit [BMS^{+}10, JH12]. Universe [ACS^{+}17]. Unobtrusive [KÓOH13]. Unorganized [AJA11, JBG19, OM13]. Unoriented [CCLN10, KJT14, WYZC13]. Unprojected [WLT^{+}17]. Unreported [GUS12]. Unrestricted [LKSD17]. Unsharp [TMHD12]. Unsigned [MDD^{+}10]. Unsteady [GKKT13, KSW^{+}12, KG19, PPF^{+}11]. Unstructured [BM16, DLD12, PSHZ^{+}15, ZSS17, Kur15]. Unsupervised [GFK^{+}19, LVV18, NJC^{+}19]. Untangle [EASKC18]. Untangling [YMJ^{+}17]. Update [GGV^{+}19]. Updating [LTX^{+}14, ZSW^{+}10b]. Uplift [CBC^{+}16]. Upper [vBE11]. Upper-Body [vBE11]. Upsampling [DER^{+}10, JH19, JWH^{+}19b, RSD^{+}12, Tok15a]. Uptake [DHS^{+}13]. Urban [ABCN10, BWK14, BSK^{+}17, KBWS13, MKMA19, MWA^{+}13, NJGW15, NGDA16, PW17, PSC10, PDV^{+}15, VAW^{+}10, VKW^{+}12, WXR^{+}16, ZSJT19, vKvLV13, GDAU14]. Usability [CKE^{+}12, GEY12, KGP^{+}12, NW13]. Usage [SO12, SSSG16, TSdSK13]. Use [CJP^{+}19, KPK10, LW18, SR19, SNJ^{+}14]. Used [SJL15]. User [AP10a, APM^{+}11, ATF12, BAAR14, BZL^{+}18, CHH^{+}19, CCH^{+}14, DKMT18, DBB^{+}18, HK12, JTRS12, KBL19, KSG^{+}15, LCUR14, MFNP13, RC19, RPA^{+}15, SWS12, SPSK13, XBL^{+}18, YPF19]. User-Assisted [BAAR14]. User-based [CHH^{+}19]. User-Centered [BZL^{+}18]. User-Controllable [AP10a]. User-Controlled [KSG^{+}15]. User-driven [MFNP13]. User-Guided [DBB^{+}18, RC19, KBL19]. Users [BCBL13]. Using [ABD10, Áfr12, AKMM11, AMTMH12, APM^{+}11, AHMAM15, ACC15, ADN^{+}17, BLP10, BLY^{+}11, BDF^{+}14, BPFG11, BPKB14, BJB^{+}18, BFG^{+}17, BMB^{+}18, BBDA10, BAO^{+}19, BMB15b, BTM^{+}19, CRGZ10, CCW12, CG17, CCC17, CKL14, CNS^{+}11, DZCC19, DSH^{+}17, DBS^{+}11, DJM12, DZC11, ECN14, FKE13, GP12, GPK^{+}12, GAWJ15, GSW12, HCDC18, HTG14, HSS17, HL13, HBO^{+}10, HYL^{+}15, JYS19, JKK^{+}18, KTN10, KÓOH13, KWÖG18, KMJE12, LLC11, LMLG15, LZL^{+}15, LJL^{+}18, LCG10, LJZX15, LMPD15, LCY^{+}11, LKF12, LS15, MEMO14, MCH13, MbMYR15, MKMA19, MMG18, MSDK12, MDBS14, MJK11, MCB16, MJL^{+}13, OJ15, PDV^{+}15, PGK10, PNVS17, POCM19, RLH^{+}17b, RLH^{+}18, RHLH18, RKSA17, RL14, RL16, SM11, SLE17, SML15, SKR^{+}14, SJ13a, SR19, SCR^{+}18, SMM13, SASF11, SSLL14, SR14, SLCL19, SJF11, SSJ^{+}10, SKMS18, TLFC16, UMK19, UWA^{+}19, WTL15, WLT^{+}17, WWH18, WHS^{+}18]. Using [WLZX19, WHL10, WLZ13, WLI^{+}12, WZ15, XL10, XSXM13, XXZC13, XHW^{+}18, XXY^{+}18, YSM19, YB18, YL10, ZWC^{+}10, ZCZL13, ZZLX17, ZZ17, ZC18, ZVE^{+}14, AMAM13, AMT^{+}12, BSH12, BCN11a, BBB^{+}18, BW13, BLD14b, BCGL18, BMM^{+}15, BHMT13, BNH10, CML^{+}12, CCM16, CLGS18, EAGA^{+}16, FR11, FSTR13, HFM16, Hd14b, HMB17, HGA^{+}10, HHGJ15, HJS^{+}17, HSK14, HKSK18, IFDN12, Jac19, JBG17, JGH11, KPNS10, KS14, KS10, KHI^{+}19, KYM19, KKB18, KBÖ^{+}14, LBK14, LPSV14, LK17, LVV18, LWL^{+}16a, LMG^{+}18b, LLLY19, LGK16, LIK19, LSWW11, LBBC14, LLC19, MSW10, MP12a, MSAP15, MSHD15, MGC^{+}16, MMV^{+}13, NNRS15, OAO11, PGKS17, RKGS18, RRS12, RMZ13, RS19, SNA17, SPB^{+}17, SRK13, STKD12, SC10, SBLC17, SPF^{+}19, SARZL10, TM13, VMTS10, WTTM15, WGK^{+}19, WZCF15, XTZ19, XSX^{+}14, YI10, ZCK17]. Utility [DCK13]. Utilizing [YKS^{+}19].

V [CWGvW19]. V-Awake [CWGvW19]. Valence [KKL16a]. Valence-Arousal [KKL16a]. Valent [RKSA17]. Validation [CMF18, KQR^{+}19, PBK10, SAMG14, SNA17]. Valley [WZL^{+}12]. Value [SHF13]. Valued [HW10, PDV^{+}15]. Vancouver [IC11]. Variability [FKRW16, HAGO19, PRW11, RJT18, SSSSW13]. Variable [AVR10, KGR^{+}16, OBGB11, SSDK12, WT11]. Variance [LPG13, SHSK16, SMJ17, YDF^{+}10, YB18]. Variants [HV10]. Variate [CDS16, HWF^{+}17, KKS^{+}12, KZZM12, PSPM12, STMT12]. Variates [SKMS18]. Variation [EBC17, HCW17, HO17]. Variational [AJC11, BF15, HW16, JTSZ10, NC10, PR19, SSW14b, WYZB14, ZZH15, NW17]. Variations [CDM^{+}17, JTRS12, LBH12b, ZLMM16, ZCOM13, LT17]. Various [KS12b, MJK17]. Varying [BRM^{+}16a, DKG15, DBS^{+}18, FL19, KGK19, LVV18, MHRK19, PR12, WRS^{+}13, XGDC17, YLRC10, ZC18, MAG19, TCM10]. Vascular [BGCP11, KBT^{+}12]. Vase [BCRA12, TOZ^{+}11]. VBTC [KGR^{+}16]. VCBM [WBP11]. VDub [GVS^{+}15]. Vector [ABCCO13, BCBSG10, BPKB14, BSEH17, CYY^{+}11, COC15, DHP^{+}19, DZC11, FKSS13, FKS^{+}10, GRT18, GKKT13, GT16b, GG17, HKD15, HFM10, HG18, HLH^{+}16a, HWK15, HRS18, JA18, JCW11, Jes16, KG19, LTKD15, LJLH19, LKG^{+}16, OGHT10, OT12, PPH12, PDV^{+}15, SW10, SFL^{+}16, SBCBG11b, Szy11, SBL12, TBP18, WRS^{+}13, WLZT18, CFGL16, Lie17]. Vector-Based [LTKD15]. Vector-Valued [PDV^{+}15]. Vectorial [PJSH15]. Vectorising [RLMB^{+}14]. Vectorization [GZH^{+}19, KWÖG18, NS19, WZG^{+}19]. Vectors [FKS^{+}10, GRT18, HS19, JKLS10, MSS^{+}10, POS^{+}11a, SRWS10, SK10, XM15]. Vega [SSB13]. Vehicle [CWL^{+}15]. Velocity [RCB^{+}17a]. Velocity-Based [RCB^{+}17a]. Ventricle [SSM12]. Verification [CWS^{+}17, WFZ^{+}15]. Versatile [ATO17, SSN18]. Versatility [SHD15]. Vertebral [ZVE^{+}14]. Vertex [KBS11b, MTAM12, NB15, WHD17]. Vertically [SM10]. Vertices [PPT^{+}19, MM18]. Vessel [MMV^{+}13, WVV11]. Vessels [LGP14]. VFX [ATO17]. Via [CDM^{+}17, Deb18, DFY14, LGW18, ADG19, BPY16, BS12b, BBIG17, BSH15, CK14, CCC^{+}14, CYI^{+}12, CK11b, COC15, CRA^{+}17, DH16, GHX^{+}17, GSZ11, GZH^{+}19, HL19, HFL12, HZF10, HG13, HKM15, IEKM16, JJKL18, JKJL18, hKTL^{+}17, LKC^{+}12, MCHW18, Man16, MJBC13, MC17, MAM14, NO17, NMR^{+}18, OMPG13, PWS12, PW17, RBC14, SXY^{+}11, SAD^{+}16, TXL^{+}19, TLRB18, TLH19, WLL^{+}17, WCB15, WGH^{+}19, XXLX14, YXHH18, YHL^{+}16, YWY10, YWHB18, ZCW^{+}19, ZZZ^{+}19]. VIAN [HBRFP19]. Video [ASL^{+}19, AWCO10, BAAR14, BKY^{+}16, BZBM^{+}16, BIZ18, BHW11, BCK^{+}12, BTS^{+}17b, BCD^{+}12, BMS^{+}12, CRC^{+}15, CCM16, CWL19, CPZ^{+}15, DZD^{+}16, DTV15, DRA10, DPD^{+}17, EWMU13, GVS^{+}15, GO10, GO15, GHK^{+}10, GCW15, GKHF14, GTK^{+}12, GPR^{+}15, HML^{+}19, HCBW19, HGB^{+}10, HZF10, HZZ11, IY10, IS15, JCK16, JWL^{+}13, KR19, KS10, KrJC^{+}11, KGAC15, KGRG17, KYC16, KHW13, KK11b, LCC^{+}17, LJH10, LP15, LJZX15, LLN^{+}14, LLCZ16, LLB^{+}10, LLX^{+}11, MD19, MBDC15, OAO11, PCDS12, PWS12, PSHZ^{+}15, PNVS17, RKGS18, RSD^{+}12, RWSG13, SGMG17, SLWSS15, SDK^{+}15, SPK10, SHS^{+}17, TX16, TZD11, TMH11, WZH13, WBM^{+}18, XL10, XHW^{+}18, YSM19, YXHH18, YPL19, ZTW^{+}12, ZZH15, ZLSW17, ZZLX17, CVCH14, EMU17, SBB14, SPCR14]. Video-Based [ASL^{+}19, BCD^{+}12, CWL19, GHK^{+}10]. Videos [BO19, CCM16, GTK^{+}12, GPM^{+}18, LJH10, LTK12, RSD^{+}12, SBE19a, WPY^{+}19, XTZ19, ZLSW17, ZCT18]. View [BGCP11, BBP10, CWL19, FNH^{+}17, GPK^{+}12, HREB11, KTKM19, hKTL^{+}17, KMB^{+}17, KBLE19, KBK13, LL19, LFYX19, NPDD11, REH^{+}11, SE19, SBD15a, WBCGH11, WLI^{+}12, XXS^{+}15, ZZH15, DGR^{+}14, GSC18, LSR17, LJL^{+}18, SM10]. View-Adaptive [REH^{+}11]. View-Dependent [GPK^{+}12, KBK13, NPDD11, SBD15a]. Viewing [ALA^{+}18, AWCO10]. Viewpoint [HVH^{+}16, LLB^{+}10, TWC^{+}16]. Views [BRM^{+}16a, CCC^{+}14, EHT18, EAGA^{+}16, HSH16, RSM^{+}16, RAmB^{+}19, SS16, SLC^{+}19]. VIMTEX [DKG15]. VIPS [ZCZL13]. Viral [SCD^{+}16]. Virtual [AJL^{+}11, AWCO10, CKE^{+}12, DDH^{+}18, FVHK17, GPMG10, GHB^{+}17, KL14, KKK18, KAS^{+}19, LLB^{+}10, MGG^{+}10a, MKMA19, MCG^{+}19, MCB16, MJL^{+}13, NNDJ12, PP10, RKRD12, RPA^{+}15, SLE17, SOC19, SRK13, STKD12, SWML10, SGYF11, STBB14, SPV^{+}10, Tok15b, VVE^{+}10, VGS^{+}19, WC16, YJD^{+}18, dSNV^{+}17, JPCC14]. VirtualDesk [FFN18]. Vis [BAO^{+}19, VMH^{+}13]. Viscosity [WKBB18]. Viscous [HLL^{+}12, SKK10, TDF^{+}15, TL19]. VisFM [LB19]. visibilities [WZC^{+}11]. Visibility [AWJ13, BELD13, BLD14b, BJFadH19, Bru19, EJFadH13, LMS^{+}16, LPK13, MAAG12, NBMJ14, OPP10, PGSD13, UFK13, WVV11, WGK^{+}19]. Visibility-Aware [BJFadH19]. Visibility-Driven [Bru19]. Visibility/Sampling [BELD13, EJFadH13, UFK13]. Visible [Hd14b]. vision [DMCN^{+}17]. vision-based [DMCN^{+}17]. Visitors [LBK14]. Vismon [BMPM12]. VisRuption [RPMO13]. VisTrails [SASF11]. Visual [AASB19, ARLC^{+}13, ARRO^{+}17, ABW^{+}15, ASB^{+}17, AHG^{+}19, ACAA^{+}19, AAS^{+}16, ALA^{+}18, BEF17, BAT11, BLY^{+}11, BH19, BDF^{+}14, BSW^{+}14, BHR^{+}19, BSBE17, BCBL13, BBBL11, BvLBS11, BTB13, BJA^{+}15, CWGvW19, CHH^{+}19, CD18, CGM19, CBK^{+}17, CKGC14, CSG^{+}18, CWB^{+}14, CWL^{+}15, CLY17, CE19, CJP^{+}19, CAB^{+}16, CLWM11, CWS^{+}17, DCK12, DI18, DPD^{+}15, DPD^{+}17, DWT^{+}11, EASKC18, ERT^{+}17, FKRW16, FVHK17, FAFM19, FEM^{+}19, GVS^{+}15, GOH^{+}10, GLG^{+}16, GHWG14, HBRFP19, HRD^{+}15, HK16, HSK^{+}10, HJM^{+}11, HS17, HMP^{+}12, HBO^{+}10, HSH16, HGRS^{+}17, HC14, HF16, HHD^{+}12, IUDN10, JWC^{+}11, JFCS17, JE13, JvdGMR19, KRD^{+}15, KCJM16, KFH10, KBHM15, KH18, KMJE12, KCA^{+}16, KvLB14, KGGP18, KLK17, KWS^{+}15, KFM^{+}19, KVD^{+}10, KKL^{+}16b, KHI^{+}19, Lav11, LCSL18, LKC^{+}12, LKZ^{+}15, LSS^{+}12, LWPL15, LB19, LJH13, LMHH14, LWT^{+}15, LJLH19, LGH^{+}17, LRB^{+}15, ME13]. Visual [MHK^{+}19b, MRL^{+}17, MMNG17, MDI^{+}18, MLD^{+}18, MHHH15, MHDG11, NWHWfD16, NNN11, NHL16, OJS^{+}11, OSR^{+}14, OJMN^{+}19, OAM^{+}18, PSPM12, PGS^{+}16, PJR^{+}14, PZDD19, PMD12, PBK10, PPBT12, PEPM12, PDW^{+}14, PH17, Pos11b, RSM^{+}16, RvdHD^{+}15, RCMM^{+}16, RCMA^{+}18, RPK^{+}12, RMM15, RWS^{+}10, RPSF15, RHM^{+}12, Ros13, SAMS^{+}17, SGS18, SPB^{+}17, SRM^{+}19, SvW13, STMT12, SA15, SPF^{+}19, SHP^{+}19, SHK15, SvL16, SBM^{+}14, TPRH11, VCL^{+}11, VHG^{+}18, WHC15, WDC^{+}10, WG11, WKS^{+}14, WSK^{+}19, WCH^{+}15, XDC^{+}13, YXX14, YPF19, ZAM^{+}16, ZCW^{+}19, vdEvW13, vLKS^{+}11, GBM^{+}19]. Visual-Interactive [BHR^{+}19, BSW^{+}14]. Visualisation [KK17, MJK11, PKE15]. Visualising [GKPL11, ST18, BRM^{+}16a]. Visualization [ARLC^{+}13, ARH12, AHR13, AKV15, AMA^{+}16, AFK^{+}14, APM^{+}11, And10, ABCN10, AH11, BHRD^{+}15, BMH^{+}12, BAT11, BLY^{+}11, BAA^{+}16, BBDW17, BBB^{+}18, BBR^{+}16, BBK^{+}18, BWH^{+}11, BHP15, BPKB14, BBBV12, BKR^{+}17, BSK^{+}17, BMPM12, BCD^{+}12, BMB^{+}18, BBBL11, BRB^{+}13, BCN11b, BM10, BJG^{+}15, BTM^{+}19, CRY11, CNCO15, CBC^{+}15, CSFP12, CYY^{+}11, CYI^{+}12, CJP^{+}19, CY11, CWM19, CY14, CCH^{+}14, CKS^{+}15, CBKK^{+}19, CWG11, CR16b, DCK13, DKG15, DAF^{+}18, DMNV12, DKMT18, DPD^{+}17, DHS^{+}13, DWT^{+}11, DvKSW12, DBS^{+}11, ELM^{+}12, EGG^{+}15, ENSD12, ELPH19, EPAS11, FKE13, FR11, FFN18, FL19, FE17, FMH16, FM15, GOPT11, GRE11, GKLS19, GSE^{+}14a, GGM12, GKPL11, GRDE10, GSE^{+}14b, GRT14, GT15, GKT16, GT16a, GT16b, HPK^{+}16, HJM^{+}11, HW10, HLH^{+}16a, HBW11, HVH^{+}16, HMP^{+}12, HVVR18, HSJW14, HYL^{+}15, HWK15, HGB^{+}10, HSH16, HRS18]. Visualization [HPvU^{+}16, HC14, HLJ^{+}13, HKL17, HGH^{+}11, JC10, JBMC10, JWC^{+}11, JGH11, JPN15, KARC15, KSW^{+}12, KKS^{+}12, KARC17, KCL^{+}18, KKTD17, KGP^{+}12, KMA^{+}19, KMJE12, KvLB14, KS18, KKF^{+}17, KASH13, KMM^{+}18, KVD^{+}10, KBT^{+}12, KFR^{+}11, KKL^{+}16b, KPK18, KHW13, KW18, LCSS19, LDB11, LH11, LBK14, Lar10, LGP14, LKEP14, LSBP18, LVPI18, LPSV14, LMS^{+}16, LCC^{+}17, LCP^{+}12, LKZ^{+}15, LT16, LCB^{+}18, LMG^{+}18a, LS16, LFK^{+}13, LFS^{+}15, LBPH10, LBH12a, LLC^{+}12, LCD10, LWT^{+}15, LSB^{+}17, LFC14, LTC18, ME13, MK11, MGM^{+}19, MRL^{+}17, MLP^{+}10, ML17, MG10, MC10a, MKP^{+}16, MVB^{+}17, MC14, MMV^{+}13, MSK14, MRL10, MP10, MHDG11, MKRE16, NW13, NNN11, NJB^{+}11, NLB^{+}13, NSM19, NHG19, OJS^{+}11, OSR^{+}14, OAJ14, ODS18, OLF^{+}14, OGHT10, OBCCG13, POS^{+}11a, PSCC18, PSCN10, PJR^{+}14, PEP^{+}11a, PVS^{+}18, PFH^{+}18, PRW11, PW12, PDC^{+}19]. Visualization [PTA^{+}11, PPF^{+}11, PEP^{+}11b, PEPM12, Pos11b, PWH11, PH13, PBC^{+}16, PKE15, PKE17, RvdHD^{+}15, RHS^{+}12, RL19, RKRD12, RRS12, RLH17a, RWS^{+}10, RPLH11, RMH^{+}18, RG19, RHM^{+}12, RPMO13, RSSL17, RMF12, SMH10, SM11, SE19, SAMG14, SCD^{+}16, SGS18, SH14b, SKR^{+}14, SMvdWvW15, SGM^{+}11, Sch11, SSSSW13, SGRT12, SEG^{+}14, STKD12, SHCB15, SPH11, SASF11, STBG12, SvL16, SJL15, SGG15, SBM^{+}14, SLSG16, SRG^{+}19, SPT14, SHS^{+}17, SMP13, SMB^{+}17, Szy11, SBL12, TFA^{+}11, TLM16, TWC^{+}16, TLFC16, TE10, TCM10, TGK^{+}17, TPRH11, ÜFE10, VR12, VM12, VMH^{+}13, WH17a, WZC^{+}11, WZL^{+}12, WK12b, WFZ^{+}15, WT17, WG11, WKS^{+}14, WGS10, WHT12, WHP^{+}11, WPH^{+}12, WAF^{+}11, WPW^{+}11, WTLY12, WCH^{+}15, WBFvL17, YMM10, YBK^{+}12, YLRC10, YWTY12, ZLM^{+}15, ZM16, ZCH^{+}17, ZLSW17, ZZLX17, ZWRH14, ZRLS19, ZOM19, ZAD15, dHvPJV14, vDHO16, vGPNB17]. Visualization [vPJtHRV12, vPGL^{+}14, vdCAvW14, vdZLBI11, CKS^{+}16, TWMSK18, DKG15, CS16, HLH^{+}16b, KM16, Kuh12, TLM16]. Visualization-Based [BSK^{+}17]. Visualization-Guided [FMH16]. Visualization/NPR [CS16, HLH^{+}16b, KM16, TLM16]. Visualizations [BDA^{+}17, BEF17, BJG^{+}15, CDA^{+}14, GT15, KLK17, MP10, MHDG11, OJS^{+}11, OJ15, PH17, SLB^{+}18, SS16, SSKB15, SSK16, SH14a, TLW^{+}19, WWL^{+}19, WVV11]. Visualize [KGM^{+}10]. Visualizing [BKW13, BHW17, CJP^{+}19, CKE^{+}12, DPF16, FR11, FSTR13, GKL17, GOB^{+}10, HJS^{+}17, KLS^{+}17, LBK14, LMK^{+}15, LKD^{+}17, LMSF19, LFC14, MW18a, MSFM16, NMSL19, OBH^{+}11, PRW11, PKRJ10, RW18, RTJ^{+}11, RGG18, SAMG14, SCD^{+}16, SAAF18, SMM13, STP17, SSSG16, VBAW15, VBW17, WRS^{+}13, YWS^{+}14, ZFAQ13, ZFA^{+}16, ZSJT19, ZLMM16]. Visually [CJP^{+}19, DW13, GLCC17, KFR18, SHS^{+}17]. Visualnostics [LKZ^{+}15]. Vivid [ZCX19]. ViviSection [KMS^{+}13]. Vivo [DHS^{+}13]. Voids [BJG^{+}15]. Volume [AFK^{+}14, AAS^{+}16, BAT11, BHP15, BHH13, BM15, BRB^{+}13, BM10, BES15, CNCO15, CCLN10, CK13, CR16b, DMNV12, DC10, ENSD12, ER18, EPAS11, FE17, Fre18, FMH16, GM19, GGM12, GHB15, IY10, IYS^{+}13, JSYR14, JKK^{+}18, KPNS10, KMS^{+}13, KMAB15, KVS^{+}14, KWN^{+}14, KUMY10, LBH12a, LCD10, LWBP14, LA15, LCDW16, LKG^{+}16, MMF10, MB18, MC10a, MRL10, NL18, ND12, PRW11, PWH11, RHS^{+}12, RRS12, RPLH11, RGG^{+}14, SGM^{+}11, ŠPBV10, SBS^{+}17, SS15b, SGG15, SJF11, SMP13, TOZ^{+}11, TWC^{+}16, TCM10, VCRG14, VGB^{+}14b, VKJ^{+}17, VHB16, WZL^{+}12, WK12b, WTL15, WCB15, WGS10, XSE14, ZM16, ZC18, ZR13, JZW14, LN17, LN18, WZC^{+}11]. Volume-Aware [TOZ^{+}11]. Volume-of-Fluid [CK13]. Volumes [AMS16, CK10a, CIE^{+}16, CLS16, DS11b, ESRT13, GMAG15, HRRR18, IYS^{+}13, JRJ11, KMS^{+}13, KYM19, NRP11, PPL13, SKZ11, SHZD17, TWC^{+}16, VMTS10, WOH13, WW11, WTYH18, XYM13, ZH14, MAG19]. Volumetric [ACAA^{+}19, FBL16, GD10, GSZ11, GSW12, JRJ11, JJKL18, JSYR14, KKK18, LZQ13, MHK^{+}19a, MRAS17, NGHJ18, SJP^{+}13, SVLL10, SKZ11, SFL19, XSE14, YCL^{+}17]. Voronoi [BPY16, ERA^{+}16, EDPB15, HHA17, LLW12, Man16, NB12, PPT^{+}19, QYZ17, QCW^{+}18, SNA17, WHWB16, XLS^{+}14, NL13, PPL13]. Vortex [GT18, HGH^{+}11, MKP^{+}16, OT12, ZDM^{+}14]. Vortical [ZYF10]. Vorticity [HL13]. Voting [ATCO^{+}10, SAD^{+}16, TWC^{+}16]. Voxel [Áfr12, BLD14a, CNS^{+}11, DKB^{+}16, GGM12]. Voxelization [Lai13, LMMCO17, YFL19]. Voxelized [HL19]. Voxels [CCC17]. VPLs [SHD15, TH17]. VR [GPM^{+}18, LCC^{+}17, XTZ19]. vs [DBS^{+}18]. Vulnerability [CKS^{+}15, KBHM15].

Wake [PHTB12]. Walk [BPVR11]. Walkability [MKMA19]. Walking [CK11a, XLL^{+}10]. Walks [SR14]. Walkthroughs [PSC10]. Wall [DJSJ19, GHB^{+}17, KKTD17, MMP16, NLB^{+}13]. Wang [DJSJ19]. Ward [GMD10]. Warping [BMS^{+}12, CSD11, CG16b, KWSH^{+}13, LSR17, NREM14, RRS12, SC10]. was [AKMM11]. Watch [RKGS18]. Water [BSW10, KL19, PHTB12]. Watercolor [PPW18]. Waterfall [EPCV15]. watermarking [YI10]. Watertight [vKvLV13]. Wave [EBR^{+}14, KL19, MMRO13, OKG^{+}10, VWH18, WND^{+}14]. Wave-Optical [VWH18]. Wavejets [BDC18]. Wavelength [WND^{+}14]. Wavelet [BLY^{+}11, CDS16, DYFX19, HA17, MS11a, MP12a, RLH^{+}17b, RLH^{+}18, WMB15]. Wavelet-based [BLY^{+}11]. Wavelets [HDL11]. WaveMap [BLY^{+}11]. Waves [JK13]. Wavy [EKFM12]. Way [TL19, YLHQ12, Aan18]. Waypoints [ZFA^{+}16]. Waypoints-Constrained [ZFA^{+}16]. Weak [AGCO13]. Weather [DPD^{+}15, DPD^{+}17, RG19]. Weathering [IEKM16, XDR11]. Web [BAO^{+}19, CKH19, MKRE16, ZCL18]. Web-based [MKRE16]. webs [DPW11]. Weight [LS10a]. Weighted [BRM^{+}16b, CD10, DWT^{+}11, HGRS^{+}17, Man16, RC18, SSE^{+}14]. Weighting [ZHD18]. Weights [MBB19, TE18]. Well [BMB15b]. Well-Separated [BMB15b]. Wetting [RKN12]. White [LM15, OVV10]. Whole [FKE13]. Whole-Cell [FKE13]. Wide [JWH^{+}19b, SHD16]. wide-angle [SHD16]. Widgets [RLP10]. Willis [MMNG17]. Willmore [HP11]. Wind [SJ13a]. Winding [LBA10]. Windows [SWS12]. Winners [Ano15a, Bru11, Can11, Eis11]. Winners/Invited [Ano15a]. Winter [MGG^{+}10a]. Wiper [LQS^{+}19]. Wire [SFLP18, ZZCJ14]. Wired [LQS^{+}19]. wise [PEP^{+}11a, ZFCO^{+}11]. within [BMS^{+}10]. Without [SW11, ABK^{+}19, SDK^{+}15, XYLY19]. Woodblock [PPW18]. Woodification [KSG^{+}15]. Word [HG18, SSSG16, WPW^{+}11]. Wordification [PTT^{+}12]. Wordles [SSS^{+}12]. words [RAmB^{+}19]. Work [AYWM14]. Workflow [CHH^{+}19, SLSG16]. Workflow-Derived [SLSG16]. Workshop [AJL^{+}11, DS11a, LSF^{+}11, PS10]. Workshops [WBP11]. World [LWB14, LTR19, PDP^{+}15, Thi11, FZS^{+}19, RHM^{+}12]. Worlds [GPMG10, MGG^{+}10a, SWML10, STBB14]. Worn [KÓOH13]. Wrinkles [KÖS^{+}15]. Write [Lar10].

X [HRRR18]. X-Rays [HRRR18]. Xplorer [CD18].

Yarn [LMMCO17]. Yarn-Level [LMMCO17]. Year [APH^{+}12]. Young [Bru11, Eis11].

Zero [LK19]. Zoom [BDF^{+}14, KKS^{+}12, TLW^{+}19]. Zoom-Independent [KKS^{+}12]. Zooming [KGRG17].

References

Andrienko:2010:SCS

[AAB^{+}10] G. Andrienko, N. Andrienko, S. Bremm, T. Schreck, T. Von Landesberger, P. Bak, and D. Keim. Space & color: Space-in-time and time-in-space self-organizing maps for exploring spatiotemporal patterns. Computer Graphics Forum, 29(3):913– 922, June 2010. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Aanjaneya:2018:EST

[Aan18] Mridul Aanjaneya. An efficient solver for two-way coupling rigid bodies with incompressible flow. Computer Graphics Forum, 37(8):59–68, December 2018. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Amirkhanov:2016:VDA

[AAS^{+}16] Alexander Amirkhanov, Artem Amirkhanov, Dietmar Salaberger, Johann Kastner, M. Eduard Gröller, and Christoph Heinzl. Volume data applications: Visual analysis of defects in glass fiber reinforced polymers for 4DCT interrupted in situ tests. Computer Graphics Forum, 35(3):201–210, June 2016. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Aumentado-Armstrong:2017:SKS

[AAS17] T. Aumentado-Armstrong and K. Siddiqi. Spectra and kernels: Stochastic heat kernel estimation on sampled manifolds. Computer Graphics Forum, 36(5):131–138, August 2017. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Apaza-Aguero:2014:PMA

[AASB14] K. Apaza-Agüero, L. Silva, and O. R. P. Bellon. Projection mapping on arbitrary cubic cell complexes. Computer Graphics Forum, 33(1):152–163, February 2014. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Abbas:2019:CVQ

[AASB19] Mostafa M. Abbas, Michaël Aupetit, Michael Sedlmair, and Halima Bensmail. ClustMe: a visual quality measure for ranking monochrome scatterplots based on cluster patterns. Computer Graphics Forum, 38(3):225–236, June 2019. CODEN CG-FODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Artusi:2011:SSR

[ABC11] Alessandro Artusi, Francesco Banterle, and Dmitry Chetverikov. A survey of specularity removal methods. Computer Graphics Forum, 30(8):2208–2230, December 2011. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Azencot:2013:SOA

[ABCCO13] Omri Azencot, Mirela Ben-Chen, Frédéric Chazal, and Maks Ovsjanikov. Surfaces: an operator approach to tangent vector field processing. Computer Graphics Forum, 32(5):73–82, August 2013. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Abeysinghe:2010:GIS

[ABCJ10] S. S. Abeysinghe, M. L. Baker, W. Chiu, and T. Ju. Geometry II: Semi-isometric registration of line features for flexible fitting of protein structures. Computer Graphics Forum, 29(7):2243– 2252, September 2010. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Andujar:2010:VLS

[ABCN10] C. Andujar, P. Brunet, A. Chica, and I. Navazo. Visualization of large-scale urban models through multi-level relief impostors. Computer Graphics Forum, 29(8):2456–2468, December 2010. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Adams:2010:IPF

[ABD10] Andrew Adams, Jongmin Baek, and Myers Abraham Davis. Images and photography: Fast high-dimensional filtering using the permutohedral lattice. Computer Graphics Forum, 29 (2):753–762, May 2010. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Ammann:2012:NPR

[ABG^{+}12] Lucas Ammann, Pascal Barla, Gaël Guennebaud, Xavier Granier, and Patrick Reuter. Non-photorealistic rendering: Surface relief analysis for illustrative shading. Computer Graphics Forum, 31(4):1481–1490, June 2012. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Angelini:2019:SEA

[ABK^{+}19] M. Angelini, Juri Buchmüller, Daniel A. Keim, Philipp Meschenmoser, and G. Santucci. SurgeryCuts: Embedding additional information in maps without occluding features. Computer Graphics Forum, 38(3):237–247, June 2019. CODEN CG-FODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Aboulhassan:2015:EPS

[ABW^{+}15] A. Aboulhassan, D. Baum, O. Wodo, B. Ganapathysubramanian, A. Amassian, and M. Hadwiger. Engineering and physical sciences: a novel framework for visual detection and exploration of performance bottlenecks in organic photovoltaic solar cell materials. Computer Graphics Forum, 34(3):401–410, June 2015. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Argudo:2018:TSR

[ACA18] O. Argudo, A. Chica, and C. Andujar. Terrain super-resolution through aerial imagery and fully convolutional networks. Computer Graphics Forum, 37(2):101–110, May 2018. CODEN CG-FODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Agus:2019:IVV

[ACAA^{+}19] M. Agus, C. Calì, A. Al-Awami, E. Gobbetti, P. Magistretti, and M. Hadwiger. Interactive volumetric visual analysis of glycogen-derived energy absorption in nanometric brain structures. Computer Graphics Forum, 38(3):427–439, June 2019. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Aristidou:2015:EAC

[ACC15] Andreas Aristidou, Panayiotis Charalambous, and Yiorgos Chrysanthou. Emotion analysis and classification: Understanding the performers’ emotions using the LMA entities. Computer Graphics Forum, 34(6):262–276, September 2015. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Aliaga:2017:BIA

[ACG^{+}17] Carlos Aliaga, Carlos Castillo, Diego Gutierrez, Miguel A. Otaduy, Jorge Lopez-Moreno, and Adrian Jarabo. BRDFs and illumination: an appearance model for textile fibers. Computer Graphics Forum, 36(4):35–45, July 2017. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Alvarez-Cortes:2016:PLC

[ACKM16] Sara Alvarez-Cortes, Timo Kunkel, and Belen Masia. Practical low-cost recovery of spectral power distributions. Computer Graphics Forum, 35(1):166–178, February 2016. CODEN CG-FODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Aristidou:2018:SSA

[ACOHS18] A. Aristidou, D. Cohen-Or, J. K. Hodgins, and A. Shamir. Self-similarity analysis for motion capture cleaning. Computer Graphics Forum, 37(2):297–309, May 2018. CODEN CG-FODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Aiger:2012:TMR

[ACOM12] Dror Aiger, Daniel Cohen-Or, and Niloy J. Mitra. Textures and materials: Repetition maximization based texture rectification. Computer Graphics Forum, 31(2pt2):439–448, May 2012. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Axelsson:2017:DSG

[ACS^{+}17] Emil Axelsson, Jonathas Costa, Cláudio Silva, Carter Emmart, Alexander Bock, and Anders Ynnerman. Dynamic scene graph: Enabling scaling, positioning, and navigation in the universe. Computer Graphics Forum, 36(3):459–468, June 2017. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Andujar:2014:IRR

[ACV^{+}14] C. Andújar, A. Chica, M. A. Vico, S. Moya, and P. Brunet. Inexpensive reconstruction and rendering of realistic roadside landscapes. Computer Graphics Forum, 33(6):101–117, September 2014. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Azencot:2019:CSM

[ADG19] Omri Azencot, Anastasia Dubrovina, and Leonidas Guibas. Consistent shape matching via coupled optimization. Computer Graphics Forum, 38(5):13–25, August 2019. CODEN CG-FODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Allahverdi:2018:LMS

[ADMAS18] K. Allahverdi, H. Djavaherpour, A. Mahdavi-Amiri, and F. Samavati. Landscaper: a modeling system for 3D printing scale models of landscapes. Computer Graphics Forum, 37 (3):439–451, June 2018. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Arora:2017:SEI

[ADN^{+}17] R. Arora, I. Darolia, V. P. Namboodiri, K. Singh, and A. Bousseau. SketchSoup: Exploratory ideation using design sketches. Computer Graphics Forum, 36(8):302–312, December 2017. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Averbuch-Elor:2016:DDI

[AECOK16] Hadar Averbuch-Elor, Daniel Cohen-Or, and Johannes Kopf. Data-driven images: Smooth image sequences for data-driven morphing. Computer Graphics Forum, 35(2):203–213, May 2016. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Averbuch-Elor:2015:ICD

[AEWQ^{+}15] Hadar Averbuch-Elor, Yunhai Wang, Yiming Qian, Minglun Gong, Johannes Kopf, Hao Zhang, and Daniel Cohen-Or. Image collections: Distilled collections from textual image queries. Computer Graphics Forum, 34(2):131–142, May 2015. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

AlBorno:2014:APB

[AFHdL14] M. Al Borno, E. Fiume, A. Hertzmann, and M. de Lasa. Animation I (physically-based): Feedback control for rotational movements in feature space. Computer Graphics Forum, 33 (2):225–233, May 2014. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Amirkhanov:2014:HLD

[AFK^{+}14] Artem Amirkhanov, Bernhard Fröhler, Johann Kastner, Eduard Gröller, and Christoph Heinzl. High level direct volume rendering: InSpectr: Multi-modal exploration, visualization, and analysis of spectral data. Computer Graphics Forum, 33 (3):91–100, June 2014. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Afra:2012:IRT

[Áfr12] Attila T. Áfra. Interactive ray tracing of large models using voxel hierarchies. Computer Graphics Forum, 31(1):75–88, February 2012. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Asafi:2013:SAW

[AGCO13] Shmuel Asafi, Avi Goren, and Daniel Cohen-Or. Shape analysis: Weak convex decomposition by lines-of-sight. Computer Graphics Forum, 32(5):23–31, August 2013. CODEN CG-FODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Auzinger:2012:ICP

[AGJ12] T. Auzinger, M. Guthe, and S. Jeschke. Image and color processing: Analytic anti-aliasing of linear functions on polytopes. Computer Graphics Forum, 31(2pt1):335–344, May 2012. CO-DEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Agethen:2019:PSP

[AGR19] Philipp Agethen, Felix Gaisbauer, and Enrico Rukzio. A probabilistic steering parameter model for deterministic motion planning algorithms. Computer Graphics Forum, 38(1):549–563, February 2019. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Auer:2011:TBV

[AH11] C. Auer and I. Hotz. Topology-based visualization: Complete tensor field topology on 2D triangulated manifolds embedded in 3D. Computer Graphics Forum, 30(3):831–840, June 2011. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Afzal:2019:SAV

[AHG^{+}19] S. Afzal, M. M. Hittawe, S. Ghani, T. Jamil, O. Knio, M. Hadwiger, and I. Hoteit. The state of the art in visual analysis approaches for ocean and atmospheric datasets. Computer Graphics Forum, 38(3):881–907, June 2019. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Andersson:2015:FSS

[AHMAM15] M. Andersson, J. Hasselgren, J. Munkberg, and T. Akenine-Möller. Filtered stochastic shadow mapping using a layered approach. Computer Graphics Forum, 34(8):119–129, December 2015. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Aigner:2013:EES

[AHR13] W. Aigner, S. Hoffmann, and A. Rind. Evaluation: Eval-Bench: a software library for visualization evaluation. Computer Graphics Forum, 32(3pt1):41–50, June 2013. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Andersson:2014:RAT

[AHTAM14] M. Andersson, J. Hasselgren, R. Toth, and T. Akenine-Möiler. Rendering: Adaptive texture space shading for stochastic rendering. Computer Graphics Forum, 33(2):341–350, May 2014. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Akinci:2012:PSR

[AIAT12] G. Akinci, M. Ihmsen, N. Akinci, and M. Teschner. Parallel surface reconstruction for particle-based fluids. Computer Graphics Forum, 31(6):1797–1809, September 2012. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Abbasinejad:2011:SSP

[AJA11] Fatemeh Abbasinejad, Pushkar Joshi, and Nina Amenta. Section 1: Surface patches from unorganized space curves. Computer Graphics Forum, 30(5):1379–1387, August 2011. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Andrews:2011:LVS

[AJC11] James Andrews, Pushkar Joshi, and Nathan Carr. A linear variational system for modelling from curves. Computer Graphics Forum, 30(6):1850–1861, September 2011. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Artusi:2011:RIS

[AJL^{+}11] Alessandro Artusi, Morwena Joly, Geneviève Lucet, Alejandro Ribes, and Denis Pitzalis. Reports: 11th International Symposium on Virtual Reality, Archaology and Cultural Heritge: 8th Eurographics Workshop on Graphics and Cultural Heritage. Computer Graphics Forum, 30(1):233, March 2011. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Akeley:2011:ITP

[Ake11] Kurt Akeley. Invited talks: Projection and parallax. Computer Graphics Forum, 30(2):xxi, April 2011. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Averkiou:2016:ADE

[AKM16] Melinos Averkiou, Vladimir G. Kim, and Niloy J. Mitra. Autocorrelation descriptor for efficient co-alignment of 3D shape collections. Computer Graphics Forum, 35(1):261–271, February 2016. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Aigner:2011:BWR

[AKMM11] W. Aigner, C. Kainz, R. Ma, and S. Miksch. Bertin was right: an empirical evaluation of indexing to compare multivariate time-series data using line plots. Computer Graphics Forum, 30(1):215–228, March 2011. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Amirkhanov:2019:BRM

[AKS^{+}19] Aleksandr Amirkhanov, Ilona Kosiuk, Peter Szmolyan, Artem Amirkhanov, Gabriel Mistelbauer, M. Eduard Gröller, and Renata G. Raidou. Book review: ManyLands: A Journey Across 4D Phase Space of Trajectories. Computer Graphics Forum, 38 (7):191–202, October 2019. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Alam:2015:GVQ

[AKV15] Md. Jawaherul Alam, Stephen G. Kobourov, and Sankar Veeramoni. Geospatial visualization: Quantitative measures for cartogram generation techniques. Computer Graphics Forum, 34 (3):351–360, June 2015. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Averkiou:2014:SCF

[AKZM14] Melinos Averkiou, Vladimir G. Kim, Youyi Zheng, and Niloy J. Mitra. Shape collections, features: ShapeSynth: Parameterizing model collections for coupled shape exploration and synthesis. Computer Graphics Forum, 33(2):125–134, May 2014. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Andrienko:2018:VVA

[ALA^{+}18] N. Andrienko, T. Lammarsch, G. Andrienko, G. Fuchs, D. Keim, S. Miksch, and A. Rind. Viewing visual analytics as model building. Computer Graphics Forum, 37(6):275–299, September 2018. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Aristidou:2018:IKT

[ALCS18] A. Aristidou, J. Lasenby, Y. Chrysanthou, and A. Shamir. Inverse kinematics techniques in computer graphics: A survey. Computer Graphics Forum, 37(6):35–58, September 2018. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Alsallakh:2016:SAS

[AMA^{+}16] Bilal Alsallakh, Luana Micallef, Wolfgang Aigner, Helwig Hauser, Silvia Miksch, and Peter Rodgers. The state-of-the-art of set visualization. Computer Graphics Forum, 35(1):234–260, February 2016. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Andersson:2013:DDS

[AMAM13] M. Andersson, J. Munkberg, and T. Akenine-Möller. Depth and distance: Stochastic depth buffer compression using generalized plane encoding. Computer Graphics Forum, 32(2pt1):103–112, May 2013. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Abdelkader:2017:TCR

[AMR^{+}17] Ahmed Abdelkader, Ahmed H. Mahmoud, Ahmad A. Rushdi, Scott A. Mitchell, John D. Owens, and Mohamed S. Ebeida. Triangulations: a constrained resampling strategy for mesh improvement. Computer Graphics Forum, 36(5):189–201, August 2017. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Adsul:2016:PVI

[AMS16] Bharat Adsul, Jinesh Machchhar, and Milind Sohoni. Parametrization and volumes: Incorporating sharp features in the general solid sweep framework. Computer Graphics Forum, 35(5):21– 31, August 2016. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Auer:2012:RTF

[AMT^{+}12] S. Auer, C. B. Macdonald, M. Treib, J. Schneider, and R. Westermann. Real-time fluid effects on surfaces using the closest point method. Computer Graphics Forum, 31(6):1909–1923, September 2012. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Akenine-Moller:2012:EDF

[AMTMH12] Tomas Akenine-Möller, Robert Toth, Jacob Munkberg, and Jon Hasselgren. Efficient depth of field rasterization using a tile test based on half-space culling. Computer Graphics Forum, 31(1): 3–a18, February 2012. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Abdul-Massih:2017:MSR

[AMYB17] M. Abdul-Massih, I. Yoo, and B. Benes. Motion style retargeting to characters with different morphologies. Computer Graphics Forum, 36(6):86–99, September 2017. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Andujar:2010:HTM

[And10] C. Andujar. Hardware: Topographic map visualization from adaptively compressed textures. Computer Graphics Forum, 29(3):1083–1092, June 2010. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Andujar:2012:ACT

[And12] C. Andujar. Adaptive compression of texture pyramids. Computer Graphics Forum, 31(6):1973–1983, September 2012. CO-DEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Anonymous:2010:REG

[Ano10a] Anonymous. Reports: 31st EUROGRAPHICS General Assembly. Computer Graphics Forum, 29(8):2602–2607, December 2010. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Anonymous:2010:RNE

[Ano10b] Anonymous. Reports: New EUROGRAPHICS Fellows. Computer Graphics Forum, 29(6):2005–2006, September 2010. CO-DEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Anonymous:2010:RRS

[Ano10c] Anonymous. Reports: Report of the Statutory Auditors to the General Meeting of the Members of Eurographics Association Geneva. Computer Graphics Forum, 29(8):2601, December 2010. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Anonymous:2010:RR

[Ano10d] Anonymous. Reports: Reviewers. Computer Graphics Forum, 29(8):2608, December 2010. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Anonymous:2011:IIIa

[Ano11a] Anonymous. Issue information: Issue information. Computer Graphics Forum, 30(1):i–ii, March 2011. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Anonymous:2011:IIIb

[Ano11b] Anonymous. Issue information: Issue information. Computer Graphics Forum, 30(8):i–ii, December 2011. CODEN CG-FODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Anonymous:2011:REG

[Ano11c] Anonymous. Reports: 32nd EUROGRAPHICS General Assembly. Computer Graphics Forum, 30(8):2452–2456, December 2011. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Anonymous:2011:RNE

[Ano11d] Anonymous. Reports: New EUROGRAPHICS Fellows. Computer Graphics Forum, 30(6):1862–1864, September 2011. CO-DEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Anonymous:2011:RRS

[Ano11e] Anonymous. Reports: Report of the Statutory Auditors to the General Meeting of the Members of Eurographics Association Geneva. Computer Graphics Forum, 30(6):1868, September 2011. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Anonymous:2011:RR

[Ano11f] Anonymous. Reviewers: Reviewers. Computer Graphics Forum, 30(8):2459, December 2011. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Anonymous:2012:IIIa

[Ano12a] Anonymous. Issue information: Issue information. Computer Graphics Forum, 31(1):i, February 2012. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Anonymous:2012:IIIb

[Ano12b] Anonymous. Issue information: Issue information. Computer Graphics Forum, 31(6):i, September 2012. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Anonymous:2012:IIIc

[Ano12c] Anonymous. Issue information: Issue information. Computer Graphics Forum, 31(8):i, December 2012. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Anonymous:2012:REG

[Ano12d] Anonymous. Reports: 33rd EUROGRAPHICS General Assembly. Computer Graphics Forum, 31(6):1987–1991, September 2012. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Anonymous:2012:RNE

[Ano12e] Anonymous. Reports: New EUROGRAPHICS Fellows. Computer Graphics Forum, 31(6):1984–1985, September 2012. CO-DEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Anonymous:2012:RRS

[Ano12f] Anonymous. Reports: Reports of the Statutory Auditors to the General Meeting of the Members of Eurographics Association, Geneva. Computer Graphics Forum, 31(6):1986, September 2012. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Anonymous:2012:RR

[Ano12g] Anonymous. Reviewers: Reviewers. Computer Graphics Forum, 31(8):2567–2568, December 2012. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Anonymous:2013:Ea

[Ano13a] Anonymous. Erratum. Computer Graphics Forum, 32(1):219, February 2013. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Anonymous:2013:Eb

[Ano13b] Anonymous. Erratum. Computer Graphics Forum, 32(6):251, September 2013. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Anonymous:2013:IIIa

[Ano13c] Anonymous. Issue information: Issue information. Computer Graphics Forum, 32(1):i–ii, February 2013. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Anonymous:2013:IIIb

[Ano13d] Anonymous. Issue information: Issue information. Computer Graphics Forum, 32(6):i–ii, September 2013. CODEN CG-FODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Anonymous:2013:IIIc

[Ano13e] Anonymous. Issue information: Issue information. Computer Graphics Forum, 32(8):i–ii, December 2013. CODEN CG-FODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Anonymous:2013:REG

[Ano13f] Anonymous. Reports: 34th EUROGRAPHICS General Assembly. Computer Graphics Forum, 32(6):246–250, September 2013. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Anonymous:2013:RNE

[Ano13g] Anonymous. Reports: New EUROGRAPHICS Fellows. Computer Graphics Forum, 32(6):243–244, September 2013. CO-DEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Anonymous:2013:RPN

[Ano13h] Anonymous. Reports: Publisher’s note. Computer Graphics Forum, 32(1):218, February 2013. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Anonymous:2013:RRS

[Ano13i] Anonymous. Reports: Report of the Statutory Auditors to the General Meeting of the Members of EUROGRAPHICS Association Geneva. Computer Graphics Forum, 32(6):245, September 2013. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Anonymous:2013:RR

[Ano13j] Anonymous. Reviewers: Reviewers. Computer Graphics Forum, 32(8):271–272, December 2013. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Anonymous:2014:IIa

[Ano14a] Anonymous. Issue information. Computer Graphics Forum, 33 (1):i–ii, February 2014. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Anonymous:2014:IIb

[Ano14b] Anonymous. Issue information. Computer Graphics Forum, 33 (6):i–ii, September 2014. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Anonymous:2014:IIc

[Ano14c] Anonymous. Issue information. Computer Graphics Forum, 33 (8):i–ii, December 2014. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Anonymous:2014:RR

[Ano14d] Anonymous. Reviewers: Reviewers. Computer Graphics Forum, 33(8):189, December 2014. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Anonymous:2015:AWI

[Ano15a] Anonymous. Award winners/invited talks: Frontmatter. Computer Graphics Forum, 34(2):i–xxiii, May 2015. CODEN CG-FODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Anonymous:2015:E

[Ano15b] Anonymous. Erratum. Computer Graphics Forum, 34(6):327, September 2015. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Anonymous:2015:Fa

[Ano15c] Anonymous. Frontmatter. Computer Graphics Forum, 34(3): i–xiv, June 2015. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Anonymous:2015:Fb

[Ano15d] Anonymous. Frontmatter. Computer Graphics Forum, 34(4): i–xi, July 2015. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Anonymous:2015:Fc

[Ano15e] Anonymous. Frontmatter. Computer Graphics Forum, 34(5): i–x, August 2015. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Anonymous:2015:F

[Ano15f] Anonymous. Frontmatter. Computer Graphics Forum, 34(7):i– xv, October 2015. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Anonymous:2015:IIa

[Ano15g] Anonymous. Issue information. Computer Graphics Forum, 34 (1):i–ii, February 2015. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Anonymous:2015:IIb

[Ano15h] Anonymous. Issue information. Computer Graphics Forum, 34 (6):i–ii, September 2015. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Anonymous:2015:II

[Ano15i] Anonymous. Issue information. Computer Graphics Forum, 34 (8):i–ii, December 2015. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Anonymous:2015:R

[Ano15j] Anonymous. Reviewers. Computer Graphics Forum, 34(8):167, December 2015. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Anonymous:2016:FMa

[Ano16a] Anonymous. Front matter. Computer Graphics Forum, 35(2): i–xxi, May 2016. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Anonymous:2016:FMb

[Ano16b] Anonymous. Front matter. Computer Graphics Forum, 35(2): xxii–xxvi, May 2016. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Anonymous:2016:FMc

[Ano16c] Anonymous. Front matter. Computer Graphics Forum, 35(3): i–xiv, June 2016. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Anonymous:2016:FMd

[Ano16d] Anonymous. Front matter. Computer Graphics Forum, 35(3): xv–xx, June 2016. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Anonymous:2016:FMe

[Ano16e] Anonymous. Front matter. Computer Graphics Forum, 35(4): i–ix, July 2016. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Anonymous:2016:FMf

[Ano16f] Anonymous. Front matter. Computer Graphics Forum, 35(5): i–ix, August 2016. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Anonymous:2016:FM

[Ano16g] Anonymous. Front matter. Computer Graphics Forum, 35(7): i–xviii, October 2016. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Anonymous:2016:IIa

[Ano16h] Anonymous. Issue information. Computer Graphics Forum, 35 (1):i–ii, February 2016. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Anonymous:2016:IIb

[Ano16i] Anonymous. Issue information. Computer Graphics Forum, 35 (6):1–3, September 2016. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Anonymous:2016:II

[Ano16j] Anonymous. Issue information. Computer Graphics Forum, 35 (8):1–3, December 2016. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Anonymous:2016:IIT

[Ano16k] Anonymous. Issue information — TOC. Computer Graphics Forum, 35(1):1, February 2016. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Anonymous:2016:RL

[Ano16l] Anonymous. Report: Lauren. Computer Graphics Forum, 35 (1):323–324, February 2016. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Anonymous:2016:R

[Ano16m] Anonymous. Reviewers. Computer Graphics Forum, 35(8): 120–121, December 2016. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Anonymous:2017:FMa

[Ano17a] Anonymous. Front matter. Computer Graphics Forum, 36(2): i–v, May 2017. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Anonymous:2017:FMb

[Ano17b] Anonymous. Front matter. Computer Graphics Forum, 36(2): i–xvii, May 2017. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Anonymous:2017:FMc

[Ano17c] Anonymous. Front matter. Computer Graphics Forum, 36(3): i–vi, June 2017. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Anonymous:2017:FMd

[Ano17d] Anonymous. Front matter. Computer Graphics Forum, 36(3): i–xv, June 2017. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Anonymous:2017:FMe

[Ano17e] Anonymous. Front matter. Computer Graphics Forum, 36(4): i–ix, July 2017. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Anonymous:2017:FMf

[Ano17f] Anonymous. Front matter. Computer Graphics Forum, 36(5): i–x, August 2017. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Anonymous:2017:FMg

[Ano17g] Anonymous. Front matter. Computer Graphics Forum, 36(7):i– xv, October 2017. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Anonymous:2017:IIa

[Ano17h] Anonymous. Issue information. Computer Graphics Forum, 36 (1):1–4, January 2017. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Anonymous:2017:IIb

[Ano17i] Anonymous. Issue information. Computer Graphics Forum, 36 (6):1–4, September 2017. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Anonymous:2017:IIc

[Ano17j] Anonymous. Issue information. Computer Graphics Forum, 36 (8):1–5, December 2017. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Anonymous:2017:R

[Ano17k] Anonymous. Reviewers. Computer Graphics Forum, 36(8): 712–713, December 2017. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Anonymous:2018:FMa

[Ano18a] Anonymous. Front matter. Computer Graphics Forum, 37(2): i–xxi, May 2018. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Anonymous:2018:FMb

[Ano18b] Anonymous. Front matter. Computer Graphics Forum, 37(2): i–v, May 2018. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Anonymous:2018:FMc

[Ano18c] Anonymous. Front matter. Computer Graphics Forum, 37(3): i–xv, June 2018. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Anonymous:2018:FMd

[Ano18d] Anonymous. Front matter. Computer Graphics Forum, 37(3): i–vi, June 2018. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Anonymous:2018:FMe

[Ano18e] Anonymous. Front matter. Computer Graphics Forum, 37(4): i–vii, July 2018. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Anonymous:2018:FMf

[Ano18f] Anonymous. Front matter. Computer Graphics Forum, 37(5): i–xi, August 2018. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Anonymous:2018:FMg

[Ano18g] Anonymous. Front matter. Computer Graphics Forum, 37(7):i– xvi, October 2018. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Anonymous:2018:FMh

[Ano18h] Anonymous. Front matter. Computer Graphics Forum, 37(8): i–vii, December 2018. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Anonymous:2018:IIa

[Ano18i] Anonymous. Issue information. Computer Graphics Forum, 37 (1):1–4, February 2018. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Anonymous:2018:IIb

[Ano18j] Anonymous. Issue information. Computer Graphics Forum, 37 (6):1–4, September 2018. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Anonymous:2019:B

[Ano19a] Anonymous. Ballet. Computer Graphics Forum, 38(1):691–692, February 2019. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Anonymous:2019:FMa

[Ano19b] Anonymous. Front matter. Computer Graphics Forum, 38(2): i–xix, May 2019. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Anonymous:2019:FMb

[Ano19c] Anonymous. Front matter. Computer Graphics Forum, 38 (2):xx–xxiii, May 2019. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Anonymous:2019:FMc

[Ano19d] Anonymous. Front matter. Computer Graphics Forum, 38(3): i–xvi, June 2019. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Anonymous:2019:FMd

[Ano19e] Anonymous. Front matter. Computer Graphics Forum, 38(3): i–vi, June 2019. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Anonymous:2019:FMe

[Ano19f] Anonymous. Front matter. Computer Graphics Forum, 38(4): i–vii, July 2019. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Anonymous:2019:FMf

[Ano19g] Anonymous. Front matter. Computer Graphics Forum, 38(5): i–x, August 2019. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Anonymous:2019:FMg

[Ano19h] Anonymous. Front matter. Computer Graphics Forum, 38(7):i– xix, October 2019. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Anonymous:2019:FMh

[Ano19i] Anonymous. Front matter. Computer Graphics Forum, 38(8): i–ix, November 2019. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Anonymous:2019:IIa

[Ano19j] Anonymous. Issue information. Computer Graphics Forum, 38 (1):1–4, February 2019. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Anonymous:2019:IIb

[Ano19k] Anonymous. Issue information. Computer Graphics Forum, 38 (6):1–4, September 2019. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Azevedo:2013:ASE

[AO13] Vinicius C. Azevedo and Manuel M. Oliveira. Animation (session 2): Efficient smoke simulation on curvilinear grids. Computer Graphics Forum, 32(7):235–244, October 2013. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

An:2010:CAH

[AP10a] Xiaobo An and Fabio Pellacini. Colour adjustment and halftoning: User-controllable color transfer. Computer Graphics Forum, 29(2):263–271, May 2010. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Attene:2010:HSR

[AP10b] Marco Attene and Giuseppe Patanè. Hierarchical structure recovery of point-sampled surfaces. Computer Graphics Forum, 29(6):1905–1920, September 2010. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Anderson:2012:DEA

[APH^{+}12] E. F. Anderson, C. E. Peters, J. Halloran, P. Every, J. Shuttleworth, F. Liarokapis, R. Lane, and M. Richards. In at the deep end: an activity-led introduction to first year creative computing. Computer Graphics Forum, 31(6):1852–1866, September 2012. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Anderson:2011:EUS

[APM^{+}11] E. W. Anderson, K. C. Potter, L. E. Matzen, J. F. Shepherd, G. A. Preston, and C. T. Silva. Evaluation: a user study of visualization effectiveness using EEG and cognitive load. Computer Graphics Forum, 30(3):791–800, June 2011. CODEN CG-FODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Archambault:2010:ERP

[APP10] Daniel Archambault, Helen C. Purchase, and Bruno Pinaud. Evaluation: The readability of path-preserving clusterings of graphs. Computer Graphics Forum, 29(3):1173–1182, June 2010. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

AlBorno:2018:RPB

[ARB^{+}18] Mazen Al Borno, Ludovic Righetti, Michael J. Black, Scott L. Delp, Eugene Fiume, and Javier Romero. Robust physicsbased motion retargeting with realistic body shapes. Computer Graphics Forum, 37(8):81–92, December 2018. CODEN CG-FODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Avril:2016:CAA

[ARG^{+}16] Quentin Avril, Sarah Ribet, Donya Ghafourzadeh, Olivier Dionne, Srinivasan Ramachandran, Martin de Lasa, Sahel Fallahdoust, and Eric Paquette. Character animation: Animation setup transfer for 3D characters. Computer Graphics Forum, 35(2):115–126, May 2016. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Aigner:2012:TSD

[ARH12] W. Aigner, A. Rind, and S. Hoffmann. Time-series data: Comparative evaluation of an interactive time-series visualization that combines quantitative data with qualitative abstractions.

Computer Graphics Forum, 31(3pt2):995–1004, June 2012. CO-DEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Abdul-Rahman:2013:TRB

[ARLC^{+}13] A. Abdul-Rahman, J. Lein, K. Coles, E. Maguire, M. Meyer, M. Wynne, C. R. Johnson, A. Trefethen, and M. Chen. Text: Rule-based visual mappings – with a case study on poetry visualization. Computer Graphics Forum, 32(3pt4):381–390, June 2013. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Arpa:2015:PIC

[ARM^{+}15] Sami Arpa, Tobias Ritschel, Karol Myszkowski, Tolga Çapın, and Hans-Peter Seidel. Purkinje images: Conveying different content for different luminance adaptations in a single image. Computer Graphics Forum, 34(1):116–126, February 2015. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Abdul-Rahman:2017:CVA

[ARRO^{+}17] A. Abdul-Rahman, G. Roe, M. Olsen, C. Gladstone, R. Whaling, N. Cronk, R. Morrissey, and M. Chen. Constructive visual analytics for text similarity detection. Computer Graphics Forum, 36(1):237–248, January 2017. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Aboulhassan:2017:CVA

[ASB^{+}17] A. Aboulhassan, R. Sicat, D. Baum, O. Wodo, and M. Hadwiger. Comparative visual analysis of structure-performance relations in complex bulk-heterojunction morphologies. Computer Graphics Forum, 36(3):329–339, June 2017. CODEN CG-FODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Arpa:2015:FHR

[ASH15] Sami Arpa, Sabine Süsstrunk, and Roger D. Hersch. Fabrication: High reliefs from 3D scenes. Computer Graphics Forum, 34(2):253–263, May 2015. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Afra:2014:SMB

[ÁSK14] Attila T. Áfra and László Szirmay-Kalos. Stackless multi-BVH traversal for CPU, MIC and GPU ray tracing. Computer Graphics Forum, 33(1):129–140, February 2014. CODEN CG-FODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Aberman:2019:DVB

[ASL^{+}19] K. Aberman, M. Shi, J. Liao, D. Lischinski, B. Chen, and D. Cohen-Or. Deep video-based performance cloning. Computer Graphics Forum, 38(2):219–233, May 2019. CODEN CG-FODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Ament:2014:RCI

[ASW14] Marco Ament, Filip Sadlo, and Daniel Weiskopf. Report: 2014 cover image: Supernova. Computer Graphics Forum, 33(1): 305–306, February 2014. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Andrysco:2010:AMT

[AT10] Nathan Andrysco and Xavier Tricoche. Acceleration: Matrix trees. Computer Graphics Forum, 29(3):963–972, June 2010. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Au:2010:SMM

[ATCO^{+}10] Oscar Kin-Chung Au, Chiew-Lan Tai, Daniel Cohen-Or, Youyi Zheng, and Hongbo Fu. Shape matching and modeling: Electors voting for fast automatic shape correspondence. Computer Graphics Forum, 29(2):645–654, May 2010. CODEN CG-FODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Au:2012:IMU

[ATF12] Oscar Kin-Chung Au, Chiew-Lan Tai, and Hongbo Fu. Interactive modeling and user interfaces: Multitouch gestures for constrained transformation of 3D objects. Computer Graphics Forum, 31(2pt3):651–660, May 2012. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Andrews:2017:GSR

[ATK17] Sheldon Andrews, Marek Teichmann, and Paul G. Kry. Geometric stiffness for real-time constrained multibody dynamics. Computer Graphics Forum, 36(2):235–246, May 2017. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Alduan:2017:DVM

[ATO17] Iván Alduán, Angel Tena, and Miguel A. Otaduy. DYVERSO: a versatile multi-phase position-based fluids solution for VFX. Computer Graphics Forum, 36(8):32–44, December 2017. CO-DEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Attene:2015:SBD

[Att15] Marco Attene. Shapes in a box: Disassembling 3D objects for efficient packing and fabrication. Computer Graphics Forum, 34(8):64–76, December 2015. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Ando:2015:FFD

[ATW15] Ryoichi Ando, Nils Thürey, and Chris Wojtan. Fluids & flows: a dimension-reduced pressure solver for liquid simulations. Computer Graphics Forum, 34(2):473–480, May 2015. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Azencot:2016:MAB

[AVBC16] Omri Azencot, Orestis Vantzos, and Mirela Ben-Chen. Mappings: Advection-based function matching on surfaces. Computer Graphics Forum, 35(5):55–64, August 2016. CODEN CG-FODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Azencot:2018:ESP

[AVBC18] Omri Azencot, Orestis Vantzos, and Mirela Ben-Chen. An explicit structure-preserving numerical scheme for EPDiff. Computer Graphics Forum, 37(5):107–119, August 2018. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Agrawal:2010:IPR

[AVR10] Amit Agrawal, Ashok Veeraraghavan, and Ramesh Raskar. Images and photography: Reinterpretable imager: Towards variable post-capture space, angle and time resolution in photography. Computer Graphics Forum, 29(2):763–772, May 2010. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Auer:2013:ASS

[AW13] S. Auer and R. Westermann. Animation (session 2): a semi-Lagrangian closest point method for deforming surfaces. Computer Graphics Forum, 32(7):207–214, October 2013. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Assa:2010:VVD

[AWCO10] J. Assa, L. Wolf, and D. Cohen-Or. Video: The virtual director: a correlation-based online viewing of human motion. Computer Graphics Forum, 29(2):595–604, May 2010. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Auzinger:2013:VAV

[AWJ13] T. Auzinger, M. Wimmer, and S. Jescke. Visibility: Analytic visibility on the GPU. Computer Graphics Forum, 32 (2pt4):409–418, May 2013. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Adams:2010:MSM

[AWO^{+}10] B. Adams, M. Wicke, M. Ovsjanikov, M. Wand, H.-P. Seidel, and L. J. Guibas. Meshless shape and motion design for multiple deformable objects. Computer Graphics Forum, 29(1):43– 59, March 2010. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Azencot:2014:SSF

[AWO^{+}14] Omri Azencot, Steffen Weißmann, Maks Ovsjanikov, Max Wardetzky, and Mirela Ben-Chen. Surfaces and shells: Functional fluids on surfaces. Computer Graphics Forum, 33(5): 237–246, August 2014. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

AlHalawani:2013:SCI

[AYLM13] Sawsan AlHalawani, Yong-Liang Yang, Han Liu, and Niloy J. Mitra. Shape construction: Interactive facades analysis and synthesis of semi-regular facades. Computer Graphics Forum, 32(2pt2):215–224, May 2013. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

AlHalawani:2014:MWM

[AYWM14] Sawsan AlHalawani, Yong-Liang Yang, Peter Wonka, and Niloy J. Mitra. Matching: What makes London work like London? Computer Graphics Forum, 33(5):157–165, August 2014. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Bauer:2016:DVS

[BAA^{+}16] A. C. Bauer, H. Abbasi, J. Ahrens, H. Childs, B. Geveci, S. Klasky, K. Moreland, P. O’Leary, V. Vishwanath, B. Whitlock, and E. W. Bethel. Distributed visualization: In situ methods, infrastructures, and applications on high performance computing platforms. Computer Graphics Forum, 35(3):577–597, June 2016. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Barringer:2017:RAE

[BAAM17] R. Barringer, M. Andersson, and T. Akenine-Möller. Ray accelerator: Efficient and flexible ray tracing on a heterogeneous architecture. Computer Graphics Forum, 36(8):166–177, December 2017. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Bai:2013:SAP

[BAAR13] Jiamin Bai, Aseem Agarwala, Maneesh Agrawala, and Ravi Ramamoorthi. Simulated and augmented photography: Automatic cinemagraph portraits. Computer Graphics Forum, 32 (4):17–25, July 2013. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Bai:2014:VTU

[BAAR14] Jiamin Bai, Aseem Agarwala, Maneesh Agrawala, and Ravi Ramamoorthi. Video and textures: User-assisted video stabilization. Computer Graphics Forum, 33(4):61–70, July 2014. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Burchett:2019:IVA

[BAO^{+}19] J. N. Burchett, D. Abramov, J. Otto, C. Artanegara, J. X. Prochaska, and A. G. Forbes. IGM–Vis: Analyzing intergalactic and circumgalactic medium absorption using quasar sightlines in a cosmic Web context. Computer Graphics Forum, 38 (3):491–504, June 2019. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Barakat:2011:VVF

[BAT11] S. Barakat, N. Andrysco, and X. Tricoche. Volume visualization: Fast extraction of high-quality crease surfaces for visual analysis. Computer Graphics Forum, 30(3):961–970, June 2011. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Birklbauer:2012:LFR

[BB12a] C. Birklbauer and O. Bimber. Light fields and reflectance: Light-field retargeting. Computer Graphics Forum, 31(2pt1): 295–303, May 2012. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Buchholz:2012:GIQ

[BB12b] Bert Buchholz and Tamy Boubekeur. Global illumination: Quantized point-based global illumination. Computer Graphics Forum, 31(4):1399–1405, June 2012. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Birklbauer:2014:IPL

[BB14] C. Birklbauer and O. Bimber. Images I: Panorama light-field imaging. Computer Graphics Forum, 33(2):43–52, May 2014. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Boughida:2017:ARN

[BB17] Malik Boughida and Tamy Boubekeur. Adding and removing noise: Bayesian collaborative denoising for Monte Carlo rendering. Computer Graphics Forum, 36(4):137–153, July 2017. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Bermano:2012:MSS

[BBAM12] Amit Bermano, Ilya Baran, Marc Alexa, and Wojciech Matusk. Making and shadowing: Shadow Pix: Multiple images from self shadowing. Computer Graphics Forum, 31(2pt3):593–602, May 2012. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Behrendt:2018:EBF

[BBB^{+}18] B. Behrendt, P. Berg, O. Beuing, B. Preim, and S. Saalfeld. Explorative blood flow visualization using dynamic line filtering based on surface features. Computer Graphics Forum, 37 (3):183–194, June 2018. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Boyandin:2011:SDV

[BBBL11] Ilya Boyandin, Enrico Bertini, Peter Bak, and Denis Lalanne. Spatiotemporal data visualization: Flowstrates: An approach for visual exploration of temporal origin-destination data. Computer Graphics Forum, 30(3):971–980, June 2011. CODEN CG-FODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Birkeland:2012:IVI

[BBBV12] Å. Birkeland, S. Bruckner, A. Brambilla, and I. Viola. Illustrative visualization: Illustrative membrane clipping. Computer Graphics Forum, 31(3pt1):905–914, June 2012. CODEN CG-FODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Bernhardt:2010:ISA

[BBCW10] Adrien Bernhardt, Loic Barthe, Marie-Paule Cani, and Brian Wyvill. Implicit surfaces and augmented reality: Implicit blending revisited. Computer Graphics Forum, 29(2):367–375, May 2010. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Buchholz:2010:BSU

[BBDA10] Bert Buchholz, Tamy Boubekeur, Doug DeCarlo, and Marc Alexa. Binary shading using appearance and geometry. Computer Graphics Forum, 29(6):1981–1992, September 2010. CO-DEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Beck:2017:TSD

[BBDW17] Fabian Beck, Michael Burch, Stephan Diehl, and Daniel Weiskopf. A taxonomy and survey of dynamic graph visualization. Computer Graphics Forum, 36(1):133–159, January 2017. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Berard:2019:PPS

[BBGB19] P. Bérard, D. Bradley, M. Gross, and T. Beeler. Practical person-specific eye rigging. Computer Graphics Forum, 38 (2):441–454, May 2019. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Barak:2013:STC

[BBH13] T. Barák, J. Bittner, and V. Havran. Shadows: Temporally coherent adaptive sampling for imperfect shadow maps. Computer Graphics Forum, 32(4):87–96, July 2013. CODEN CG-FODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Bermano:2017:MLL

[BBIG17] Amit H. Bermano, Markus Billeter, Daisuke Iwai, and Anselm Grundhöfer. Makeup lamps: Live augmentation of human faces via projection. Computer Graphics Forum, 36(2):311–323, May 2017. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Behrisch:2018:QMI

[BBK^{+}18] M. Behrisch, M. Blumenschein, N. W. Kim, L. Shao, M. El-Assady, J. Fuchs, D. Seebacher, A. Diehl, U. Brandes, H. Pfister, T. Schreck, D. Weiskopf, and D. A. Keim. Quality metrics for information visualization. Computer Graphics Forum, 37 (3):625–662, June 2018. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Balashova:2019:LSB

[BBK^{+}19] Elena Balashova, Amit H. Bermano, Vladimir G. Kim, Stephen DiVerdi, Aaron Hertzmann, and Thomas Funkhouser. Learning a stroke-based representation for fonts. Computer Graphics Forum, 38(1):429–442, February 2019. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Boyandin:2012:TSD

[BBL12] Ilya Boyandin, Enrico Bertini, and Denis Lalanne. Timeseries data: a qualitative study on the exploration of temporal changes in flow maps with animation and small-multiples. Computer Graphics Forum, 31(3pt2):1005–1014, June 2012. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Busking:2010:NAS

[BBP10] Stef Busking, Charl P. Botha, and Frits H. Post. Navigating abstract spaces: Dynamic multi-view exploration of shape spaces. Computer Graphics Forum, 29(3):973–982, June 2010. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Behrisch:2016:IVT

[BBR^{+}16] Michael Behrisch, Benjamin Bach, Nathalie Henry Riche, Tobias Schreck, and Jean-Daniel Fekete. Information visualization techniques: Matrix reordering methods for table and network visualization. Computer Graphics Forum, 35(3):693–716, June 2016. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Benard:2011:SAR

[BBT11] Pierre Bénard, Adrien Bousseau, and Joëlle Thollot. State-of-the-art report on temporal coherence for stylized animations. Computer Graphics Forum, 30(8):2367–2386, December 2011. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Bang:2015:CIR

[BCB^{+}15] Seungbae Bang, Byungkuk Choi, Roger Blanco i Ribera, Meekyoung Kim, Sung-Hee Lee, and Junyong Noh. Characters: Interactive rigging with intuitive tools. Computer Graphics Forum, 34(7):123–132, October 2015. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Biasotti:2016:RTA

[BCBB16] S. Biasotti, A. Cerri, A. Bronstein, and M. Bronstein. Recent trends, applications, and perspectives in 3D shape similarity assessment. Computer Graphics Forum, 35(6):87–119, September 2016. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Boukhelifa:2013:EEV

[BCBL13] N. Boukhelifa, W. Cancino, A. Bezerianos, and E. Lutton. Evaluation: Evolutionary visual exploration: Evaluation with expert users. Computer Graphics Forum, 32(3pt1):31–40, June 2013. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Ben-Chen:2010:SDK

[BCBSG10] Mirela Ben-Chen, Adrian Butscher, Justin Solomon, and Leonidas Guibas. Symmetry: On discrete Killing vector fields and patterns on surfaces. Computer Graphics Forum, 29(5): 1701–1711, July 2010. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Banterle:2012:LMS

[BCCS12] Francesco Banterle, Massimiliano Corsini, Paolo Cignoni, and Roberto Scopigno. A low-memory, straightforward and fast bilateral filter through subsampling in spatial domain. Computer Graphics Forum, 31(1):19–32, February 2012. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Bezerianos:2010:GGS

[BCD^{+}10] A. Bezerianos, F. Chevalier, P. Dragicevic, N. Elmqvist, and J. D. Fekete. Graphs: GraphDice: a system for exploring multivariate social networks. Computer Graphics Forum, 29 (3):863–872, June 2010. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Borgo:2012:SAR

[BCD^{+}12] R. Borgo, M. Chen, B. Daubney, E. Grundy, G. Heidemann, B. Höferlin, M. Höferlin, H. Leitte, D. Weiskopf, and X. Xie. State of the art report on video-based graphics and video visualization. Computer Graphics Forum, 31(8):2450–2477, December 2012. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Banterle:2013:IPS

[BCD^{+}13] Francesco Banterle, Marco Callieri, Matteo Dellepiane, Massimiliano Corsini, Fabio Pellacini, and Roberto Scopigno. Image processing (session 6): EnvyDepth: An interface for recovering local natural illumination from environment maps. Computer Graphics Forum, 32(7):411–420, October 2013. CODEN CG-FODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Bonneel:2018:MSM

[BCGL18] Nicolas Bonneel, David Coeurjolly, Pierre Gueth, and Jacques-Olivier Lachaud. Mumford–Shah mesh processing using the Ambrosio–Tortorelli functional. Computer Graphics Forum, 37 (7):75–85, October 2018. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Biasotti:2013:SAP

[BCGS13] S. Biasotti, A. Cerri, D. Giorgi, and M. Spagnuolo. Shape analysis: PHOG: Photometric and geometric functions for textured shape retrieval. Computer Graphics Forum, 32(5):13–22, August 2013. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

BlancoiRibera:2012:IEP

[BCK^{+}12] Roger Blanco i Ribera, Sungwoo Choi, Younghui Kim, JungJin Lee, and Junyong Noh. Image editing and processing: Video panorama for 2D to 3D conversion. Computer Graphics Forum, 31(7pt2):2213–2222, September 2012. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Baden:2018:MR

[BCK18] Alex Baden, Keenan Crane, and Misha Kazhdan. Möbius registration. Computer Graphics Forum, 37(5):211–220, August 2018. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Bickel:2018:SAS

[BCMP18] Bernd Bickel, Paolo Cignoni, Luigi Malomo, and Nico Pietroni. State of the art on stylized fabrication. Computer Graphics Forum, 37(6):325–342, September 2018. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Bando:2011:IEM

[BCN11a] Yosuke Bando, Bing-Yu Chen, and Tomoyuki Nishita. Image enhancement: Motion deblurring from a single image using circular sensor motion. Computer Graphics Forum, 30(7):1869– 1878, September 2011. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Brosz:2011:ODF

[BCN11b] John Brosz, Sheelagh Carpendale, and Miguel A. Nacenta. Overview & detail, Focus+Context visualization: The undistort lens. Computer Graphics Forum, 30(3):881–890, June 2011. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Bousseau:2011:RMO

[BCRA11] Adrien Bousseau, Emmanuelle Chapoulie, Ravi Ramamoorthi, and Maneesh Agrawala. Reflectance and materials: Optimizing environment maps for material depiction. Computer Graphics Forum, 30(4):1171–1180, June 2011. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Bousseau:2012:RSV

[BCRA12] Adrien Bousseau, Emmanuelle Chapoulie, Ravi Ramamoorthi, and Maneesh Agrawala. Reports: Shiny vase, translucent candle and metallic sculpture. Computer Graphics Forum, 31(1): 217–218, February 2012. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Bertails-Descoubes:2012:CCS

[BD12] Florence Bertails-Descoubes. Crowds, cloths, and shape deformation: Super-clothoids. Computer Graphics Forum, 31 (2pt2):509–518, May 2012. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Billen:2016:SLS

[BD16] Niels Billen and Philip Dutré. Sampling: Line sampling for direct illumination. Computer Graphics Forum, 35(4):45–55, July 2016. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Bach:2017:DFT

[BDA^{+}17] B. Bach, P. Dragicevic, D. Archambault, C. Hurter, and S. Carpendale. A descriptive framework for temporal data visualizations based on generalized space–time cubes. Computer Graphics Forum, 36(6):36–61, September 2017. CODEN CG-FODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Bearzi:2018:WLF

[BDC18] Yohann Béarzi, Julie Digne, and Raphaëlle Chaine. Wavejets: a local frequency framework for shape details amplification. Computer Graphics Forum, 37(2):13–24, May 2018. CODEN CG-FODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Behrisch:2014:VAV

[BDF^{+}14] Michael Behrisch, James Davey, Fabian Fischer, Olivier Thonnard, Tobias Schreck, Daniel Keim, and Jörn Kohlhammer. Visual analysis: Visual analysis of sets of heterogeneous matrices using projection-based distance functions and semantic zoom. Computer Graphics Forum, 33(3):411–420, June 2014. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Boulch:2014:RPP

[BdM14] Alexandre Boulch, Martin de La Gorce, and Renaud Marlet. Reconstruction: Piecewise-planar 3D reconstruction with edge and corner regularization. Computer Graphics Forum, 33(5): 55–64, August 2014. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Bouaziz:2012:MSS

[BDS^{+}12] Sofien Bouaziz, Mario Deuss, Yuliy Schwartzburg, Thibaut Weise, and Mark Pauly. Modeling: Shape-up: Shaping discrete geometry with projections. Computer Graphics Forum, 31(5): 1657–1667, August 2012. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Bauszat:2015:GIG

[BEEM15] Pablo Bauszat, Martin Eisemann, Elmar Eisemann, and Marcus Magnor. Global illumination: General and robust error estimation and reconstruction for Monte Carlo rendering. Computer Graphics Forum, 34(2):597–608, May 2015. CODEN CG-FODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Badam:2017:UEV

[BEF17] Sriram Karthik Badam, Niklas Elmqvist, and Jean-Daniel Fekete. UI elements and visualizations for supporting progressive visual analytics. Computer Graphics Forum, 36(3):491– 502, June 2017. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Bauszat:2015:SBM

[BEJM15] P. Bauszat, M. Eisemann, S. John, and M. Magnor. Samplebased manifold filtering for interactive global illumination and depth of field. Computer Graphics Forum, 34(1):265–276, February 2015. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Boscaini:2015:SMS

[BEKB15] Davide Boscaini, Davide Eynard, Drosos Kourounis, and Michael M. Bronstein. Shape manipulation: Shape-from-operator: Recovering shapes from intrinsic operators. Computer Graphics Forum, 34(2):265–274, May 2015. CODEN CG-FODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Billen:2013:VSP

[BELD13] Niels Billen, Björn Engelen, Ares Lagae, and Philip Dutré. Visibility/sampling: Probabilistic visibility evaluation for direct illumination. Computer Graphics Forum, 32(4):39–47, July 2013. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Bauszat:2011:AGI

[BEM11] Pablo Bauszat, Martin Eisemann, and Marcus Magnor. Accelerating global illumination: Guided image filtering for interactive high-quality global illumination. Computer Graphics Forum, 30(4):1361–1368, June 2011. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Bussler:2015:VRP

[BES15] M. Bußler, T. Ertl, and F. Sadlo. Volume rendering: Photoelasticity raycasting. Computer Graphics Forum, 34(3):141–150, June 2015. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Bender:2014:ISR

[BET14] Jan Bender, Kenny Erleben, and Jeff Trinkle. Interactive simulation of rigid body dynamics in computer graphics. Computer Graphics Forum, 33(1):246–270, February 2014. CODEN CG-FODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Belyaev:2015:VPB

[BF15] Alexander G. Belyaev and Pierre-Alain Fayolle. On variational and PDE-based distance function approximations. Computer Graphics Forum, 34(8):104–118, December 2015. CODEN CG-FODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Bogl:2017:CPR

[BFG^{+}17] M. Bögl, P. Filzmoser, T. Gschwandtner, T. Lammarsch, R. A. Leite, S. Miksch, and A. Rind. Cycle plot revisited: Multivariate outlier detection using a distance-based abstraction. Computer Graphics Forum, 36(3):227–238, June 2017. CODEN CG-FODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Bermano:2017:SAM

[BFR17] Amit H. Bermano, Thomas Funkhouser, and Szymon Rusinkiewicz. State of the art in methods and representations for fabricationaware design. Computer Graphics Forum, 36(2):509–535, May 2017. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Baer:2011:EPE

[BGCP11] Alexandra Baer, Rocco Gasteiger, Douglas Cunningham, and Bernhard Preim. Evaluation: Perceptual evaluation of ghosted view techniques for the exploration of vascular structures and embedded flow. Computer Graphics Forum, 30(3):811–820, June 2011. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Boye:2010:GLS

[BGS10] S. Boyé, G. Guennebaud, and C. Schlick. Geometry I: Least squares subdivision surfaces. Computer Graphics Forum, 29(7): 2021–2028, September 2010. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Bubnik:2015:AHL

[BH15] V. Bubník and V. Havran. Agile hardware: Light chisel: 6DOF pen tracking. Computer Graphics Forum, 34(2):325–336, May 2015. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Battle:2019:CEV

[BH19] Leilani Battle and Jeffrey Heer. Characterizing exploratory visual analysis: a literature review and evaluation of analytic provenance in Tableau. Computer Graphics Forum, 38(3):145– 159, June 2019. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Bittner:2013:FIB

[BHH13] Jiří Bittner, Michal Hapala, and Vlastimil Havran. Fast insertion-based optimization of bounding volume hierarchies. Computer Graphics Forum, 32(1):85–100, February 2013. CO-DEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Boulch:2013:SAS

[BHMT13] A. Boulch, S. Houllier, R. Marlet, and O. Tournaire. Shape analysis: Semantizing complex 3D scenes using constrained attribute grammars. Computer Graphics Forum, 32(5):33–42, August 2013. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Beyer:2015:SAG

[BHP15] Johanna Beyer, Markus Hadwiger, and Hanspeter Pfister. State-of-the-art in GPU-based large-scale volume visualization. Computer Graphics Forum, 34(8):13–37, December 2015. CO-DEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Bae:2017:UIC

[BHR17] Juhee Bae, Tove Helldin, and Maria Riveiro. Understanding indirect causal relationships in node-link graphs. Computer Graphics Forum, 36(3):411–421, June 2017. CODEN CG-FODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Bernard:2019:VIP

[BHR^{+}19] Jürgen Bernard, Marco Hutter, Heiko Reinemuth, Hendrik Pfeifer, Christian Bors, and Jörn Kohlhammer. Visualinteractive preprocessing of multivariate time series data. Computer Graphics Forum, 38(3):401–412, June 2019. CODEN CG-FODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Bach:2015:BVS

[BHRD^{+}15] B. Bach, N. Henry-Riche, T. Dwyer, T. Madhyastha, J-D. Fekete, and T. Grabowski. Biomedical visualization: Small MultiPiles: Piling time to explore temporal patterns in dynamic networks. Computer Graphics Forum, 34(3):31–40, June 2015. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Berkiten:2017:LDT

[BHS^{+}17] Sema Berkiten, Maciej Halber, Justin Solomon, Chongyang Ma, Hao Li, and Szymon Rusinkiewicz. Learning detail transfer based on geometric features. Computer Graphics Forum, 36 (2):361–373, May 2017. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Baker:2019:EIQ

[BHT19] A. H. Baker, D. M. Hammerling, and T. L. Turton. Evaluating image quality measures to assess the impact of lossy data compression applied to climate simulation data. Computer Graphics Forum, 38(3):517–528, June 2019. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Burkhart:2010:AFP

[BHU10a] D. Burkhart, B. Hamann, and G. Umlauf. Adaptive and feature-preserving subdivision for high-quality tetrahedral meshes. Computer Graphics Forum, 29(1):117–127, March 2010. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Burkhart:2010:FDO

[BHU10b] D. Burkhart, B. Hamann, and G. Umlauf. FEM and differential operators: Iso-geometric finite element analysis based on Catmull–Clark subdivision solids. Computer Graphics Forum, 29(5):1575–1584, July 2010. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Bie:2011:IVE

[BHW11] Xiaohui Bie, Haoda Huang, and Wencheng Wang. Image & video editing: Real time edit propagation by efficient sampling. Computer Graphics Forum, 30(7):2041–2048, September 2011. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Burch:2017:VST

[BHW17] M. Burch, M. Hlawatsch, and D. Weiskopf. Visualizing a sequence of a thousand graphs (or even more). Computer Graphics Forum, 36(3):261–271, June 2017. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Bikker:2012:IDL

[Bik12] J. Bikker. Improving data locality for efficient in-core path tracing. Computer Graphics Forum, 31(6):1936–1947, September 2012. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Bertholet:2018:TCM

[BIZ18] P. Bertholet, A. E. Ichim, and M. Zwicker. Temporally consistent motion segmentation from RGB-D video. Computer Graphics Forum, 37(6):118–134, September 2018. CODEN CG-FODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Buchmuller:2015:TVA

[BJA^{+}15] J. Buchmüller, H. Janetzko, G. Andrienko, N. Andrienko, G. Fuchs, and D. A. Keim. Traffic: Visual analytics for exploring local impact of air traffic. Computer Graphics Forum, 34(3):181–190, June 2015. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Bhatia:2018:IIT

[BJB^{+}18] H. Bhatia, N. Jain, A. Bhatele, Y. Livnat, J. Domke, V. Pascucci, and P.-T. Bremer. Interactive investigation of traffic congestion on fat-tree networks using Tree Scope. Computer Graphics Forum, 37(3):561–572, June 2018. CODEN CG-FODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Brandt:2019:VAP

[BJFadH19] Sascha Brandt, Claudius Jähn, Matthias Fischer, and Friedhelm Meyer auf der Heide. Visibility-aware progressive farthest point sampling on the GPU. Computer Graphics Forum, 38(7): 413–424, October 2019. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Byska:2015:BVM

[BJG^{+}15] J. Byška, A. Jurčík, M. E. Gröller, I. Viola, and B. Kozlíková. Biomedical visualization: MoleCollar and tunnel heat map visualizations for conveying spatio-temporo-chemical properties across and along protein voids. Computer Graphics Forum, 34(3):1–10, June 2015. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Baran:2012:MSM

[BKB^{+}12] Ilya Baran, Philipp Keller, Derek Bradley, Stelian Coros, Wojciech Jarosz, Derek Nowrouzezahrai, and Markus Gross. Making and shadowing: Manufacturing layered attenuators for multiple prescribed shadow images. Computer Graphics Forum, 31 (2pt3):603–610, May 2012. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Benes:2017:RAP

[BKD^{+}17] J. Beneš, T. Kelly, F. Děchtěrenko, J. Křivánek, and P. Müller. On realism of architectural procedural models. Computer Graphics Forum, 36(2):225–234, May 2017. CODEN CG-FODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Bemana:2019:LPI

[BKM^{+}19] M. Bemana, J. Keinert, K. Myszkowski, M. Bätz, M. Ziegler, H.-P. Seidel, and T. Ritschel. Learning to predict image-based rendering artifacts with respect to a hidden reference image. Computer Graphics Forum, 38(7):579–589, October 2019. CO-DEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Bonneel:2017:IDI

[BKPB17] Nicolas Bonneel, Balazs Kovacs, Sylvain Paris, and Kavita Bala. Intrinsic decompositions for image editing. Computer Graphics Forum, 36(2):593–609, May 2017. CODEN CG-FODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Blascheck:2017:VET

[BKR^{+}17] T. Blascheck, K. Kurzhals, M. Raschke, M. Burch, D. Weiskopf, and T. Ertl. Visualization of eye tracking data: a taxonomy and survey. Computer Graphics Forum, 36(8):260–284, December 2017. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Burch:2013:STA

[BKW13] Michael Burch, Andreas Kull, and Daniel Weiskopf. Space and time: AOI rivers for visualizing dynamic eye gaze frequencies. Computer Graphics Forum, 32(3pt3):281–290, June 2013. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Bazin:2016:IVP

[BKY^{+}16] J.-C. Bazin, C. Plüss (Kuster), G. Yu, T. Martin, A. Jacobson, and M. Gross. Images and video: Physically based video editing. Computer Graphics Forum, 35(7):421–429, October 2016. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Bukenberger:2018:HQM

[BL18] Dennis R. Bukenberger and Hendrik P. A. Lensch. Hierarchical quad meshing of 3D scanned surfaces. Computer Graphics Forum, 37(5):135–145, August 2018. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Baert:2014:CCS

[BLD14a] J. Baert, A. Lagae, and Ph. Dutré. Out-of-core construction of sparse voxel octrees. Computer Graphics Forum, 33(6):220– 227, September 2014. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Billen:2014:VPV

[BLD14b] Niels Billen, Ares Lagae, and Philip Dutré. Visibility: Probabilistic visibility evaluation using geometry proxies. Computer Graphics Forum, 33(4):143–152, July 2014. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Bommes:2011:MPS

[BLK11] David Bommes, Timm Lempfer, and Leif Kobbelt. Mesh processing and surface reconstruction: Global structure optimization of quadrilateral meshes. Computer Graphics Forum, 30 (2):375–384, April 2011. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Baran:2010:SMM

[BLP10] Ilya Baran, Jaakko Lehtinen, and Jovan Popović. Shape matching and modeling: Sketching clothoid splines using shortest paths. Computer Graphics Forum, 29(2):655–664, May 2010. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Bommes:2013:QMG

[BLP^{+}13] David Bommes, Bruno Lévy, Nico Pietroni, Enrico Puppo, Claudio Silva, Marco Tarini, and Denis Zorin. Quad-mesh generation and processing: a survey. Computer Graphics Forum, 32(6):51–76, September 2013. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Bo:2017:FDP

[BLS^{+}17] Z. Bo, L. Lu, A. Sharf, Y. Xia, O. Deussen, and B. Chen. Fabrication and design: Printable 3D trees. Computer Graphics Forum, 36(7):29–40, October 2017. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Budninskiy:2017:SKS

[BLTD17] Max Budninskiy, Beibei Liu, Yiying Tong, and Mathieu Desbrun. Spectra and kernels: Spectral affine-kernel embeddings.

Computer Graphics Forum, 36(5):117–129, August 2017. CO-DEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Benard:2010:DNP

[BLV^{+}10] P. Bénard, A. Lagae, P. Vangorp, S. Lefebvre, G. Drettakis, and J. Thollot. A dynamic noise primitive for coherent stylization. Computer Graphics Forum, 29(4):1497–1506, June 2010. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Benhabiles:2011:LBE

[BLVD11] Halim Benhabiles, Guillaume Lavoué, Jean-Philippe Vandeborre, and Mohamed Daoudi. Learning boundary edges for 3Dmesh segmentation. Computer Graphics Forum, 30(8):2170– 2182, December 2011. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Bowers:2011:RTR

[BLW11] John C. Bowers, Jonathan Leahey, and Rui Wang. Ray tracing and real-time rendering: a ray tracing approach to diffusion curves. Computer Graphics Forum, 30(4):1345–1352, June 2011. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Barlowe:2011:BDV

[BLY^{+}11] Scott Barlowe, Yujie Liu, Jing Yang, Dennis R. Livesay, Donald J. Jacobs, James Mottonen, and Deeptak Verma. Biological data visualization: WaveMap: Interactively discovering features from protein flexibility matrices using wavelet-based visual analytics. Computer Graphics Forum, 30(3):1001–1010, June 2011. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Bruckner:2010:VVI

[BM10] Stefan Bruckner and Torsten Möller. Volume visualization: Isosurface similarity maps. Computer Graphics Forum, 29(3):773– 782, June 2010. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Boulch:2012:PSF

[BM12] Alexandre Boulch and Renaud Marlet. Point sets: Fast and robust normal estimation for point clouds with sharp features. Computer Graphics Forum, 31(5):1765–1774, August 2012. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Bittner:2015:RTR

[BM15] J. Bittner and D. Meister. Real-time rendering & quantization: T-SAH: Animation optimized bounding volume hierarchies. Computer Graphics Forum, 34(2):527–536, May 2015. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Boulch:2016:DPD

[BM16] Alexandre Boulch and Renaud Marlet. Differential properties: Deep learning for robust normal estimation in unstructured point clouds. Computer Graphics Forum, 35(5):281–290, August 2016. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Bus:2015:GII

[BMB15a] N. Bus, N. H. Mustafa, and V. Biri. Global illumination: IlluminationCut. Computer Graphics Forum, 34(2):561–573, May 2015. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Bus:2015:GIU

[BMB15b] N. Bus, N. H. Mustafa, and V. Biri. Global illumination using well-separated pair decomposition. Computer Graphics Forum, 34(8):88–103, December 2015. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Borgo:2018:IVE

[BMB^{+}18] R. Borgo, L. Micallef, B. Bach, F. McGee, and B. Lee. Information visualization evaluation using crowdsourcing. Computer Graphics Forum, 37(3):573–595, June 2018. CODEN CG-FODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Bako:2019:ODI

[BMDS19] Steve Bako, Mark Meyer, Tony DeRose, and Pradeep Sen. Offline deep importance sampling for Monte Carlo path tracing. Computer Graphics Forum, 38(7):527–542, October 2019. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Bak:2012:GVS

[BMH^{+}12] P. Bak, M. Marder, S. Harary, A. Yaeli, and H. J. Ship. Geospatial visualization: Scalable detection of spatiotemporal encounters in historical movement data. Computer Graphics Forum, 31(3pt1):915–924, June 2012. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Boscaini:2015:DSS

[BMM^{+}15] D. Boscaini, J. Masci, S. Melzi, M. M. Bronstein, U. Castellani, and P. Vandergheynst. Descriptors and shape synthesis: Learning class-specific descriptors for deformable shapes using localized spectral convolutional networks. Computer Graphics Forum, 34(5):13–23, August 2015. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Bender:2014:SPB

[BMO^{+}14] Jan Bender, Matthias Müller, Miguel A. Otaduy, Matthias Teschner, and Miles Macklin. A survey on position-based simulation methods in computer graphics. Computer Graphics Forum, 33(6):228–251, September 2014. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Booshehrian:2012:VDS

[BMPM12] Maryam Booshehrian, Torsten Möller, Randall M. Peterman, and Tamara Munzner. Visualization for decision support: Vismon: Facilitating analysis of trade-offs, uncertainty, and sensitivity in fisheries management decision making. Computer Graphics Forum, 31(3pt3):1235–1244, June 2012. CODEN CG-FODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Boscaini:2016:GMA

[BMR^{+}16] D. Boscaini, J. Masci, E. Rodolà, M. M. Bronstein, and D. Cremers. Geometric modeling: Anisotropic diffusion descriptors. Computer Graphics Forum, 35(2):431–441, May 2016. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Butkiewicz:2010:SCA

[BMS^{+}10] Thomas Butkiewicz, Ross K. Meentemeyer, Douglas A. Shoemaker, Remco Chang, Zachary Wartell, and William Ribarsky. Space & color: Alleviating the modifiable areal unit problem within probe-based geospatial analyses. Computer Graphics Forum, 29(3):923–932, June 2010. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Bowles:2012:IVR

[BMS^{+}12] Huw Bowles, Kenny Mitchell, Robert W. Sumner, Jeremy Moore, and Markus Gross. Image & video retargeting: Iterative image warping. Computer Graphics Forum, 31(2pt1):237– 246, May 2012. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Birsak:2014:MNA

[BMWW14] Michael Birsak, Przemysław Musialski, Peter Wonka, and Michael Wimmer. Men and nature: Automatic generation of tourist brochures. Computer Graphics Forum, 33(2):449–458, May 2014. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Bruneton:2012:MMR

[BN12] Eric Bruneton and Fabrice Neyret. Massive models: Real-time realistic rendering and lighting of forests. Computer Graphics Forum, 31(2pt1):373–382, May 2012. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Bruneton:2010:RRT

[BNH10] Eric Bruneton, Fabrice Neyret, and Nicolas Holzschuch. Rendering I: Real-time realistic ocean lighting using seamless transitions from geometry to BRDF. Computer Graphics Forum, 29(2):487–496, May 2010. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Blumer:2016:RRR

[BNH^{+}16] Adrian Blumer, Jan Novák, Ralf Habel, Derek Nowrouzezahrai, and Wojciech Jarosz. Realistic rendering: Reduced aggregate scattering operators for path tracing. Computer Graphics Forum, 35(7):461–473, October 2016. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Bitterli:2015:SPM

[BNJ15] Benedikt Bitterli, Jan Novák, and Wojciech Jarosz. Shadows: Portal-masked environment map sampling. Computer Graphics Forum, 34(4):13–19, July 2015. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Bekos:2019:ELT

[BNN19] Michael A. Bekos, Benjamin Niedermann, and Martin Nöllenburg. External labeling techniques: a taxonomy and survey. Computer Graphics Forum, 38(3):833–860, June 2019. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Brandes:2013:ITG

[BNRS13] Ulrik Brandes, Bobo Nick, Brigitte Rockstroh, and Astrid Steffen. InfoVis techniques: Gestaltlines. Computer Graphics Forum, 32(3pt2):171–180, June 2013. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Backes:2019:PBA

[BO19] Marcos H. Backes and Manuel M. Oliveira. A PatchMatchbased approach for matte propagation in videos. Computer Graphics Forum, 38(7):651–662, October 2019. CODEN CG-FODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Birklbauer:2013:RRG

[BOB13] C. Birklbauer, S. Opelt, and O. Bimber. Rendering: Rendering gigaray light fields. Computer Graphics Forum, 32(2pt4):469– 478, May 2013. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Bilgili:2011:GBR

[BÖK11] Ahmet Bilgili, Aydn Öztürk, and Murat Kurt. A general BRDF representation based on tensor decomposition. Computer Graphics Forum, 30(8):2427–2439, December 2011. CO-DEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Besuievsky:2013:CLP

[BP13] Gonzalo Besuievsky and Gustavo Patow. Customizable LoD for procedural architecture. Computer Graphics Forum, 32(8): 26–34, December 2013. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Bruneau:2017:EEA

[BP17] J. Bruneau and J. Pettré. EACS: Effective avoidance combination strategy. Computer Graphics Forum, 36(8):108–122, December 2017. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Bosch:2019:CIB

[BP19] Carles Bosch and Gustavo Patow. Controllable image-based transfer of flow phenomena. Computer Graphics Forum, 38(1): 274–285, February 2019. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Beacco:2016:SRT

[BPA16] A. Beacco, N. Pelechano, and C. Andújar. A survey of realtime crowd rendering. Computer Graphics Forum, 35(8):32–50, December 2016. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Berger:2011:PSA

[BPFG11] W. Berger, H. Piringer, P. Filzmoser, and E. Gröller. Parameter spaces and analysis: Uncertainty-aware exploration of continuous parameter spaces using multivariate prediction. Computer Graphics Forum, 30(3):911–920, June 2011. CODEN CG-FODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Bhatia:2014:FVE

[BPKB14] H. Bhatia, V. Pascucci, R. M. Kirby, and P.-T. Bremer. Flow visualization: Extracting features from time-dependent vector fields using internal reference frames. Computer Graphics Forum, 33(3):21–30, June 2014. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Barla:2018:CBM

[BPV18] P. Barla, R. Pacanowski, and P. Vangorp. A composite BRDF model for hazy gloss. Computer Graphics Forum, 37(4):55–66, July 2018. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Benjamin:2011:MSP

[BPVR11] William Benjamin, Andrew Wood Polk, S. V. N. Vishwanathan, and Karthik Ramani. Mesh segmentation & processing: Heat walk: Robust salient segmentation of non-rigid shapes. Computer Graphics Forum, 30(7):2097–2106, September 2011. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Barton:2014:GAR

[BPW14] M. Bartoň, H. Pottmann, and J. Wallner. Geometry acquisition, reconstruction and analysis: Detection and reconstruction of freeform sweeps. Computer Graphics Forum, 33(2):23–32, May 2014. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Bennett:2016:VAP

[BPY16] H. Bennett, E. Papadopoulou, and C. Yap. Voronoi et al.: Planar minimization diagrams via subdivision with applications to anisotropic Voronoi diagrams. Computer Graphics Forum, 35(5):229–247, August 2016. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Bramon:2013:FIT

[BRB^{+}13] R. Bramon, M. Ruiz, A. Bardera, I. Boada, M. Feixas, and M. Sbert. Foundations: an information-theoretic observation channel for volume visualization. Computer Graphics Forum, 32(3pt4):411–420, June 2013. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Bashford-Rogers:2012:SCA

[BRDC12] Thomas Bashford-Rogers, Kurt Debattista, and Alan Chalmers. A significance cache for accelerating global illumination. Computer Graphics Forum, 31(6):1837–1851, September 2012. CO-DEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Beecham:2016:CVI

[BRM^{+}16a] R. Beecham, C. Rooney, S. Meier, J. Dykes, A. Slingsby, C. Turkay, J. Wood, and B. L. W. Wong. Coordinated views and interaction design: Faceted views of varying emphasis (FaVVEs): a framework for visualising multi-perspective small multiples. Computer Graphics Forum, 35(3):241–249, June 2016. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Bitterli:2016:FRN

[BRM^{+}16b] Benedikt Bitterli, Fabrice Rousselle, Bochang Moon, José A. Iglesias-Guitián, David Adler, Kenny Mitchell, Wojciech Jarosz, and Jan Novák. Faster rendering: Nonlinearly weighted first-order regression for denoising Monte Carlo renderings.

Computer Graphics Forum, 35(4):107–117, July 2016. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Bruckner:2011:AWE

[Bru11] Stefan Bruckner. Award winners: Eurographics Young researcher award. Computer Graphics Forum, 30(2):xix, April 2011. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Bruckner:2019:DVD

[Bru19] Stefan Bruckner. Dynamic visibility-driven molecular surfaces. Computer Graphics Forum, 38(2):317–329, May 2019. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Birsak:2018:SAT

[BRWM18] Michael Birsak, Florian Rist, Peter Wonka, and Przemyslaw Musialski. String art: Towards computational fabrication of string images. Computer Graphics Forum, 37(2):263–274, May 2018. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Berger:2012:CGG

[BS12a] Matthew Berger and Claudio T. Silva. Computational geometry and geometry processing: Medial kernels. Computer Graphics Forum, 31(2pt4):795–804, May 2012. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Berger:2012:CNM

[BS12b] Matthew Berger and Claudio T. Silva. Correspondence: Nonrigid matching of undersampled shapes via medial diffusion. Computer Graphics Forum, 31(5):1587–1596, August 2012. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Barrielle:2019:RPD

[BS19] V. Barrielle and N. Stoiber. Realtime performance-driven physical simulation for facial animation. Computer Graphics Forum, 38(1):151–166, February 2019. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Beacco:2011:FAO

[BSAP11] A. Beacco, B. Spanlang, C. Andujar, and N. Pelechano. A flexible approach for output-sensitive rendering of animated characters. Computer Graphics Forum, 30(8):2328–2340, December 2011. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Blascheck:2017:VCE

[BSBE17] Tanja Blascheck, Markus Schweizer, Fabian Beck, and Thomas Ertl. Visual comparison of eye movement patterns. Computer Graphics Forum, 36(3):87–97, June 2017. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Barrielle:2016:FMB

[BSC16] Vincent Barrielle, Nicolas Stoiber, and Cédric Cagniart. Faces & motion: BlendForces: a dynamic framework for facial animation. Computer Graphics Forum, 35(2):341–352, May 2016. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Bobenrieth:2018:RFS

[BSCH18] Cédric Bobenrieth, Hyewon Seo, Frédéric Cordier, and Arash Habibi. Reconstructing flowers from sketches. Computer Graphics Forum, 37(7):167–178, October 2018. CODEN CG-FODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Brandt:2017:SPT

[BSEH17] Christopher Brandt, Leonardo Scandolo, Elmar Eisemann, and Klaus Hildebrandt. Spectral processing of tangential vector fields. Computer Graphics Forum, 36(6):338–353, September 2017. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Bagher:2012:MAA

[BSH12] M. M. Bagher, C. Soler, and N. Holzschuch. Material appearance: Accurate fitting of measured reflectances using a shifted Gamma micro-facet distribution. Computer Graphics Forum, 31(4):1509–1518, June 2012. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Brandt:2015:SMO

[BSH15] Christopher Brandt, Hans-Peter Seidel, and Klaus Hildebrandt. Splines & meshes: Optimal spline approximation via ℓ_{0}- minimization. Computer Graphics Forum, 34(2):617–626, May 2015. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Barton:2013:SMC

[BSK^{+}13] M. Bartoň, L. Shi, M. Kilian, J. Wallner, and H. Pottmann. Shape modeling: Circular arc snakes and kinematic surface generation. Computer Graphics Forum, 32(2pt1):1–10, May 2013. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Barnett:2016:CCS

[BSK16] Adam Barnett, Hubert P. H. Shum, and Taku Komura. Coordinated crowd simulation with topological scene analysis. Computer Graphics Forum, 35(6):120–132, September 2016. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Bock:2017:VBA

[BSK^{+}17] Alexander Bock, Åsa Svensson, Alexander Kleiner, Jonas Lundberg, and Timo Ropinski. A visualization-based analysis system for urban search & rescue mission planning support. Computer Graphics Forum, 36(6):148–159, September 2017. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Bukenberger:2018:SCC

[BSL18] Dennis R. Bukenberger, Katharina Schwarz, and Hendrik P. A. Lensch. Stereo-consistent contours in object space. Computer Graphics Forum, 37(1):301–312, February 2018. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Benes:2011:PMG

[BŠMM11] B. Beneš, O. Št’ava, R. Měch, and G. Miller. Procedural modelling: Guided procedural modeling. Computer Graphics Forum, 30(2):325–334, April 2011. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Bagar:2010:SRL

[BSW10] Florian Bagar, Daniel Scherzer, and Michael Wimmer. Scattering and refraction: a layered particle-based fluid model for real-time rendering of water. Computer Graphics Forum, 29(4): 1383–1389, June 2010. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Bachthaler:2012:TMF

[BSW^{+}12] Sven Bachthaler, Filip Sadlo, Rudolf Weeber, Sofia Kantorovich, Christian Holm, and Daniel Weiskopf. Topology: Magnetic flux topology of 2D point dipoles. Computer Graphics Forum, 31(3pt1):955–964, June 2012. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Bernard:2014:DRV

[BSW^{+}14] Jürgen Bernard, Martin Steiger, Sven Widmer, Hendrik Lücke-Tieke, Thorsten May, and Jörn Kohlhammer. Data relationships: Visual-interactive exploration of interesting multivariate relations in mixed research data sets. Computer Graphics Forum, 33(3):291–300, June 2014. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Besancon:2019:HTT

[BSY^{+}19] Lonni Besançon, Mickael Sereno, Lingyun Yu, Mehdi Ammi, and Tobias Isenberg. Hybrid touch/tangible spatial 3D data selection. Computer Graphics Forum, 38(3):553–567, June 2019. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Broeksema:2013:VAM

[BTB13] Bertjan Broeksema, Alexandru C. Telea, and Thomas Baudel. Visual analysis of multi-dimensional categorical data sets. Computer Graphics Forum, 32(8):158–169, December 2013. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Byska:2019:ALM

[BTM^{+}19] J. Byška, T. Trautner, S. M. Marques, J. Damborský, B. Kozlíková, and M. Waldner. Analysis of long molecular dynamics simulations using interactive focus+context visualization. Computer Graphics Forum, 38(3):441–453, June 2019. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Bouaziz:2013:FAS

[BTP13] Sofien Bouaziz, Andrea Tagliasacchi, and Mark Pauly. Fitting and approximation: Sparse iterative closest point. Computer Graphics Forum, 32(5):113–123, August 2013. CODEN CG-FODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Berger:2017:SSR

[BTS^{+}17a] Matthew Berger, Andrea Tagliasacchi, Lee M. Seversky, Pierre Alliez, Gaël Guennebaud, Joshua A. Levine, Andrei Sharf, and Claudio T. Silva. A survey of surface reconstruction from point clouds. Computer Graphics Forum, 36(1):301–329, January 2017. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Bonneel:2017:DML

[BTS^{+}17b] Nicolas Bonneel, James Tompkin, Deqing Sun, Oliver Wang, Kalyan Sunkavalli, Sylvain Paris, and Hanspeter Pfister. Depth to motion, lens and filters: Consistent video filtering for camera arrays. Computer Graphics Forum, 36(2):397–407, May 2017. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Bharaj:2012:AAR

[BTST12] Gaurav Bharaj, Thorsten Thormählen, Hans-Peter Seidel, and Christian Theobalt. Animation: Automatically rigging multicomponent characters. Computer Graphics Forum, 31(2pt4): 755–764, May 2012. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Boehs:2019:SIL

[BV19] G. E. Boehs and M. L. H. Vieira. Style invariant locomotion classification for character control. Computer Graphics Forum, 38(1):537–548, February 2019. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Bremm:2011:PSA

[BvLBS11] Sebastian Bremm, Tatiana von Landesberger, Jürgen Bernard, and Tobias Schreck. Parameter spaces and analysis: Assisted descriptor selection based on visual comparative data analysis. Computer Graphics Forum, 30(3):891–900, June 2011. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Brandt:2016:CSG

[BvTH16] Christopher Brandt, Christoph von Tycowicz, and Klaus Hildebrandt. Curves & surfaces: Geometric flows of curves in shape space for processing motion of deformable objects. Computer Graphics Forum, 35(2):295–305, May 2016. CODEN CG-FODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Bhattacharya:2013:SCI

[BW13] A. Bhattacharya and R. Wenger. Surfaces: Constructing isosurfaces with sharp edges and corners using cube merging. Computer Graphics Forum, 32(3pt1):11–20, June 2013. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Barros:2017:SHS

[BW17] R. S. Barros and M. Walter. Synthesis of human skin pigmentation disorders. Computer Graphics Forum, 36(1):330–344, January 2017. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Beketayev:2011:MDV

[BWH^{+}11] K. Beketayev, G. H. Weber, M. Haranczyk, P.-T. Bremer, M. Hlawitschka, and B. Hamann. Molecular data visualization: Topology-based visualization of transformation pathways in complex chemical systems. Computer Graphics Forum, 30 (3):663–672, June 2011. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Benes:2014:PMU

[BWK14] Jan Beneš, Alexander Wilkie, and Jaroslav Křivánek. Procedural modelling of urban road networks. Computer Graphics Forum, 33(6):132–142, September 2014. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Berner:2011:SAS

[BWM^{+}11] Alexander Berner, Michael Wand, Niloy J. Mitra, Daniel Mewes, and Hans-Peter Seidel. Shape analysis: Shape analysis with subspace symmetries. Computer Graphics Forum, 30 (2):277–286, April 2011. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Batty:2010:FST

[BXH10] Christopher Batty, Stefan Xenos, and Ben Houston. Fluids and smoke: Tetrahedral embedded boundary methods for accurate and flexible adaptive fluids. Computer Graphics Forum, 29 (2):695–704, May 2010. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Bai:2016:CAD

[BYL16] Yunfei Bai, Wenhao Yu, and C. Karen Liu. Cloth & animation: Dexterous manipulation of cloth. Computer Graphics Forum, 35(2):523–532, May 2016. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Back:2018:FGA

[BYM18] Jonghee Back, Sung-Eui Yoon, and Bochang Moon. Feature generation for adaptive gradient-domain path tracing. Computer Graphics Forum, 37(7):65–74, October 2018. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Ben-Zvi:2016:LDV

[BZBM^{+}16] N. Ben-Zvi, J. Bento, M. Mahler, J. Hodgins, and A. Shamir. Line-drawing video stylization. Computer Graphics Forum, 35 (6):18–32, September 2016. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Bernard:2018:TUC

[BZL^{+}18] Jürgen Bernard, Matthias Zeppelzauer, Markus Lehmann, Martin Müller, and Michael Sedlmair. Towards user-centered active learning algorithms. Computer Graphics Forum, 37(3): 121–132, June 2018. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Chung:2016:CGH

[CAB^{+}16] David H. S. Chung, Daniel Archambault, Rita Borgo, Darren J. Edwards, Robert S. Laramee, and Min Chen. Charts and glyphs: How ordered is it? on the perceptual orderability of visual channels. Computer Graphics Forum, 35(3):131–140, June 2016. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Comino:2018:SAN

[CACB18] M. Comino, C. Andujar, A. Chica, and P. Brunet. Sensor-aware normal estimation for point clouds from 3D range scans. Computer Graphics Forum, 37(5):233–243, August 2018. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Campen:2017:PSQ

[Cam17] M. Campen. Partitioning surfaces into quadrilateral patches: A survey. Computer Graphics Forum, 36(8):567–588, December 2017. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Cani:2011:AWE

[Can11] Marie-Paule Cani. Award winners: Eurographics outstanding technical contribution award. Computer Graphics Forum, 30 (2):xviii, April 2011. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Cashman:2012:BCC

[Cas12] Thomas J. Cashman. Beyond Catmull–Clark? A survey of advances in subdivision surface methods. Computer Graphics Forum, 31(1):42–61, February 2012. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Cherchi:2019:SPP

[CAS^{+}19] G. Cherchi, P. Alliez, R. Scateni, M. Lyon, and D. Bommes. Selective padding for polycube-based hexahedral meshing. Computer Graphics Forum, 38(1):580–591, February 2019. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Chen:2018:E

[CB18] Min Chen and Bedrich Benes. Editorial. Computer Graphics Forum, 37(1):5–6, February 2018. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Chen:2019:X

[CB19] Min Chen and Bedrich Benes. 2019 editorial v2. Computer Graphics Forum, 38(1):5–6, February 2019. CODEN CG-FODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Cano:2015:GVM

[CBC^{+}15] R. G. Cano, K. Buchin, T. Castermans, A. Pieterse, W. Sonke, and B. Speckmann. Geospatial visualization: Mosaic drawings and cartograms. Computer Graphics Forum, 34(3):361–370, June 2015. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Cordonnier:2016:TFL

[CBC^{+}16] Guillaume Cordonnier, Jean Braun, Marie-Paule Cani, Bedrich Benes, Éric Galin, Adrien Peytavie, and Éric Guérin. Terrains & fluids: Large scale terrain generation from tectonic uplift and fluvial erosion. Computer Graphics Forum, 35(2):165–175, May 2016. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Chandrasegaran:2017:IVA

[CBK^{+}17] Senthil Chandrasegaran, Sriram Karthik Badam, Lorraine Kisselburgh, Karthik Ramani, and Niklas Elmqvist. Integrating visual analytics support for grounded theory practice in qualitative text analysis. Computer Graphics Forum, 36(3):201–212, June 2017. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Cornel:2019:IVF

[CBKK^{+}19] D. Cornel, A. Buttinger-Kreuzhuber, A. Konev, Z. Horváth, M. Wimmer, R. Heidrich, and J. Waser. Interactive visualization of flood and heavy rain simulations. Computer Graphics Forum, 38(3):25–39, June 2019. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Chu:2019:SFT

[CBLW19] Lei Chu, Pengbo Bo, Yang Liu, and Wenping Wang. Surface fairing towards regular principal curvature line networks. Computer Graphics Forum, 38(7):115–125, October 2019. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Claici:2017:MPI

[CBSS17] S. Claici, M. Bessmeltsev, S. Schaefer, and J. Solomon. Maps and parametrization: Isometry-aware preconditioning for mesh parameterization. Computer Graphics Forum, 36(5):37–47, August 2017. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Chiang:2016:RTP

[CBTB16] Matt Jen-Yuan Chiang, Benedikt Bitterli, Chuck Tappan, and Brent Burley. Rendering techniques: a practical and controllable hair and fur model for production path tracing. Computer Graphics Forum, 35(2):275–283, May 2016. CODEN CG-FODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Claux:2014:SIC

[CBV^{+}14] Frédéric Claux, Loïc Barthe, David Vanderhaeghe, Jean-Pierre Jessel, and Mathias Paulin. Surfaces II: Crack-free rendering of dynamically tesselated B-rep models. Computer Graphics Forum, 33(2):263–272, May 2014. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Charalambous:2014:PCG

[CC14] P. Charalambous and Y. Chrysanthou. The PAG crowd: a graph based approach for efficient data-driven crowd simulation. Computer Graphics Forum, 33(8):95–108, December 2014. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Chen:2014:RDM

[CCC^{+}14] Shao-Chi Chen, Hsin-Yi Chen, Yi-Ling Chen, Hsin-Mu Tsai, and Bing-Yu Chen. Reconstruction and depth: Making infront-of cars transparent: Sharing first-person-views via dashcam. Computer Graphics Forum, 33(7):289–297, October 2014. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Chen:2017:QIU

[CCC17] Y.-Y. Chen, Y.-J. Chen, and S.-Y. Chien. Quantizing intersections using compact voxels. Computer Graphics Forum, 36(6): 76–85, September 2017. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Conati:2014:DSE

[CCH^{+}14] C. Conati, G. Carenini, E. Hoque, B. Steichen, and D. Toker. Design study: Evaluating the impact of user characteristics and different layouts on an interactive visualization for decision making. Computer Graphics Forum, 33(3):371–380, June 2014. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Choi:2013:RSI

[CCI13] B. Choi, B. Chang, and I. Ihm. Rendering (session 2): Improving memory space efficiency of kd-tree for real-time ray tracing. Computer Graphics Forum, 32(7):335–344, October 2013. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Chen:2010:GBO

[CCLN10] Yi-Ling Chen, Bing-Yu Chen, Shang-Hong Lai, and Tomoyuki Nishita. Geometry I: Binary orientation trees for volume and surface reconstruction from unoriented point clouds. Computer Graphics Forum, 29(7):2011–2019, September 2010. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Chang:2016:VIV

[CCM16] Chia-Sheng Chang, Hung-Kuo Chu, and Niloy J. Mitra. Video: Interactive videos: Plausible video editing using sparse structure points. Computer Graphics Forum, 35(2):489–500, May 2016. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Chermain:2019:GRB

[CCM19] Xavier Chermain, Frédéric Claux, and Stéphane Mérillou. Glint rendering based on a multiple-scattering patch BRDF. Computer Graphics Forum, 38(4):27–37, July 2019. CODEN CG-FODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Cho:2012:IEP

[CCTL12] Sunghyun Cho, Hojin Cho, Yu-Wing Tai, and Seungyong Lee. Image editing and processing: Registration based non-uniform motion deblurring. Computer Graphics Forum, 31(7pt2):2183– 2192, September 2012. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Chen:2012:GPI

[CCW12] Zhonggui Chen, Juan Cao, and Wenping Wang. Geometry processing: Isotropic surface remeshing using constrained centroidal Delaunay mesh. Computer Graphics Forum, 31(7pt1): 2077–2085, September 2012. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Cazals:2010:RIM

[CD10] Frédéric Cazals and Tom Dreyfus. Reconstruction II: Multiscale geometric modeling of ambiguous shapes with toleranced balls and compoundly weighted α-shapes. Computer Graphics Forum, 29(5):1713–1722, July 2010. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Cavallo:2018:TXS

[CD18] Marco Cavallo and Çağatay Demiralp. Track Xplorer: a system for visual analysis of sensor-based motor activity predictions. Computer Graphics Forum, 37(3):339–349, June 2018. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Chaturvedi:2014:GBM

[CDA^{+}14] S. Chaturvedi, C. Dunne, Z. Ashktorab, R. Zachariah, and B. Shneiderman. Group-in-a-Box meta-layouts for topological clusters and attribute-based groups: Space-efficient visualizations of network communities and their ties. Computer Graphics Forum, 33(8):52–68, December 2014. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Ceylan:2017:DSV

[CDM^{+}17] Duygu Ceylan, Minh Dang, Niloy J. Mitra, Boris Neubert, and Mark Pauly. Discovering structured variations via template matching. Computer Graphics Forum, 36(1):76–88, January 2017. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Crane:2010:DGT

[CDS10] Keenan Crane, Mathieu Desbrun, and Peter Schröder. Differential geometry: Trivial connections on discrete surfaces. Computer Graphics Forum, 29(5):1525–1533, July 2010. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Cioaca:2016:GBW

[CDS16] Teodor Cioaca, Bogdan Dumitrescu, and Mihai-Sorin Stupariu. Graph-based wavelet representation of multi-variate terrain data. Computer Graphics Forum, 35(1):44–58, February 2016. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Cazals:2014:GGA

[CDSS14] F. Cazals, T. Dreyfus, S. Sachdeva, and N. Shah. Greedy geometric algorithms for collection of balls, with applications to geometric approximation and molecular coarse-graining. Computer Graphics Forum, 33(6):1–17, September 2014. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Chen:2019:OFS

[CE19] Min Chen and David S. Ebert. An ontological framework for supporting the design and evaluation of visual analytics systems. Computer Graphics Forum, 38(3):131–144, June 2019. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Cordonnier:2018:IGT

[CEG^{+}18] G. Cordonnier, P. Ecormier, E. Galin, J. Gain, B. Benes, and M.-P. Cani. Interactive generation of time-evolving, snowcovered landscapes with avalanches. Computer Graphics Forum, 37(2):497–509, May 2018. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Cetinaslan:2019:PBS

[Cet19] Ozan Cetinaslan. Position-based simulation of elastic models on the GPU with energy aware Gauss–Seidel algorithm. Computer Graphics Forum, 38(8):41–52, November 2019. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Cai:2018:SED

[CEX^{+}18] X. Cai, K. Efstathiou, X. Xie, Y. Wu, Y. Shi, and L. Yu. A study of the effect of doughnut chart parameters on proportion estimation accuracy. Computer Graphics Forum, 37(6):300– 312, September 2018. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Chen:2016:MSM

[CFB16] Xue Chen, Jieqing Feng, and Dominique Bechmann. Mesh sequence morphing. Computer Graphics Forum, 35(1):179–190, February 2016. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Coeurjolly:2016:RPS

[CFGL16] David Coeurjolly, Marion Foare, Pierre Gueth, and Jacques-Olivier Lachaud. Reconstruction: Piecewise smooth reconstruction of normal vector field on digital data. Computer Graphics Forum, 35(7):157–167, October 2016. CODEN CG-FODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Civit-Flores:2014:RTD

[CFS14] O. Civit-Flores and A. Susín. Robust treatment of degenerate elements in interactive corotational FEM simulations. Computer Graphics Forum, 33(6):298–309, September 2014. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Chen:2012:PSP

[CG12] Renjie Chen and Craig Gotsman. Point sets: Parallel blue-noise sampling by constrained farthest point optimization. Computer Graphics Forum, 31(5):1775–1785, August 2012. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Cai:2016:IPA

[CG16a] Yiqi Cai and Xiaohu Guo. Image processing: Anisotropic superpixel generation based on Mahalanobis distance. Computer Graphics Forum, 35(7):199–207, October 2016. CODEN CG-FODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Chen:2016:IEP

[CG16b] Renjie Chen and Craig Gotsman. Image editing & processing: Generalized as-similar-as-possible warping with applications in digital photography. Computer Graphics Forum, 35(2):81–92, May 2016. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Chen:2016:MCT

[CG16c] Renjie Chen and Craig Gotsman. Mappings: Complex transfinite barycentric mappings with similarity kernels. Computer Graphics Forum, 35(5):41–53, August 2016. CODEN CG-FODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Chen:2017:APC

[CG17] Renjie Chen and Craig Gotsman. Approximating planar conformal maps using regular polygonal meshes. Computer Graphics Forum, 36(8):629–642, December 2017. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Chen:2019:GAT

[CG19] Jia Chen and M. Gopi. Geometry aware tori decomposition. Computer Graphics Forum, 38(2):331–341, May 2019. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Chen:2013:NOM

[CGBG13] Jiazhou Chen, Gaël Guennebaud, Pascal Barla, and Xavier Granier. Non-oriented MLS gradient fields. Computer Graphics Forum, 32(8):98–109, December 2013. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Chen:2018:EPG

[CGH18] Renjie Chen, Craig Gotsman, and Kai Hormann. Efficient path generation with reduced coordinates. Computer Graphics Forum, 37(5):37–48, August 2018. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Ceneda:2019:RGA

[CGM19] Davide Ceneda, Theresia Gschwandtner, and Silvia Miksch. A review of guidance approaches in visual data analysis: a multifocal perspective. Computer Graphics Forum, 38(3):861–879, June 2019. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Ciccone:2016:MFC

[CGS16] L. Ciccone, M. Guay, and R. Sumner. Modeling: Flow curves: an intuitive interface for coherent scene deformation. Computer Graphics Forum, 35(7):247–256, October 2016. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Carr:2015:MMM

[CGT^{+}15] Hamish Carr, Zhao Geng, Julien Tierny, Amit Chattopadhyay, and Aaron Knoll. Multi-modal and multi-field: Fiber surfaces: Generalizing isosurfaces to bivariate data. Computer Graphics Forum, 34(3):241–250, June 2015. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Courbet:2011:TPM

[CH11] Clément Courbet and Céline Hudelot. Taylor prediction for mesh geometry compression. Computer Graphics Forum, 30 (1):139–151, March 2011. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Cashman:2012:ACE

[CH12] Thomas J. Cashman and Kai Hormann. Animation: a continuous, editable representation for deforming mesh sequences with separate signals for time, pose and shape. Computer Graphics Forum, 31(2pt4):735–744, May 2012. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Chapiro:2014:IOS

[CHA^{+}14] Alexandre Chapiro, Simon Heinzle, Tunç Ozan Aydın, Steven Poulakos, Matthias Zwicker, Aljosa Smolic, and Markus Gross. Images I: Optimizing stereo-to-multiview conversion for autostereoscopic displays. Computer Graphics Forum, 33(2):63– 72, May 2014. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Cashman:2019:UBV

[CHH^{+}19] Dylan Cashman, Shah Rukh Humayoun, Florian Heimerl, Kendall Park, Subhajit Das, John Thompson, Bahador Saket, Abigail Mosca, John Stasko, Alex Endert, Michael Gleicher, and Remco Chang. A user-based visual analytics workflow for exploratory model analysis. Computer Graphics Forum, 38 (3):185–199, June 2019. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Campen:2013:SPA

[CHK13] Marcel Campen, Martin Heistermann, and Leif Kobbelt. Surfaces: Practical anisotropic geodesy. Computer Graphics Forum, 32(5):63–71, August 2013. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Cadik:2013:IPS

[ČHM^{+}13] Martin Čadík, Robert Herzog, Rafał Mantiuk, Radosław Mantiuk, Karol Myszkowski, and Hans-Peter Seidel. Image processing (session 6): Learning to predict localized distortions in rendered images. Computer Graphics Forum, 32(7):401–410, October 2013. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Chu:2019:BRT

[CHWL19] Yiyao Chu, Fei Hou, Wencheng Wang, and Lei Li. Book review: Topology Preserving Simplification of Medial Axes in 3D Models. Computer Graphics Forum, 38(7):607–616, October 2019. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Campen:2016:PVS

[CIE^{+}16] Marcel Campen, Moritz Ibing, Hans-Christian Ebke, Denis Zorin, and Leif Kobbelt. Parametrization and volumes: Scaleinvariant directional alignment of surface parametrizations. Computer Graphics Forum, 35(5):1–10, August 2016. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Cornelis:2014:APB

[CIPT14] Jens Cornelis, Markus Ihmsen, Andreas Peer, and Matthias Teschner. Animation I (physically-based): IISPH–FLIP for incompressible fluids. Computer Graphics Forum, 33(2):255–262, May 2014. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Cao:2014:SCI

[CJFH14] Yan-Pei Cao, Tao Ju, Zhao Fu, and Shi-Min Hu. Shapes and cryptography: Interactive image-guided modeling of extruded shapes. Computer Graphics Forum, 33(7):101–110, October 2014. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Choi:2019:VNV

[CJP^{+}19] Jinho Choi, Sanghun Jung, Deok Gun Park, Jaegul Choo, and Niklas Elmqvist. Visualizing for the non-visual: Enabling the visually impaired to use visualization. Computer Graphics Forum, 38(3):249–260, June 2019. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Celarek:2019:QEL

[CJWL19] A. Celarek, W. Jakob, M. Wimmer, and J. Lehtinen. Quantifying the error of light transport algorithms. Computer Graphics Forum, 38(4):111–121, July 2019. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Cao:2017:ESF

[CJXH17] Y-P. Cao, T. Ju, J. Xu, and S-M. Hu. Extracting sharp features from RGB–D images. Computer Graphics Forum, 36(8): 138–152, December 2017. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Chen:2012:IAE

[CJZW12] Xiaowu Chen, Xin Jin, Qinping Zhao, and Hongyu Wu. Image analysis and editing: Artistic illumination transfer for portraits. Computer Graphics Forum, 31(4):1425–1434, June 2012. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Campen:2010:CPB

[CK10a] Marcel Campen and Leif Kobbelt. Collisions: Polygonal boundary evaluation of Minkowski sums and swept volumes. Computer Graphics Forum, 29(5):1613–1622, July 2010. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Campen:2010:MPE

[CK10b] Marcel Campen and Leif Kobbelt. Mesh processing: Exact and robust (self-)intersections for polygonal meshes. Computer Graphics Forum, 29(2):397–406, May 2010. CODEN CG-FODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Campen:2011:IGP

[CK11a] Marcel Campen and Leif Kobbelt. Interactive geometric processing: Walking on broken mesh: Defect-tolerant geodesic distances and parameterizations. Computer Graphics Forum, 30 (2):623–632, April 2011. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Chuang:2011:FMC

[CK11b] Ming Chuang and Michael Kazhdan. Fast mean-curvature flow via finite-elements tracking. Computer Graphics Forum, 30(6): 1750–1760, September 2011. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Cho:2013:FGA

[CK13] Junghyun Cho and Hyeong-Seok Ko. Fluids: Geometry-aware volume-of-fluid method. Computer Graphics Forum, 32(2pt3): 379–388, May 2013. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Campen:2014:QLE

[CK14] M. Campen and L. Kobbelt. Quad layout embedding via aligned parameterization. Computer Graphics Forum, 33(8): 69–81, December 2014. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Coffey:2012:IDU

[CKE^{+}12] D. Coffey, F. Korsakov, M. Ewert, H. Hagh-Shenas, L. Thorson, A. Ellingson, D. Nuckley, and D. F. Keefe. Interaction design and usability: Visualizing motion data in virtual reality: Understanding the roles of animation, interaction, and static presentation. Computer Graphics Forum, 31(3pt3):1215–1224, June 2012. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Charalambous:2014:CAD

[CKGC14] Panayiotis Charalambous, Ioannis Karamouzas, Stephen J. Guy, and Yiorgos Chrysanthou. Character animation: a datadriven framework for visual crowd analysis. Computer Graphics Forum, 33(7):41–50, October 2014. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Conlen:2019:CAA

[CKH19] Matthew Conlen, Alex Kale, and Jeffrey Heer. Capture & analysis of active reading behaviors for interactive articles on the Web. Computer Graphics Forum, 38(3):687–698, June 2019. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Choi:2011:MCS

[CKHL11] Myung Geol Choi, Manmyung Kim, Kyung Lyul Hyun, and Jehee Lee. Motion capture, simulation, and manipulation: Deformable motion: Squeezing into cluttered environments. Computer Graphics Forum, 30(2):445–453, April 2011. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Christensen:2018:PMJ

[CKK18] Per Christensen, Andrew Kensler, and Charlie Kilpatrick. Progressive multi-jittered sample sequences. Computer Graphics Forum, 37(4):21–33, July 2018. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Cho:2014:CIS

[CKL14] Hojin Cho, Seon Joo Kim, and Seungyong Lee. Color and imaging: Single-shot high dynamic range imaging using coded electronic shutter. Computer Graphics Forum, 33(7):329–338, October 2014. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Cornel:2015:GVV

[CKS^{+}15] D. Cornel, A. Konev, B. Sadransky, Z. Horváth, E. Gröller, and J. Waser. Geospatial visualization: Visualization of objectcentered vulnerability to possible flood hazards. Computer Graphics Forum, 34(3):331–340, June 2015. CODEN CG-FODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Cornel:2016:GDV

[CKS^{+}16] D. Cornel, A. Konev, B. Sadransky, Z. Horváth, A. Brambilla, I. Viola, and J. Waser. Geospatial data visualization: Composite flow maps. Computer Graphics Forum, 35(3):461–470, June 2016. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Chen:2012:STF

[CLC12] Ying-Chieh Chen, Su Ian Eugene Lei, and Chun-Fa Chang. Spatio-temporal filtering of indirect lighting for interactive global illumination. Computer Graphics Forum, 31(1):189–201, February 2012. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Chen:2011:MRA

[CLCL11] Yi-Ling Chen, Tung-Ying Lee, Bing-Yu Chen, and Shang-Hong Lai. Model reconstruction & analysis: Bipartite polar classification for surface reconstruction. Computer Graphics Forum, 30(7):2003–2010, September 2011. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Calian:2018:FOL

[CLG^{+}18] Dan A. Calian, Jean-François Lalonde, Paulo Gotardo, Tomas Simon, Iain Matthews, and Kenny Mitchell. From faces to outdoor light probes. Computer Graphics Forum, 37(2):51–61, May 2018. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Cui:2018:FDY

[CLGS18] Y. R. Cui, Q. Liu, C. Y. Gao, and Z. Su. FashionGAN: Display your fashion design using Conditional Generative Adversarial Nets. Computer Graphics Forum, 37(7):109–119, October 2018. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Chen:2015:DPD

[CLJ^{+}15] Hsin-I Chen, Tse-Ju Lin, Xiao-Feng Jian, I-Chao Shen, and Bing-Yu Chen. Drawing and painting: Data-driven handwriting synthesis in a conjoined manner. Computer Graphics Forum, 34(7):235–244, October 2015. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Corsini:2013:PMS

[CLL^{+}13] M. Corsini, M. C. Larabi, G. Lavoué, O. Petřík, L. Váša, and K. Wang. Perceptual metrics for static and dynamic triangle meshes. Computer Graphics Forum, 32(1):101–125, February 2013. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Chiu:2015:DPT

[CLLC15] Chun-Chia Chiu, Yi-Hsiang Lo, Ruen-Rone Lee, and Hung-Kuo Chu. Drawing and painting: Tone- and feature-aware circular scribble art. Computer Graphics Forum, 34(7):225–234, October 2015. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Cherchi:2016:PVP

[CLS16] Gianmarco Cherchi, Marco Livesu, and Riccardo Scateni. Parametrization and volumes: Polycube simplification for coarse layouts of surfaces and volumes. Computer Graphics Forum, 35(5):11–20, August 2016. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Crnovrsanin:2011:GNV

[CLWM11] Tarik Crnovrsanin, Isaac Liao, Yingcai Wu, and Kwan-Liu Ma. Graphs & networks: Visual recommendations for network navigation. Computer Graphics Forum, 30(3):1081–1090, June 2011. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Chen:2017:SMV

[CLY17] Siming Chen, Lijing Lin, and Xiaoru Yuan. Social media visual analytics. Computer Graphics Forum, 36(3):563–587, June 2017. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Clausen:2018:AVS

[CMF18] O. Clausen, R. Marroquim, and A. Fuhrmann. Acquisition and validation of spectral ground truth data for predictive rendering of rough surfaces. Computer Graphics Forum, 37(4):1–12, July 2018. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Ceylan:2012:CBF

[CML^{+}12] Duygu Ceylan, Niloy J. Mitra, Hao Li, Thibaut Weise, and Mark Pauly. Cities and buildings: Factored facade acquisition using symmetric line arrangements. Computer Graphics Forum, 31(2pt3):671–680, May 2012. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Cai:2015:VRR

[CNCO15] Li-Le Cai, Binh P. Nguyen, Chee-Kong Chui, and Sim-Heng Ong. Volume rendering: Rule-enhanced transfer function generation for medical volume visualization. Computer Graphics Forum, 34(3):121–130, June 2015. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Choi:2013:IPS

[CNKI13] Myung Geol Choi, Seung-Tak Noh, Taku Komura, and Takeo Igarashi. Image processing (session 1): Dynamic comics for hierarchical abstraction of 3D animation data. Computer Graphics Forum, 32(7):1–9, October 2013. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Crassin:2011:LRI

[CNS^{+}11] Cyril Crassin, Fabrice Neyret, Miguel Sainz, Simon Green, and Elmar Eisemann. Lighting & rendering: Interactive indirect illumination using voxel cone tracing. Computer Graphics Forum, 30(7):1921–1930, September 2011. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Corman:2015:CMC

[COC15] Etienne Corman, Maks Ovsjanikov, and Antonin Chambolle. Correspondence and matching: Continuous matching via vector field flow. Computer Graphics Forum, 34(5):129–139, August 2015. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Carriere:2015:DSS

[COO15] Mathieu Carrière, Steve Y. Oudot, and Maks Ovsjanikov. Descriptors and shape synthesis: Stable topological signatures for points on 3D shapes. Computer Graphics Forum, 34(5):1–12, August 2015. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Coll:2010:ASM

[CP10] N. Coll and T. Paradinas. Accurate simplification of multi-chart textured models. Computer Graphics Forum, 29(6):1842–1853, September 2010. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Coll:2011:CM

[CP11] Narcís Coll and Teresa Paradinas. Compact models. Computer Graphics Forum, 30(1):187–198, March 2011. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Cortial:2019:PTP

[CPGG19] Y. Cortial, A. Peytavie, E. Galin, and E. Guérin. Procedural tectonic planets. Computer Graphics Forum, 38(2):1–11, May 2019. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Cheng:2015:IVD

[CPZ^{+}15] M. M. Cheng, V. A. Prisacariu, S. Zheng, P. H. S. Torr, and C. Rother. Image and video: DenseCut: Densely connected CRFs for realtime GrabCut. Computer Graphics Forum, 34 (7):193–201, October 2015. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Chen:2016:E

[CR16a] Min Chen and Richard. Editorial. Computer Graphics Forum, 35(1):3, February 2016. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Csebfalvi:2016:VDV

[CR16b] Balázs Csébfalvi and Gergely Rácz. Volume data visualization: Retailoring box splines to lattices for highly isotropic volume representations. Computer Graphics Forum, 35(3):411– 420, June 2016. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Chen:2017:E

[CR17] Min Chen and Richard. Editorial. Computer Graphics Forum, 36(1):5, January 2017. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Cosmo:2017:CPM

[CRA^{+}17] L. Cosmo, E. Rodolà, A. Albarelli, F. Mémoli, and D. Cremers. Consistent partial matching of shape collections via sparse modeling. Computer Graphics Forum, 36(1):209–221, January 2017. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Casas:2015:IVM

[CRC^{+}15] Dan Casas, Christian Richardt, John Collomosse, Christian Theobalt, and Adrian Hilton. Image and video: 4D model flow: Precomputed appearance alignment for real-time 4D video interpolation. Computer Graphics Forum, 34(7):173–182, October 2015. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Cao:2010:SRI

[CRGZ10] Chen Cao, Zhong Ren, Baining Guo, and Kun Zhou. Scattering and refraction: Interactive rendering of non-constant, refractive media using the ray equations of gradient-index optics. Computer Graphics Forum, 29(4):1375–1382, June 2010. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Caban:2011:EEV

[CRY11] J. J. Caban, P. Rheingans, and T. Yoo. Evaluation: an evaluation of visualization techniques to illustrate statistical deformation models. Computer Graphics Forum, 30(3):821–830, June 2011. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Cardona:2016:VNT

[CS16] L. Cardona and S. Saito. Visualization/NPR: Temporally coherent and artistically intended stylization of feature lines extracted from 3D models. Computer Graphics Forum, 35(7): 137–146, October 2016. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Chen:2013:RCF

[CSaLM13] Desai Chen, Pitchaya Sitthi-amorn, Justin T. Lan, and Wojciech Matusik. Reality: Computing and fabricating multiplanar models. Computer Graphics Forum, 32(2pt3):305–315, May 2013. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Cheng:2018:LSI

[CSC^{+}18] Dachuan Cheng, Jian Shi, Yanyun Chen, Xiaoming Deng, and Xiaopeng. Zhang. Learning scene illumination by pairwise photos from rear and front mobile cameras. Computer Graphics Forum, 37(7):213–221, October 2018. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Chaurasia:2011:GRS

[CSD11] Gaurav Chaurasia, Olga Sorkine, and George Drettakis. Geometry for rendering: Silhouette-aware warping for image-based rendering. Computer Graphics Forum, 30(4):1223–1232, June 2011. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Csebfalvi:2018:FCR

[Csé18] Balázs Csébfalvi. Fast Catmull–Rom spline interpolation for high-quality texture sampling. Computer Graphics Forum, 37 (2):455–462, May 2018. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Carnecky:2012:IVM

[CSFP12] R. Carnecky, B. Schindler, R. Fuchs, and R. Peikert. Illustrative visualization: Multi-layer illustrative dense flow visualization. Computer Graphics Forum, 31(3pt1):895–904, June 2012. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Chegini:2018:IVE

[CSG^{+}18] Mohammad Chegini, Lin Shao, Robert Gregor, Dirk J. Lehmann, Keith Andrews, and Tobias Schreck. Interactive visual exploration of local patterns in large scatterplot spaces. Computer Graphics Forum, 37(3):99–109, June 2018. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Carr:2010:RSB

[CSLG10] Nathan Carr, Faramarz Savamati, Joe Laviola, and Cindy Grimm. Reports: Sketch-Based Interfaces and Modeling 2009 Co-Sponsored by Eurographics and ACM SIGGRAPH New Orleans, Louisiana, August 1–2, 2009. Computer Graphics Forum, 29(1):249, March 2010. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Chlumsky:2018:ICM

[CSŠ18] V. Chlumský, J. Sloup, and I. Šimeček. Improved corners with multi-channel signed distance fields. Computer Graphics Forum, 37(1):273–287, February 2018. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Calakli:2011:MRA

[CT11] F. Calakli and G. Taubin. Model reconstruction & analysis: SSD: Smooth signed distance surface reconstruction. Computer Graphics Forum, 30(7):1993–2002, September 2011. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Clarberg:2010:CGS

[CTHAM10] Petrik Clarberg, Robert Toth, Jon Hasselgren, and Tomas Akenine-Möller. Computer graphics systems: an optimizing compiler for automatic shader bounding. Computer Graphics Forum, 29(4):1259–1268, June 2010. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Chan:2013:ASS

[CTL13] Jacky C. P. Chan, Jeff K. T. Tang, and Howard Leung. Animation (session 1): Synthesizing two-character interactions by merging captured interaction samples with their spacetime relationships. Computer Graphics Forum, 32(7):41–50, October 2013. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Casas:2014:IIV

[CVCH14] Dan Casas, Marco Volino, John Collomosse, and Adrian Hilton. Images III: 4D video textures for interactive character appearance. Computer Graphics Forum, 33(2):371–380, May 2014. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Caillaud:2016:GPC

[CVDL16] F. Caillaud, V. Vidal, F. Dupont, and G. Lavoué. Geometry: Progressive compression of arbitrary textured meshes. Computer Graphics Forum, 35(7):475–484, October 2016. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Chambers:2015:CMA

[CVJ15] Erin Wolf Chambers and Mikael Vejdemo-Johansson. Computing minimum area homologies. Computer Graphics Forum, 34 (6):13–21, September 2015. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Chen:2014:CPV

[CWB^{+}14] M. Chen, S. Walton, K. Berger, J. Thiyagalingam, B. Duffy, H. Fang, C. Holloway, and A. E. Trefethen. Cognition and perception: Visual multiplexing. Computer Graphics Forum, 33(3):241–250, June 2014. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Correll:2011:TDV

[CWG11] M. Correll, M. Witmore, and M. Gleicher. Text and document visualization: Exploring collections of tagged text for literary scholarship. Computer Graphics Forum, 30(3):731–740, June 2011. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Caballero:2019:VAV

[CWGvW19] Humberto S. Garcia Caballero, Michel A. Westenberg, Binyam Gebre, and Jarke J. van Wijk. V-Awake: a visual analytics approach for correcting sleep predictions from deep learning models. Computer Graphics Forum, 38(3):1–12, June 2019. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Chen:2015:TIV

[CWL^{+}15] Yi-Cheng Chen, Yu-Shuen Wang, Wen-Chieh Lin, Wei-Xiang Huang, and I-Chen Lin. Traffic: Interactive visual analysis for vehicle detector data. Computer Graphics Forum, 34(3):171– 180, June 2015. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Chen:2019:DVB

[CWL19] Mingjia Chen, Changbo Wang, and Ligang Liu. Deep videobased performance synthesis from sparse multi-view capture. Computer Graphics Forum, 38(7):543–554, October 2019. CO-DEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Chou:2019:PPV

[CWM19] Jia-Kai Chou, Yang Wang, and Kwan-Liu Ma. Privacy preserving visualization: a study on event sequence data. Computer Graphics Forum, 38(1):340–355, February 2019. CODEN CG-FODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Cypko:2017:VVC

[CWS^{+}17] M. A. Cypko, J. Wojdziak, M. Stoehr, B. Kirchner, B. Preim, A. Dietz, H. U. Lemke, and S. Oeltze-Jafra. Visual verification of cancer staging for therapy decision support. Computer Graphics Forum, 36(3):109–120, June 2017. CODEN CG-FODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Chen:2011:EDF

[CWW^{+}11] Jiating Chen, Bin Wang, Yuxiang Wang, Ryan S. Overbeck, Jun-Hai Yong, and Wenping Wang. Efficient depth-of-field rendering with adaptive sampling and multiscale reconstruction. Computer Graphics Forum, 30(6):1667–1680, September 2011. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Chen:2011:SIS

[CWY11] Jiating Chen, Bin Wang, and Jun-Hai Yong. Sampling: Improved stochastic progressive photon mapping with Metropolis sampling. Computer Graphics Forum, 30(4):1205–1213, June 2011. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Chou:2011:TDV

[CY11] J.-K. Chou and C.-K. Yang. Text and document visualization: PaperVis: Literature review made easy. Computer Graphics Forum, 30(3):721–730, June 2011. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Claudio:2014:GVM

[CY14] Pio Claudio and Sung-Eui Yoon. Geographic visualization: Metro transit-centric visualization for city tour planning. Computer Graphics Forum, 33(3):271–280, June 2014. CODEN CG-FODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Celikcan:2015:EBR

[CYC15] Ufuk Celikcan, Ilker O. Yaz, and Tolga Capin. Example-based retargeting of human motion to arbitrary mesh models. Computer Graphics Forum, 34(1):216–227, February 2015. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Choi:2012:AIR

[CYI^{+}12] M. G. Choi, K. Yang, T. Igarashi, J. Mitani, and J. Lee. Animation and interaction: Retrieval and visualization of human motion data via stick figures. Computer Graphics Forum, 31 (7pt1):2057–2065, September 2012. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Chen:2011:NPR

[CYY^{+}11] Cheng-Kai Chen, Shi Yan, Hongfeng Yu, Nelson Max, and Kwan-Liu Ma. Non-photorealistic rendering & GPU computing: An illustrative visualization framework for 3D vector fields. Computer Graphics Forum, 30(7):1941–1951, September 2011. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Chen:2011:MCS

[CZY11] Fan Chen, Ye Zhao, and Zhi Yuan. Motion capture, simulation, and manipulation: Langevin particle: a self-adaptive Lagrangian primitive for flow simulation enhancement. Computer Graphics Forum, 30(2):435–444, April 2011. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Chen:2018:PBF

[CZZ^{+}18] Wei Chen, Fei Zhu, Jing Zhao, Sheng Li, and Guoping Wang. Peridynamics-based fracture animation for elastoplastic solids. Computer Graphics Forum, 37(1):112–124, February 2018. CO-DEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Deschaintre:2019:FSC

[DAD^{+}19] Valentin Deschaintre, Miika Aittala, Fredo Durand, George Drettakis, and Adrien Bousseau. Flexible SVBRDF capture with a multi-image deep network. Computer Graphics Forum, 38(4):1–13, July 2019. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Dasgupta:2018:HFS

[DAF^{+}18] Aritra Dasgupta, Dustin L. Arendt, Lyndsey R. Franklin, Pak Chung Wong, and Kristin A. Cook. Human factors in streaming data analysis: Challenges and opportunities for information visualization. Computer Graphics Forum, 37(1):254– 272, February 2018. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Dinev:2018:UGL

[DBB^{+}18] D. Dinev, T. Beeler, D. Bradley, M. Bächer, H. Xu, and L. Kavan. User-guided lip correction for facial performance capture. Computer Graphics Forum, 37(8):93–101, December 2018. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Deng:2013:SME

[DBD^{+}13] Bailin Deng, Sofien Bouaziz, Mario Deuss, Juyong Zhang, Yuliy Schwartzburg, and Mark Pauly. Shape modeling: Exploring local modifications for constrained meshes. Computer Graphics Forum, 32(2pt1):11–20, May 2013. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Duvenhage:2011:EBP

[DBK11] B. Duvenhage, K. Bouatouch, and D. G. Kourie. Extending backward polygon beam tracing to glossy scattering surfaces. Computer Graphics Forum, 30(6):1825–1836, September 2011. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Dorn:2015:RTA

[DBLW15] Jonathan Dorn, Connelly Barnes, Jason Lawrence, and Westley Weimer. Rendering: Towards automatic band-limited procedural shaders. Computer Graphics Forum, 34(7):77–87, October 2015. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Drocourt:2011:SDV

[DBS^{+}11] Y. Drocourt, R. Borgo, K. Scharrer, T. Murray, S. I. Bevan, and M. Chen. Spatiotemporal data visualization: Temporal visualization of boundary-based geo-information using radial projection. Computer Graphics Forum, 30(3):981–990, June 2011. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Debattista:2018:FRV

[DBS^{+}18] K. Debattista, K. Bugeja, S. Spina, T. Bashford-Rogers, and V. Hulusic. Frame rate vs resolution: a subjective evaluation of spatiotemporal perceived quality under varying computational budgets. Computer Graphics Forum, 37(1):363–374, February 2018. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Du:2010:VRC

[DC10] Zhiyan Du and Yi-Jen Chiang. Volume rendering: Out-of-core simplification and crack-free LOD volume rendering for irregular grids. Computer Graphics Forum, 29(3):873–882, June 2010. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Darles:2011:SOS

[DCGG11] E. Darles, B. Crespin, D. Ghazanfarpour, and J. C. Gonzato. A survey of ocean simulation and rendering techniques in computer graphics. Computer Graphics Forum, 30(1):43–60, March 2011. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Dasgupta:2012:UCV

[DCK12] Aritra Dasgupta, Min Chen, and Robert Kosara. Uncertainty: Conceptualizing visual uncertainty in parallel coordinates. Computer Graphics Forum, 31(3pt2):1015–1024, June 2012. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Dasgupta:2013:MPU

[DCK13] Aritra Dasgupta, Min Chen, and Robert Kosara. Measuring privacy and utility in privacy-preserving visualization. Computer Graphics Forum, 32(8):35–47, December 2013. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Dang:2014:PMS

[DCNP14] Minh Dang, Duygu Ceylan, Boris Neubert, and Mark Pauly. Procedural Modeling I: SAFE: Structure-aware facade editing. Computer Graphics Forum, 33(2):83–93, May 2014. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Digne:2014:SSS

[DCV14] Julie Digne, Raphaëlle Chaine, and Sébastien Valette. Surfaces I: Self-similarity for accurate compression of point sampled surfaces. Computer Graphics Forum, 33(2):155–164, May 2014. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Dong:2019:SGA

[DCYG19] Xiao Dong, Zhonggui Chen, Junfeng Yao, and Xiaohu Guo. Superpixel generation by agglomerative clustering with quadratic error minimization. Computer Graphics Forum, 38(1):405–416, February 2019. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Doukakis:2018:ARA

[DDH^{+}18] E. Doukakis, K. Debattista, C. Harvey, T. Bashford-Rogers, and A. Chalmers. Audiovisual resource allocation for bimodal virtual environments. Computer Graphics Forum, 37(1):172– 183, February 2018. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Danerek:2017:DGS

[DDÖ^{+}17] R. Daněřek, E. Dibra, C. Öztireli, R. Ziegler, and M. Gross. DeepGarment: 3D garment shape estimation from a single image. Computer Graphics Forum, 36(2):269–280, May 2017. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Debattista:2018:AST

[Deb18] K. Debattista. Application-specific tone mapping via genetic programming. Computer Graphics Forum, 37(1):439–450, February 2018. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Didyk:2010:RIP

[DER^{+}10] Piotr Didyk, Elmar Eisemann, Tobias Ritschel, Karol Myszkowski, and Hans-Peter Seidel. Rendering II: Perceptuallymotivated real-time temporal upsampling of 3D content for high-refresh-rate displays. Computer Graphics Forum, 29(2): 713–722, May 2010. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

dEon:2011:RME

[dFH^{+}11] Eugene d’Eon, Guillaume François, Martin Hill, Joe Letteri, and Jean-Marie Aubry. Reflectance and materials: an energyconserving hair reflectance model. Computer Graphics Forum, 30(4):1181–1187, June 2011. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

DeFloriani:2015:SAR

[DFIM15] Leila De Floriani, Ulderico Fugacci, Federico Iuricich, and Paola Magillo. State of the art reports: Morse complexes for shape segmentation and homological analysis: discrete models and algorithms. Computer Graphics Forum, 34(2):761–785, May 2015. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Du:2014:RSP

[DFY14] Wenjun Du, Jieqing Feng, and Baoguang Yang. Rendering: Sub-pixel anti-aliasing via triangle-based geometry reconstruction. Computer Graphics Forum, 33(7):81–90, October 2014. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Derzapf:2012:DFP

[DG12] E. Derzapf and M. Guthe. Dependency-free parallel progressive meshes. Computer Graphics Forum, 31(8):2288–2302, December 2012. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

deGoes:2011:SOT

[dGCSAD11] Fernando de Goes, David Cohen-Steiner, Pierre Alliez, and Mathieu Desbrun. Section 8: an optimal transport approach to robust reconstruction and simplification of 2D shapes. Computer Graphics Forum, 30(5):1593–1602, August 2011. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Derzapf:2011:PMN

[DGGK11] E. Derzapf, B. Ganster, M. Guthe, and R. Klein. Procedural modeling & natural phenomena: River networks for instant procedural planets. Computer Graphics Forum, 30(7):2031– 2040, September 2011. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

deGoes:2014:CFD

[dGLB^{+}14] Fernando de Goes, Beibei Liu, Max Budninskiy, Yiying Tong, and Mathieu Desbrun. Coordinates and fields: Discrete 2tensor fields on triangulations. Computer Graphics Forum, 33 (5):13–24, August 2014. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Dagenais:2017:DPE

[DGP17] F. Dagenais, J. Gagnon, and E. Paquette. Detail-preserving explicit mesh projection and topology matching for particle-based fluids. Computer Graphics Forum, 36(8):444–457, December 2017. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Dey:2012:PSF

[DGQ^{+}12] T. K. Dey, X. Ge, Q. Que, I. Safa, L. Wang, and Y. Wang. Point sets: Feature-preserving reconstruction of singular surfaces. Computer Graphics Forum, 31(5):1787–1796, August 2012. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

DiBenedetto:2014:MNE

[DGR^{+}14] M. Di Benedetto, F. Ganovelli, M. Balsa Rodriguez, A. Jaspe Villanueva, R. Scopigno, and E. Gobbetti. Men and nature: ExploreMaps: Efficient construction and ubiquitous exploration of panoramic view graphs of complex 3D environments. Computer Graphics Forum, 33(2):459–468, May 2014. CODEN CG-FODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

deGroot:2014:IDI

[dGWB^{+}14] Erwin de Groot, Brian Wyvill, Loïc Barthe, Ahmad Nasri, and Paul Lalonde. Implicit decals: Interactive editing of repetitive patterns on surfaces. Computer Graphics Forum, 33(1): 141–151, February 2014. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Deng:2014:SSP

[DH14] Chongyang Deng and Kai Hormann. Surfaces and shells: Pseudo-spline subdivision surfaces. Computer Graphics Forum, 33(5):227–236, August 2014. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

DiBartolomeo:2016:TSD

[DH16] Marco Di Bartolomeo and Yifan Hu. Time series data and sequences: There is more to streamgraphs than movies: Better aesthetics via ordering and lassoing. Computer Graphics Forum, 35(3):341–350, June 2016. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Dupuy:2015:MAE

[DHI^{+}15] Jonathan Dupuy, Eric Heitz, Jean-Claude Iehl, Pierre Poulin, and Victor Ostromoukhov. Material appearance: Extracting microfacet-based BRDF parameters from arbitrary materials with power iterations. Computer Graphics Forum, 34(4):21– 30, July 2015. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Dokter:2019:HRC

[DHP^{+}19] Mark Dokter, Jozef Hladky, Mathias Parger, Dieter Schmalstieg, Hans-Peter Seidel, and Markus Steinberger. Hierarchical rasterization of curved primitives for vector graphics rendering on the GPU. Computer Graphics Forum, 38(2):93–103, May 2019. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Diepenbrock:2013:MVC

[DHS^{+}13] S. Diepenbrock, S. Hermann, M. Schäfers, M. Kuhlmann, and K. Hinrichs. Medical visualization: Comparative visualization of tracer uptake in in vivo small animal PET/CT imaging of the carotid arteries. Computer Graphics Forum, 32(3pt2):241– 250, June 2013. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

deHoon:2014:BMF

[dHvPJV14] N. de Hoon, R. van Pelt, A. Jalba, and A. Vilanova. BioMedical I: 4D MRI flow coupled to physics-based fluid simulation for blood-flow visualization. Computer Graphics Forum, 33 (3):121–130, June 2014. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Devkota:2018:CDV

[DI18] Sabin Devkota and Katherine E. Isaacs. CFGExplorer: Designing a visual control flow analytics system around basic program analysis operations. Computer Graphics Forum, 37(3):453–464, June 2018. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Du:2012:IEP

[DJM12] Hui Du, Xiaogang Jin, and Xiaoyang Mao. Image editing and processing: Digital camouflage images using two-scale decomposition. Computer Graphics Forum, 31(7pt2):2203–2212, September 2012. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Derouet-Jourdan:2019:GSW

[DJSJ19] Alexandre Derouet-Jourdan, Marc Salvati, and Théo Jonchier. Generating stochastic wall patterns on-the-fly with Wang tiles. Computer Graphics Forum, 38(2):255–264, May 2019. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Dado:2016:TCG

[DKB^{+}16] Bas Dado, Timothy R. Kol, Pablo Bauszat, Jean-Marc Thiery, and Elmar Eisemann. Texturing & compression: Geometry and attribute compression for voxel scenes. Computer Graphics Forum, 35(2):397–407, May 2016. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Dasgupta:2015:GVV

[DKG15] A. Dasgupta, R. Kosara, and L. Gosink. Geospatial visualization: VIMTEX: a V isualization I nterface for M ultivariate, T ime-varying, geological data Exploration. Computer Graphics Forum, 34(3):341–350, June 2015. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Dachsbacher:2014:SRR

[DKH^{+}14] Carsten Dachsbacher, Jaroslav Křivánek, Miloš Hašan, Adam Arbree, Bruce Walter, and Jan Novák. Scalable realistic rendering with many-light methods. Computer Graphics Forum, 33(1):88–104, February 2014. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Dammertz:2010:PPL

[DKL10] H. Dammertz, A. Keller, and H. P. A. Lensch. Progressive point-light-based global illumination. Computer Graphics Forum, 29(8):2504–2515, December 2010. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Didimo:2018:VFU

[DKMT18] Walter Didimo, Evgenios M. Kornaropoulos, Fabrizio Montecchiani, and Ioannis G. Tollis. A visualization framework and user studies for overloaded orthogonal drawings. Computer Graphics Forum, 37(1):288–300, February 2018. CODEN CG-FODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Dabala:2014:IMR

[DKR^{+}14] Ł. Da̧bała, P. Kellnhofer, T. Ritschel, P. Didyk, K. Templin, K. Myszkowski, P. Rokita, and H.-P. Seidel. Images I: Manipulating refractive and reflective binocular disparity. Computer Graphics Forum, 33(2):53–62, May 2014. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Deul:2018:DPB

[DKWB18] Crispin Deul, Tassilo Kugelstadt, Marcel Weiler, and Jan Bender. Direct position-based solver for stiff rods. Computer Graphics Forum, 37(6):313–324, September 2018. CODEN CG-FODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Du:2016:RTA

[DKY16] P. Du, Y. J. Kim, and S. E. Yoon. Ray tracing/appearance capture: TSS BVHs: Tetrahedron swept sphere BVHs for ray tracing subdivision surfaces. Computer Graphics Forum, 35(7): 279–288, October 2016. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Davis:2012:LFR

[DLD12] Abe Davis, Marc Levoy, and Fredo Durand. Light fields and reflectance: Unstructured light fields. Computer Graphics Forum, 31(2pt1):305–314, May 2012. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Deng:2012:GMB

[DLGY12] Liqun Deng, Howard Leung, Naijie Gu, and Yang Yang. Generalized model-based human motion recognition with body partition index maps. Computer Graphics Forum, 31(1):202–215, February 2012. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Dey:2010:MPH

[DLL^{+}10] T. K. Dey, K. Li, C. Luo, P. Ranjan, I. Safa, and Y. Wang. Matching: Persistent heat signature for pose-oblivious matching of incomplete models. Computer Graphics Forum, 29(5): 1545–1554, July 2010. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Dey:2010:RIL

[DLS10] T. K. Dey, J. A. Levine, and A. Slatton. Reconstruction II: Localized Delaunay refinement for sampling and meshing. Computer Graphics Forum, 29(5):1723–1732, July 2010. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Du:2018:FAI

[DLY^{+}18] Xingyi Du, Xiaohan Liu, Dong-Ming Yan, Caigui Jiang, Juntao Ye, and Hui Zhang. Field-aligned isotropic surface remeshing. Computer Graphics Forum, 37(6):343–357, September 2018. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Digne:2010:RHF

[DMAL10] J. Digne, J.-M. Morel, N. Audfray, and C. Lartigue. Reconstruction I: High fidelity scan merging. Computer Graphics Forum, 29(5):1643–1651, July 2010. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Dutra:2017:GBS

[DMCN^{+}17] T. B. Dutra, R. Marques, J. B. Cavalcante-Neto, C. A. Vidal, and J. Pettré. Gradient-based steering for vision-based crowd simulation algorithms. Computer Graphics Forum, 36(2):337– 348, May 2017. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Diaz:2012:VVR

[DMNV12] J. Díaz, E. Monclús, I. Navazo, and P. Vázquez. Visualization and volume rendering: Adaptive cross-sections of anatomical models. Computer Graphics Forum, 31(7pt2):2155–2164, September 2012. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Dobos:2014:SCF

[DMS14] Jozef Doboš, Niloy J. Mitra, and Anthony Steed. Shape collections, features: 3D timeline: Reverse engineering of a part-based provenance from consecutive 3D models. Computer Graphics Forum, 33(2):135–144, May 2014. CODEN CG-FODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Digne:2011:SSM

[DMSL11] Julie Digne, Jean-Michel Morel, Charyar-Mehdi Souzani, and Claire Lartigue. Scale space meshing of raw data point sets. Computer Graphics Forum, 30(6):1630–1642, September 2011. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Diehl:2015:EPS

[DPD^{+}15] A. Diehl, L. Pelorosso, C. Delrieux, C. Saulo, J. Ruiz, M. E. Gröller, and S. Bruckner. Engineering and physical sciences: Visual analysis of spatio-temporal data: Applications in weather forecasting. Computer Graphics Forum, 34(3):381– 390, June 2015. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Diehl:2017:VVA

[DPD^{+}17] A. Diehl, L. Pelorosso, C. Delrieux, K. Matković, J. Ruiz, M. E. Gröller, and S. Bruckner. Video and visualization: Albero: a visual analytics approach for probabilistic weather forecasting. Computer Graphics Forum, 36(7):135–144, October 2017. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Dang:2016:NGT

[DPF16] T. N. Dang, N. Pendar, and A. G. Forbes. Networks and graphs 1: TimeArcs: Visualizing fluctuations in dynamic networks. Computer Graphics Forum, 35(3):61–69, June 2016. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Deng:2011:SFW

[DPW11] B. Deng, H. Pottmann, and J. Wallner. Section 1: Functional webs for freeform architecture. Computer Graphics Forum, 30 (5):1369–1378, August 2011. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Dalmau:2010:BSI

[DRA10] Oscar Dalmau, Mariano Rivera, and Teresa Alarcón. Bayesian scheme for interactive colourization, recolourization and image/video editing. Computer Graphics Forum, 29(8):2372– 2386, December 2010. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Delame:2012:CGM

[DRF12] T. Delamé, C. Roudet, and D. Faudot. Computational geometry: From a medial surface to a mesh. Computer Graphics Forum, 31(5):1637–1646, August 2012. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Dextras-Romagnino:2019:SIR

[DRM19] K. Dextras-Romagnino and T. Munzner. Segmentifier: Interactive refinement of clickstream data. Computer Graphics Forum, 38(3):623–634, June 2019. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Dinkla:2015:GDA

[DRW15] K. Dinkla, N. Henry Riche, and M. A. Westenberg. Graphs: Dual adjacency matrix: Exploring link groups in dense networks. Computer Graphics Forum, 34(3):311–320, June 2015. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Daoudi:2011:REW

[DS11a] Mohamed Daoudi and Tobias Schreck. Reports: Eurographics 2010 Workshop on 3D Object Retrieval (EG 3DOR’10) in cooperation with ACM SIGGRAPH. Computer Graphics Forum, 30(1):229–230, March 2011. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Dey:2011:SLD

[DS11b] Tamal K. Dey and Andrew G. Slatton. Section 2: Localized Delaunay refinement for volumes. Computer Graphics Forum, 30(5):1417–1426, August 2011. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Ding:2017:IID

[DSH^{+}17] Shouhong Ding, Bin Sheng, Xiaonan Hou, Zhifeng Xie, and Lizhuang Ma. Intrinsic image decomposition using multi-scale measurements and sparsity. Computer Graphics Forum, 36(6): 251–261, September 2017. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Dym:2015:CGH

[DSL15] N. Dym, A. Shtengel, and Y. Lipman. Curves and graphs: Homotopic morphing of planar curves. Computer Graphics Forum, 34(5):239–251, August 2015. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

daSilva:2017:TRA

[dSNV^{+}17] D. B. da Silva, R. F. Nunes, C. A. Vidal, J. B. Cavalcante-Neto, P. G. Kry, and V. B. Zordan. Tunable robustness: an artificial contact strategy with virtual actuator control for balance. Computer Graphics Forum, 36(8):499–510, December 2017. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Dessein:2017:SAM

[DSWH17] A. Dessein, W. A. P. Smith, R. C. Wilson, and E. R. Hancock. Symmetry-aware mesh segmentation into uniform overlapping patches. Computer Graphics Forum, 36(8):95–107, December 2017. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Dobashi:2010:AMC

[DSY10] Yoshinori Dobashi, Yusuke Shinzo, and Tsuyoshi Yamamoto. Acquisition: Modeling of clouds from a single photograph. Computer Graphics Forum, 29(7):2083–2090, September 2010. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Dhillon:2014:IDB

[DTS^{+}14] D. S. Dhillon, J. Teyssier, M. Single, I. Gaponenko, M. C. Milinkovitch, and M. Zwicker. Interactive diffraction from biological nanostructures. Computer Graphics Forum, 33(8):177– 188, December 2014. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Dai:2015:IVP

[DTV15] D. Dai, R. Timofte, and L. Van Gool. Image and video processing: Jointly optimized regressors for image super-resolution.

Computer Graphics Forum, 34(2):95–104, May 2015. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Duchowski:2014:PEC

[Duc14] A. T. Duchowski. τϵχ́νη Photons: Evolution of a course in data structures. Computer Graphics Forum, 33(1):294–304, February 2014. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Dinkla:2012:GDK

[DvKSW12] Kasper Dinkla, Marc J. van Kreveld, Bettina Speckmann, and Michel A. Westenberg. Graphs and diagrams: Kelp diagrams: Point set membership visualization. Computer Graphics Forum, 31(3pt1):875–884, June 2012. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Desrichard:2019:GIS

[DVP19] François Desrichard, David Vanderhaeghe, and Mathias Paulin. Global illumination shadow layers. Computer Graphics Forum, 38(4):183–191, July 2019. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Diamanti:2014:CFD

[DVPSH14] Olga Diamanti, Amir Vaxman, Daniele Panozzo, and Olga Sorkine-Hornung. Coordinates and fields: Designing N-PolyVector fields with complex polynomials. Computer Graphics Forum, 33(5):1–11, August 2014. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Demir:2013:SPH

[DW13] I. Demir and R. Westermann. Surfaces: Progressive highquality response surfaces for visually guided sensitivity analysis. Computer Graphics Forum, 32(3pt1):21–30, June 2013. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

DeCoro:2010:RDB

[DWR10] Christopher DeCoro, Tim Weyrich, and Szymon Rusinkiewicz. Rendering: Density-based outlier rejection in Monte Carlo rendering. Computer Graphics Forum, 29(7):2119–2125, September 2010. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Dinkla:2011:HLS

[DWT^{+}11] K. Dinkla, M. A. Westenberg, H. M. Timmerman, S. A. F. T. van Hijum, and J. J. Van Wijk. Hierarchies and large-scale visualization: Comparison of multiple weighted hierarchies: Visual analytics for microbe community profiling. Computer Graphics Forum, 30(3):1141–1150, June 2011. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Ding:2019:WFO

[DYFX19] Hong Ding, Qingan Yan, Gang Fu, and Chunxia Xiao. Wavelet flow: Optical flow guided wavelet facial image fusion. Computer Graphics Forum, 38(7):663–674, October 2019. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Deussen:2015:E

[DZ15] Oliver Deussen and Hao (Richard) Zhang. Editorial. Computer Graphics Forum, 34(1):1, February 2015. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Duan:2011:FAT

[DZC11] Qi Duan, Jianmin Zheng, and Jianfei Cai. Flexible and accurate transparent-object matting and compositing using refractive vector field. Computer Graphics Forum, 30(6):1812–1824, September 2011. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Deng:2019:SBS

[DZCC19] Teng Deng, Jianmin Zheng, Jianfei Cai, and Tat-Jen Cham. Shading-based surface recovery using subdivision-based representation. Computer Graphics Forum, 38(1):417–428, February 2019. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Dabala:2016:IVE

[DZD^{+}16] Ł. Da̧bała, M. Ziegler, P. Didyk, F. Zilly, J. Keinert, K. Myszkowski, H.-P. Seidel, P. Rokita, and T. Ritschel. Images and video: Efficient multi-image correspondences for online light field video processing. Computer Graphics Forum, 35 (7):401–410, October 2016. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

El-Assady:2016:TDD

[EAGA^{+}16] Mennatallah El-Assady, Valentin Gold, Carmela Acevedo, Christopher Collins, and Daniel Keim. Text and document data: ConToVi: Multi-party conversation exploration using topic-space views. Computer Graphics Forum, 35(3):431–440, June 2016. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

El-Assady:2017:NNE

[EASG^{+}17] Mennatallah El-Assady, Rita Sevastjanova, Bela Gipp, Daniel Keim, and Christopher Collins. NEREx: Named-entity relationship exploration in multi-party conversations. Computer Graphics Forum, 36(3):213–225, June 2017. CODEN CG-FODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

El-Assady:2018:TMR

[EASKC18] Mennatallah El-Assady, Rita Sevastjanova, Daniel Keim, and Christopher Collins. ThreadReconstructor: Modeling replychains to untangle conversational text through visual analytics. Computer Graphics Forum, 37(3):351–365, June 2018. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Ezuz:2017:SAV

[EBC17] Danielle Ezuz and Mirela Ben-Chen. Shape analysis and variation: Deblurring and denoising of maps between shapes. Computer Graphics Forum, 36(5):165–174, August 2017. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Eisemann:2012:GGP

[EBGM12] M. Eisemann, P. Bauszat, S. Guthe, and M. Magnor. Geometry: Geometry presorting for implicit object space partitioning. Computer Graphics Forum, 31(4):1445–1454, June 2012. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Elek:2014:SWR

[EBR^{+}14] Oskar Elek, Pablo Bauszat, Tobias Ritschel, Marcus Magnor, and Hans-Peter Seidel. Spectral and wave rendering: Spectral ray differentials. Computer Graphics Forum, 33(4):113–122, July 2014. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Eom:2014:CAD

[ECN14] Haegwang Eom, Byungkuk Choi, and Junyong Noh. Character animation: Data-driven reconstruction of human locomotion using a single Smartphone. Computer Graphics Forum, 33(7): 11–19, October 2014. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Edwards:2015:SMA

[EDPB15] John Edwards, Eric Daniel, Valerio Pascucci, and Chandrajit Bajaj. Shape manipulation: Approximating the generalized Voronoi diagram of closely spaced objects. Computer Graphics Forum, 34(2):299–309, May 2015. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Elshehaly:2015:MMM

[EGG^{+}15] M. Elshehaly, D. Gračanin, M. Gad, H. G. Elmongui, and K. Matković. Multi-modal and multi-field: Interactive fusion and tracking for multi-modal spatial data visualization. Computer Graphics Forum, 34(3):251–260, June 2015. CODEN CG-FODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Ezuz:2019:ECB

[EHA^{+}19] D. Ezuz, B. Heeren, O. Azencot, M. Rumpf, and M. Ben-Chen. Elastic correspondence between triangle meshes. Computer Graphics Forum, 38(2):121–134, May 2019. CODEN CG-FODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Engel:2013:ATH

[EHH^{+}13] D. Engel, M. Hummel, F. Hoepel, K. Bein, A. Wexler, C. Garth, B. Hamann, and H. Hagen. Applications: Towards highdimensional data analysis in air quality research. Computer Graphics Forum, 32(3pt1):101–110, June 2013. CODEN CG-FODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Ellis:2019:BRR

[EHH19] A. Ellis, W. Hunt, and J. Hart. Book review: Real-Time Analytic Antialiased Text for 3-D Environments. Computer Graphics Forum, 38(8):23–32, November 2019. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Eckert:2018:CFD

[EHT18] M.-L. Eckert, W. Heidrich, and N. Thuerey. Coupled fluid density and motion from single views. Computer Graphics Forum, 37(8):47–58, December 2018. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Endo:2016:DDI

[EIKM16] Yuki Endo, Satoshi Iizuka, Yoshihiro Kanamori, and Jun Mitani. Data-driven images: DeepProp: Extracting deep features from a single image for edit propagation. Computer Graphics Forum, 35(2):189–201, May 2016. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Eisemann:2011:AWE

[Eis11] Elmar Eisemann. Award winners: Eurographics Young researcher award. Computer Graphics Forum, 30(2):xx, April 2011. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Eikel:2013:VSS

[EJFadH13] Benjamin Eikel, Claudius Jähn, Matthias Fischer, and Friedhelm Meyer auf der Heide. Visibility/sampling: Spherical visibility sampling. Computer Graphics Forum, 32(4):49–58, July 2013. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Eynard:2014:IIL

[EKB14] D. Eynard, A. Kovnatsky, and M. M. Bronstein. Images II: Laplacian colormaps: a framework for structure-preserving color transformations. Computer Graphics Forum, 33(2):215– 224, May 2014. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Endo:2012:IAE

[EKFM12] Yuki Endo, Yoshihiro Kanamori, Yukio Fukui, and Jun Mitani. Image analysis and editing: Matting and compositing for Fresnel reflection on wavy surfaces. Computer Graphics Forum, 31 (4):1435–1443, June 2012. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Eisenberger:2019:DFS

[ELC19] M. Eisenberger, Z. Lähner, and D. Cremers. Divergence-free shape correspondence by deformation. Computer Graphics Forum, 38(5):1–12, August 2019. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Edmunds:2012:FVA

[ELM^{+}12] M. Edmunds, R. S. Laramee, R. Malki, I. Masters, T. N. Croft, G. Chen, and E. Zhang. Flow visualization: Automatic stream surface seeding: a feature centered approach. Computer Graphics Forum, 31(3pt2):1095–1104, June 2012. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Engelke:2019:API

[ELPH19] Wito Engelke, Kai Lawonn, Bernhard Preim, and Ingrid Hotz. Autonomous particles for interactive flow visualization. Computer Graphics Forum, 38(1):248–259, February 2019. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Ebeida:2013:PPS

[EMA^{+}13] Mohamed S. Ebeida, Ahmed H. Mahmoud, Muhammad A. Awad, Mohammed A. Mohammed, Scott A. Mitchell, Alexander Rand, and John D. Owens. Pixels and points: Sifted disks. Computer Graphics Forum, 32(2pt4):509–518, May 2013. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Ebeida:2012:CGG

[EMP^{+}12] Mohamed S. Ebeida, Scott A. Mitchell, Anjul Patney, Andrew A. Davidson, and John D. Owens. Computational geometry and geometry processing: a simple algorithm for maximal Poisson-disk sampling in high dimensions. Computer Graphics Forum, 31(2pt4):785–794, May 2012. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Eilertsen:2017:CRT

[EMU17] G. Eilertsen, R. K. Mantiuk, and J. Unger. A comparative review of tone-mapping algorithms for high dynamic range video. Computer Graphics Forum, 36(2):565–592, May 2017. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Ecormier-Nocca:2019:ASM

[ENMGC19] Pierre Ecormier-Nocca, Pooran Memari, James Gain, and Marie-Paule Cani. Accurate synthesis of multi-class disk distributions. Computer Graphics Forum, 38(2):157–168, May 2019. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Eisenacher:2013:GIS

[ENSB13] Christian Eisenacher, Gregory Nichols, Andrew Selle, and Brent Burley. Global illumination: Sorted deferred shading for production path tracing. Computer Graphics Forum, 32 (4):125–132, July 2013. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Engelhardt:2012:VVR

[ENSD12] Thomas Engelhardt, Jan Novák, Thorsten-W. Schmidt, and Carsten Dachsbacher. Visualization and volume rendering: Approximate bias compensation for rendering scenes with heterogeneous participating media. Computer Graphics Forum, 31 (7pt2):2145–2154, September 2012. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Escalera:2011:VIG

[EPAS11] S. Escalera, A. Puig, O. Amoros, and M. Salamó. Visualization: Intelligent GPGPU classification in volume visualization: a framework based on error-correcting output codes. Computer Graphics Forum, 30(7):2107–2115, September 2011. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Emilien:2015:IPM

[EPCV15] Arnaud Emilien, Pierre Poulin, Marie-Paule Cani, and Ulysse Vimont. Interactive procedural modelling of coherent waterfall scenes. Computer Graphics Forum, 34(6):22–35, September 2015. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Englund:2018:QQA

[ER18] R. Englund and T. Ropinski. Quantitative and qualitative analysis of the perception of semi-transparent structures in direct volume rendering. Computer Graphics Forum, 37(6):174–187, September 2018. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Ebeida:2016:VAD

[ERA^{+}16] Mohamed S. Ebeida, Ahmad A. Rushdi, Muhammad A. Awad, Ahmed H. Mahmoud, Dong-Ming Yan, Shawn A. English, John D. Owens, Chandrajit L. Bajaj, and Scott A. Mitchell. Voronoi et al.: Disk density tuning of a maximal random packing. Computer Graphics Forum, 35(5):259–269, August 2016. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Engel:2011:PSA

[ERHH11] D. Engel, R. Rosenbaum, B. Hamann, and H. Hagen. Parameter spaces and analysis: Structural decomposition trees. Computer Graphics Forum, 30(3):921–930, June 2011. CODEN CG-FODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Endert:2017:SAI

[ERT^{+}17] A. Endert, W. Ribarsky, C. Turkay, B. L. William Wong, I. Nabney, I. Díaz Blanco, and F. Rossi. The state of the art in integrating machine learning into visual analytics. Computer Graphics Forum, 36(8):458–486, December 2017. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Ezuz:2017:MPG

[ESKBC17] Danielle Ezuz, Justin Solomon, Vladimir G. Kim, and Mirela Ben-Chen. Maps and parametrization: GWCNN: a metric alignment layer for deep shape analysis. Computer Graphics Forum, 36(5):49–57, August 2017. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Engelhardt:2014:LCS

[ESKD14] Thomas Engelhardt, Thorsten-Walther Schmidt, Jan Kautz, and Carsten Dachsbacher. Low-cost subpixel rendering for diverse displays. Computer Graphics Forum, 33(1):199–209, February 2014. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Elhayek:2015:OHM

[ESKT15] A. Elhayek, C. Stoll, K. I. Kim, and C. Theobalt. Outdoor human motion capture by simultaneous optimization of pose and camera parameters. Computer Graphics Forum, 34(6):86–98, September 2015. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Esturo:2013:VGS

[ESRT13] Janick Martinez Esturo, Maik Schulze, Christian Rössl, and Holger Theisel. In volumes: Global selection of stream surfaces. Computer Graphics Forum, 32(2pt1):113–122, May 2013. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Echevarria:2013:SAP

[EWK^{+}13] Jose I. Echevarria, Gregg Wilensky, Aravind Krishnaswamy, Byungmoon Kim, and Diego Gutierrez. Simulated and augmented photography: Computational simulation of alternative photographic processes. Computer Graphics Forum, 32(4):7– 16, July 2013. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Eilertsen:2013:IPS

[EWMU13] Gabriel Eilertsen, Robert Wanat, Rafał K. Mantiuk, and Jonas Unger. Image processing (session 4): Evaluation of tone mapping operators for HDR-video. Computer Graphics Forum, 32 (7):275–284, October 2013. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Fried:2017:REI

[FACO17] O. Fried, S. Avidan, and D. Cohen-Or. Representing and editing images: Patch2Vec: Globally consistent image patch representation. Computer Graphics Forum, 36(7):183–194, October 2017. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Filipov:2019:CVE

[FAFM19] V. Filipov, A. Arleo, P. Federico, and S. Miksch. CV3: Visual exploration, assessment, and comparison of CVs. Computer Graphics Forum, 38(3):107–118, June 2019. CODEN CG-FODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Ferstl:2016:FSN

[FAW^{+}16] Florian Ferstl, Ryoichi Ando, Chris Wojtan, Rüdiger Westermann, and Nils Thuerey. Fluid simulation: Narrow band FLIP for liquid simulations. Computer Graphics Forum, 35(2):225– 232, May 2016. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Frohlich:2011:EDD

[FB11] Stefan Fröhlich and Mario Botsch. Example-driven deformations based on discrete shells. Computer Graphics Forum, 30 (8):2246–2257, December 2011. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Fu:2016:TME

[FBL16] Xiao-Ming Fu, Chong-Yang Bai, and Yang Liu. Textures/ mapping: Efficient volumetric PolyCube-map construction. Computer Graphics Forum, 35(7):97–106, October 2016. CO-DEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Faure:2010:RMI

[FC10] William Faure and Chun-Fa Chang. Rendering: Metalights: Improved interleaved shading. Computer Graphics Forum, 29 (7):2109–2117, September 2010. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Fu:2016:ESD

[FCS^{+}16] Qiang Fu, Xiaowu Chen, Xiaoyu Su, Jia Li, and Hongbo Fu. Editing, sketch & drawing: Structure-adaptive shape editing for man-made objects. Computer Graphics Forum, 35(2):27– 36, May 2016. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Fried:2015:ICI

[FDH^{+}15] O. Fried, S. DiVerdi, M. Halber, E. Sizikova, and A. Finkelstein. Image collections: IsoMatch: Creating informative grid layouts. Computer Graphics Forum, 34(2):155–166, May 2015. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Finckh:2014:NOL

[FDL14] M. Finckh, H. Dammertz, and H. P. A. Lensch. On near optimal lattice quantization of multi-dimensional data points. Computer Graphics Forum, 33(1):271–281, February 2014. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Frey:2017:FBT

[FE17] S. Frey and T. Ertl. Flow-based temporal selection for interactive volume visualization. Computer Graphics Forum, 36(8):

153–165, December 2017. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Frohler:2019:VTA

[FEM^{+}19] B. Fröhler, T. Elberfeld, T. Möller, H. C. Hege, J. Weissenböck, J. De Beenhouwer, J. Sijbers, J. Kastner, and C. Heinzl. A visual tool for the analysis of algorithms for tomographic fiber reconstruction in materials science. Computer Graphics Forum, 38(3):273–283, June 2019. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Filho:2018:VCE

[FFN18] J. A. Wagner Filho, C. M. D. S. Freitas, and L. Nedel. VirtualDesk: a comfortable and efficient immersive information visualization approach. Computer Graphics Forum, 37(3):415– 426, June 2018. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Fyffe:2016:FMN

[FGT^{+}16] G. Fyffe, P. Graham, B. Tunwattanapong, A. Ghosh, and P. Debevec. Faces & motion: Near-instant capture of highresolution facial geometry and reflectance. Computer Graphics Forum, 35(2):353–363, May 2016. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Fan:2018:FAC

[FH18] Chaoran Fan and Helwig Hauser. Fast and accurate CNNbased brushing in scatterplots. Computer Graphics Forum, 37 (3):111–120, June 2018. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Fechteler:2019:BRM

[FHE19] P. Fechteler, A. Hilsmann, and P. Eisert. Book review: Markerless Multiview Motion Capture with 3D Shape Model Adaptation. Computer Graphics Forum, 38(6):91–109, September 2019. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Fang:2018:TAM

[FHHJ18] Yu Fang, Yuanming Hu, Shi-Min Hu, and Chenfanfu Jiang. A temporally adaptive material point method with regional time stepping. Computer Graphics Forum, 37(8):195–204, December 2018. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Fyffe:2011:MCS

[FHW^{+}11] Graham Fyffe, Tim Hawkins, Chris Watts, Wan-Chun Ma, and Paul Debevec. Motion capture, simulation, and manipulation: Comprehensive facial performance capture. Computer Graphics Forum, 30(2):425–434, April 2011. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Falk:2013:AVM

[FKE13] Martin Falk, Michael Krone, and Thomas Ertl. Atomistic visualization of mesoscopic whole-cell simulations using ray-casted instancing. Computer Graphics Forum, 32(8):195–206, December 2013. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Fuchs:2013:DFF

[FKR13] Martin Fuchs, Markus Kächele, and Szymon Rusinkiewicz. Design and fabrication of faceted mirror arrays for light field capture. Computer Graphics Forum, 32(8):246–257, December 2013. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Ferstl:2016:PFV

[FKRW16] F. Ferstl, M. Kanzler, M. Rautenhaus, and R. Westermann. Prediction and forecasting: Visual analysis of spatial variability and global correlations in ensembles of iso-contours. Computer Graphics Forum, 35(3):221–230, June 2016. CODEN CG-FODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Fuchs:2010:VTT

[FKS^{+}10] Raphael Fuchs, Jan Kemmler, Benjamin Schindler, Jürgen Waser, Filip Sadlo, Helwig Hauser, and Ronald Peikert. Vectors & tensors: Toward a Lagrangian vector field topology. Computer Graphics Forum, 29(3):1163–1172, June 2010. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Ferreira:2013:VFV

[FKSS13] Nivan Ferreira, James T. Klosowski, Carlos E. Scheidegger, and Cláudio T. Silva. Vector fields: Vector field k-means: Clustering trajectories by fitting multiple vector fields. Computer Graphics Forum, 32(3pt2):201–210, June 2013. CODEN CG-FODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Fofonov:2019:PFS

[FL19] A. Fofonov and L. Linsen. Projected field similarity for comparative visualization of multi-run multi-field time-varying spatial data. Computer Graphics Forum, 38(1):286–299, February 2019. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Fiser:2014:ISC

[FLJ^{+}14] J. Fišer, M. Lukáč, O. Jamriška, M. Čadík, Y. Gingold, P. Asente, and D. Sýkora. Illustration and stylization: Color me noisy: Example-based rendering of hand-colored animations with temporal noise control. Computer Graphics Forum, 33(4): 1–10, July 2014. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Fan:2011:IGP

[FLL11] Lubin Fan, Ligang Lic, and Kun Liu. Interactive geometric processing: Paint mesh cutting. Computer Graphics Forum, 30 (2):603–612, April 2011. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Fujisawa:2015:SVE

[FM15] Makoto Fujisawa and Kenjiro T. Miura. Simulation and visualization: an efficient boundary handling with a modified density calculation for SPH. Computer Graphics Forum, 34(7): 155–162, October 2015. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Fulton:2019:LSD

[FMD^{+}19] Lawson Fulton, Vismay Modi, David Duvenaud, David I. W. Levin, and Alec Jacobson. Latent-space dynamics for reduced deformable simulation. Computer Graphics Forum, 38(2):379– 391, May 2019. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Frohler:2016:VDA

[FMH16] B. Fröhler, T. Möller, and C. Heinzl. Volume data applications: GEMSe: Visualization-guided exploration of multi-channel segmentation algorithms. Computer Graphics Forum, 35(3):191– 200, June 2016. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Fyffe:2017:MVS

[FNH^{+}17] G. Fyffe, K. Nagano, L. Huynh, S. Saito, J. Busch, A. Jones, H. Li, and P. Debevec. Multi-view stereo on consistent face topology. Computer Graphics Forum, 36(2):295–309, May 2017. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Fortunato:2012:CPC

[FO12] Horacio E. Fortunato and Manuel M. Oliveira. Computational photography: Coding depth through mask structure. Computer Graphics Forum, 31(2pt2):459–468, May 2012. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Fratarcangeli:2015:BMS

[FP15] M. Fratarcangeli and F. Pellacini. Bodies in motion: Scalable partitioning for parallel position based dynamics. Computer Graphics Forum, 34(2):405–413, May 2015. CODEN CG-FODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Faridul:2016:CMR

[FPC^{+}16] H. Sheikh Faridul, T. Pouli, C. Chamaret, J. Stauder, E. Reinhard, D. Kuzovkin, and A. Tremeau. Colour mapping: a review of recent methods, extensions and applications. Computer Graphics Forum, 35(1):59–88, February 2016. CODEN CG-FODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Ferdosi:2011:MDV

[FR11] Bilkis J. Ferdosi and Jos B. T. M. Roerdink. Multidimensional data visualization: Visualizing high-dimensional structures by dimension ordering and filtering using subspace analysis. Computer Graphics Forum, 30(3):1121–1130, June 2011. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Frey:2018:STC

[Fre18] S. Frey. Spatio-Temporal contours from deep volume raycasting. Computer Graphics Forum, 37(3):513–524, June 2018. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Frisvad:2014:PDS

[FSES14] Jeppe Revall Frisvad, Lars Schjøth, Kenny Erleben, and Jon Sporring. Photon differential splatting for rendering caustics. Computer Graphics Forum, 33(6):252–263, September 2014. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Fioravante:2013:BSV

[FSTR13] M. Fioravante, A. Shook, I. Thorpe, and P. Rheingans. Biological shape: Visualizing motional correlations in molecular dynamics using geometric deformations. Computer Graphics Forum, 32(3pt3):311–320, June 2013. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Filip:2014:RTB

[FV14] J. Filip and R. Vávra. Rendering: Template-based sampling of anisotropic BRDFs. Computer Graphics Forum, 33(7):91–99, October 2014. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Firmin:2015:CDM

[FvdP15] M. Firmin and M. van de Panne. Controller design for multiskilled bipedal characters. Computer Graphics Forum, 34(8): 50–63, December 2015. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Filip:2017:PVP

[FVHK17] J. Filip, R. Vávra, M. Havlíček, and M. Krupička. Predicting visual perception of material structure in virtual environments. Computer Graphics Forum, 36(1):89–100, January 2017. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Flotynski:2017:OBR

[FW17] Jakub Flotyński and Krzysztof Walczak. Ontology-based representation and modelling of synthetic 3D content: a stateof-the-art review. Computer Graphics Forum, 36(8):329–353, December 2017. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Fan:2011:SHG

[FWPS11] Wenshan Fan, Bin Wang, Jean-Claude Paul, and Jiaguang Sun. Section 3: a hierarchical grid based framework for fast collision detection. Computer Graphics Forum, 30(5):1451–1459, August 2011. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Fang:2019:LES

[FWSH19] Xiaonan Fang, Miao Wang, Ariel Shamir, and Shi-Min Hu. Learning explicit smoothing kernels for joint image filtering. Computer Graphics Forum, 38(7):181–190, October 2019. CO-DEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Fan:2013:GSM

[FWX^{+}13] Lubin Fan, Ruimin Wang, Linlin Xu, Jiansong Deng, and Ligang Liu. Geometry (session 1): Modeling by drawing with shadow guidance. Computer Graphics Forum, 32(7):157–166, October 2013. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Fu:2019:SHR

[FZS^{+}19] Gang Fu, Qing Zhang, Chengfang Song, Qifeng Lin, and Chunxia Xiao. Specular highlight removal for real-world images. Computer Graphics Forum, 38(7):253–263, October 2019. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Ghosh:2010:SCF

[GAK10] Deboshmita Ghosh, Nina Amenta, and Michael Kazhdan. Symmetry: Closed-form blending of local symmetries. Computer Graphics Forum, 29(5):1681–1688, July 2010. CODEN CG-FODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Gamito:2016:SSA

[Gam16] Manuel N. Gamito. Sampling: Solid angle sampling of disk and cylinder lights. Computer Graphics Forum, 35(4):25–36, July 2016. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Gribel:2017:TCQ

[GAM17] C. J. Gribel and T. Akenine-Möller. Time-continuous quasi-Monte Carlo ray tracing. Computer Graphics Forum, 36(6): 354–367, September 2017. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Guerrero:2015:PSM

[GAWJ15] Paul Guerrero, Thomas Auzinger, Michael Wimmer, and Stefan Jeschke. Partial shape matching using transformation parameter similarity. Computer Graphics Forum, 34(1):239–252, February 2015. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Gois:2010:RSD

[GB10] João Paulo Gois and Gustavo C. Buscaglia. Resampling strategies for deforming MLS surfaces. Computer Graphics Forum, 29(6):1969–1980, September 2010. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Gyulassy:2014:TSD

[GBG^{+}14] A. Gyulassy, P. T. Bremer, R. Grout, H. Kolla, J. Chen, and V. Pascucci. Topology: Stability of dissipation elements: a case study in combustion. Computer Graphics Forum, 33(3): 51–60, June 2014. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Gehre:2018:ICC

[GBKS18] A. Gehre, M. Bronstein, L. Kobbelt, and J. Solomon. Interactive curve constrained functional maps. Computer Graphics Forum, 37(5):1–12, August 2018. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Gobbo:2019:TTT

[GBM^{+}19] B. Gobbo, D. Balsamo, M. Mauri, P. Bajardi, A. Panisson, and P. Ciuccarelli. Topic tomographies (TopTom): a visual approach to distill information from media streams. Computer Graphics Forum, 38(3):609–621, June 2019. CODEN CG-FODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Guarnera:2019:TDC

[GBS19] G. C. Guarnera, S. Bianco, and R. Schettini. Turning a digital camera into an absolute 2D tele-colorimeter. Computer Graphics Forum, 38(1):73–86, February 2019. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Goldade:2016:FSP

[GBW16] Ryan Goldade, Christopher Batty, and Chris Wojtan. Fluid simulation: a practical method for high-resolution embedded liquid surfaces. Computer Graphics Forum, 35(2):233–242, May 2016. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Guo:2018:PBA

[GCGP18] Jie Guo, Yanjun Chen, Yanwen Guo, and Jingui Pan. A physically-based appearance model for special effect pigments. Computer Graphics Forum, 37(4):67–76, July 2018. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Gao:2017:DDS

[GCLX17] Lin Gao, Shu-Yu Chen, Yu-Kun Lai, and Shihong Xia. Datadriven shape interpolation and morphing editing. Computer Graphics Forum, 36(8):19–31, December 2017. CODEN CG-FODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Giraudot:2013:RNA

[GCSA13] Simon Giraudot, David Cohen-Steiner, and Pierre Alliez. Reconstruction: Noise-adaptive shape reconstruction from raw point sets. Computer Graphics Forum, 32(5):229–238, August 2013. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Ghanem:2015:IVP

[GCW15] Bernard Ghanem, Yuanhao Cao, and Peter Wonka. Image and video processing: Designing camera networks by convex quadratic programming. Computer Graphics Forum, 34(2):69– 80, May 2015. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Guo:2014:CAL

[GCY^{+}14] Shihui Guo, Jian Chang, Xiaosong Yang, Wencheng Wang, and Jianjun Zhang. Character animation: Locomotion skills for insects with sample-based controller. Computer Graphics Forum, 33(7):31–40, October 2014. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Goldberg:2012:PMD

[GCZ^{+}12] Chen Goldberg, Tao Chen, Fang-Lue Zhang, Ariel Shamir, and Shi-Min Hu. Photo manipulation: Data-driven object manipulation in images. Computer Graphics Forum, 31(2pt1):265–274, May 2012. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Gilet:2010:TGI

[GD10] G. Gilet and J-M. Dischler. Texture generation: an imagebased approach for stochastic volumetric and procedural details. Computer Graphics Forum, 29(4):1411–1419, June 2010. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Ganestam:2016:RTS

[GD16] Per Ganestam and Michael Doggett. Rendering techniques: SAH guided spatial split partitioning for fast BVH construction. Computer Graphics Forum, 35(2):285–293, May 2016. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Gdawiec:2017:IFI

[Gda17] K. Gdawiec. Inversion fractals and iteration processes in the generation of aesthetic patterns. Computer Graphics Forum, 36 (1):35–45, January 2017. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Garcia-Dorado:2014:AID

[GDAU14] I. Garcia-Dorado, D. G. Aliaga, and S. V. Ukkusuri. Animation III: Designing large-scale interactive traffic animations for urban modeling. Computer Graphics Forum, 33(2):411–420, May 2014. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Gilet:2012:PTM

[GDG12] G. Gilet, J-M. Dischler, and D. Ghazanfarpour. Parameterization and texture mapping: Multi-scale assemblage for procedural texturing. Computer Graphics Forum, 31(7pt1):2117–2126, September 2012. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Guerin:2016:TFS

[GDGP16] Eric Guérin, Julie Digne, Eric Galin, and Adrien Peytavie. Terrains & fluids: Sparse representation of terrains for procedural modeling. Computer Graphics Forum, 35(2):177–187, May 2016. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Gautron:2013:RSB

[GDML13] Pascal Gautron, Cyril Delalandre, Jean-Eudes Marvie, and Pascal Lecocq. Rendering (session 2): Boundary-aware extinction mapping. Computer Graphics Forum, 32(7):305–314, October 2013. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Ghani:2012:IDU

[GEY12] S. Ghani, N. Elmqvist, and J. S. Yi. Interaction design and usability: Perception of animated node-link diagrams for dynamic graphs. Computer Graphics Forum, 31(3pt3):1205–1214, June 2012. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Garces:2017:ILF

[GEZ^{+}17] Elena Garces, Jose I. Echevarria, Wen Zhang, Hongzhi Wu, Kun Zhou, and Diego Gutierrez. Intrinsic light field images. Computer Graphics Forum, 36(8):589–599, December 2017. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Groueix:2019:UCC

[GFK^{+}19] Thibault Groueix, Matthew Fisher, Vladimir G. Kim, Bryan C. Russell, and Mathieu Aubry. Unsupervised cycle-consistent deformation for shape matching. Computer Graphics Forum, 38 (5):123–133, August 2019. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Gunther:2014:DCS

[GG14] Tobias Günther and Thorsten Grosch. Distributed out-of-core stochastic progressive photon mapping. Computer Graphics Forum, 33(6):154–166, September 2014. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Gunther:2015:ACI

[GG15] Tobias Günther and Thorsten Grosch. Appearance changes and images: Consistent scene editing by progressive difference images. Computer Graphics Forum, 34(4):41–51, July 2015. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Gunther:2017:FII

[GG17] Tobias Günther and Markus Gross. Flow-induced inertial steady vector field topology. Computer Graphics Forum, 36 (2):143–152, May 2017. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Guarnera:2016:SAR

[GGG^{+}16a] D. Guarnera, G. C. Guarnera, A. Ghosh, C. Denk, and M. Glencross. State of the art reports: BRDF representation and acquisition. Computer Graphics Forum, 35(2):625– 650, May 2016. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Guerin:2016:MEM

[GGG^{+}16b] Eric Guérin, Eric Galin, François Grosbellet, Adrien Peytavie, and Jean-David Génevaux. Modeling: Efficient modeling of entangled details for natural scenes. Computer Graphics Forum, 35(7):257–267, October 2016. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Gobbetti:2012:VVC

[GGM12] Enrico Gobbetti, José Antonio Iglesias Guitián, and Fabio Marton. Volume visualization: COVRA: a compression-domain output-sensitive volume rendering architecture based on a sparse representation of voxel blocks. Computer Graphics Forum, 31(3pt4):1315–1324, June 2012. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Genevaux:2015:TMF

[GGP^{+}15] Jean-David Génevaux, Eric Galin, Adrien Peytavie, Eric Guérin, Cyril Briquet, François Grosbellet, and Bedrich Benes. Terrain modelling from feature primitives. Computer Graphics Forum, 34(6):198–210, September 2015. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Galin:2019:RDT

[GGP^{+}19] Eric Galin, Eric Guérin, Adrien Peytavie, Guillaume Cordonnier, Marie-Paule Cani, Bedrich Benes, and James Gain. A review of digital terrain modeling. Computer Graphics Forum, 38(2):553–577, May 2019. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Gagnon:2019:DUD

[GGV^{+}19] Jonathan Gagnon, Julián E. Guzmán, Valentin Vervondel, François Dagenais, David Mould, and Eric Paquette. Distribution update of deformable patches for texture synthesis on the free surface of fluids. Computer Graphics Forum, 38(7): 491–500, October 2019. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Gu:2015:RTR

[GHB15] Yan Gu, Yong He, and Guy E. Blelloch. Ray tracing: Ray specialized contraction on bounding volume hierarchies. Computer Graphics Forum, 34(7):309–318, October 2015. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Glasser:2017:VIC

[GHB^{+}17] S. Glaßer, T. Hoffmann, A. Boese, S. Voß, T. Kalinski, M. Skalej, and B. Preim. Virtual inflation of the cerebral artery wall for the integrated exploration of OCT and histology data. Computer Graphics Forum, 36(8):57–68, December 2017. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Germann:2010:VAB

[GHK^{+}10] Marcel Germann, Alexander Hornung, Richard Keiser, Remo Ziegler, Stephan Würmlin, and Markus Gross. Video: Articulated billboards for video-based rendering. Computer Graphics Forum, 29(2):585–594, May 2010. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Grottel:2014:VAT

[GHWG14] S. Grottel, J. Heinrich, D. Weiskopf, and S. Gumhold. Visual analysis of trajectories in multi-dimensional state spaces. Computer Graphics Forum, 33(6):310–321, September 2014. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Gao:2017:MEH

[GHX^{+}17] Xifeng Gao, Jin Huang, Kaoji Xu, Zherong Pan, Zhigang Deng, and Guoning Chen. Meshing: Evaluating hex-mesh quality metrics via correlation analysis. Computer Graphics Forum, 36 (5):105–116, August 2017. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Grundhofer:2018:RAP

[GI18] A. Grundhöfer and D. Iwai. Recent advances in projection mapping algorithms, hardware and applications. Computer Graphics Forum, 37(2):653–675, May 2018. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Giachetti:2018:ECR

[Gia18] A. Giachetti. Effective characterization of relief patterns. Computer Graphics Forum, 37(5):83–92, August 2018. CODEN CG-FODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Gassenbauer:2011:LRI

[GKB^{+}11] Václav Gassenbauer, Jaroslav Křivánek, Kadi Bouatouch, Christian Bouville, and Mickaël Ribardière. Lighting & rendering: Improving performance and accuracy of local PCA. Computer Graphics Forum, 30(7):1903–1910, September 2011. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Gao:2012:PMC

[GKB12] Junhong Gao, Seon Joo Kim, and Michael S. Brown. Photo manipulation: Creating picture legends for group photos. Computer Graphics Forum, 31(2pt1):257–264, May 2012. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Gieseke:2014:VTI

[GKHF14] L. Gieseke, S. Koch, J.-U. Hahn, and M. Fuchs. Video and textures: Interactive parameter retrieval for two-tone procedural textures. Computer Graphics Forum, 33(4):71–79, July 2014. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Gunther:2013:VFM

[GKKT13] Tobias Günther, Alexander Kuhn, Benjamin Kutz, and Holger Theisel. Vector fields: Mass-dependent integral curves in unsteady vector fields. Computer Graphics Forum, 32(3pt2):211– 220, June 2013. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Giesen:2017:SPV

[GKL17] J. Giesen, L. Kühne, and P. Lucas. Sclow plots: Visualizing empty space. Computer Graphics Forum, 36(3):145–155, June 2017. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Giesen:2019:VSD

[GKLS19] Joachim Giesen, Julien Klaus, Sören Laue, and Ferdinand Schreck. Visualization support for developing a matrix calculus algorithm: a case study. Computer Graphics Forum, 38 (3):351–361, June 2019. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Guerrero:2018:PLL

[GKOM18] Paul Guerrero, Yanir Kleiman, Maks Ovsjanikov, and Niloy J. Mitra. PCPN et learning local shape properties from raw point clouds. Computer Graphics Forum, 37(2):75–85, May 2018. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Graham:2011:BDV

[GKPL11] Martin Graham, Jessie Kennedy, Trevor Paterson, and Andy Law. Biological data visualization: Visualising errors in animal pedigree genotype data. Computer Graphics Forum, 30 (3):1011–1020, June 2011. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Georgiev:2012:LRI

[GKPS12] Iliyan Georgiev, Jaroslav Křivánek, Stefan Popov, and Philipp Slusallek. Lighting and rendering: Importance caching for complex illumination. Computer Graphics Forum, 31(2pt3):701– 710, May 2012. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Gunther:2016:FVM

[GKT16] Tobias Günther, Alexander Kuhn, and Holger Theisel. Flow visualization: MCFTLE: Monte Carlo rendering of finite-time Lyapunov exponent fields. Computer Graphics Forum, 35(3):

381–390, June 2016. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Garanzha:2010:RTC

[GL10a] Kirill Garanzha and Charles Loop. Ray tracing and culling: Fast ray sorting and breadth-first packet traversal for GPU ray tracing. Computer Graphics Forum, 29(2):289–298, May 2010. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Georgiev:2010:RPC

[GL10b] T. Georgiev and A. Lumsdaine. Reducing plenoptic camera artifacts. Computer Graphics Forum, 29(6):1955–1968, September 2010. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Giachetti:2012:SAR

[GL12] A. Giachetti and C. Lovato. Shape analysis: Radial symmetry detection and shape characterization with the multiscale area projection transform. Computer Graphics Forum, 31(5): 1669–1678, August 2012. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Gain:2017:EIC

[GLCC17] J. Gain, H. Long, G. Cordonnier, and M.-P. Cani. EcoBrush: Interactive control of visually consistent large-scale ecosystems. Computer Graphics Forum, 36(2):63–73, May 2017. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Galvane:2018:DPC

[GLCC18] Q. Galvane, C. Lino, M. Christie, and R. Cozot. Directing the photography: Combining cinematic rules, indirect light controls and lighting-by-example. Computer Graphics Forum, 37(7):45– 53, October 2018. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Gratzl:2016:SHE

[GLG^{+}16] S. Gratzl, A. Lex, N. Gehlenborg, N. Cosgrove, and M. Streit. Story, history, and evolution: From visual exploration to storytelling and back again. Computer Graphics Forum, 35(3):491– 500, June 2016. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Guo:2012:IEP

[GLGW12] Y. W. Guo, M. Liu, T. T. Gu, and W. P. Wang. Image editing and processing: Improving photo composition elegantly: Considering image similarity during composition optimization. Computer Graphics Forum, 31(7pt2):2193–2202, September 2012. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Gao:2013:SCD

[GLHH13] Lin Gao, Yu-Kun Lai, Qi-Xing Huang, and Shi-Min Hu. Shape correspondences: a data-driven approach to realistic shape morphing. Computer Graphics Forum, 32(2pt4):449–457, May 2013. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Gehre:2016:CSA

[GLK16] Anne Gehre, Isaak Lim, and Leif Kobbelt. Curves & surfaces: Adapting feature curve networks to a prescribed scale. Computer Graphics Forum, 35(2):319–330, May 2016. CODEN CG-FODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Gehre:2018:FCC

[GLK18] Anne Gehre, Isaak Lim, and Leif Kobbelt. Feature curve cocompletion in noisy data. Computer Graphics Forum, 37(2): 1–12, May 2018. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Gurung:2011:MPS

[GLLR11] Topraj Gurung, Daniel Laney, Peter Lindstrom, and Jarek Rossignac. Mesh processing and surface reconstruction: SQuad: Compact representation for triangle meshes. Computer Graphics Forum, 30(2):355–364, April 2011. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Galerne:2017:TN

[GLM17] B. Galerne, A. Leclaire, and L. Moisan. Texton noise. Computer Graphics Forum, 36(8):205–218, December 2017. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Guo:2016:MDC

[GLX^{+}16] Xuekun Guo, Juncong Lin, Kai Xu, Siddhartha Chaudhuri, and Xiaogang Jin. Modeling and design: CustomCut: On-demand extraction of customized 3D parts with 2D sketches. Computer Graphics Forum, 35(5):89–100, August 2016. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Ganter:2019:ARC

[GM19] D. Ganter and M. Manzke. An analysis of region clustered BVH volume rendering on GPU. Computer Graphics Forum, 38(8): 13–21, November 2019. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Gerhards:2015:RTR

[GMAG15] J. Gerhards, F. Mora, L. Aveneau, and D. Ghazanfarpour. Real-time rendering & quantization: Partitioned shadow volumes. Computer Graphics Forum, 34(2):549–559, May 2015. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Geisler-Moroder:2010:CGT

[GMD10] David Geisler-Moroder and Arne Dür. Computer graphics theory: a new Ward BRDF model with bounded albedo. Computer Graphics Forum, 29(4):1391–1398, June 2010. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Garces:2012:IAE

[GMLMG12] Elena Garces, Adolfo Munoz, Jorge Lopez-Moreno, and Diego Gutierrez. Image analysis and editing: Intrinsic images by clustering. Computer Graphics Forum, 31(4):1415–1424, June 2012. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Glondu:2012:MSE

[GMM^{+}12] L. Glondu, L. Muguercia, M. Marchal, C. Bosch, H. Rushmeier, G. Dumont, and G. Drettakis. Material synthesis: Examplebased fractured appearance. Computer Graphics Forum, 31(4): 1547–1556, June 2012. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Gain:2015:PPM

[GMM15] J. Gain, B. Merry, and P. Marais. Procedural and parametric modeling: Parallel, realistic and controllable terrain synthesis. Computer Graphics Forum, 34(2):105–116, May 2015. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Gastal:2010:VSS

[GO10] Eduardo S. L. Gastal and Manuel M. Oliveira. Video: Shared sampling for real-time Alpha matting. Computer Graphics Forum, 29(2):575–584, May 2010. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Gastal:2015:IVP

[GO15] Eduardo S. L. Gastal and Manuel M. Oliveira. Image and video processing: High-order recursive filtering of non-uniformly sampled signals for image and video processing. Computer Graphics Forum, 34(2):81–93, May 2015. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Gretarsson:2010:GSV

[GOB^{+}10] Brynjar Gretarsson, John O’Donovan, Svetlin Bostandjiev, Christopher Hall, and Tobias Höllerer. Graphs: SmallWorlds: Visualizing social recommendations. Computer Graphics Forum, 29(3):833–842, June 2010. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Glasser:2010:ATF

[GOH^{+}10] S. Glaßer, S. Oeltze, A. Hennemuth, C. Kubisch, A. Mahnken, S. Wilhelmsen, and B. Preim. Automatic transfer function specification for visual emphasis of coronary artery plaque. Computer Graphics Forum, 29(1):191–201, March 2010. CO-DEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Germer:2011:FVL

[GOPT11] T. Germer, M. Otto, R. Peikert, and H. Theisel. Flow visualization: Lagrangian coherent structures with guaranteed material separation. Computer Graphics Forum, 30(3):761–770, June 2011. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Gove:2019:RSF

[Gov19] R. Gove. A random sampling O(n) force-calculation algorithm for graph layouts. Computer Graphics Forum, 38(3):739–751, June 2019. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Geijtenbeek:2012:ICA

[GP12] T. Geijtenbeek and N. Pronost. Interactive character animation using simulated physics: a state-of-the-art review. Computer Graphics Forum, 31(8):2492–2515, December 2012. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Gonzalez:2016:CIT

[GP16] F. Gonzalez and G. Patow. Continuity and interpolation techniques for computer graphics. Computer Graphics Forum, 35 (1):309–322, February 2016. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Gkaravelis:2018:LOD

[GP18] Anastasios Gkaravelis and Georgios Papaioannou. Light optimization for detail highlighting. Computer Graphics Forum, 37 (7):37–44, October 2018. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Grosbellet:2016:EOA

[GPG^{+}16] François Grosbellet, Adrien Peytavie, Éric Guérin, Éric Galin, Stéphane Mérillou, and Bedrich Benes. Environmental objects for authoring procedural scenes. Computer Graphics Forum, 35 (1):296–308, February 2016. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Galin:2011:PMN

[GPGB11] Eric Galin, Adrien Peytavie, Eric Guérin, and Bedřich Beneš. Procedural modeling & natural phenomena: Authoring hierarchical road networks. Computer Graphics Forum, 30(7):2021– 2030, September 2011. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Grittmann:2018:ECR

[GPGSK18] Pascal Grittmann, Arsène Pérard-Gayot, Philipp Slusallek, and Jaroslav Křivánek. Efficient caustic rendering with lightweight photon mapping. Computer Graphics Forum, 37(4):133–142, July 2018. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Germann:2012:LFR

[GPK^{+}12] Marcel Germann, Tiberiu Popa, Richard Keiser, Remo Ziegler, and Markus Gross. Light fields and reflectance: Novel-view synthesis of outdoor sport events using an adaptive view-dependent geometry. Computer Graphics Forum, 31(2pt1):325–333, May 2012. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Guo:2018:NUF

[GPM^{+}18] J. Guo, Q. K. Pei, G. L. Ma, L. Liu, and X. Y. Zhang. A new uniform format for 360 VR videos. Computer Graphics Forum, 37(7):245–253, October 2018. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Galin:2010:VWP

[GPMG10] E. Galin, A. Peytavie, N. Maréchal, and E. Guérin. Virtual worlds: Procedural generation of roads. Computer Graphics Forum, 29(2):429–438, May 2010. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Gourmel:2010:RTC

[GPP^{+}10] Olivier Gourmel, Anthony Pajot, Mathias Paulin, Loïc Barthe, and Pierre Poulin. Ray tracing and culling: Fitted BVH for fast raytracing of metaballs. Computer Graphics Forum, 29 (2):281–288, May 2010. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Gryaditskaya:2015:EVM

[GPR^{+}15] Yulia Gryaditskaya, Tania Pouli, Erik Reinhard, Karol Myszkowski, and Hans-Peter Seidel. Exposure and video: Motion aware exposure bracketing for HDR video. Computer Graphics Forum, 34(4):119–130, July 2015. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Gryaditskya:2014:RSB

[GPRS14] Yulia Gryaditskya, Tania Pouli, Erik Reinhard, and Hans-Peter Seidel. Rendering: Sky based light metering for high dynamic range images. Computer Graphics Forum, 33(7):61–69, October 2014. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Groller:2011:E

[GR11] Eduard Gröller and Holly Rushmeier. Editorial. Computer Graphics Forum, 30(1):1–2, March 2011. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Gruson:2013:RSE

[GRC13] A. Gruson, M. Ribardière, and R. Cozot. Rendering (session 1): Eye-centered color adaptation in global illumination. Computer Graphics Forum, 32(7):111–120, October 2013. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Grottel:2010:ACC

[GRDE10] Sebastian Grottel, Guido Reina, Carsten Dachsbacher, and Thomas Ertl. Acceleration: Coherent culling and shading for large molecular dynamics visualization. Computer Graphics Forum, 29(3):953–962, June 2010. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Ghani:2011:ODF

[GRE11] S. Ghani, N. Henry Riche, and N. Elmqvist. Overview & detail, focus+context visualization: Dynamic insets for context-aware graph navigation. Computer Graphics Forum, 30(3):861–870, June 2011. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Groller:2011:CC

[Grö11] M. Eduard Gröller. Capstone: Capstone. Computer Graphics Forum, 30(3):xiv, June 2011. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Grosso:2016:RCT

[Gro16] Roberto Grosso. Reconstruction: Construction of topologically correct and manifold isosurfaces. Computer Graphics Forum, 35(5):187–196, August 2016. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Gilles:2010:CAS

[GRP10] B. Gilles, L. Revéret, and D. K. Pai. Creating and animating subject-specific anatomical models. Computer Graphics Forum, 29(8):2340–2351, December 2010. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Gruendl:2016:TSD

[GRPF16] Henning Gruendl, Patrick Riehmann, Yves Pausch, and Bernd Froehlich. Time series data and sequences: Time-series plots integrated in parallel-coordinates displays. Computer Graphics Forum, 35(3):321–330, June 2016. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Gunther:2016:GIS

[GRR^{+}16] Tobias Günther, Kai Rohmer, Christian Rössl, Thorsten Grosch, and Holger Theisel. Global illumination: Stylized caustics: Progressive rendering of animated caustics. Computer Graphics Forum, 35(2):243–252, May 2016. CODEN CG-FODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Gunther:2014:VHO

[GRT14] Tobias Günther, Christian Rössl, and Holger Theisel. Visualization: Hierarchical opacity optimization for sets of 3D line fields. Computer Graphics Forum, 33(2):507–516, May 2014. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Gerrits:2018:APV

[GRT18] Tim Gerrits, Christian Rössl, and Holger Theisel. An approximate parallel vectors operator for multiple vector fields. Computer Graphics Forum, 37(3):315–326, June 2018. CODEN CG-FODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Gerrits:2019:TGU

[GRT19] Tim Gerrits, Christian Rössl, and Holger Theisel. Towards glyphs for uncertain symmetric second-order tensors. Computer Graphics Forum, 38(3):325–336, June 2019. CODEN CG-FODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Gafni:2014:RMC

[GS14] Niv Gafni and Andrei Sharf. Reconstruction: 3D motion completion in crowded scenes. Computer Graphics Forum, 33(5): 65–74, August 2014. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Golla:2018:PLM

[GSC18] Björn Golla, Hans-Peter Seidel, and Renjie Chen. Piecewise linear mapping optimization based on the complex view. Computer Graphics Forum, 37(7):233–243, October 2018. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Guingo:2017:MTB

[GSDC17] Geoffrey Guingo, Basile Sauvage, Jean-Michel Dischler, and Marie-Paule Cani. Materials and textures: Bi-layer textures: a model for synthesis and deformation of composite textures. Computer Graphics Forum, 36(4):111–122, July 2017. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Ganapathi-Subramanian:2018:MLS

[GSDG18] Vignesh Ganapathi-Subramanian, Olga Diamanti, and Leonidas J. Guibas. Modular latent spaces for shape correspondences. Computer Graphics Forum, 37(5):199–210, August 2018. CO-DEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Gladisch:2014:GVS

[GSE^{+}14a] S. Gladisch, H. Schumann, M. Ernst, G. Füllen, and C. Tominski. Graph visualization: Semi-automatic editing of graphs with customized layouts. Computer Graphics Forum, 33(3):381–390, June 2014. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Gunther:2014:FVO

[GSE^{+}14b] Tobias Günther, Maik Schulze, Janick Martinez Esturo, Christian Rössl, and Holger Theisel. Flow visualization: Opacity optimization for surfaces. Computer Graphics Forum, 33(3): 11–20, June 2014. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Gain:2010:AAC

[GSHM10] J. Gain, W. Straßer, A. Hardy, and P. Marais. Afrigraph: The African Computer Graphics Association and its activities. Computer Graphics Forum, 29(1):1, March 2010. CODEN CG-FODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Gao:2019:FPO

[GSP19] Xifeng Gao, Hanxiao Shen, and Daniele Panozzo. Feature preserving octree-based hexahedral meshing. Computer Graphics Forum, 38(5):135–149, August 2019. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Gunther:2014:TMC

[GST14] David Günther, Joseph Salmon, and Julien Tierny. Topology: Mandatory critical points of 2D uncertain scalar fields. Computer Graphics Forum, 33(3):31–40, June 2014. CODEN CG-FODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Ganapathi-Subramanian:2016:FCS

[GSTOG16] V. Ganapathi-Subramanian, B. Thibert, M. Ovsjanikov, and L. Guibas. Functional correspondence: Stable region correspondences between non-isometric shapes. Computer Graphics Forum, 35(5):121–133, August 2016. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Gunther:2012:EDE

[GSW12] D. Günther, H.-P. Seidel, and T. Weinkauf. Extraction of dominant extremal structures in volumetric data using separatrix persistence. Computer Graphics Forum, 31(8):2554–2566, December 2012. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Gregson:2011:SAH

[GSZ11] James Gregson, Alla Sheffer, and Eugene Zhang. Section 2: All-hex mesh generation via volumetric PolyCube deformation. Computer Graphics Forum, 30(5):1407–1416, August 2011. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Gunther:2015:FVF

[GT15] Tobias Günther and Holger Theisel. Flow visualization: Finitetime mass separation for comparative visualizations of inertial particles. Computer Graphics Forum, 34(3):471–480, June 2015. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Gunther:2016:FVS

[GT16a] Tobias Günther and Holger Theisel. Flow visualization: Source inversion by forward integration in inertial flows. Computer Graphics Forum, 35(3):371–380, June 2016. CODEN CG-FODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Gunther:2016:VIS

[GT16b] Tobias Günther and Holger Theisel. Visualization: Inertial steady 2D vector field topology. Computer Graphics Forum, 35(2):455–466, May 2016. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Gunther:2018:SAV

[GT18] Tobias Günther and Holger Theisel. The state of the art in vortex extraction. Computer Graphics Forum, 37(6):149–173, September 2018. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Graham:2013:RMB

[GTB^{+}13] Paul Graham, Borom Tunwattanapong, Jay Busch, Xueming Yu, Andrew Jones, Paul Debevec, and Abhijeet Ghosh. Reality: Measurement-based synthesis of facial microgeometry. Computer Graphics Forum, 32(2pt3):335–344, May 2013. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Guy:2014:SSS

[GTB14] Emilie Guy, Jean-Marc Thiery, and Tamy Boubekeur. Surfaces I: SimSelect: Similarity-based selection for 3D surfaces. Computer Graphics Forum, 33(2):165–173, May 2014. CODEN CG-FODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Gunther:2017:DOO

[GTG17] Tobias Günther, Holger Theisel, and Markus Gross. Decoupled opacity optimization for points, lines and surfaces. Computer Graphics Forum, 36(2):153–162, May 2017. CODEN CG-FODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Granados:2012:IVR

[GTK^{+}12] M. Granados, J. Tompkin, K. Kim, O. Grau, J. Kautz, and C. Theobalt. Image & video retargeting: How not to be seen — object removal from videos of crowded scenes. Computer Graphics Forum, 31(2pt1):219–228, May 2012. CODEN CG-FODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Gao:2018:PCP

[GTL^{+}18] Chengying Gao, Mengyue Tang, Xiangguo Liang, Zhuo Su, and Changqing Zou. PencilArt: a chromatic penciling style generation framework. Computer Graphics Forum, 37(6):395–409, September 2018. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Gallo:2012:CPM

[GTM^{+}12] O. Gallo, M. Tico, R. Manduchi, N. Gelfand, and K. Pulli. Computational photography: Metering for exposure stacks. Computer Graphics Forum, 31(2pt2):479–488, May 2012. CO-DEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Goferman:2010:SCP

[GTZM10] Stas Goferman, Ayellet Tal, and Lihi Zelnik-Manor. Stitching and compositing: Puzzle-like collage. Computer Graphics Forum, 29(2):459–468, May 2010. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Guillen:2017:GTA

[GUK^{+}17] Ibón Guillén, Carlos Ureña, Alan King, Marcos Fajardo, Iliyan Georgiev, Jorge López-Moreno, and Adrian Jarabo. Geometric techniques: Area-preserving parameterizations for spherical ellipses. Computer Graphics Forum, 36(4):179–187, July 2017. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Garcia:2012:DSU

[GUS12] R. García, C. Ureña, and M. Sbert. Description and solution of an unreported intrinsic bias in photon mapping density estimation with constant kernel. Computer Graphics Forum, 31 (1):33–41, February 2012. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Garrido:2015:AAF

[GVS^{+}15] P. Garrido, L. Valgaerts, H. Sarmadi, I. Steiner, K. Varanasi, P. Pérez, and C. Theobalt. All about faces: VDub: Modifying face video of actors for plausible visual alignment to a dubbed audio track. Computer Graphics Forum, 34(2):193–204, May 2015. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Gal:2010:SCS

[GWO^{+}10] Ran Gal, Yonatan Wexler, Eyal Ofek, Hugues Hoppe, and Daniel Cohen-Or. Stitching and compositing: Seamless Montage for texturing models. Computer Graphics Forum, 29(2):

479–486, May 2010. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Guo:2019:DLD

[GZH^{+}19] Yi Guo, Zhuming Zhang, Chu Han, Wenbo Hu, Chengze Li, and Tien-Tsin Wong. Deep line drawing vectorization via line subdivision and topology reconstruction. Computer Graphics Forum, 38(7):81–90, October 2019. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Hahne:2011:IEE

[HA11] Uwe Hahne and Marc Alexa. Image enhancement: Exposure fusion for time-of-flight imaging. Computer Graphics Forum, 30(7):1887–1894, September 2011. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Horacsek:2017:CSB

[HA17] J. J. Horacsek and U. R. Alim. Compactly supported biorthogonal wavelet bases on the body centered cubic lattice. Computer Graphics Forum, 36(3):35–45, June 2017. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Herholz:2019:ECS

[HA19] Philipp Herholz and Marc Alexa. Efficient computation of smoothed exponential maps. Computer Graphics Forum, 38 (6):79–90, September 2019. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Huang:2019:BRL

[HAGO19] Ruqi Huang, Panos Achlioptas, Leonidas Guibas, and Maks Ovsjanikov. Book review: Limit Shapes — A Tool for Understanding Shape Differences and Variability in 3D Model Collections. Computer Graphics Forum, 38(5):187–202, August 2019. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Heitz:2019:DMC

[HB19] E. Heitz and L. Belcour. Distributing Monte Carlo errors as a blue noise in screen space by permuting pixel seeds between frames. Computer Graphics Forum, 38(4):149–158, July 2019. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Hildebrand:2012:MSC

[HBA12] Kristian Hildebrand, Bernd Bickel, and Marc Alexa. Making and shadowing: crdbrd: Shape fabrication by sliding planar slices. Computer Graphics Forum, 31(2pt3):583–592, May 2012. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Hadrich:2017:IMA

[HBDP17] Torsten Hädrich, Bedrich Benes, Oliver Deussen, and Sören Pirk. Interactive modeling and authoring of climbing plants. Computer Graphics Forum, 36(2):49–61, May 2017. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Hoffswell:2018:SHL

[HBH18] Jane Hoffswell, Alan Borning, and Jeffrey Heer. SetCoLa: High-level constraints for graph layout. Computer Graphics Forum, 37(3):537–548, June 2018. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Hu:2017:SRH

[HBLB17] Liwen Hu, Derek Bradley, Hao Li, and Thabo Beeler. Simulation-ready hair capture. Computer Graphics Forum, 36 (2):281–294, May 2017. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Hillaire:2010:UVA

[HBO^{+}10] S. Hillaire, G. Breton, N. Ouarti, R. Cozot, and A. Lécuyer. Using a visual attention model to improve gaze tracking systems in interactive 3D applications. Computer Graphics Forum, 29(6):1830–1841, September 2010. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Harvey:2018:OSR

[HBRD^{+}18] Carlo Harvey, Thomas Bashford-Rogers, Kurt Debattista, Efstratios Doukakis, and Alan Chalmers. Olfaction and selective rendering. Computer Graphics Forum, 37(1):350–362, February 2018. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Halter:2019:VVA

[HBRFP19] Gaudenz Halter, Rafael Ballester-Ripoll, Barbara Flueckiger, and Renato Pajarola. VIAN: a visual annotation tool for film analysis. Computer Graphics Forum, 38(3):119–129, June 2019. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Happa:2012:CHP

[HBRW^{+}12] Jassim Happa, Tom Bashford-Rogers, Alexander Wilkie, Alessandro Artusi, Kurt Debattista, and Alan Chalmers. Cultural heritage predictive rendering. Computer Graphics Forum, 31(6):1823–1836, September 2012. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Heinrich:2011:VAP

[HBW11] J. Heinrich, S. Bachthaler, and D. Weiskopf. 2D visualization by aggregation: Progressive splatting of continuous scatterplots and parallel coordinates. Computer Graphics Forum, 30 (3):653–662, June 2011. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Hoque:2014:DTV

[HC14] E. Hoque and G. Carenini. Document and text visualization: ConVis: a visual text analytic system for exploring blog conversations. Computer Graphics Forum, 33(3):221–230, June 2014. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Herzog:2012:PGN

[HČA^{+}12] Robert Herzog, Martin Čadík, Tunç O. Aydčin, Kwang In Kim, Karol Myszkowski, and Hans-P. Seidel. Perceptual graphics: NoRM: No-reference image quality metric for realistic image synthesis. Computer Graphics Forum, 31(2pt3):545–554, May 2012. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Halperin:2019:CSA

[HCBW19] Tavi Halperin, Harel Cain, Ofir Bibi, and Michael Werman. Clear skies ahead: Towards real-time automatic sky replacement in video. Computer Graphics Forum, 38(2):207–218, May 2019. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Hamdi-Cherif:2018:SRP

[HCDC18] Azzouz Hamdi-Cherif, Julie Digne, and Raphaëlle Chaine. Super-resolution of point set surfaces using local similarities.

Computer Graphics Forum, 37(1):60–70, February 2018. CO-DEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

He:2014:SAM

[HCGW14] Shuiqing He, Yi-King Choi, Yanwen Guo, and Wenping Wang. Spectral analysis on medial axis of 2D shapes. Computer Graphics Forum, 33(8):109–120, December 2014. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Habel:2013:SSM

[HCJ13] Ralf Habel, Per H. Christensen, and Wojciech Jarosz. Surface and subsurface materials: Photon beam diffusion: a hybrid Monte Carlo method for subsurface scattering. Computer Graphics Forum, 32(4):27–37, July 2013. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Huang:2018:SFM

[HCO18] R. Huang, F. Chazal, and M. Ovsjanikov. On the stability of functional maps and shape difference operators. Computer Graphics Forum, 37(1):145–158, February 2018. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Huang:2016:MIR

[HCSC16] Chun-Kai Huang, Yi-Ling Chen, I-Chao Shen, and Bing-Yu Chen. Matching and interpolation: Retargeting 3D objects and scenes with a general framework. Computer Graphics Forum, 35(7):33–42, October 2016. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Hefetz:2017:SAV

[HCW17] Eden Fedida Hefetz, Edward Chien, and Ofir Weber. Shape analysis and variation: Fast planar harmonic deformations with alternating tangential projections. Computer Graphics Forum, 36(5):175–188, August 2017. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Hefetz:2019:SMF

[HCW19] Eden Fedida Hefetz, Edward Chien, and Ofir Weber. A subspace method for fast locally injective harmonic mapping. Computer Graphics Forum, 38(2):105–119, May 2019. CODEN CG-FODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Hanika:2014:REM

[HD14a] Johannes Hanika and Carsten Dachsbacher. Rendering: Efficient Monte Carlo rendering with realistic lenses. Computer Graphics Forum, 33(2):323–332, May 2014. CODEN CG-FODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Heitz:2014:SIS

[Hd14b] E. Heitz and E. d’Eon. Sampling: Importance sampling microfacet-based BSDFs using the distribution of visible normals. Computer Graphics Forum, 33(4):103–112, July 2014. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Harvey:2017:MMP

[HDBRC17] Carlo Harvey, Kurt Debattista, Thomas Bashford-Rogers, and Alan Chalmers. Multi-modal perception for selective rendering. Computer Graphics Forum, 36(1):172–183, January 2017. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Hanika:2015:LPM

[HDF15] Johannes Hanika, Marc Droske, and Luca Fascione. Light paths: Manifold next event estimation. Computer Graphics Forum, 34(4):87–97, July 2015. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Hanika:2011:IEE

[HDL11] Johannes Hanika, Holger Dammertz, and Hendrik Lensch. Image enhancement: Edge-optimized À-trous wavelets for local contrast enhancement with robust denoising. Computer Graphics Forum, 30(7):1879–1886, September 2011. CODEN CG-FODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Hruda:2019:ECR

[HDV19] L. Hruda, J. Dvořák, and L. Váša. On evaluating consensus in RANSAC surface registration. Computer Graphics Forum, 38 (5):175–186, August 2019. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Han:2016:CAD

[HENfSYS16] Daseong Han, Haegwang Eom, Junyong Noh, and Joseph S. Shin (formerly Sung Yong Shin). Cloth & animation: Dataguided model predictive control based on smoothed contact dynamics. Computer Graphics Forum, 35(2):533–543, May 2016. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Hurter:2012:GDG

[HET12] C. Hurter, O. Ersoy, and A. Telea. Graphs and diagrams: Graph bundling by kernel density estimation. Computer Graphics Forum, 31(3pt1):865–874, June 2012. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Herholz:2016:LTP

[HEV^{+}16] Sebastian Herholz, Oskar Elek, Jiří Vorba, Hendrik Lensch, and Jaroslav Křivánek. Light transport: Product importance sampling for light transport path guiding. Computer Graphics Forum, 35(4):67–77, July 2016. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Huber:2015:BHC

[HEW15] M. Huber, B. Eberhardt, and D. Weiskopf. Boundary handling at cloth-fluid contact. Computer Graphics Forum, 34(1):14–25, February 2015. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Huberman:2016:RLV

[HF16] Inbar Huberman and Raanan Fattal. Reducing lateral visual biases in displays. Computer Graphics Forum, 35(8):19–31, December 2016. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Hilsmann:2013:MPS

[HFE13] A. Hilsmann, P. Fechteler, and P. Eisert. Models: Pose space image based rendering. Computer Graphics Forum, 32(2pt3): 265–274, May 2013. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Hu:2012:SAC

[HFL12] Ruizhen Hu, Lubin Fan, and Ligang Liu. Shape analysis: Cosegmentation of 3D shapes via subspace clustering. Computer Graphics Forum, 31(5):1703–1713, August 2012. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Havran:2010:BTF

[HFM10] V. Havran, J. Filip, and K. Myszkowski. Bidirectional texture function compression based on multi-level vector quantization. Computer Graphics Forum, 29(1):175–190, March 2010. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Havran:2016:CNP

[HFM16] V. Havran, J. Filip, and K. Myszkowski. Capturing nature: Perceptually motivated BRDF comparison using single image. Computer Graphics Forum, 35(4):1–12, July 2016. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Huang:2013:SCM

[HG13] Qi-Xing Huang and Leonidas Guibas. Shape collections and maps: Consistent shape maps via semidefinite programming. Computer Graphics Forum, 32(5):177–186, August 2013. CO-DEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Heimerl:2018:IAW

[HG18] F. Heimerl and M. Gleicher. Interactive analysis of word vector embeddings. Computer Graphics Forum, 37(3):253–265, June 2018. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Hnaidi:2010:MFB

[HGA^{+}10] Houssam Hnaidi, Eric Guérin, Samir Akkouche, Adrien Peytavie, and Eric Galin. Modeling: Feature based terrain generation using diffusion equation. Computer Graphics Forum, 29(7):2179–2186, September 2010. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Hoferlin:2010:AVV

[HGB^{+}10] M. Höferlin, E. Grundy, R. Borgo, D. Weiskopf, M. Chen, I. W. Griffiths, and W. Griffiths. Applications: Video visualization for snooker skill training. Computer Graphics Forum, 29(3): 1053–1062, June 2010. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Hummel:2011:CCI

[HGH^{+}11] Mathias Hummel, Christoph Garth, Bernd Hamann, Hans Hagen, and Kenneth I. Joy. CGF cover image 2011: Illustrative visualization of a vortex breakdown bubble. Computer Graphics Forum, 30(1):235–236, March 2011. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Hua:2017:GDP

[HGNH17] Binh-Son Hua, Adrien Gruson, Derek Nowrouzezahrai, and Toshiya Hachisuka. Gradient-domain photon density estimation. Computer Graphics Forum, 36(2):31–38, May 2017. CO-DEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Henz:2018:DJD

[HGO18] Bernardo Henz, Eduardo S. L. Gastal, and Manuel M. Oliveira. Deep joint design of color filter arrays and demosaicing. Computer Graphics Forum, 37(2):389–399, May 2018. CODEN CG-FODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Hua:2019:SGD

[HGP^{+}19] Binh-Son Hua, Adrien Gruson, Victor Petitjean, Matthias Zwicker, Derek Nowrouzezahrai, Elmar Eisemann, and Toshiya Hachisuka. A survey on gradient-domain rendering. Computer Graphics Forum, 38(2):455–472, May 2019. CODEN CG-FODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Hofmann:2017:VEG

[HGRS^{+}17] J. Hofmann, M. Größler, M. Rubio-Sánchez, P.-P. Pichler, and D. J. Lehmann. Visual exploration of global trade networks with time-dependent and weighted hierarchical edge bundles on GPU. Computer Graphics Forum, 36(3):273–282, June 2017. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Hoberock:2010:AIF

[HH10] Jared Hoberock and John C. Hart. Arbitrary importance functions for Metropolis light transport. Computer Graphics Forum, 29(6):1993–2003, September 2010. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Hapala:2011:RTT

[HH11] M. Hapala and V. Havran. Review: kd-tree traversal algorithms for ray tracing. Computer Graphics Forum, 30(1):199– 213, March 2011. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Herholz:2017:DDV

[HHA17] Philipp Herholz, Felix Haase, and Marc Alexa. Diffusion diagrams: Voronoi cells and centroids from diffusion. Computer Graphics Forum, 36(2):163–175, May 2017. CODEN CG-FODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Huettenberger:2013:TTM

[HHC^{+}13] L. Huettenberger, C. Heine, H. Carr, G. Scheuermann, and C. Garth. Topology: Towards multifield scalar topology based on Pareto optimality. Computer Graphics Forum, 32(3pt3): 341–350, June 2013. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Huang:2018:RIC

[HHCJ18] Z. Y. Huang, M. Holloway, N. Carr, and T. Ju. Repairing inconsistent curve networks on non-parallel cross-sections. Computer Graphics Forum, 37(2):25–35, May 2018. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Hulusic:2012:ARA

[HHD^{+}12] Vedad Hulusic, Carlo Harvey, Kurt Debattista, Nicolas Tsingos, Steve Walker, David Howard, and Alan Chalmers. Acoustic rendering and auditory–visual cross-modal perception and interaction. Computer Graphics Forum, 31(1):102–131, February 2012. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Hostettler:2015:SDB

[HHGJ15] Rafael Hostettler, Ralf Habel, Markus Gross, and Wojciech Jarosz. Stylization: Dispersion-based color projection using masked prisms. Computer Graphics Forum, 34(7):329–338, October 2015. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Hullin:2012:OPO

[HHH12] Matthias B. Hullin, Johannes Hanika, and Wolfgang Heidrich. Optics: Polynomial optics: a construction kit for efficient raytracing of lens systems. Computer Graphics Forum, 31(4):1375– 1383, June 2012. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Havran:2019:ISR

[HHNČ19] V. Havran, J. Hošek, Š. Němcová, and J. Čáp. Increasing the spatial resolution of BTF measurement with scheimpflug imaging. Computer Graphics Forum, 38(1):592–609, February 2019. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Huang:2012:RST

[HHRZ12] Xuezhen Huang, Qiming Hou, Zhong Ren, and Kun Zhou. Rendering systems and techniques: Scalable programmable motion effects on GPUs. Computer Graphics Forum, 31(7pt2):2259– 2266, September 2012. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Hosu:2014:IIL

[HHS14] Vlad Hosu, Mai Lan Ha, and Terence Sim. Images II: Light montage for perceptual image enhancement. Computer Graphics Forum, 33(2):185–194, May 2014. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Hao:2011:TDV

[HJM^{+}11] M. C. Hao, H. Janetzko, S. Mittelstädt, W. Hill, U. Dayal, D. A. Keim, M. Marwah, and R. K. Sharma. Temporal data visualization: a visual analytics approach for peak-preserving prediction of large seasonal time series. Computer Graphics Forum, 30(3):691–700, June 2011. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Hummel:2017:VPM

[HJS^{+}17] M. Hummel, L. Jöckel, J. Schäfer, M. W. Hlawitschka, and C. Garth. Visualizing probabilistic multi-phase fluid simulation data using a sampling approach. Computer Graphics Forum, 36(3):469–477, June 2017. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Habbecke:2012:IMU

[HK12] Martin Habbecke and Leif Kobbelt. Interactive modeling and user interfaces: Linear analysis of nonlinear constraints for interactive geometric modeling. Computer Graphics Forum, 31 (2pt3):641–650, May 2012. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Han:2016:DSO

[HK16] D. Han and J. Keyser. Deformable & soft objects: Effect of low-level visual details in perception of deformation. Computer Graphics Forum, 35(2):375–383, May 2016. CODEN CG-FODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Herholz:2015:QPP

[HKA15] Philipp Herholz, Jan Eric Kyprianidis, and Marc Alexa. Quads and polygons: Perfect Laplacians for polygon meshes. Computer Graphics Forum, 34(5):211–218, August 2015. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Hanika:2015:LPI

[HKD15] Johannes Hanika, Anton Kaplanyan, and Carsten Dachsbacher. Light paths: Improved half vector space light transport. Computer Graphics Forum, 34(4):65–74, July 2015. CODEN CG-FODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Hullman:2017:FCP

[HKL17] Jessica Hullman, Robert Kosara, and Heidi Lam. Finding a clear path: Structuring strategies for visualization sequences. Computer Graphics Forum, 36(3):365–375, June 2017. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Huang:2015:DSS

[HKM15] Haibin Huang, Evangelos Kalogerakis, and Benjamin Marlin. Descriptors and shape synthesis: Analysis and synthesis of 3D shape families via deep-learned generative models of surfaces. Computer Graphics Forum, 34(5):25–38, August 2015. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Hwang:2018:RTL

[HKSK18] Jaepyung Hwang, Jongmin Kim, Il Hong Suh, and Taesoo Kwon. Real-time locomotion controller using an Inverted-Pendulum-based abstract model. Computer Graphics Forum, 37(2):287–296, May 2018. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Kim:2017:CSS

[hKTL^{+}17] Seong heum Kim, Yu-Wing Tai, Joon-Young Lee, Jaesik Park, and In So Kweon. Category-specific salient view selection via deep convolutional neural networks. Computer Graphics Forum, 36(8):313–328, December 2017. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Habel:2012:CPP

[HKW12] Ralf Habel, Michael Kudenov, and Michael Wimmer. Computational photography: Practical spectral photography. Computer Graphics Forum, 31(2pt2):449–458, May 2012. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

He:2013:SCT

[HL13] S. He and R. W. H. Lau. Synthetic controllable turbulence using robust second vorticity confinement. Computer Graphics Forum, 32(1):27–35, February 2013. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Hergel:2014:CPC

[HL14] Jean Hergel and Sylvain Lefebvre. Cutting and printing: Clean color: Improving multi-filament 3D prints. Computer Graphics Forum, 33(2):469–478, May 2014. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Hergel:2015:FFM

[HL15] Jean Hergel and Sylvain Lefebvre. Fabrication: 3D fabrication of 2D mechanisms. Computer Graphics Forum, 34(2):229–238, May 2015. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Hu:2018:MSS

[HL18] Chen-Hui Hu and Wen-Chieh Lin. Motion sickness simulation based on sensorimotor control. Computer Graphics Forum, 37 (2):475–484, May 2018. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Hao:2019:CVP

[HL19] Yue Hao and Jyh-Ming Lien. Compacting voxelized polyhedra via tree stacking. Computer Graphics Forum, 38(7):323–333, October 2019. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Heine:2016:FES

[HLH^{+}16a] C. Heine, H. Leitte, M. Hlawitschka, F. Iuricich, L. De Floriani, G. Scheuermann, H. Hagen, and C. Garth. Feature extraction for scalar and vector fields: A survey of topology-based methods in visualization. Computer Graphics Forum, 35(3):643–667, June 2016. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Huang:2016:VNT

[HLH^{+}16b] H. Huang, D. Lischinski, Z. Hao, M. Gong, M. Christie, and D. Cohen-Or. Visualization/NPR: Trip synopsis: 60km in 60sec. Computer Graphics Forum, 35(7):107–116, October 2016. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Hughes:2013:ICK

[HLJ^{+}13] D. M. Hughes, I. S. Lim, M. W. Jones, A. Knoll, and B. Spencer. InK-Compact: In-kernel stream compaction and its application to multi-kernel data visualization on generalpurpose GPUs. Computer Graphics Forum, 32(6):178–188, September 2013. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

He:2012:LPS

[HLL^{+}12] Xiaowei He, Ning Liu, Sheng Li, Hongan Wang, and Guoping Wang. Local Poisson SPH for viscous incompressible fluids. Computer Graphics Forum, 31(6):1948–1958, September 2012. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Hyun:2016:CAM

[HLL16] Kyunglyul Hyun, Kyungho Lee, and Jehee Lee. Character animation: Motion grammars for character animation. Computer Graphics Forum, 35(2):103–113, May 2016. CODEN CG-FODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Hu:2019:SSB

[HLL^{+}19] Chen-Hui Hu, Chien-Ying Lee, Yen-Ting Liou, Feng-Yu Sung, and Wen-Chieh Lin. Skiing simulation based on skill-guided motion planning. Computer Graphics Forum, 38(6):66–78, September 2019. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Hullin:2011:SAD

[HLR^{+}11] Matthias B. Hullin, Hendrik P. A. Lensch, Ramesh Raskar, Hans-Peter Seidel, and Ivo Ihrke. Surface appearance: Dynamic display of BRDFs. Computer Graphics Forum, 30(2):475–483, April 2011. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

He:2012:PTM

[HLS12] Lei He, Charles Loop, and Scott Schaefer. Parameterization and texture mapping: Improving the parameterization of approximate subdivision surfaces. Computer Graphics Forum, 31 (7pt1):2127–2134, September 2012. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Hennessey:2015:GII

[HM15] James W. Hennessey and Niloy J. Mitra. Geometry and images: an image degradation model for depth-augmented image editing. Computer Graphics Forum, 34(5):191–199, August 2015. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Herholz:2015:FAF

[HMA15] Philipp Herholz, Wojciech Matusik, and Marc Alexa. Fabrication: Approximating free-form geometry with height fields for manufacturing. Computer Graphics Forum, 34(2):239–251, May 2015. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Hendrich:2017:PBC

[HMB17] J. Hendrich, D. Meister, and J. Bittner. Parallel BVH construction using progressive hierarchical refinement. Computer Graphics Forum, 36(2):487–494, May 2017. CODEN CG-FODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Hajisharif:2019:LFV

[HML^{+}19] Saghi Hajisharif, Ehsan Miandji, Per Larsson, Kiet Tran, and Jonas Unger. Light field video compression and real time rendering. Computer Graphics Forum, 38(7):265–276, October 2019. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Ha:2013:APS

[HMLP13] Sehoon Ha, Jim McCann, C. Karen Liu, and Jovan Popović. Animation: Physics storyboards. Computer Graphics Forum, 32(2pt2):133–142, May 2013. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Herghelegiu:2012:MVB

[HMP^{+}12] P. C. Herghelegiu, V. Manta, R. Perin, S. Bruckner, and E. Gröller. Medical visualization: Biopsy planner — visual analysis for needle pathway planning in deep seated brain tumor biopsy. Computer Graphics Forum, 31(3pt2):1085–1094, June 2012. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Hermosilla:2019:DLL

[HMRR19] P. Hermosilla, S. Maisch, T. Ritschel, and T. Ropinski. Deeplearning the latent space of light transport. Computer Graphics Forum, 38(4):207–217, July 2019. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Heckel:2013:SBE

[HMTH13] Frank Heckel, Jan H. Moltz, Christian Tietjen, and Horst K. Hahn. Sketch-based editing tools for tumour segmentation in 3D medical images. Computer Graphics Forum, 32(8):144–157, December 2013. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Herzog:2015:CML

[HMW^{+}15] Robert Herzog, Daniel Mewes, Michael Wand, Leonidas Guibas, and Hans-Peter Seidel. Correspondence and matching: LeSSS: Learned shared semantic spaces for relating multimodal representations of 3D shapes. Computer Graphics Forum, 34(5):141–151, August 2015. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Han:2014:APB

[HNJ^{+}14] Daseong Han, Junyong Noh, Xiaogang Jin, Joseph S. Shin, and Sung Y. Shin. Animation I (physically-based): On-line real-time physics-based predictive motion control with balance recovery. Computer Graphics Forum, 33(2):245–254, May 2014. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Huang:2017:SAV

[HO17] Ruqi Huang and Maks Ovsjanikov. Shape analysis and variation: Adjoint map representation for shape analysis and matching. Computer Graphics Forum, 36(5):151–163, August 2017. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Ha:2019:GOR

[HOB^{+}19] Saerom Ha, Sojin Oh, Jonghee Back, Sung-Eui Yoon, and Bochang Moon. Gradient outlier removal for gradient-domain path tracing. Computer Graphics Forum, 38(2):245–253, May 2019. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Holzschuch:2015:ACS

[Hol15] N. Holzschuch. Accurate computation of single scattering in participating media with refractive boundaries. Computer Graphics Forum, 34(6):48–59, September 2015. CODEN CG-FODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Hildebrandt:2011:SAL

[HP11] Klaus Hildebrandt and Konrad Polthier. Section 5: On approximation of the Laplace–Beltrami operator and the Willmore energy of surfaces. Computer Graphics Forum, 30(5):1513–1520, August 2011. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Hong:2010:SSD

[HPH10] Q. Youn Hong, Sang Il Park, and Jessica K. Hodgins. Shape segmentation: a data-driven segmentation for the shoulder complex. Computer Graphics Forum, 29(2):537–544, May 2010. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Hall:2016:IVT

[HPK^{+}16] K. Wm. Hall, C. Perin, P. G. Kusalik, C. Gutwin, and S. Carpendale. Information visualization techniques: Formalizing emphasis in information visualization. Computer Graphics Forum, 35(3):717–737, June 2016. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Hendrich:2019:RCA

[HPMB19] J. Hendrich, A. Pospíšil, D. Meister, and J. Bittner. Ray classification for accelerated BVH traversal. Computer Graphics Forum, 38(4):49–56, July 2019. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Hollt:2016:BDV

[HPvU^{+}16] T. Höllt, N. Pezzotti, V. van Unen, F. Koning, E. Eisemann, B. Lelieveldt, and A. Vilanova. Biological data visualization: Cytosplore: Interactive immune cell phenotyping for large single-cell datasets. Computer Graphics Forum, 35(3):171–180, June 2016. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Hou:2015:TBB

[HQH15] Fei Hou, Hong Qin, and Aimin Hao. Trivariate biharmonic B-splines. Computer Graphics Forum, 34(6):36–47, September 2015. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Huang:2010:SCS

[HR10] Fu-Chung Huang and Ravi Ramamoorthi. Sparse computing: Sparsely precomputing the light transport matrix for real-time rendering. Computer Graphics Forum, 29(4):1335–1345, June 2010. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Han:2015:THV

[HRD^{+}15] Y. Han, A. Rozga, N. Dimitrova, G. D. Abowd, and J. Stasko. Text & humanities: Visual analysis of proximal temporal relationships of social and communicative behaviors. Computer Graphics Forum, 34(3):51–60, June 2015. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Hollander:2011:GRM

[HREB11] Matthias Hollander, Tobias Ritschel, Elmar Eisemann, and Tamy Boubekeur. Geometry for rendering: ManyLoDs: Parallel many-view level-of-detail selection for real-time global illumination. Computer Graphics Forum, 30(4):1233–1240, June 2011. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Henzler:2018:SIT

[HRRR18] Phlipp Henzler, Volker Rasche, Timo Ropinski, and Tobias Ritschel. Single-image tomography: 3D volumes from 2D cranial X-Rays. Computer Graphics Forum, 37(2):377–388, May 2018. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Heeren:2014:SSE

[HRS^{+}14] B. Heeren, M. Rumpf, P. Schröder, M. Wardetzky, and B. Wirth. Surfaces and shells: Exploring the geometry of the space of shells. Computer Graphics Forum, 33(5):247–256, August 2014. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Heeren:2016:MDS

[HRS^{+}16] Behrend Heeren, Martin Rumpf, Peter Schröder, Max Wardetzky, and Benedikt Wirth. Modeling and design: Splines in the space of shells. Computer Graphics Forum, 35(5):111–120, August 2016. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Hofmann:2018:VVF

[HRS18] Lutz Hofmann, Bastian Rieck, and Filip Sadlo. Visualization of 4D vector field topology. Computer Graphics Forum, 37 (3):301–313, June 2018. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Heeren:2012:FTD

[HRWW12] B. Heeren, M. Rumpf, M. Wardetzky, and B. Wirth. Flows: Time-discrete geodesics in the space of shells. Computer Graphics Forum, 31(5):1755–1764, August 2012. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Heinzl:2017:SVC

[HS17] C. Heinzl and S. Stappen. STAR: Visual computing in materials science. Computer Graphics Forum, 36(3):647–666, June 2017. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Hofmann:2019:DVO

[HS19] Lutz Hofmann and Filip Sadlo. The dependent vectors operator. Computer Graphics Forum, 38(3):261–272, June 2019. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Hlawatsch:2013:ITS

[HSBW13] M. Hlawatsch, F. Sadlo, M. Burch, and D. Weiskopf. InfoVis techniques: Scale-stack bar charts. Computer Graphics Forum, 32(3pt2):181–190, June 2013. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Hoffswell:2016:CVI

[HSH16] Jane Hoffswell, Arvind Satyanarayan, and Jeffrey Heer. Coordinated views and interaction design: Visual debugging techniques for reactive data visualization. Computer Graphics Forum, 35(3):271–280, June 2016. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Hlawatsch:2014:VPG

[HSJW14] M. Hlawatsch, F. Sadlo, H. Jang, and D. Weiskopf. Visualization: Pathline glyphs. Computer Graphics Forum, 33(2):497– 506, May 2014. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Hao:2010:AVA

[HSK^{+}10] M. C. Hao, R. K. Sharma, D. A. Keim, U. Dayal, C. Patel, and R. Vennelakanti. Application of visual analytics for thermal state management in large data centres. Computer Graphics Forum, 29(6):1895–1904, September 2010. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Hwang:2014:CAE

[HSK14] J. Hwang, I. H. Suh, and T. Kwon. Character animation: Editing and synthesizing two-character motions using a coupled inverted pendulum model. Computer Graphics Forum, 33(7):21– 30, October 2014. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Ho:2013:AST

[HSmCY13] Edmond S. L. Ho, Hubert P. H. Shum, Yiu ming Cheung, and P. C. Yuen. Animation (session 1): Topology aware data-driven inverse kinematics. Computer Graphics Forum, 32(7):61–70, October 2013. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Haubenwallner:2017:SUG

[HSS17] Karl Haubenwallner, Hans-Peter Seidel, and Markus Steinberger. ShapeGenetics: Using genetic algorithms for procedural modeling. Computer Graphics Forum, 36(2):213–223, May 2017. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Hladky:2019:TSS

[HSS19] J. Hladky, H. P. Seidel, and M. Steinberger. Tessellated shading streaming. Computer Graphics Forum, 38(4):171–182, July 2019. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Hu:2018:FRA

[HSvK18] R. Hu, M. Savva, and O. van Kaick. Functionality representations and applications for shape analysis. Computer Graphics Forum, 37(2):603–624, May 2018. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Harary:2011:NPN

[HT11] Gur Harary and Ayellet Tal. Natural phenomena: The natural 3D spiral. Computer Graphics Forum, 30(2):237–246, April 2011. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Harary:2014:RFP

[HTG14] Gur Harary, Ayellet Tal, and Eitan Grinspun. Reconstruction: Feature-preserving surface completion using four points. Computer Graphics Forum, 33(5):45–54, August 2014. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Hua:2017:BDP

[Hua17] H. Hua. A bi-directional procedural model for architectural design. Computer Graphics Forum, 36(8):219–231, December 2017. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Holten:2010:MDE

[HV10] Danny Holten and Jarke J. Van Wijk. Multivariate data: Evaluation of cluster identification performance for different PCP variants. Computer Graphics Forum, 29(3):793–802, June 2010. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Heinrich:2016:BDV

[HVH^{+}16] J. Heinrich, J. Vuong, C. J. Hammang, A. Wu, M. Rittenbruch, J. Hogan, M. Brereton, and S. I. O’Donoghue. Biological data visualization: Evaluating viewpoint entropy for ribbon representation of protein structure. Computer Graphics Forum, 35 (3):181–190, June 2016. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Hollt:2019:FEH

[HVP^{+}19] T. Höllt, A. Vilanova, N. Pezzotti, B. P. F. Lelieveldt, and H. Hauser. Focus+context exploration of hierarchical embeddings. Computer Graphics Forum, 38(3):569–579, June 2019. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Hermosilla:2018:GIM

[HVVR18] P. Hermosilla, P. Vázquez, A. Vinacua, and T. Ropinski. A general illumination model for molecular visualization. Computer Graphics Forum, 37(3):367–378, June 2018. CODEN CG-FODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Harvey:2010:NAS

[HW10] William Harvey and Yusu Wang. Navigating abstract spaces: Topological landscape ensembles for visualization of scalarvalued functions. Computer Graphics Forum, 29(3):993–1002, June 2010. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Hafner:2016:VIF

[HW16] David Hafner and Joachim Weickert. Variational image fusion with optimal local contrast. Computer Graphics Forum, 35(1): 100–112, February 2016. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Haegler:2010:PTT

[HWA^{+}10] Simon Haegler, Peter Wonka, Stefan Müller Arisona, Luc Van Gool, and Pascal Müller. Procedural textures and texture atlases: Grammar-based encoding of facades. Computer Graphics Forum, 29(4):1479–1487, June 2010. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Huang:2017:MVG

[HWF^{+}17] Jing Huang, Qi Wang, Marco Fratarcangeli, Ke Yan, and Catherine Pelachaud. Multi-variate Gaussian-based inverse kinematics. Computer Graphics Forum, 36(8):418–428, December 2017. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Heinzle:2010:RIM

[HWK^{+}10] Simon Heinzle, Johanna Wolf, Yoshihiro Kanamori, Tim Weyrich, Tomoyuki Nishita, and Markus Gross. Rendering II: Motion blur for EWA surface splatting. Computer Graphics Forum, 29(2):733–742, May 2010. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Hochstetter:2015:FVV

[HWK15] H. Hochstetter, M. Wurm, and A. Kolb. Flow visualization: Vector field visualization of advective-diffusive flows. Computer Graphics Forum, 34(3):481–490, June 2015. CODEN CG-FODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Huang:2018:TIR

[HWL18] Yifei Huang, Changbo Wang, and Chenhui Li. Translucent image recoloring through homography estimation. Computer Graphics Forum, 37(7):421–432, October 2018. CODEN CG-FODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Han:2019:RTG

[HWU^{+}19] Mengjiao Han, Ingo Wald, Will Usher, Qi Wu, Feng Wang, Valerio Pascucci, Charles D. Hansen, and Chris R. Johnson. Ray tracing generalized tube primitives: Method and applications. Computer Graphics Forum, 38(3):467–478, June 2019. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Ho:2015:FVE

[HYL^{+}15] Hsin-Yang Ho, I-Cheng Yeh, Yu-Chi Lai, Wen-Chieh Lin, and Fu-Yin Cherng. Flow visualization: Evaluating 2D flow visualization using eye tracking. Computer Graphics Forum, 34 (3):501–510, June 2015. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Huang:2014:PDS

[HYZ^{+}14] Zhichao Huang, Junfeng Yao, Zichun Zhong, Yang Liu, and Xiaohu Guo. Particles and deformation: Sparse localized decomposition of deformation gradients. Computer Graphics Forum, 33(7):239–248, October 2014. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Huang:2010:IVV

[HZF10] Hua Huang, Lei Zhang, and Tian-Nan Fu. Image and video I: Video painting via motion layer manipulation. Computer Graphics Forum, 29(7):2055–2064, September 2010. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Huang:2019:SMS

[HZL19] Zixuan Huang, Jinghuai Zhang, and Jing Liao. Style mixer: Semantic-aware multi-style transfer network. Computer Graphics Forum, 38(7):469–480, October 2019. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Huang:2014:CIL

[HZMH14] H.-Z. Huang, S.-H. Zhang, R. R. Martin, and S.-M. Hu. Color and imaging: Learning natural colors for image recoloring. Computer Graphics Forum, 33(7):299–308, October 2014. CO-DEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Huang:2018:QSR

[HZN^{+}18] Jingwei Huang, Yichao Zhou, Matthias Niessner, Jonathan Richard Shewchuk, and Leonidas J. Guibas. QuadriFlow: a scalable and robust method for quadrangulation. Computer Graphics Forum, 37(5):147–160, August 2018. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Heeren:2018:PGA

[HZRS18] B. Heeren, C. Zhang, M. Rumpf, and W. Smith. Principal geodesic analysis in the space of discrete shells. Computer Graphics Forum, 37(5):173–184, August 2018. CODEN CG-FODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Huang:2011:IVE

[HZZ11] Hua Huang, Lei Zhang, and Hong-Chao Zhang. Image & video editing: RepSnapping: Efficient image cutout for repeated scene elements. Computer Graphics Forum, 30(7):2059–2066, September 2011. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Ihmsen:2011:PSI

[IABT11] Markus Ihmsen, Nadir Akinci, Markus Becker, and Matthias Teschner. A parallel SPH implementation on multi-core CPUs. Computer Graphics Forum, 30(1):99–112, March 2011. CO-DEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Isenberg:2011:RCA

[IC11] Tobias Isenberg and Douglas Cunningham. Reports: Computational Aesthetics 2011 in Vancouver, Canada, August 5–7, 2011, Sponsored by Eurographics, in Collaboration with ACM SIGGRAPH. Computer Graphics Forum, 30(8):2457–2458, December 2011. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Isenberg:2010:RCA

[ID10] Tobias Isenberg and Neil Dodgson. Reports: Computational Aesthetics 2010 in London, England, June 14–15, 2010, sponsored by Eurographics, in collaboration with ACM SIG-GRAPH. Computer Graphics Forum, 29(6):2007, September 2010. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Inglis:2017:PDO

[IEGT17] T. Inglis, M.-L. Eckert, J. Gregson, and N. Thuerey. Primaldual optimization for fluids. Computer Graphics Forum, 36(8): 354–368, December 2017. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Iizuka:2014:ORB

[IEH^{+}14] S. Iizuka, Y. Endo, M. Hirose, Y. Kanamori, J. Mitani, and Y. Fukui. Object repositioning based on the perspective in a single image. Computer Graphics Forum, 33(8):157–166, December 2014. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Iizuka:2014:RDE

[IEK^{+}14] S. Iizuka, Y. Endo, Y. Kanamori, J. Mitani, and Y. Fukui. Reconstruction and depth: Efficient depth propagation for constructing a layered depth image from a single image. Computer Graphics Forum, 33(7):279–288, October 2014. CODEN CG-FODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Iizuka:2016:IGS

[IEKM16] Satoshi Iizuka, Yuki Endo, Yoshihiro Kanamori, and Jun Mitani. Images & geometry: Single image weathering via exemplar propagation. Computer Graphics Forum, 35(2):501–509, May 2016. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Iwasaki:2012:LRR

[IFDN12] Kei Iwasaki, Wataru Furuya, Yoshinori Dobashi, and Tomoyuki Nishita. Lighting and rendering: Real-time rendering of dynamic scenes under all-frequency lighting using integral spherical Gaussian. Computer Graphics Forum, 31(2pt3):727–734, May 2012. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Inger:2013:ILA

[IFL13] Yaron Inger, Zeev Farbman, and Dani Lischinski. Illumination: Locally adaptive products for all-frequency relighting. Computer Graphics Forum, 32(2pt1):73–82, May 2013. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Iglesias-Guitian:2015:SSB

[IGAJG15] Jose A. Iglesias-Guitian, Carlos Aliaga, Adrian Jarabo, and Diego Gutierrez. Sampling & skins: a biophysically-based model of the optical properties of skin aging. Computer Graphics Forum, 34(2):45–55, May 2015. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Iglesias-Guitian:2016:ERP

[IGMK^{+}16] Jose A. Iglesias-Guitian, Bochang Moon, Charalampos Koniaris, Eric Smolikowski, and Kenny Mitchell. Efficient rendering: Pixel history linear models for real-time temporal filtering. Computer Graphics Forum, 35(7):363–372, October 2016. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Ize:2011:RTR

[IH11] Thiago Ize and Charles Hansen. Ray tracing: RTSAH traversal order for occlusion rays. Computer Graphics Forum, 30(2):297– 305, April 2011. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Inoue:2019:LTE

[IIX^{+}19] N. Inoue, D. Ito, N. Xu, J. Yang, B. Price, and T. Yamasaki. Learning to trace: Expressive line drawing generation from photographs. Computer Graphics Forum, 38(7):69–80, October 2019. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Ihrke:2010:TSO

[IKL^{+}10] Ivo Ihrke, Kiriakos N. Kutulakos, Hendrik P. A. Lensch, Marcus Magnor, and Wolfgang Heidrich. Transparent and specular object reconstruction. Computer Graphics Forum, 29(8):2400– 2426, December 2010. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Ilcik:2015:CRL

[IMAW15] Martin Ilčík, Przemysław Musialski, Thomas Auzinger, and Michael Wimmer. Cities & roads: Layer-based procedural design of Façades. Computer Graphics Forum, 34(2):205–216, May 2015. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Iwasaki:2014:RIC

[IMDN14] K. Iwasaki, K. Mizutani, Y. Dobasbi, and T. Nisbita. Rendering: Interactive cloth rendering of microcylinder appearance model under environment lighting. Computer Graphics Forum, 33(2):333–340, May 2014. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Im:2013:FPG

[IPKK13] Jaeho Im, Hanwook Park, Jong-Hyun Kim, and Chang-Hun Kim. Fluids: a particle-grid method for opaque ice formation. Computer Graphics Forum, 32(2pt3):371–377, May 2013. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Innamorati:2017:RED

[IRWM17] Carlo Innamorati, Tobias Ritschel, Tim Weyrich, and Niloy J. Mitra. Rendering is everywhere: Decomposing single images for layered photo retouching. Computer Graphics Forum, 36 (4):15–25, July 2017. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Ilan:2015:SDD

[IS15] S. Ilan and A. Shamir. A survey on data-driven video completion. Computer Graphics Forum, 34(6):60–85, September 2015. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Iwamoto:2015:CML

[ISYM15] Naoya Iwamoto, Hubert P. H. Shum, Longzhi Yang, and Shigeo Morishima. Characters: Multi-layer lattice model for real-time dynamic character deformation. Computer Graphics Forum, 34 (7):99–109, October 2015. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Iwasaki:2010:SFP

[IUDN10] K. Iwasaki, H. Uchida, Y. Dobashi, and T. Nishita. Simulation: Fast particle-based visual simulation of ice melting. Computer Graphics Forum, 29(7):2215–2223, September 2010. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Ijiri:2010:IVI

[IY10] Takashi Ijiri and Hideo Yokota. Image and video II: Contourbased interface for refining volume segmentation. Computer Graphics Forum, 29(7):2153–2160, September 2010. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Ijiri:2013:VBH

[IYS^{+}13] Takashi Ijiri, Shin Yoshizawa, Yu Sato, Masaaki Ito, and Hideo Yokota. In volumes: Bilateral Hermite radial basis functions for contour-based volume segmentation. Computer Graphics Forum, 32(2pt1):123–132, May 2013. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Jarabo:2018:BRV

[JA18] Adrian Jarabo and Victor Arellano. Bidirectional rendering of vector light transport. Computer Graphics Forum, 37(6): 96–105, September 2018. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Jacobson:2017:DSG

[Jac17] Alec Jacobson. Design and segmentation: Generalized matryoshka: Computational design of nesting objects. Computer Graphics Forum, 36(5):27–35, August 2017. CODEN CG-FODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Jacobson:2019:RDA

[Jac19] Alec Jacobson. RodSteward: A design-to-assembly system for fabrication using 3d-printed joints and precision-cut rods. Computer Graphics Forum, 38(7):765–774, October 2019. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Joan-Arinyo:2010:GEG

[JAP10] Robert Joan-Arinyo and João Madeiras Pereira. Guest editorial: Guest editorial. Computer Graphics Forum, 29(6):1743– 1744, September 2010. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Jakob:2017:PPP

[JBG17] Johannes Jakob, Christoph Buchenau, and Michael Guthe. Parallel processing: a parallel approach to compression and decompression of triangle meshes using the GPU. Computer Graphics Forum, 36(5):71–80, August 2017. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Jakob:2019:PGC

[JBG19] J. Jakob, C. Buchenau, and M. Guthe. Parallel globally consistent normal orientation of raw unorganized point clouds. Computer Graphics Forum, 38(5):163–173, August 2019. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Janicke:2010:VSS

[JBMC10] H. Jänicke, R. Borgo, J. S. D. Mason, and M. Chen. Visualization: SoundRiver: Semantically-rich sound illustration. Computer Graphics Forum, 29(2):357–366, May 2010. CODEN CG-FODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Janicke:2010:ESB

[JC10] H. Jänicke and M. Chen. Evaluation: a salience-based quality metric for visualization. Computer Graphics Forum, 29(3): 1183–1192, June 2010. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Jeon:2013:ASC

[JCK^{+}13] Inyong Jeon, Kwang-Jin Choi, Tae-Yong Kim, Bong-Ouk Choi, and Hyeong-Seok Ko. Animation (session 1): Constrainable multigrid for cloth. Computer Graphics Forum, 32(7):31–39, October 2013. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Jeon:2016:VMC

[JCK16] Daniel S. Jeon, Inchang Choi, and Min H. Kim. Video: Multisampling compressive video spectroscopy. Computer Graphics Forum, 35(2):467–477, May 2016. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Jin:2014:IIR

[JCT14] Sou-Young Jin, Ho-Jin Choi, and Yu-Wing Tai. Images II: a randomized algorithm for natural object colorization. Computer Graphics Forum, 33(2):205–214, May 2014. CODEN CG-FODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Jeschke:2011:NPR

[JCW11] S. Jeschke, D. Cline, and P. Wonka. Non-photorealistic rendering: Estimating color and texture parameters for vector graphics. Computer Graphics Forum, 30(2):523–532, April 2011. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Javed:2013:IAE

[JE13] W. Javed and N. Elmqvist. Interactive analysis: ExPlates: Spatializing interactive analysis to scaffold visual exploration.

Computer Graphics Forum, 32(3pt4):441–450, June 2013. CO-DEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Jarosz:2019:OAS

[JEK^{+}19] Wojciech Jarosz, Afnan Enayet, Andrew Kensler, Charlie Kilpatrick, and Per Christensen. Orthogonal array sampling for Monte Carlo rendering. Computer Graphics Forum, 38(4):135– 147, July 2019. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Jeschke:2016:IEP

[Jes16] Stefan Jeschke. Image editing & processing: Generalized diffusion curves: an improved vector representation for smoothshaded images. Computer Graphics Forum, 35(2):71–79, May 2016. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Jimenez:2012:ICP

[JESG12] Jorge Jimenez, Jose I. Echevarria, Tiago Sousa, and Diego Gutierrez. Image and color processing: SMAA: Enhanced subpixel morphological antialiasing. Computer Graphics Forum, 31(2pt1):355–364, May 2012. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Janicke:2017:VTA

[JFCS17] S. Jänicke, G. Franzini, M. F. Cheema, and G. Scheuermann. Visual text analysis in digital humanities. Computer Graphics Forum, 36(6):226–250, September 2017. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Jendersie:2019:MMR

[JG19] Johannes Jendersie and Thorsten Grosch. Microfacet model regularization for robust light transport. Computer Graphics Forum, 38(4):39–47, July 2019. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Jia:2011:SNC

[JGH11] Yuntao Jia, Michael Garland, and John C. Hart. Social network clustering and visualization using hierarchical edge bundles. Computer Graphics Forum, 30(8):2314–2327, December 2011. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Jang:2012:FPD

[JH12] Hanyoung Jang and JungHyun Han. Feature-preserving displacement mapping with graphics processing unit (GPU) tessellation. Computer Graphics Forum, 31(6):1880–1894, September 2012. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Jankowski:2015:AIE

[JH15] J. Jankowski and M. Hachet. Advances in interaction with 3D environments. Computer Graphics Forum, 34(1):152–190, February 2015. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Jakob:2019:LDF

[JH19] Wenzel Jakob and Johannes Hanika. A low-dimensional function space for efficient spectral upsampling. Computer Graphics Forum, 38(2):147–155, May 2019. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Jin:2014:MRA

[JHT14] Y. Jin, J. Huang, and R. Tong. Maps: Remeshing-assisted optimization for locally injective mappings. Computer Graphics Forum, 33(5):269–279, August 2014. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Jeon:2018:SRR

[JJKL18] Junho Jeon, Jinwoong Jung, Jungeon Kim, and Seungyong Lee. Semantic reconstruction: Reconstruction of semantically segmented 3D meshes via volumetric semantic fusion. Computer Graphics Forum, 37(7):25–35, October 2018. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Jeong:2013:ASC

[JK13] SoHyeon Jeong and Chang-Hun Kim. Animation (session 2): Combustion waves on the point set surface. Computer Graphics Forum, 32(7):225–234, October 2013. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Jeruzalski:2018:CAO

[JKJL18] Timothy Jeruzalski, John Kanji, Alec Jacobson, and David I. W. Levin. Collision-aware and online compression of rigid body simulations via integrated error minimization. Computer Graphics Forum, 37(8):11–20, December 2018. CODEN CG-FODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Jung:2018:FIB

[JKK^{+}18] Y. Jung, J. Kim, A. Kumar, D. D. Feng, and M. Fulham. Feature of interest-based direct volume rendering using contextual saliency-driven ray profile analysis. Computer Graphics Forum, 37(6):5–19, September 2018. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Jeong:2013:RTD

[JKL13] Yuna Jeong, Kangtae Kim, and Sungkil Lee. Real-time defocus rendering with level of detail and sub-sample blur. Computer Graphics Forum, 32(6):126–134, September 2013. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Joo:2016:LTS

[JKL^{+}16] Hyuntae Joo, Soonhyeon Kwon, Sangmin Lee, Elmar Eisemann, and Sungkil Lee. Looking through surfaces: Efficient ray tracing through aspheric lenses and imperfect Bokeh synthesis. Computer Graphics Forum, 35(4):99–105, July 2016. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Jin:2018:AMM

[JKL18] Taeil Jin, Meekyoung Kim, and Sung-Hee Lee. Aura mesh: Motion retargeting to preserve the spatial relationships between skinned characters. Computer Graphics Forum, 37(2):311–320, May 2018. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Jankun-Kelly:2010:VTE

[JKLS10] T. J. Jankun-Kelly, Y. S. Lanka, and J. E. Swan II. Vectors & tensors: an evaluation of glyph perception for real symmetric traceless tensor properties. Computer Graphics Forum, 29(3): 1133–1142, June 2010. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Jang:2017:AGR

[JL17] Deok-Kyeong Jang and Sung-Hee Lee. Analyzing geometries: Regression-based landmark detection on dynamic human models. Computer Graphics Forum, 36(7):73–82, October 2017.

CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Jung:2018:HMS

[JL18] Seunghwan Jung and Sung-Hee Lee. Hair modeling and simulation by style. Computer Graphics Forum, 37(2):355–363, May 2018. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Jin:2019:BRL

[JL19] Sam Jin and Sung-Hee Lee. Book review: Lighting Layout Optimization for 3D Indoor Scenes. Computer Graphics Forum, 38(7):733–743, October 2019. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Jeon:2016:TMS

[JLKL16] Junho Jeon, Hyunjoon Lee, Henry Kang, and Seungyong Lee. Textures/mapping: Scale-aware structure-preserving texture filtering. Computer Graphics Forum, 35(7):77–86, October 2016. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Ji:2010:MBM

[JLW10] Zhongping Ji, Ligang Liu, and Yigang Wang. Modeling: Bmesh: a modeling system for base meshes of 3D articulated shapes. Computer Graphics Forum, 29(7):2169–2177, September 2010. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Jung:2015:GIR

[JMD15] Jin Woo Jung, Gary Meyer, and Ralph DeLong. Global illumination: Robust statistical pixel estimation. Computer Graphics Forum, 34(2):585–596, May 2015. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Jarabo:2015:RET

[JMV^{+}15] Adrian Jarabo, Belen Masia, Andreas Velten, Christopher Barsi, Ramesh Raskar, and Diego Gutierrez. Relativistic effects for time-resolved light transport. Computer Graphics Forum, 34(8):1–12, December 2015. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Jordao:2014:PMI

[JPCC14] K. Jordao, J. Pettré, M. Christie, and M.-P. Cani. Procedural Modeling II: Crowd sculpting: A space–time sculpting method for populating virtual environments. Computer Graphics Forum, 33(2):351–360, May 2014. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Jeong:2013:SMC

[JPK13] SoHyeon Jeong, Si-Hyung Park, and Chang-Hun Kim. Simulation of morphology changes in drying leaves. Computer Graphics Forum, 32(1):204–215, February 2013. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Joia:2015:HDV

[JPN15] P. Joia, F. Petronetto, and L. G. Nonato. High-dimensional visualization: Uncovering representative groups in multidimensional projections. Computer Graphics Forum, 34(3):281–290, June 2015. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Jones:2016:LTP

[JR16] Nathaniel L. Jones and Christoph F. Reinhart. Light transport: Parallel multiple-bounce irradiance caching. Computer Graphics Forum, 35(4):57–66, July 2016. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Jakob:2011:RAV

[JRJ11] Wenzel Jakob, Christian Regg, and Wojciech Jarosz. Realistic appearance and volumes: Progressive expectationmaximization for hierarchical volumetric photon mapping. Computer Graphics Forum, 30(4):1287–1297, June 2011. CO-DEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Janicke:2010:MCL

[JS10] H. Jänicke and G. Scheuermann. Measuring complexity in Lagrangian and Eulerian flow descriptions. Computer Graphics Forum, 29(6):1783–1794, September 2010. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Jeong:2013:CDP

[JSH^{+}13] W. Jeong, J. Schneider, A. Hansen, M. Lee, S. G. Turney, B. E. Faulkner-Jones, J. L. Hecht, R. Najarian, E. Yee, J. W. Lichtman, and H. Pfister. A collaborative digital pathology system for multi-touch mobile and desktop computing platforms. Computer Graphics Forum, 32(6):227–242, September 2013. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Ji:2014:RRT

[JSLW14] Zhongping Ji, Xianfang Sun, Shi Li, and Yigang Wang. Reconstruction: Real-time bas-relief generation from depth-and-normal maps on GPU. Computer Graphics Forum, 33(5):75– 83, August 2014. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Jonsson:2014:SVI

[JSYR14] Daniel Jönsson, Erik Sundén, Anders Ynnerman, and Timo Ropinski. A survey of volumetric illumination techniques for interactive volume rendering. Computer Graphics Forum, 33 (1):27–51, February 2014. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Jain:2012:IMU

[JTRS12] Arjun Jain, Thorsten Thormählen, Tobias Ritschel, and Hans-Peter Seidel. Interactive modeling and user interfaces: Exploring shape variations by 3D-model decomposition and partbased recombination. Computer Graphics Forum, 31(2pt3): 631–640, May 2012. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Jacobson:2010:FDO

[JTSZ10] Alec Jacobson, Elif Tosun, Olga Sorkine, and Denis Zorin. FEM and differential operators: Mixed finite elements for variational surface modeling. Computer Graphics Forum, 29(5):1565–1574, July 2010. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Ji:2019:VED

[JvdGMR19] C. Ji, J. J. van de Gronde, N. M. Maurits, and J. B. T. M. Roerdink. Visual exploration of dynamic multichannel EEG coherence networks. Computer Graphics Forum, 38(1):507–520, February 2019. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Jarabo:2012:PGC

[JVS^{+}12] Adrian Jarabo, Tom Van Eyck, Veronica Sundstedt, Kavita Bala, Diego Gutierrez, and Carol O’Sullivan. Perceptual graphics: Crowd light: Evaluating the perceived fidelity of illuminated dynamic scenes. Computer Graphics Forum, 31(2pt3): 565–574, May 2012. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Janicke:2011:FVV

[JWC^{+}11] Heike Jänicke, Thomas Weidner, David Chung, Robert S. Laramee, Peter Townsend, and Min Chen. Flow visualization: Visual reconstructability as a quality metric for flow visualization. Computer Graphics Forum, 30(3):781–790, June 2011. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Jankowai:2019:RES

[JWH19a] Jochen Jankowai, Bei Wang, and Ingrid Hotz. Robust extraction and simplification of 2D symmetric tensor field topology. Computer Graphics Forum, 38(3):337–349, June 2019. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Jung:2019:WGS

[JWH^{+}19b] A. Jung, A. Wilkie, J. Hanika, W. Jakob, and C. Dachsbacher. Wide gamut spectral upsampling with fluorescence. Computer Graphics Forum, 38(4):87–96, July 2019. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Ju:2013:IPS

[JWL^{+}13] Jinlong Ju, Jue Wang, Yebin Liu, Haoqian Wang, and Qionghai Dai. Image processing (session 4): a progressive tri-level segmentation approach for topology-change-aware video matting. Computer Graphics Forum, 32(7):245–253, October 2013. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Jacobson:2012:DSS

[JWS12] Alec Jacobson, Tino Weinkauf, and Olga Sorkine. Deformation: Smooth shape-aware functions with controlled extrema. Computer Graphics Forum, 31(5):1577–1586, August 2012. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Jiang:2014:MFH

[JWWP14] Caigui Jiang, Jun Wang, Johannes Wallner, and Helmut Pottmann. Meshing: Freeform honeycomb structures. Computer Graphics Forum, 33(5):185–194, August 2014. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Jeong:2019:SPC

[JYS19] Taehong Jeong, Myunghyun Yang, and Hyun Joon Shin. Succinct palette and color model generation and manipulation using hierarchical representation. Computer Graphics Forum, 38 (7):1–10, October 2019. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Jimenez:2015:SSS

[JZJ^{+}15] Jorge Jimenez, Károly Zsolnai, Adrian Jarabo, Christian Freude, Thomas Auzinger, Xian-Chun Wu, Javier von der Pahlen, Michael Wimmer, and Diego Gutierrez. Separable subsurface scattering. Computer Graphics Forum, 34(6):188–197, September 2015. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Jinsp:2014:AID

[JZW14] Shuo Jinsp, Yunbo Zhang, and Charlie C. L. Wang. Animation III: Deformation with enforced metrics on length, area and volume. Computer Graphics Forum, 33(2):429–438, May 2014. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Jia:2018:CGP

[JZYP18] Shiyu Jia, Weizhong Zhang, Xiaokang Yu, and Zhenkuan Pan. CPU–GPU parallel framework for real-time interactive cutting of adaptive octree-based deformable objects. Computer Graphics Forum, 37(1):45–59, February 2018. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Kanjanabose:2015:HDV

[KARC15] Rassadarie Kanjanabose, Alfie Abdul-Rahman, and Min Chen. High-dimensional visualization: a multi-task comparative study on scatter plots and parallel coordinates plots. Computer Graphics Forum, 34(3):261–270, June 2015. CODEN CG-FODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Kijmongkolchai:2017:EMS

[KARC17] Natchaya Kijmongkolchai, Alfie Abdul-Rahman, and Min Chen. Empirically measuring soft knowledge in visualization. Computer Graphics Forum, 36(3):73–85, June 2017. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Koulieris:2019:NED

[KAS^{+}19] G. A. Koulieris, K. Akşit, M. Stengel, R. K. Mantiuk, K. Mania, and C. Richardt. Near-eye display and tracking technologies for virtual and augmented reality. Computer Graphics Forum, 38 (2):493–519, May 2019. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Kratz:2013:VAS

[KASH13] A. Kratz, C. Auer, M. Stommel, and I. Hotz. Visualization and analysis of second-order tensors: Moving beyond the symmetric positive-definite case. Computer Graphics Forum, 32(1):49–74, February 2013. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Kim:2019:DFG

[KAT^{+}19] Byungsoo Kim, Vinicius C. Azevedo, Nils Thuerey, Theodore Kim, Markus Gross, and Barbara Solenthaler. Deep fluids: a generative network for parameterized fluid simulations. Computer Graphics Forum, 38(2):59–70, May 2019. CODEN CG-FODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Kazhdan:2015:NMG

[Kaz15] M. Kazhdan. Numerical methods for geometry processing: Fast and exact (Poisson) solvers on symmetric geometries. Computer Graphics Forum, 34(5):153–165, August 2015. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Kuhnert:2012:IHS

[KB12] T. Kühnert and G. Brunnett. Imaging and hair: Fur shading and modification based on cone step mapping. Computer Graphics Forum, 31(7pt1):2011–2018, September 2012. CO-DEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Kovnatsky:2013:SCC

[KBB^{+}13] A. Kovnatsky, M. M. Bronstein, A. M. Bronstein, K. Glashoff, and R. Kimmel. Shape correspondences: Coupled quasiharmonic bases. Computer Graphics Forum, 32(2pt4):439–448, May 2013. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Kozlov:2017:EFB

[KBB^{+}17] Yeara Kozlov, Derek Bradley, Moritz Bächer, Bernhard Thomaszewski, Thabo Beeler, and Markus Gross. Enriching facial blendshape rigs with physical simulation. Computer Graphics Forum, 36(2):75–84, May 2017. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Kronander:2015:SAR

[KBG^{+}15] Joel Kronander, Francesco Banterle, Andrew Gardner, Ehsan Miandji, and Jonas Unger. State of the art reports: Photorealistic rendering of mixed reality scenes. Computer Graphics Forum, 34(2):643–665, May 2015. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Kerzner:2015:EPS

[KBHM15] Ethan Kerzner, Lee A. Butler, Charles Hansen, and Miriah Meyer. Engineering and physical sciences: a shot at visual vulnerability analysis. Computer Graphics Forum, 34(3):391– 400, June 2015. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Kim:2010:RTC

[KBK^{+}10] Tae-Joon Kim, Yongyoung Byun, Yongjin Kim, Bochang Moon, Seungyong Lee, and Sung-Eui Yoon. Ray tracing and culling: HCCMeshes: Hierarchical-culling oriented compact meshes. Computer Graphics Forum, 29(2):299–308, May 2010. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Krecklau:2013:RVD

[KBK13] Lars Krecklau, Janis Born, and Leif Kobbelt. Rendering: Viewdependent realtime rendering of procedural facades with high geometric detail. Computer Graphics Forum, 32(2pt4):479–488, May 2013. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Kolb:2010:TFC

[KBKL10] A. Kolb, E. Barth, R. Koch, and R. Larsen. Time-of-flight cameras in computer graphics. Computer Graphics Forum, 29 (1):141–159, March 2010. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Kappe:2019:ADC

[KBL19] C. P. Kappe, M. Böttinger, and H. Leitte. Analysis of decadal climate predictions with user-guided hierarchical ensemble clustering. Computer Graphics Forum, 38(3):505–515, June 2019. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Kol:2019:MSM

[KBLE19] Timothy R. Kol, Pablo Bauszat, Sungkil Lee, and Elmar Eisemann. MegaViews: Scalable many-view rendering with concurrent scene-view hierarchy traversal. Computer Graphics Forum, 38(1):235–247, February 2019. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Kuster:2014:GAR

[KBÖ^{+}14] Claudia Kuster, Jean-Charles Bazin, Cengiz Öztireli, Teng Deng, Tobias Martin, Tiberiu Popa, and Markus Gross. Geometry acquisition, reconstruction and analysis: Spatio-temporal geometry fusion for multiple hybrid cameras using moving least squares surfaces. Computer Graphics Forum, 33(2):1–10, May 2014. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Kalojanov:2011:RTT

[KBS11a] Javor Kalojanov, Markus Billeter, and Philipp Slusallek. Ray tracing: Two-level grids for ray tracing on GPUs. Computer Graphics Forum, 30(2):307–314, April 2011. CODEN CG-FODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Kavan:2011:ETL

[KBS11b] L. Kavan, A. W. Bargteil, and P.-P. Sloan. Efficient textures: Least squares vertex baking. Computer Graphics Forum, 30(4): 1319–1326, June 2011. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Kretschmer:2012:MVR

[KBT^{+}12] Jan Kretschmer, Thomas Beck, Christian Tietjen, Bernhard Preim, and Marc Stamminger. Medical visualization: Reliable adaptive modelling of vascular structures with non-circular cross-sections. Computer Graphics Forum, 31(3pt2):1055–1064, June 2012. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Kohler:2017:SCP

[KBvP^{+}17] Benjamin Köhler, Silvia Born, Roy F. P. van Pelt, Anja Hennemuth, Uta Preim, and Bernhard Preim. A survey of cardiac 4D PC–MRI data processing. Computer Graphics Forum, 36 (6):5–35, September 2017. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Kalojanov:2012:CME

[KBW^{+}12] Javor Kalojanov, Martin Bokeloh, Michael Wand, Leonidas Guibas, Hans-Peter Seidel, and Philipp Slusallek. Correspondence: Microtiles: Extracting building blocks from correspondences. Computer Graphics Forum, 31(5):1597–1606, August 2012. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Kerber:2013:SSD

[KBWS13] J. Kerber, M. Bokeloh, M. Wand, and H.-P. Seidel. Scalable symmetry detection for urban scenes. Computer Graphics Forum, 32(1):3–15, February 2013. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Ko:2016:ASV

[KCA^{+}16] S. Ko, I. Cho, S. Afzal, C. Yau, J. Chae, A. Malik, K. Beck, Y. Jang, W. Ribarsky, and D. S. Ebert. Applications: a survey on visual analysis approaches for financial data. Computer Graphics Forum, 35(3):599–617, June 2016. CODEN CG-FODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Kindlmann:2018:REE

[KCH^{+}18] G. Kindlmann, C. Chiw, T. Huynh, A. Gyulassy, J. Reppy, and P.-T. Bremer. Rendering and extracting extremal features in 3D fields. Computer Graphics Forum, 37(3):525–536, June 2018. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Kim:2019:DAT

[KCH19] Younghoon Kim, Michael Correll, and Jeffrey Heer. Designing animated transitions to convey aggregate operations. Computer Graphics Forum, 38(3):541–551, June 2019. CODEN CG-FODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Kamaleswaran:2016:TSD

[KCJM16] Rishikesan Kamaleswaran, Christopher Collins, Andrew James, and Carolyn McGregor. Time series data and sequences: PhysioEx: Visual analysis of physiological event streams. Computer Graphics Forum, 35(3):331–340, June 2016. CODEN CG-FODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Kim:2018:PPA

[KCL^{+}18] Minjeong Kim, Minsuk Choi, Sunwoong Lee, Jian Tang, Haesun Park, and Jaegul Choo. PixelSNE: Pixel-aligned stochastic neighbor embedding for efficient 2D visualization with screenresolution precision. Computer Graphics Forum, 37(3):267–276, June 2018. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Kaplanyan:2013:IPS

[KD13] Anton S. Kaplanyan and Carsten Dachsbacher. Illumination: Path space regularization for holistic and robust light transport. Computer Graphics Forum, 32(2pt1):63–72, May 2013. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Kolar:2017:SET

[KDC17] M. Kolář, K. Debattista, and A. Chalmers. A subjective evaluation of texture synthesis methods. Computer Graphics Forum, 36(2):189–198, May 2017. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Koulieris:2014:GRC

[KDCM14] G. A. Koulieris, G. Drettakis, D. Cunningham, and K. Mania. GPU rendering: C-LOD: Context-aware material level-of-detail applied to mobile graphics. Computer Graphics Forum, 33(4): 41–49, July 2014. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Kuhn:2014:TLC

[KER^{+}14] A. Kuhn, W. Engelke, C. Rössl, M. Hadwiger, and H. Theisel. Time line cell tracking for the approximation of Lagrangian coherent structures with subgrid accuracy. Computer Graphics Forum, 33(1):222–234, February 2014. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Kulla:2012:SIS

[KF12] Christopher Kulla and Marcos Fajardo. Sampling: Importance sampling techniques for path tracing in participating media. Computer Graphics Forum, 31(4):1519–1528, June 2012. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Kravtsov:2010:EIS

[KFA^{+}10] D. Kravtsov, O. Fryazinov, V. Adzhiev, A. Pasko, and P. Comninos. Embedded implicit stand-ins for animated meshes: a case of hybrid modelling. Computer Graphics Forum, 29(1):128–140, March 2010. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Kravtsov:2014:CMB

[KFA^{+}14] Denis Kravtsov, Oleg Fryazinov, Valery Adzhiev, Alexander Pasko, and Peter Comninos. Controlled metamorphosis between skeleton-driven animated polyhedral meshes of arbitrary topologies. Computer Graphics Forum, 33(1):64–72, February 2014. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Kehrer:2010:MDB

[KFH10] Johannes Kehrer, Peter Filzmoser, and Helwig Hauser. Multivariate data: Brushing moments in interactive visual analysis. Computer Graphics Forum, 29(3):813–822, June 2010. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Kleiman:2013:SCM

[KFLCO13] Yanir Kleiman, Noa Fish, Joel Lanir, and Daniel Cohen-Or. Shape collections and maps: Dynamic maps for exploring and browsing shapes. Computer Graphics Forum, 32(5):187–196, August 2013. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Kottravel:2019:VAC

[KFM^{+}19] S. Kottravel, M. Falk, T. Bin Masood, M. Linares, and I. Hotz. Visual analysis of charge flow networks for complex morphologies. Computer Graphics Forum, 38(3):479–489, June 2019. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Krone:2011:MDV

[KFR^{+}11] M. Krone, M. Falk, S. Rehm, J. Pleiss, and T. Ertl. Molecular data visualization: Interactive exploration of protein cavities. Computer Graphics Forum, 30(3):673–682, June 2011. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Kreiser:2018:VSM

[KFR18] J. Kreiser, J. Freedman, and T. Ropinski. Visually supporting multiple needle placement in irreversible electroporation interventions. Computer Graphics Forum, 37(6):59–71, September 2018. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Karasik:2019:OPS

[KFW19] E. Karasik, R. Fattal, and M. Werman. Object partitioning for support-free 3D-printing. Computer Graphics Forum, 38 (2):305–316, May 2019. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Koyama:2018:DIL

[KG18] Yuki Koyama and Masataka Goto. Decomposing images into layers with advanced color blending. Computer Graphics Forum, 37(7):397–407, October 2018. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Kim:2019:RRF

[KG19] Byungsoo Kim and Tobias Günther. Robust reference frame extraction from unsteady 2D vector fields with convolutional neural networks. Computer Graphics Forum, 38(3):285–295, June 2019. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Kim:2015:ICC

[KGAC15] Jongdae Kim, Charles Gray, Paul Asente, and John Collomosse. Image collections: Comprehensible video thumbnails.

Computer Graphics Forum, 34(2):167–177, May 2015. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Kohler:2018:VQA

[KGGP18] B. Köhler, M. Grothoff, M. Gutberlet, and B. Preim. Visual and quantitative analysis of great arteries’ blood flow jets in cardiac 4D PC-MRI data. Computer Graphics Forum, 37 (3):195–204, June 2018. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Kneiphof:2019:RTI

[KGK19] Tom Kneiphof, Tim Golla, and Reinhard Klein. Real-time image-based lighting of microfacet BRDFs with varying iridescence. Computer Graphics Forum, 38(4):77–85, July 2019. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Kuss:2010:FLE

[KGM^{+}10] A. Kuß, M. Gensel, B. Meyer, V. J. Dercksen, and S. Prohaska. Fiber & lines: Effective techniques to visualize filamentsurface relationships. Computer Graphics Forum, 29(3):1003– 1012, June 2010. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Klein:2012:IDU

[KGP^{+}12] Tijmen Klein, Florimond Guéniat, Luc Pastur, Frédéric Vernier, and Tobias Isenberg. Interaction design and usability: a design study of direct-touch interaction for exploratory 3D scientific visualization. Computer Graphics Forum, 31(3pt3): 1225–1234, June 2012. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Krajcevski:2016:TCV

[KGR^{+}16] P. Krajcevski, A. Golas, K. Ramani, M. Shebanow, and D. Manocha. Texturing & compression: VBTC: GPU-friendly variable block size texture encoding. Computer Graphics Forum, 35(2):409–418, May 2016. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Kumar:2017:ZAA

[KGRG17] Moneish Kumar, Vineet Gandhi, Remi Ronfard, and Michael Gleicher. Zooming on all actors: Automatic focus+context split screen video generation. Computer Graphics Forum, 36(2):455– 465, May 2017. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Kim:2018:AET

[KH18] Younghoon Kim and Jeffrey Heer. Assessing effects of task and data distribution on the effectiveness of visual encodings. Computer Graphics Forum, 37(3):157–167, June 2018. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Kazhdan:2019:AMG

[KH19] Misha Kazhdan and Hugues Hoppe. An adaptive multi-grid solver for applications in computer graphics. Computer Graphics Forum, 38(1):138–150, February 2019. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Krueger:2019:BEL

[KHI^{+}19] Robert Krueger, Qi Han, Nikolay Ivanov, Sanae Mahtal, Dennis Thom, Hanspeter Pfister, and Thomas Ertl. Bird’s-eye — large-scale visual analytics of city dynamics using social location data. Computer Graphics Forum, 38(3):595–607, June 2019. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Khoury:2012:LGD

[KHKS12] Marc Khoury, Yifan Hu, Shankar Krishnan, and Carlos Scheidegger. Large graphs: Drawing large graphs by low-rank stress majorization. Computer Graphics Forum, 31(3pt1):975–984, June 2012. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Kurzhals:2013:EEA

[KHW13] K. Kurzhals, M. Höferlin, and D. Weiskopf. Evaluation: Evaluation of attention-guiding video visualization. Computer Graphics Forum, 32(3pt1):51–60, June 2013. CODEN CG-FODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Kim:2015:NMG

[Kim15] Theodore Kim. Numerical methods for geometry processing: Quaternion Julia set shape optimization. Computer Graphics Forum, 34(5):167–176, August 2015. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Kustra:2014:RSM

[KJT14] J. Kustra, A. Jalba, and A. Telea. Robust segmentation of multiple intersecting manifolds from unoriented noisy point clouds. Computer Graphics Forum, 33(1):73–87, February 2014. CO-DEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Krecklau:2011:PMP

[KK11a] Lars Krecklau and Leif Kobbelt. Procedural modelling: Procedural modeling of interconnected structures. Computer Graphics Forum, 30(2):335–344, April 2011. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Kyprianidis:2011:IVI

[KK11b] Jan Eric Kyprianidis and Henry Kang. Image and video: Image and video abstraction by coherence-enhancing filtering. Computer Graphics Forum, 30(2):593–602, April 2011. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Kohout:2014:RTM

[KK14] J. Kohout and M. Kukačka. Real-time modelling of fibrous muscle. Computer Graphics Forum, 33(8):1–15, December 2014. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Kerracher:2017:CEV

[KK17] N. Kerracher and J. Kennedy. Constructing and evaluating visualisation task classifications: Process and considerations. Computer Graphics Forum, 36(3):47–59, June 2017. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Kim:2018:PAP

[KK18] Goanghun Kim and Hyeong-Seok Ko. A practical approach to Physically-based reproduction of diffusive cosmetics. Computer Graphics Forum, 37(7):223–232, October 2018. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Kugelstadt:2018:FCF

[KKB18] T. Kugelstadt, D. Koschier, and J. Bender. Fast corotated FEM using operator splitting. Computer Graphics Forum, 37 (8):149–160, December 2018. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Kanazawa:2016:DSO

[KKBJ16] Angjoo Kanazawa, Shahar Kovalsky, Ronen Basri, and David Jacobs. Deformable & soft objects: Learning 3D deformation of animals from 2D images. Computer Graphics Forum, 35 (2):365–374, May 2016. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Kezurer:2015:CMT

[KKBL15] Itay Kezurer, Shahar Z. Kovalsky, Ronen Basri, and Yaron Lipman. Correspondence and matching: Tight relaxation of quadratic matching. Computer Graphics Forum, 34(5):115– 128, August 2015. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Kozlikova:2017:VBS

[KKF^{+}17] B. Kozlíková, M. Krone, M. Falk, N. Lindow, M. Baaden, D. Baum, I. Viola, J. Parulek, and H.-C. Hege. Visualization of biomolecular structures: State of the art revisited. Computer Graphics Forum, 36(8):178–204, December 2017. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Kim:2018:DDO

[KKK18] Yeojin Kim, Byungmoon Kim, and Young J. Kim. Dynamic deep octree for high-resolution volumetric painting in virtual reality. Computer Graphics Forum, 37(7):179–190, October 2018. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Kim:2016:IPI

[KKL16a] Hye-Rin Kim, Henry Kang, and In-Kwon Lee. Image processing: Image recoloring with valence-arousal emotion model. Computer Graphics Forum, 35(7):209–216, October 2016. CO-DEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Krone:2016:BVV

[KKL^{+}16b] M. Krone, B. Kozlíková, N. Lindow, M. Baaden, D. Baum, J. Parulek, H.-C. Hege, and I. Viola. Biomedical visualization: Visual analysis of biomolecular cavities: State of the art. Computer Graphics Forum, 35(3):527–551, June 2016. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Kim:2019:GTM

[KKPL19] Jungeon Kim, Hyomin Kim, Jaesik Park, and Seungyong Lee. Global texture mapping for dynamic objects. Computer Graphics Forum, 38(7):697–705, October 2019. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Kuznetsov:2018:DAS

[KKR18] Alexandr Kuznetsov, Nima Khademi Kalantari, and Ravi Ramamoorthi. Deep adaptive sampling for low sample count rendering. Computer Graphics Forum, 37(4):35–44, July 2018. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Khlebnikov:2012:MDM

[KKS^{+}12] R. Khlebnikov, B. Kainz, M. Steinberger, M. Streit, and D. Schmalstieg. Multi-dimensional and multi-variate data: Procedural texture synthesis for zoom-independent visualization of multivariate data. Computer Graphics Forum, 31(3pt4):1355– 1364, June 2012. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Kerbl:2017:HBQ

[KKS^{+}17] Bernhard Kerbl, Michael Kenzel, Dieter Schmalstieg, Hans-Peter Seidel, and Markus Steinberger. Hierarchical bucket queuing for fine-grained priority scheduling on the GPU. Computer Graphics Forum, 36(8):232–246, December 2017. CO-DEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Kerbl:2015:SMI

[KKSS15] Bernhard Kerbl, Denis Kalkofen, Markus Steinberger, and Dieter Schmalstieg. Shape manipulation: Interactive disassembly planning for complex objects. Computer Graphics Forum, 34 (2):287–297, May 2015. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Kister:2017:GCS

[KKTD17] U. Kister, K. Klamka, C. Tominski, and R. Dachselt. GraSp: Combining spatially-aware mobile devices and a display wall for graph visualization and interaction. Computer Graphics Forum, 36(3):503–514, June 2017. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Kang:2014:CAE

[KL14] Changgu Kang and Sung-Hee Lee. Character animation: Environment-adaptive contact poses for virtual characters. Computer Graphics Forum, 33(7):1–10, October 2014. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Kim:2019:RUF

[KL19] Jong-Hyun Kim and Jung Lee. Retracted: Unified framework for efficient and enriched water animation with surface and wave foams. Computer Graphics Forum, 38(1):690, February 2019. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Korman:2015:PAS

[KLAB15] Simon Korman, Roee Litman, Shai Avidan, and Alex Bronstein. Probably approximately symmetric: Fast rigid symmetry detection with global guarantees. Computer Graphics Forum, 34(1):2–13, February 2015. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Kuo:2015:SPC

[KLC^{+}15] Ming-Hsun Kuo, You-En Lin, Hung-Kuo Chu, Ruen-Rone Lee, and Yong-Liang Yang. Stylization: Pixel 2Brick: Constructing brick sculptures from pixel art. Computer Graphics Forum, 34 (7):339–348, October 2015. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Kim:2010:SMT

[KLCF10] Vladimir G. Kim, Yaron Lipman, Xiaobai Chen, and Thomas Funkhouser. Symmetry: Möbius transformations for global intrinsic symmetry analysis. Computer Graphics Forum, 29(5): 1689–1700, July 2010. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Kong:2017:IEV

[KLK17] Ha-Kyung Kong, Zhicheng Liu, and Karrie Karahalios. Internal and external visual cue preferences for visualizations in presentations. Computer Graphics Forum, 36(3):515–525, June 2017. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Kalojanov:2019:SBS

[KLMK19] Javor Kalojanov, Isaak Lim, Niloy Mitra, and Leif Kobbelt. String-based synthesis of structured shapes. Computer Graphics Forum, 38(2):27–36, May 2019. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Kerzner:2017:GVC

[KLS^{+}17] E. Kerzner, A. Lex, C. L. Sigulinsky, T. Urness, B. W. Jones, R. E. Marc, and M. Meyer. Graffinity: Visualizing connectivity in large graphs. Computer Graphics Forum, 36(3):251–260, June 2017. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Kermani:2016:SLS

[KLTZ16] Z. Sadeghipour Kermani, Z. Liao, P. Tan, and H. Zhang. Structures: Learning 3D scene synthesis from annotated RGB-D images. Computer Graphics Forum, 35(5):197–206, August 2016. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Kaufman:2012:CAA

[KLW12] Liad Kaufman, Dani Lischinski, and Michael Werman. Content-aware automatic photo enhancement. Computer Graphics Forum, 31(8):2528–2540, December 2012. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Kita:2016:VNA

[KM16] N. Kita and K. Miyata. Visualization/NPR: Aesthetic rating and color suggestion for color palettes. Computer Graphics Forum, 35(7):127–136, October 2016. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Klein:2019:PGV

[KMA^{+}19] T. Klein, P. Mindek, L. Autin, D. S. Goodsell, A. J. Olson, E. M. Gröller, and I. Viola. Parallel generation and visualization of bacterial genome structures. Computer Graphics Forum, 38(7):57–68, October 2019. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Karimov:2015:VAC

[KMAB15] A. Karimov, G. Mistelbauer, T. Auzinger, and S. Bruckner. Volume analysis and classification: Guided volume editing based on histogram dissimilarity. Computer Graphics Forum, 34(3):91–100, June 2015. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Klaudiny:2017:RTM

[KMB^{+}17] Martin Klaudiny, Steven McDonagh, Derek Bradley, Thabo Beeler, and Kenny Mitchell. Real-time multi-view facial capture with synthetic training. Computer Graphics Forum, 36 (2):325–336, May 2017. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Khalil:2017:SFP

[KMD^{+}17] J. El Sayeh Khalil, A. Munteanu, L. Denis, P. Lambert, and R. Van de Walle. Scalable feature-preserving irregular mesh coding. Computer Graphics Forum, 36(6):275–290, September 2017. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Kim:2013:GSG

[KMHG13] Young Min Kim, Niloy J. Mitra, Qixing Huang, and Leonidas Guibas. Geometry (session 1): Guided real-time scanning of indoor objects. Computer Graphics Forum, 32(7):177–186, October 2013. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Ko:2012:VDS

[KMJE12] S. Ko, R. Maciejewski, Y. Jang, and D. S. Ebert. Visualization for decision support: MarketAnalyzer: An interactive visual analytics system for analyzing competitive advantage using point of sale data. Computer Graphics Forum, 31(3pt3):1245– 1254, June 2012. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Krecklau:2012:CBP

[KMK12] Lars Krecklau, Christopher Manthei, and Leif Kobbelt. Cities and buildings: Procedural interpolation of historical city maps. Computer Graphics Forum, 31(2pt3):691–700, May 2012. CO-DEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Kreiser:2018:SFB

[KMM^{+}18] J. Kreiser, M. Meuschke, G. Mistelbauer, B. Preim, and T. Ropinski. A survey of flattening-based medical visualization techniques. Computer Graphics Forum, 37(3):597–624, June 2018. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Karimov:2013:VVS

[KMS^{+}13] A. Karimov, G. Mistelbauer, J. Schmidt, P. Mindek, E. Schmidt, T. Sharipov, S. Bruckner, and E. Gröller. Volumes: ViviSection: Skeleton-based volume editing. Computer Graphics Forum, 32(3pt4):461–470, June 2013. CODEN CG-FODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Kratt:2018:SGS

[KNH^{+}18] J. Kratt, T. Niese, R. Hu, H. Huang, S. Pirk, A. Sharf, D. Cohen-Or, and O. Deussen. Sketching in gestalt space: Interactive shape abstraction through perceptual reasoning. Computer Graphics Forum, 37(6):188–204, September 2018. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Kaspar:2015:CTS

[KNL^{+}15] Alexandre Kaspar, Boris Neubert, Dani Lischinski, Mark Pauly, and Johannes Kopf. Colors and textures: Self tuning texture optimization. Computer Graphics Forum, 34(2):349–359, May 2015. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Kleiman:2019:RSB

[KO19] Yanir Kleiman and Maks Ovsjanikov. Robust structure-based shape correspondence. Computer Graphics Forum, 38(1):7–20, February 2019. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Kelly:2013:MSS

[KÓOH13] P. Kelly, C. Ó Conaire, N. E. O’Connor, and J. Hodgins. Motion synthesis for sports using unobtrusive lightweight body-worn and environment sensing. Computer Graphics Forum, 32(8): 48–60, December 2013. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Kim:2015:AAF

[KÖS^{+}15] Hyeon-Joong Kim, A. Cengiz Öztireli, Il-Kyu Shin, Markus Gross, and Soo-Mi Choi. All about faces: Interactive generation of realistic facial wrinkles from sketchy drawings. Computer Graphics Forum, 34(2):179–191, May 2015. CODEN CG-FODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Karciauskas:2011:SRB

[KP11] Kȩstutis Karčiauskas and Jörg Peters. Section 1: Rational bicubic G^{2} splines for design with basic shapes. Computer Graphics Forum, 30(5):1389–1395, August 2011. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Karciauskas:2015:QPC

[KP15] KȨstutis Karčiauskas and Jörg Peters. Quads and polygons: Can bi-cubic surfaces be class A? Computer Graphics Forum, 34(5):229–238, August 2015. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Karciauskas:2018:NCG

[KP18] Kȩstutis Karčiauskas and Jörg Peters. A new class of guided C^{2} subdivision surfaces combining good shape with nested refinement. Computer Graphics Forum, 37(6):84–95, September 2018. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Karciauskas:2019:HQR

[KP19] K. Karčiauskas and J. Peters. High quality refinable G-splines for locally quad-dominant meshes with T-gons. Computer Graphics Forum, 38(5):151–161, August 2019. CODEN CG-FODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Kerr:2010:ESM

[KPD10] William B. Kerr, Fabio Pellacini, and Jonathan D. Denning. Editing of shadows and materials: BendyLights: Artistic control of direct illumination by curving light rays. Computer Graphics Forum, 29(4):1451–1459, June 2010. CODEN CG-FODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Kohler:2016:RCF

[KPG^{+}16] Benjamin Köhler, Uta Preim, Matthias Grothoff, Matthias Gutberlet, Katharina Fischbach, and Bernhard Preim. Robust cardiac function assessment in 4D PC–MRI data of the aorta and pulmonary artery. Computer Graphics Forum, 35 (1):32–43, February 2016. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Krecklau:2010:GUN

[KPK10] L. Krecklau, D. Pavic, and L. Kobbelt. Generalized use of nonterminal symbols for procedural modeling. Computer Graphics Forum, 29(8):2291–2303, December 2010. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Kucher:2018:SAS

[KPK18] Kostiantyn Kucher, Carita Paradis, and Andreas Kerren. The state of the art in sentiment visualization. Computer Graphics Forum, 37(1):71–96, February 2018. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Kang:2010:FSH

[KPNS10] Nahyup Kang, Jinho Park, Junyong Noh, and Sung Yong Shin. Fluids and smoke: a hybrid approach to multiple fluid simulation using volume fractions. Computer Graphics Forum, 29 (2):685–694, May 2010. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Kabul:2011:OCA

[KPRN11] Ilknur Kabul, Stephen M. Pizer, Julian Rosenman, and Marc Niethammer. An optimal control approach for texture metamorphosis. Computer Graphics Forum, 30(8):2341–2353, December 2011. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Koerner:2014:FLD

[KPS^{+}14] D. Koerner, J. Portsmouth, F. Sadlo, T. Ertl, and B. Eberhardt. Flux-limited diffusion for multiple scattering in participating media. Computer Graphics Forum, 33(6):178–189, September 2014. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Kakar:2019:DEV

[KQR^{+}19] T. Kakar, X. Qin, E. A. Rundensteiner, L. Harrison, S. K. Sahoo, and S. De. DIVA: Exploration and validation of hypothesized drug-drug interactions. Computer Graphics Forum, 38(3):95–106, June 2019. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Kalantari:2019:DHV

[KR19] Nima Khademi Kalantari and Ravi Ramamoorthi. Deep HDR video from sequences with alternating exposures. Computer Graphics Forum, 38(2):193–205, May 2019. CODEN CG-FODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Kider:2011:NPF

[KRB11] Joseph T. Kider, Jr., Samantha Raja, and Norman I. Badler. Natural phenomena: Fruit senescence and decay simulation. Computer Graphics Forum, 30(2):257–266, April 2011. CO-DEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Kairam:2015:GRV

[KRD^{+}15] S. Kairam, N. H. Riche, S. Drucker, R. Fernandez, and J. Heer. Graphs: Refinery: Visual exploration of large, heterogeneous networks through associative browsing. Computer Graphics Forum, 34(3):301–310, June 2015. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Kim:2011:IVE

[KrJC^{+}11] Younghui Kim, Hwi ryong Jung, Sungwoo Choi, Jungjin Lee, and Junyong Noh. Image & video editing: a single image representation model for efficient stereoscopic image creation. Computer Graphics Forum, 30(7):2067–2076, September 2011. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Kellnhofer:2015:PSM

[KRM^{+}15] Petr Kellnhofer, Tobias Ritschel, Karol Myszkowski, Elmar Eisemann, and Hans-Peter Seidel. Perception and simplification: Modeling luminance perception at absolute threshold. Computer Graphics Forum, 34(4):155–164, July 2015. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Kruiger:2017:GLS

[KRM^{+}17] J. F. Kruiger, P. E. Rauber, R. M. Martins, A. Kerren, S. Kobourov, and A. C. Telea. Graph layouts by t-SNE. Computer Graphics Forum, 36(3):283–294, June 2017. CODEN CG-FODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Kellnhofer:2013:EPO

[KRMS13] Petr Kellnhofer, Tobias Ritschel, Karol Myszkowski, and Hans-Peter Seidel. Expressiveness and perception: Optimizing disparity for motion in depth. Computer Graphics Forum, 32 (4):143–152, July 2013. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Klehm:2015:SAR

[KRP^{+}15] Oliver Klehm, Fabrice Rousselle, Marios Papas, Derek Bradley, Christophe Hery, Bernd Bickel, Wojciech Jarosz, and Thabo Beeler. State of the art reports: Recent advances in facial appearance capture. Computer Graphics Forum, 34(2):709– 733, May 2015. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Krone:2013:BSI

[KRS^{+}13] M. Krone, G. Reina, C. Schulz, T. Kulschewski, J. Pleiss, and T. Ertl. Biological shape: Interactive extraction and tracking of biomolecular surface features. Computer Graphics Forum, 32(3pt3):331–340, June 2013. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Kim:2010:IVI

[KS10] Mikyung Kim and Hyun Joon Shin. Image and video II: an example-based approach to synthesize artistic strokes using graphs. Computer Graphics Forum, 29(7):2145–2152, September 2010. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Kalantari:2011:SEC

[KS11] Nima Khademi Kalantari and Pradeep Sen. Sampling: Efficient computation of blue noise point sets through importance sampling. Computer Graphics Forum, 30(4):1215–1221, June 2011. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Kalantari:2012:SFG

[KS12a] Nima Khademi Kalantari and Pradeep Sen. Sampling: Fast generation of approximate blue noise point sets. Computer Graphics Forum, 31(4):1529–1535, June 2012. CODEN CG-FODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Kyota:2012:AFG

[KS12b] Fumihito Kyota and Suguru Saito. Animation: Fast grasp synthesis for various shaped objects. Computer Graphics Forum, 31(2pt4):765–774, May 2012. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Kalantari:2013:IRN

[KS13a] Nima Khademi Kalantari and Pradeep Sen. Illumination: Removing the noise in Monte Carlo rendering with general image denoising algorithms. Computer Graphics Forum, 32(2pt1):93– 102, May 2013. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Karloff:2013:GME

[KS13b] Howard Karloff and Kenneth E. Shirley. Graphs: Maximum entropy summary trees. Computer Graphics Forum, 32(3pt1): 71–80, June 2013. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Kang:2014:PDI

[KS14] Nahyup Kang and Donghoon Sagong. Particles and deformation: Incompressible SPH using the divergence-free condition. Computer Graphics Forum, 33(7):219–228, October 2014. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Kohek:2018:ILS

[KS18] Štefan Kohek and Damjan Strnad. Interactive large-scale procedural forest construction and visualization based on particle flow simulation. Computer Graphics Forum, 37(1):389–402, February 2018. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Kazhdan:2012:FCM

[KSBC12] Michael Kazhdan, Jake Solomon, and Mirela Ben-Chen. Flows: Can mean-curvature flow be modified to be non-singular? Computer Graphics Forum, 31(5):1745–1754, August 2012. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Kim:2019:RTH

[KSCN19] Jaedong Kim, Hyunggoog Seo, Seunghoon Cha, and Junyong Noh. Real-time human shadow removal in a front projection system. Computer Graphics Forum, 38(1):443–454, February 2019. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Kosinka:2014:SSC

[KSD14a] J. Kosinka, M. A. Sabin, and N. A. Dodgson. Subdivision surfaces with creases and truncated multiple knot lines. Computer Graphics Forum, 33(1):118–128, February 2014. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Kosinka:2014:SSS

[KSD14b] J. Kosinka, M. A. Sabin, and N. A. Dodgson. Surfaces and shells: Semi-sharp creases on subdivision curves and surfaces. Computer Graphics Forum, 33(5):217–226, August 2014. CO-DEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Kratt:2015:CTW

[KSG^{+}15] J. Kratt, M. Spicker, A. Guayaquil, M. Fiser, S. Pirk, O. Deussen, J. C. Hart, and B. Benes. Colors and textures: Woodification: User-controlled cambial growth modeling. Computer Graphics Forum, 34(2):361–372, May 2015. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Kurtek:2013:SCL

[KSKL13] Sebastian Kurtek, Anuj Srivastava, Eric Klassen, and Hamid Laga. Shape correspondences: Landmark-guided elastic shape analysis of spherically-parameterized surfaces. Computer Graphics Forum, 32(2pt4):429–438, May 2013. CODEN CG-FODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Kavan:2010:SDA

[KSO10] L. Kavan, P.-P. Sloan, and C. O’Sullivan. Shape deformation and animation: Fast and efficient skinning of animated meshes. Computer Graphics Forum, 29(2):327–336, May 2010. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Kim:2018:DMB

[KSP^{+}18] Beomseok Kim, Hyeongseok Son, Seong-Jin Park, Sunghyun Cho, and Seungyong Lee. Defocus and motion blur detection with deep contextual features. Computer Graphics Forum, 37 (7):277–288, October 2018. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Kopta:2015:MCL

[KSS^{+}15] D. Kopta, K. Shkurko, J. Spjut, E. Brunvand, and A. Davis. Memory considerations for low energy ray tracing. Computer Graphics Forum, 34(1):47–59, February 2015. CODEN CG-FODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Karch:2012:FVV

[KSW^{+}12] Grzegorz Karol Karch, Filip Sadlo, Daniel Weiskopf, Claus-Dieter Munz, and Thomas Ertl. Flow visualization: Visualization of advection-diffusion in unsteady fluid flow. Computer Graphics Forum, 31(3pt2):1105–1114, June 2012. CODEN CG-FODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Kato:2019:SVM

[KTKM19] Y. Kato, S. Tanaka, Y. Kanamori, and J. Mitani. Single-view modeling of layered origami with plausible outer shape. Computer Graphics Forum, 38(7):629–640, October 2019. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Kakimoto:2010:ESU

[KTN10] M. Kakimoto, T. Tatsukawa, and T. Nishita. An eyeglass simulator using conoid tracing. Computer Graphics Forum, 29(8): 2427–2437, December 2010. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Kontkanen:2011:AGI

[KTO11] Janne Kontkanen, Eric Tabellion, and Ryan S. Overbeck. Accelerating global illumination: Coherent out-of-core pointbased global illumination. Computer Graphics Forum, 30(4): 1353–1360, June 2011. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Kruger:2013:IAT

[KTW^{+}13] Robert Krüger, Dennis Thom, Michael Wörner, Harald Bosch, and Thomas Ertl. Interactive analysis: TrajectoryLenses – a set-based filtering and exploration technique for long-term trajectory data. Computer Graphics Forum, 32(3pt4):451–460, June 2013. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Kuhlen:2012:RES

[Kuh12] Torsten Kuhlen. Reports: 2011 Eurographics Symposium on Parallel Graphics and Visualization. Computer Graphics Forum, 31(1):216, February 2012. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Kronander:2010:VRE

[KUMY10] Joel Kronander, Jonas Unger, Torsten Möller, and Anders Ynnerman. Volume rendering: Estimation and modeling of actual numerical errors in volume rendering. Computer Graphics Forum, 29(3):893–902, June 2010. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Kurlin:2015:CGO

[Kur15] V. Kurlin. Curves and graphs: a one-dimensional homologically persistent skeleton of an unstructured point cloud in any metric space. Computer Graphics Forum, 34(5):253–262, August 2015. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Krekel:2010:MVV

[KVD^{+}10] Peter R. Krekel, Edward R. Valstar, Jurriaan De Groot, Frits H. Post, Rob G. H. H. Nelissen, and Charl P. Botha. Medical visualization: Visual analysis of multi-joint kinematic data. Computer Graphics Forum, 29(3):1123–1132, June 2010. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Kochtchi:2014:DTV

[KvLB14] A. Kochtchi, T. von Landesberger, and C. Biemann. Document and text visualization: Networks of names: Visual exploration and semi-automatic tagging of social networks from newspaper articles. Computer Graphics Forum, 33(3):211–220, June 2014. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Khlebnikov:2014:LLD

[KVS^{+}14] Rostislav Khlebnikov, Philip Voglreiter, Markus Steinberger, Bernhard Kainz, and Dieter Schmalstieg. Low-level direct volume rendering / filtering: Parallel irradiance caching for interactive Monte-Carlo direct volume rendering. Computer Graphics Forum, 33(3):61–70, June 2014. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Kurzhals:2018:EVD

[KW18] Kuno Kurzhals and Daniel Weiskopf. Exploring the visualization design space with repertory grids. Computer Graphics Forum, 37(3):133–144, June 2018. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Kerber:2012:CAR

[KWC^{+}12] J. Kerber, M. Wang, J. Chang, J. J. Zhang, A. Belyaev, and H.-P. Seidel. Computer assisted relief generation — a survey. Computer Graphics Forum, 31(8):2363–2377, December 2012. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Kachkaev:2014:DRG

[KWD14] A. Kachkaev, J. Wood, and J. Dykes. Data relationships: Glyphs for exploring crowd-sourced subjective survey classification. Computer Graphics Forum, 33(3):311–320, June 2014. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Kelly:2015:PPM

[KWM15] T. Kelly, P. Wonka, and P. Mueller. Procedural and parametric modeling: Interactive dimensioning of parametric models. Computer Graphics Forum, 34(2):117–129, May 2015. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Knoll:2014:LLD

[KWN^{+}14] Aaron Knoll, Ingo Wald, Paul Navratil, Anne Bowen, Khairi Reda, Michael E. Papka, and Kelly Gaither. Low-level direct volume rendering / filtering: RBF volume ray casting on multicore and manycore CPUs. Computer Graphics Forum, 33(3): 71–80, June 2014. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Kim:2018:SSL

[KWÖG18] Byungsoo Kim, Oliver Wang, A. Cengiz Öztireli, and Markus Gross. Semantic segmentation for line drawing vectorization using neural networks. Computer Graphics Forum, 37(2):329– 338, May 2018. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Kothur:2015:TST

[KWS^{+}15] P. Köthur, C. Witt, M. Sips, N. Marwan, S. Schinkel, and D. Dransch. Time-series and topology: Visual analytics for correlation-based comparison of time series ensembles. Computer Graphics Forum, 34(3):411–420, June 2015. CODEN CG-FODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Kalojanov:2016:SMB

[KWS16] Javor Kalojanov, Michael Wand, and Philipp Slusallek. Shape M&M: Building construction sets by tiling grammar simplification. Computer Graphics Forum, 35(2):13–25, May 2016. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Kaufmann:2013:IFE

[KWSH^{+}13] Peter Kaufmann, Oliver Wang, Alexander Sorkine-Hornung, Olga Sorkine-Hornung, Aljoscha Smolic, and Markus Gross. Images: Finite element image warping. Computer Graphics Forum, 32(2pt1):31–39, May 2013. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Kurz:2014:SAT

[KWW^{+}14] C. Kurz, X. Wu, M. Wand, T. Thormählen, P. Kohli, and H.-P. Seidel. Symmetry-aware template deformation and fitting. Computer Graphics Forum, 33(6):205–219, September 2014. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Kuo:2016:IVF

[KYC16] Ming-Hsun Kuo, Yong-Liang Yang, and Hung-Kuo Chu. Images and video: Feature-aware pixel art animation. Computer Graphics Forum, 35(7):411–420, October 2016. CODEN CG-FODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Kim:2014:CIP

[KYKL14] Hye-Rin Kim, Min-Joon Yoo, Henry Kang, and In-Kwon Lee. Color and imaging: Perceptually-based color assignment. Computer Graphics Forum, 33(7):309–318, October 2014. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Kuge:2019:RTI

[KYM19] Takahiro Kuge, Tatsuya Yatagawa, and Shigeo Morishima. Real-time indirect illumination of emissive inhomogeneous volumes using layered polygonal area lights. Computer Graphics Forum, 38(7):449–460, October 2019. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Kaiser:2019:SSG

[KZB19] Adrien Kaiser, Jose Alonso Ybanez Zepeda, and Tamy Boubekeur. A survey of simple geometric primitives detection methods for captured 3D data. Computer Graphics Forum, 38 (1):167–196, February 2019. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Kuang:2012:MDM

[KZZM12] X. Kuang, H. Zhang, S. Zhao, and M. J. McGuffin. Multidimensional and multi-variate data: Tracing tuples across dimensions: a comparison of scatterplots and parallel coordinate plots. Computer Graphics Forum, 31(3pt4):1365–1374, June 2012. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Lespinats:2011:CSC

[LA11] Sylvain Lespinats and Michaël Aupetit. CheckViz: Sanity check and topological clues for linear and non-linear mappings. Computer Graphics Forum, 30(1):113–125, March 2011. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Lafarge:2013:SCS

[LA13] Florent Lafarge and Pierre Alliez. Shape construction: Surface reconstruction through point set structuring. Computer Graphics Forum, 32(2pt2):225–234, May 2013. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Liu:2015:VAC

[LA15] Xiaoyang Liu and Usman R. Alim. Volume analysis and classification: Compressive volume rendering. Computer Graphics Forum, 34(3):101–110, June 2015. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Lehmann:2012:SCR

[LAE^{+}12] Dirk J. Lehmann, Georgia Albuquerque, Martin Eisemann, Marcus Magnor, and Holger Theisel. Selecting coherent and relevant plots in large scatterplot matrices. Computer Graphics Forum, 31(6):1895–1908, September 2012. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Liu:2012:AIT

[LAFT12] Jingbo Liu, Oscar Kin-Chung Au, Hongbo Fu, and Chiew-Lan Tai. Animation and interaction: Two-finger gestures for 6DOF manipulation of 3D objects. Computer Graphics Forum, 31 (7pt1):2047–2055, September 2012. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Laine:2013:GTT

[Lai13] Samuli Laine. Geometry and topology: a topological approach to voxelization. Computer Graphics Forum, 32(4):77–86, July 2013. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Laramee:2010:HWV

[Lar10] Robert S. Laramee. How to write a visualization research paper: a starting point. Computer Graphics Forum, 29(8):2363–2371, December 2010. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Law:2011:PBA

[LAS^{+}11] Alvin J. Law, Daniel G. Aliaga, Behzad Sajadi, Aditi Majumder, and Zygmunt Pizlo. Perceptually based appearance modification for compliant appearance editing. Computer Graphics Forum, 30(8):2288–2300, December 2011. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Lavoue:2011:SMM

[Lav11] Guillaume Lavoué. Section 3: a multiscale metric for 3D mesh visual quality assessment. Computer Graphics Forum, 30(5):

1427–1437, August 2011. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Li:2019:VVA

[LB19] Chenhui Li and George Baciu. VisFM: Visual analysis of image feature matchings. Computer Graphics Forum, 38(1):46–58, February 2019. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Lambert:2010:GWR

[LBA10] A. Lambert, R. Bourqui, and D. Auber. Graphs: Winding roads: Routing edges into bundles. Computer Graphics Forum, 29(3):853–862, June 2010. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Litman:2014:MSL

[LBBC14] Roee Litman, Alex Bronstein, Michael Bronstein, and Umberto Castellani. Matching: Supervised learning of bag-of-features shape descriptors using sparse coding. Computer Graphics Forum, 33(5):127–136, August 2014. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Lambert:2016:MMR

[LBG16] Thibaud Lambert, Pierre Bénard, and Gaël Guennebaud. Meshes: Multi-resolution meshes for feature-aware hardware tessellation. Computer Graphics Forum, 35(2):253–262, May 2016. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Lindow:2012:VVI

[LBH12a] N. Lindow, D. Baum, and H.-C. Hege. Volume visualization: Interactive rendering of materials and biological structures on atomic and nanoscopic scale. Computer Graphics Forum, 31 (3pt4):1325–1334, June 2012. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Lindow:2012:PGP

[LBH12b] Norbert Lindow, Daniel Baum, and Hans-Christian Hege. Perceptual graphics: Perceptually linear parameter variations. Computer Graphics Forum, 31(2pt3):535–544, May 2012. CO-DEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Liu:2016:HDD

[LBJ^{+}16] S. Liu, P.-T Bremer, J. J. Jayaraman, B. Wang, B. Summa, and V. Pascucci. High-dimensional data: The Grassmannian atlas: A general framework for exploring linear projections of high-dimensional data. Computer Graphics Forum, 35(3):1–10, June 2016. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Lanir:2014:GVV

[LBK14] J. Lanir, P. Bak, and T. Kuflik. Geographic visualization: Visualizing proximity-based spatiotemporal behavior of museum visitors using Tangram diagrams. Computer Graphics Forum, 33(3):261–270, June 2014. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Lindow:2010:AAV

[LBPH10] Norbert Lindow, Daniel Baum, Steffen Prohaska, and Hans-Christian Hege. Acceleration: Accelerated visualization of dynamic molecular surfaces. Computer Graphics Forum, 29(3): 943–952, June 2010. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Leimer:2018:SRI

[LBRM18] Kurt Leimer, Michael Birsak, Florian Rist, and Przemyslaw Musialski. Sit & relax: Interactive design of body-supporting surfaces. Computer Graphics Forum, 37(7):349–359, October 2018. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Li:2018:CSC

[LCB^{+}18] Mingzhao Li, Farhana Choudhury, Zhifeng Bao, Hanan Samet, and Timos Sellis. ConcaveCubes: Supporting cluster-based geographical visualization in large data scale. Computer Graphics Forum, 37(3):217–228, June 2018. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Lee:2017:VVH

[LCC^{+}17] Wei-Tse Lee, Hsin-I Chen, Ming-Shiuan Chen, I-Chao Shen, and Bing-Yu Chen. Video and visualization: High-resolution 360 video foveated stitching for real-time VR. Computer Graphics Forum, 36(7):115–123, October 2017. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Lee:2018:ERS

[LCC^{+}18] Yin-Hsuan Lee, Yu-Kai Chang, Yu-Lun Chang, I-Chen Lin, Yu-Shuen Wang, and Wen-Chieh Lin. Enhancing the realism of sketch and painted portraits with adaptable patches. Computer Graphics Forum, 37(1):214–225, February 2018. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Lei:2013:IPS

[LCCC13] Su-Ian Eugene Lei, Ying-Chieh Chen, Hsiang-Ting Chen, and Chun-Fa Chang. Image processing (session 3): Interactive physics-based ink splattering art creation. Computer Graphics Forum, 32(7):147–156, October 2013. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Liu:2010:VRV

[LCD10] Baoquan Liu, Gordon J. Clapworthy, and Feng Dong. Volume reconstruction and visualization: Multi-layer depth peeling by single-pass rasterisation for faster isosurface raytracing on GPUs. Computer Graphics Forum, 29(3):1231–1240, June 2010. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Liu:2016:VDA

[LCDW16] Baoquan Liu, Gordon J. Clapworthy, Feng Dong, and Enhua Wu. Volume data applications: Parallel marching blocks: A practical isosurfacing algorithm for large data on many-core architectures. Computer Graphics Forum, 35(3):211–220, June 2016. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Liang:2010:CAH

[LCG10] Chia-Kai Liang, Wei-Chao Chen, and Natasha Gelfand. Colour adjustment and halftoning: TouchTone: Interactive local image adjustment using point-and-swipe. Computer Graphics Forum, 29(2):253–261, May 2010. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Lv:2012:SPM

[LCHB12] Jiajun Lv, Xinlei Chen, Jin Huang, and Hujun Bao. Shape processing and modeling: Semi-supervised mesh segmentation and labeling. Computer Graphics Forum, 31(7pt2):2241–2248, September 2012. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Liu:2010:GIS

[LCLJ10] L. Liu, E. W. Chambers, D. Letscher, and T. Ju. Geometry II: a simple and robust thinning algorithm on cell complexes. Computer Graphics Forum, 29(7):2253–2260, September 2010. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Luo:2018:PMP

[LCM^{+}18] Linbo Luo, Cheng Chai, Jianfeng Ma, Suiping Zhou, and Wentong Cai. ProactiveCrowd: Modelling proactive steering behaviours for agent-based crowd simulation. Computer Graphics Forum, 37(1):375–388, February 2018. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Lopez:2019:CND

[LCMP19] Axel López, François Chaumette, Eric Marchand, and Julien Pettré. Character navigation in dynamic environments based on optical flow. Computer Graphics Forum, 38(2):181–192, May 2019. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Legg:2012:VDS

[LCP^{+}12] P. A. Legg, D. H. S. Chung, M. L. Parry, M. W. Jones, R. Long, I. W. Griffiths, and M. Chen. Visualization for decision support: MatchPad: Interactive glyph-based visualization for real-time sports performance analysis. Computer Graphics Forum, 31 (3pt4):1255–1264, June 2012. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Lerner:2010:ACD

[LCSCO10] Alon Lerner, Yiorgos Chrysanthou, Ariel Shamir, and Daniel Cohen-Or. Animation: Context-dependent crowd evaluation. Computer Graphics Forum, 29(7):2197–2206, September 2010. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Lavoue:2018:VAR

[LCSL18] Guillaume Lavoué, Frédéric Cordier, Hyewon Seo, and Mohamed-Chaker Larabi. Visual attention for rendered 3D shapes. Computer Graphics Forum, 37(2):191–203, May 2018. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

LYi:2019:TUR

[LCSS19] Sehi L’Yi, Youli Chang, DongHwa Shin, and Jinwook Seo. Toward understanding representation methods in visualization recommendations through scatterplot construction tasks. Computer Graphics Forum, 38(3):201–211, June 2019. CODEN CG-FODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Liu:2014:ISA

[LCUR14] Yiming Liu, Michael Cohen, Matt Uyttendaele, and Szymon Rusinkiewicz. Illustration and stylization: AutoStyle: Automatic style transfer from image collections to users’ images. Computer Graphics Forum, 33(4):21–31, July 2014. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Liu:2010:SCO

[LCWCO10] Ligang Liu, Renjie Chen, Lior Wolf, and Daniel Cohen-Or. Stitching and compositing: Optimizing photo composition. Computer Graphics Forum, 29(2):469–478, May 2010. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Liu:2011:SCC

[LCY^{+}11] Dongquan Liu, Quan Chen, Jun Yu, Huiqin Gu, Dacheng Tao, and Hock Soon Seah. Stroke correspondence construction using manifold learning. Computer Graphics Forum, 30(8):2194– 2207, December 2011. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Lu:2019:RSS

[LCY^{+}19] Jia-Ming Lu, Xiao-Song Chen, Xiao Yan, Chen-Feng Li, Ming Lin, and Shi-Min Hu. A rigging-skinning scheme to control fluid simulation. Computer Graphics Forum, 38(7):501–512, October 2019. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Lambert:2011:BDV

[LDB11] A. Lambert, J. Dubois, and R. Bourqui. Biological data visualization: Pathway preserving representation of metabolic networks. Computer Graphics Forum, 30(3):1021–1030, June 2011. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Laurent:2016:LTF

[LDdLRB16] Gilles Laurent, Cyril Delalandre, Grégoire de La Rivière, and Tamy Boubekeur. Light transport: Forward light cuts: a scalable approach to real-time global illumination. Computer Graphics Forum, 35(4):79–88, July 2016. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Lancelle:2019:CMB

[LDG19] M. Lancelle, P. Dogan, and M. Gross. Controlling motion blur in synthetic long time exposures. Computer Graphics Forum, 38(2):393–403, May 2019. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Li:2015:RDA

[LDGN15] Yangyan Li, Angela Dai, Leonidas Guibas, and Matthias Nießner. Reconstruction: Database-assisted object retrieval for real-time 3D reconstruction. Computer Graphics Forum, 34(2):435–446, May 2015. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Lan:2010:ACB

[LDW^{+}10] Yanxiang Lan, Yue Dong, Jiaping Wang, Xin Tong, and Baining Guo. Acquisition: Condenser-based instant reflectometry. Computer Graphics Forum, 29(7):2091–2098, September 2010. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Lopez:2010:AMC

[LDY10] Luis D. Lopez, Yuanyuan Ding, and Jingyi Yu. Acquisition: Modeling complex unfoliaged trees from a sparse set of images. Computer Graphics Forum, 29(7):2075–2082, September 2010. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Lee:2013:SAP

[LE13] Sungkil Lee and Elmar Eisemann. Simulated and augmented photography: Practical real-time lens-flare rendering. Computer Graphics Forum, 32(4):1–6, July 2013. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Liao:2017:RGB

[LEE17] Jingtang Liao, Martin Eisemann, and Elmar Eisemann. Reconstruction and generation based on RGBD images: Split-depth image generation and optimization. Computer Graphics Forum, 36(7):175–182, October 2017. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Livesu:2017:MPO

[LEM^{+}17] Marco Livesu, Stefano Ellero, Jonàs Martínez, Sylvain Lefebvre, and Marco Attene. From 3D models to 3D prints: an overview of the processing pipeline. Computer Graphics Forum, 36(2):537–564, May 2017. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Lee:2010:SMR

[LeYO^{+}10] Yongjoon Lee, Sung eui Yoon, Seungwoo Oh, Duksu Kim, and Sunghee Choi. Simulation: Multi-resolution cloth simulation. Computer Graphics Forum, 29(7):2225–2232, September 2010. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Lessig:2010:CGT

[LF10] Christian Lessig and Eugene Fiume. Computer graphics theory: On the effective dimension of light transport. Computer Graphics Forum, 29(4):1399–1403, June 2010. CODEN CG-FODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Larboulette:2011:RSC

[LF11] Caroline Larboulette and François Faure. Reports: Symposium on Computer Animation 2010. Computer Graphics Forum, 30 (6):1867, September 2011. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Lukac:2015:DPB

[LFA^{+}15] M. Lukáč, J. Fišer, P. Asente, J. Lu, E. Shechtman, and D. Sýkora. Drawing and painting: Brushables: Example-based edge-aware directional texture painting. Computer Graphics Forum, 34(7):257–267, October 2015. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Loorak:2014:GVP

[LFC14] M. Hosseinkhani Loorak, P. W. L. Fong, and S. Carpendale. Graph visualization: Papilio: Visualizing Android application permissions. Computer Graphics Forum, 33(3):391–400, June 2014. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Lin:2013:FSS

[LFK^{+}13] Sharon Lin, Julie Fortuna, Chinmay Kulkarni, Maureen Stone, and Jeffrey Heer. Foundations: Selecting semantically-resonant colors for data visualization. Computer Graphics Forum, 32 (3pt4):401–410, June 2013. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Lindholm:2015:HDV

[LFS^{+}15] S. Lindholm, M. Falk, E. Sundén, A. Bock, A. Ynnerman, and T. Ropinski. Hybrid data visualization based on depth complexity histogram analysis. Computer Graphics Forum, 34(1): 74–85, February 2015. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Liao:2019:PMV

[LFYX19] Jie Liao, Yanping Fu, Qingan Yan, and Chunxia Xiao. Pyramid multi-view stereo with local consistency. Computer Graphics Forum, 38(7):335–346, October 2019. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Levi:2013:SAS

[LG13] Zohar Levi and Craig Gotsman. Sketching: ArtiSketch: a system for articulated sketch modeling. Computer Graphics Forum, 32(2pt2):235–244, May 2013. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Lawonn:2019:SIT

[LG19] Kai Lawonn and Tobias Günther. Stylized image triangulation. Computer Graphics Forum, 38(1):221–234, February 2019. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Liang:2019:GIP

[LGC^{+}19] Y. Liang, Y. Gan, M. Chen, D. Gutierrez, and A. Muñoz. Generic interactive pixel-level image editing. Computer Graphics Forum, 38(7):23–34, October 2019. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Landes:2013:GTS

[LGH13] Pierre-Edouard Landes, Bruno Galerne, and Thomas Hurtut. Geometry and topology: a shape-aware model for discrete texture synthesis. Computer Graphics Forum, 32(4):67–76, July 2013. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Lu:2017:SAP

[LGH^{+}17] Yafeng Lu, Rolando Garcia, Brett Hansen, Michael Gleicher, and Ross Maciejewski. The state-of-the-art in predictive visual analytics. Computer Graphics Forum, 36(3):539–562, June 2017. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Lim:2016:SIS

[LGK16] Isaak Lim, Anne Gehre, and Leif Kobbelt. Structures: Identifying style of 3D shapes using deep metric learning. Computer Graphics Forum, 35(5):207–215, August 2016. CODEN CG-FODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Lawonn:2014:ASV

[LGP14] Kai Lawonn, Rocco Gasteiger, and Bernhard Preim. Adaptive surface visualization of vessels with animated blood flow. Computer Graphics Forum, 33(8):16–27, December 2014. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Luo:2018:UDA

[LGW18] Chuanjiang Luo, Xiaoyin Ge, and Yusu Wang. Uniformization and density adaptation for point cloud data via graph Laplacian. Computer Graphics Forum, 37(1):325–337, February 2018. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Liu:2016:SMF

[LGZ^{+}16] Tiantian Liu, Ming Gao, Lifeng Zhu, Eftychios Sifakis, and Ladislav Kavan. Shape M&M: Fast and robust inversion-free shape manipulation. Computer Graphics Forum, 35(2):1–11, May 2016. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Lampe:2011:VAC

[LH11] O. Daae Lampe and H. Hauser. 2D visualization by aggregation: Curve density estimates. Computer Graphics Forum, 30 (3):633–642, June 2011. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Lin:2013:EPI

[LHH^{+}13] Wen-Chieh Lin, Tsung-Shian Huang, Tan-Chi Ho, Yueh-Tse Chen, and Jung-Hong Chuang. Expressiveness and perception: Interactive lighting design with hierarchical light representation. Computer Graphics Forum, 32(4):133–142, July 2013. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Liu:2018:CSC

[LHNS18] Xiaotong Liu, Yifan Hu, Stephen North, and Han-Wei Shen. CorrelatedMultiples: Spatially coherent small multiples with constrained multi-dimensional scaling. Computer Graphics Forum, 37(1):7–18, February 2018. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Lhuillier:2017:SAE

[LHT17] A. Lhuillier, C. Hurter, and A. Telea. State of the art in edge and trail bundling techniques. Computer Graphics Forum, 36 (3):619–645, June 2017. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Lieng:2017:REI

[Lie17] Henrik Lieng. Representing and editing images: a probabilistic framework for component-based vector graphics. Computer Graphics Forum, 36(7):195–205, October 2017. CODEN CG-FODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Lim:2019:CDP

[LIK19] Isaak Lim, Moritz Ibing, and Leif Kobbelt. A convolutional decoder for point clouds using adaptive instance normalization. Computer Graphics Forum, 38(5):99–108, August 2019. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Levine:2012:TQB

[LJB^{+}12] J. A. Levine, S. Jadhav, H. Bhatia, V. Pascucci, and P.-T. Bremer. Topology: a quantized boundary representation of 2D flows. Computer Graphics Forum, 31(3pt1):945–954, June 2012. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Limper:2013:GSP

[LJBA13] M. Limper, Y. Jung, J. Behr, and M. Alexa. Geometry (session 1): The POP buffer: Rapid progressive clustering by geometry quantization. Computer Graphics Forum, 32(7):197–206, October 2013. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Liu:2017:SKD

[LJC17] Hsueh-Ti Derek Liu, Alec Jacobson, and Keenan Crane. Spectra and kernels: a Dirac operator for extrinsic shape analysis. Computer Graphics Forum, 36(5):139–149, August 2017. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Li:2010:IVI

[LJH10] Yong Li, Tao Ju, and Shi-Min Hu. Image and video I: Instant propagation of sparse edits on images and videos. Computer Graphics Forum, 29(7):2049–2054, September 2010. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Liu:2013:FIR

[LJH13] Zhicheng Liu, Biye Jiang, and Jeffrey Heer. Foundations: im-Mens: Real-time visual querying of big data. Computer Graphics Forum, 32(3pt4):421–430, June 2013. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Lemercier:2012:CCS

[LJK^{+}12] S. Lemercier, A. Jelic, R. Kulpa, J. Hua, J. Fehrenbach, P. Degond, C. Appert-Rolland, S. Donikian, and J. Pettré. Crowds, cloths, and shape deformation: Realistic following behaviors for crowd simulation. Computer Graphics Forum, 31(2pt2):489– 498, May 2012. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Lee:2017:STD

[LJKL17] Hyunjoon Lee, Junho Jeon, Junho Kim, and Seungyong Lee. Structure-texture decomposition of images with interval gradient. Computer Graphics Forum, 36(6):262–274, September 2017. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Li:2018:SPE

[LJL^{+}18] Kun Li, Nianhong Jiao, Yebin Liu, Yangang Wang, and Jingyu Yang. Shape and pose estimation for closely interacting persons using multi-view images. Computer Graphics Forum, 37(7): 361–371, October 2018. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Liu:2019:LSC

[LJLH19] Yang Liu, Eunice Jun, Qisheng Li, and Jeffrey Heer. Latent space cartography: Visual analysis of vector space embeddings. Computer Graphics Forum, 38(3):67–78, June 2019. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Liang:2015:IVM

[LJZX15] Lingyu Liang, Lianwen Jin, Xin Zhang, and Yong Xu. Image and video: Multiple facial image editing using edge–aware PDE learning. Computer Graphics Forum, 34(7):203–212, October 2015. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Laine:2010:ILA

[LK10] Samuli Laine and Tero Karras. Indirect lighting and ambient occlusion: Two methods for fast ray-cast ambient occlusion. Computer Graphics Forum, 29(4):1325–1333, June 2010. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Laine:2011:SSS

[LK11] Samuli Laine and Tero Karras. Sampling: Stratified sampling for stochastic transparency. Computer Graphics Forum, 30(4): 1197–1204, June 2011. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Larsson:2013:FAF

[LK13] Thomas Larsson and Linus Källberg. Fitting and approximation: Fast and robust approximation of smallest enclosing balls in arbitrary dimensions. Computer Graphics Forum, 32(5):93– 101, August 2013. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Lee:2017:DML

[LK17] Yoonsang Lee and Taesoo Kwon. Depth to motion, lens and filters: Performance-based biped control using a consumer depth camera. Computer Graphics Forum, 36(2):387–395, May 2017. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Lee:2019:DPP

[LK19] Sing Chun Lee and Misha Kazhdan. Dense point-to-point correspondences between genus-zero shapes. Computer Graphics Forum, 38(5):27–37, August 2019. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Lee:2012:TDI

[LKC^{+}12] Hanseung Lee, Jaeyeon Kihm, Jaegul Choo, John Stasko, and Haesun Park. Text and documents: iVisClustering: an interactive visual document clustering via topic modeling. Computer Graphics Forum, 31(3pt3):1155–1164, June 2012. CODEN CG-FODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Liu:2017:CEV

[LKD^{+}17] Zhicheng Liu, Bernard Kerr, Mira Dontcheva, Justin Grover, Matthew Hoffman, and Alan Wilson. CoreFlow: Extracting and visualizing branching patterns from event sequences. Computer Graphics Forum, 36(3):527–538, June 2017. CODEN CG-FODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Lawonn:2014:MVL

[LKEP14] Kai Lawonn, Michael Krone, Thomas Ertl, and Bernhard Preim. Molecular visualization: Line integral convolution for real-time illustration of molecular surface shape and salient regions. Computer Graphics Forum, 33(3):181–190, June 2014. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Liu:2012:CFS

[LKF12] Tianqiang Liu, Vladimir G. Kim, and Thomas Funkhouser. Correspondence: Finding surface correspondences using symmetry axis curves. Computer Graphics Forum, 31(5):1607– 1616, August 2012. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Ljung:2016:FES

[LKG^{+}16] Patric Ljung, Jens Krüger, Eduard Groller, Markus Hadwiger, Charles D. Hansen, and Anders Ynnerman. Feature extraction for scalar and vector fields: State of the art in transfer functions for direct volume rendering. Computer Graphics Forum, 35 (3):669–691, June 2016. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Lieng:2017:CIS

[LKSD17] Henrik Lieng, Jiří Kosinka, Jingjing Shen, and Neil A. Dodgson. A colour interpolation scheme for topologically unrestricted gradient meshes. Computer Graphics Forum, 36(6):112–121, September 2017. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Lehmann:2015:HDV

[LKZ^{+}15] Dirk J. Lehmann, Fritz Kemmler, Tatsiana Zhyhalava, Marco Kirschke, and Holger Theisel. High-dimensional visualization: Visualnostics: Visual guidance pictograms for analyzing projections of high-dimensional data. Computer Graphics Forum, 34(3):291–300, June 2015. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Liang:2018:TDP

[LL18] Junbang Liang and Ming C. Lin. Time-domain parallelization for accelerating cloth simulation. Computer Graphics Forum, 37(8):21–34, December 2018. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Le:2019:AFC

[LL19] Hoang Le and Feng Liu. Appearance flow completion for novel view synthesis. Computer Graphics Forum, 38(7):555–565, October 2019. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Lipski:2010:VVC

[LLB^{+}10] C. Lipski, C. Linz, K. Berger, A. Sellent, and M. Magnor. Virtual video camera: Image-based viewpoint navigation through space and time. Computer Graphics Forum, 29(8):2555–2568, December 2010. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Lagae:2010:SPN

[LLC^{+}10a] A. Lagae, S. Lefebvre, R. Cook, T. DeRose, G. Drettakis, D. S. Ebert, J. P. Lewis, K. Perlin, and M. Zwicker. A survey of procedural noise functions. Computer Graphics Forum, 29(8): 2579–2600, December 2010. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Li:2010:ETM

[LLC10b] H. Li, W. K. Leow, and I.-S. Chiu. Elastic tubes: Modeling elastic deformation of hollow tubes. Computer Graphics Forum, 29(6):1770–1782, September 2010. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Lee:2011:IER

[LLC11] Kun-Ting Lee, Sheng-Jie Luo, and Bing-Yu Chen. Image enhancement: Rephotography using image collections. Computer Graphics Forum, 30(7):1895–1901, September 2011. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Lipsa:2012:VPS

[LLC^{+}12] Dan R. Lipşa, Robert S. Laramee, Simon J. Cox, Jonathan C. Roberts, Rick Walker, Michelle A. Borkin, and Hanspeter Pfister. Visualization for the physical sciences. Computer Graphics Forum, 31(8):2317–2347, December 2012. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Lin:2013:IPS

[LLC13] Yi-Shan Lin, Sheng-Jie Luo, and Bing-Yu Chen. Image processing (session 3): Artistic QR code embellishment. Computer Graphics Forum, 32(7):137–146, October 2013. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Liu:2015:CDE

[LLC^{+}15] Le Liu, Xiaosheng Li, Yanyun Chen, Xuehui Liu, Jian J. Zhang, and Enhua Wu. Collision detection: an efficient feathering system with collision control. Computer Graphics Forum, 34(7): 279–288, October 2015. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Lo:2019:GCS

[LLC19] Yi-Hsiang Lo, Ruen-Rone Lee, and Hung-Kuo Chu. Generating color scribble images using multi-layered monochromatic Strokes dithering. Computer Graphics Forum, 38(2):265–276, May 2019. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Lin:2016:VSV

[LLCZ16] Kaimo Lin, Shuaicheng Liu, Loong-Fah Cheong, and Bing Zeng. Video: Seamless video stitching from hand-held camera inputs. Computer Graphics Forum, 35(2):479–487, May 2016. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Lister:2010:KPC

[LLD10] W. Lister, R. G. Laycock, and A. M. Day. A key-pose caching system for rendering an animated crowd in real-time. Computer Graphics Forum, 29(8):2304–2312, December 2010. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Lasram:2012:TMP

[LLD12] Anass Lasram, Sylvain Lefebvre, and Cyrille Damez. Textures and materials: Procedural texture preview. Computer Graphics Forum, 31(2pt2):413–420, May 2012. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Li:2019:BRG

[LLLY19] Kun Li, Jingying Liu, Yu-Kun Lai, and Jingyu Yang. Book review: Generating 3D Faces using Multi-column Graph Convolutional Networks. Computer Graphics Forum, 38(7):215–224, October 2019. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Lienhard:2017:DTR

[LLM^{+}17] Stefan Lienhard, Cheryl Lau, Pascal Müller, Peter Wonka, and Mark Pauly. Design transformations for rule-based procedural modeling. Computer Graphics Forum, 36(2):39–48, May 2017. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Liao:2014:VTS

[LLN^{+}14] Jing Liao, Rodolfo S. Lima, Diego Nehab, Hugues Hoppe, and Pedro V. Sander. Video and textures: Semi-automated video morphing. Computer Graphics Forum, 33(4):51–60, July 2014. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Luo:2013:IPS

[LLSC13] Sheng-Jie Luo, Chin-Yu Lin, I-Chao Shen, and Bing-Yu Chen. Image processing (session 2): Stroke-guided image synthesis for skeletal structure editing. Computer Graphics Forum, 32 (7):71–78, October 2013. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Lu:2012:CGG

[LLW12] Lin Lu, Bruno Lévy, and Wenping Wang. Computational geometry and geometry processing: Centroidal Voronoi tessellation of line segments and graphs. Computer Graphics Forum, 31 (2pt4):775–784, May 2012. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Liu:2011:IVI

[LLX^{+}11] Yong-Jin Liu, Xi Luo, Yu-Ming Xuan, Wen-Feng Chen, and Xiao-Lan Fu. Image and video: Image retargeting quality assessment. Computer Graphics Forum, 30(2):583–592, April 2011. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Li:2010:CAH

[LM10] Hua Li and David Mould. Colour adjustment and halftoning: Contrast-aware halftoning. Computer Graphics Forum, 29 (2):273–280, May 2010. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Li:2015:SCE

[LM15] Hua Li and David Mould. Stylization: Contrast-enhanced black and white images. Computer Graphics Forum, 34(7):319–328, October 2015. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Lichtenberg:2018:ARS

[LMA^{+}18] N. Lichtenberg, R. Menges, V. Ageev, A. A. Paul George, P. Heimer, D. Imhof, and K. Lawonn. Analyzing residue surface proximity to interpret molecular dynamics. Computer Graphics Forum, 37(3):379–390, June 2018. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Li:2018:DRT

[LMG^{+}18a] S. Li, N. Marsaglia, C. Garth, J. Woodring, J. Clyne, and H. Childs. Data reduction techniques for simulation, visualization and data analysis. Computer Graphics Forum, 37(6): 422–447, September 2018. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Li:2018:PRC

[LMG^{+}18b] Yuqi Li, Aditi Majumder, M. Gopi, Chong Wang, and Jieyu Zhao. Practical radiometric compensation for projection display on textured surfaces using a multidimensional model. Computer Graphics Forum, 37(2):365–375, May 2018. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Lopez-Moreno:2013:MLS

[LMGH^{+}13] Jorge Lopez-Moreno, Elena Garces, Sunil Hadap, Erik Reinhard, and Diego Gutierrez. Multiple light source estimation in a single image. Computer Graphics Forum, 32(8):170–182, December 2013. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Liu:2014:SCS

[LMHH14] Bin Liu, Ralph R. Martin, Ji-Wu Huang, and Shi-Min Hu. Shapes and cryptography: Structure aware visual cryptography. Computer Graphics Forum, 33(7):141–150, October 2014. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Liu:2015:EPS

[LMK^{+}15] L. Liu, M. Mirzangar, R. M. Kirby, R. Whitaker, and D. H. House. Engineering and physical sciences: Visualizing timespecific hurricane predictions, with uncertainty, from storm path ensembles. Computer Graphics Forum, 34(3):371–380, June 2015. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Liu:2015:ISC

[LMLF15] Tianqiang Liu, Jim McCann, Wilmot Li, and Thomas Funkhouser. Images & scenes: Composition-aware scene optimization for product images. Computer Graphics Forum, 34 (2):13–24, May 2015. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Li:2015:AHC

[LMLG15] Yuqi Li, Aditi Majumder, Dongming Lu, and M. Gopi. Agile hardware: Content-independent multi-spectral display using superimposed projections. Computer Graphics Forum, 34 (2):337–348, May 2015. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Lauterbach:2010:MPG

[LMM10] C. Lauterbach, Q. Mo, and D. Manocha. Mesh processing: gProximity: Hierarchical GPU-based operations for collision and distance queries. Computer Graphics Forum, 29(2):419– 428, May 2010. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Lopez-Moreno:2017:SGV

[LMMCO17] Jorge Lopez-Moreno, David Miraut, Gabriel Cirio, and Miguel A. Otaduy. Sparse GPU voxelization of yarn-level cloth. Computer Graphics Forum, 36(1):22–34, January 2017. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Lewiner:2010:SSS

[LMP^{+}10] Thomas Lewiner, Vinícius Mello, Adelailson Peixoto, Sinésio Pesco, and Hélio Lopes. Surface and space structures: Fast generation of pointerless octree duals. Computer Graphics Forum, 29(5):1661–1669, July 2010. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Lawonn:2013:BSS

[LMP13] K. Lawonn, T. Moench, and B. Preim. Biological shape: Streamlines for illustrative real-time rendering. Computer Graphics Forum, 32(3pt3):321–330, June 2013. CODEN CG-FODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Lindemeier:2015:AHH

[LMPD15] Thomas Lindemeier, Jens Metzner, Lena Pollak, and Oliver Deussen. Agile hardware: Hardware-based non-photorealistic rendering using a painting robot. Computer Graphics Forum, 34(2):311–323, May 2015. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Livesu:2016:MSD

[LMPS16] Marco Livesu, Alessandro Muntoni, Enrico Puppo, and Riccardo Scateni. Modeling: Skeleton-driven adaptive hexahedral meshing of tubular shapes. Computer Graphics Forum, 35(7): 237–246, October 2016. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

LeMuzic:2016:BDV

[LMS^{+}16] M. Le Muzic, P. Mindek, J. Sorger, L. Autin, D. S. Goodsell, and I. Viola. Biological data visualization: Visibility equalizer cutaway visualization of mesoscopic biological models. Computer Graphics Forum, 35(3):161–170, June 2016. CODEN CG-FODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Liu:2019:VCT

[LMSF19] Xiaopei Liu, Maneesh Mishra, Martin Skote, and Chi-Wing Fu. On visualizing continuous turbulence scales. Computer Graphics Forum, 38(1):300–315, February 2019. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Lancelle:2016:FAC

[LMSG16] M. Lancelle, T. Martin, B. Solenthaler, and M. Gross. Fabrication: Anaglyph caustics with motion parallax. Computer Graphics Forum, 35(7):343–352, October 2016. CODEN CG-FODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Lawonn:2019:GOA

[LMW^{+}19] Kai Lawonn, Monique Meuschke, Ralph Wickenhöfer, Bernhard Preim, and Klaus Hildebrandt. A geometric optimization approach for the detection and segmentation of multiple aneurysms. Computer Graphics Forum, 38(3):413–425, June 2019. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Loubet:2017:HMV

[LN17] Guillaume Loubet and Fabrice Neyret. Hybrid mesh-volume LoDs for all-scale pre-filtering of complex 3D assets. Computer Graphics Forum, 36(2):431–442, May 2017. CODEN CG-FODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Loubet:2018:NMM

[LN18] Guillaume Loubet and Fabrice Neyret. A new microflake model with microscopic self-shadowing for accurate volume downsampling. Computer Graphics Forum, 37(2):111–121, May 2018. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Lescoat:2018:SDD

[LOM^{+}18] Thibault Lescoat, Maks Ovsjanikov, Pooran Memari, Jean-Marc Thiery, and Tamy Boubekeur. A survey on data-driven dictionary-based methods for 3D modeling. Computer Graphics Forum, 37(2):577–601, May 2018. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Li:2015:EVR

[LP15] Han Li and Pieter Peers. Exposure and video: Radiometric transfer: Example-based radiometric linearization of photographs. Computer Graphics Forum, 34(4):109–118, July 2015. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Liktor:2014:SFR

[LPD14] G. Liktor, M. Pan, and C. Dachsbacher. Surfaces: Fractional reyes-style adaptive tessellation for continuous level of detail. Computer Graphics Forum, 33(7):191–198, October 2014. CO-DEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Liu:2010:GDH

[LPG10] Yang Liu, Balakrishnan Prabhakaran, and Xiaohu Guo. Geometry I: Dirichlet harmonic shape compression with feature preservation for parameterized surfaces. Computer Graphics Forum, 29(7):2039–2048, September 2010. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Lu:2013:RSS

[LPG13] H. Lu, R. Pacanowski, and X. Granier. Rendering (session 1): Second-order approximation for variance reduction in multiple importance sampling. Computer Graphics Forum, 32(7): 131–136, October 2013. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Lin:2019:SEB

[LPG19] Y. Lin, P. Peers, and A. Ghosh. On-site example-based material appearance acquisition. Computer Graphics Forum, 38(4):15– 25, July 2019. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Li:2015:CDD

[LPH^{+}15] Siwang Li, Zherong Pan, Jin Huang, Hujun Bao, and Xiaogang Jin. Collision detection: Deformable objects collision handling with fast convergence. Computer Graphics Forum, 34(7): 269–278, October 2015. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Lee:2013:VPT

[LPK13] Sung-Ho Lee, Taejung Park, and Chang-Hun Kim. Visibility: Primitive trees for precomputed distance queries. Computer Graphics Forum, 32(2pt4):419–428, May 2013. CODEN CG-FODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Luan:2018:DPH

[LPSB18] Fujun Luan, Sylvain Paris, Eli Shechtman, and Kavita Bala. Deep painterly harmonization. Computer Graphics Forum, 37 (4):95–106, July 2018. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

LeMuzic:2014:BIV

[LPSV14] M. Le Muzic, J. Parulek, A. K. Stavrum, and I. Viola. BioMedical I: Illustrative visualization of molecular reactions using omniscient intelligence and passive agents. Computer Graphics Forum, 33(3):141–150, June 2014. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Liang:2019:RWI

[LQS^{+}19] X. Liang, B. Qiu, Z. Su, C. Gao, X. Shi, and R. Wang. Rain wiper: An incremental randomly wired network for single image deraining. Computer Graphics Forum, 38(7):159–169, October 2019. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Luboschik:2015:TST

[LRB^{+}15] M. Luboschik, M. Röhlig, A. T. Bittig, N. Andrienko, H. Schumann, and C. Tominski. Time-series and topology: Featuredriven visual analytics of chaotic parameter-dependent movement. Computer Graphics Forum, 34(3):421–430, June 2015. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Litany:2016:FCN

[LRB^{+}16] O. Litany, E. Rodolà, A. M. Bronstein, M. M. Bronstein, and D. Cremers. Functional correspondence: Non-rigid puzzles. Computer Graphics Forum, 35(5):135–143, August 2016. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Litany:2017:FSP

[LRBB17] O. Litany, E. Rodolà, A. M. Bronstein, and M. M. Bronstein. Fully spectral partial shape matching. Computer Graphics Forum, 36(2):247–258, May 2017. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Landreneau:2010:PBW

[LS10a] Eric Landreneau and Scott Schaefer. Poisson-based weight reduction of animated meshes. Computer Graphics Forum, 29(6): 1945–1954, September 2010. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Landreneau:2010:SSS

[LS10b] Eric Landreneau and Scott Schaefer. Surface and space structures: Scales and scale-like structures. Computer Graphics Forum, 29(5):1653–1660, July 2010. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Ljubovic:2015:IPI

[LS15] V. Ljubovic and H. Supic. Improving performance of image retrieval based on fuzzy colour histograms by using hybrid colour model and genetic algorithm. Computer Graphics Forum, 34 (8):77–87, December 2015. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Liebmann:2016:FVC

[LS16] T. Liebmann and G. Scheuermann. Flow visualization: Critical points of Gaussian-distributed scalar fields on simplicial grids. Computer Graphics Forum, 35(3):361–370, June 2016. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Liu:2017:IVM

[LSB^{+}17] L. Liu, D. Silver, K. Bemis, D. Kang, and E. Curchitser. Illustrative visualization of mesoscale ocean eddies. Computer Graphics Forum, 36(3):447–458, June 2017. CODEN CG-FODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Lawonn:2018:SMM

[LSBP18] K. Lawonn, N. N. Smit, K. Bühler, and B. Preim. A survey on multimodal medical data visualization. Computer Graphics Forum, 37(1):413–438, February 2018. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Laga:2011:REW

[LSF^{+}11] Hamid Laga, Tobias Schreck, Alfredo Ferreira, Afzal Godil, Ioannis Pratikakis, and Remco Veltkamp. Reports: Eurographics 2011 Workshop on 3D Object Retrieval (EG 3DOR’2011) in Cooperation with ACM SIGGRAPH: Lluandudno, UK, April 10, 2011. Computer Graphics Forum, 30(6):1865–1866, September 2011. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Lipp:2019:LEP

[LSL^{+}19] M. Lipp, M. Specht, C. Lau, P. Wonka, and P. Müller. Local editing of procedural models. Computer Graphics Forum, 38 (2):13–25, May 2019. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Liktor:2015:SSS

[LSMD15] G. Liktor, S. Spassov, G. Mückl, and C. Dachsbacher. Shadows: Stochastic soft shadow mapping. Computer Graphics Forum, 34(4):1–11, July 2015. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Lienhard:2014:PMI

[LSN^{+}14] S. Lienhard, M. Specht, B. Neubert, M. Pauly, and P. Müller. Procedural Modeling II: Thumbnail galleries for procedural models. Computer Graphics Forum, 33(2):361–370, May 2014. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Leimkuhler:2017:REM

[LSR17] Thomas Leimkühler, Hans-Peter Seidel, and Tobias Ritschel. Rendering is everywhere: Minimal warping: Planning incremental novel-view synthesis. Computer Graphics Forum, 36(4): 1–14, July 2017. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Lex:2012:VAS

[LSS^{+}12] A. Lex, M. Streit, H.-J. Schulz, C. Partl, D. Schmalstieg, P. J. Park, and N. Gehlenborg. Visual analysis in science and engineering: StratomeX: Visual analysis of large-scale heterogeneous genomics data for cancer subtype characterization. Computer Graphics Forum, 31(3pt3):1175–1184, June 2012. CO-DEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Lipp:2011:PMI

[LSWW11] M. Lipp, D. Scherzer, P. Wonka, and M. Wimmer. Procedural modelling: Interactive modeling of city layouts using layers of procedural content. Computer Graphics Forum, 30(2):345–354, April 2011. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Lin:2018:AMM

[LSZ^{+}18] Minmin Lin, Tianjia Shao, Youyi Zheng, Zhong Ren, Yanlin Weng, and Yin Yang. Automatic mechanism modeling from a single image with CNNs. Computer Graphics Forum, 37(7): 337–348, October 2018. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Leifman:2012:TMM

[LT12] George Leifman and Ayellet Tal. Textures and materials: Mesh colorization. Computer Graphics Forum, 31(2pt2):421–430, May 2012. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Lehmann:2016:VGP

[LT16] Dirk J. Lehmann and Holger Theisel. Visualization: General projective maps for multidimensional data projection. Computer Graphics Forum, 35(2):443–453, May 2016. CODEN CG-FODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Laga:2017:DSM

[LT17] Hamid Laga and Hedi Tabia. Design and segmentation: Modeling and exploring co-variations in the geometry and configuration of man-made 3D shape families. Computer Graphics Forum, 36(5):13–25, August 2017. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Legrand:2019:FQH

[LTB19] Hélène Legrand, Jean-Marc Thiery, and Tamy Boubekeur. Filtered quadrics for high-speed geometry smoothing and clustering. Computer Graphics Forum, 38(1):663–677, February 2019. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Loorak:2018:CII

[LTC18] M. Loorak, M. Tory, and S. Carpendale. ChangeCatcher: Increasing inter-author awareness for visualization development. Computer Graphics Forum, 37(3):51–62, June 2018. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Lieng:2012:PMI

[LTK12] Henrik Lieng, James Tompkin, and Jan Kautz. Photo manipulation: Interactive multi-perspective imagery from photos and videos. Computer Graphics Forum, 31(2pt1):285–293, May 2012. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Lieng:2015:SCV

[LTKD15] Henrik Lieng, Flora Tasse, Jiří Kosinka, and Neil A. Dodgson. Shading curves: Vector-based drawing with explicit gradient control. Computer Graphics Forum, 34(6):228–239, September 2015. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Liu:2019:BRL

[LTR19] Yifan Liu, Ruolan Tang, and Daniel Ritchie. Book review: Learning Style Compatibility Between Objects in a Real-World 3D Asset Database. Computer Graphics Forum, 38(7):775–784, October 2019. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Liu:2014:RDA

[LTX^{+}14] Zhenbao Liu, Sicong Tang, Weiwei Xu, Shuhui Bu, Junwei Han, and Kun Zhou. Reconstruction and depth: Automatic 3D indoor scene updating with RGBD cameras. Computer Graphics Forum, 33(7):269–278, October 2014. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Liu:2010:GCM

[LVJ10] Y. Liu, O. Veksler, and O. Juan. Generating classic mosaics with graph cuts. Computer Graphics Forum, 29(8):2387–2399, December 2010. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Lawonn:2018:SSB

[LVPI18] Kai Lawonn, Ivan Viola, Bernhard Preim, and Tobias Isenberg. A survey of surface-based illustrative rendering for visualization. Computer Graphics Forum, 37(6):205–234, September 2018. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Lettry:2018:UDS

[LVV18] L. Lettry, K. Vanhoey, and L. Van Gool. Unsupervised deep single-image intrinsic decomposition using illumination-varying image sequences. Computer Graphics Forum, 37(7):409–419, October 2018. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Liu:2015:SCR

[LVW^{+}15] Han Liu, Ulysse Vimont, Michael Wand, Marie-Paule Cani, Stefanie Hahmann, Damien Rohmer, and Niloy J. Mitra. Shape collections: Replaceable substructures for efficient part-based modeling. Computer Graphics Forum, 34(2):503–513, May 2015. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Li:2017:CMS

[LW17] Lei Li and Wencheng Wang. Contracting medial surfaces isotropically for fast extraction of centred curve skeletons. Computer Graphics Forum, 36(8):529–539, December 2017. CO-DEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Li:2018:IUL

[LW18] Lei Li and Wencheng Wang. Improved use of LOP for curve skeleton extraction. Computer Graphics Forum, 37(7):313–323, October 2018. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Li:2014:ASM

[LWB14] C. Li, P. J. Willis, and M. Brown. Appearance stylization of Manhattan world buildings. Computer Graphics Forum, 33(1): 15–26, February 2014. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Liu:2014:HLD

[LWBP14] S. Liu, B. Wang, P.-T. Bremer, and V. Pascucci. High level direct volume rendering: Distortion-guided structure-driven interactive exploration of high-dimensional data. Computer Graphics Forum, 33(3):101–110, June 2014. CODEN CG-FODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Laurijssen:2010:ILA

[LWDB10] J. Laurijssen, R. Wang, Ph. Dutré, and B. J. Brown. Indirect lighting and ambient occlusion: Fast estimation and rendering of indirect highlights. Computer Graphics Forum, 29(4):1305– 1313, June 2010. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Li:2016:FTM

[LWL^{+}16a] Hao Li, Guowei Wan, Honghua Li, Andrei Sharf, Kai Xu, and Baoquan Chen. Fitting and tracking: Mobility fitting using 4D RANSAC. Computer Graphics Forum, 35(5):79–88, August 2016. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Li:2016:IPR

[LWL^{+}16b] Kai Li, Jue Wang, Yebin Liu, Li Xu, and Qionghai Dai. Image processing: Re-compositable panoramic selfie with robust multi-frame segmentation and stitching. Computer Graphics Forum, 35(7):227–236, October 2016. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Laurijssen:2011:PCG

[LWLD11] Jurgen Laurijssen, Rui Wang, Ares Lagae, and Philip Dutré. Pre-computed gathering of multi-bounce glossy reflections. Computer Graphics Forum, 30(8):2270–2278, December 2011. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Lukasczyk:2017:NTG

[LWM^{+}17] Jonas Lukasczyk, Gunther Weber, Ross Maciejewski, Christoph Garth, and Heike Leitte. Nested tracking graphs. Computer Graphics Forum, 36(3):12–22, June 2017. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Li:2015:BMB

[LWPL15] Weizi Li, David Wolinski, Julien Pettré, and Ming C. Lin. Bodies in motion: Biologically-inspired visual simulation of insect swarms. Computer Graphics Forum, 34(2):425–434, May 2015. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Li:2013:MCR

[LWS^{+}13] Guannan Li, Chenglei Wu, Carsten Stoll, Yebin Liu, Kiran Varanasi, Qionghai Dai, and Christian Theobalt. Models: Capturing relightable human performances under general uncontrolled illumination. Computer Graphics Forum, 32(2pt3):275– 284, May 2013. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Lin:2016:TME

[LWS^{+}16] Ting-Hao Lin, Der-Lor Way, Zen-Chung Shih, Wen-Kai Tai, and Chin-Chen Chang. Textures/mapping: an efficient structure-aware bilateral texture filtering for image smoothing. Computer Graphics Forum, 35(7):57–66, October 2016. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Liebmann:2018:HCC

[LWS18] Tom Liebmann, Gunther H. Weber, and Gerik Scheuermann. Hierarchical correlation clustering in multiple 2D scalar fields. Computer Graphics Forum, 37(3):1–12, June 2018. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Liu:2015:HDV

[LWT^{+}15] S. Liu, B. Wang, J. J. Thiagarajan, P.-T. Bremer, and V. Pascucci. High-dimensional visualization: Visual exploration of high-dimensional data through subspace analysis and dynamic projections. Computer Graphics Forum, 34(3):271–280, June 2015. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Lv:2015:CSI

[LWX^{+}15] P. Lv, P. J. Wang, W. W. Xu, J. X. Chai, M. M. Zhang, Z. G. Pan, and M. L. Xu. Characters: a suggestive interface for sketch-based character posing. Computer Graphics Forum, 34 (7):111–121, October 2015. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Liu:2011:NLB

[LWY^{+}11] Baoquan Liu, Li-Yi Wei, Xu Yang, Chongyang Ma, Ying-Qing Xu, Baining Guo, and Enhua Wu. Non-linear beam tracing on a GPU. Computer Graphics Forum, 30(8):2156–2169, December 2011. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Liu:2011:ENL

[LXFW11] Xuecheng Liu, Shihong Xia, Yiwen Fan, and Zhaoqi Wang. Exploring non-linear relationship of blendshape facial animation. Computer Graphics Forum, 30(6):1655–1666, September 2011. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Liu:2019:ABO

[LXY19] Yifan Liu, Kun Xu, and Ling-Qi Yan. Adaptive BRDF-Oriented multiple importance sampling of many lights. Computer Graphics Forum, 38(4):123–133, July 2019. CODEN CG-FODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Liu:2015:BMI

[LYG15] Libin Liu, KangKang Yin, and Baining Guo. Bodies in motion: Improving sampling-based motion control. Computer Graphics Forum, 34(2):415–423, May 2015. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Lo:2010:MSB

[LZ10] Wan-Yen Lo and Matthias Zwicker. Motion synthesis: Bidirectional search for interactive motion synthesis. Computer Graphics Forum, 29(2):563–573, May 2010. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Luan:2017:MTF

[LZB17] Fujun Luan, Shuang Zhao, and Kavita Bala. Materials and textures: Fiber-level on-the-fly procedural textiles. Computer Graphics Forum, 36(4):123–135, July 2017. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Li:2018:NLL

[LZFH18] Xianzhi Li, Lei Zhu, Chi-Wing Fu, and Pheng-Ann Heng. Nonlocal low-rank normal filtering for mesh denoising. Computer Graphics Forum, 37(7):155–166, October 2018. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Li:2015:NLI

[LZL^{+}15] Wei Li, Lei Zhao, Zhijie Lin, Duanqing Xu, and Dongming Lu. Non-local image inpainting using low-rank matrix completion. Computer Graphics Forum, 34(6):111–122, September 2015. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Li:2013:FDG

[LZQ13] Bo Li, Xin Zhao, and Hong Qin. Four-dimensional geometry lens: a novel volumetric magnification approach. Computer Graphics Forum, 32(8):122–133, December 2013. CODEN CG-FODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Li:2013:CSA

[LZW^{+}13] H. Li, H. Zhang, Y. Wang, J. Cao, A. Shamir, and D. Cohen-Or. Curve style analysis in a set of shapes. Computer Graphics Forum, 32(6):77–88, September 2013. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Li:2017:FDR

[LZY^{+}17] Wei Li, Anzong Zheng, Lihua You, Xiaosong Yang, Jianjun Zhang, and Ligang Liu. Fabrication and design: Rib-reinforced shell structure. Computer Graphics Forum, 36(7):15–27, October 2017. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Mora:2012:LVE

[MAAG12] F. Mora, L. Aveneau, O. Apostu, and D. Ghazanfarpour. Lazy visibility evaluation for exact soft shadows. Computer Graphics Forum, 31(1):132–145, February 2012. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Marton:2019:FGA

[MAG19] Fabio Marton, Marco Agus, and Enrico Gobbetti. A framework for GPU-accelerated exploration of massive time-varying rectilinear scalar volumes. Computer Graphics Forum, 38(3): 53–66, June 2019. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Mellado:2014:PCS

[MAM14] Nicolas Mellado, Dror Aiger, and Niloy J. Mitra. Point clouds: Super 4PCS fast global pointcloud registration via smart indexing. Computer Graphics Forum, 33(5):205–215, August 2014. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Manak:2016:VAE

[Man16] M. Manak. Voronoi et al.: Exploration of empty space among spherical obstacles via additively weighted Voronoi diagram. Computer Graphics Forum, 35(5):249–258, August 2016. CO-DEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Meister:2018:PRB

[MB18] D. Meister and J. Bittner. Parallel reinsertion for bounding volume hierarchy optimization. Computer Graphics Forum, 37 (2):463–473, May 2018. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Marques:2019:OSW

[MBB19] Ricardo Marques, Christian Bouville, and Kadi Bouatouch. Optimal sample weights for hemispherical integral quadratures. Computer Graphics Forum, 38(1):59–72, February 2019. CO-DEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Melo:2015:ETM

[MBDC15] M. Melo, M. Bessa, K. Debattista, and A. Chalmers. Evaluation of tone-mapping operators for HDR video under different ambient luminance levels. Computer Graphics Forum, 34(8): 38–49, December 2015. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Machado:2016:SCP

[MBES16] Gustavo Machado, Sebastian Boblest, Thomas Ertl, and Filip Sadlo. Structures, clusters, and patterns: Space-time bifurcation lines for extraction of 2D Lagrangian coherent structures. Computer Graphics Forum, 35(3):91–100, June 2016. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Marvie:2012:GPG

[MBG^{+}12] Jean-Eudes Marvie, Cyprien Buron, Pascal Gautron, Patrice Hirtzlin, and Gaël Sourimant. Geometry processing: GPU shape grammars. Computer Graphics Forum, 31(7pt1):2087– 2095, September 2012. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Mattausch:2015:RTR

[MBJ^{+}15] O. Mattausch, J. Bittner, A. Jaspe, E. Gobbetti, M. Wimmer, and R. Pajarola. Real-time rendering & quantization: CHC+RT: Coherent hierarchical culling for ray tracing. Computer Graphics Forum, 34(2):537–548, May 2015. CODEN CG-FODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Mantiuk:2013:LEG

[MBM13] R. Mantiuk, B. Bazyluk, and R. K. Mantiuk. Looking and editing: Gaze-driven object tracking for real time rendering. Computer Graphics Forum, 32(2pt2):163–173, May 2013. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Masoodian:2015:SFC

[MbMYR15] M. Masoodian, A. b. Mohd Yusof, and B. Rogers. Supporting focus and context awareness in 3D modelling tasks using multi-layered displays. Computer Graphics Forum, 34(6):1–12, September 2015. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Marques:2013:SFP

[MBR^{+}13] R. Marques, C. Bouville, M. Ribardière, L. P. Santos, and K. Bouatouch. Spherical Fibonacci point sets for illumination integrals. Computer Graphics Forum, 32(8):134–143, December 2013. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Marnerides:2018:EDC

[MBRHD18] D. Marnerides, T. Bashford-Rogers, J. Hatchett, and K. Debattista. ExpandNet: a deep convolutional neural network for high dynamic range expansion from low dynamic range content. Computer Graphics Forum, 37(2):37–49, May 2018. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Miguel:2012:CCS

[MBT^{+}12] E. Miguel, D. Bradley, B. Thomaszewski, B. Bickel, W. Matusik, M. A. Otaduy, and S. Marschner. Crowds, cloths, and shape deformation: Data-driven estimation of cloth simulation models. Computer Graphics Forum, 31(2pt2):519–528, May 2012. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Meredith:2010:VRV

[MC10a] Jeremy S. Meredith and Hank Childs. Volume reconstruction and visualization: Visualization and analysis-oriented reconstruction of material interfaces. Computer Graphics Forum, 29 (3):1241–1250, June 2010. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Michael:2010:FIF

[MC10b] D. Michael and Y. Chrysanthou. Fullsphere irradiance factorization for real-time all-frequency illumination for dynamic scenes. Computer Graphics Forum, 29(8):2516–2529, December 2010. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Mishchenko:2014:PST

[MC14] O. Mishchenko and R. Crawfis. On perception of semitransparent streamlines for three-dimensional flow visualization. Computer Graphics Forum, 33(1):210–221, February 2014. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Mattei:2017:PCD

[MC17] E. Mattei and A. Castrodad. Point cloud denoising via moving RPCA. Computer Graphics Forum, 36(8):123–137, December 2017. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Mandad:2019:ECS

[MC19] M. Mandad and M. Campen. Exact constraint satisfaction for truly seamless parametrization. Computer Graphics Forum, 38 (2):135–145, May 2019. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Merabti:2016:VDU

[MCB16] B. Merabti, M. Christie, and K. Bouatouch. A virtual director using hidden Markov models. Computer Graphics Forum, 35 (8):51–67, December 2016. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

McCloud:2011:KCM

[McC11] Scott McCloud. Keynote: Comics: a medium in transition. Computer Graphics Forum, 30(3):xiii, June 2011. CODEN CG-FODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

McDonald:2010:TQC

[McD10] J. McDonald. Teaching quaternions is not complex. Computer Graphics Forum, 29(8):2447–2455, December 2010. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Mendes:2019:SVO

[MCG^{+}19] D. Mendes, F. M. Caputo, A. Giachetti, A. Ferreira, and J. Jorge. A survey on 3D virtual object manipulation: From the desktop to immersive virtual environments. Computer Graphics Forum, 38(1):21–45, February 2019. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Marras:2013:EIA

[MCH13] S. Marras, T. J. Cashman, and K. Hormann. Efficient interpolation of articulated shapes using mixed shape spaces. Computer Graphics Forum, 32(8):258–270, December 2013. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Ma:2018:PIO

[MCHW18] Y. Ma, Z. Chen, W. Hu, and W. Wang. Packing irregular objects in 3D space via hybrid optimization. Computer Graphics Forum, 37(5):49–59, August 2018. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Muller:2018:CJ

[MCJM18] Matthias Müller, Nuttapong Chentanez, Stefan Jeschke, and Miles Macklin. Cable joints. Computer Graphics Forum, 37 (8):1–10, December 2018. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Martin:2012:CGG

[MCM^{+}12] Tobias Martin, Guoning Chen, Suraj Musuvathy, Elaine Cohen, and Charles Hansen. Computational geometry and geometry processing: Generalized swept mid-structure for polygonal models. Computer Graphics Forum, 31(2pt4):805–814, May 2012. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Ma:2019:RTF

[MD19] L. Ma and Z. Deng. Real-time facial expression transformation for monocular RGB video. Computer Graphics Forum, 38(1): 470–481, February 2019. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Medeiros:2014:RUP

[MDBS14] Esdras Medeiros, Harish Doraiswamy, Matthew Berger, and Claudio T. Silva. Rendering: Using physically based rendering to benchmark structured light scanners. Computer Graphics Forum, 33(7):71–80, October 2014. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Mullen:2010:RIS

[MDD^{+}10] Patrick Mullen, Fernando De Goes, Mathieu Desbrun, David Cohen-Steiner, and Pierre Alliez. Reconstruction II: Signing the unsigned: Robust surface reconstruction from raw pointsets. Computer Graphics Forum, 29(5):1733–1741, July 2010. CO-DEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Miao:2018:DDS

[MDI^{+}18] H. Miao, E. De Llano, T. Isenberg, M. E. Gröller, I. Barišić, and I. Viola. DimSUM: Dimension and scale unifying map for visual abstraction of DNA origami structures. Computer Graphics Forum, 37(3):403–413, June 2018. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

MacNeil:2013:VMM

[ME13] S. MacNeil and N. Elmqvist. Visualization mosaics for multivariate visual exploration. Computer Graphics Forum, 32(6): 38–50, September 2013. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Miyamoto:2017:FDS

[MEKM17] Emi Miyamoto, Yuki Endo, Yoshihiro Kanamori, and Jun Mitani. Fabrication and design: Semi-automatic conversion of 3D shape into flat-foldable polygonal model. Computer Graphics Forum, 36(7):41–50, October 2017. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

MachadoeSilva:2014:ISR

[MEMO14] R. Machado e Silva, C. Esperança, R. Marroquim, and A. A. F. Oliveira. Image space rendering of point clouds using the HPR operator. Computer Graphics Forum, 33(1):178–189, February 2014. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Merigot:2011:SMA

[Mér11] Quentin Mérigot. Section 8: a multiscale approach to optimal transport. Computer Graphics Forum, 30(5):1583–1592, August 2011. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Munoz:2011:SAB

[MES^{+}11] Adolfo Munoz, Jose I. Echevarria, Francisco J. Seron, Jorge Lopez-Moreno, Mashhuda Glencross, and Diego Gutierrez. Surface appearance: BSSRDF estimation from single images. Computer Graphics Forum, 30(2):455–464, April 2011. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Munoz:2011:CBS

[MESG11] Adolfo Munoz, Jose I. Echevarria, Francisco J. Seron, and Diego Gutierrez. Convolution-based simulation of homogeneous subsurface scattering. Computer Graphics Forum, 30(8):2279– 2287, December 2011. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Meulemans:2019:EOO

[Meu19] W. Meulemans. Efficient optimal overlap removal: Algorithms and experiments. Computer Graphics Forum, 38(3):713–723, June 2019. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Molchanov:2013:PCR

[MFL13] Vladimir Molchanov, Alexey Fofonov, and Lars Linsen. Projections: Continuous representation of projected attribute spaces of multifields over any spatial sampling. Computer Graphics Forum, 32(3pt3):301–310, June 2013. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Mamani:2013:PUD

[MFNP13] G. M. H. Mamani, F. M. Fatore, L. G. Nonato, and F. V. Paulovich. Projections: User-driven feature space transformation. Computer Graphics Forum, 32(3pt3):291–299, June 2013. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Markus:2015:FRI

[MFPA15] Nenad Markuš, Marco Fratarcangeli, Igor S. Pandžić, and Jörgen Ahlberg. Fast rendering of image mosaics and ASCII art. Computer Graphics Forum, 34(6):251–261, September 2015. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Mojzik:2018:HFU

[MFW18] M. Mojzík, A. Fichet, and A. Wilkie. Handling fluorescence in a uni-directional spectral path tracer. Computer Graphics Forum, 37(4):77–94, July 2018. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Menzel:2010:VPT

[MG10] Nicolas Menzel and Michael Guthe. Visualization and perception: Towards perceptual simplification of models with arbitrary materials. Computer Graphics Forum, 29(7):2261–2270, September 2010. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Mellado:2012:SAG

[MGB^{+}12] Nicolas Mellado, Gaël Guennebaud, Pascal Barla, Patrick Reuter, and Christophe Schlick. Shape analysis: Growing least squares for the analysis of manifolds in scale-space. Computer Graphics Forum, 31(5):1691–1701, August 2012. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Mindek:2014:MSS

[MGB14] P. Mindek, M. E. Gröller, and S. Bruckner. Managing spatial selections with contextual snapshots. Computer Graphics Forum, 33(8):132–144, December 2014. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Milliez:2016:TMP

[MGC^{+}16] A. Milliez, M. Guay, M.-P. Cani, M. Gross, and R. W. Sumner. Textures/mapping: Programmable animation texturing using motion stamps. Computer Graphics Forum, 35(7):67–75, October 2016. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Marechal:2010:VWH

[MGG^{+}10a] N. Maréchal, E. Guérin, E. Galin, S. Mérillou, and N. Mérillou. Virtual worlds: Heat transfer simulation for modeling realistic winter sceneries. Computer Graphics Forum, 29(2):449–458, May 2010. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Martinek:2010:SDO

[MGG10b] M. Martinek, R. Grosso, and G. Greiner. A shape descriptor for 3D objects based on rotational symmetry. Computer Graphics Forum, 29(8):2328–2339, December 2010. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Marco:2019:PTP

[MGJ^{+}19] Julio Marco, Ibón Guillén, Wojciech Jarosz, Diego Gutierrez, and Adrian Jarabo. Progressive transient photon beams. Computer Graphics Forum, 38(6):19–30, September 2019. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

McGee:2019:SAM

[MGM^{+}19] F. McGee, M. Ghoniem, G. Melançon, B. Otjacques, and B. Pinaud. The state of the art in multilayer network visualization. Computer Graphics Forum, 38(6):125–149, September 2019. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Muller:2017:MCA

[MGN17] Thomas Müller, Markus Gross, and Jan Novák. Monte Carlo after coffee: Practical path guiding for efficient light-transport simulation. Computer Graphics Forum, 36(4):91–100, July 2017. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Macedo:2011:HRB

[MGV11] I. Macêdo, J. P. Gois, and L. Velho. Hermite radial basis functions implicits. Computer Graphics Forum, 30(1):27–42, March 2011. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Ma:2018:ISS

[MGY^{+}18] Lei Ma, Jianwei Guo, Dong-Ming Yan, Hanqiu Sun, and Yanyun Chen. Instant stippling on 3D scenes. Computer Graphics Forum, 37(7):255–266, October 2018. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Middendorf:2013:RSP

[MH13] L. Middendorf and Ch. Haubelt. Rendering (session 2): a programmable graphics processor based on partial stream rewriting. Computer Graphics Forum, 32(7):325–334, October 2013. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

McNeill:2017:GTM

[MH17] Graham McNeill and Scott A. Hale. Generating tile maps. Computer Graphics Forum, 36(3):435–445, June 2017. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Maria:2017:CCS

[MHA17] M. Maria, S. Horna, and L. Aveneau. Constrained convex space partition for ray tracing in architectural environments. Computer Graphics Forum, 36(1):288–300, January 2017. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Meng:2016:SID

[MHD16] Johannes Meng, Johannes Hanika, and Carsten Dachsbacher. Sampling: Improving the Dwivedi sampling scheme. Computer Graphics Forum, 35(4):37–44, July 2016. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Muigg:2011:VAV

[MHDG11] Philipp Muigg, Markus Hadwiger, Helmut Doleisch, and Eduard Gröller. 2D visualization by aggregation: Visual coherence for large-scale line-plot visualizations. Computer Graphics Forum, 30(3):643–652, June 2011. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Moritz:2015:THP

[MHHH15] Dominik Moritz, Daniel Halperin, Bill Howe, and Jeffrey Heer. Text & humanities: Perfopticon: Visual query analysis for distributed databases. Computer Graphics Forum, 34(3):71–80, June 2015. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Martschinke:2019:ATS

[MHK^{+}19a] J. Martschinke, S. Hartnagel, B. Keinert, K. Engel, and M. Stamminger. Adaptive temporal sampling for volumetric path tracing of medical data. Computer Graphics Forum, 38 (4):67–76, July 2019. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Mathisen:2019:IID

[MHK^{+}19b] A. Mathisen, T. Horak, C. N. Klokmose, K. Grønbæk, and N. Elmqvist. InsideInsights: Integrating data-driven reporting in collaborative visual analytics. Computer Graphics Forum, 38 (3):649–661, June 2019. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Merzbach:2019:LFS

[MHRK19] S. Merzbach, M. Hermann, M. Rump, and R. Klein. Learned fitting of spatially varying BRDFs. Computer Graphics Forum, 38(4):193–205, July 2019. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Ma:2014:SAD

[MHS^{+}14] Chongyang Ma, Haibin Huang, Alla Sheffer, Evangelos Kalogerakis, and Rui Wang. Surfaces I: Analogy-driven 3D style transfer. Computer Graphics Forum, 33(2):175–184, May 2014. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Moon:2017:NRB

[MIGMM17] Bochang Moon, Jose A. Iglesias-Guitian, Steven McDonagh, and Kenny Mitchell. Noise reduction on G-buffers for Monte Carlo filtering. Computer Graphics Forum, 36(8):600–612, December 2017. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Mattausch:2013:LEF

[MIW13] O. Mattausch, T. Igarashi, and M. Wimmer. Looking and editing: Freeform shadow boundary editing. Computer Graphics Forum, 32(2pt2):175–184, May 2013. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Martin:2013:ENL

[MJBC13] Tobias Martin, Pushkar Joshi, Miklós Bergou, and Nathan Carr. Efficient non-linear optimization via multi-scale gradient filtering. Computer Graphics Forum, 32(6):89–100, September 2013. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Moritz:2017:TST

[MJH^{+}17] Joep Moritz, Stuart James, Tom S. F. Haines, Tobias Ritschel, and Tim Weyrich. Texture stationarization: Turning photos into tileable textures. Computer Graphics Forum, 36(2):177– 188, May 2017. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Menk:2011:VTU

[MJK11] Christoffer Menk, Eduard Jundt, and Reinhard Koch. Visualisation techniques for using spatial augmented reality in the design process of a car. Computer Graphics Forum, 30(8):2354– 2366, December 2011. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Manak:2017:IAC

[MJK17] M. Manak, L. Jirkovsky, and I. Kolingerova. Interactive analysis of connolly surfaces for various probes. Computer Graphics Forum, 36(6):160–172, September 2017. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Moon:2013:RID

[MJL^{+}13] Bochang Moon, Jong Yun Jun, JongHyeob Lee, Kunho Kim, Toshiya Hachisuka, and Sung-Eui Yoon. Robust image denoising using a virtual flash image for Monte Carlo ray tracing. Computer Graphics Forum, 32(1):139–151, February 2013. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Marino:2011:MDV

[MK11] Joseph Marino and Arie Kaufman. Medical data visualization: Prostate cancer visualization from MR imagery and MR spectroscopy. Computer Graphics Forum, 30(3):1051–1060, June 2011. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Mittelstadt:2015:EDE

[MK15] Sebastian Mittelstädt and Daniel A. Keim. Evaluation and design: Efficient contrast effect compensation with personalized perception models. Computer Graphics Forum, 34(3):211–220, June 2015. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Mathew:2019:UWD

[MKMA19] C. D. Tharindu Mathew, Paulo R. Knob, Soraia Raupp Musse, and Daniel G. Aliaga. Urban walkability design using virtual population simulation. Computer Graphics Forum, 38(1): 455–469, February 2019. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Meuschke:2016:FVS

[MKP^{+}16] M. Meuschke, B. Köhler, U. Preim, B. Preim, and K. Lawonn. Flow visualization: Semi-automatic vortex flow classification in 4D PC-MRI data of the aorta. Computer Graphics Forum, 35 (3):351–360, June 2016. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

McGraw:2011:ASC

[MKR11] T. McGraw, T. Kawai, and J. Richards. Allometric scaling for character design. Computer Graphics Forum, 30(1):153– 168, March 2011. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Mwalongo:2016:DVS

[MKRE16] F. Mwalongo, M. Krone, G. Reina, and T. Ertl. Distributed visualization: State-of-the-art report in Web-based visualization. Computer Graphics Forum, 35(3):553–575, June 2016. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Meyer:2012:DPD

[MKSS12] Quirin Meyer, Benjamin Keinert, Gerd Sußner, and Marc Stamminger. Data-parallel decompression of triangle mesh topology. Computer Graphics Forum, 31(8):2541–2553, December 2012. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Miandji:2015:SSC

[MKU15] Ehsan Miandji, Joel Kronander, and Jonas Unger. Sampling & skins: Compressive image reconstruction in reduced union of subspaces. Computer Graphics Forum, 34(2):33–44, May 2015. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

McNabb:2017:SSS

[ML17] Liam McNabb and Robert S. Laramee. Survey of surveys (SoS) — mapping the landscape of survey papers in information visualization. Computer Graphics Forum, 36(3):589–617, June 2017. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Miranda:2018:TLD

[MLD^{+}18] Fabio Miranda, Marcos Lage, Harish Doraiswamy, Charlie Mydlarz, Justin Salamon, Yitzchak Lockerman, Juliana Freire, and Claudio T. Silva. Time lattice: a data structure for the interactive visual analysis of large time series. Computer Graphics Forum, 37(3):23–35, June 2018. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Monch:2013:IMS

[MLK^{+}13] Tobias Mönch, Kai Lawonn, Christoph Kubisch, Rüdiger Westermann, and Bernhard Preim. Interactive mesh smoothing for medical applications. Computer Graphics Forum, 32(8):110– 121, December 2013. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

McLoughlin:2010:TDI

[MLP^{+}10] Tony McLoughlin, Robert S. Laramee, Ronald Peikert, Frits H. Post, and Min Chen. Over two decades of integration-based, geometric flow visualization. Computer Graphics Forum, 29(6): 1807–1829, September 2010. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Ma:2018:SSO

[MM18] Yue Ma and Weiyin Ma. Subdivision schemes with optimal bounded curvature near extraordinary vertices. Computer Graphics Forum, 37(7):455–467, October 2018. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Ma:2019:SSQ

[MM19] Yue Ma and Weiyin Ma. Subdivision schemes for quadrilateral meshes with the least polar artifact in extraordinary regions. Computer Graphics Forum, 38(7):127–139, October 2019. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Maximo:2010:VRH

[MMF10] A. Maximo, R. Marroquim, and R. Farias. Volume rendering: Hardware-assisted projected tetrahedra. Computer Graphics Forum, 29(3):903–912, June 2010. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Merry:2010:SCS

[MMG10] B. Merry, P. Marais, and J. Gain. Simplifying character skins with analytic error metrics. Computer Graphics Forum, 29 (1):13–24, March 2010. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Mattausch:2018:RUS

[MMG18] Oliver Mattausch, Maxim Makhinya, and Orcun Goksel. Realistic ultrasound simulation of complex surface models using interactive Monte-Carlo path tracing. Computer Graphics Forum, 37(1):202–213, February 2018. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Miao:2017:VQC

[MMNG17] H. Miao, G. Mistelbauer, C. Našel, and M. E. Gröller. Visual quantification of the Circle of Willis: An automated identification and standardized representation. Computer Graphics Forum, 36(6):393–404, September 2017. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Miguel:2016:DSO

[MMO16] Eder Miguel, David Miraut, and Miguel A. Otaduy. Deformable & soft objects: Modeling and estimation of energy-based hyperelastic objects. Computer Graphics Forum, 35(2):385–396, May 2016. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Mura:2016:RPP

[MMP16] C. Mura, O. Mattausch, and R. Pajarola. Reconstruction: Piecewise-planar reconstruction of multi-room interiors with arbitrary wall arrangements. Computer Graphics Forum, 35(7): 179–188, October 2016. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Musbach:2013:FWM

[MMRO13] A. Musbach, G. W. Meyer, F. Reitich, and S. H. Oh. Full wave modelling of light propagation and reflection. Computer Graphics Forum, 32(6):24–37, September 2013. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Mistelbauer:2013:MVV

[MMV^{+}13] G. Mistelbauer, A. Morar, A. Varchola, R. Schernthaner, I. Baclija, A. Köchl, A. Kanitsar, S. Bruckner, and E. Gröller. Medical visualization: Vessel visualization using curvicircular feature aggregation. Computer Graphics Forum, 32(3pt2):231– 240, June 2013. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Marra:2017:RCI

[MNP^{+}17] Alessia Marra, Maurizio Nitti, Marios Papas, Thomas Müller, Markus Gross, Wojciech Jarosz, and Jan ovák. Report: 2017 cover image: Mixing bowl. Computer Graphics Forum, 36(1): 345–346, January 2017. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Machado:2010:SCR

[MO10] Gustavo M. Machado and Manuel M. Oliveira. Space & color: Real-time temporal-coherent color contrast enhancement for dichromats. Computer Graphics Forum, 29(3):933–942, June 2010. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Muhler:2010:MVR

[MP10] K. Mühler and B. Preim. Medical visualization: Reusable visualizations and animations for surgery planning. Computer Graphics Forum, 29(3):1103–1112, June 2010. CODEN CG-FODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Mavridis:2012:PTM

[MP12a] Pavlos Mavridis and Georgios Papaioannou. Parameterization and texture mapping: Texture compression using wavelet decomposition. Computer Graphics Forum, 31(7pt1):2107–2116, September 2012. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

McCann:2012:CPS

[MP12b] J. McCann and N. S. Pollard. Computational photography: Soft stacking. Computer Graphics Forum, 31(2pt2):469–478, May 2012. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Munoz-Pandiella:2017:RTS

[MPBM^{+}17] I. Muñoz-Pandiella, C. Bosch, N. Mérillou, X. Pueyo, and S. Mérillou. Real-time solar exposure simulation in complex cities. Computer Graphics Forum, 36(8):554–566, December 2017. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Masia:2012:POC

[MPCG12] Belen Masia, Lara Presa, Adrian Corrales, and Diego Gutierrez. Perceptually optimized coded apertures for defocus deblurring. Computer Graphics Forum, 31(6):1867–1879, September 2012. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Mattausch:2014:GAR

[MPM^{+}14] Oliver Mattausch, Daniele Panozzo, Claudio Mura, Olga Sorkine-Hornung, and Renato Pajarola. Geometry acquisition, reconstruction and analysis: Object detection and classification from large-scale cluttered indoor scans. Computer Graphics Forum, 33(2):11–21, May 2014. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Marcias:2013:MAA

[MPP^{+}13] Giorgio Marcias, Nico Pietroni, Daniele Panozzo, Enrico Puppo, and Olga Sorkine-Hornung. Meshing: Animation-aware quadrangulation. Computer Graphics Forum, 32(5):167–175, August 2013. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Mitra:2013:SGE

[MPWC13] Niloy J. Mitra, Mark Pauly, Michael Wand, and Duygu Ceylan. Symmetry in 3D geometry: Extraction and applications. Computer Graphics Forum, 32(6):1–23, September 2013. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Melvaer:2012:GPC

[MR12] Eivind Lyche Melvær and Martin Reimers. Geodesic polar coordinates on polygonal meshes. Computer Graphics Forum, 31(8):2423–2435, December 2012. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Maisch:2017:SAM

[MR17] S. Maisch and T. Ropinski. Spatial adjacency maps for translucency simulation under general illumination. Computer Graphics Forum, 36(2):443–453, May 2017. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Mueller-Roemer:2017:PPT

[MRAS17] J. S. Mueller-Roemer, C. Altenhofen, and A. Stork. Parallel processing: Ternary sparse matrix representation for volumetric mesh subdivision and processing on GPUs. Computer Graphics Forum, 36(5):59–69, August 2017. CODEN CG-FODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Melzi:2018:LMH

[MRCB18] S. Melzi, E. Rodolà, U. Castellani, and M. M. Bronstein. Localized manifold harmonics for spectral shape analysis. Computer Graphics Forum, 37(6):20–34, September 2018. CODEN CG-FODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

McIntosh:2012:ESB

[MRD12] L. McIntosh, B. E. Riecke, and S. DiPaola. Efficiently simulating the Bokeh of polygonal apertures in a post-process depth of field shader. Computer Graphics Forum, 31(6):1810–1822, September 2012. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Molchanov:2010:VRV

[MRL10] Vladimir Molchanov, Paul Rosenthal, and Lars Linsen. Volume reconstruction and visualization: Non-iterative second-order approximation of signed distance functions for any isosurface representation. Computer Graphics Forum, 29(3):1211–1220, June 2010. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

McKenna:2017:VNF

[MRL^{+}17] S. McKenna, N. Henry Riche, B. Lee, J. Boy, and M. Meyer. Visual narrative flow: Exploring factors shaping data visualization story reading experiences. Computer Graphics Forum, 36 (3):377–387, June 2017. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Madaras:2018:SST

[MRM^{+}18] M. Madaras, A. Riečický, M. Mesároš, M. Stuchlík, and M. Piovarči. Skeletex: Skeleton-texture co-representation for topology-driven real-time interchange and manipulation of surface regions. Computer Graphics Forum, 37(7):325–336, October 2018. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Mehta:2012:MAA

[MRMH12] Soham Uday Mehta, Ravi Ramamoorthi, Mark Meyer, and Christophe Hery. Material appearance: Analytic tangent irradiance environment maps for anisotropic surfaces. Computer Graphics Forum, 31(4):1501–1508, June 2012. CODEN CG-FODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Morvan:2012:EIB

[MRS12] T. Morvan, M. Reimers, and E. Samset. Efficient imagebased proximity queries with object-space precision. Computer Graphics Forum, 31(1):62–74, February 2012. CODEN CG-FODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Mueller-Roemer:2018:GBP

[MRS18] J. S. Mueller-Roemer and A. Stork. GPU-based polynomial finite element matrix assembly for simplex meshes. Computer Graphics Forum, 37(7):443–454, October 2018. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Manson:2010:DGM

[MS10a] Josiah Manson and Scott Schaefer. Differential geometry: Moving least squares coordinates. Computer Graphics Forum, 29 (5):1517–1524, July 2010. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Manson:2010:ISA

[MS10b] Josiah Manson and Scott Schaefer. Implicit surfaces and augmented reality: Isosurfaces over simplicial partitions of multiresolution grids. Computer Graphics Forum, 29(2):377–385, May 2010. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Manson:2011:QSW

[MS11a] J. Manson and S. Schaefer. Quality and scalability: Wavelet rasterization. Computer Graphics Forum, 30(2):395–404, April 2011. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Manson:2011:HDL

[MS11b] Josiah Manson and Scott Schaefer. Hierarchical deformation of locally rigid meshes. Computer Graphics Forum, 30(8):2387– 2396, December 2011. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Manson:2012:GPA

[MS12] Josiah Manson and Scott Schaefer. Geometry: Parameterizationaware MIP-mapping. Computer Graphics Forum, 31(4):1455– 1463, June 2012. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Manson:2013:PPA

[MS13] Josiah Manson and Scott Schaefer. Pixels and points: Analytic rasterization of curves with polynomial filters. Computer Graphics Forum, 32(2pt4):499–507, May 2013. CODEN CG-FODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Manson:2014:GRB

[MS14] Josiah Manson and Scott Schaefer. GPU rendering: Bilinear accelerated filter approximation. Computer Graphics Forum, 33(4):33–40, July 2014. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Manson:2016:FRF

[MS16] Josiah Manson and Peter-Pike Sloan. Faster rendering: Fast filtering of reflection probes. Computer Graphics Forum, 35 (4):119–127, July 2016. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Mavridis:2015:SMO

[MSAP15] P. Mavridis, I. Sipiran, A. Andreadis, and G. Papaioannou. Shape and mesh: Object completion using k-sparse optimization. Computer Graphics Forum, 34(7):13–21, October 2015. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

May:2012:LGU

[MSDK12] T. May, M. Steiger, J. Davey, and J. Kohlhammer. Large graphs: Using signposts for navigation in large graphs. Computer Graphics Forum, 31(3pt2):985–994, June 2012. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

McKenna:2016:NGB

[MSFM16] S. McKenna, D. Staheli, C. Fulcher, and M. Meyer. Networks and graphs 2: BubbleNet: a cyber security dashboard for visualizing patterns. Computer Graphics Forum, 35(3):281–290, June 2016. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Milliez:2018:HTS

[MSGT18] A. Milliez, R. W. Sumner, M. Gross, and B. Thomaszewski. HairControl: a tracking solution for directable hair simulation. Computer Graphics Forum, 37(8):115–123, December 2018. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Meng:2015:MAP

[MSHD15] Johannes Meng, Florian Simon, Johannes Hanika, and Carsten Dachsbacher. Material appearance: Physically meaningful rendering using tristimulus colours. Computer Graphics Forum, 34(4):31–40, July 2015. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Mittelstadt:2014:CPM

[MSK14] S. Mittelstädt, A. Stoffel, and D. A. Keim. Cognition and perception: Methods for compensating contrast effects in information visualization. Computer Graphics Forum, 33(3):231–240, June 2014. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Meyer:2010:CGT

[MSS^{+}10] Quirin Meyer, Jochen Süßmuth, Gerd Sußner, Marc Stamminger, and Günther Greiner. Computer graphics theory: On floating-point normal vectors. Computer Graphics Forum, 29 (4):1405–1409, June 2010. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Manson:2011:MPS

[MSS11] J. Manson, J. Smith, and S. Schaefer. Mesh processing and surface reconstruction: Contouring discrete indicator functions. Computer Graphics Forum, 30(2):385–393, April 2011. CO-DEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Mattausch:2010:HQS

[MSW10] Oliver Mattausch, Daniel Scherzer, and Michael Wimmer. High-quality screen-space ambient occlusion using temporal coherence. Computer Graphics Forum, 29(8):2492–2503, December 2010. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Mennens:2019:SGL

[MSW19] Robin J. P. Mennens, Roeland Scheepens, and Michel A. Westenberg. A stable graph layout algorithm for processes. Computer Graphics Forum, 38(3):725–737, June 2019. CODEN CG-FODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Mattausch:2012:GTI

[MSWI12] Oliver Mattausch, Daniel Scherzer, Michael Wimmer, and Takeo Igarashi. Geometry: Tessellation-independent smooth shadow boundaries. Computer Graphics Forum, 31(4):1465– 1470, June 2012. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Mason:2018:FSL

[MSZ^{+}18] I. Mason, S. Starke, H. Zhang, H. Bilen, and T. Komura. Fewshot learning of homogeneous human locomotion styles. Computer Graphics Forum, 37(7):143–153, October 2018. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Munkberg:2012:OPV

[MTAM12] Jacob Munkberg, Robert Toth, and Tomas Akenine-Möller. Optics: Per-vertex defocus blur for stochastic rasterization. Computer Graphics Forum, 31(4):1385–1389, June 2012. CO-DEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Mantiuk:2012:CFS

[MTM12] Rafał K. Mantiuk, Anna Tomaszewska, and Radosław Mantiuk. Comparison of four subjective methods for image quality assessment. Computer Graphics Forum, 31(8):2478–2491, December 2012. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Muratet:2011:EFP

[MTVJ11] M. Muratet, P. Torguet, F. Viallet, and J. P. Jessel. Experimental feedback on Prog&Play: a serious game for programming practice. Computer Graphics Forum, 30(1):61–73, March 2011. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Munoz:2014:HOR

[Muñ14] Adolfo Muñoz. Higher order ray marching. Computer Graphics Forum, 33(8):167–176, December 2014. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Meuschke:2017:GBC

[MVB^{+}17] M. Meuschke, S. Voß, O. Beuing, B. Preim, and K. Lawonn. Glyph-based comparative stress tensor visualization in cerebral aneurysms. Computer Graphics Forum, 36(3):99–108, June 2017. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Munkberg:2014:RDL

[MVH^{+}14] Jacob Munkberg, Karthik Vaidyanathan, Jon Hasselgren, Petrik Clarberg, and Tomas Akenine-Möller. Reconstruction and denoising: Layered reconstruction for defocus and motion blur. Computer Graphics Forum, 33(4):81–92, July 2014. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Ma:2014:PMG

[MVLS14] Chongyang Ma, Nicholas Vining, Sylvain Lefebvre, and Alla Sheffer. Procedural Modeling I: Game level layout from design specification. Computer Graphics Forum, 33(2):95–104, May 2014. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Martinez:2011:SSC

[MVPG11] J. Martinez, M. Vigo, and N. Pla-Garcia. Section 7: Skeleton computation of orthogonal polyhedra. Computer Graphics Forum, 30(5):1573–1582, August 2011. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Manzi:2016:RTR

[MVZ16] M. Manzi, D. Vicini, and M. Zwicker. Rendering techniques: Regularizing image reconstruction for gradient-domain rendering with feature patches. Computer Graphics Forum, 35(2): 263–273, May 2016. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Maletz:2011:RTR

[MW11] David Maletz and Rui Wang. Ray tracing and real-time rendering: Importance point projection for GPU-based final gathering. Computer Graphics Forum, 30(4):1327–1336, June 2011. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Mazurek:2018:VEQ

[MW18a] Michael Mazurek and Manuela Waldner. Visualizing expanded query results. Computer Graphics Forum, 37(3):87–98, June 2018. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Mirzargar:2018:RCL

[MW18b] Mahsa Mirzargar and Ross T. Whitaker. Representative consensus from limited-size ensembles. Computer Graphics Forum, 37(3):13–22, June 2018. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Musialski:2013:SUR

[MWA^{+}13] P. Musialski, P. Wonka, D. G. Aliaga, M. Wimmer, L. van Gool, and W. Purgathofer. A survey of urban reconstruction. Computer Graphics Forum, 32(6):146–177, September 2013. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Milliez:2013:SMM

[MWCS13] A. Milliez, M. Wand, M.-P. Cani, and H.-P. Seidel. Shape modeling: Mutable elastic models for sculpting structured shapes.

Computer Graphics Forum, 32(2pt1):21–30, May 2013. CO-DEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Manteaux:2017:APB

[MWN^{+}17] P.-L. Manteaux, C. Wojtan, R. Narain, S. Redon, F. Faure, and M.-P. Cani. Adaptive physically based models in computer graphics. Computer Graphics Forum, 36(6):312–337, September 2017. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Meyer:2010:APT

[MWS^{+}10] M. Meyer, B. Wong, M. Styczynski, T. Munzner, and H. Pfister. Applications: Pathline: a tool for comparative functional genomics. Computer Graphics Forum, 29(3):1043–1052, June 2010. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Ma:2016:IPA

[MWS^{+}16] Li-Qian Ma, Jue Wang, Eli Shechtman, Kalyan Sunkavalli, and Shi-Min Hu. Image processing: Appearance harmonization for single image shadow removal. Computer Graphics Forum, 35 (7):189–197, October 2016. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Mizuno:2019:OSA

[MWTI19] Kazuyo Mizuno, Hsiang-Yun Wu, Shigeo Takahashi, and Takeo Igarashi. Optimizing stepwise animation in dynamic set diagrams. Computer Graphics Forum, 38(3):13–24, June 2019. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Musialski:2012:CBI

[MWW12] Przemysław Musialski, Michael Wimmer, and Peter Wonka. Cities and buildings: Interactive coherence-based Façade modeling. Computer Graphics Forum, 31(2pt3):661–670, May 2012. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Mukherjee:2016:RID

[MWW16] R. Mukherjee, X. Wu, and H. Wang. Reconstruction: Incremental deformation subspace reconstruction. Computer Graphics Forum, 35(7):169–178, October 2016. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Ma:2019:TRD

[MYGY19] Li-Ke Ma, Zeshi Yang, Baining Guo, and KangKang Yin. Towards robust direction invariance in character animation. Computer Graphics Forum, 38(7):235–242, October 2019. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Ma:2016:ASL

[MYLZ16] Guanghui Ma, Juntao Ye, Jituo Li, and Xiaopeng Zhang. Anisotropic strain limiting for quadrilateral and triangular cloth meshes. Computer Graphics Forum, 35(1):89–99, February 2016. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Nalbach:2017:LSD

[NAM^{+}17] O. Nalbach, E. Arabadzhiyska, D. Mehta, H.-P. Seidel, and T. Ritschel. Lighting and shading: Deep shading: Convolutional neural networks for screen space shading. Computer Graphics Forum, 36(4):65–78, July 2017. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Nocaj:2012:GDC

[NB12] Arlind Nocaj and Ulrik Brandes. Graphs and diagrams: Computing Voronoi treemaps: Faster, simpler, and resolutionindependent. Computer Graphics Forum, 31(3pt1):855–864, June 2012. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Noel:2015:RTS

[NB15] Laurent Noël and Venceslas Biri. Ray tracing: Skeleton based vertex connection resampling for bidirectional path tracing. Computer Graphics Forum, 34(7):299–307, October 2015. CO-DEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Nishida:2018:PMB

[NBA18] Gen Nishida, Adrien Bousseau, and Daniel G. Aliaga. Procedural modeling of a building from a single image. Computer Graphics Forum, 37(2):415–429, May 2018. CODEN CG-FODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Nishida:2019:MPI

[NBA19] G. Nishida, A. Bousseau, and D. G. Aliaga. Multi-pose interactive linkage design. Computer Graphics Forum, 38(2):277–289, May 2019. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Nguyen:2011:SOA

[NBCW^{+}11] Andy Nguyen, Mirela Ben-Chen, Katarzyna Welnicka, Yinyu Ye, and Leonidas Guibas. Section 4: an optimization approach to improving collections of shape maps. Computer Graphics Forum, 30(5):1481–1491, August 2011. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Naderi:2018:LPB

[NBH18a] Kourosh Naderi, Amin Babadi, and Perttu Hämäläinen. Learning physically based humanoid climbing movements. Computer Graphics Forum, 37(8):69–80, December 2018. CODEN CG-FODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Nasikun:2018:FAL

[NBH18b] Ahmad Nasikun, Christopher Brandt, and Klaus Hildebrandt. Fast approximation of Laplace–Beltrami eigenproblems. Computer Graphics Forum, 37(5):121–134, August 2018. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Nucha:2017:CCT

[NBHN17] Girijanandan Nucha, Georges-Pierre Bonneau, Stefanie Hahmann, and Vijay Natarajan. Computing contour trees for 2D piecewise polynomial functions. Computer Graphics Forum, 36 (3):23–33, June 2017. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Nowrouzezahrai:2014:VSS

[NBMJ14] Derek Nowrouzezahrai, Ilya Baran, Kenny Mitchell, and Wojciech Jarosz. Visibility silhouettes for semi-analytic spherical integration. Computer Graphics Forum, 33(1):105–117, February 2014. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Nielsen:2010:FSI

[NC10] Michael B. Nielsen and Brian B. Christensen. Fluids and smoke: Improved variational guiding of smoke animations. Computer Graphics Forum, 29(2):705–712, May 2010. CODEN CG-FODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Nian:2016:MIP

[NC16] Xianshun Nian and Falai Chen. Matching and interpolation: Planar shape interpolation based on Teichmüller mapping. Computer Graphics Forum, 35(7):43–56, October 2016. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Novak:2012:MMR

[ND12] Jan Novák and Carsten Dachsbacher. Massive models: Rasterized bounding volume hierarchies. Computer Graphics Forum, 31(2pt2):403–412, May 2012. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Nguyen:2014:RDR

[NDD14] H. H. Nguyen, B. Desbenoit, and M. Daniel. Reconstruction and depth: Realistic road path reconstruction from GIS data. Computer Graphics Forum, 33(7):259–268, October 2014. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Newson:2017:SFG

[NDG17] A. Newson, J. Delon, and B. Galerne. A stochastic film grain model for resolution-independent rendering. Computer Graphics Forum, 36(8):684–699, December 2017. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Nishida:2016:EDP

[NGDA16] G. Nishida, I. Garcia-Dorado, and D. G. Aliaga. Exampledriven procedural urban roads. Computer Graphics Forum, 35 (6):5–17, September 2016. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Novak:2018:MCM

[NGHJ18] Jan Novák, Iliyan Georgiev, Johannes Hanika, and Wojciech Jarosz. Monte Carlo methods for volumetric light transport simulation. Computer Graphics Forum, 37(2):551–576, May 2018. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Nader:2014:SAM

[NGM14] G. Nader, G. Guennebaud, and N. Mellado. Surfaces: Adaptive multi-scale analysis for point-based surface editing. Computer Graphics Forum, 33(7):171–179, October 2014. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Nusrat:2019:TTT

[NHG19] S. Nusrat, T. Harbig, and N. Gehlenborg. Tasks, techniques, and tools for genomic data visualization. Computer Graphics Forum, 38(3):781–805, June 2019. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Ngo:2016:NGV

[NHL16] Quynh Quang Ngo, Marc-Thorsten Hütt, and Lars Linsen. Networks and graphs 2: Visual analysis of governing topological structures in excitable network dynamics. Computer Graphics Forum, 35(3):301–310, June 2016. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Nabata:2016:RRE

[NIDN16] K. Nabata, K. Iwasaki, Y. Dobashi, and T. Nishita. Realistic rendering: an error estimation framework for many-light rendering. Computer Graphics Forum, 35(7):431–439, October 2016. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Neugebauer:2011:MDV

[NJB^{+}11] Mathias Neugebauer, Gabor Janiga, Oliver Beuing, Martin Skalej, and Bernhard Preim. Medical data visualization: Anatomy-guided multi-level exploration of blood flow in cerebral aneurysms. Computer Graphics Forum, 30(3):1041–1050, June 2011. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Ni:2019:UDL

[NJC^{+}19] Lixia Ni, Haiyong Jiang, Jianfei Cai, Jianmin Zheng, Haifeng Li, and Xu Liu. Unsupervised dense light field reconstruction with occlusion awareness. Computer Graphics Forum, 38(7): 425–436, October 2019. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Nan:2015:CRT

[NJGW15] Liangliang Nan, Caigui Jiang, Bernard Ghanem, and Peter Wonka. Cities & roads: Template assembly for detailed urban reconstruction. Computer Graphics Forum, 34(2):217–228, May 2015. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Nusrat:2016:ASA

[NK16] Sabrina Nusrat and Stephen Kobourov. Applications: The state of the art in cartograms. Computer Graphics Forum, 35 (3):619–642, June 2016. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Nguyen:2014:CII

[NKB14] R. M. H. Nguyen, S. J. Kim, and M. S. Brown. Color and imaging: Illuminant aware gamut-based color transfer. Computer Graphics Forum, 33(7):319–328, October 2014. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Niessner:2016:RTR

[NKF^{+}16] M. Nießner, B. Keinert, M. Fisher, M. Stamminger, C. Loop, and H. Schäfer. Real-time rendering techniques with hardware tessellation. Computer Graphics Forum, 35(1):113–137, February 2016. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Nguyen:2010:ESM

[NKLN10] Chuong H. Nguyen, Min-Ho Kyung, Joo-Haeng Lee, and Seung-Woo Nam. Editing of shadows and materials: a PCA decomposition for real-time BRDF editing and relighting with global illumination. Computer Graphics Forum, 29(4):1469– 1478, June 2010. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Nakamura:2016:FID

[NKSI16] Morihiro Nakamura, Yuki Koyama, Daisuke Sakamoto, and Takeo Igarashi. Fabrication: an interactive design system of free-formed bamboo-copters. Computer Graphics Forum, 35 (7):323–332, October 2016. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Nivoliers:2013:FAA

[NL13] V. Nivoliers and B. Lévy. Fitting and approximation: Approximating functions on a mesh with restricted Voronoï diagrams. Computer Graphics Forum, 32(5):83–92, August 2013. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Nirel:2018:FPV

[NL18] D. Nirel and D. Lischinski. Fast penetration volume for rigid bodies. Computer Graphics Forum, 37(2):239–250, May 2018. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Neugebauer:2013:MVA

[NLB^{+}13] M. Neugebauer, K. Lawonn, O. Beuing, P. Berg, G. Janiga, and B. Preim. Medical visualization: AmniVis – a system for qualitative exploration of near-wall hemodynamics in cerebral aneurysms. Computer Graphics Forum, 32(3pt3):251–260, June 2013. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Nah:2014:SSA

[NM14] Jae-Ho Nah and Dinesh Manocha. SATO: Surface area traversal order for shadow ray tracing. Computer Graphics Forum, 33 (6):167–177, September 2014. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Nunes:2019:AML

[NMM^{+}19] Augusto L. P. Nunes, Anderson Maciel, Gary W. Meyer, Nigel W. John, Gladimir V. G. Baranoski, and Marcelo Walter. Appearance modelling of living human tissues. Computer Graphics Forum, 38(6):43–65, September 2019. CODEN CG-FODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Nogneng:2018:IFM

[NMR^{+}18] D. Nogneng, S. Melzi, E. Rodolà, U. Castellani, M. Bronstein, and M. Ovsjanikov. Improved functional mappings via product preservation. Computer Graphics Forum, 37(2):179–190, May 2018. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Nobre:2019:SAV

[NMSL19] C. Nobre, M. Meyer, M. Streit, and A. Lex. The state of the art in visualizing multivariate networks. Computer Graphics Forum, 38(3):807–832, June 2019. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Novak:2012:GIP

[NNDJ12] Jan Novák, Derek Nowrouzezahrai, Carsten Dachsbacher, and Wojciech Jarosz. Global illumination: Progressive virtual beam lights. Computer Graphics Forum, 31(4):1407–1413, June 2012. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Nagaraj:2011:MDV

[NNN11] Suthambhara Nagaraj, Vijay Natarajan, and Ravi S. Nanjundiah. Multidimensional data visualization: a gradient-based comparison measure for visual analysis of multifield data. Computer Graphics Forum, 30(3):1101–1110, June 2011. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Nguyen:2015:ICG

[NNRS15] Chuong H. Nguyen, Oliver Nalbach, Tobias Ritschel, and Hans-Peter Seidel. Image collections: Guiding image manipulations using shape-appearance subspaces from co-alignment of image collections. Computer Graphics Forum, 34(2):143–154, May 2015. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Nogneng:2017:IDP

[NO17] Dorian Nogneng and Maks Ovsjanikov. Informative descriptor preservation via commutativity for shape matching. Computer Graphics Forum, 36(2):259–267, May 2017. CODEN CG-FODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Nguyen:2017:CRS

[NPCB17] R. M. H. Nguyen, B. Price, S. Cohen, and M. S. Brown. Coloring, rendering, and sampling: Group-theme recoloring for multi-image color consistency. Computer Graphics Forum, 36 (7):83–92, October 2017. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Neubert:2011:IMV

[NPDD11] B. Neubert, S. Pirk, O. Deussen, and C. Dachsbacher. Improved model- and view-dependent pruning of large botanical scenes. Computer Graphics Forum, 30(6):1708–1718, September 2011. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Nichols:2010:SOI

[NPW10] Greg Nichols, Rajeev Penmatsa, and Chris Wyman. Shadows and order independent transparency: Interactive, multiresolution image-space rendering for dynamic area lighting. Computer Graphics Forum, 29(4):1279–1288, June 2010. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Noguera:2014:OGS

[NREM14] José M. Noguera, Antonio J. Rueda, Miguel A. Espada, and Máximo Martín. Optimized generation of stereoscopic CGI films by 3D image warping. Computer Graphics Forum, 33(8): 145–156, December 2014. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Nguyen:2012:TMM

[NRM^{+}12] Chuong H. Nguyen, Tobias Ritschel, Karol Myszkowski, Elmar Eisemann, and Hans-Peter Seidel. Textures and materials: 3D material style transfer. Computer Graphics Forum, 31 (2pt2):431–438, May 2012. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Nieser:2011:SCC

[NRP11] M. Nieser, U. Reitebuch, and K. Polthier. Section 2: Cube Cover — parameterization of 3D volumes. Computer Graphics Forum, 30(5):1397–1406, August 2011. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Noll:2011:ETE

[NS11] T. Nöll and D. Strieker. Efficient textures: Efficient packing of arbitrary shaped charts for automatic texture Atlas generation. Computer Graphics Forum, 30(4):1309–1317, June 2011. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Najgebauer:2019:IBF

[NS19] Patryk Najgebauer and Rafał Scherer. Inertia-based fast vectorization of line drawings. Computer Graphics Forum, 38(7): 203–213, October 2019. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Nan:2014:RDD

[NSC14] Liangliang Nan, Andrei Sharf, and Baoquan Chen. Reconstruction and depth: 2d-d lifting for shape reconstruction. Computer Graphics Forum, 33(7):249–258, October 2014. CODEN CG-FODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Navarro:2011:MBR

[NSG11] Fernando Navarro, Francisco J. Serón, and Diego Gutierrez. Motion blur rendering: State of the art. Computer Graphics Forum, 30(1):3–26, March 2011. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Neuroth:2019:IVS

[NSM19] T. A. Neuroth, F. Sauer, and K. L. Ma. An interactive visualization system for large sets of phase space trajectories. Computer Graphics Forum, 38(3):297–309, June 2019. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Nalbach:2017:PCR

[NSR17] O. Nalbach, H.-P. Seidel, and T. Ritschel. Practical capture and reproduction of phosphorescent appearance. Computer Graphics Forum, 36(2):409–420, May 2017. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Nguyen:2013:LEM

[NSRS13] Chuong H. Nguyen, Daniel Scherzer, Tobias Ritschel, and Hans-Peter Seidel. Looking and editing: Material editing in complex scenes by surface light field manipulation and reflectance optimization. Computer Graphics Forum, 32(2pt2): 185–194, May 2013. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Noris:2012:SSS

[NSS^{+}12] G. Noris, D. Sýkora, A. Shamir, S. Coros, B. Whited, M. Simmons, A. Hornung, M. Gross, and R. Sumner. Smart scribbles for sketch segmentation. Computer Graphics Forum, 31(8):

2516–2527, December 2012. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Neumann:2013:MCS

[NVH^{+}13] T. Neumann, K. Varanasi, N. Hasler, M. Wacker, M. Magnor, and C. Theobalt. Models: Capture and statistical modeling of arm-muscle deformations. Computer Graphics Forum, 32 (2pt3):285–294, May 2013. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Neumann:2014:CFC

[NVT^{+}14] T. Neumann, K. Varanasi, C. Theobalt, M. Magnor, and M. Wacker. Coordinates and fields: Compressed manifold modes for mesh processing. Computer Graphics Forum, 33 (5):35–44, August 2014. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Nafari:2013:TAV

[NW13] M. Nafari and C. Weaver. Text: Augmenting visualization with natural language translation of interaction: a usability study. Computer Graphics Forum, 32(3pt4):391–400, June 2013. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Nash:2017:DSS

[NW17] C. Nash and C. K. I. Williams. Design and segmentation: The shape variational autoencoder: a deep generative model of partsegmented 3D objects. Computer Graphics Forum, 36(5):1–12, August 2017. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Nader:2016:GVC

[NWHWfD16] G. Nader, K. Wang, F. Hétroy-Wheeler, and f. Dupont. Geometry: Visual contrast sensitivity and discrimination for 3D meshes and their applications. Computer Graphics Forum, 35 (7):497–506, October 2016. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Ni:2018:FAL

[NZH^{+}18] Saifeng Ni, Zichun Zhong, Jin Huang, Wenping Wang, and Xiaohu Guo. Field-aligned and lattice-guided tetrahedral meshing. Computer Graphics Forum, 37(5):161–172, August 2018.

CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Okoe:2014:CPG

[OAJ14] Mershack Okoe, Sayeed Safayet Alam, and Radu Jianu. Cognition and perception: a gaze-enabled graph visualization to improve graph reading tasks. Computer Graphics Forum, 33 (3):251–260, June 2014. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Oster:2018:FTF

[OAM^{+}18] T. Oster, A. Abdelsamie, M. Motejat, T. Gerrits, C. Rössl, D. Thévenin, and H. Theisel. On-the-fly tracking of flame surfaces for the visual analysis of combustion processes. Computer Graphics Forum, 37(6):358–369, September 2018. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Okabe:2011:ACF

[OAO11] Makoto Okabe, Ken Anjyor, and Rikio Onai. Animation: Creating fluid animation from a single image using video database. Computer Graphics Forum, 30(7):1973–1982, September 2011. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Ovsjanikov:2013:AVM

[OBCCG13] M. Ovsjanikov, M. Ben-Chen, F. Chazal, and L. Guibas. Analysis and visualization of maps between shapes. Computer Graphics Forum, 32(6):135–145, September 2013. CODEN CG-FODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Olano:2011:ETV

[OBGB11] M. Olano, D. Baker, W. Griffin, and J. Barczak. Efficient textures: Variable bit rate GPU texture decompression. Computer Graphics Forum, 30(4):1299–1308, June 2011. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Obermaier:2011:VSA

[OBH^{+}11] H. Obermaier, M. I. Billen, H. Hagen, M. Hering-Bertram, and B. Hamann. Visualizing strain anisotropy in mantle flow fields. Computer Graphics Forum, 30(8):2301–2313, December 2011. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Ono:2018:BTS

[ODS18] Jorge P. Ono, Carlos Dietrich, and Claudio T. Silva. Baseball timeline: Summarizing baseball plays into a static visualization. Computer Graphics Forum, 37(3):491–501, June 2018. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Otto:2010:VUV

[OGHT10] Mathias Otto, Tobias Germer, Hans-Christian Hege, and Holger Theisel. Visualization: Uncertain 2D vector field topology. Computer Graphics Forum, 29(2):347–356, May 2010. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Ottley:2019:FCL

[OGW19] Alvitta Ottley, Roman Garnett, and Ran Wan. Follow the clicks: Learning and anticipating mouse interactions during exploratory data analysis. Computer Graphics Forum, 38(3):41– 52, June 2019. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Ovsjanikov:2011:SCN

[OHG11] Maks Ovsjanikov, Qi-Xing Huang, and Leonidas Guibas. Section 5: a condition number for non-rigid shape matching. Computer Graphics Forum, 30(5):1503–1512, August 2011. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Okoe:2015:EGG

[OJ15] Mershack Okoe and Radu Jianu. Evaluation of graphs: GraphUnit: Evaluating interactive graph visualizations using crowdsourcing. Computer Graphics Forum, 34(3):451–460, June 2015. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Oeltze-Jafra:2019:GVE

[OJMN^{+}19] S. Oeltze-Jafra, M. Meuschke, M. Neugebauer, S. Saalfeld, K. Lawonn, G. Janiga, H.-C. Hege, S. Zachow, and B. Preim. Generation and visual exploration of medical flow data: Survey, research trends and future challenges. Computer Graphics Forum, 38(1):87–125, February 2019. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Oelke:2011:ODF

[OJS^{+}11] Daniela Oelke, Halldor Janetzko, Svenja Simon, Klaus Neuhaus, and Daniel A. Keim. Overview & detail, focus+context visualization: Visual boosting in pixel-based visualizations. Computer Graphics Forum, 30(3):871–880, June 2011. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Orthmann:2012:TBA

[OK12] Jens Orthmann and Andreas Kolb. Temporal blending for adaptive SPH. Computer Graphics Forum, 31(8):2436–2449, December 2012. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Ozturk:2010:CBB

[ÖKB10] Aydın Öztürk, Murat Kurt, and Ahmet Bilgili. A copula-based BRDF model. Computer Graphics Forum, 29(6):1795–1806, September 2010. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Oh:2010:RRW

[OKG^{+}10] Se Baek Oh, Sriram Kashyap, Rohit Garg, Sharat Chandran, and Ramesh Raskar. Rendering I: Rendering wave effects with augmented light field. Computer Graphics Forum, 29(2):507– 516, May 2010. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Oelke:2013:TFM

[OKK13] D. Oelke, D. Kokkinakis, and D. A. Keim. Text: Fingerprint matrices: Uncovering the dynamics of social networks in prose literature. Computer Graphics Forum, 32(3pt4):371–380, June 2013. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Oesau:2016:PSD

[OLA16] Sven Oesau, Florent Lafarge, and Pierre Alliez. Planar shape detection and regularization in tandem. Computer Graphics Forum, 35(1):203–215, February 2016. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Oster:2014:ASR

[OLF^{+}14] Timo Oster, Dirk J. Lehmann, Gordon Fru, Holger Theisel, and Dominique Thévenin. Applications: Sparse representation and visualization for direct numerical simulation of premixed combustion. Computer Graphics Forum, 33(3):321–330, June 2014. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Ohrhallinger:2013:EAD

[OM13] S. Ohrhallinger and S. Mudur. An efficient algorithm for determining an aesthetic shape connecting unorganized 2D points. Computer Graphics Forum, 32(8):72–88, December 2013. CO-DEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Ovsjanikov:2010:MOP

[OMMG10] Maks Ovsjanikov, Quentin Mérigot, Facundo Mémoli, and Leonidas Guibas. Matching: One point isometric matching with the heat kernel. Computer Graphics Forum, 29(5):1555– 1564, July 2010. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Ovsjanikov:2013:SAS

[OMPG13] Maks Ovsjanikov, Quentin Mérigot, Viorica Pātrāucean, and Leonidas Guibas. Shape analysis: Shape matching via quotient spaces. Computer Graphics Forum, 32(5):1–11, August 2013. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Ohrhallinger:2016:RCR

[OMW16] S. Ohrhallinger, S. A. Mitchell, and M. Wimmer. Reconstruction: Curve reconstruction with many fewer samples. Computer Graphics Forum, 35(5):167–176, August 2016. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Ou:2010:CGS

[OP10] Jiawei Ou and Fabio Pellacini. Computer graphics systems: SafeGI: Type checking to improve correctness in rendering system implementation. Computer Graphics Forum, 29(4):1269– 1277, June 2010. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Obert:2010:ESM

[OPP10] Juraj Obert, Fabio Pellacini, and Sumanta Pattanaik. Editing of shadows and materials: Visibility editing for all-frequency shadow design. Computer Graphics Forum, 29(4):1441–1449, June 2010. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Oster:2018:CLS

[ORT18] T. Oster, C. Rössl, and H. Theisel. Core lines in 3D secondorder tensor fields. Computer Graphics Forum, 37(3):327–337, June 2018. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Oelke:2014:DTV

[OSR^{+}14] D. Oelke, H. Strobelt, C. Rohrdantz, I. Gurevych, and O. Deussen. Document and text visualization: Comparative exploration of document collections: a visual analytics approach. Computer Graphics Forum, 33(3):201–210, June 2014. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Ogaki:2011:RTR

[OT11] Shinji Ogaki and Yusuke Tokuyoshi. Ray tracing and realtime rendering: Direct ray tracing of Phong tessellation. Computer Graphics Forum, 30(4):1337–1344, June 2011. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Otto:2012:UVA

[OT12] Mathias Otto and Holger Theisel. Uncertainty: Vortex analysis in uncertain vector fields. Computer Graphics Forum, 31(3pt2): 1035–1044, June 2012. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Okura:2015:ACI

[OVB^{+}15] Fumio Okura, Kenneth Vanhoey, Adrien Bousseau, Alexei A. Efros, and George Drettakis. Appearance changes and images: Unifying color and texture transfer for predictive appearance manipulation. Computer Graphics Forum, 34(4):53–63, July 2015. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Otten:2010:FLI

[OVV10] Ron Otten, Anna Vilanova, and Huub Van De Wetering. Fiber & lines: Illustrative White matter fiber bundles. Computer Graphics Forum, 29(3):1013–1022, June 2010. CODEN CG-FODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Ohrhallinger:2019:FCN

[OW19] S. Ohrhallinger and M. Wimmer. FitConnect: Connecting noisy 2D samples by fitted neighbourhoods. Computer Graphics Forum, 38(1):126–137, February 2019. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Ou:2012:SII

[OXKP12] Jiawei Ou, Feng Xie, Parashar Krishnamachari, and Fabio Pellacini. Sampling: ISHair: Importance sampling for hair scattering. Computer Graphics Forum, 31(4):1537–1545, June 2012. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Otsu:2018:RSR

[OYH18] H. Otsu, M. Yamamoto, and T. Hachisuka. Reproducing spectral reflectances from tristimulus colours. Computer Graphics Forum, 37(6):370–381, September 2018. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Patane:2016:SAR

[Pat16] Giuseppe Patané. State of the art reports: STAR — Laplacian spectral kernels and distances for geometry processing and shape analysis. Computer Graphics Forum, 35(2):599–624, May 2016. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Patane:2017:AEC

[Pat17] Giuseppe Patané. Accurate and efficient computation of Laplacian spectral distances and kernels. Computer Graphics Forum, 36(1):184–196, January 2017. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Papazov:2011:SDS

[PB11] C. Papazov and D. Burschka. Section 5: Deformable 3D shape registration based on local similarity transforms. Computer Graphics Forum, 30(5):1493–1502, August 2011. CODEN CG-FODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Pokrass:2013:SCS

[PBB^{+}13] J. Pokrass, A. M. Bronstein, M. M. Bronstein, P. Sprechmann, and G. Sapiro. Shape correspondences: Sparse modeling of intrinsic correspondences. Computer Graphics Forum, 32(2pt4): 459–468, May 2013. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Preim:2016:BVS

[PBC^{+}16] Bernhard Preim, Alexandra Baer, Douglas Cunningham, Tobias Isenberg, and Timo Ropinski. Biomedical visualization: a survey of perceptually motivated 3D visualization of medical image data. Computer Graphics Forum, 35(3):501–525, June 2016. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Petitjean:2018:SGS

[PBE18] Victor Petitjean, Pablo Bauszat, and Elmar Eisemann. Spectral gradient sampling for path tracing. Computer Graphics Forum, 37(4):45–53, July 2018. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Piringer:2010:NAS

[PBK10] H. Piringer, W. Berger, and J. Krasser. Navigating abstract spaces: HyperMoVal: Interactive visual validation of regression models for real-time simulation. Computer Graphics Forum, 29 (3):983–992, June 2010. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Pages:2015:SSM

[PBMG15] R. Pagés, D. Berjón, F. Morán, and N. García. Seamless, static multi-texturing of 3D meshes. Computer Graphics Forum, 34 (1):228–238, February 2015. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Pajot:2011:RTC

[PBPP11] Anthony Pajot, Loïc Barthe, Mathias Paulin, and Pierre Poulin. Ray tracing: Combinatorial bidirectional path-tracing for efficient hybrid CPU/GPU rendering. Computer Graphics Forum, 30(2):315–324, April 2011. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Peng:2012:MMG

[PC12] Chao Peng and Yong Cao. Massive models: a GPU-based approach for massive model rendering with frame-to-frame coherence. Computer Graphics Forum, 31(2pt2):393–402, May 2012. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Park:2016:MIE

[PCBL16] Hangil Park, Youngjin Cho, Seungbae Bang, and Sung-Hee Lee. Matching and interpolation: an Eulerian approach for constructing a map between surfaces with different topologies. Computer Graphics Forum, 35(7):11–19, October 2016. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Palma:2016:DGT

[PCBS16] Gianpaolo Palma, Paolo Cignoni, Tamy Boubekeur, and Roberto Scopigno. Detection of geometric temporal changes in point clouds. Computer Graphics Forum, 35(6):33–45, September 2016. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Palma:2012:MAS

[PCDS12] Gianpaolo Palma, Marco Callieri, Matteo Dellepiane, and Roberto Scopigno. Material appearance: a statistical method for SVBRDF approximation from video sequences in general lighting conditions. Computer Graphics Forum, 31(4):1491– 1500, June 2012. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Pavic:2010:HB

[PCK10] Darko Pavić, Marcel Campen, and Leif Kobbelt. Hybrid booleans. Computer Graphics Forum, 29(1):75–87, March 2010. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Pouli:2011:SIS

[PCR11] Tania Pouli, Douglas W. Cunningham, and Erik Reinhard. A survey of image statistics relevant to computer graphics. Computer Graphics Forum, 30(6):1761–1788, September 2011. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Perrier:2018:SLD

[PCX^{+}18] Hélène Perrier, David Coeurjolly, Feng Xie, Matt Pharr, Pat Hanrahan, and Victor Ostromoukhov. Sequences with lowdiscrepancy blue-noise 2-D projections. Computer Graphics Forum, 37(2):339–353, May 2018. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Petrelli:2016:PRL

[PD16] Alioscia Petrelli and Luigi Di Stefano. Pairwise registration by local orientation cues. Computer Graphics Forum, 35(6):59–72, September 2016. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Pintus:2019:SAM

[PDC^{+}19] R. Pintus, T. G. Dulecha, I. Ciortan, E. Gobbetti, and A. Giachetti. State-of-the-art in multi-light image collections for surface visualization and analysis. Computer Graphics Forum, 38 (3):909–934, June 2019. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Peytavie:2019:PR

[PDG^{+}19] A. Peytavie, T. Dupont, E. Guérin, Y. Cortial, B. Benes, J. Gain, and E. Galin. Procedural riverscapes. Computer Graphics Forum, 38(7):35–46, October 2019. CODEN CG-FODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Papas:2014:LRP

[PdMJ14] Marios Papas, Krystle de Mesa, and Henrik Wann Jensen. Light and reflectance: a physically-based BSDF for modeling the appearance of paper. Computer Graphics Forum, 33(4):133–142, July 2014. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Panozzo:2015:FTM

[PDP^{+}15] Daniele Panozzo, Olga Diamanti, Sylvain Paris, Marco Tarini, Evgeni Sorkine, and Olga Sorkine-Hornung. Fabrication: Texture mapping real-world objects with hydrographics. Computer Graphics Forum, 34(5):65–75, August 2015. CODEN CG-FODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Poco:2015:TET

[PDV^{+}15] Jorge Poco, Harish Doraiswamy, Huy. T. Vo, João L. D. Comba, Juliana Freire, and Cláudio. T. Silva. Traffic: Exploring traffic dynamics in urban environments using vector-valued functions. Computer Graphics Forum, 34(3):161–170, June 2015. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Poco:2014:ASV

[PDW^{+}14] J. Poco, A. Dasgupta, Y. Wei, W. Hargrove, C. Schwalm, R. Cook, E. Bertini, and C. Silva. Applications: Similarity-Explorer: a visual inter-comparison tool for multifaceted climate data. Computer Graphics Forum, 33(3):341–350, June 2014. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Paulovich:2011:MDV

[PEP^{+}11a] F. V. Paulovich, D. M. Eler, J. Poco, C. P. Botha, R. Minghim, and L. G. Nonato. Multidimensional data visualization: Piece wise laplacian-based projection for interactive data exploration and organization. Computer Graphics Forum, 30(3):1091–1100, June 2011. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Poco:2011:MDV

[PEP^{+}11b] J. Poco, R. Etemadpour, F. V. Paulovich, T. V. Long, P. Rosenthal, M. C. F. Oliveira, L. Linsen, and R. Minghim. Multidimensional data visualization: a framework for exploring multidimensional data with 3D projections. Computer Graphics Forum, 30(3):1111–1120, June 2011. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Poco:2012:MVE

[PEPM12] Jorge Poco, Danilo M. Eler, Fernando V. Paulovich, and Rosane Minghim. Medical visualization: Employing 2D projections for fast visual exploration of large fiber tracking data. Computer Graphics Forum, 31(3pt2):1075–1084, June 2012. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Peters:2010:LEL

[Pet10] Thomas J. Peters. Letters to the editors: Letters to the Editors. Computer Graphics Forum, 29(6):2008–2009, September 2010. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Phan:2015:DPF

[PFC15] H. Q. Phan, H. Fu, and A. B. Chan. Drawing and painting: FlexyFont: Learning transferring rules for flexible typeface synthesis. Computer Graphics Forum, 34(7):245–256, October 2015. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Pezzotti:2018:MVE

[PFH^{+}18] Nicola Pezzotti, Jean-Daniel Fekete, Thomas Höllt, Boudewijn P. F. Lelieveldt, Elmar Eisemann, and Anna Vilanova. Multiscale visualization and exploration of large bipartite graphs. Computer Graphics Forum, 37(3):549–560, June 2018. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Peer:2018:ISF

[PGBT18] Andreas Peer, Christoph Gissler, Stefan Band, and Matthias Teschner. An implicit SPH formulation for incompressible linearly elastic solids. Computer Graphics Forum, 37(6):135–148, September 2018. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Peytavie:2010:RGH

[PGGM10] Adrien Peytavie, Eric Galin, Jérôme Grosjean, and Stéphane Mérillou. Reports: a grasping hand, made of small stones. Computer Graphics Forum, 29(1):250–251, March 2010. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Poranne:2010:SRU

[PGK10] R. Poranne, C. Gotsman, and D. Keren. 3D surface reconstruction using a generalized distance function. Computer Graphics Forum, 29(8):2479–2491, December 2010. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Perard-Gayot:2017:GRT

[PGKS17] Arsène Pérard-Gayot, Javor Kalojanov, and Philipp Slusallek. GPU ray tracing using irregular grids. Computer Graphics Forum, 36(2):477–486, May 2017. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Partl:2016:NGP

[PGS^{+}16] C. Partl, S. Gratzl, M. Streit, A. M. Wassermann, H. Pfister, D. Schmalstieg, and A. Lex. Networks and graphs 1: Pathfinder: Visual analysis of paths in graphs. Computer Graphics Forum, 35(3):71–80, June 2016. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Popov:2013:VAQ

[PGSD13] Stefan Popov, Iliyan Georgiev, Philipp Slusallek, and Carsten Dachsbacher. Visibility: Adaptive quantization visibility caching. Computer Graphics Forum, 32(2pt4):399–408, May 2013. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Pintore:2019:AMC

[PGVG19] Giovanni Pintore, Fabio Ganovelli, Alberto Jaspe Villanueva, and Enrico Gobbetti. Automatic modeling of cluttered multiroom floor plans from panoramic images. Computer Graphics Forum, 38(7):347–358, October 2019. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Pothkow:2013:UNM

[PH13] Kai Pöthkow and Hans-Christian Hege. Uncertainty: Nonparametric models for uncertainty visualization. Computer Graphics Forum, 32(3pt2):131–140, June 2013. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Poco:2017:REV

[PH17] Jorge Poco and Jeffrey Heer. Reverse-engineering visualizations: Recovering visual encodings from chart images. Computer Graphics Forum, 36(3):353–363, June 2017. CODEN CG-FODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Pajak:2011:QSS

[PHE^{+}11] Dawid Paja̧k, Robert Herzog, Elmar Eisemann, Karol Myszkowski, and Hans-Peter Seidel. Quality and scalability: Scalable remote rendering with depth and motion-flow augmented streaming. Computer Graphics Forum, 30(2):415–424, April 2011. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Peng:2010:GFO

[PHK^{+}10] Jingliang Peng, Yan Huang, C.-C. Jay Kuo, Ilya Eckstein, and M. Gopi. Geometry I: Feature oriented progressive lossless mesh coding. Computer Graphics Forum, 29(7):2029–2038, September 2010. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Pezzotti:2016:HDD

[PHL^{+}16] N. Pezzotti, T. Höllt, B. Lelieveldt, E. Eisemann, and A. Vilanova. High-dimensional data: Hierarchical stochastic neighbor embedding. Computer Graphics Forum, 35(3):21–30, June 2016. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Pajak:2014:IIP

[PHM^{+}14] Dawid Paja̧k, Robert Herzog, Radosław Mantiuk, Piotr Didyk, Elmar Eisemann, Karol Myszkowski, and Kari Pulli. Images II: Perceptual depth compression for stereo applications. Computer Graphics Forum, 33(2):195–204, May 2014. CODEN CG-FODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Pan:2012:AIW

[PHTB12] Zherong Pan, Jin Huang, Yiying Tong, and Hujun Bao. Animation and interaction: Wake synthesis for shallow water equation. Computer Graphics Forum, 31(7pt1):2029–2036, September 2012. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Papas:2011:IGG

[PJJ^{+}11] Marios Papas, Wojciech Jarosz, Wenzel Jakob, Szymon Rusinkiewicz, Wojciech Matusik, and Tim Weyrich. Intelligent graphics: Goal-based caustics. Computer Graphics Forum, 30 (2):503–511, April 2011. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Parulek:2014:CLD

[PJR^{+}14] Julius Parulek, Daniel Jönsson, Timo Ropinski, Stefan Bruckner, Anders Ynnerman, and Ivan Viola. Continuous levelsof-detail and visual abstraction for seamless molecular visualization. Computer Graphics Forum, 33(6):276–287, September 2014. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Prevost:2015:VFR

[PJSH15] Romain Prévost, Wojciech Jarosz, and Olga Sorkine-Hornung. A vectorial framework for ray traced diffusion curves. Computer Graphics Forum, 34(1):253–264, February 2015. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Pavic:2010:IPT

[PK10] Darko Pavić and Leif Kobbelt. Images and photography: Twocolored pixels. Computer Graphics Forum, 29(2):743–752, May 2010. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Prada:2015:GIU

[PK15] F. Prada and M. Kazhdan. Geometry and images: Unconditionally stable shock filters for image and geometry processing. Computer Graphics Forum, 34(5):201–210, August 2015. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Pretorius:2015:BVC

[PKE15] A. J. Pretorius, I. A. Khan, and R. J. Errington. Biomedical visualization: Cell lineage visualisation. Computer Graphics Forum, 34(3):21–30, June 2015. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Pretorius:2017:SVL

[PKE17] A. J. Pretorius, I. A. Khan, and R. J. Errington. A survey of visualization for live cell imaging. Computer Graphics Forum, 36(1):46–63, January 2017. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Potter:2010:MDV

[PKRJ10] K. Potter, J. Kniss, R. Riesenfeld, and C. R. Johnson. Multivariate data: Visualizing summary statistics and uncertainty. Computer Graphics Forum, 29(3):823–832, June 2010. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Pabst:2010:CFS

[PKS10] Simon Pabst, Artur Koch, and Wolfgang Straßer. Collisions: Fast and scalable CPU/GPU collision detection for rigid and deformable surfaces. Computer Graphics Forum, 29(5):1605– 1612, July 2010. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Piuze:2011:NPG

[PKS11] Emmanuel Piuze, Paul G. Kry, and Kaleem Siddiqi. Natural phenomena: Generalized helicoids for modeling hair geometry. Computer Graphics Forum, 30(2):247–256, April 2011. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Park:2011:FSS

[PLL11] Jong Pil Park, Kang Hoon Lee, and Jehee Lee. Finding syntactic structures from human motion data. Computer Graphics Forum, 30(8):2183–2193, December 2011. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Parakkat:2016:RCT

[PM16] Amal Dev Parakkat and Ramanathan Muthuganapathy. Reconstruction: Crawl through neighbors: a simple curve reconstruction algorithm. Computer Graphics Forum, 35(5):177–186, August 2016. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Pratapa:2019:HHM

[PM19] S. Pratapa and D. Manocha. HMLFC: Hierarchical motioncompensated light field compression for interactive rendering. Computer Graphics Forum, 38(8):1–12, November 2019. CO-DEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Pinaud:2012:GRP

[PMD12] B. Pinaud, G. Melançon, and J. Dubois. Graphs and relationships: PORGY: a visual graph rewriting environment for complex systems. Computer Graphics Forum, 31(3pt4):1265– 1274, June 2012. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Pejsa:2013:ASP

[PMG13] Tomislav Pejsa, Bilge Mutlu, and Michael Gleicher. Animation: Stylized and performative gaze for character animation. Computer Graphics Forum, 32(2pt2):143–152, May 2013. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Prashnani:2017:PBA

[PNVS17] Ekta Prashnani, Maneli Noorkami, Daniel Vaquero, and Pradeep Sen. A phase-based approach for animating images using video examples. Computer Graphics Forum, 36(6):303– 311, September 2017. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Pratapa:2019:TFT

[POCM19] S. Pratapa, T. Olson, A. Chalfin, and D. Manocha. TexNN: Fast texture encoding using neural networks. Computer Graphics Forum, 38(1):328–339, February 2019. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Poranne:2013:IPO

[POG13] Roi Poranne, Elena Ovreiu, and Craig Gotsman. Interactive planarization and optimization of 3D meshes. Computer Graphics Forum, 32(1):152–163, February 2013. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Pagot:2011:FVE

[POS^{+}11a] C. Pagot, D. Osmari, F. Sadlo, D. Weiskopf, T. Ertl, and J. Comba. Flow visualization: Efficient parallel vectors feature extraction from higher-order data. Computer Graphics Forum, 30(3):751–760, June 2011. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Post:2011:ITD

[Pos11b] Frits H. Post. Invited talks: Data visualization: Featuring interactive visual analysis. Computer Graphics Forum, 30(2): xxiii, April 2011. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Pejsa:2010:SAE

[PP10] T. Pejsa and I. S. Pandzic. State of the art in example-based motion synthesis for virtual characters in interactive applications. Computer Graphics Forum, 29(1):202–226, March 2010. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Panozzo:2011:IHQ

[PP11] Daniele Panozzo and Enrico Puppo. Implicit hierarchical quaddominant meshes. Computer Graphics Forum, 30(6):1617– 1629, September 2011. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Piringer:2012:VAS

[PPBT12] H. Piringer, S. Pajer, W. Berger, and H. Teichmann. Visual analysis in science and engineering: Comparative visual analysis of 2D function ensembles. Computer Graphics Forum, 31 (3pt3):1195–1204, June 2012. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Pobitzer:2011:SAT

[PPF^{+}11] Armin Pobitzer, Ronald Peikert, Raphael Fuchs, Benjamin Schindler, Alexander Kuhn, Holger Theisel, Krešimir Matković, and Helwig Hauser. The state of the art in topology-based visualization of unsteady flow. Computer Graphics Forum, 30(6): 1789–1811, September 2011. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Paris:2019:DS

[PPG^{+}19] A. Paris, A. Peytavie, E. Guérin, O. Argudo, and E. Galin. Desertscape simulation. Computer Graphics Forum, 38(7):47– 55, October 2019. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Petz:2012:UPL

[PPH12] Christoph Petz, Kai Pöthkow, and Hans-Christian Hege. Uncertainty: Probabilistic local features in uncertain vector fields with spatial correlation. Computer Graphics Forum, 31(3pt2): 1045–1054, June 2012. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Petronetto:2013:MFD

[PPH^{+}13] F. Petronetto, A. Paiva, E. S. Helou, D. E. Stewart, and L. G. Nonato. Mesh-free discrete Laplace–Beltrami operator. Computer Graphics Forum, 32(6):214–226, September 2013. CO-DEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Paille:2013:FAF

[PPL13] Gilles-Philippe Paillé, Pierre Poulin, and Bruno Lévy. Fitting and approximation: Fitting polynomial volumes to surface meshes with Voronoï squared distance minimization. Computer Graphics Forum, 32(5):103–112, August 2013. CODEN CG-FODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Pietroni:2016:GTF

[PPM^{+}16] Nico Pietroni, Enrico Puppo, Giorgio Marcias, Roberto Roberto, and Paolo Cignoni. Geometry: Tracing field-coherent quad layouts. Computer Graphics Forum, 35(7):485–496, October 2016. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Peethambaran:2019:ILV

[PPT^{+}19] J. Peethambaran, A. D. Parakkat, A. Tagliasacchi, R. Wang, and R. Muthuganapathy. Incremental labelling of Voronoi vertices for shape reconstruction. Computer Graphics Forum, 38 (1):521–536, February 2019. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Panotopoulou:2018:WWP

[PPW18] A. Panotopoulou, S. Paris, and E. Whiting. Watercolor woodblock printing with image analysis. Computer Graphics Forum, 37(2):275–286, May 2018. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Pintus:2016:SGA

[PPY^{+}16] Ruggero Pintus, Kazim Pal, Ying Yang, Tim Weyrich, Enrico Gobbetti, and Holly Rushmeier. A survey of geometric analysis in cultural heritage. Computer Graphics Forum, 35(1):4–31, February 2016. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Pereira:2012:MSG

[PR12] Thiago Pereira and Szymon Rusinkiewicz. Material synthesis: Gamut mapping spatially varying reflectance with an improved BRDF similarity metric. Computer Graphics Forum, 31(4): 1557–1566, June 2012. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Paman:2019:VAR

[PR19] Ashish Paman and Ramsharan Rangarajan. A variational approach to registration with local exponential coordinates. Computer Graphics Forum, 38(1):482–506, February 2019. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Popov:2015:LPP

[PRDD15] Stefan Popov, Ravi Ramamoorthi, Fredo Durand, and George Drettakis. Light paths: Probabilistic connections for bidirectional path tracing. Computer Graphics Forum, 34(4):75–86, July 2015. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Payan:2015:SRT

[PRS15] F. Payan, C. Roudet, and B. Sauvage. Semi-regular triangle remeshing: a comprehensive study. Computer Graphics Forum, 34(1):86–102, February 2015. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Pfaffelmoser:2011:VVV

[PRW11] Tobias Pfaffelmoser, Matthias Reitinger, and Rüdiger Westermann. Volume visualization: Visualizing the positional and geometrical variability of isosurfaces in uncertain scalar fields. Computer Graphics Forum, 30(3):951–960, June 2011. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Pratikakis:2010:REW

[PS10] Ioannis Pratikakis and Michela Spagnuolo. Reports: Eurographics 2009 Workshop on 3D Object Retrieval (EG 3DOR’09) in Cooperation with ACM SIGGRAPH Munich, Germany March 29, 2009. Computer Graphics Forum, 29(1):247–248, March 2010. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Pina:2010:BTD

[PSC10] J. L. Pina, F. Seron, and E. Cerezo. BqR-Tree: a data structure for flights and walkthroughs in urban scenes with mobile elements. Computer Graphics Forum, 29(6):1745–1755, September 2010. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Palma:2018:EVD

[PSCC18] Gianpaolo Palma, Manuele Sabbadin, Massimiliano Corsini, and Paolo Cignoni. Enhanced visualization of detected 3D geometric differences. Computer Graphics Forum, 37(1):159–171, February 2018. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Park:2010:SVM

[PSCN10] Jinho Park, Yeongho Seol, Frederic Cordier, and Junyong Noh. A smoke visualization model for capturing surface-like features. Computer Graphics Forum, 29(8):2352–2362, December 2010. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Popa:2010:RGC

[PSDB^{+}10] T. Popa, I. South-Dickinson, D. Bradley, A. Sheffer, and W. Heidrich. Reconstruction I: Globally consistent space–time reconstruction. Computer Graphics Forum, 29(5):1633–1642, July 2010. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Perazzi:2015:IVP

[PSHZ^{+}15] F. Perazzi, A. Sorkine-Hornung, H. Zimmer, P. Kaufmann, O. Wang, S. Watson, and M. Gross. Image and video processing: Panoramic video from unstructured camera arrays. Computer Graphics Forum, 34(2):57–68, May 2015. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Poulenard:2018:TFO

[PSO18] Adrien Poulenard, Primoz Skraba, and Maks Ovsjanikov. Topological function optimization for continuous shape matching. Computer Graphics Forum, 37(5):13–25, August 2018. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Pegoraro:2010:SRC

[PSP10] Vincent Pegoraro, Mathias Schott, and Steven G. Parker. Scattering and refraction: a closed-form solution to single scattering for general phase functions and light distributions. Computer Graphics Forum, 29(4):1365–1374, June 2010. CODEN CG-FODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Pal:2014:IIC

[PSP^{+}14] Kazim Pal, Christian Schüller, Daniele Panozzo, Olga Sorkine-Hornung, and Tim Weyrich. Images III: Content-aware surface parameterization for interactive restoration of historical documents. Computer Graphics Forum, 33(2):401–409, May 2014. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Paiva:2012:MDM

[PSPM12] Jose Gustavo S. Paiva, William Robson Schwartz, Helio Pedrini, and Rosane Minghim. Multi-dimensional and multivariate data: Semi-supervised dimensionality reduction based on partial least squares for visual analysis of high dimensional data. Computer Graphics Forum, 31(3pt4):1345–1354, June 2012. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Pobitzer:2011:FVE

[PTA^{+}11] A. Pobitzer, M. Tutkun, Ø. Andreassen, R. Fuchs, R. Peikert, and H. Hauser. Flow visualization: Energy-scale aware feature extraction for flow visualization. Computer Graphics Forum, 30 (3):771–780, June 2011. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Patney:2010:CGS

[PTO10] Anjul Patney, Stanley Tzeng, and John D. Owens. Computer graphics systems: Fragment-parallel composite and filter. Computer Graphics Forum, 29(4):1251–1258, June 2010. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Pietroni:2015:SMS

[PTP^{+}15] Nico Pietroni, Davide Tonelli, Enrico Puppo, Maurizio Froli, Roberto Scopigno, and Paolo Cignoni. Splines & meshes: Statics aware grid shells. Computer Graphics Forum, 34(2):627– 641, May 2015. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Paulovich:2012:TDS

[PTT^{+}12] Fernando V. Paulovich, Franklina M. B. Toledo, Guilherme P. Telles, Rosane Minghim, and Luis Gustavo Nonato. Text and documents: Semantic wordification of document collections. Computer Graphics Forum, 31(3pt3):1145–1153, June 2012. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Pal:2013:RIE

[PTW13] Kazim Pal, Melissa Terras, and Tim Weyrich. Reality: Interactive exploration and flattening of deformed historical documents. Computer Graphics Forum, 32(2pt3):327–334, May 2013. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Perin:2018:SAS

[PVS^{+}18] C. Perin, R. Vuillemot, C. D. Stolper, J. T. Stasko, J. Wood, and S. Carpendale. State of the art of sports data visualization. Computer Graphics Forum, 37(3):663–686, June 2018. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Pfaffelmoser:2012:UVG

[PW12] Tobias Pfaffelmoser and Rüdiger Westermann. Uncertainty: Visualization of global correlation structures in uncertain 2D scalar fields. Computer Graphics Forum, 31(3pt2):1025–1034, June 2012. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Peng:2013:MOA

[PW13] Chi-Han Peng and Peter Wonka. Mesh optimization and applications: Connectivity editing for quad-dominant meshes. Computer Graphics Forum, 32(5):43–52, August 2013. CODEN CG-FODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Peethambaran:2017:EUF

[PW17] Jiju Peethambaran and Ruisheng Wang. Enhancing urban façades via LiDAR-based sculpting. Computer Graphics Forum, 36(8):511–528, December 2017. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Pothkow:2011:VVP

[PWH11] Kai Pöthkow, Britta Weber, and Hans-Christian Hege. Volume visualization: Probabilistic marching cubes. Computer Graphics Forum, 30(3):931–940, June 2011. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Panozzo:2012:IVR

[PWS12] Daniele Panozzo, Ofir Weber, and Olga Sorkine. Image & video retargeting: Robust image retargeting via axis-aligned deformation. Computer Graphics Forum, 31(2pt1):229–236, May 2012. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Parvinzamir:2019:MPV

[PZDD19] F. Parvinzamir, Y. Zhao, Z. Deng, and F. Dong. MyEvents: a personal visual analytics approach for mining key events and knowledge discovery in support of personal reminiscence. Computer Graphics Forum, 38(1):647–662, February 2019. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Qin:2018:LBO

[QCW^{+}18] Hongxing Qin, Yi Chen, Yunhai Wang, Xiaoyang Hong, Kangkang Yin, and Hui Huang. Laplace–Beltrami operator on point clouds based on anisotropic Voronoi diagram. Computer Graphics Forum, 37(6):106–117, September 2018. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Quinan:2019:EID

[QPCRM19] P. S. Quinan, L. M. Padilla, S. H. Creem-Regehr, and M. Meyer. Examining implicit discretization in spectral schemes. Computer Graphics Forum, 38(3):363–374, June 2019. CODEN CG-FODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Qiu:2019:TPH

[QWZ^{+}19] H. Qiu, C. Wang, H. Zhu, X. Zhu, J. Gu, and X. Han. Twophase hair image synthesis by self-enhancing generative model. Computer Graphics Forum, 38(7):403–412, October 2019. CO-DEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Qin:2017:MFM

[QYZ17] Yipeng Qin, Hongchuan Yu, and Jianjun Zhang. Meshing: Fast and memory-efficient Voronoi diagram construction on triangle meshes. Computer Graphics Forum, 36(5):93–104, August 2017. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Roberts:2019:MVD

[RAmB^{+}19] J. C. Roberts, H. Al-maneea, P. W. S. Butcher, R. Lew, G. Rees, N. Sharma, and A. Frankenberg-Garcia. Multiple views: different meanings and collocated words. Computer Graphics Forum, 38(3):79–93, June 2019. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Rodriguez:2015:THA

[RAMG15] M. Balsa Rodriguez, M. Agus, F. Marton, and E. Gobbetti. Text & humanities: Adaptive recommendations for enhanced non-linear exploration of annotated 3D objects. Computer Graphics Forum, 34(3):41–50, June 2015. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Rebain:2019:LLS

[RAV^{+}19] Daniel Rebain, Baptiste Angles, Julien Valentin, Nicholas Vining, Jiju Peethambaran, Shahram Izadi, and Andrea Tagliasacchi. LSMAT least squares medial axis transform. Computer Graphics Forum, 38(6):5–18, September 2019. CODEN CG-FODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Rensink:2010:EPC

[RB10] Ronald A. Rensink and Gideon Baldridge. Evaluation: The perception of correlation in scatterplots. Computer Graphics Forum, 29(3):1203–1210, June 2010. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Rodola:2014:SSR

[RBC14] E. Rodolà, S. Rota Bulò, and D. Cremers. Segmentation and structure: Robust region detection via consensus segmentation of deformable shapes. Computer Graphics Forum, 33(5):97– 106, August 2014. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Rodriguez:2018:ERI

[RBDD18] Simon Rodriguez, Adrien Bousseau, Fredo Durand, and George Drettakis. Exploiting repetitions for image-based rendering of facades. Computer Graphics Forum, 37(4):119–131, July 2018. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Ribardiere:2017:SSD

[RBMS17] M. Ribardière, B. Bringier, D. Meneveaux, and L. Simonot. STD: Student’s t-distribution of slopes for microfacet based BSDFs. Computer Graphics Forum, 36(2):421–429, May 2017. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Rostami:2019:SDD

[RBRY19] R. Rostami, F. S. Bashiri, B. Rostami, and Z. Yu. A survey on data-driven 3D shape descriptors. Computer Graphics Forum, 38(1):356–393, February 2019. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Racz:2018:CWB

[RC18] Gergely Ferenc Rácz and Balázs Csébfalvi. Cosine-weighted B-spline interpolation on the face-centered cubic lattice. Computer Graphics Forum, 37(3):503–511, June 2018. CODEN CG-FODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Reed:2019:UGF

[RC19] K. Reed and D. Cosker. User-guided facial animation through an evolutionary interface. Computer Graphics Forum, 38(6): 165–176, September 2019. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Ribardiere:2011:ARI

[RCB11] Mickaël Ribardière, Samuel Carré, and Kadi Bouatouch. Adaptive records for irradiance caching. Computer Graphics Forum, 30(6):1603–1616, September 2011. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Ren:2017:GMU

[RCB^{+}17a] Z. Ren, P. Charalambous, J. Bruneau, Q. Peng, and J. Pettré. Group modeling: a unified velocity-based approach. Computer Graphics Forum, 36(8):45–56, December 2017. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Rodola:2017:PFC

[RCB^{+}17b] E. Rodolà, L. Cosmo, M. M. Bronstein, A. Torsello, and D. Cremers. Partial functional correspondence. Computer Graphics Forum, 36(1):222–236, January 2017. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Reiner:2014:CPD

[RCM^{+}14] Tim Reiner, Nathan Carr, Radomír Měch, Ondřej Št’ava, Carsten Dachsbacher, and Gavin Miller. Cutting and printing: Dual-color mixing for fused deposition modeling printers. Computer Graphics Forum, 33(2):479–486, May 2014. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Raidou:2018:BRV

[RCMA^{+}18] R. G. Raidou, O. Casares-Magaz, A. Amirkhanov, V. Moiseenko, L. P. Muren, J. P. Einck, A. Vilanova, and M. E. Gröller. Bladder Runner: Visual analytics for the exploration of RT-induced bladder toxicity in a cohort study. Computer Graphics Forum, 37(3):205–216, June 2018. CODEN CG-FODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Raidou:2016:PFV

[RCMM^{+}16] R. G. Raidou, O. Casares-Magaz, L. P. Muren, U. A. van der Heide, J. Rørvik, M. Breeuwer, and A. Vilanova. Prediction and forecasting: Visual analysis of tumor control models for prediction of radiotherapy response. Computer Graphics Forum, 35 (3):231–240, June 2016. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Rushmeier:2012:E

[RD12] Holly Rushmeier and Oliver Deussen. Editorial. Computer Graphics Forum, 31(1):1–2, February 2012. CODEN CG-FODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Rushmeier:2013:E

[RD13] Holly Rushmeier and Oliver Deussen. Editorial. Computer Graphics Forum, 32(1):1–2, February 2013. CODEN CG-FODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Rushmeier:2014:Ea

[RD14a] Holly Rushmeier and Oliver Deussen. Editorial. Computer Graphics Forum, 33(1):1, February 2014. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Rushmeier:2014:Eb

[RD14b] Holly Rushmeier and Oliver Deussen. Editorial. Computer Graphics Forum, 33(1):1, February 2014. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Ritschel:2012:SAI

[RDGK12] Tobias Ritschel, Carsten Dachsbacher, Thorsten Grosch, and Jan Kautz. The state of the art in interactive global illumination. Computer Graphics Forum, 31(1):160–188, February 2012. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Rosman:2013:PCS

[RDK13] G. Rosman, A. Dubrovina, and R. Kimmel. Patch-collaborative spectral point-cloud denoising. Computer Graphics Forum, 32 (8):1–12, December 2013. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Ritschel:2012:PGC

[RE12] Tobias Ritschel and Elmar Eisemann. Perceptual graphics: a computational model of afterimages. Computer Graphics Forum, 31(2pt3):529–534, May 2012. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Ritschel:2011:MIS

[REH^{+}11] Tobias Ritschel, Elmar Eisemann, Inwoo Ha, James D. K. Kim, and Hans-Peter Seidel. Making imperfect shadow maps viewadaptive: High-quality global illumination in large dynamic scenes. Computer Graphics Forum, 30(8):2258–2269, December 2011. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Renka:2016:TSM

[Ren16] Robert J. Renka. Two simple methods for improving a triangle mesh surface. Computer Graphics Forum, 35(6):46–58, September 2016. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Rumman:2015:PBS

[RF15] Nadine Abu Rumman and Marco Fratarcangeli. Position-based skinning for soft articulated characters. Computer Graphics Forum, 34(6):240–250, September 2015. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Roesch:2019:VNN

[RG19] Isabelle Roesch and Tobias Günther. Visualization of neural network predictions for weather forecasting. Computer Graphics Forum, 38(1):209–220, February 2019. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Raymond:2014:ROB

[RGB^{+}14] Boris Raymond, Gaël Guennebaud, Pascal Barla, Romain Pacanowski, and Xavier Granier. Rendering: Optimizing BRDF orientations for the manipulation of anisotropic highlights. Computer Graphics Forum, 33(2):313–321, May 2014. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Rodriguez:2014:SAC

[RGG^{+}14] M. Balsa Rodríguez, E. Gobbetti, J. A. Iglesias Guitián, M. Makhinya, F. Marton, R. Pajarola, and S. K. Suter. Stateof-the-art in compressed GPU-Based direct volume rendering. Computer Graphics Forum, 33(6):77–100, September 2014. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Reich:2015:PSI

[RGG15] Andreas Reich, Tobias Günther, and Thorsten Grosch. Perception and simplification: Illumination-driven mesh reduction for accelerating light transport simulations. Computer Graphics Forum, 34(4):165–174, July 2015. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Rojo:2018:VPS

[RGG18] Irene Baeza Rojo, Markus Gross, and Tobias Günther. Visualizing the phase space of heterogeneous inertial particles in 2D flows. Computer Graphics Forum, 37(3):289–300, June 2018. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Radiano:2018:SAC

[RGM^{+}18] O. Radiano, Y. Graber, M. Mahler, L. Sigal, and A. Shamir. Story albums: Creating fictional stories from personal photograph sets. Computer Graphics Forum, 37(1):19–31, February 2018. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Reif:2010:EAR

[RGSK10] R. Reif, W. A. Günthner, B. Schwerdtfeger, and G. Klinker. Evaluation of an augmented reality supported picking system under practical conditions. Computer Graphics Forum, 29(1): 2–12, March 2010. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Rossing:2012:PMR

[RHL12] Christoph Rößing, Johannes Hanika, and Hendrik Lensch. Photo manipulation: Real-time disparity map-based pictorial depth cue enhancement. Computer Graphics Forum, 31(2pt1): 275–284, May 2012. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Ren:2018:CDC

[RHLH18] Bo Ren, Jiahui Huang, Ming C. Lin, and Shi-Min Hu. Controllable dendritic crystal simulation using orientation field. Computer Graphics Forum, 37(2):485–495, May 2018. CODEN CG-FODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Rohrdantz:2012:GVW

[RHM^{+}12] Christian Rohrdantz, Michael Hund, Thomas Mayer, Bernhard Wälchli, and Daniel A. Keim. Geospatial visualization: The World’s languages explorer: Visual analysis of language features in genealogical and areal contexts. Computer Graphics Forum, 31(3pt1):935–944, June 2012. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Ranieri:2012:VVR

[RHS^{+}12] Nicola Ranieri, Simon Heinzle, Quinn Smithwick, Daniel Reetz, Lanny S. Smoot, Wojciech Matusik, and Markus Gross. Visualization and volume rendering: Multi-layered automultiscopic displays. Computer Graphics Forum, 31(7pt2):2135– 2143, September 2012. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Roberts:2017:RRF

[RI17] D. A. T. Roberts and I. Ivrissimtzis. Reevaluating reconstruction filters for path-searching tasks in 3D. Computer Graphics Forum, 36(6):291–302, September 2017. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Rainer:2019:NBC

[RJGW19] Gilles Rainer, Wenzel Jakob, Abhijeet Ghosh, and Tim Weyrich. Neural BTF compression and interpolation. Computer Graphics Forum, 38(2):235–244, May 2019. CODEN CG-FODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Ritchie:2018:EBA

[RJT18] Daniel Ritchie, Sarah Jobalia, and Anna Thomas. Examplebased authoring of procedural modeling programs with structural and continuous variability. Computer Graphics Forum, 37 (2):401–413, May 2018. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Rump:2010:SCS

[RK10] Martin Rump and Reinhard Klein. Sparse computing: Spectralization: Reconstructing spectra from sparse data. Computer Graphics Forum, 29(4):1347–1354, June 2010. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Rachavarapu:2018:WEV

[RKGS18] Kranthi Kumar Rachavarapu, Moneish Kumar, Vineet Gandhi, and Ramanathan Subramanian. Watch to edit: Video retargeting using gaze. Computer Graphics Forum, 37(2):205–215, May 2018. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Rungjiratananon:2010:CSM

[RKN10] W. Rungjiratananon, Y. Kanamori, and T. Nishita. Chain shape matching for simulating complex hairstyles. Computer Graphics Forum, 29(8):2438–2446, December 2010. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Rungjiratananon:2012:IHW

[RKN12] W. Rungjiratananon, Y. Kanamori, and T. Nishita. Imaging and hair: Wetting effects in hair simulation. Computer Graphics Forum, 31(7pt1):1993–2002, September 2012. CODEN CG-FODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Reinert:2016:ERP

[RKR^{+}16] Bernhard Reinert, Johannes Kopf, Tobias Ritschel, Eduardo Cuervo, David Chu, and Hans-Peter Seidel. Efficient rendering: Proxy-guided image-based rendering for mobile devices. Computer Graphics Forum, 35(7):353–362, October 2016. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Reiner:2012:LRS

[RKRD12] Tim Reiner, Anton Kaplanyan, Marcel Reinhard, and Carsten Dachsbacher. Lighting and rendering: Selective inspection and interactive visualization of light transport in virtual scenes. Computer Graphics Forum, 31(2pt3):711–718, May 2012. CO-DEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Richter:2017:CMV

[RKSA17] Ronald Richter, Jan Eric Kyprianidis, Boris Springborn, and Marc Alexa. Constrained modelling of 3-valent meshes using a hyperbolic deformation metric. Computer Graphics Forum, 36(6):62–75, September 2017. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Rieck:2014:SAM

[RL14] B. Rieck and H. Leitte. Structural analysis of multivariate point clouds using simplicial chains. Computer Graphics Forum, 33 (8):28–37, December 2014. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Rieck:2015:TST

[RL15] B. Rieck and H. Leitte. Time-series and topology: Persistent homology for the evaluation of dimensionality reduction schemes. Computer Graphics Forum, 34(3):431–440, June 2015. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Rieck:2016:SCP

[RL16] B. Rieck and H. Leitte. Structures, clusters, and patterns: Exploring and comparing clusterings of multivariate data sets using persistent homology. Computer Graphics Forum, 35(3):81– 90, June 2016. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Rees:2019:SIV

[RL19] D. Rees and R. S. Laramee. A survey of information visualization books. Computer Graphics Forum, 38(1):610–646, February 2019. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Rodola:2019:FMR

[RLB^{+}19] E. Rodolà, Z. Lähner, A. M. Bronstein, M. M. Bronstein, and J. Solomon. Functional maps representation on product manifolds. Computer Graphics Forum, 38(1):678–689, February 2019. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Ritchie:2015:SCG

[RLGH15] Daniel Ritchie, Sharon Lin, Noah D. Goodman, and Pat Hanrahan. Shape collections: Generating design suggestions under tight constraints with gradient-based probabilistic programming. Computer Graphics Forum, 34(2):515–526, May 2015. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Ren:2017:SAT

[RLH17a] Donghao Ren, Bongshin Lee, and Tobias Höllerer. Stardust: Accessible and transparent GPU support for information visualization rendering. Computer Graphics Forum, 36(3):179–188, June 2017. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Ren:2017:REI

[RLH^{+}17b] Xiaohua Ren, Luan Lyu, Xiaowei He, Yanci Zhang, and Enhua Wu. Representing and editing images: Efficient gradientdomain compositing using an approximate curl-free wavelet projection. Computer Graphics Forum, 36(7):207–215, October 2017. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Ren:2018:BWS

[RLH^{+}18] Xiaohua Ren, Luan Lyu, Xiaowei He, Wei Cao, Zhixin Yang, Bin Sheng, Yanci Zhang, and Enhua Wu. Biorthogonal wavelet surface reconstruction using partial integrations. Computer Graphics Forum, 37(7):13–24, October 2018. CODEN CG-FODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Richardt:2014:ISV

[RLMB^{+}14] C. Richardt, J. Lopez-Moreno, A. Bousseau, M. Agrawala, and G. Drettakis. Illustration and stylization: Vectorising bitmaps into semi-transparent gradient layers. Computer Graphics Forum, 33(4):11–19, July 2014. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Riche:2010:EUI

[RLP10] Nathalie Henry Riche, Bongshin Lee, and Catherine Plaisant. Evaluation: Understanding interactive legends: a comparative evaluation with standard widgets. Computer Graphics Forum, 29(3):1193–1202, June 2010. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Ruiz:2014:CPM

[RLYL14] Conrado R. Ruiz, Jr., Sang N. Le, Jinze Yu, and Kok-Lim Low. Cutting and printing: Multi-style paper pop-up designs from 3D models. Computer Graphics Forum, 33(2):487–496, May 2014. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Rodola:2017:RPM

[RMC17] E. Rodolà, M. Moeller, and D. Cremers. Regularized pointwise map recovery from functional correspondence. Computer Graphics Forum, 36(8):700–711, December 2017. CODEN CG-FODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Rufiange:2012:THV

[RMF12] Sébastien Rufiange, Michael J. McGuffin, and Christopher P. Fuhrman. TreeMatrix: a hybrid visualization of compound graphs. Computer Graphics Forum, 31(1):89–101, February 2012. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Rodrigues:2018:PBM

[RMG18] Rui S. V. Rodrigues, José F. M. Morgado, and Abel J. P. Gomes. Part-based mesh segmentation: a survey. Computer Graphics Forum, 37(6):235–274, September 2018. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Rocha:2018:IMV

[RMH^{+}18] A. Rocha, R. C. R. Mota, H. Hamdi, U. R. Alim, and M. Costa Sousa. Illustrative multivariate visualization for geological modelling. Computer Graphics Forum, 37(3):465–477, June 2018. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Renoust:2015:GDV

[RMM15] B. Renoust, G. Melançon, and T. Munzner. Graphs: Detangler: Visual analytics for multiplex networks. Computer Graphics Forum, 34(3):321–330, June 2015. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Rousselle:2013:RSR

[RMZ13] Fabrice Rousselle, Marco Manzi, and Matthias Zwicker. Rendering (session 1): Robust denoising using feature and color information. Computer Graphics Forum, 32(7):121–130, October 2013. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Ray:2010:IS

[RNLL10] Nicolas Ray, Vincent Nivoliers, Sylvain Lefebvre, and Bruno Lévy. Invisible seams. Computer Graphics Forum, 29(4):1489– 1496, June 2010. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Roveri:2017:GPS

[RÖG17] Riccardo Roveri, A. Cengiz Öztireli, and Markus Gross. General point sampling with adaptive density and correlations. Computer Graphics Forum, 36(2):107–117, May 2017. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Roveri:2015:DSS

[RÖM^{+}15] Riccardo Roveri, A. Cengiz Öztireli, Sebastian Martin, Barbara Solenthaler, and Markus Gross. Descriptors and shape synthesis: Example based repetitive structure synthesis. Computer Graphics Forum, 34(5):39–52, August 2015. CODEN CG-FODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Roveri:2018:PCP

[RÖPG18] Riccardo Roveri, A. Cengiz Öztireli, Ioana Pandele, and Markus Gross. PointProNets: Consolidation of point clouds with convolutional neural networks. Computer Graphics Forum, 37(2):87–99, May 2018. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Rosen:2013:PVA

[Ros13] Paul Rosen. Performance: a visual approach to investigating shared and global memory behavior of CUDA kernels. Computer Graphics Forum, 32(3pt2):161–170, June 2013. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Ruhland:2015:REG

[RPA^{+}15] K. Ruhland, C. E. Peters, S. Andrist, J. B. Badler, N. I. Badler, M. Gleicher, B. Mutlu, and R. McDonnell. A review of eye gaze in virtual agents, social robotics and HCI: Behaviour generation, user interaction and perception. Computer Graphics Forum, 34(6):299–326, September 2015. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Riviere:2016:MSR

[RPG16] J. Riviere, P. Peers, and A. Ghosh. Mobile surface reflectometry. Computer Graphics Forum, 35(1):191–202, February 2016. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Reh:2012:VAS

[RPK^{+}12] A. Reh, B. Plank, J. Kastner, E. Gröller, and C. Heinzl. Visual analysis in science and engineering: Porosity maps — interactive exploration and visual analysis of porosity in carbon fiber reinforced polymers. Computer Graphics Forum, 31(3pt3): 1185–1194, June 2012. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Rieder:2011:MDV

[RPLH11] Christian Rieder, Stephan Palmer, Florian Link, and Horst K. Hahn. Medical data visualization: a shader framework for rapid prototyping of GPU-based volume rendering. Computer Graphics Forum, 30(3):1031–1040, June 2011. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Rosenthal:2013:AVI

[RPMO13] P. Rosenthal, L. Pfeiffer, N. H. Müller, and P. Ohler. Applications: VisRuption: Intuitive and efficient visualization of temporal airline disruption data. Computer Graphics Forum, 32 (3pt1):81–90, June 2013. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Ruiz:2017:MBC

[RPPD17] A. L. Cruz Ruiz, C. Pontonnier, N. Pronost, and G. Dumont. Muscle-based control for character animation. Computer Graphics Forum, 36(6):122–147, September 2017. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Riehmann:2015:THV

[RPSF15] P. Riehmann, M. Potthast, B. Stein, and B. Froehlich. Text & humanities: Visual assessment of alleged plagiarism cases.

Computer Graphics Forum, 34(3):61–70, June 2015. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Ren:2019:SRF

[RPWO19] Jing Ren, Mikhail Panine, Peter Wonka, and Maks Ovsjanikov. Structured regularization of functional map computations. Computer Graphics Forum, 38(5):39–53, August 2019. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Razafindrazaka:2015:QPP

[RRP15] Faniry H. Razafindrazaka, Ulrich Reitebuch, and Konrad Polthier. Quads and polygons: Perfect matching quad layouts for manifold meshes. Computer Graphics Forum, 34(5): 219–228, August 2015. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Reinert:2012:VVR

[RRS12] Bernhard Reinert, Tobias Ritschel, and Hans-Peter Seidel. Visualization and volume rendering: Homunculus warping: Conveying importance using self-intersection-free nonhomogeneous mesh deformation. Computer Graphics Forum, 31 (7pt2):2165–2171, September 2012. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Reinert:2016:PBN

[RRSG16] Bernhard Reinert, Tobias Ritschel, Hans-Peter Seidel, and Iliyan Georgiev. Projective blue-noise sampling. Computer Graphics Forum, 35(1):285–295, February 2016. CODEN CG-FODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Ruckert:2019:ESS

[RS19] D. Rückert and M. Stamminger. An efficient solution to structured optimization problems using recursive matrices. Computer Graphics Forum, 38(8):33–39, November 2019. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Richardt:2012:IVR

[RSD^{+}12] Christian Richardt, Carsten Stoll, Neil A. Dodgson, Hans-Peter Seidel, and Christian Theobalt. Image & video retargeting: Coherent spatiotemporal filtering, upsampling and rendering of RGBZ videos. Computer Graphics Forum, 31(2pt1):247– 256, May 2012. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Ruiters:2010:TGP

[RSK10] Roland Ruiters, Ruwen Schnabel, and Reinhard Klein. Texture generation: Patch-based texture interpolation. Computer Graphics Forum, 29(4):1421–1429, June 2010. CODEN CG-FODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Ruiters:2012:LFR

[RSK12] Roland Ruiters, Christopher Schwartz, and Reinhard Klein. Light fields and reflectance: Data driven surface reflectance from sparse and irregular samples. Computer Graphics Forum, 31(2pt1):315–324, May 2012. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Ruiters:2013:SEB

[RSK13] Roland Ruiters, Christopher Schwartz, and Reinhard Klein. Synthesis: Example-based interpolation and synthesis of bidirectional texture functions. Computer Graphics Forum, 32 (2pt3):361–370, May 2013. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Rados:2016:CVI

[RSM^{+}16] S. Radoš, R. Splechtna, K. Matković, M. – Duras, E. Gröller, and H. Hauser. Coordinated views and interaction design: Towards quantitative visual analytics with structured brushing and linked statistics. Computer Graphics Forum, 35(3):251– 260, June 2016. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Rubio-Sanchez:2017:ARA

[RSSL17] M. Rubio-Sánchez, A. Sanchez, and D. J. Lehmann. Adaptable radial axes plots for improved multivariate data visualization. Computer Graphics Forum, 36(3):389–399, June 2017. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Reda:2011:GNV

[RTJ^{+}11] Khairi Reda, Chayant Tantipathananandh, Andrew Johnson, Jason Leigh, and Tanya Berger-Wolf. Graphs & networks: Visualizing the evolution of community structures in dynamic social networks. Computer Graphics Forum, 30(3):1061–1070, June 2011. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Rhodin:2014:AIM

[RTK^{+}14] Helge Rhodin, James Tompkin, Kwang In Kim, Kiran Varanasi, Hans-Peter Seidel, and Christian Theobalt. Animation II (motion data processing): Interactive motion mapping for real-time character control. Computer Graphics Forum, 33(2):273–282, May 2014. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Rustamov:2010:DGB

[Rus10] Raif M. Rustamov. Differential geometry: Barycentric coordinates on surfaces. Computer Graphics Forum, 29(5):1507–1516, July 2010. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Rustamov:2011:SMB

[Rus11] Raif M. Rustamov. Section 6: Multiscale biharmonic kernels. Computer Graphics Forum, 30(5):1521–1531, August 2011. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Raidou:2015:BVV

[RvdHD^{+}15] R. G. Raidou, U. A. van der Heide, C. V. Dinh, G. Ghobadi, J. F. Kallehauge, M. Breeuwer, and A. Vilanova. Biomedical visualization: Visual analytics for the exploration of tumor tissue characterization. Computer Graphics Forum, 34(3):11–20, June 2015. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Raj:2018:VMD

[RW18] M. Raj and R. T. Whitaker. Visualizing multidimensional data with order statistics. Computer Graphics Forum, 37(3):277– 287, June 2018. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Rieder:2010:MVV

[RWS^{+}10] Christian Rieder, Andreas Weihusen, Christian Schumann, Stephan Zidowitz, and Heinz-Otto Peitgen. Medical visualization: Visual support for interactive post-interventional assessment of radiofrequency ablation therapy. Computer Graphics Forum, 29(3):1093–1102, June 2010. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Ruegg:2013:IDS

[RWSG13] Jan Rüegg, Oliver Wang, Aljoscha Smolic, and Markus Gross. Images: DuctTake: Spatiotemporal video compositing. Computer Graphics Forum, 32(2pt1):51–61, May 2013. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Reichl:2016:MEI

[RWW16] F. Reichl, J. Weiss, and R. Westermann. Memory-efficient interactive online reconstruction from depth image streams. Computer Graphics Forum, 35(8):108–119, December 2016. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Remil:2017:AGD

[RXX^{+}17] O. Remil, Q. Xie, X. Xie, K. Xu, and J. Wang. Analyzing geometries: Data-driven sparse priors of 3D shapes. Computer Graphics Forum, 36(7):63–72, October 2017. CODEN CG-FODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Ren:2010:MSH

[RZS10] Cheng Ren, Liming Zhao, and Alla Safonova. Motion synthesis: Human motion synthesis with optimization-based graphs. Computer Graphics Forum, 29(2):545–554, May 2010. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Sedlmair:2015:EDD

[SA15] M. Sedlmair and M. Aupetit. Evaluation and design: Datadriven evaluation of visual quality measures. Computer Graphics Forum, 34(3):201–210, June 2015. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Simonetto:2011:GNI

[SAAB11] Paolo Simonetto, Daniel Archambault, David Auber, and Romain Bourqui. Graphs & networks: ImPrEd: an improved force-directed algorithm that prevents nodes from crossing edges. Computer Graphics Forum, 30(3):1071–1080, June 2011. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Shurkhovetskyy:2018:DAV

[SAAF18] G. Shurkhovetskyy, N. Andrienko, G. Andrienko, and G. Fuchs. Data abstraction for visualizing large time series. Computer Graphics Forum, 37(1):125–144, February 2018. CODEN CG-FODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Shi:2016:SSO

[SAD^{+}16] Zeyun Shi, Pierre Alliez, Mathieu Desbrun, Hujun Bao, and Jin Huang. Structures: Symmetry and orbit detection via Liealgebra voting. Computer Graphics Forum, 35(5):217–227, August 2016. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Shtof:2013:SGS

[SAG^{+}13] A. Shtof, A. Agathos, Y. Gingold, A. Shamir, and D. Cohen-Or. Sketching: Geosemantic snapping for sketch-based modeling. Computer Graphics Forum, 32(2pt2):245–253, May 2013. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Sarikaya:2014:MVV

[SAMG14] A. Sarikaya, D. Albers, J. Mitchell, and M. Gleicher. Molecular visualization: Visualizing validation of protein surface classifiers. Computer Graphics Forum, 33(3):171–180, June 2014. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Sacha:2017:DVA

[SAMS^{+}17] D. Sacha, F. Al-Masoudi, M. Stein, T. Schreck, D. A. Keim, G. Andrienko, and H. Janetzko. Dynamic visual abstraction of soccer movement. Computer Graphics Forum, 36(3):305– 315, June 2017. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Sitthi-Amorn:2010:ESM

[SARZL10] Pitchaya Sitthi-Amorn, Fabiano Romeiro, Todd Zickler, and Jason Lawrence. Editing of shadows and materials: Interactive editing of lighting and materials using a bivariate BRDF representation. Computer Graphics Forum, 29(4):1461–1468, June 2010. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Silva:2011:UVP

[SASF11] Cláudio T. Silva, Erik Anderson, Emanuele Santos, and Juliana Freire. Using VisTrails and provenance for teaching scientific visualization. Computer Graphics Forum, 30(1):75–84, March 2011. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Shier:2013:ARF

[SB13] John Shier and Paul Bourke. An algorithm for random fractal filling of space. Computer Graphics Forum, 32(8):89–97, December 2013. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Stoppel:2019:LFL

[SB19] S. Stoppel and S. Bruckner. LinesLab: A flexible low-cost approach for the generation of physical monochrome art. Computer Graphics Forum, 38(6):110–124, September 2019. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Schedl:2014:ICE

[SBB14] D. C. Schedl, C. Birklbauer, and O. Bimber. Images I: Coded exposure HDR light-field video recording. Computer Graphics Forum, 33(2):33–42, May 2014. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Shapira:2014:MCC

[SBC14] Nitzan Shapira and Mirela Ben-Chen. Maps: Cross-collection map inference by intrinsic alignment of shape spaces. Computer Graphics Forum, 33(5):281–290, August 2014. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Segall:2016:MIC

[SBC16] Aviv Segall and Mirela Ben-Chen. Mappings: Iterative closest conformal maps between planar domains. Computer Graphics Forum, 35(5):33–40, August 2016. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Song:2017:SRP

[SBC^{+}17] Jaewon Song, Roger Blanco i Ribera, Kyungmin Cho, Mi You, J. P. Lewis, Byungkuk Choi, and Junyong Noh. Sparse rig parameter optimization for character animation. Computer Graphics Forum, 36(2):85–94, May 2017. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Solomon:2011:MPS

[SBCBG11a] Justin Solomon, Mirela Ben-Chen, Adrian Butscher, and Leonidas Guibas. Mesh processing and surface reconstruction: Discovery of intrinsic primitives on triangle meshes. Computer Graphics Forum, 30(2):365–374, April 2011. CODEN CG-FODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Solomon:2011:SKP

[SBCBG11b] Justin Solomon, Mirela Ben-Chen, Adrian Butscher, and Leonidas Guibas. Section 6: As-killing-as-possible vector fields for planar deformation. Computer Graphics Forum, 30(5): 1543–1552, August 2011. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Scholz:2015:RTI

[SBD15a] Manuel Scholz, Jan Bender, and Carsten Dachsbacher. Realtime isosurface extraction with view-dependent level of detail and applications. Computer Graphics Forum, 34(1):103–115, February 2015. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Spritzer:2015:EGT

[SBD^{+}15b] Andre Suslik Spritzer, Jeremy Boy, Pierre Dragicevic, Jean-Daniel Fekete, and Carla Maria Dal Sasso Freitas. Evaluation of graphs: Towards a smooth design process for static communicative node-link diagrams. Computer Graphics Forum, 34 (3):461–470, June 2015. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Scandolo:2016:ERM

[SBE16a] Leonardo Scandolo, Pablo Bauszat, and Elmar Eisemann. Efficient rendering: Merged multiresolution hierarchies for shadow map compression. Computer Graphics Forum, 35(7):383–390, October 2016. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Scandolo:2016:SCM

[SBE16b] Leonardo Scandolo, Pablo Bauszat, and Elmar Eisemann. Shadows: Compressed multiresolution hierarchies for highquality precomputed shadows. Computer Graphics Forum, 35 (2):331–340, May 2016. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Salamon:2019:SST

[SBE19a] Nestor Z. Salamon, Markus Billeter, and Elmar Eisemann. ShutterApp: Spatio-temporal exposure control for videos. Computer Graphics Forum, 38(7):675–683, October 2019. CO-DEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Scandolo:2019:GGL

[SBE19b] Leonardo Scandolo, Pablo Bauszat, and Elmar Eisemann. Gradient-guided local disparity editing. Computer Graphics Forum, 38(1):394–404, February 2019. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Schollmeyer:2015:ROI

[SBF15] Andre Schollmeyer, Andrey Babanin, and Bernd Froehlich. Rendering: Order-independent transparency for programmable deferred shading pipelines. Computer Graphics Forum, 34(7): 67–76, October 2015. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Schertler:2017:CRC

[SBG17] Nico Schertler, Manfred Buchroithner, and Stefan Gumhold. Chamber recognition in cave data sets. Computer Graphics Forum, 36(2):375–386, May 2017. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Szymczak:2012:FVN

[SBL12] Andrzej Szymczak and Nicholas Brunhart-Lupo. Flow visualization: Nearly recurrent components in 3D piecewise constant vector fields. Computer Graphics Forum, 31(3pt3):1115–1124, June 2012. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Sher:2017:ESR

[SBLC17] Varshita Sher, Karen G. Bemis, Ilaria Liccardi, and Min Chen. An empirical study on the reliability of perceiving correlation indices using scatterplots. Computer Graphics Forum, 36(3): 61–72, June 2017. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Stava:2010:SMM

[ŠBM^{+}10] O. Št’ava, B. Beneš, R. Měch, D. G. Aliaga, and P. Krištof. Shape matching and modeling: Inverse procedural modeling by automatic generation of L-systems. Computer Graphics Forum, 29(2):665–674, May 2010. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Steiger:2014:GVV

[SBM^{+}14] Martin Steiger, Jürgen Bernard, Sebastian Mittelstädt, Hendrik Lücke-Tieke, Daniel Keim, Thorsten May, and Jörn Kohlhammer. Graph visualization: Visual analysis of timeseries similarities for anomaly detection in sensor networks. Computer Graphics Forum, 33(3):401–410, June 2014. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Sanchez-Banderas:2018:SRD

[SBO18] Rosa M. Sánchez-Banderas and Miguel A. Otaduy. Strain rate dissipation for elastic deformations. Computer Graphics Forum, 37(8):161–170, December 2018. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Solteszova:2017:OSF

[SBS^{+}17] Veronika Solteszova, Åsmund Birkeland, Sergej Stoppel, Ivan Viola, and Stefan Bruckner. Output-sensitive filtering of streaming volume data. Computer Graphics Forum, 36(1): 249–262, January 2017. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Seo:2010:CTM

[SC10] H. Seo and F. Cordier. Constrained texture mapping using image warping. Computer Graphics Forum, 29(1):160–174, March 2010. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Seo:2016:MIS

[SC16] H. Seo and F. Cordier. Matching and interpolation: Spatial matching of animated meshes. Computer Graphics Forum, 35 (7):21–32, October 2016. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Sheng:2011:PPG

[SCCN11] Yu Sheng, Barbara Cutler, Chao Chen, and Joshua Nasman. Perception: Perceptual global illumination cancellation in complex projection environments. Computer Graphics Forum, 30 (4):1261–1268, June 2011. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Sarikaya:2016:BDV

[SCD^{+}16] A. Sarikaya, M. Correli, J. M. Dinis, D. H. O’Connor, and M. Gleicher. Biological data visualization: Visualizing cooccurrence of events in populations of viral genome sequences. Computer Graphics Forum, 35(3):151–160, June 2016. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Sun:2010:MFG

[SCF10] Jian Sun, Xiaobai Chen, and Thomas A. Funkhouser. Matching: Fuzzy geodesics and consistent sparse correspondences for: Deformable shapes. Computer Graphics Forum, 29(5):1535– 1544, July 2010. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Schultz:2011:TBV

[Sch11] T. Schultz. Topology-based visualization: Topological features in 2D symmetric higher-order tensor fields. Computer Graphics Forum, 30(3):841–850, June 2011. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Schmidt:2013:SSP

[Sch13] R. Schmidt. Sketching: Stroke parameterization. Computer Graphics Forum, 32(2pt2):255–263, May 2013. CODEN CG-FODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Sellan:2019:SGP

[SCM^{+}19] Silvia Sellán, Herng Yi Cheng, Yuming Ma, Mitchell Dembowski, and Alec Jacobson. Solid geometry processing on deconstructed domains. Computer Graphics Forum, 38(1):564– 579, February 2019. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Sandim:2016:FSB

[SCN^{+}16] Marcos Sandim, Douglas Cedrim, Luis Gustavo Nonato, Paulo Pagliosa, and Afonso Paiva. Fluid simulation: Boundary detection in particle-based fluids. Computer Graphics Forum, 35 (2):215–224, May 2016. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Scopigno:2017:DFT

[SCP^{+}17] R. Scopigno, P. Cignoni, N. Pietroni, M. Callieri, and M. Dellepiane. Digital fabrication techniques for cultural heritage: A survey. Computer Graphics Forum, 36(1):6–21, January 2017. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Sun:2019:RRG

[SCQ^{+}19] Y. Sun, H. Chen, J. Qin, H. Li, M. Wei, and H. Zong. Reliable rolling-guided point normal filtering for surface texture removal. Computer Graphics Forum, 38(7):721–732, October 2019. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Serra:2018:EGF

[SCR^{+}18] J. Serra, O. Cetinaslan, S. Ravikumar, V. Orvalho, and D. Cosker. Easy generation of facial animation using motion graphs. Computer Graphics Forum, 37(1):97–111, February 2018. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Schlomer:2010:TGS

[SD10a] Thomas Schlömer and Oliver Deussen. Texture generation: Semi-stochastic tilings for example-based texture synthesis. Computer Graphics Forum, 29(4):1431–1439, June 2010. CO-DEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Sen:2010:SCC

[SD10b] Pradeep Sen and Soheil Darabi. Sparse computing: Compressive estimation for signal integration in rendering. Computer Graphics Forum, 29(4):1355–1363, June 2010. CODEN CG-FODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Santos:2018:DAC

[SDA^{+}18] B. Sousa Santos, J.-M. Dischler, V. Adzhiev, E. F. Anderson, A. Ferko, O. Fryazinov, M. Ilčík, I. Ilčíková, P. Slavik, V. Sundstedt, L. Svobodova, M. Wimmer, and J. Zara. Distinctive approaches to computer graphics education. Computer Graphics Forum, 37(1):403–412, February 2018. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Stuyck:2017:RTO

[SDHD17] Tuur Stuyck, Fang Da, Sunil Hadap, and Philip Dutré. Realtime oil painting on mobile hardware. Computer Graphics Forum, 36(8):69–79, December 2017. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Steinert:2011:GSC

[SDHL11] B. Steinert, H. Dammertz, J. Hanika, and H. P. A. Lensch. General spectral camera lens simulation. Computer Graphics Forum, 30(6):1643–1654, September 2011. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Sheng:2015:IVE

[SDK^{+}15] Kekai Sheng, Weiming Dong, Yan Kong, Xing Mei, Jilin Li, Chengjie Wang, Feiyue Huang, and Bao-Gang Hu. Image and video: Evaluating the quality of face alignment without ground truth. Computer Graphics Forum, 34(7):213–223, October 2015. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Sung:2018:LFS

[SDKG18] Minhyuk Sung, Anastasia Dubrovina, Vladimir G. Kim, and Leonidas Guibas. Learning fuzzy set representations of partial shapes on dual embedding spaces. Computer Graphics Forum, 37(5):71–81, August 2018. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Selgrad:2015:FML

[SDMS15] K. Selgrad, C. Dachsbacher, Q. Meyer, and M. Stamminger. Filtering multi-layer shadow maps for accurate soft shadows. Computer Graphics Forum, 34(1):205–215, February 2015. CO-DEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Shinya:2016:ERM

[SDS^{+}16] Mikio Shinya, Yoshinori Dobashi, Michio Shiraishi, Motonobu Kawashima, and Tomoyuki Nishita. Efficient rendering: Multiple scattering approximation in heterogeneous media by narrow beam distributions. Computer Graphics Forum, 35(7):373–382, October 2016. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Saket:2019:IMV

[SE19] Bahador Saket and Alex Endert. Investigating the manual view specification and visualization by demonstration paradigms for visualization construction. Computer Graphics Forum, 38(3): 663–674, June 2019. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Schulze:2014:FVS

[SEG^{+}14] M. Schulze, J. Martinez Esturo, T. Günther, C. Rössl, H.-P. Seidel, T. Weinkauf, and H. Theisel. Flow visualization: Sets of globally optimal stream surfaces for flow visualization. Computer Graphics Forum, 33(3):1–10, June 2014. CODEN CG-FODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Svetachov:2010:FLD

[SEI10] Pjotr Svetachov, Maarten H. Everts, and Tobias Isenberg. Fiber & lines: DTI in context: Illustrating brain fiber tracts in situ. Computer Graphics Forum, 29(3):1023–1032, June 2010. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Sanchez:2015:CFC

[SFFP15] Mathieu Sanchez, Oleg Fryazinov, Pierre-Alain Fayolle, and Alexander Pasko. Convolution filtering of continuous signed distance fields for polygonal meshes. Computer Graphics Forum, 34(6):277–288, September 2015. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Shen:2016:GHF

[SFL^{+}16] Zhongwei Shen, Xianzhong Fang, Xinguo Liu, Hujun Bao, and Jin Huang. Geometry: Harmonic functions for rotational symmetry vector fields. Computer Graphics Forum, 35(7):507–516, October 2016. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Su:2019:PFF

[SFL19] Jian-Ping Su, Xiao-Ming Fu, and Ligang Liu. Practical foldover-free volumetric mapping construction. Computer Graphics Forum, 38(7):287–297, October 2019. CODEN CG-FODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Summa:2018:FLW

[SFLP18] B. Summa, N. Faraj, C. Licorish, and V. Pascucci. Flexible live-wire: Image segmentation with floating anchors. Computer Graphics Forum, 37(2):321–328, May 2018. CODEN CG-FODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Shen:2013:SES

[SFY13] Li Shen, Jieqing Feng, and Baoguang Yang. Shadows: Exponential soft shadow mapping. Computer Graphics Forum, 32 (4):107–116, July 2013. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Scopigno:2010:E

[SG10] Roberto Scopigno and Eduard Gröller. Editorial. Computer Graphics Forum, 29(1):iii–iv, March 2010. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Solomon:2013:SCM

[SGB13] Justin Solomon, Leonidas Guibas, and Adrian Butscher. Shape collections and maps: Dirichlet energy for analysis and synthesis of soft maps. Computer Graphics Forum, 32(5):197–206, August 2013. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Schmaltz:2010:EH

[SGBW10] Christian Schmaltz, Pascal Gwosdek, Andrés Bruhn, and Joachim Weickert. Electrostatic halftoning. Computer Graphics Forum, 29(8):2313–2327, December 2010. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Stengel:2016:FRA

[SGEM16] Michael Stengel, Steve Grogorick, Martin Eisemann, and Marcus Magnor. Faster rendering: Adaptive image-space sampling for gaze-contingent real-time rendering. Computer Graphics Forum, 35(4):129–139, July 2016. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Staib:2015:VRV

[SGG15] Joachim Staib, Sebastian Grottel, and Stefan Gumhold. Volume rendering: Visualization of particle-based data with transparency and ambient occlusion. Computer Graphics Forum, 34 (3):151–160, June 2015. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Staib:2016:HDD

[SGG16] J. Staib, S. Grottel, and S. Gumhold. High-dimensional data: Enhancing scatterplots with multi-dimensional focal blur. Computer Graphics Forum, 35(3):11–20, June 2016. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Schott:2011:VVD

[SGM^{+}11] Mathias Schott, A. V. Pascal Grosset, Tobias Martin, Vincent Pegoraro, Sean T. Smith, and Charles D. Hansen. Volume visualization: Depth of field effects for interactive direct volume rendering. Computer Graphics Forum, 30(3):941–950, June 2011. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Serrano:2017:CSC

[SGMG17] Ana Serrano, Elena Garces, Belen Masia, and Diego Gutierrez. Convolutional sparse coding for capturing high-speed video content. Computer Graphics Forum, 36(8):380–389, December 2017. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Schulze:2012:VPS

[SGRT12] Maik Schulze, Tobias Germer, Christian Rössl, and Holger Theisel. Visualization and perception: Stream surface parametrization by flow-orthogonal front lines. Computer Graphics Forum, 31(5):1725–1734, August 2012. CODEN CG-FODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Sipiran:2014:SCA

[SGS14] Ivan Sipiran, Robert Gregor, and Tobias Schreck. Shapes and cryptography: Approximate symmetry detection in partial 3D meshes. Computer Graphics Forum, 33(7):131–140, October 2014. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Sarikaya:2018:DFS

[SGS18] A. Sarikaya, M. Gleicher, and D. A. Szafir. Design factors for summary visualization in visual analytics. Computer Graphics Forum, 37(3):145–156, June 2018. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Summa:2015:CTP

[SGSP15] B. Summa, A. A. Gooch, G. Scorzelli, and V. Pascucci. Colors and textures: Paint and click: Unified interactions for image boundaries. Computer Graphics Forum, 34(2):385–393, May 2015. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Schmaltz:2012:MCA

[SGW12] C. Schmaltz, P. Gwosdek, and J. Weickert. Multi-class anisotropic electrostatic halftoning. Computer Graphics Forum, 31(6):1924–1935, September 2012. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Shen:2011:IGP

[SGYF11] Li Shen, Gaël Guennebaud, Baoguang Yang, and Jieqing Feng. Intelligent graphics: Predicted virtual soft shadow maps with high quality filtering. Computer Graphics Forum, 30(2):493– 502, April 2011. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Satyanarayan:2014:DSA

[SH14a] Arvind Satyanarayan and Jeffrey Heer. Design study: Authoring narrative visualizations with ellipsis. Computer Graphics Forum, 33(3):361–370, June 2014. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Satyanarayan:2014:DSL

[SH14b] Arvind Satyanarayan and Jeffrey Heer. Design study: Lyra: an interactive visualization design environment. Computer Graphics Forum, 33(3):351–360, June 2014. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Shen:2015:SVG

[SHCB15] Zhongwei Shen, Jin Huang, Wei Chen, and Hujun Bao. Simulation and visualization: Geometrically exact simulation of inextensible ribbon. Computer Graphics Forum, 34(7):145–154, October 2015. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Simon:2015:GIR

[SHD15] Florian Simon, Johannes Hanika, and Carsten Dachsbacher. Global illumination: Rich-VPLs for improving the versatility of many-light methods. Computer Graphics Forum, 34(2):575– 584, May 2015. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Schrade:2016:LTS

[SHD16] Emanuel Schrade, Johannes Hanika, and Carsten Dachsbacher. Looking through surfaces: Sparse high-degree polynomials for wide-angle lenses. Computer Graphics Forum, 35(4):89–97, July 2016. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Sheehy:2012:CGN

[She12] Donald R. Sheehy. Computational geometry: New bounds on the size of optimal meshes. Computer Graphics Forum, 31(5): 1627–1635, August 2012. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Schneider:2013:DCB

[SHF13] T. Schneider, K. Hormann, and M. S. Floater. Deformations and coordinates: Bijective composite mean value mappings. Computer Graphics Forum, 32(5):137–146, August 2013. CO-DEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Serrano:2016:HIC

[SHG^{+}16] Ana Serrano, Felix Heide, Diego Gutierrez, Gordon Wetzstein, and Belen Masia. HDR imaging: Convolutional sparse coding for high dynamic range imaging. Computer Graphics Forum, 35(2):153–163, May 2016. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Shen:2016:IEP

[SHJ^{+}16] Xiaoyong Shen, Aaron Hertzmann, Jiaya Jia, Sylvain Paris, Brian Price, Eli Shechtman, and Ian Sachs. Image editing & processing: Automatic portrait segmentation for image stylization. Computer Graphics Forum, 35(2):93–102, May 2016. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Skau:2015:EDE

[SHK15] Drew Skau, Lane Harrison, and Robert Kosara. Evaluation and design: an evaluation of the impact of visual embellishments in bar charts. Computer Graphics Forum, 34(3):221–230, June 2015. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Sharf:2014:MTI

[SHL^{+}14] Andrei Sharf, Hui Huang, Cheng Liang, Jiapei Zhang, Baoquan Chen, and Minglun Gong. Mobility-trees for indoor scenes manipulation. Computer Graphics Forum, 33(1):2–14, February 2014. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Simons:2019:AVA

[SHP^{+}19] G. Simons, S. Herholz, V. Petitjean, T. Rapp, M. Ament, H. Lensch, C. Dachsbacher, M. Eisemann, and E. Eisemann. Applying visual analytics to physically based rendering. Computer Graphics Forum, 38(1):197–208, February 2019. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Shah:2018:DLB

[SHQL18] C. Shah, D. Hyde, H. Qu, and P. Levis. Distributing and load balancing sparse fluid simulations. Computer Graphics Forum, 37(8):35–46, December 2018. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Sofer:2013:ASI

[SHS13] Yerry Sofer, Tal Hassner, and Andrei Sharf. Animation (session 1): Interactive learning for point-cloud motion segmentation. Computer Graphics Forum, 32(7):51–60, October 2013. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Sung:2017:VVE

[SHS^{+}17] Ching-Ying Sung, Xun-Yi Huang, Yicong Shen, Fu-Yin Cherng, Wen-Chieh Lin, and Hao-Chuan Wang. Video and visualization: Exploring online learners’ interactive dynamics by visually analyzing their time-anchored comments. Computer Graphics Forum, 36(7):145–155, October 2017. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Sbert:2016:RRV

[SHSK16] M. Sbert, V. Havran, and L. Szirmay-Kalos. Realistic rendering: Variance analysis of multi-sample and one-sample multiple importance sampling. Computer Graphics Forum, 35(7): 451–460, October 2016. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Shen:2018:DDC

[SHW^{+}18] Yijun Shen, Joseph Henry, He Wang, Edmond S. L. Ho, Taku Komura, and Hubert P. H. Shum. Data-driven crowd motion control with multi-touch gestures. Computer Graphics Forum, 37(6):382–394, September 2018. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Simon:2017:MCA

[SHZD17] Florian Simon, Johannes Hanika, Tobias Zirr, and Carsten Dachsbacher. Monte Carlo after coffee: Line integration for rendering heterogeneous emissive volumes. Computer Graphics Forum, 36(4):101–110, July 2017. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Selino:2013:LSE

[SJ13a] A. Selino and M. D. Jones. Large and small eddies matter: Animating trees in wind using coarse fluid simulation and synthetic turbulence. Computer Graphics Forum, 32(1):75–84, February 2013. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Spencer:2013:IPP

[SJ13b] B. Spencer and M. W. Jones. Illumination: Photon parameterisation for robust relaxation constraints. Computer Graphics Forum, 32(2pt1):83–92, May 2013. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Spencer:2013:RCI

[SJ13c] Ben Spencer and Mark W. Jones. Reports: 2013 cover image: Prism. Computer Graphics Forum, 32(1):216–217, February 2013. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Salesin:2019:CPL

[SJ19] Katherine Salesin and Wojciech Jarosz. Combining point and line samples for direct illumination. Computer Graphics Forum, 38(4):159–169, July 2019. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Stoffel:2017:IAR

[SJB^{+}17] Florian Stoffel, Wolfgang Jentner, Michael Behrisch, Johannes Fuchs, and Daniel Keim. Interactive ambiguity resolution of named entities in fictional literature. Computer Graphics Forum, 36(3):189–200, June 2017. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Sun:2011:AMR

[SJF11] Chuan Sun, Imran Junejo, and Hassan Foroosh. Animation: Motion retrieval using low-rank subspace decomposition of motion volume. Computer Graphics Forum, 30(7):1953–1962, September 2011. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Spencer:2015:VTU

[SJL15] B. Spencer, M. W. Jones, and I. S. Lim. A visualization tool used to develop new photon mapping techniques. Computer Graphics Forum, 34(1):127–140, February 2015. CODEN CG-FODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Sacht:2013:MCV

[SJP^{+}13] Leonardo Sacht, Alec Jacobson, Daniele Panozzo, Christian Schüller, and Olga Sorkine-Hornung. Meshing: Consistent volumetric discretizations inside self-intersecting surfaces. Computer Graphics Forum, 32(5):147–156, August 2013. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Sykora:2019:SFE

[SJT^{+}19] D. Sýkora, O. Jamriška, O. Texler, J. Fišer, M. Lukáč, J. Lu, and E. Shechtman. StyleBlit: Fast example-based stylization with local guidance. Computer Graphics Forum, 38(2):83–91, May 2019. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Sunkel:2011:SAL

[SJW^{+}11] M. Sunkel, S. Jansen, M. Wand, E. Eisemann, and H.-P. Seidel. Shape analysis: Learning line features in 3D geometry. Computer Graphics Forum, 30(2):267–276, April 2011. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Sunkel:2013:SCC

[SJWS13] M. Sunkel, S. Jansen, M. Wand, and H.-P. Seidel. Shape construction: a correlated parts model for object detection in large 3D scans. Computer Graphics Forum, 32(2pt2):205–214, May 2013. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Schultz:2010:VTM

[SK10] T. Schultz and G. Kindlmann. Vectors & tensors: a maximum enhancing higher-order tensor glyph. Computer Graphics Forum, 29(3):1143–1152, June 2010. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Seltzer:2016:CGG

[SK16a] Nicholas Seltzer and Gordon Kindlmann. Charts and glyphs: Glyphs for asymmetric second-order 2D tensors. Computer Graphics Forum, 35(3):141–150, June 2016. CODEN CG-FODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Skau:2016:CGA

[SK16b] Drew Skau and Robert Kosara. Charts and glyphs: Arcs, angles, or areas: Individual data encodings in pie and donut charts. Computer Graphics Forum, 35(3):121–130, June 2016. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Sibbing:2017:BLD

[SK17] Dominik Sibbing and Leif Kobbelt. Building a large database of facial movements for deformation model-based 3D face tracking. Computer Graphics Forum, 36(8):285–301, December 2017. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Szirmay-Kalos:2017:ULT

[SKGM^{+}17] László Szirmay-Kalos, Iliyan Georgiev, Milán Magdics, Balázs Molnár, and Dávid Légrády. Unbiased light transport estimators for inhomogeneous participating media. Computer Graphics Forum, 36(2):9–19, May 2017. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Shiny:2010:FSH

[SKK10] Seung-Ho Shiny, Hyeong Ryeol Kamz, and Chang-Hun Kimx. Fluids and smoke: Hybrid simulation of miscible mixing with viscous fingering. Computer Graphics Forum, 29(2):675–683, May 2010. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Steinberger:2014:PMP

[SKK^{+}14a] Markus Steinberger, Michael Kenzel, Bernhard Kainz, Jörg Müller, Wonka Peter, and Dieter Schmalstieg. Procedural Modeling I: Parallel generation of architecture on the GPU. Computer Graphics Forum, 33(2):73–82, May 2014. CODEN CG-FODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Steinberger:2014:PMF

[SKK^{+}14b] Markus Steinberger, Michael Kenzel, Bernhard Kainz, Peter Wonka, and Dieter Schmalstieg. Procedural Modeling I: Onthe-fly generation and rendering of infinite cities on the GPU. Computer Graphics Forum, 33(2):105–114, May 2014. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Szirmay-Kalos:2018:MSI

[SKMS18] László Szirmay-Kalos, Milán Magdics, and Mateu Sbert. Multiple scattering in inhomogeneous participating media using Rao–Blackwellization and control variates. Computer Graphics Forum, 37(2):63–74, May 2018. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Schafer:2015:LPD

[SKNS15] H. Schäfer, B. Keinert, M. Nießner, and M. Stamminger. Local painting and deformation of meshes on the GPU. Computer Graphics Forum, 34(1):26–35, February 2015. CODEN CG-FODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Schuller:2013:DCL

[SKPSH13] Christian Schüller, Ladislav Kavan, Daniele Panozzo, and Olga Sorkine-Hornung. Deformations and coordinates: Locally injective mappings. Computer Graphics Forum, 32(5):125–135, August 2013. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Scharnowski:2014:MVC

[SKR^{+}14] K. Scharnowski, M. Krone, G. Reina, T. Kulschewski, J. Pleiss, and T. Ertl. Molecular visualization: Comparative visualization of molecular surfaces using deformable models. Computer Graphics Forum, 33(3):191–200, June 2014. CODEN CG-FODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Sons:2014:RSJ

[SKSS14] Kristian Sons, Felix Klein, Jan Sutter, and Philipp Slusallek. Rendering: shade.js: Adaptive material descriptions. Computer Graphics Forum, 33(7):51–60, October 2014. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Szirmay-Kalos:2011:FPS

[SKTM11] László Szirmay-Kalos, Balázs Tóth, and Milán Magdics. Free path sampling in high resolution inhomogeneous participating media. Computer Graphics Forum, 30(1):85–97, March 2011. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Son:2013:IPS

[SKWL13] Minjung Son, Byungmoon Kim, Gregg Wilensky, and Seungyong Lee. Image processing (session 5): Still-frame simulation for fire effects of images. Computer Graphics Forum, 32(7): 295–304, October 2013. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Schroder:2011:RAV

[SKZ11] K. Schroder, R. Klein, and A. Zinke. Realistic appearance and volumes: a volumetric approach to predictive rendering of fabrics. Computer Graphics Forum, 30(4):1277–1286, June 2011. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Schroder:2013:NLI

[SKZ13] K. Schröder, R. Klein, and A. Zinke. Non-local image reconstruction for efficient computation of synthetic bidirectional texture functions. Computer Graphics Forum, 32(8):61–71, December 2013. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Schroeder:2011:AAE

[SKZF11] Craig Schroeder, Nipun Kwatra, Wen Zheng, and Ron Fedkiw. Animation: Asynchronous evolution for fully-implicit and semiimplicit time integration. Computer Graphics Forum, 30(7): 1983–1992, September 2011. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Sivaks:2011:NMT

[SL11] Eliyahu Sivaks and Dani Lischinski. On neighbourhood matching for texture-by-numbers. Computer Graphics Forum, 30(1): 127–138, March 2011. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Salinas:2015:SAM

[SLA15] D. Salinas, F. Lafarge, and P. Alliez. Structure-aware mesh decimation. Computer Graphics Forum, 34(6):211–227, September 2015. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Saalfeld:2018:GQE

[SLB^{+}18] P. Saalfeld, M. Luz, P. Berg, B. Preim, and S. Saalfeld. Guidelines for quantitative evaluation of medical visualizations on the example of 3D aneurysm surface comparisons. Computer Graphics Forum, 37(1):226–238, February 2018. CODEN CG-FODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Su:2019:IPP

[SLC^{+}19] Rung-De Su, Zhe-Yo Liao, Li-Chi Chen, Ai-Ling Tung, and Yu-Shuen Wang. Imitating popular photos to select views for an indoor scene. Computer Graphics Forum, 38(7):141–148, October 2019. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Son:2019:NPI

[SLCL19] Hyeongseok Son, Gunhee Lee, Sunghyun Cho, and Seungyong Lee. Naturalness-preserving image tone enhancement using generative adversarial networks. Computer Graphics Forum, 38(7):277–285, October 2019. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Sendik:2019:WFM

[SLCO19] O. Sendik, D. Lischinski, and D. Cohen-Or. What’s in a face? Metric learning for face characterization. Computer Graphics Forum, 38(2):405–416, May 2019. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Simoes:2017:GDA

[SLD^{+}17] Tiago Simões, Daniel Lopes, Sérgio Dias, Francisco Fernandes, João Pereira, Joaquim Jorge, Chandrajit Bajaj, and Abel Gomes. Geometric detection algorithms for cavities on protein surfaces in molecular graphics: a survey. Computer Graphics Forum, 36(8):643–683, December 2017. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Salamon:2017:ADS

[SLE17] Nestor Z. Salamon, Marcel Lancelle, and Elmar Eisemann. Art, design, and sketching: Computational light painting using a virtual exposure. Computer Graphics Forum, 36(2):1–8, May 2017. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Scandolo:2018:QBF

[SLE18] Leonardo Scandolo, Sungkil Lee, and Elmar Eisemann. Quadbased Fourier transform for efficient diffraction synthesis. Computer Graphics Forum, 37(4):167–176, July 2018. CODEN CG-FODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Soni:2018:PGP

[SLH^{+}18] Utkarsh Soni, Yafeng Lu, Brett Hansen, Helen C. Purchase, Stephen Kobourov, and Ross Maciejewski. The perception of graph properties in graph layouts. Computer Graphics Forum, 37(3):169–181, June 2018. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Shen:2012:GPS

[SLHC12] Liang-Tsen Shen, Sheng-Jie Luo, Chun-Kai Huang, and Bing-Yu Chen. Geometry processing: SD models: Super-deformed character models. Computer Graphics Forum, 31(7pt1):2067– 2075, September 2012. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Son:2014:IIA

[SLKL14] Minjung Son, Yunjin Lee, Henry Kang, and Seungyong Lee. Images III: Art-photographic detail enhancement. Computer Graphics Forum, 33(2):391–400, May 2014. CODEN CG-FODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Stitz:2016:SHE

[SLSG16] H. Stitz, S. Luger, M. Streit, and N. Gehlenborg. Story, history, and evolution: AVOCADO: Visualization of workflow-derived data provenance for reproducible biomedical research. Computer Graphics Forum, 35(3):481–490, June 2016. CODEN CG-FODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Sevilla-Lara:2015:EVS

[SLWSS15] L. Sevilla-Lara, J. Wulff, K. Sunkavalli, and E. Shechtman. Exposure and video: Smooth loops from unconstrained video. Computer Graphics Forum, 34(4):99–107, July 2015. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Su:2019:DCF

[SLZ^{+}19] Kh Su, Cc Li, Ym Zhou, X Xu, and Xf Gu. Discrete Calabi flow: A unified conformal parameterization method. Computer Graphics Forum, 38(7):707–720, October 2019. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Sajadi:2010:HSM

[SM10] Behzad Sajadi and Aditi Majumder. Hardware: Scalable multiview registration for multi-projector displays on vertically extruded surfaces. Computer Graphics Forum, 29(3):1063–1072, June 2010. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Sajadi:2011:HLS

[SM11] B. Sajadi and A. Majumder. Hierarchies and large-scale visualization: Automatic registration of multi-projector domes using a single uncalibrated camera. Computer Graphics Forum, 30 (3):1161–1170, June 2011. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Sakurai:2014:MNP

[SM14a] K. Sakurai and K. Miyata. Modelling of non-periodic aggregates having a pile structure. Computer Graphics Forum, 33 (1):190–198, February 2014. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Sidorov:2014:MLR

[SM14b] K. A. Sidorov and A. D. Marshall. Matching: Learnt realtime meshless simulation. Computer Graphics Forum, 33(5): 147–156, August 2014. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Swoboda:2017:VQI

[SMB^{+}17] N. Swoboda, J. Moosburner, S. Bruckner, J. Y. Yu, B. J. Dickson, and K. Bühler. Visualization and quantification for interactive analysis of neural connectivity in Drosophila. Computer Graphics Forum, 36(1):160–171, January 2017. CODEN CG-FODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Sussmuth:2010:SRB

[SMG10] J. Süßmuth, Q. Meyer, and G. Greiner. Surface reconstruction based on hierarchical floating radial basis functions. Computer Graphics Forum, 29(6):1854–1864, September 2010. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Saad:2010:MVP

[SMH10] Ahmed Saad, Torsten Möller, and Ghassan Hamarneh. Medical visualization: ProbExplorer: Uncertainty-guided exploration and editing of probabilistic medical image segmentation. Computer Graphics Forum, 29(3):1113–1122, June 2010. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Singh:2017:MCB

[SMJ17] Gurprit Singh, Bailey Miller, and Wojciech Jarosz. Monte Carlo before coffee: Variance and convergence analysis of Monte Carlo line and segment sampling. Computer Graphics Forum, 36(4): 79–89, July 2017. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Samii:2015:DDA

[SML15] A. Samii, R. Měch, and Z. Lin. Data-driven automatic cropping using semantic composition search. Computer Graphics Forum, 34(1):141–151, February 2015. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Sigovan:2013:PVL

[SMM13] Carmen Sigovan, Chris W. Muelder, and Kwan-Liu Ma. Performance: Visualizing large-scale parallel communication traces using a particle animation technique. Computer Graphics Forum, 32(3pt2):141–150, June 2013. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Suter:2013:PTT

[SMP13] S. K. Suter, M. Makhynia, and R. Pajarola. Performance: TAMRESH – tensor approximation multiresolution hierarchy for interactive volume visualization. Computer Graphics Forum, 32(3pt2):151–160, June 2013. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Steiner:2017:ISA

[SMS^{+}17] B. Steiner, E. Mousavian, F. Mehdizadeh Saradj, M. Wimmer, and P. Musialski. Integrated structural–architectural design for interactive planning. Computer Graphics Forum, 36(8):80–94, December 2017. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Soler:2018:CRP

[SMSH18] Carlota Soler, Tobias Martin, and Olga Sorkine-Hornung. Cosserat rods with projective dynamics. Computer Graphics Forum, 37(8):137–147, December 2018. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Shao:2017:IRL

[SMSL17] L. Shao, A. Mahajan, T. Schreck, and D. J. Lehmann. Interactive regression lens for exploring scatter plots. Computer Graphics Forum, 36(3):157–166, June 2017. CODEN CG-FODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Scheepens:2015:TRV

[SMvdWvW15] Roeland Scheepens, Steffen Michels, Huub van de Wetering, and Jarke J. van Wijk. Traffic: Rationale visualization for safety and security. Computer Graphics Forum, 34(3):191–200, June 2015. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Shivashankar:2012:TPC

[SN12] Nithin Shivashankar and Vijay Natarajan. Topology: Parallel computation of 3D Morse–Smale complexes. Computer Graphics Forum, 31(3pt1):965–974, June 2012. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Sainlot:2017:MRV

[SNA17] M. Sainlot, V. Nivoliers, and D. Attali. Meshing: Restricting Voronoi diagrams to meshes using corner validation. Computer Graphics Forum, 36(5):81–91, August 2017. CODEN CG-FODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Solomon:2012:CSM

[SNB^{+}12] Justin Solomon, Andy Nguyen, Adrian Butscher, Mirela Ben-Chen, and Leonidas Guibas. Correspondence: Soft maps between surfaces. Computer Graphics Forum, 31(5):1617–1626, August 2012. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Subr:2014:SEA

[SNJ^{+}14] Kartic Subr, Derek Nowrouzezahrai, Wojciech Jarosz, Jan Kautz, and Kenny Mitchell. Sampling: Error analysis of estimators that use combinations of stochastic sampling strategies for direct illumination. Computer Graphics Forum, 33(4):93– 102, July 2014. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Scherzer:2012:GIP

[SNRS12] Daniel Scherzer, Chuong H. Nguyen, Tobias Ritschel, and Hans-Peter Seidel. Global illumination: Pre-convolved radiance caching. Computer Graphics Forum, 31(4):1391–1397, June 2012. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Stefanoski:2010:SFE

[SO10] Nikolče Stefanoski and Jörn Ostermann. SPC: Fast and efficient scalable predictive coding of animated meshes. Computer Graphics Forum, 29(1):101–116, March 2010. CODEN CG-FODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Shapira:2012:ICP

[SO12] L. Shapira and B. Oicherman. Image and color processing: Black is green: Adaptive color transformation for reduced ink usage. Computer Graphics Forum, 31(2pt1):365–372, May 2012. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Singh:2019:ASC

[SÖA^{+}19] Gurprit Singh, Cengiz Öztireli, Abdalla G. M. Ahmed, David Coeurjolly, Kartic Subr, Oliver Deussen, Victor Ostromoukhov, Ravi Ramamoorthi, and Wojciech Jarosz. Analysis of sample correlations for Monte Carlo rendering. Computer Graphics Forum, 38(2):473–491, May 2019. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Santesteban:2019:LBA

[SOC19] Igor Santesteban, Miguel A. Otaduy, and Dan Casas. Learningbased animation of clothing for virtual try-on. Computer Graphics Forum, 38(2):355–366, May 2019. CODEN CG-FODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Schwartzburg:2013:RFA

[SP13] Yuliy Schwartzburg and Mark Pauly. Reality: Fabricationaware design with intersecting planar pieces. Computer Graphics Forum, 32(2pt3):317–326, May 2013. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Schikora:2017:VAC

[SPB^{+}17] Christoph M. Schikora, Markus Plack, Rainer Bornemann, Peter Haring Bolívar, and Andreas Kolb. Visual analysis of confocal Raman spectroscopy data using cascaded transfer function design. Computer Graphics Forum, 36(3):239–249, June 2017. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Sabbadin:2019:HDR

[SPB^{+}19] Manuele Sabbadin, Gianpaolo Palma, Francesco Banterle, Tamy Boubekeur, and Paolo Cignoni. High dynamic range point clouds for real-time relighting. Computer Graphics Forum, 38(7):513–525, October 2019. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Solteszova:2010:VRM

[ŠPBV10] Veronika Šoltészová, Daniel Patel, Stefan Bruckner, and Ivan Viola. Volume rendering: a multidirectional occlusion shading model for direct volume rendering. Computer Graphics Forum, 29(3):883–891, June 2010. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Shahrian:2014:IIT

[SPCR14] E. Shahrian, B. Price, S. Cohen, and D. Rajan. Images III: Temporally coherent and spatially accurate video matting. Computer Graphics Forum, 33(2):381–390, May 2014. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Simari:2014:SFS

[SPD14] Patricio Simari, Giulia Picciau, and Leila De Floriani. Surfaces: Fast and scalable mesh superfacets. Computer Graphics Forum, 33(7):181–190, October 2014. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Shimizu:2019:ESL

[SPF^{+}19] Evan Shimizu, Sylvain Paris, Matt Fisher, Ersin Yumer, and Kayvon Fatahalian. Exploratory stage lighting design using visual objectives. Computer Graphics Forum, 38(2):417–429, May 2019. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Shin:2011:HLS

[SPH11] HyunJu Shin, GwangHyun Park, and JungHyun Han. Hierarchies and large-scale visualization: Tablorer — an interactive tree visualization system for tablet PCs. Computer Graphics Forum, 30(3):1131–1140, June 2011. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Sibbing:2010:IVI

[SPK10] Dominik Sibbing, Darko Pavić, and Leif Kobbelt. Image and video II: Image synthesis for branching structures. Computer Graphics Forum, 29(7):2135–2144, September 2010. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Stava:2014:IPM

[SPK^{+}14] O. Stava, S. Pirk, J. Kratt, B. Chen, R. Měch, O. Deussen, and B. Benes. Inverse procedural modelling of trees. Computer Graphics Forum, 33(6):118–131, September 2014. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Schmidt:2016:SAA

[SPN^{+}16] Thorsten-Walther Schmidt, Fabio Pellacini, Derek Nowrouzezahrai, Wojciech Jarosz, and Carsten Dachsbacher. State of the art in artistic editing of appearance, lighting and material. Computer Graphics Forum, 35(1):216–233, February 2016. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Subr:2013:IAB

[SPSK13] Kartic Subr, Sylvain Paris, Cyril Soler, and Jan Kautz. Images: Accurate binary image selection from inaccurate user input. Computer Graphics Forum, 32(2pt1):41–50, May 2013. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Stojakovic:2014:VCP

[SPT14] V. Stojaković, S. Popov, and B. Tepavčević. Visualization of the centre of projection geometrical locus in a single image. Computer Graphics Forum, 33(1):52–63, February 2014. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Soares:2010:DHI

[SPV^{+}10] L. P. Soares, F. Pires, R. Varela, R. Bastos, N. Carvalho, F. Gaspar, and M. S. Dias. Designing a highly immersive interactive environment: The virtual mine. Computer Graphics Forum, 29(6):1756–1769, September 2010. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Sinha:2014:MSK

[SR14] A. Sinha and K. Ramani. Multi-scale kernels using random walks. Computer Graphics Forum, 33(1):164–177, February 2014. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Serpa:2019:FUT

[SR19] Ygor Rebouças Serpa and Maria Andréia Formico Rodrigues. Flexible use of temporal and spatial reasoning for fast and scalable CPU Broad-Phase collision detection using KD-trees.

Computer Graphics Forum, 38(1):260–273, February 2019. CO-DEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Shaheen:2016:SSA

[SRG16] Sara Shaheen, Alyn Rockwood, and Bernard Ghanem. SAR: Stroke authorship recognition. Computer Graphics Forum, 35 (6):73–86, September 2016. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Stoiber:2019:NDN

[SRG^{+}19] C. Stoiber, A. Rind, F. Grassinger, R. Gutounig, E. Goldgruber, M. Sedlmair, Š. Emrich, and W. Aigner. netflower: Dynamic network visualization for data journalists. Computer Graphics Forum, 38(3):699–711, June 2019. CODEN CG-FODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Scherbaum:2011:IGC

[SRH^{+}11] Kristina Scherbaum, Tobias Ritschel, Matthias Hullin, Thorsten Thormählen, Volker Blanz, and Hans-Peter Seidel. Intelligent graphics: Computer-suggested facial makeup. Computer Graphics Forum, 30(2):485–492, April 2011. CODEN CG-FODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Schreck:2017:IPT

[SRH17] Camille Schreck, Damien Rohmer, and Stefanie Hahmann. Interactive paper tearing. Computer Graphics Forum, 36(2):95– 106, May 2017. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Schwartz:2013:RSL

[SRK13] Christopher Schwartz, Roland Ruiters, and Reinhard Klein. Rendering (session 2): Level-of-detail streaming and rendering using bidirectional sparse virtual texture functions. Computer Graphics Forum, 32(7):345–354, October 2013. CODEN CG-FODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Schlachter:2019:SAR

[SRM^{+}19] M. Schlachter, R. G. Raidou, L. P. Muren, B. Preim, P. M. Putora, and K. Bühler. State-of-the-art report: Visual computing in radiation therapy planning. Computer Graphics Forum, 38(3):753–779, June 2019. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Schneider:2010:VTT

[SRWS10] Dominic Schneider, Wieland Reich, Alexander Wiebel, and Gerik Scheuermann. Vectors & tensors: Topology aware stream surfaces. Computer Graphics Forum, 29(3):1153–1161, June 2010. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Smith:2015:SMS

[SS15a] J. Smith and S. Schaefer. Splines & meshes: Selective degree elevation for multi-sided Bézier patches. Computer Graphics Forum, 34(2):609–615, May 2015. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Soundararajan:2015:VAC

[SS15b] K. P. Soundararajan and T. Schultz. Volume analysis and classification: Learning probabilistic transfer functions: a comparative study of classifiers. Computer Graphics Forum, 34(3):111– 120, June 2015. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Sadana:2016:CVI

[SS16] R. Sadana and J. Stasko. Coordinated views and interaction design: Designing multiple coordinated visualizations for tablets. Computer Graphics Forum, 35(3):261–270, June 2016. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Sin:2013:VNL

[SSB13] F. S. Sin, D. Schroeder, and J. Barbič. Vega: Non-linear FEM deformable object simulator. Computer Graphics Forum, 32 (1):36–48, February 2013. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Shen:2014:CIT

[SSB^{+}14] Fangyang Shen, Kalyan Sunkavalli, Nicolas Bonneel, Szymon Rusinkiewicz, Hanspeter Pfister, and Xin Tong. Color and imaging: Time-lapse photometric stereo and applications. Computer Graphics Forum, 33(7):359–367, October 2014. CO-DEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Stoffel:2012:TDD

[SSDK12] A. Stoffel, H. Strobelt, O. Deussen, and D. A. Keim. Text and documents: Document thumbnails with variable text scaling. Computer Graphics Forum, 31(3pt3):1165–1173, June 2012. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Silva:2014:WDT

[SSE^{+}14] Luis F. Silva, Luiz F. Scheidegger, Tiago Etiene, João L. D. Comba, Luis G. Nonato, and Cláudio T. Silva. A weighted Delaunay triangulation framework for merging triangulations in a connectivity oblivious fashion. Computer Graphics Forum, 33(6):18–30, September 2014. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Schertler:2017:TGO

[SSG17] Nico Schertler, Bogdan Savchynskyy, and Stefan Gumhold. Towards globally optimal normal orientations for large point clouds. Computer Graphics Forum, 36(1):197–208, January 2017. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Sun:2017:BIA

[SSGM17] Tiancheng Sun, Ana Serrano, Diego Gutierrez, and Belen Masia. BRDFs and illumination: Attribute-preserving gamut mapping of measured BRDFs. Computer Graphics Forum, 36 (4):47–54, July 2017. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Sykora:2010:NPR

[SSJ^{+}10] D. Sýkora, D. Sedlacek, S. Jinchao, J. Dingliana, and S. Collins. Non-photorealistic rendering: Adding depth to cartoons using sparse depth (in)equalities. Computer Graphics Forum, 29 (2):615–623, May 2010. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Saket:2016:NGC

[SSK16] Bahador Saket, Carlos Scheidegger, and Stephen Kobourov. Networks and graphs 1: Comparing node-link and node-link-group visualizations from an enjoyment perspective. Computer Graphics Forum, 35(3):41–50, June 2016. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Saket:2015:EGM

[SSKB15] Bahador Saket, Carlos Scheidegger, Stephen G. Kobourov, and Katy Börner. Evaluation of graphs: Map-based visualizations increase recall accuracy of data. Computer Graphics Forum, 34 (3):441–450, June 2015. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Silvennoinen:2014:OSU

[SSLL14] Ari Silvennoinen, Hannu Saransaari, Samuli Laine, and Jaakko Lehtinen. Occluder simplification using planar sections. Computer Graphics Forum, 33(1):235–245, February 2014. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Silva:2012:EDP

[SSM12] Samuel Silva, Beatriz Sousa Santos, and Joaquim Madeira. Exploring different parameters to assess left ventricle global and regional functional analysis from coronary CT angiography. Computer Graphics Forum, 31(1):146–159, February 2012. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Soler:2018:VPM

[SSN18] Cyril Soler, Kartic Subr, and Derek Nowrouzezahrai. A versatile parameterization for measured material manifolds. Computer Graphics Forum, 37(2):135–144, May 2018. CODEN CG-FODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Schmitz:2010:EQC

[SSO^{+}10] L. Schmitz, L. F. Scheidegger, D. K. Osmari, C. A. Dietrich, and J. L. D. Comba. Efficient and quality contouring algorithms on the GPU. Computer Graphics Forum, 29(8):2569–2578, December 2010. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Strobelt:2012:TDR

[SSS^{+}12] H. Strobelt, M. Spicker, A. Stoffel, D. Keim, and O. Deussen. Text and documents: Rolled-out wordles: a heuristic method for overlap removal of 2D data representatives. Computer Graphics Forum, 31(3pt3):1135–1144, June 2012. CODEN CG-FODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Szafir:2016:TDD

[SSSG16] D. A. Szafir, D. Stuffer, Y. Sohail, and M. Gleicher. Text and document data: TextDNA: Visualizing word usage with configurable colorfields. Computer Graphics Forum, 35(3):421–430, June 2016. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Schultz:2013:UHH

[SSSSW13] T. Schultz, L. Schlaffke, B. Schölkopf, and T. Schmidt-Wilcke. Uncertainty: HiFiVE: a Hilbert space embedding of fiber variability estimates for uncertainty modeling and visualization. Computer Graphics Forum, 32(3pt1):121–130, June 2013. CO-DEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Saikia:2014:TEB

[SSW14a] Himangshu Saikia, Hans-Peter Seidel, and Tino Weinkauf. Topology: Extended branch decomposition graphs: Structural comparison of scalar data. Computer Graphics Forum, 33(3): 41–50, June 2014. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Schroers:2014:PCV

[SSW14b] C. Schroers, S. Setzer, and J. Weickert. Point clouds: a variational taxonomy for surface reconstruction from oriented points. Computer Graphics Forum, 33(5):195–204, August 2014. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Simon:2018:HHL

[ST18] P. M. Simon and C. Turkay. Hunting high and low: Visualising shifting correlations in financial markets. Computer Graphics Forum, 37(3):479–490, June 2018. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Smelik:2014:SPM

[STBB14] Ruben M. Smelik, Tim Tutenel, Rafael Bidarra, and Bedrich Benes. A survey on procedural modelling for virtual worlds. Computer Graphics Forum, 33(6):31–50, September 2014. CO-DEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Skouras:2012:VSC

[STBG12] Mélina Skouras, Bernhard Thomaszewski, Bernd Bickel, and Markus Gross. Visualization and simulation: Computational design of rubber balloons. Computer Graphics Forum, 31 (2pt4):835–844, May 2012. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Song:2016:RBS

[STC^{+}16] Dan Song, Ruofeng Tong, Jian Chang, Xiaosong Yang, Min Tang, and Jian Jun Zhang. Reconstruction: 3D body shapes estimation from dressed-human silhouettes. Computer Graphics Forum, 35(7):147–156, October 2016. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Steinberg:2019:ASI

[Ste19] S. Steinberg. Analytic spectral integration of birefringenceinduced iridescence. Computer Graphics Forum, 38(4):97–110, July 2019. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Schumacher:2016:MDS

[STG16] Christian Schumacher, Bernhard Thomaszewski, and Markus Gross. Modeling and design: Stenciling: Designing structurallysound surfaces with decorative patterns. Computer Graphics Forum, 35(5):101–110, August 2016. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Semmo:2012:IVI

[STKD12] Amir Semmo, Matthias Trapp, Jan Eric Kyprianidis, and Jürgen Döllner. Illustrative visualization: Interactive visualization of generalized virtual 3D city models using level-of-abstraction transitions. Computer Graphics Forum, 31(3pt1): 885–894, June 2012. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Sedlmair:2012:MDM

[STMT12] M. Sedlmair, A. Tatu, T. Munzner, and M. Tory. Multidimensional and multi-variate data: a taxonomy of visual cluster separation factors. Computer Graphics Forum, 31(3pt4): 1335–1344, June 2012. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Summa:2017:VUG

[STP17] B. Summa, J. Tierny, and V. Pascucci. Visualizing the uncertainty of graph-based 2D segmentation with min-path stability. Computer Graphics Forum, 36(3):133–143, June 2017. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Shrinivasan:2010:MDS

[SV10] Yedendra B. Shrinivasan and Jarke. J. Van Wijk. Multivariate data: Supporting exploratory analysis with the select & Slice table. Computer Graphics Forum, 29(3):803–812, June 2010. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Salvi:2014:SSI

[SV14] Péter Salvi and Tamás Várady. Surfaces: G^{2} surface interpolation over general topology curve networks. Computer Graphics Forum, 33(7):151–160, October 2014. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Shoham:2019:HFM

[SVBC19] Meged Shoham, Amir Vaxman, and Mirela Ben-Chen. Hierarchical functional maps between subdivision surfaces. Computer Graphics Forum, 38(5):55–73, August 2019. CODEN CG-FODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Slingsby:2016:GDV

[SvL16] A. Slingsby and E. van Loon. Geospatial data visualization: Exploratory visual analysis for animal movement ecology. Computer Graphics Forum, 35(3):471–480, June 2016. CODEN CG-FODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Salvi:2010:SOI

[SVLL10] Marco Salvi, Kiril Vidimče, Andrew Lauritzen, and Aaron Lefohn. Shadows and order independent transparency: Adaptive volumetric shadow maps. Computer Graphics Forum, 29 (4):1289–1296, June 2010. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Schrijvers:2013:IAV

[SvW13] Okke Schrijvers and Jarke J. van Wijk. Interactive analysis: Visual explanation of the complexity in Julia sets. Computer Graphics Forum, 32(3pt4):431–440, June 2013. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Solomon:2012:DFD

[SVWG12] Justin Solomon, Etienne Vouga, Max Wardetzky, and Eitan Grinspun. Deformation: Flexible developable surfaces. Computer Graphics Forum, 31(5):1567–1576, August 2012. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Sadlo:2010:TDD

[SW10] F. Sadlo and D. Weiskopf. Time-dependent 2-D vector field topology: An approach inspired by Lagrangian coherent structures. Computer Graphics Forum, 29(1):88–100, March 2010. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Sanftmann:2011:PAS

[SW11] H. Sanftmann and D. Weiskopf. Perception: Anaglyph stereo without ghosting. Computer Graphics Forum, 30(4):1251–1259, June 2011. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Saikia:2017:GFT

[SW17] H. Saikia and T. Weinkauf. Global feature tracking and similarity estimation in time-dependent scalar fields. Computer Graphics Forum, 36(3):1–11, June 2017. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

StJean:2016:TSD

[SWG16] C. St.Jean, C. Ware, and R. Gamble. Time series data and sequences: Dynamic change arcs to explore model forecasts. Computer Graphics Forum, 35(3):311–320, June 2016. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Sewall:2010:VWC

[SWML10] J. Sewall, D. Wilkie, P. Merrell, and M. C. Lin. Virtual worlds: Continuum traffic simulation. Computer Graphics Forum, 29 (2):439–448, May 2010. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Scherzer:2011:SRT

[SWP11] Daniel Scherzer, Michael Wimmer, and Werner Purgathofer. A survey of real-time hard shadow mapping methods. Computer Graphics Forum, 30(1):169–186, March 2011. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Steinberger:2012:IMU

[SWS12] Markus Steinberger, Manuela Waldner, and Dieter Schmalstieg. Interactive modeling and user interfaces: Interactive self-organizing Windows. Computer Graphics Forum, 31(2pt3): 621–630, May 2012. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Sato:2018:ENB

[SWT^{+}18] T. Sato, C. Wojtan, N. Thuerey, T. Igarashi, and R. Ando. Extended narrow band FLIP for liquid simulations. Computer Graphics Forum, 37(2):169–177, May 2018. CODEN CG-FODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Song:2019:SAS

[SXLS19] Chengfang Song, Chunxia Xiao, Ling Lei, and Haigang Sui. Scale-adaptive structure-preserving texture filtering. Computer Graphics Forum, 38(7):149–158, October 2019. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Shao:2011:MRA

[SXY^{+}11] Tianjia Shao, Weiwei Xu, Kangkang Yin, Jingdong Wang, Kun Zhou, and Baining Guo. Model reconstruction & analysis: Discriminative sketch-based 3D model retrieval via robust shape matching. Computer Graphics Forum, 30(7):2011–2020, September 2011. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Sahillioglu:2011:SCF

[SY11] Y. Sahillioğlu and Y. Yemez. Section 4: Coarse-to-fine combinatorial matching for dense isometric shape correspondence. Computer Graphics Forum, 30(5):1461–1470, August 2011. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Sahillioglu:2012:SPM

[SY12a] Y. Sahillioğlu and Y. Yemez. Shape processing and modeling: Scale normalization for isometric shape matching. Computer Graphics Forum, 31(7pt2):2233–2240, September 2012. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Seversky:2012:GPG

[SY12b] L. M. Seversky and L. Yin. Geometry processing: a global parity measure for incomplete point cloud data. Computer Graphics Forum, 31(7pt1):2097–2106, September 2012. CODEN CG-FODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Sahillioglu:2013:CFI

[SY13] Y. Sahillioğlu and Y. Yemez. Coarse-to-fine isometric shape correspondence by tracking symmetric flips. Computer Graphics Forum, 32(1):177–189, February 2013. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Sahillioglu:2014:PDC

[SY14a] Y. Sahillioğlu and Y. Yemez. Partial 3-D correspondence from shape extremities. Computer Graphics Forum, 33(6):63–76, September 2014. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Sahillioglu:2014:SCM

[SY14b] Y. Sahillioğlu and Y. Yemez. Shapes and cryptography: Multiple shape correspondence by dynamic programming. Computer Graphics Forum, 33(7):121–130, October 2014. CODEN CG-FODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Sheng:2010:ISA

[SYC10] Yu Sheng, Theodore C. Yapo, and Barbara Cutler. Implicit surfaces and augmented reality: Global illumination compensation for spatially augmented reality. Computer Graphics Forum, 29 (2):387–396, May 2010. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Salman:2010:RFP

[SYM10] Nader Salman, Mariette Yvinec, and Quentin Merigot. Reconstruction I: Feature preserving mesh generation from 3D point clouds. Computer Graphics Forum, 29(5):1623–1632, July 2010. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Scherzer:2012:TCM

[SYM^{+}12] Daniel Scherzer, Lei Yang, Oliver Mattausch, Diego Nehab, Pedro V. Sander, Michael Wimmer, and Elmar Eisemann. Temporal coherence methods in real-time rendering. Computer Graphics Forum, 31(8):2378–2408, December 2012. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Shi:2013:GSP

[SYT^{+}13] K.-L. Shi, J.-H. Yong, L. Tang, J.-G. Sun, and J.-C. Paul. Geometry (session 2): Polar NURBS surface with curvature continuity. Computer Graphics Forum, 32(7):363–370, October 2013. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Sun:2018:CCC

[SZ18] Lanyin Sun and Chungang Zhu. Curvature continuity conditions between adjacent toric surface patches. Computer Graphics Forum, 37(7):469–477, October 2018. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Shao:2015:SFF

[SZMTW15] X. Shao, Z. Zhou, N. Magnenat-Thalmann, and W. Wu. Stable and fast fluid–solid coupling for incompressible SPH. Computer Graphics Forum, 34(1):191–204, February 2015. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Szymczak:2011:TBV

[Szy11] Andrzej Szymczak. Topology-based visualization: Stable Morse decompositions for piecewise constant vector fields on surfaces. Computer Graphics Forum, 30(3):851–860, June 2011. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Tonneau:2016:CAC

[TAAP^{+}16] Steve Tonneau, Rami Ali Al-Ashqar, Julien Pettré, Taku Komura, and Nicolas Mansard. Character animation: Character contact re-positioning under large environment deformation. Computer Graphics Forum, 35(2):127–138, May 2016. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Tursun:2015:SAR

[TAEE15] Okan Tarhan Tursun, Ahmet Oğuz Akyüz, Aykut Erdem, and Erkut Erdem. State of the art reports: The state of the art in HDR deghosting: a survey and evaluation. Computer Graphics Forum, 34(2):683–707, May 2015. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Tursun:2016:HIO

[TAEE16] Okan Tarhan Tursun, Ahmet Oğuz Akyüz, Aykut Erdem, and Erkut Erdem. HDR imaging: an objective deghosting quality metric for HDR images. Computer Graphics Forum, 35(2):139– 152, May 2016. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Takayama:2019:DSM

[Tak19] Kenshi Takayama. Dual sheet meshing: an interactive approach to robust hexahedralization. Computer Graphics Forum, 38(2): 37–48, May 2019. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Tagliasacchi:2012:FMC

[TAOZ12] Andrea Tagliasacchi, Ibraheem Alhashim, Matt Olson, and Hao Zhang. Flows: Mean curvature skeletons. Computer Graphics Forum, 31(5):1735–1744, August 2012. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Tao:2012:ICP

[TBKP12] Michael Tao, Jiamin Bai, Pushmeet Kohli, and Sylvain Paris. Image and color processing: SimpleFlow: a non-iterative, sublinear optical flow algorithm. Computer Graphics Forum, 31 (2pt1):345–353, May 2012. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Thony:2018:LSP

[TBP18] Matthias Thöny, Markus Billeter, and Renato Pajarola. Largescale pixel-precise deferred vector maps. Computer Graphics Forum, 37(1):338–349, February 2018. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Thiery:2012:SPM

[TBTB12] Jean-Marc Thiery, Bert Buchholz, Julien Tierny, and Tamy Boubekeur. Shape processing and modeling: Analytic curve skeletons for 3D surface modeling and processing. Computer Graphics Forum, 31(7pt2):2223–2232, September 2012. CO-DEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Tevs:2011:SMM

[TBW^{+}11] A. Tevs, A. Berner, M. Wand, I. Ihrke, and H.-P. Seidel. Shape matching and manipulation: Intrinsic shape matching by planned landmark sampling. Computer Graphics Forum, 30 (2):543–552, April 2011. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Turmukhambetov:2015:ISD

[TCGK15] Daniyar Turmukhambetov, Neill D. F. Campbell, Dan B. Goldman, and Jan Kautz. Interactive sketch-driven image synthesis. Computer Graphics Forum, 34(8):130–142, December 2015. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Tang:2012:AIR

[TCLK12] Jeff K. T. Tang, Jacky C. P. Chan, Howard Leung, and Taku Komura. Animation: Interaction retrieval by spacetime proximity graphs. Computer Graphics Forum, 31(2pt4):745–754, May 2012. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Tikhonova:2010:VVE

[TCM10] A. Tikhonova, C. D. Correa, and K.-L. Ma. Volume visualization: an exploratory technique for coherent visualization of time-varying volume data. Computer Graphics Forum, 29 (3):783–792, June 2010. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Thonat:2018:TSI

[TDDD18] Theo Thonat, Abdelaziz Djelouah, Fredo Durand, and George Drettakis. Thin structures in image based rendering. Computer Graphics Forum, 37(4):107–118, July 2018. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Takahashi:2015:FFI

[TDF^{+}15] Tetsuya Takahashi, Yoshinori Dobashi, Issei Fujishiro, Tomoyuki Nishita, and Ming C. Lin. Fluids & flows: Implicit formulation for SPH-based viscous fluids. Computer Graphics Forum, 34(2):493–502, May 2015. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Templin:2014:CIP

[TDMS14] Krzysztof Templin, Piotr Didyk, Karol Myszkowski, and Hans-Peter Seidel. Color and imaging: Perceptually-motivated stereoscopic film grain. Computer Graphics Forum, 33(7): 349–358, October 2014. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Takahashi:2018:EHI

[TDNL18] Tetsuya Takahashi, Yoshinori Dobashi, Tomoyuki Nishita, and Ming C. Lin. An efficient hybrid incompressible SPH solver with interface handling for boundary conditions. Computer Graphics Forum, 37(1):313–324, February 2018. CODEN CG-FODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Tagliasacchi:2016:SAR

[TDS^{+}16] Andrea Tagliasacchi, Thomas Delame, Michela Spagnuolo, Nina Amenta, and Alexandru Telea. State of the art reports: 3D skeletons: A state-of-the-art report. Computer Graphics Forum, 35(2):573–597, May 2016. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Telea:2010:GIB

[TE10] A. Telea and O. Ersoy. Graphs: Image-based edge bundles: Simplified visualization of large graphs. Computer Graphics Forum, 29(3):843–852, June 2010. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Thiery:2018:ARE

[TE18] J.-M. Thiery and E. Eisemann. ARAPLBS: Robust and efficient elasticity-based optimization of weights and skeleton joints for linear blend skinning with parametrized bones. Computer Graphics Forum, 37(1):32–44, February 2018. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Tang:2015:RHM

[TF15] Yizhi Tang and Jieqing Feng. Registration: Hierarchical multiview rigid registration. Computer Graphics Forum, 34(5):77– 87, August 2015. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Tam:2011:PSA

[TFA^{+}11] G. K. L. Tam, H. Fang, A. J. Aubrey, P. W. Grant, P. L. Rosin, D. Marshall, and M. Chen. Parameter spaces and analysis: Visualization of time-series data in parameter space for understanding facial dynamics. Computer Graphics Forum, 30 (3):901–910, June 2011. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Tominski:2017:ILV

[TGK^{+}17] C. Tominski, S. Gladisch, U. Kister, R. Dachselt, and H. Schumann. Interactive lenses for visualization: an extended survey. Computer Graphics Forum, 36(6):173–200, September 2017. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Tasse:2012:ETB

[TGM12] F. P. Tasse, J. Gain, and P. Marais. Enhanced texture-based terrain synthesis on graphics hardware. Computer Graphics Forum, 31(6):1959–1972, September 2012. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Tokuyoshi:2017:LSS

[TH17] Yusuke Tokuyoshi and Takahiro Harada. Lighting and shading: Stochastic light culling for VPLs on GGX microsurfaces. Computer Graphics Forum, 36(4):55–63, July 2017. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Thierry:2011:ITN

[Thi11] Guillaume Thierry. Invited talks: Neuromatrix: The world is embodied in our brain. Computer Graphics Forum, 30(2):xxii, April 2011. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Tsukamoto:2017:DOF

[TIK17] J. Tsukamoto, D. Iwai, and K. Kashima. Distributed optimization framework for shadow removal in multi-projection systems. Computer Graphics Forum, 36(8):369–379, December 2017. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Timonen:2012:LCI

[Tim12] Ville Timonen. Low-complexity intervisibility in height fields. Computer Graphics Forum, 31(8):2348–2362, December 2012. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Timonen:2013:SLS

[Tim13] Ville Timonen. Shadows: Line-sweep ambient obscurance. Computer Graphics Forum, 32(4):97–105, July 2013. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Takahashi:2016:FMS

[TL16] Tetsuya Takahashi and Ming C. Lin. Fluids: a multilevel SPH solver with unified solid boundary handling. Computer Graphics Forum, 35(7):517–526, October 2016. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Takahashi:2019:GCV

[TL19] Tetsuya Takahashi and Ming C. Lin. A geometrically consistent viscous fluid solver with two-way fluid-solid coupling. Computer Graphics Forum, 38(2):49–58, May 2019. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Tarameshloo:2016:NGU

[TLFC16] E. Tarameshloo, M. Hosseinkhani Loorak, P. W. L. Fong, and S. Carpendale. Networks and graphs 2: Using visualization to explore original and anonymized LBSN data. Computer Graphics Forum, 35(3):291–300, June 2016. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Tu:2019:PPS

[TLH19] Peihan Tu, Dani Lischinski, and Hui Huang. Point pattern synthesis via irregular convolution. Computer Graphics Forum, 38(5):109–122, August 2019. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Tanahashi:2016:VNS

[TLM16] Yuzuru Tanahashi, Nick Leaf, and Kwan-Liu Ma. Visualization/NPR: a study on designing effective introductory materials for information visualization. Computer Graphics Forum, 35(7):117–126, October 2016. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Thiagarajan:2018:EHD

[TLRB18] J. J. Thiagarajan, S. Liu, K. N. Ramamurthy, and P.-T. Bremer. Exploring high-dimensional structure via axis-aligned decomposition of linear projections. Computer Graphics Forum, 37(3):241–251, June 2018. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Tao:2019:KIP

[TLW^{+}19] Wenbo Tao, Xiaoyu Liu, Yedi Wang, Leilani Battle, Çağatay Demiralp, Remco Chang, and Michael Stonebraker. Kyrix: Interactive pan/zoom visualizations at scale. Computer Graphics Forum, 38(3):529–540, June 2019. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Tao:2013:PPS

[TM13] Michael W. Tao and Jitendra Malik. Pixels and points: Sharpening out of focus images using high-frequency transfer. Computer Graphics Forum, 32(2pt4):489–498, May 2013. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Trentacoste:2011:IVB

[TMH11] Matthew Trentacoste, Rafal Mantiuk, and Wolfgang Heidrich. Image and video: Blur-aware image downsampling. Computer Graphics Forum, 30(2):573–582, April 2011. CODEN CG-FODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Trentacoste:2012:PGU

[TMHD12] Matthew Trentacoste, Ratal Mantiuk, Wolfgang Heidrich, and Florian Dufrot. Perceptual graphics: Unsharp masking, countershading and halos: Enhancements or artifacts? Computer Graphics Forum, 31(2pt3):555–564, May 2012. CODEN CG-FODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Tam:2014:MEA

[TMRL14] Gary K. L. Tam, Ralph R. Martin, Paul L. Rosin, and Yu-Kun Lai. Matching: an efficient approach to correspondences between multiple non-rigid parts. Computer Graphics Forum, 33(5):137–146, August 2014. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Tokuyoshi:2015:SLA

[Tok15a] Y. Tokuyoshi. Specular lobe-aware filtering and upsampling for interactive indirect illumination. Computer Graphics Forum, 34 (6):135–147, September 2015. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Tokuyoshi:2015:RVS

[Tok15b] Yusuke Tokuyoshi. Rendering: Virtual spherical Gaussian lights for real-time glossy indirect illumination. Computer Graphics Forum, 34(7):89–98, October 2015. CODEN CG-FODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Tagliasacchi:2011:SVV

[TOZ^{+}11] Andrea Tagliasacchi, Matt Olson, Hao Zhang, Ghassan Hamarneh, and Daniel Cohen-Or. Section 7: VASE: Volumeaware surface evolution for surface reconstruction from incomplete point clouds. Computer Graphics Forum, 30(5):1563– 1571, August 2011. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Tarini:2010:MPP

[TPC^{+}10] Marco Tarini, Nico Pietroni, Paolo Cignoni, Daniele Panozzo, and Enrico Puppo. Mesh processing: Practical quad mesh simplification. Computer Graphics Forum, 29(2):407–418, May 2010. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Turkay:2011:TDV

[TPRH11] C. Turkay, J. Parulek, N. Reuter, and H. Hauser. Temporal data visualization: Interactive visual analysis of temporal cluster structures. Computer Graphics Forum, 30(3):711–720, June 2011. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Takayama:2014:MPB

[TPSH14a] Kenshi Takayama, Daniele Panozzo, and Olga Sorkine-Hornung. Meshing: Pattern-based quadrangulation for N-sided patches. Computer Graphics Forum, 33(5):177–184, August 2014. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Tarini:2014:AIA

[TPSH14b] Marco Tarini, Daniele Panozzo, and Olga Sorkine-Hornung. Animation III: Accurate and efficient lighting for skinned models. Computer Graphics Forum, 33(2):421–428, May 2014. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Treib:2012:MMI

[TRAW12] Marc Treib, Florian Reichl, Stefan Auer, and Rüdiger Westermann. Massive models: Interactive editing of GigaSample terrain fields. Computer Graphics Forum, 31(2pt2):383–392, May 2012. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Tao:2016:FTN

[TSB16] Michael Tao, Justin Solomon, and Adrian Butscher. Fitting and tracking: Near-isometric level set tracking. Computer Graphics Forum, 35(5):65–77, August 2016. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Tokuyoshi:2013:RSA

[TSdSK13] Yusuke Tokuyoshi, Takashi Sekine, Tiago da Silva, and Takashi Kanai. Rendering (session 2): Adaptive ray-bundle tracing with memory usage prediction: Efficient global illumination in large scenes. Computer Graphics Forum, 32(7):315–324, October 2013. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Tomcin:2014:APB

[TSK14] Robin Tomcin, Dominik Sibbing, and Leif Kobbelt. Animation I (physically-based): Efficient enforcement of hard articulation constraints in the presence of closed loops and contacts. Computer Graphics Forum, 33(2):235–244, May 2014. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Tsai:2016:FAH

[TSPP16] Yun-Ta Tsai, Markus Steinberger, Dawid Paja̧k, and Kari Pulli. Fast ANN for high-quality collaborative filtering. Computer Graphics Forum, 35(1):138–151, February 2016. CODEN CG-FODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Takayama:2011:IGP

[TSS^{+}11] Kenshi Takayama, Ryan Schmidt, Karan Singh, Takeo Igarashi, Tamy Boubekeur, and Olga Sorkine. Interactive geometric processing: GeoBrush: Interactive mesh geometry cloning. Computer Graphics Forum, 30(2):613–622, April 2011. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Tagliasacchi:2015:RRA

[TST^{+}15] Andrea Tagliasacchi, Matthias Schröder, Anastasia Tkach, Sofien Bouaziz, Mario Botsch, and Mark Pauly. Registration: Robust articulated-ICP for real-time hand tracking. Computer Graphics Forum, 34(5):101–114, August 2015. CODEN CG-FODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Tang:2019:CDS

[TSW^{+}19] Keke Tang, Peng Song, Xiaofei Wang, Bailin Deng, Chi-Wing Fu, and Ligang Liu. Computational design of steady 3D dissection puzzles. Computer Graphics Forum, 38(2):291–303, May 2019. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Thiery:2012:CCB

[TTB12] Jean-Marc Thiery, Julien Tierny, and Tamy Boubekeur. CageR: Cage-based reverse engineering of animated 3D shapes. Computer Graphics Forum, 31(8):2303–2316, December 2012. CO-DEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Tang:2013:ASG

[TTN^{+}13] Min Tang, Ruofeng Tong, Rahul Narain, Chang Meng, and Dinesh Manocha. Animation (session 1): a GPU-based streaming algorithm for high-resolution cloth simulation. Computer Graphics Forum, 32(7):21–30, October 2013. CODEN CG-FODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Timonen:2010:RIS

[TW10] Ville Timonen and Jan Westerholm. Rendering II: Scalable height field self-shadowing. Computer Graphics Forum, 29 (2):723–731, May 2010. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Tao:2016:VDV

[TWC^{+}16] Yubo Tao, Qirui Wang, Wei Chen, Yingcai Wu, and Hai Lin. Volume data visualization: Similarity voting based viewpoint selection for volumes. Computer Graphics Forum, 35(3):391– 400, June 2016. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Thiyagalingam:2013:UCP

[TWD^{+}13] Jeyarajan Thiyagalingam, Simon Walton, Brian Duffy, Anne Trefethen, and Min Chen. Uncertainty: Complexity plots. Computer Graphics Forum, 32(3pt1):111–120, June 2013. CO-DEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Torsney-Weir:2018:HIV

[TWMSK18] T. Torsney-Weir, T. Möller, M. Sedlmair, and R. M. Kirby. Hypersliceplorer: Interactive visualization of shapes in multiple dimensions. Computer Graphics Forum, 37(3):229–240, June 2018. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Takahashi:2011:MSP

[TWS^{+}11] Shigeo Takahashi, Hsiang-Yun Wu, Seow Hui Saw, Chun-Cheng Lin, and Hsu-Chun Yen. Mesh segmentation & processing: Optimized topological surgery for unfolding 3D meshes. Computer Graphics Forum, 30(7):2077–2086, September 2011. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Torsney-Weir:2017:SSM

[TWSM17] T. Torsney-Weir, M. Sedlmair, and T. Möller. Sliceplorer: 1D slices for multi-dimensional continuous functions. Computer Graphics Forum, 36(3):167–177, June 2017. CODEN CG-FODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Tang:2016:CAC

[TWT^{+}16] Min Tang, Huamin Wang, Le Tang, Ruofeng Tong, and Dinesh Manocha. Cloth & animation: CAMA: Contact-aware matrix assembly with unified collision handling for GPU-based cloth simulation. Computer Graphics Forum, 35(2):511–521, May 2016. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Tian:2016:IVF

[TX16] Lihui Tian and Shuangjiu Xiao. Images and video: Facial feature exaggeration according to social psychology of face perception. Computer Graphics Forum, 35(7):391–399, October 2016. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Tan:2019:ICP

[TXGW19] Xuehan Tan, Panpan Xu, Shihui Guo, and Wencheng Wang. Image composition of partially occluded objects. Computer Graphics Forum, 38(7):641–650, October 2019. CODEN CG-FODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Tang:2019:FGL

[TXL^{+}19] Shusen Tang, Zeqing Xia, Zhouhui Lian, Yingmin Tang, and Jianguo Xiao. FontRNN: Generating large-scale Chinese fonts via recurrent neural network. Computer Graphics Forum, 38 (7):567–577, October 2019. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Tong:2011:IVE

[TZD11] Ruo-Feng Tong, Yun Zhang, and Meng Ding. Image & video editing: Video brush: a novel interface for efficient video cutout. Computer Graphics Forum, 30(7):2049–2057, September 2011. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Um:2014:PDA

[UBH14] Kiwon Um, Seungho Baek, and JungHyun Han. Particles and deformation: Advanced hybrid particle-grid method with subgrid particle correction. Computer Graphics Forum, 33(7): 209–218, October 2014. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Uffinger:2010:VIH

[ÜFE10] Markus Üffinger, Steffen Frey, and Thomas Ertl. Visualization: Interactive high-quality visualization of higher-order finite elements. Computer Graphics Forum, 29(2):337–346, May 2010. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Urena:2013:VSA

[UFK13] Carlos Ureña, Marcos Fajardo, and Alan King. Visibility/ sampling: an area-preserving parametrization for spherical rectangles. Computer Graphics Forum, 32(4):59–66, July 2013. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Urena:2018:SSP

[UG18] Carlos Ureña and Iliyan Georgiev. Stratified sampling of projected spherical caps. Computer Graphics Forum, 37(4):13–20, July 2018. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Um:2018:LSM

[UHT18] Kiwon Um, Xiangyu Hu, and Nils Thuerey. Liquid splash modeling with neural networks. Computer Graphics Forum, 37(8): 171–182, December 2018. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Untereiner:2015:CCR

[UKCB15] L. Untereiner, P. Kraemer, D. Cazier, and D. Bechmann. CPH: a compact representation for hierarchical meshes generated by primal refinement. Computer Graphics Forum, 34(8):155–166, December 2015. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Ulu:2019:SDU

[UMK19] E. Ulu, J. McCann, and L. B. Kara. Structural design using Laplacian shells. Computer Graphics Forum, 38(5):85–98, August 2019. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Urbano:2010:TMO

[UMM^{+}10] C. Urbano, L. Magalhães, J. Moura, M. Bessa, A. Marcos, and A. Chalmers. Tone mapping operators on small screen devices: An evaluation study. Computer Graphics Forum, 29(8):2469– 2478, December 2010. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Umenhoffer:2011:NPR

[USSK11] Tamás Umenhoffer, László Szécsi, and László Szirmay-Kalos. Non-photorealistic rendering: Hatching for motion picture production. Computer Graphics Forum, 30(2):533–542, April 2011. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Ureta:2016:FIM

[UTZ16] Francisca Gil Ureta, Chelsea Tymms, and Denis Zorin. Fabrication: Interactive modeling of mechanical objects. Computer Graphics Forum, 35(5):145–155, August 2016. CODEN CG-FODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Usher:2019:SRT

[UWA^{+}19] Will Usher, Ingo Wald, Jefferson Amstutz, Johannes Günther, Carson Brownlee, and Valerio Pascucci. Scalable ray tracing using the Distributed FrameBuffer. Computer Graphics Forum, 38(3):455–466, June 2019. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Vicini:2019:DDM

[VAN^{+}19] D. Vicini, D. Adler, J. Novák, F. Rousselle, and B. Burley. Denoising deep Monte Carlo renderings. Computer Graphics Forum, 38(1):316–327, February 2019. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Vanegas:2010:MAB

[VAW^{+}10] C. A. Vanegas, D. G. Aliaga, P. Wonka, P. Müller, P. Waddell, and B. Watson. Modelling the appearance and behaviour of urban spaces. Computer Graphics Forum, 29(1):25–42, March 2010. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Vaxman:2012:MMP

[Vax12] Amir Vaxman. Modeling: Modeling polyhedral meshes with affine maps. Computer Graphics Forum, 31(5):1647–1656, August 2012. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Vaxman:2014:PFP

[Vax14] Amir Vaxman. A projective framework for polyhedral mesh modelling. Computer Graphics Forum, 33(8):121–131, December 2014. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Vasa:2014:AIM

[VB14a] L. Váša and G. Brunnett. Animation II (motion data processing): Rate-distortion optimized compression of motion capture data. Computer Graphics Forum, 33(2):283–292, May 2014. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Vasa:2014:CFE

[VB14b] Libor Váša and Guido Brunnett. Coordinates and fields: Efficient encoding of texture coordinates guided by mesh geometry. Computer Graphics Forum, 33(5):25–34, August 2014. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Vehlow:2015:VEC

[VBAW15] C. Vehlow, F. Beck, P. Auwärter, and D. Weiskopf. Visualizing the evolution of communities in dynamic graphs. Computer Graphics Forum, 34(1):277–288, February 2015. CODEN CG-FODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

vanBasten:2011:AFS

[vBE11] B. J. H. van Basten and A. Egges. Animation: Flexible splicing of upper-body motion spaces on locomotion. Computer Graphics Forum, 30(7):1963–1971, September 2011. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Vinkler:2013:MPH

[VBHH13] Marek Vinkler, Jiří Bittner, Vlastimil Havran, and Michal Hapala. Massively parallel hierarchical scene processing with applications in rendering. Computer Graphics Forum, 32(8): 13–25, December 2013. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Vehlow:2017:VGS

[VBW17] Corinna Vehlow, Fabian Beck, and Daniel Weiskopf. Visualizing group structures in graphs: a survey. Computer Graphics Forum, 36(6):201–225, September 2017. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Vaxman:2016:SAR

[VCD^{+}16] Amir Vaxman, Marcel Campen, Olga Diamanti, Daniele Panozzo, David Bommes, Klaus Hildebrandt, and Mirela Ben-Chen. State of the art reports: Directional field synthesis, design, and processing. Computer Graphics Forum, 35(2):545– 572, May 2016. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Vangorp:2011:PPV

[VCL^{+}11] P. Vangorp, G. Chaurasia, P.-Y. Laffont, R. W. Fleming, and G. Drettakis. Perception: Perception of visual artifacts in image-based rendering of Façades. Computer Graphics Forum, 30(4):1241–1250, June 2011. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Vad:2014:LLD

[VCRG14] Viktor Vad, Balázs Csébfalvi, Peter Rautek, and Eduard Gröller. Low-level direct volume rendering / filtering: Towards an unbiased comparison of CC, BCC, and FCC lattices in terms of prealiasing. Computer Graphics Forum, 33(3):81–90, June 2014. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Vasa:2018:EPC

[VD18] L. Váša and J. Dvořák. Error propagation control in Laplacian mesh compression. Computer Graphics Forum, 37(5):61–70, August 2018. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

vanderCorput:2014:BIV

[vdCAvW14] Paul van der Corput, Johan Arends, and Jarke J. van Wijk. BioMedical II: Visualization of medicine prescription behavior. Computer Graphics Forum, 33(3):161–170, June 2014. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

vanderCorput:2016:HDD

[vdCvW16] Paul van der Corput and Jarke J. van Wijk. High-dimensional data: Exploring items and features with I^{F}, F^{I}-tables. Computer Graphics Forum, 35(3):31–40, June 2016. CODEN CG-FODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

vanderCorput:2017:CPI

[vdCvW17] Paul van der Corput and Jarke J. van Wijk. Comparing personal image collections with PICTuReVis. Computer Graphics Forum, 36(3):295–304, June 2017. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

vandenElzen:2013:ITS

[vdEvW13] Stef van den Elzen and Jarke J. van Wijk. InfoVis techniques: Small multiples, large singles: A new approach for visual data exploration. Computer Graphics Forum, 32(3pt2): 191–200, June 2013. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

vanDijk:2016:GDV

[vDHO16] Thomas C. van Dijk, Jan-Henrik Haunert, and Johannes Oehrlein. Geospatial data visualization: Location-dependent generalization of road networks based on equivalent destinations. Computer Graphics Forum, 35(3):451–460, June 2016. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

vandeKerkhof:2019:DAG

[vdKdJP^{+}19] Mees van de Kerkhof, Tim de Jong, Raphael Parment, Maarten Löffler, Amir Vaxman, and Marc van Kreveld. Design and automated generation of Japanese picture puzzles. Computer Graphics Forum, 38(2):343–353, May 2019. CODEN CG-FODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

vanDortmont:2019:CEE

[vDvdEvW19] M. A. M. M. van Dortmont, S. van den Elzen, and J. J. van Wijk. ChronoCorrelator: Enriching events with time series. Computer Graphics Forum, 38(3):387–399, June 2019. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

vanderZwan:2011:MDV

[vdZLBI11] Matthew van der Zwan, Wouter Lueks, Henk Bekker, and Tobias Isenberg. Molecular data visualization: Illustrative molecular visualization with continuous abstraction. Computer Graphics Forum, 30(3):683–690, June 2011. CODEN CG-FODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Vasilakis:2014:AIM

[VF14] Andreas A. Vasilakis and Ioannis Fudos. Animation II (motion data processing): Pose partitioning for multi-resolution segmentation of arbitrary mesh animations. Computer Graphics Forum, 33(2):293–302, May 2014. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Vavra:2016:RTA

[VF16] R. Vávra and J. Filip. Ray tracing/appearance capture: Minimal sampling for effective acquisition of anisotropic BRDFs. Computer Graphics Forum, 35(7):299–309, October 2016. CO-DEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

vFestenberg:2011:DBS

[vFG11] Niels v. Festenberg and Stefan Gumhold. Diffusion-based snow cover generation. Computer Graphics Forum, 30(6):1837–1849, September 2011. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Vanek:2014:SSC

[VGB14a] J. Vanek, J. A. G. Galicia, and B. Benes. Segmentation and structure: Clever support: Efficient support structure generation for digital fabrication. Computer Graphics Forum, 33(5): 117–125, August 2014. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Vanek:2014:PPV

[VGB^{+}14b] J. Vanek, J. A. Garcia Galicia, B. Benes, R. Měch, N. Carr, O. Stava, and G. S. Miller. PackMerger: a 3D print volume optimizer. Computer Graphics Forum, 33(6):322–332, September 2014. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

vanGarderen:2017:MDO

[vGPNB17] M. van Garderen, B. Pampel, A. Nocaj, and U. Brandes. Minimum-displacement overlap removal for geo-referenced data visualization. Computer Graphics Forum, 36(3):423–433, June 2017. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Vibert:2019:SVR

[VGS^{+}19] N. Vibert, A. Gruson, H. Stokholm, T Mortensen, W. Jarosz, T. Hachisuka, and D. Nowrouzezahrai. Scalable virtual ray lights rendering for participating media. Computer Graphics Forum, 38(4):57–65, July 2019. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Vinkler:2015:RED

[VH15] M. Vinkler and V. Havran. Register efficient dynamic memory allocator for GPUs. Computer Graphics Forum, 34(8):143–154, December 2015. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Vinkler:2016:PCB

[VHB16] Marek Vinkler, Vlastimil Havran, and Jiří Bittner. Performance comparison of bounding volume hierarchies and kd-trees for GPU ray tracing. Computer Graphics Forum, 35(8):68–79, December 2016. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Vazquez:2018:VAP

[VHG^{+}18] P. Vázquez, P. Hermosilla, V. Guallar, J. Estrada, and A. Vinacua. Visual analysis of protein-ligand interactions. Computer Graphics Forum, 37(3):391–402, June 2018. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Verstraaten:2018:LHR

[VK18] T. W. Verstraaten and J. Kosinka. Local and hierarchical refinement for subdivision gradient meshes. Computer Graphics Forum, 37(7):373–383, October 2018. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Viitanen:2017:GTF

[VKJ^{+}17] T. Viitanen, M. Koskela, P. Jääskeläinen, K. Immonen, and J. Takala. Geometric techniques: Fast hardware construction and refitting of quantized bounding volume hierarchies. Computer Graphics Forum, 36(4):167–178, July 2017. CODEN CG-FODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

vanKaick:2011:SMM

[vKTS^{+}11] Oliver van Kaick, Andrea Tagliasacchi, Oana Sidi, Hao Zhang, Daniel Cohen-Or, Lior Wolf, and Ghassan Hamarneh. Shape matching and manipulation: Prior knowledge for part correspondence. Computer Graphics Forum, 30(2):553–562, April 2011. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

vanKreveld:2011:SSS

[vKvLV11] M. van Kreveld, T. van Lankveld, and R. C. Veltkamp. Section 7: On the shape of a set of points and lines in the plane. Computer Graphics Forum, 30(5):1553–1562, August 2011. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

vanKreveld:2013:RWS

[vKvLV13] M. van Kreveld, T. van Lankveld, and R. C. Veltkamp. Reconstruction: Watertight scenes from urban LiDAR and planar surfaces. Computer Graphics Forum, 32(5):217–228, August 2013. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Vanegas:2012:CBP

[VKW^{+}12] Carlos A. Vanegas, Tom Kelly, Basil Weber, Jan Halatsch, Daniel G. Aliaga, and Pascal Müller. Cities and buildings: Procedural generation of parcels in urban modeling. Computer Graphics Forum, 31(2pt3):681–690, May 2012. CODEN CG-FODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

vanKaick:2013:BMP

[vKZH13] Oliver van Kaick, Hao Zhang, and Ghassan Hamarneh. Bilateral maps for partial matching. Computer Graphics Forum, 32 (6):189–200, September 2013. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

vanKaick:2011:SSC

[vKZHCO11] Oliver van Kaick, Hao Zhang, Ghassan Hamarneh, and Daniel Cohen-Or. A survey on shape correspondence. Computer Graphics Forum, 30(6):1681–1707, September 2011. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

vonLandesberger:2011:VAL

[vLKS^{+}11] T. von Landesberger, A. Kuijper, T. Schreck, J. Kohlhammer, J. J. Van Wijk, J.-D. Fekete, and D. W. Fellner. Visual analysis of large graphs: State-of-the-art and future research challenges. Computer Graphics Forum, 30(6):1719–1749, September 2011. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Viau:2012:GRC

[VM12] C. Viau and M. J. McGuffin. Graphs and relationships: ConnectedCharts: Explicit visualization of relationships between data graphics. Computer Graphics Forum, 31(3pt4):1285–1294, June 2012. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Volke:2013:TDV

[VMH^{+}13] S. Volke, M. Middendorf, M. Hlawitschka, J. Kasten, D. Zeckzer, and G. Scheuermann. Topology: dPSO-Vis: Topology-based visualization of discrete particle swarm optimization. Computer Graphics Forum, 32(3pt3):351–360, June 2013. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Vasa:2014:SCD

[VMHB14] L. Váša, S. Marras, K. Hormann, and G. Brunnett. Surfaces I: Compressing dynamic meshes with geometric Laplacians. Computer Graphics Forum, 33(2):145–154, May 2014. CODEN CG-FODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

vonMarcard:2017:SIP

[vMRBPM17] T. von Marcard, B. Rosenhahn, M. J. Black, and G. Pons-Moll. Sparse inertial poser: Automatic 3D human pose estimation from sparse IMUs. Computer Graphics Forum, 36(2):349–360, May 2017. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Vogiannou:2010:CEB

[VMTS10] Athanasios Vogiannou, Konstantinos Moustakas, Dimitrios Tzovaras, and Michael G. Strintzis. Collisions: Enhancing bounding volumes using support plane mappings for collision detection. Computer Graphics Forum, 29(5):1595–1604, July 2010. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Vo:2010:HSE

[VOS^{+}10] Huy T. Vo, Daniel K. Osmari, Brian Summa, João L. D. Comba, Valerio Pascucci, and Cláudio T. Silva. Hardware: Streaming-enabled parallel dataflow architecture for multicore systems. Computer Graphics Forum, 29(3):1073–1082, June 2010. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Vasa:2011:SOP

[VP11] L. Váša and O. Petřík. Section 3: Optimising perceived distortion in lossy encoding of dynamic meshes. Computer Graphics Forum, 30(5):1439–1449, August 2011. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

vanPelt:2014:BCB

[vPGL^{+}14] R. van Pelt, R. Gasteiger, K. Lawonn, M. Meuschke, and B. Preim. BioMedical I: Comparative blood flow visualization for cerebral aneurysm treatment assessment. Computer Graphics Forum, 33(3):131–140, June 2014. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

vanPelt:2012:MVV

[vPJtHRV12] R. F. P. van Pelt, S. S. A. M. Jacobs, B. M. ter Haar Romeny, and A. Vilanova. Medical visualization: Visualization of 4D blood-flow fields by spatiotemporal hierarchical clustering. Computer Graphics Forum, 31(3pt2):1065–1074, June 2012. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Vasa:2012:VPD

[VR12] Libor Váša and Jan Rus. Visualization and perception: Dihedral angle mesh error: a fast perception correlated distortion measure for fixed connectivity triangle meshes. Computer Graphics Forum, 31(5):1715–1724, August 2012. CODEN CG-FODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Vimont:2017:DGH

[VRBC17] Ulysse Vimont, Damien Rohmer, Antoine Begault, and Marie-Paule Cani. Deformation grammars: Hierarchical constraint preservation under deformation. Computer Graphics Forum, 36 (8):429–443, December 2017. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Vasa:2010:GDL

[VS10] Libor Váša and Václav Skala. Geometry-driven local neighbourhood based predictors for dynamic mesh compression. Computer Graphics Forum, 29(6):1921–1933, September 2010.

CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Vanhoey:2013:RFP

[VSG^{+}13] K. Vanhoey, B. Sauvage, O. Génevaux, F. Larue, and J.-M. Dischler. Robust fitting on poorly sampled data for surface light field rendering and image relighting. Computer Graphics Forum, 32(6):101–112, September 2013. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Varady:2016:CSM

[VSK16] Tamás Várady, Péter Salvi, and György Karikó. Curves & surfaces: a multi-sided Bézier patch with a simple control structure. Computer Graphics Forum, 35(2):307–317, May 2016. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

vonTycowicz:2011:CBC

[vTKP11] Christoph von Tycowicz, Felix Kälberer, and Konrad Polthier. Context-based coding of adaptive multiresolution meshes. Computer Graphics Forum, 30(8):2231–2245, December 2011. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Vergne:2011:NPR

[VVC^{+}11] Romain Vergne, David Vanderhaeghe, Jiazhou Chen, Pascal Barla, Xavier Granier, and Christophe Schlick. Nonphotorealistic rendering: Implicit brushes for stylized linebased rendering. Computer Graphics Forum, 30(2):513–522, April 2011. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

VanWelbergen:2010:RTA

[VVE^{+}10] H. Van Welbergen, B. J. H. Van Basten, A. Egges, Zs. M. Ruttkay, and M. H. Overmars. Real time animation of virtual humans: a trade-off between naturalness and control. Computer Graphics Forum, 29(8):2530–2554, December 2010. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Vasa:2016:DPM

[VVP^{+}16] L. Váša, P. Vaněček, M. Prantl, V. Skorkovská, P. Martínek, and I. Kolingerová. Differential properties: Mesh statistics for robust curvature estimation. Computer Graphics Forum, 35(5): 271–280, August 2016. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Velinov:2018:RTR

[VWH18] Z. Velinov, S. Werner, and M. B. Hullin. Real-time rendering of wave-optical effects on scratched surfaces. Computer Graphics Forum, 37(2):123–134, May 2018. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Woodring:2011:HLS

[WAF^{+}11] J. Woodring, J. Ahrens, J. Figg, J. Wendelberger, S. Habib, and K. Heitmann. Hierarchies and large-scale visualization: Insitu sampling of a large-scale particle simulation for interactive visualization and analysis. Computer Graphics Forum, 30(3): 1151–1160, June 2011. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Wang:2019:ATD

[WBAI19] Xiyao Wang, Lonni Besançon, Mehdi Ammi, and Tobias Isenberg. Augmenting tactile 3D data navigation with pressure sensing. Computer Graphics Forum, 38(3):635–647, June 2019. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Weber:2011:SCV

[WBCGH11] O. Weber, M. Ben-Chen, C. Gotsman, and K. Hormann. Section 6: a complex view of barycentric mappings. Computer Graphics Forum, 30(5):1533–1542, August 2011. CODEN CG-FODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Wunderlich:2017:VDU

[WBFvL17] M. Wunderlich, K. Ballweg, G. Fuchs, and T. von Landesberger. Visualization of delay uncertainty and its impact on train trip planning: a design study. Computer Graphics Forum, 36(3):317–328, June 2017. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Wood:2018:GFA

[WBM^{+}18] Erroll Wood, Tadas Baltrušaitis, Louis-Philippe Morency, Peter Robinson, and Andreas Bulling. GazeDirector: Fully articulated eye gaze redirection in video. Computer Graphics Forum, 37(2):217–225, May 2018. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Wiebel:2011:REW

[WBP11] Alexander Wiebel, Charl Botha, and Bernhard Preim. Reports: Eurographics Workshops VCBM 2008 and 2010. Computer Graphics Forum, 30(1):231–232, March 2011. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Weber:2013:EGD

[WBS^{+}13] Daniel Weber, Jan Bender, Markus Schnoes, André Stork, and Dieter Fellner. Efficient GPU data structures and methods to solve sparse linear systems in dynamics applications. Computer Graphics Forum, 32(1):16–26, February 2013. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Wenner:2013:SSM

[WBSH^{+}13] Simon Wenner, Jean-Charles Bazin, Alexander Sorkine-Hornung, Changil Kim, and Markus Gross. Synthesis: Scalable music: Automatic music retargeting and synthesis. Computer Graphics Forum, 32(2pt3):345–354, May 2013. CODEN CG-FODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Wiewel:2019:LSP

[WBT19] S. Wiewel, M. Becher, and N. Thuerey. Latent space physics: Towards learning the temporal evolution of fluid flow. Computer Graphics Forum, 38(2):71–82, May 2019. CODEN CG-FODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Wong:2014:CSC

[WC14] Sai-Keung Wong and Yu-Chun Cheng. Continuous self-collision detection for deformable surfaces interacting with solid models. Computer Graphics Forum, 33(6):143–153, September 2014. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Wong:2016:PAM

[WC16] Sai-Keung Wong and Kai-Chun Chen. A procedural approach to modelling virtual climbing plants with tendrils. Computer Graphics Forum, 35(8):5–18, December 2016. CODEN CG-FODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Wei:2015:VAC

[WCB15] Tzu-Hsuan Wei, Chun-Ming Chen, and Ayan Biswas. Volume analysis and classification: Efficient local histogram searching via bitmap indexing. Computer Graphics Forum, 34(3):81–90, June 2015. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Webanck:2018:PC

[WCGG18] A. Webanck, Y. Cortial, E. Guérin, and E. Galin. Procedural cloudscapes. Computer Graphics Forum, 37(2):431–442, May 2018. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Wu:2015:SVE

[WCH^{+}15] Feiran Wu, Guoning Chen, Jin Huang, Yubo Tao, and Wei Chen. Simulation and visualization: EasyXplorer: A flexible visual exploration approach for multivariate spatial data. Computer Graphics Forum, 34(7):163–172, October 2015. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Wong:2015:RSI

[WCM15] Yu-Shiang Wong, Hung-Kuo Chu, and Niloy J. Mitra. Reconstruction: SmartAnnotator an interactive tool for annotating indoor RGBD images. Computer Graphics Forum, 34(2):447– 457, May 2015. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Wang:2015:SPS

[WCT^{+}15] Weiming Wang, Haiyuan Chao, Jing Tong, Zhouwang Yang, Xin Tong, Hang Li, Xiuping Liu, and Ligang Liu. Saliencypreserving slicing optimization for effective 3D printing. Computer Graphics Forum, 34(6):148–160, September 2015. CO-DEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Weng:2013:GSR

[WCX^{+}13] Yanlin Weng, Menglei Chai, Weiwei Xu, Yiying Tong, and Kun Zhou. Geometry (session 2): As-rigid-as-possible distance field metamorphosis. Computer Graphics Forum, 32(7):381–389, October 2013. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Weiss:2011:SDH

[WD11] Kenneth Weiss and Leila De Floriani. Simplex and diamond hierarchies: Models and applications. Computer Graphics Forum, 30(8):2127–2155, December 2011. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Winkler:2010:SDA

[WDAH10] T. Winkler, J. Drieseberg, M. Alexa, and K. Hormann. Shape deformation and animation: Multi-scale geometry interpolation. Computer Graphics Forum, 29(2):309–318, May 2010. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Wang:2010:AIV

[WDC^{+}10] Xiaoyu Wang, Wenwen Dou, Shen-En Chen, William Ribarsky, and Remco Chang. Applications: an interactive visual analytics system for bridge management. Computer Graphics Forum, 29 (3):1033–1042, June 2010. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Wu:2013:CBC

[WDK^{+}13] Fuzhang Wu, Weiming Dong, Yan Kong, Xing Mei, Jean-Claude Paul, and Xiaopeng Zhang. Content-based colour transfer. Computer Graphics Forum, 32(1):190–203, February 2013. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Wang:2012:GRS

[WDM^{+}12] X. Wang, W. Dou, Z. Ma, J. Villalobos, Y. Chen, T. Kraft, and W. Ribarsky. Graphs and relationships: I-SI: Scalable architecture for analyzing latent topical-level information from social media data. Computer Graphics Forum, 31(3pt4):1275– 1284, June 2012. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Wu:2011:SAS

[WDR11] Hongzhi Wu, Julie Dorsey, and Holly Rushmeier. Surface appearance: a sparse parametric mixture model for BTF compression, editing and rendering. Computer Graphics Forum, 30 (2):465–473, April 2011. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Weller:2017:KPC

[WDZ17] René Weller, Nicole Debowski, and Gabriel Zachmann. kDet: Parallel constant time collision detection for polygonal objects. Computer Graphics Forum, 36(2):131–141, May 2017. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Wang:2017:SLB

[WESW17] Z. Wang, J. Martinez Esturo, H.-P. Seidel, and T. Weinkauf. Stream line-based pattern search in flows. Computer Graphics Forum, 36(8):7–18, December 2017. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Wu:2018:MBL

[WFLW18] Wenming Wu, Lubin Fan, Ligang Liu, and Peter Wonka. MIQP-based layout design for building interiors. Computer Graphics Forum, 37(2):511–521, May 2018. CODEN CG-FODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Wang:2015:FVV

[WFZ^{+}15] Y. H. Wang, C. R. Fan, J. Zhang, T. Niu, S. Zhang, and J. R. Jiang. Forecast verification and visualization based on Gaussian mixture model co-estimation. Computer Graphics Forum, 34 (6):99–110, September 2015. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Ward:2011:TDV

[WG11] Matthew O. Ward and Zhenyu Guo. Temporal data visualization: Visual exploration of time-series data with shape space projections. Computer Graphics Forum, 30(3):701–710, June 2011. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Weiss:2012:LRS

[WG12] Maayan Weiss and Thorsten Grosch. Lighting and rendering: Stochastic progressive photon mapping for dynamic scenes. Computer Graphics Forum, 31(2pt3):719–726, May 2012. CO-DEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Wang:2018:KFM

[WGBS18] L. Wang, A. Gehre, M. M. Bronstein, and J. Solomon. Kernel functional maps. Computer Graphics Forum, 37(5):27–36, August 2018. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Wei:2019:MDC

[WGH^{+}19] M. Wei, X. Guo, J. Huang, H. Xie, H. Zong, R. Kwan, F. L. Wang, and J. Qin. Mesh defiltering via cascaded geometry recovery. Computer Graphics Forum, 38(7):591–605, October 2019. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Wolski:2019:STR

[WGK^{+}19] K. Wolski, D. Giunchi, S. Kinuwaki, P. Didyk, K. Myszkowski, A. Steed, and R. K. Mantiuk. Selecting texture resolution using a task-specific visibility metric. Computer Graphics Forum, 38 (7):685–696, October 2019. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Wolinski:2014:AIM

[WGO^{+}14] D. Wolinski, S. J. Guy, A.-H. Olivier, M. Lin, D. Manocha, and J. Pettré. Animation II (motion data processing): Parameter estimation and comparative evaluation of crowd simulations. Computer Graphics Forum, 33(2):303–312, May 2014. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Weinkauf:2010:VRV

[WGS10] Tino Weinkauf, Yotam Gingold, and Olga Sorkine. Volume reconstruction and visualization: Topology-based smoothing of 2D scalar fields with C^{1}-continuity. Computer Graphics Forum, 29(3):1221–1230, June 2010. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Wang:2019:HHC

[WGW^{+}19] Zihan Wang, Neng Gao, Xin Wang, Ji Xiang, Daren Zha, and Linghui Li. HidingGAN: High capacity information hiding with generative adversarial network. Computer Graphics Forum, 38 (7):393–401, October 2019. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Wan:2017:UFV

[WH17a] Y. Wan and C. Hansen. Uncertainty footprint: Visualization of nonuniform behavior of iterative algorithms applied to 4D cell tracking. Computer Graphics Forum, 36(3):479–489, June 2017. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Wang:2017:AGG

[WH17b] Hui Wang and Hui Huang. Analyzing geometries: Group representation of global intrinsic symmetries. Computer Graphics Forum, 36(7):51–61, October 2017. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Wang:2013:GIF

[WHB^{+}13] Beibei Wang, Jing Huang, Bert Buchholz, Xiangxu Meng, and Tamy Boubekeur. Global illumination: Factorized point based global illumination. Computer Graphics Forum, 32(4):117–123, July 2013. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Walny:2015:EDE

[WHC15] J. Walny, S. Huron, and S. Carpendale. Evaluation and design: an exploratory study of data sketching for visual representation. Computer Graphics Forum, 34(3):231–240, June 2015. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Weber:2017:MVN

[WHD17] Pascal Weber, Johannes Hanika, and Carsten Dachsbacher. Multiple vertex next event estimation for lighting in dense, forward-scattering media. Computer Graphics Forum, 36(2): 21–30, May 2017. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Wetzstein:2010:OIP

[WHL10] Gordon Wetzstein, Wolfgang Heidrich, and David Luebke. Optical image processing using light modulation displays. Computer Graphics Forum, 29(6):1934–1944, September 2010. CO-DEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Wang:2019:SMF

[WHLW19] Zhenni Wang, Tze Yui Ho, Chi Sing Leung, and Eric. W. M. Wong. Seamless mipmap filtering for dual paraboloid maps. Computer Graphics Forum, 38(7):437–448, October 2019. CO-DEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Williams:2011:SDV

[WHP^{+}11] S. Williams, M. Hecht, M. Petersen, R. Strelitz, M. Maltrud, J. Ahrens, M. Hlawitschka, and B. Hamann. Spatiotemporal data visualization: Visualization and analysis of eddies in a global ocean simulation. Computer Graphics Forum, 30(3):991– 1000, June 2011. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Wang:2018:TEC

[WHS^{+}18] Yong Wang, Hammad Haleem, Conglei Shi, Yanhong Wu, Xun Zhao, Siwei Fu, and Huamin Qu. Towards easy comparison of local businesses using online reviews. Computer Graphics Forum, 37(3):63–74, June 2018. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Weinkauf:2012:VSA

[WHT12] T. Weinkauf, H.-C. Hege, and H. Theisel. Visualization and simulation: Advected tangent curves: a general scheme for characteristic curves of flow fields. Computer Graphics Forum, 31(2pt4):825–834, May 2012. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Wang:2016:HAR

[WHWB16] L. Wang, F. Hétroy-Wheeler, and E. Boyer. A hierarchical approach for regular centroidal Voronoi tessellations. Computer Graphics Forum, 35(1):152–165, February 2016. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Weiss:2013:TPD

[WIFD13] Kenneth Weiss, Federico Iuricich, Riccardo Fellegara, and Leila De Floriani. Topology: a primal/dual representation for discrete Morse complexes on tetrahedral meshes. Computer Graphics Forum, 32(3pt3):361–370, June 2013. CODEN CG-FODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Wetzstein:2011:CPI

[WILH11] Gordon Wetzstein, Ivo Ihrke, Douglas Lanman, and Wolfgang Heidrich. Computational plenoptic imaging. Computer Graphics Forum, 30(8):2397–2426, December 2011. CODEN CG-FODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Wang:2013:MOA

[WJB^{+}13] J. Wang, C. Jiang, P. Bompas, J. Wallner, and H. Pottmann. Mesh optimization and applications: Discrete line congruences for shading and lighting. Computer Graphics Forum, 32(5):53– 62, August 2013. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Wang:2014:INA

[WJDZ14] Xinjie Wang, Xiaogang Jin, Zhigang Deng, and Linling Zhou. Inherent noise-aware insect swarm simulation. Computer Graphics Forum, 33(6):51–62, September 2014. CODEN CG-FODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Weber:2016:CNP

[WJG^{+}16] Y. Weber, V. Jolivet, G. Gilet, K. Nanko, and D. Ghazanfarpour. Capturing nature: a phenomenological model for throughfall rendering in real-time. Computer Graphics Forum, 35(4):13–23, July 2016. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Wang:2012:CCS

[WK12a] He Wang and Taku Komura. Crowds, cloths, and shape deformation: Manipulation of flexible objects by geodesic control. Computer Graphics Forum, 31(2pt2):499–508, May 2012. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Wang:2012:VVI

[WK12b] Lei Wang and Arie Kaufman. Volume visualization: Importance driven automatic color design for direct volume rendering. Computer Graphics Forum, 31(3pt4):1305–1314, June 2012. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Welch:2017:MSD

[WK17] E. Welch and S. Kobourov. Measuring symmetry in drawings of graphs. Computer Graphics Forum, 36(3):341–351, June 2017. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Weiler:2018:PCI

[WKBB18] Marcel Weiler, Dan Koschier, Magnus Brand, and Jan Bender. A physically consistent implicit viscosity solver for SPH fluids. Computer Graphics Forum, 37(2):145–155, May 2018. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Wang:2015:GID

[WKM15] Tuanfeng Y. Wang, Pushmeet Kohli, and Niloy J. Mitra. Geometry and images: Dynamic SfM: Detecting scene changes from image pairs. Computer Graphics Forum, 34(5):177–189, August 2015. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Waser:2014:GVM

[WKS^{+}14] J. Waser, A. Konev, B. Sadransky, Z. Horváth, H. Ribičić, R. Carnecky, P. Kluding, and B. Schindler. Geographic visualization: Many plans: Multidimensional ensembles for visual decision support in flood management. Computer Graphics Forum, 33(3):281–290, June 2014. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Wang:2010:MMB

[WL10] S. F. Wang and S. H. Lai. Modeling: Manifold-based 3D face caricature generation with individualized facial feature extraction. Computer Graphics Forum, 29(7):2161–2168, September 2010. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Wu:2012:IHP

[WLI^{+}12] Di Wu, Yebin Liu, Ivo Ihrke, Qionghai Dai, and Christian Theobalt. Imaging and hair: Performance capture of highspeed motion using staggered multi-view recording. Computer Graphics Forum, 31(7pt1):2019–2028, September 2012. CO-DEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Wang:2018:SMG

[WLJ^{+}18] Guan Wang, Hamid Laga, Jinyuan Jia, Ning Xie, and Hedi Tabia. Statistical modeling of the 3D geometry and topology of botanical trees. Computer Graphics Forum, 37(5):185–198, August 2018. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Wang:2017:ICD

[WLL^{+}17] Fei Wang, Shujin Lin, Xiaonan Luo, Hefeng Wu, Ruomei Wang, and Fan Zhou. Interaction and creation: a data-driven approach for sketch-based 3D shape retrieval via similar drawingstyle recommendation. Computer Graphics Forum, 36(7):157– 166, October 2017. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Wenger:2013:IPS

[WLM13] Stephan Wenger, Dirk Lorenz, and Marcus Magnor. Image processing (session 2): Fast image-based modeling of astronomical nebulae. Computer Graphics Forum, 32(7):93–100, October 2013. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Wang:2017:LDS

[WLN^{+}17] Yunhai Wang, Jingting Li, Feiping Nie, Holger Theisel, Minglun Gong, and Dirk J. Lehmann. Linear discriminative star coordinates for exploring class and cluster separation of high dimensional data. Computer Graphics Forum, 36(3):401– 410, June 2017. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Wei:2013:SEI

[WLS13] Tzu-Hsuan Wei, Teng-Yok Lee, and Han-Wei Shen. Surfaces: Evaluating isosurfaces with level-set-based information maps. Computer Graphics Forum, 32(3pt1):1–10, June 2013. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Wang:2012:LSR

[WLT12] Y. Wang, B. Liu, and Y. Tong. Linear surface reconstruction from discrete fundamental forms on triangle meshes. Computer Graphics Forum, 31(8):2277–2287, December 2012. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Wang:2017:ERS

[WLT^{+}17] Tongtong Wang, Zhihua Liu, Min Tang, Ruofeng Tong, and Dinesh Manocha. Efficient and reliable self-collision culling using unprojected normal cones. Computer Graphics Forum, 36(8): 487–498, December 2017. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Wang:2019:IGA

[WLX19] Yili Wang, Yifan Liu, and Kun Xu. An improved geometric approach for palette-based image decomposition and recoloring. Computer Graphics Forum, 38(7):11–22, October 2019. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Wong:2013:MBB

[WLZ13] T. H. Wong, G. Leach, and F. Zambetta. Modelling bending behaviour in cloth simulation using hysteresis. Computer Graphics Forum, 32(8):183–194, December 2013. CODEN CG-FODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Wang:2017:AMA

[WLZH17] Yupan Wang, Guiqing Li, Zhichao Zeng, and Huayun He. Articulated-motion-aware sparse localized decomposition. Computer Graphics Forum, 36(8):247–259, December 2017. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Wang:2018:VFM

[WLZT18] Y. Wang, B. Liu, K. Zhou, and Y. Tong. Vector field map representation for near conformal surface correspondence. Computer Graphics Forum, 37(6):72–83, September 2018. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Wei:2019:SIR

[WLZX19] Jinjiang Wei, Chengjiang Long, Hua Zou, and Chunxia Xiao. Shadow inpainting and removal using generative adversarial networks with slice convolutions. Computer Graphics Forum, 38(7):381–392, October 2019. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Wang:2015:RTR

[WMB15] Beibei Wang, Xiangxu Meng, and Tamy Boubekeur. Real-time rendering and filtering: Wavelet point-based global illumination. Computer Graphics Forum, 34(4):143–153, July 2015. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Weber:2015:FFC

[WMRSF15] Daniel Weber, Johannes Mueller-Roemer, André Stork, and Dieter Fellner. Fluids & flows: a cut-cell geometric multigrid Poisson solver for fluid simulation. Computer Graphics Forum, 34 (2):481–491, May 2015. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Wang:2019:BRI

[WMXC19] Wencheng Wang, Junhui Ma, Panpan Xu, and Yiyao Chu. Book review: Intrinsic Symmetry Detection on 3D Models with Skeleton-guided Combination of Extrinsic Symmetries. Computer Graphics Forum, 38(7):617–628, October 2019. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Weber:2012:SAC

[WMZ12] Ofir Weber, Ashish Myles, and Denis Zorin. Shape analysis: Computing extremal quasiconformal maps. Computer Graphics Forum, 31(5):1679–1689, August 2012. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Wilkie:2014:SWR

[WND^{+}14] A. Wilkie, S. Nawaz, M. Droske, A. Weidlich, and J. Hanika. Spectral and wave rendering: Hero wavelength spectral sampling. Computer Graphics Forum, 33(4):123–131, July 2014. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Whited:2010:NPR

[WNS^{+}10] Brian Whited, Gioacchino Noris, Maryann Simmons, Robert W. Sumner, Markus Gross, and Jarek Rossignac. Non-photorealistic rendering: BetweenIT: An interactive tool for tight inbetweening. Computer Graphics Forum, 29(2):605–614, May 2010. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Wan:2013:VSB

[WOH13] Y. Wan, H. Otsuna, and C. Hansen. Volumes: Synthetic brainbows. Computer Graphics Forum, 32(3pt4):471–480, June 2013. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Weber:2012:BC

[WPG12] Ofir Weber, Roi Poranne, and Craig Gotsman. Biharmonic coordinates. Computer Graphics Forum, 31(8):2409–2422, December 2012. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Wolff:2018:PS

[WPGSH18] Katja Wolff, Roi Poranne, Oliver Glauser, and Olga Sorkine-Hornung. Packable springs. Computer Graphics Forum, 37 (2):251–262, May 2018. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Williams:2012:FVI

[WPH^{+}12] S. Williams, M. Petersen, M. Hecht, M. Maltrud, J. Patchett, J. Ahrens, and B. Hamann. Flow visualization: Interface exchange as an indicator for eddy heat transport. Computer Graphics Forum, 31(3pt3):1125–1134, June 2012. CODEN CG-FODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Wun:2016:CGC

[WPHC16] Tiffany Wun, Jennifer Payne, Samuel Huron, and Sheelagh Carpendale. Charts and glyphs: Comparing bar chart authoring with Microsoft Excel and tangible tiles. Computer Graphics Forum, 35(3):111–120, June 2016. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Wampler:2013:AAL

[WPP13] Kevin Wampler, Jovan Popović, and Zoran Popović. Animation: Animal locomotion controllers from scratch. Computer Graphics Forum, 32(2pt2):153–162, May 2013. CODEN CG-FODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Wu:2011:TDV

[WPW^{+}11] Yingcai Wu, Thomas Provan, Furu Wei, Shixia Liu, and Kwan-Liu Ma. Text and document visualization: Semantic-preserving word clouds by seam carving. Computer Graphics Forum, 30 (3):741–750, June 2011. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Wang:2019:LGM

[WPY^{+}19] He Wang, Sören Pirk, Ersin Yumer, Vladimir G. Kim, Ozan Sener, Srinath Sridhar, and Leonidas J. Guibas. Learning a generative model for multi-step human-object interactions from videos. Computer Graphics Forum, 38(2):367–378, May 2019. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Weier:2016:RTA

[WRK^{+}16] Martin Weier, Thorsten Roth, Ernst Kruijff, André Hinkenjann, Arsène Pérard-Gayot, Philipp Slusallek, and Yongmin Li. Ray tracing/appearance capture: Foveated real-time ray tracing for head-mounted displays. Computer Graphics Forum, 35 (7):289–298, October 2016. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Wang:2013:VFV

[WRS^{+}13] B. Wang, P. Rosen, P. Skraba, H. Bhatia, and V. Pascucci. Vector fields: Visualizing robustness of critical points for 2D timevarying vector fields. Computer Graphics Forum, 32(3pt2):221– 230, June 2013. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Wang:2013:JTT

[WSCP13] Lili Wang, Yulong Shi, Yi Chen, and Voicu Popescu. Justin-time texture synthesis. Computer Graphics Forum, 32(1): 126–138, February 2013. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Wesslen:2019:IEV

[WSK^{+}19] R. Wesslen, S. Santhanam, A. Karduni, I. Cho, S. Shaikh, and W. Dou. Investigating effects of visual anchors on decisionmaking about misinformation. Computer Graphics Forum, 38 (3):161–171, June 2019. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Weier:2017:PDA

[WSR^{+}17] M. Weier, M. Stengel, T. Roth, P. Didyk, E. Eisemann, M. Eisemann, S. Grogorick, A. Hinkenjann, E. Kruijff, M. Magnor, K. Myszkowski, and P. Slusallek. Perception-driven accelerated rendering. Computer Graphics Forum, 36(2):611–643, May 2017. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Windheuser:2011:SLS

[WSSC11] Thomas Windheuser, Ulrich Schlickwei, Frank R. Schimdt, and Daniel Cremers. Section 4: Large-scale integer linear programming for orientation preserving 3 D shape matching. Computer Graphics Forum, 30(5):1471–1480, August 2011. CODEN CG-FODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Wong:2011:NPR

[WT11] Fernando J. Wong and Shigeo Takahashi. Non-photorealistic rendering & GPU computing: a graph-based approach to continuous line illustrations with variable levels of detail. Computer Graphics Forum, 30(7):1931–1939, September 2011. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Wang:2017:GSV

[WT17] Chaoli Wang and Jun Tao. Graphs in scientific visualization: a survey. Computer Graphics Forum, 36(1):263–287, January 2017. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Wu:2013:STS

[WTH^{+}13] Hsiang-Yun Wu, Shigeo Takahashi, Daichi Hirono, Masatoshi Arikawa, Chun-Cheng Lin, and Hsu-Chun Yen. Space and time: Spatially efficient design of annotated metro maps. Computer Graphics Forum, 32(3pt3):261–270, June 2013. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Wang:2015:VRE

[WTL15] Qichao Wang, Yubo Tao, and Hai Lin. Volume rendering: Edge-aware volume smoothing using L_{0} gradient minimization. Computer Graphics Forum, 34(3):131–140, June 2015. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Wu:2013:EPG

[WTLL13] Yi-Chian Wu, Yu-Ting Tsai, Wen-Chieh Lin, and Wen-Hsin Li. Expressiveness and perception: Generating pointillism paintings based on Seurat’s color composition. Computer Graphics Forum, 32(4):153–162, July 2013. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Wu:2012:GVT

[WTLY12] Hsiang-Yun Wu, Shigeo Takahashi, Chun-Cheng Lin, and Hsu-Chun Yen. Geospatial visualization: Travel-route-centered metro map layout and annotation. Computer Graphics Forum, 31(3pt1):925–934, June 2012. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Wang:2018:EBB

[WTMT18] Xinlei Wang, Min Tang, Dinesh Manocha, and Ruofeng Tong. Efficient BVH-based collision detection scheme with ordering and restructuring. Computer Graphics Forum, 37(2):227–237, May 2018. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Wang:2015:CDT

[WTTM15] Zhendong Wang, Min Tang, Ruofeng Tong, and Dinesh Manocha. Collision detection: TightCCD: Efficient and robust continuous collision detection using tight error bounds. Computer Graphics Forum, 34(7):289–298, October 2015. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Wu:2018:FFS

[WTYH18] Kui Wu, Nghia Truong, Cem Yuksel, and Rama Hoetzlein. Fast fluid simulations with sparse volumes on the GPU. Computer Graphics Forum, 37(2):157–167, May 2018. CODEN CG-FODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Willems:2011:EEV

[WVV11] Niels Willems, Huub Van De Wetering, and Jarke J. Van Wijk. Evaluation: Evaluation of the visibility of vessel movement features in trajectory visualizations. Computer Graphics Forum, 30(3):801–810, June 2011. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Wilkie:2011:RAV

[WW11] A. Wilkie and A. Weidlich. Realistic appearance and volumes: a physically plausible model for light emission from glowing solid objects. Computer Graphics Forum, 30(4):1269–1276, June 2011. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Wang:2016:FBO

[WW16] L. Wang and E. Whiting. Fabrication: Buoyancy optimization for computational fabrication. Computer Graphics Forum, 35 (2):49–58, May 2016. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Wu:2015:SPB

[WWD15] Jun Wu, Rüdiger Westermann, and Christian Dick. A survey of physically based simulation of cuts in deformable bodies. Computer Graphics Forum, 34(6):161–187, September 2015. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Wang:2018:PMN

[WWF^{+}18] Zhendong Wang, Longhua Wu, Marco Fratarcangeli, Min Tang, and Huamin Wang. Parallel multigrid for nonlinear cloth simulation. Computer Graphics Forum, 37(7):131–141, October 2018. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Wang:2010:RRT

[WWH^{+}10] Yajun Wang, Jiaping Wang, Nicolas Holzschuch, Kartic Subr, Jun-Hai Yong, and Baining Guo. Rendering I: Real-time rendering of heterogeneous translucent objects with arbitrary shapes. Computer Graphics Forum, 29(2):497–506, May 2010. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Wu:2014:PDR

[WWH^{+}14] Xiaokun Wu, Michael Wand, Klaus Hildebrandt, Pushmeet Kohli, and Hans-Peter Seidel. Particles and deformation: Realtime symmetry-preserving deformation. Computer Graphics Forum, 33(7):229–238, October 2014. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Wang:2018:FGI

[WWH18] Beibei Wang, Lu Wang, and Nicolas Holzschuch. Fast global illumination with discrete stochastic microfacets using a filterable model. Computer Graphics Forum, 37(7):55–64, October 2018. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Weng:2013:IPS

[WWL^{+}13] Yanlin Weng, Lvdi Wang, Xiao Li, Menglei Chai, and Kun Zhou. Image processing (session 2): Hair interpolation for portrait morphing. Computer Graphics Forum, 32(7):79–84, October 2013. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Waldin:2019:CCM

[WWL^{+}19] N. Waldin, M. Waldner, M. Le Muzic, E. Gröller, D. S. Goodsell, L. Autin, A. J. Olson, and I. Viola. Cuttlefish: Color mapping for dynamic multi-scale visualizations. Computer Graphics Forum, 38(6):150–164, September 2019. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Wheatland:2015:SAR

[WWS^{+}15] Nkenge Wheatland, Yingying Wang, Huaguang Song, Michael Neff, Victor Zordan, and Sophie Jörg. State of the art reports: State of the art in hand and finger modeling and animation. Computer Graphics Forum, 34(2):735–760, May 2015. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Widmer:2016:RRD

[WWT^{+}16] S. Widmer, D. Wodniok, D. Thul, S. Guthe, and M. Goesele. Realistic rendering: Decoupled space and time sampling of motion and defocus blur for unified rendering of transparent and opaque objects. Computer Graphics Forum, 35(7):441–450, October 2016. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Waldin:2017:FOE

[WWV17] N. Waldin, M. Waldner, and I. Viola. Flicker observer effect: Guiding attention through high frequency flicker in images. Computer Graphics Forum, 36(2):467–476, May 2017. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Wang:2011:SAS

[WXL^{+}11] Y. Wang, K. Xu, J. Li, H. Zhang, A. Shamir, L. Liu, Z. Cheng, and Y. Xiong. Shape analysis: Symmetry hierarchy of manmade objects. Computer Graphics Forum, 30(2):287–296, April 2011. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Wang:2013:RCL

[WXL^{+}13] Jun Wang, Kai Xu, Ligang Liu, Junjie Cao, Shengjun Liu, Zeyun Yu, and Xianfeng David Gu. Reconstruction: Consolidation of low-quality point clouds from outdoor scenes. Computer Graphics Forum, 32(5):207–216, August 2013. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Wang:2016:MAM

[WXR^{+}16] J. Wang, Y. Xu, O. Remil, X. Xie, N. Ye, C. Yi, and M. Wei. Modeling: Automatic modeling of urban facades from raw Li-DAR point data. Computer Graphics Forum, 35(7):269–278, October 2016. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Wu:2018:CCM

[WXW18] Q. Wu, K. Xu, and J. Wang. Constructing 3D CSG models from 3D raw point clouds. Computer Graphics Forum, 37(5): 221–232, August 2018. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Wang:2013:EBE

[WYD^{+}13] C. Wang, Y. Yue, F. Dong, Y. Tao, X. Ma, G. Clapworthy, and X. Ye. Enhancing Bayesian estimators for removing camera shake. Computer Graphics Forum, 32(6):113–125, September 2013. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Wu:2017:ARN

[WYKR17] Lifan Wu, Ling-Qi Yan, Alexandr Kuznetsov, and Ravi Ramamoorthi. Adding and removing noise: Multiple axis-aligned filters for rendering of combined distribution effects. Computer Graphics Forum, 36(4):155–166, July 2017. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Wu:2013:FNP

[WYY13] Xiaoyue Wu, Xubo Yang, and Yang Yang. Fluids: a novel projection technique with detail capture and shape correction for smoke simulation. Computer Graphics Forum, 32(2pt4):389– 397, May 2013. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Wang:2014:VTS

[WYZB14] Rui Wang, Yinhui Yang, Hongxin Zhang, and Hujun Bao. Variational tree synthesis. Computer Graphics Forum, 33(8):82–94, December 2014. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Wang:2013:FPS

[WYZC13] J. Wang, Z. Yu, W. Zhu, and J. Cao. Feature-preserving surface reconstruction from unoriented, noisy point data. Computer Graphics Forum, 32(1):164–176, February 2013. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Wu:2015:AIA

[WZ15] Hongzhi Wu and Kun Zhou. AppFusion: Interactive appearance acquisition using a Kinect sensor. Computer Graphics Forum, 34(6):289–298, September 2015. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Wang:2011:VEO

[WZC^{+}11] Yunhai Wang, Jian Zhang, Wei Chen, Huai Zhang, and Xuebin Chi. Visualization: Efficient opacity specification based on feature visibilities in direct volume rendering. Computer Graphics Forum, 30(7):2117–2126, September 2011. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Wu:2015:SMM

[WZCF15] Xiaoqun Wu, Jianmin Zheng, Yiyu Cai, and Chi-Wing Fu. Shape and mesh: Mesh denoising using extended ROF model with L_{1} fidelity. Computer Graphics Forum, 34(7):35–45, October 2015. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Wei:2019:FAQ

[WZG^{+}19] Guangshun Wei, Yuanfeng Zhou, Xifeng Gao, Qian Ma, Shiqing Xin, and Ying He. Field-aligned quadrangulation for image vectorization. Computer Graphics Forum, 38(7):171–180, October 2019. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Wang:2013:IPS

[WZH13] Zhong-Qiang Wang, Lei Zhang, and Hua Huang. Image processing (session 4): Multiplane video stabilization. Computer Graphics Forum, 32(7):265–273, October 2013. CODEN CG-FODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Wang:2016:FIS

[WZK16] W. M. Wang, C. Zanni, and L. Kobbelt. Fabrication: Improved surface quality in 3D printing by optimizing the printing direction. Computer Graphics Forum, 35(2):59–70, May 2016. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Wang:2014:GGE

[WZKP14] Lili Wang, Shiheng Zhou, Wei Ke, and Voicu Popescu. GEARS: a general and efficient algorithm for rendering shadows. Computer Graphics Forum, 33(6):264–275, September 2014. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Wang:2012:VVA

[WZL^{+}12] Yunhai Wang, Jian Zhang, Dirk J. Lehmann, Holger Theisel, and Xuebin Chi. Volume visualization: Automating transfer function design with valley cell-based clustering of 2D density plots. Computer Graphics Forum, 31(3pt4):1295–1304, June 2012. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Wang:2017:CRS

[WZL^{+}17] Dong Wang, Changqing Zou, Guiqing Li, Chengying Gao, Zhuo Su, and Ping Tan. Coloring, rendering, and sampling: L_{0} gradient-preserving color transfer. Computer Graphics Forum, 36(7):93–103, October 2017. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Xue:2013:IPS

[XADR13] S. Xue, A. Agarwala, J. Dorsey, and H. Rushmeier. Image processing (session 4): Learning and applying color styles from feature films. Computer Graphics Forum, 32(7):255–264, October 2013. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Xu:2018:CCE

[XBL^{+}18] Shenyu Xu, Chris Bryan, Jianping Kelvin Li, Jian Zhao, and Kwan-Liu Ma. Chart constellations: Effective chart summarization for collaborative and multi-user analyses. Computer Graphics Forum, 37(3):75–86, June 2018. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Xue:2010:NPR

[XCDR10] Su Xue, Xuejin Chen, Julie Dorsey, and Holly Rushmeier. Nonphotorealistic rendering: Printed patterns for enhanced shape perception of papercraft models. Computer Graphics Forum, 29(2):625–634, May 2010. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Xu:2013:GVA

[XDC^{+}13] Panpan Xu, Fan Du, Nan Cao, Conglei Shi, Hong Zhou, and Huamin Qu. Graphs: Visual analysis of set relations in a graph. Computer Graphics Forum, 32(3pt1):61–70, June 2013. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Xue:2011:RMS

[XDR11] Su Xue, Julie Dorsey, and Holly Rushmeier. Reflectance and materials: Stone weathering in a photograph. Computer Graphics Forum, 30(4):1189–1196, June 2011. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Xu:2018:EPS

[XDY18] Qun-Ce Xu, Bailin Deng, and Yong-Liang Yang. Ellipsoid packing structures on freeform surfaces. Computer Graphics Forum, 37(7):87–95, October 2018. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Xu:2017:HMV

[XGDC17] Kaoji Xu, Xifeng Gao, Zhigang Deng, and Guoning Chen. Hexahedral meshing with varying element sizes. Computer Graphics Forum, 36(8):540–553, December 2017. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Xu:2018:RHM

[XHC^{+}18] Liyou Xu, Xiaowei He, Wei Chen, Sheng Li, and Guoping Wang. Reformulating hyperelastic materials with peridynamic modeling. Computer Graphics Forum, 37(7):121–130, October 2018. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Xu:2018:DVS

[XHW^{+}18] Sen-Zhe Xu, Jun Hu, Miao Wang, Tai-Jiang Mu, and Shi-Min Hu. Deep video stabilization using adversarial networks. Computer Graphics Forum, 37(7):267–276, October 2018. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Xu:2017:DDS

[XKHK17] Kai Xu, Vladimir G. Kim, Qixing Huang, and Evangelos Kalogerakis. Data-driven shape analysis and processing. Computer Graphics Forum, 36(1):101–132, January 2017. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Xiao:2010:IVE

[XL10] Chunxia Xiao and Meng Liu. Image and video I: Efficient meanshift clustering using Gaussian KD-tree. Computer Graphics Forum, 29(7):2065–2073, September 2010. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Xiao:2013:CSR

[XLH^{+}13] Yi Xiao, Chi Sing Leung, Tze Yui Ho, Liang Wan, and Tien Tsing Wong. Concentric spherical representation for omnidirectional soft shadow. Computer Graphics Forum, 32(6):201–213, September 2013. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Xu:2010:ALL

[XLL^{+}10] Mingliang Xu, Huansen Li, Pei Lv, Wenzhi Chen, Gengdai Liu, Pengyu Zhu, and Zhigeng Pan. Animation: L4RW: Lazinessbased realistic real-time responsive rebalance in walking. Computer Graphics Forum, 29(7):2187–2196, September 2010. CO-DEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Xu:2014:SPS

[XLS^{+}14] Chunxu Xu, Yong-Jin Liu, Qian Sun, Jinyan Li, and Ying He. Surfaces: Polyline-sourced geodesic Voronoi diagrams on triangle meshes. Computer Graphics Forum, 33(7):161–170, October 2014. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Xu:2015:SMP

[XM15] Ling Xu and David Mould. Shape and mesh: Procedural tree modeling with guiding vectors. Computer Graphics Forum, 34 (7):47–56, October 2015. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Xu:2014:HLD

[XSE14] X. Xu, E. Sakhaee, and A. Entezari. High level direct volume rendering: Volumetric data reduction in a compressed sensing framework. Computer Graphics Forum, 33(3):111–120, June 2014. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Xu:2013:IPS

[XSQ13] X. Xu, H. S. Seah, and C. K. Quah. Image processing (session 5): Animated 3D line drawings with temporal coherence. Computer Graphics Forum, 32(7):285–294, October 2013. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Xu:2014:SST

[XSX^{+}14] Weiwei Xu, Zhouxu Shi, Mingliang Xu, Kun Zhou, Jingdong Wang, Bin Zhou, Jinrong Wang, and Zhenming Yuan. Segmentation and structure: Transductive 3D shape segmentation using sparse reconstruction. Computer Graphics Forum, 33(5): 107–115, August 2014. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Xiao:2013:FSR

[XSXM13] Chunxia Xiao, Ruiyun She, Donglin Xiao, and Kwan-Liu Ma. Fast shadow removal using adaptive multi-scale illumination transfer. Computer Graphics Forum, 32(8):207–218, December 2013. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Xiao:2019:GCE

[XTZ19] J. Y. Xiao, J. T. Tang, and X. Y. Zhang. A generalized cubemap for encoding 360^{o} VR videos using polynomial approximation. Computer Graphics Forum, 38(7):359–367, October 2019. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Xu:2015:RRR

[XWB15] Chao Xu, Rui Wang, and Hujun Bao. Rendering: Realtime rendering glossy to glossy reflections in screen space. Computer Graphics Forum, 34(7):57–66, October 2015. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Xia:2013:ASO

[XWG^{+}13] Yang Xia, Jiangqin Wu, Pengcheng Gao, Yuan Lin, and Tianjiao Mao. Animation (session 1): Ontology-based model for Chinese calligraphy synthesis. Computer Graphics Forum, 32 (7):11–20, October 2013. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Xiao:2015:OBG

[XWL^{+}15] Yi Xiao, Liang Wan, Chi Sing Leung, Yu-Kun Lai, and Tien-Tsin Wong. Optimization-based gradient mesh colour transfer. Computer Graphics Forum, 34(6):123–134, September 2015. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Xu:2015:CCF

[XWY^{+}15] Mingliang Xu, Yunpeng Wu, Yangdong Ye, Illes Farkas, Hao Jiang, and Zhigang Deng. Collective crowd formation transform with mutual information-based runtime feedback. Computer Graphics Forum, 34(1):60–73, February 2015. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Xiao:2019:CBF

[XWY19] Xiangyun Xiao, Hui Wang, and Xubo Yang. A CNN-based flow correction method for fast preview. Computer Graphics Forum, 38(2):431–440, May 2019. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Xu:2017:BIR

[XWZB17] Chao Xu, Rui Wang, Shuang Zhao, and Hujun Bao. BRDFs and illumination: Real-time linear BRDF MIP-mapping. Computer Graphics Forum, 36(4):27–34, July 2017. CODEN CG-FODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Xie:2014:SSS

[XXLX14] Zhige Xie, Kai Xu, Ligang Liu, and Yueshan Xiong. Segmentation and structure: 3D shape segmentation and labeling via extreme learning machine. Computer Graphics Forum, 33(5):85– 95, August 2014. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Xie:2013:SDC

[XXM^{+}13] Xiaohua Xie, Kai Xu, Niloy J. Mitra, Daniel Cohen-Or, Wenyong Gong, Qi Su, and Baoquan Chen. Sketch-to-Design: Context-based part assembly. Computer Graphics Forum, 32 (8):233–245, December 2013. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Xie:2015:SMP

[XXS^{+}15] Zhige Xie, Kai Xu, Wen Shan, Ligang Liu, Yueshan Xiong, and Hui Huang. Shape and mesh: Projective feature learning for 3D shapes with multi-view depth images. Computer Graphics Forum, 34(7):1–11, October 2015. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Xu:2018:OGN

[XXY^{+}18] Jiamin Xu, Weiwei Xu, Yin Yang, Zhigang Deng, and Hujun Bao. Online global Non-rigid registration for 3D object reconstruction using Consumer-level depth cameras. Computer Graphics Forum, 37(7):1–12, October 2018. CODEN CG-FODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Xiao:2013:IPS

[XXZC13] Chunxia Xiao, Donglin Xiao, Ling Zhang, and Lin Chen. Image processing (session 6): Efficient shadow removal using subregion matching illumination transfer. Computer Graphics Forum, 32(7):421–430, October 2013. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Xu:2019:ASR

[XYLY19] Qun-Ce Xu, Dong-Ming Yan, Wenbin Li, and Yong-Liang Yang. Anisotropic surface remeshing without obtuse angles. Computer Graphics Forum, 38(7):755–763, October 2019. CO-DEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Xie:2013:VIR

[XYM13] Jinrong Xie, Hongfeng Yu, and Kwan-Liu Ma. Volumes: Interactive ray casting of geodesic grids. Computer Graphics Forum, 32(3pt4):481–490, June 2013. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Xing:2013:RSL

[XZP^{+}13] Guanyu Xing, Xuehong Zhou, Qunsheng Peng, Yanli Liu, and Xueying Qin. Rendering (session 1): Lighting simulation of augmented outdoor scene based on a legacy photograph. Computer Graphics Forum, 32(7):101–110, October 2013. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Yang:2018:APS

[YB18] Y. Yang and C. Barnes. Approximate program smoothing using mean-variance statistics, with application to procedural shader bandlimiting. Computer Graphics Forum, 37(2):443–454, May 2018. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Yim:2012:VSC

[YBK^{+}12] D. Yim, G. V. G. Baranoski, B. W. Kimmel, T. F. Chen, and E. Miranda. Visualization and simulation: a cell-based light interaction model for human blood. Computer Graphics Forum, 31(2pt4):845–854, May 2012. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Yoshizawa:2010:FGB

[YBY10] Shin Yoshizawa, Alexander Belyaev, and Hideo Yokota. Fast Gauss bilateral filtering. Computer Graphics Forum, 29(1):60– 74, March 2010. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Yao:2019:UNN

[YC19] L. Yao and A. Chyau. A unified neural network for panoptic segmentation. Computer Graphics Forum, 38(7):461–468, October 2019. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Yang:2017:RGB

[YCL^{+}17] Sheng Yang, Kang Chen, Minghua Liu, Hongbo Fu, and Shi-Min Hu. Reconstruction and generation based on RGBD images: Saliency-aware real-time volumetric fusion for object reconstruction. Computer Graphics Forum, 36(7):167–174, October 2017. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Yao:2017:FDM

[YCXW17] Miaojun Yao, Zhili Chen, Weiwei Xu, and Huamin Wang. Fabrication and design: Modeling, evaluation and optimization of interlocking Shell pieces. Computer Graphics Forum, 36(7):1– 13, October 2017. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Yang:2010:RVS

[YDF^{+}10] Baoguang Yang, Zhao Dong, Jieqing Feng, Hans-Peter Seidel, and Jan Kautz. Rendering: Variance soft shadow mapping. Computer Graphics Forum, 29(7):2127–2134, September 2010. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Ye:2019:ICD

[YDP19] Wenjie Ye, Yue Dong, and Pieter Peers. Interactive curation of datasets for training and refining generative models. Computer Graphics Forum, 38(7):369–380, October 2019. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Ye:2018:UDF

[YDT^{+}18] Zi Ye, Olga Diamanti, Chengcheng Tang, Leonidas Guibas, and Tim Hoffmann. A unified discrete framework for intrinsic and extrinsic Dirac operators for geometry processing. Computer Graphics Forum, 37(5):93–106, August 2018. CODEN CG-FODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Yang:2019:CSP

[YFL19] Yang Yang, Xiao-Ming Fu, and Ligang Liu. Computing surface PolyCube-Maps by constrained voxelization. Computer Graphics Forum, 38(7):299–309, October 2019. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Yang:2012:SPM

[YFW12] Wenwu Yang, Jieqing Feng, and Xun Wang. Shape processing and modeling: Structure preserving manipulation and interpolation for multi-element 2D shapes. Computer Graphics Forum, 31(7pt2):2249–2258, September 2012. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Yin:2011:NQP

[YFWR11] Xuetao Yin, John Femiani, Peter Wonka, and Anshuman Razdan. A new QEM for parametrization of raster images. Computer Graphics Forum, 30(8):2440–2451, December 2011. CO-DEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Yan:2014:MNF

[YGCO^{+}14] Feilong Yan, Minglun Gong, Daniel Cohen-Or, Oliver Deussen, and Baoquan Chen. Men and nature: Flower reconstruction from a single photo. Computer Graphics Forum, 33(2):439– 447, May 2014. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Yan:2014:MBN

[YGJ^{+}14] Dong-Ming Yan, Jianwei Guo, Xiaohong Jia, Xiaopeng Zhang, and Peter Wonka. Meshing: Blue-noise remeshing with farthest point optimization. Computer Graphics Forum, 33(5):167–176, August 2014. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Yang:2013:GST

[YH13] Yong-Liang Yang and Qi-Xing Huang. Geometry (session 1): TrayGen: Arranging objects for exhibition and packaging. Computer Graphics Forum, 32(7):187–195, October 2013. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Yang:2010:SOI

[YHGT10] Jason C. Yang, Justin Hensley, Holger Grün, and Nicolas Thibieroz. Shadows and order independent transparency: Realtime concurrent linked list construction on the GPU. Computer Graphics Forum, 29(4):1297–1304, June 2010. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Yin:2016:FPR

[YHL^{+}16] Kangxue Yin, Hui Huang, Pinxin Long, Alexei Gaissinski, Minglun Gong, and Andrei Sharf. Full 3D plant reconstruction via intrusive acquisition. Computer Graphics Forum, 35 (1):272–284, February 2016. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Yang:2010:FDO

[YI10] Y. Yang and I. Ivrissimtzis. FEM and differential operators: Polygonal mesh watermarking using Laplacian coordinates. Computer Graphics Forum, 29(5):1585–1593, July 2010. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Yue:2011:LRT

[YIC^{+}11] Yonghao Yue, Kei Iwasaki, Bing-Yu Chen, Yoshinori Dobashi, and Tomoyuki Nishita. Lighting & rendering: Toward optimal space partitioning for unbiased, adaptive free path sampling of inhomogeneous participating media. Computer Graphics Forum, 30(7):1911–1919, September 2011. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Yue:2012:MSP

[YIC^{+}12] Yonghao Yue, Kei Iwasaki, Bing-Yu Chen, Yoshinori Dobashi, and Tomoyuki Nishita. Making and shadowing: Pixel art with refracted light by rearrangeable sticks. Computer Graphics Forum, 31(2pt3):575–582, May 2012. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Yang:2018:MGV

[YJD^{+}18] Yalong Yang, Bernhard Jenny, Tim Dwyer, Kim Marriott, Haohui Chen, and Maxime Cordeil. Maps and globes in virtual reality. Computer Graphics Forum, 37(3):427–438, June 2018. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Yang:2012:NPR

[YKM12] Heekyung Yang, Yunmi Kwon, and Kyungha Min. Nonphotorealistic rendering: a stylized approach for pencil drawing from photographs. Computer Graphics Forum, 31(4):1471– 1480, June 2012. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Yau:2019:BDA

[YKS^{+}19] C. Yau, M. Karimzadeh, C. Surakitbanharn, N. Elmqvist, and D. S. Ebert. Bridging the data analysis communication gap utilizing a three-component summarized line graph. Computer Graphics Forum, 38(3):375–386, June 2019. CODEN CG-FODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Ye:2010:MSS

[YL10] Yuting Ye and C. Karen Liu. Motion synthesis: Synthesis of responsive motion using a dynamic model. Computer Graphics Forum, 29(2):555–562, May 2010. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Yu:2011:MSP

[YL11] W. Yu and X. Li. Mesh segmentation & processing: Computing 3D shape guarding and star decomposition. Computer Graphics Forum, 30(7):2087–2096, September 2011. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Yan:2018:MSI

[YLCH18] X. Yan, C-F. Li, X-S. Chen, and S-M. Hu. MPM simulation of interacting fluids and solids. Computer Graphics Forum, 37(8): 183–193, December 2018. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Ye:2018:SIS

[YLD^{+}18] Wenjie Ye, Xiao Li, Yue Dong, Pieter Peers, and Xin Tong. Single image surface appearance modeling with self-augmented CNNs and inexact supervision. Computer Graphics Forum, 37 (7):201–211, October 2018. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Yi:2018:TGM

[YLG^{+}18] Lei Yi, Hongjun Li, Jianwei Guo, Oliver Deussen, and Xiaopeng Zhang. Tree growth modelling constrained by growth equations. Computer Graphics Forum, 37(1):239–253, February 2018. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Yuan:2014:MCC

[YLH^{+}14] Chunqiang Yuan, Xiaohui Liang, Shiyu Hao, Yue Qi, and Qinping Zhao. Modelling Cumulus cloud shape from a single image. Computer Graphics Forum, 33(6):288–297, September 2014. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Yang:2012:AIR

[YLHQ12] Lipeng Yang, Shuai Li, Aimin Hao, and Hong Qin. Animation and interaction: Realtime two-way coupling of meshless fluids and nonlinear FEM. Computer Graphics Forum, 31(7pt1): 2037–2046, September 2012. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Yang:2014:PDH

[YLHQ14] Lipeng Yang, Shuai Li, Aimin Hao, and Hong Qin. Particles and deformation: Hybrid particle-grid modeling for multi-scale droplet/spray simulation. Computer Graphics Forum, 33(7): 199–208, October 2014. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Yoo:2015:CTC

[YLL15] M.-J. Yoo, I.-K. Lee, and S. Lee. Colors and textures: Color sequence preserving decolorization. Computer Graphics Forum, 34(2):373–383, May 2015. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Yang:2015:RSN

[YLLL15] Jingyu Yang, Ke Li, Kun Li, and Yu-Kun Lai. Registration: Sparse non-rigid registration of 3D shapes. Computer Graphics Forum, 34(5):89–99, August 2015. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Yu:2010:VPA

[YLRC10] Li Yu, Aidong Lu, William Ribarsky, and Wei Chen. Visualization and perception: Automatic animation for time-varying data visualization. Computer Graphics Forum, 29(7):2271– 2280, September 2010. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Yuksel:2019:RTM

[YLT19] Cem Yuksel, Sylvain Lefebvre, and Marco Tarini. Rethinking texture mapping. Computer Graphics Forum, 38(2):535–551, May 2019. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Yuan:2016:GMS

[YLX^{+}16] Qing Yuan, Guiqing Li, Kai Xu, Xudong Chen, and Hui Huang. Geometric modeling: Space–time co-segmentation of articulated point cloud sequences. Computer Graphics Forum, 35 (2):419–429, May 2016. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Ye:2017:SAU

[YMJ^{+}17] Juntao Ye, Guanghui Ma, Liguo Jiang, Lan Chen, Jituo Li, Gang Xiong, Xiaopeng Zhang, and Min Tang. Simulation and animation: a unified cloth untangling framework through discrete collision detection. Computer Graphics Forum, 36(7): 217–228, October 2017. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Yan:2010:VPP

[YMM10] S. Yan, N. Max, and K.-L. Ma. Visualization and perception: Polygonal surface advection applied to strange attractors. Computer Graphics Forum, 29(7):2281–2290, September 2010. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Yu:2010:MPM

[YMS10] J. Yu, L. McMillan, and P. Sturm. Multi-perspective modelling, rendering and imaging. Computer Graphics Forum, 29(1):227– 246, March 2010. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Yoshiyasu:2014:MCP

[YMYK14] Yusuke Yoshiyasu, Wan-Chun Ma, Eiichi Yoshida, and Fumio Kanehiro. Maps: As-conformal-as-possible surface registration. Computer Graphics Forum, 33(5):257–267, August 2014. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Younesy:2013:AIA

[YNM^{+}13] H. Younesy, C. B. Nielsen, T. Möller, O. Alder, R. Cullum, M. C. Lorincz, M. M. Karimi, and S. J. M. Jones. Applications: an interactive analysis and exploration tool for epigenomic data. Computer Graphics Forum, 32(3pt1):91–100, June 2013. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Yen:2019:EUS

[YPF19] Chi-Hsien Eric Yen, Aditya Parameswaran, and Wai-Tat Fu. An exploratory user study of visual causality analysis. Computer Graphics Forum, 38(3):173–184, June 2019. CODEN CG-FODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Yu:2019:FSS

[YPL19] Ri Yu, Hwangpil Park, and Jehee Lee. Figure skating simulation from video. Computer Graphics Forum, 38(7):225–234, October 2019. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Yan:2019:FBC

[YS19] Z. Yan and S. Schaefer. A family of barycentric coordinates for co-dimension 1 manifolds with simplicial facets. Computer Graphics Forum, 38(5):75–83, August 2019. CODEN CG-FODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Yang:2018:FSF

[YSC^{+}18] Wenwu Yang, Hock-Soon Seah, Quan Chen, Hong-Ze Liew, and Daniel Sýkora. FTP-SC: Fuzzy topology preserving stroke correspondence. Computer Graphics Forum, 37(8):125–135, December 2018. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Yan:2019:PDV

[YSM19] Dingkun Yan, Yun Sheng, and Xiaoyang Mao. Pencil drawing video rendering using convolutional networks. Computer Graphics Forum, 38(7):91–102, October 2019. CODEN CG-FODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Yuksel:2015:SSS

[Yuk15] Cem Yuksel. Sampling & skins: Sample elimination for generating Poisson disk sample sets. Computer Graphics Forum, 34(2):25–32, May 2015. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Yao:2010:ILA

[YWC^{+}10] Chunhui Yao, Bin Wang, Bin Chan, Junhai Yong, and Jean-Claude Paul. Indirect lighting and ambient occlusion: Multiimage based photon tracing for interactive global illumination of dynamic scenes. Computer Graphics Forum, 29(4):1315– 1324, June 2010. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Yang:2019:CPB

[YWF^{+}19] B. Yang, T. Wei, X. Fang, Z. Deng, F. W. B. Li, Y. Ling, and X. Wang. A color-pair based approach for accurate color harmony estimation. Computer Graphics Forum, 38(7):481–490, October 2019. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Yuan:2018:RSS

[YWHB18] Yazhen Yuan, Rui Wang, Tianlei Hu, and Hujun Bao. Runtime shader simplification via instant search in reduced optimization space. Computer Graphics Forum, 37(4):143–154, July 2018. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Yang:2015:FRS

[YWM15] Yong-Liang Yang, Jun Wang, and Niloy J. Mitra. Fabrication: Reforming shapes for material-aware fabrication. Computer Graphics Forum, 34(5):53–64, August 2015. CODEN CG-FODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Yates:2014:DRV

[YWS^{+}14] A. Yates, A. Webb, M. Sharpnack, H. Chamberlin, K. Huang, and R. Machiraju. Data relationships: Visualizing multidimensional data with glyph SPLOMs. Computer Graphics Forum, 33(3):301–310, June 2014. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Yu:2012:VSE

[YWTY12] Jihun Yu, Chris Wojtan, Greg Turk, and Chee Yap. Visualization and simulation: Explicit mesh surfaces for particle based fluids. Computer Graphics Forum, 31(2pt4):815–824, May 2012. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Yu:2010:RRT

[YWY10] Xuan Yu, Rui Wang, and Jingyi Yu. Rendering: Real-time depth of field rendering via dynamic light field generation and filtering. Computer Graphics Forum, 29(7):2099–2107, September 2010. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Yang:2018:UAV

[YXHH18] Hao Yang, Ke Xie, Shengqiu Huang, and Hui Huang. Uncut aerial video via a single sketch. Computer Graphics Forum, 37 (7):191–199, October 2018. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Yan:2014:CIF

[YXX14] Qingan Yan, Zhan Xu, and Chunxia Xiao. Color and imaging: Fast feature-oriented visual connection for large image collections. Computer Graphics Forum, 33(7):339–348, October 2014. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Yang:2018:DDM

[YYG18] Baorong Yang, Junfeng Yao, and Xiaohu Guo. DMAT: Deformable medial axis transform for animated mesh approximation. Computer Graphics Forum, 37(7):301–311, October 2018. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Yan:2016:MIG

[YYL^{+}16] Qingan Yan, Long Yang, Chao Liang, Huajun Liu, Ruimin Hu, and Chunxia Xiao. Matching and interpolation: Geometrically based linear iterative clustering for quantitative feature correspondence. Computer Graphics Forum, 35(7):1–10, October 2016. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Yang:2018:CSS

[YYZZ18] Lingchen Yang, Lumin Yang, Mingbo Zhao, and Youyi Zheng. Controlling stroke size in fast style transfer with recurrent convolutional neural network. Computer Graphics Forum, 37(7): 97–107, October 2018. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Yuan:2018:CDT

[YZC18] Ye Yuan, Changxi Zheng, and Stelian Coros. Computational design of transformables. Computer Graphics Forum, 37(8): 103–113, December 2018. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Yao:2017:CSS

[YZL17] Guilin Yao, Zhijie Zhao, and Shaohui Liu. A comprehensive survey on sampling-based image matting. Computer Graphics Forum, 36(8):613–628, December 2017. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Yan:2012:RST

[YZXW12] Ling-Qi Yan, Yahan Zhou, Kun Xu, and Rui Wang. Rendering systems and techniques: Accurate translucent material rendering under spherical Gaussian lights. Computer Graphics Forum, 31(7pt2):2267–2276, September 2012. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Zirr:2015:FVV

[ZAD15] Tobias Zirr, Marco Ament, and Carsten Dachsbacher. Flow visualization: Visualization of coherent structures of light transport. Computer Graphics Forum, 34(3):491–500, June 2015. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Zhang:2016:TDD

[ZAM^{+}16] Jiawei Zhang, Benjamin Ahlbrand, Abish Malik, Junghoon Chae, Zhiyu Min, Sungahn Ko, and David S. Ebert. Text and document data: a visual analytics framework for microblog data analysis at multiple scales of aggregation. Computer Graphics Forum, 35(3):441–450, June 2016. CODEN CG-FODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Zhang:2017:TBL

[ZBM^{+}17] Xiaopeng Zhang, Guanbo Bao, Weiliang Meng, Marc Jaeger, Hongjun Li, Oliver Deussen, and Baoquan Chen. Tree branch level of detail models for forest navigation. Computer Graphics Forum, 36(8):402–417, December 2017. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Zanni:2013:SII

[ZBQC13] C. Zanni, A. Bernhardt, M. Quiblier, and M.-P. Cani. SCALeinvariant integral surfaces. Computer Graphics Forum, 32(8): 219–232, December 2013. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Zimmer:2011:QSF

[ZBW11] Henning Zimmer, Andrés Bruhn, and Joachim Weickert. Quality and scalability: Freehand HDR imaging of moving scenes with simultaneous resolution enhancement. Computer Graphics Forum, 30(2):405–414, April 2011. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Zhou:2018:KTS

[ZC18] Bo Zhou and Yi-Jen Chiang. Key time steps selection for large-scale time-varying volume datasets using an informationtheoretic storyboard. Computer Graphics Forum, 37(3):37–49, June 2018. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Zimmer:2012:MSR

[ZCBK12] Henrik Zimmer, Marcel Campen, David Bommes, and Leif Kobbelt. Making and shadowing: Rationalization of trianglebased point-folding structures. Computer Graphics Forum, 31 (2pt3):611–620, May 2012. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Zhang:2014:EBP

[ZCC14] Xiaoyan Zhang, Martin Constable, and Kap Luk Chan. Exemplar-based portrait photograph enhancement as informed by portrait paintings. Computer Graphics Forum, 33(8):38–51, December 2014. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Zhou:2013:IPS

[ZCF^{+}13] Bin Zhou, Xiaowu Chen, Qiang Fu, Kan Guo, and Ping Tan. Image processing (session 2): Garment modeling from a single image. Computer Graphics Forum, 32(7):85–91, October 2013. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Zhang:2017:ODV

[ZCH^{+}17] C. Zhang, M. W. A. Caan, T. Höllt, E. Eisemann, and A. Vilanova. Overview + detail visualization for ensembles of diffusion tensors. Computer Graphics Forum, 36(3):121–132, June 2017. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Zhao:2017:COI

[ZCK17] X. Zhao, M. G. Choi, and T. Komura. Character-object interaction retrieval using the interaction bisector surface. Computer Graphics Forum, 36(2):119–129, May 2017. CODEN CG-FODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Zhao:2018:MFC

[ZCL18] Nanxuan Zhao, Ying Cao, and Rynson W. H. Lau. Modeling fonts in context: Font prediction on Web designs. Computer Graphics Forum, 37(7):385–395, October 2018. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Zheng:2014:SCF

[ZCOAM14] Youyi Zheng, Daniel Cohen-Or, Melinos Averkiou, and Niloy J. Mitra. Shape collections, features: Recurring part arrangements in shape collections. Computer Graphics Forum, 33 (2):115–124, May 2014. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Zheng:2013:SCS

[ZCOM13] Youyi Zheng, Daniel Cohen-Or, and Niloy J. Mitra. Shape construction: Smart variations: Functional substructures for part compatibility. Computer Graphics Forum, 32(2pt2):195– 204, May 2013. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Zhao:2018:CSR

[ZCT18] M. Zhao, W. Cai, and S. J. Turner. CLUST: Simulating realistic crowd behaviour by mining pattern from crowd videos. Computer Graphics Forum, 37(1):184–201, February 2018. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Zeng:2015:ISH

[ZCW^{+}15] Qiong Zeng, Wenzheng Chen, Huan Wang, Changhe Tu, Daniel Cohen-Or, Dani Lischinski, and Baoquan Chen. Images & scenes: Hallucinating stereoscopy from a single image. Computer Graphics Forum, 34(2):1–12, May 2015. CODEN CG-FODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Zhao:2019:OOI

[ZCW^{+}19] Xun Zhao, Weiwei Cui, Yanhong Wu, Haidong Zhang, Huamin Qu, and Dongmei Zhang. Oui! Outlier interpretation on multidimensional data via visual analytics. Computer Graphics Forum, 38(3):213–224, June 2019. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Zhou:2019:AGV

[ZCX19] J. Zhou, X. Chen, and Y. Xu. Automatic generation of vivid LEGO architectural sculptures. Computer Graphics Forum, 38 (6):31–42, September 2019. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Zhang:2013:IPS

[ZCZL13] J. Zhang, X. H. Chen, Y. Zhao, and H. Li. Image processing (session 6): an efficient and scalable image filtering framework using VIPS fusion. Computer Graphics Forum, 32(7):391–400, October 2013. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Zhao:2016:RDA

[ZDJ16] Yandan Zhao, Hui Du, and Xiaogang Jin. Recognitiondifficulty-aware hidden images based on clue-map. Computer Graphics Forum, 35(8):80–94, December 2016. CODEN CG-FODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Zhang:2014:BTP

[ZDM^{+}14] L. Zhang, Q. Deng, R. Machiraju, A. Rangarajan, D. Thompson, D. K. Walters, and H.-W. Shen. Boosting techniques for physics-based vortex detection. Computer Graphics Forum, 33 (1):282–293, February 2014. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Zhang:2015:SMG

[ZDZ^{+}15] Wangyu Zhang, Bailin Deng, Juyong Zhang, Sofien Bouaziz, and Ligang Liu. Shape and mesh: Guided mesh normal filtering. Computer Graphics Forum, 34(7):23–34, October 2015. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Zeng:2016:VWC

[ZFA^{+}16] W. Zeng, C.-W. Fu, S. Müller Arisona, A. Erath, and H. Qu. Visualizing waypoints-constrained origin-destination patterns for massive transportation data. Computer Graphics Forum, 35(8):95–107, December 2016. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Zeng:2013:STV

[ZFAQ13] Wei Zeng, Chi-Wing Fu, Stefan Müller Arisona, and Huamin Qu. Space and time: Visualizing interchange patterns in massive movement data. Computer Graphics Forum, 32(3pt3):271– 280, June 2013. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Zheng:2011:SMM

[ZFCO^{+}11] Youyi Zheng, Hongbo Fu, Daniel Cohen-Or, Oscar Kin-Chung Au, and Chiew-Lan Tai. Shape matching and manipulation: Component-wise controllers for structure-preserving shape manipulation. Computer Graphics Forum, 30(2):563–572, April 2011. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Zhang:2016:TMR

[ZFE16] L. Zhang, S. Fenney, and F. Escribano. Textures/mapping: Real-time texture synthesis and concurrent random-access rendering for low-cost GPU chip design. Computer Graphics Forum, 35(7):87–95, October 2016. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Zheng:2017:RBF

[ZFG^{+}17] Qian Zheng, Xiaochen Fan, Minglun Gong, Andrei Sharf, Oliver Deussen, and Hui Huang. 4D reconstruction of blooming flowers. Computer Graphics Forum, 36(6):405–417, September 2017. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Zhu:2016:IPN

[ZFJ^{+}16] Lei Zhu, Chi-Wing Fu, Yueming Jin, Mingqiang Wei, Jing Qin, and Pheng-Ann Heng. Image processing: Non-local sparse and low-rank regularization for structure-preserving image smoothing. Computer Graphics Forum, 35(7):217–226, October 2016. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Zhang:2012:IEP

[ZH12] Lei Zhang and Hua Huang. Image editing and processing: Hierarchical narrative collage for digital photo album. Computer Graphics Forum, 31(7pt2):2173–2181, September 2012.

CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Zhou:2014:BIG

[ZH14] L. Zhou and C. Hansen. BioMedical II: GuideME: Slice-guided semiautomatic multivariate exploration of volumes. Computer Graphics Forum, 33(3):151–160, June 2014. CODEN CG-FODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Zirr:2018:RWF

[ZHD18] Tobias Zirr, Johannes Hanika, and Carsten Dachsbacher. Reweighting firefly samples for improved finite-sample Monte Carlo estimates. Computer Graphics Forum, 37(6):410–421, September 2018. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Zheng:2015:SMS

[ZHH^{+}15] Qian Zheng, Zhuming Hao, Hui Huang, Kai Xu, Hao Zhang, Daniel Cohen-Or, and Baoquan Chen. Shape manipulation: Skeleton-intrinsic symmetrization of shapes. Computer Graphics Forum, 34(2):275–286, May 2015. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Zhu:2015:RAA

[ZHK15] Lifeng Zhu, Xiaoyan Hu, and Ladislav Kavan. Reconstruction: Adaptable anatomical models for realistic bone motion reconstruction. Computer Graphics Forum, 34(2):459–471, May 2015. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Zhang:2018:BTM

[ZHLW18] Zhuming Zhang, Xinghong Hu, Xueting Liu, and Tien-Tsin Wong. Binocular tone mapping with improved overall contrast and local details. Computer Graphics Forum, 37(7):433–442, October 2018. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Zhai:2017:IMI

[ZHQH17] Xiao Zhai, Fei Hou, Hong Qin, and Aimin Hao. Inverse modelling of incompressible gas flow in subspace. Computer Graphics Forum, 36(6):100–111, September 2017. CODEN CG-FODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Zhu:2013:GSS

[ZIM13] L. Zhu, T. Igarashi, and J. Mitani. Geometry (session 1): Soft folding. Computer Graphics Forum, 32(7):167–176, October 2013. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Zhou:2018:CIT

[ZJ18] Qingnan Zhou and Alec Jacobson. 2018 cover image: Thingi10K. Computer Graphics Forum, 37(1):451–452, February 2018. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Zou:2013:MAT

[ZJC13] Ming Zou, Tao Ju, and Nathan Carr. Meshing: an algorithm for triangulating multiple 3D polygons. Computer Graphics Forum, 32(5):157–166, August 2013. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Zwicker:2015:SAR

[ZJL^{+}15] M. Zwicker, W. Jarosz, J. Lehtinen, B. Moon, R. Ramamoorthi, F. Rousselle, P. Sen, C. Soler, and S.-E. Yoon. State of the art reports: Recent advances in adaptive sampling and reconstruction for Monte Carlo rendering. Computer Graphics Forum, 34(2):667–681, May 2015. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Zhou:2016:FDS

[ZKWG16] Yahan Zhou, Evangelos Kalogerakis, Rui Wang, and Ian R. Grosse. Fabrication: Direct shape optimization for strengthening 3D printable objects. Computer Graphics Forum, 35(7): 333–342, October 2016. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Zheng:2016:ESD

[ZLDM16] Youyi Zheng, Han Liu, Julie Dorsey, and Niloy J. Mitra. Editing, sketch & drawing: Smart Canvas: Context-inferred interpretation of sketches for preparatory design studies. Computer Graphics Forum, 35(2):37–48, May 2016. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Zhou:2013:ASI

[ZLKW13] Yahan Zhou, Zhaoliang Lun, Evangelos Kalogerakis, and Rui Wang. Animation (session 2): Implicit integration for particlebased simulation of elasto-plastic solids. Computer Graphics Forum, 32(7):215–223, October 2013. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Zhou:2013:SES

[ZLL13] Shizhe Zhou, Anass Lasram, and Sylvain Lefebvre. Synthesis: By-example synthesis of curvilinear structured patterns. Computer Graphics Forum, 32(2pt3):355–360, May 2013. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Zhang:2015:SVQ

[ZLM^{+}15] Tianxiang Zhang, Sheng Li, Dinesh Manocha, Guoping Wang, and Hanqiu Sun. Simulation and visualization: Quadratic contact energy model for multi-impact simulation. Computer Graphics Forum, 34(7):133–144, October 2015. CODEN CG-FODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Zhang:2016:SCP

[ZLMM16] Y. Zhang, W. Luo, E. A. Mack, and R. Maciejewski. Structures, clusters, and patterns: Visualizing the impact of geographical variations on multivariate clustering. Computer Graphics Forum, 35(3):101–110, June 2016. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Zhang:2017:VVP

[ZLSW17] Xuaner Zhang, Joon-Young Lee, Kalyan Sunkavalli, and Zhaowen Wang. Video and visualization: Photometric stabilization for fast-forward videos. Computer Graphics Forum, 36 (7):105–113, October 2017. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Zhu:2015:EBM

[ZLW15] Fei Zhu, Sheng Li, and Guoping Wang. Example-based materials in Laplace–Beltrami shape space. Computer Graphics Forum, 34(1):36–46, February 2015. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Zhang:2016:FDD

[ZLW^{+}16] Xiaoting Zhang, Xinyi Le, Zihao Wu, Emily Whiting, and Charlie C. L. Wang. Fabrication: Data-driven bending elasticity design by shell thickness. Computer Graphics Forum, 35 (5):157–166, August 2016. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Zhao:2019:PSS

[ZLW^{+}19] Hui Zhao, Xuan Li, Wencheng Wang, Xiaoling Wang, Shaodong Wang, Na Lei, and Xiangfeng Gu. Polycube shape space. Computer Graphics Forum, 38(7):311–322, October 2019. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Zhang:2017:SAM

[ZLYL17] Zili Zhang, Xiaohui Liang, Chunqiang Yuan, and Frederick W. B. Li. Simulation and animation: Modeling cumulus cloud scenes from high-resolution satellite images. Computer Graphics Forum, 36(7):229–238, October 2017. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Zhu:2018:GAI

[ZLZ^{+}18] Xiaobin Zhu, Zhuangzi Li, Xiaoyu Zhang, Haisheng Li, Ziyu Xue, and Lei Wang. Generative adversarial image superresolution through deep dense skip connections. Computer Graphics Forum, 37(7):289–300, October 2018. CODEN CG-FODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Zhang:2016:VDV

[ZM16] Y. Zhang and K.-L. Ma. Volume data visualization: Decoupled shading for real-time heterogeneous volume illumination. Computer Graphics Forum, 35(3):401–410, June 2016. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Zhang:2019:DIE

[ZNZ19] Qing Zhang, Yongwei Nie, and Wei-Shi Zheng. Dual illumination estimation for robust exposure correction. Computer Graphics Forum, 38(7):243–252, October 2019. CODEN CG-FODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Zhang:2018:FPA

[ZOA^{+}18] Yunjin Zhang, Marta Ortin, Victor Arellano, Rui Wang, Diego Gutierrez, and Hujun Bao. On-the-fly power-aware rendering.

Computer Graphics Forum, 37(4):155–166, July 2018. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Zhi:2019:LLE

[ZOM19] Qiyu Zhi, Alvitta Ottley, and Ronald Metoyer. Linking and layout: Exploring the integration of text and visualization in storytelling. Computer Graphics Forum, 38(3):675–685, June 2019. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Zhuo:2013:MFS

[ZR13] Wei Zhuo and Jarek Rossignac. Models: Fleshing: Spine-driven bending with local volume preservation. Computer Graphics Forum, 32(2pt3):295–304, May 2013. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Zirr:2019:DFD

[ZR19] Tobias Zirr and Tobias Ritschel. Distortion-free displacement mapping. Computer Graphics Forum, 38(8):53–62, November 2019. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Zimmer:2015:RTR

[ZRJ^{+}15] Henning Zimmer, Fabrice Rousselle, Wenzel Jakob, Oliver Wang, David Adler, Wojciech Jarosz, Olga Sorkine-Hornung, and Alexander Sorkine-Hornung. Real-time rendering and filtering: Path-space motion estimation and decomposition for robust animation filtering. Computer Graphics Forum, 34(4):131– 142, July 2015. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Zheng:2019:VEB

[ZRLS19] Boyan Zheng, Bastian Rieck, Heike Leitte, and Filip Sadlo. Visualization of equivalence in 2D bivariate fields. Computer Graphics Forum, 38(3):311–323, June 2019. CODEN CG-FODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Zollhofer:2018:SAR

[ZSG^{+}18] Michael Zollhöfer, Patrick Stotko, Andreas Görlitz, Christian Theobalt, Matthias Nießner, Reinhard Klein, and Andreas Kolb. State of the art on 3D reconstruction with RGB-D cameras. Computer Graphics Forum, 37(2):625–652, May 2018.

CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Zeng:2019:RAE

[ZSJT19] W. Zeng, Q. Shen, Y. Jiang, and A. Telea. Route-aware edge bundling for visualizing origin-destination trails in urban traffic. Computer Graphics Forum, 38(3):581–593, June 2019. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Zhou:2017:ACS

[ZSL^{+}17] Yang Zhou, Huajie Shi, Dani Lischinski, Minglun Gong, Johannes Kopf, and Hui Huang. Analysis and controlled synthesis of inhomogeneous textures. Computer Graphics Forum, 36 (2):199–212, May 2017. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Zayer:2017:GAS

[ZSS17] Rhaleb Zayer, Markus Steinberger, and Hans-Peter Seidel. A GPU-adapted structure for unstructured grids. Computer Graphics Forum, 36(2):495–507, May 2017. CODEN CG-FODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Zheng:2010:SMM

[ZST^{+}10] Q. Zheng, A. Sharf, A. Tagliasacchi, B. Chen, H. Zhang, A. Sheffer, and D. Cohen-Or. Shape matching and modeling: Consensus skeleton for non-rigid space–time registration. Computer Graphics Forum, 29(2):635–644, May 2010. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Zadravec:2010:SSS

[ZSW10a] Mirko Zadravec, Alexander Schiftner, and Johannes Wallner. Surface and space structures: Designing quad-dominant meshes with planar faces. Computer Graphics Forum, 29(5):1671–1679, July 2010. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Zhou:2010:GIF

[ZSW^{+}10b] Yuanfeng Zhou, Feng Sun, Wenping Wang, Jiaye Wang, and Caiming Zhang. Geometry II: Fast updating of Delaunay triangulation of moving points by bi-cell filtering. Computer Graphics Forum, 29(7):2233–2242, September 2010. CODEN CG-FODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Zheng:2010:SSM

[ZT10] Youyi Zheng and Chiew-Lan Tai. Shape segmentation: Mesh decomposition with cross-boundary brushes. Computer Graphics Forum, 29(2):527–535, May 2010. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Zollhofer:2018:SAM

[ZTG^{+}18] M. Zollhöfer, J. Thies, P. Garrido, D. Bradley, T. Beeler, P. Pérez, M. Stamminger, M. Nießner, and C. Theobalt. State of the art on monocular 3D face reconstruction, tracking, and applications. Computer Graphics Forum, 37(2):523–550, May 2018. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Zhang:2012:IHS

[ZTW^{+}12] Qing Zhang, Jing Tong, Huamin Wang, Zhigeng Pan, and Ruigang Yang. Imaging and hair: Simulation guided hair dynamics modeling from video. Computer Graphics Forum, 31 (7pt1):2003–2010, September 2012. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Zhang:2010:SMP

[ZVD10] H. Zhang, O. Van Kaick, and R. Dyer. Spectral mesh processing. Computer Graphics Forum, 29(6):1865–1894, September 2010. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Zukic:2014:RDS

[ZVE^{+}14] Dženan Zukić, Aleš Vlasák, Jan Egger, Daniel Hořínek, Christopher Nimsky, and Andreas Kolb. Robust detection and segmentation for diagnosis of vertebral diseases using routine MR images. Computer Graphics Forum, 33(6):190–204, September 2014. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Zhang:2010:SSM

[ZWC^{+}10] Juyong Zhang, Chunlin Wu, Jianfei Cai, Jianmin Zheng, and Xue cheng Tai. Shape segmentation: Mesh snapping: Robust interactive mesh cutting using fast geodesic curvature flow. Computer Graphics Forum, 29(2):517–526, May 2010. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Zielasko:2016:NGI

[ZWHK16] D. Zielasko, B. Weyers, B. Hentschel, and T. W. Kuhlen. Networks and graphs 1: Interactive 3D force-directed edge bundling. Computer Graphics Forum, 35(3):51–60, June 2016. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Zhang:2019:CAR

[ZWJ^{+}19] Yu-Wei Zhang, Jing Wu, Zhongping Ji, Mingqiang Wei, and Caiming Zhang. Computer-assisted relief modelling: a comprehensive survey. Computer Graphics Forum, 38(2):521–534, May 2019. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Zhao:2014:ALL

[ZWRH14] Kaiyu Zhao, Matthew O. Ward, Elke A. Rundensteiner, and Huong N. Higgins. Applications: LoVis: Local pattern visualization for model refinement. Computer Graphics Forum, 33 (3):331–340, June 2014. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Zhang:2017:SOT

[ZWXL17] Junsong Zhang, Yu Wang, Weiyi Xiao, and Zhenshan Luo. Synthesizing ornamental typefaces. Computer Graphics Forum, 36 (1):64–75, January 2017. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Zhu:2013:GSC

[ZWY^{+}13] Lei Zhu, Mingqiang Wei, Jinze Yu, Weiming Wang, Jing Qin, and Pheng-Ann Heng. Geometry (session 2): Coarse-to-fine normal filtering for feature-preserving mesh denoising based on isotropic subneighborhoods. Computer Graphics Forum, 32(7): 371–380, October 2013. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Zhou:2010:SDA

[ZXTD10] Kun Zhou, Weiwei Xu, Yiying Tong, and Mathieu Desbrun. Shape deformation and animation: Deformation transfer to multi-component objects. Computer Graphics Forum, 29(2): 319–325, May 2010. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Zhu:2019:SSM

[ZXW19] Junqiu Zhu, Yanning Xu, and Lu Wang. A stationary SVBRDF material modeling method based on discrete microsurface. Computer Graphics Forum, 38(7):745–754, October 2019. CO-DEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Zhao:2017:SCT

[ZXZ^{+}17] Haiming Zhao, Weiwei Xu, Kun Zhou, Yin Yang, Xiaogang Jin, and Hongzhi Wu. Stress-constrained thickness optimization for shell object fabrication. Computer Graphics Forum, 36(6):368– 380, September 2017. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Zhu:2010:SCP

[ZYF10] Bo Zhu, Xubo Yang, and Ye Fan. Simulation: Creating and preserving vortical details in SPH fluid. Computer Graphics Forum, 29(7):2207–2214, September 2010. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Zhang:2013:GSS

[ZYF13] Zhiyuan Zhang, KangKang Yin, and Kelvin W. C. Foong. Geometry (session 2): Symmetry robust descriptor for non-rigid surface matching. Computer Graphics Forum, 32(7):355–362, October 2013. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Zheng:2017:NDD

[ZZ17] Quan Zheng and Changwen Zheng. NeuroLens: Data-driven camera lens simulation using neural networks. Computer Graphics Forum, 36(8):390–401, December 2017. CODEN CG-FODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Zheng:2019:LIS

[ZZ19] Quan Zheng and Matthias Zwicker. Learning to importance sample in primary sample space. Computer Graphics Forum, 38(2):169–179, May 2019. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Zhuang:2014:SCA

[ZZCJ14] Yixin Zhuang, Ming Zou, Nathan Carr, and Tao Ju. Shapes and cryptography: Anisotropic geodesics for live-wire mesh segmentation. Computer Graphics Forum, 33(7):111–120, October 2014. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Zhang:2015:IVE

[ZZH15] Lei Zhang, Yu-Hang Zhang, and Hua Huang. Image and video: Efficient variational light field view synthesis for making stereoscopic 3D images. Computer Graphics Forum, 34(7):183–191, October 2015. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Zhu:2017:DEM

[ZZL^{+}17] Fei Zhu, Jing Zhao, Sheng Li, Yong Tang, and Guoping Wang. Dynamically enriched MPM for invertible elasticity. Computer Graphics Forum, 36(6):381–392, September 2017. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Zhang:2017:VVV

[ZZLX17] Ling Zhang, Yao Zhu, Bin Liao, and Chunxia Xiao. Video and visualization: Video shadow removal using spatio-temporal illumination transfer. Computer Graphics Forum, 36(7):125–134, October 2017. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Zhang:2015:BMR

[ZZT15] Wenjing Zhang, Jianmin Zheng, and Nadia Magnenat Thalmann. Bodies in motion: Real-time subspace integration for example-based elastic material. Computer Graphics Forum, 34 (2):395–404, May 2015. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Zhang:2016:FAB

[ZZWC16] Yu-Wei Zhang, Caiming Zhang, Wenping Wang, and Yanzhao Chen. Fabrication: Adaptive bas-relief generation from 3D object under illumination. Computer Graphics Forum, 35(7): 311–321, October 2016. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Zheng:2019:ASU

[ZZZ^{+}19] Lintao Zheng, Chenyang Zhu, Jiazhao Zhang, Hang Zhao, Hui Huang, Matthias Niessner, and Kai Xu. Active scene understanding via online semantic reconstruction. Computer Graph-REFERENCES 493

ics Forum, 38(7):103–114, October 2019. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).